Page 1  MITSFS Seriesdex (Quick List of Series)  September 24, 2007
ENTRIES  SERIES  AUTHOR
2  ABACAR  ERENBERGER, TIMOTHY
1  ABARAT  BARKER, CLIVE
1  ABC WARRIORS  MILLS, PAT/MITCHELL, ALAN
2  ABHORSEN  NIX, GARTH
1  ACCOMPLICE  AYLETT, STEVE
9  ACORNA  authorship varies
2  ACORNA’S CHILDREN  MCCAFFREY, ANNE/SCARBOROUGH, ELIZABETH ANN
2  ADDAMS FAMILY  authorship varies
5  ADEPT  KURTZ, KATHERINE/HARRIS, DEBORAH TURNER
6  ADVERSARY  WILSON, F. PAUL
1  ADVOCACY  LEE, SHARON/MILLER, STEVE
3  AEGYPT  CROWLEY, JOHN
3  AELWYN  CHARRETTE, ROBERT N.
4  AFRIKORPS  DOLAN, BILL
1  AGE OF CONAN ANOK  YORK, J. STEVEN
3  AGE OF CONAN KERN  COLEMAN, LOREN L.
3  AGE OF CONAN MARAUDERS  MARIOTTE, JEFF
3  AGE OF CONAN SOLDIER’S QUEST  KNAAK, RICHARD A.
1  AGE OF DINOSAURS  RIVKIN, J. F.
2  AGE OF DISCOVERY  STACKPOLE, MICHAEL A.
2  AGE OF FIRE  KNIGHT, E. E.
3  AGE OF MISRULE  CHADBOURN, MARK
3  AGE OF THE FIVE  CANAVAN, TRUDI
4  AGE OF UNREASON  KEYES, J. GREGORY
4  AGENT OF TERRA  MADDOCK, LARRY
2  AGGAR  WOLFE, CHRIS ANNE
2  AGGRESSOR SIX  MCCARTHY, WIL
2  AGHAMORE  NEASON, REBECCA
2  AIAH  WILLIAMS, WALTER JON
3  AISLING GREY  MACALISTER, KATIE
2  ALAMUT  TARR, JUDITH
7  ALASTOR  VANCE, JACK
3  ALDABRESHIN COMPASS  MCKENNA, JULIET E.
4  ALDAIR  BARRETT, NEAL JR.
2  ALDERLY  GARNER, ALAN
3  ALEUTIAN  JONES, GWYNETH
3  ALEX BENEDICT  MCDEVITT, JACK
3  ALEXIA AND TIRAN  GREEN, SHARON
4  ALFAR  BOYER, ELIZABETH H.
3  ALIEN CHRONICLES  CHESTER, DEBORAH
1  ALIEN HUNTRESS  SHOWALTER, GENA
9  ALIEN NATION  authorship varies
4  ALIEN SPEEDWAY  authorship varies
4  ALIEN STARS  MITCHELL, ELIZABETH
2  ALIEN TRACE  MAJOR, H. M.
15  ALIENS  authorship varies
3  ALIENS VS. PREDATOR  authorship varies
3  ALLAN QUATERMAIN  HAGGARD, H. RIDER
3  ALLIANCE OF LIGHT  WURTS, JANNY
1  ALO NUDGER  LUTZ, JOHN
3  ALPHA CENTAURI  ELY, MICHAEL
3  ALTA  YOLEN, JANE
8  ALVIN MAKER  CARD, ORSON SCOTT
15  AMBER  authorship varies
2  AMBERGRIS  VANDERMEER, JEFF
3  AMBERMERE  PIERCE, J. CALVIN
4  AMERICA 2040  INNES, EVAN
4  AMTRAK WARS  TILLEY, PATRICK
1  ANCIENT DARKNESS  PAVER, MICHELLE
4  ANDROMEDA  HOYLE, FRED/ELLIOT, JOHN
5  ANDROMEDA (GENE RODDENBERRY’S)  authorship varies
29  ANGEL  authorship varies
3  ANGEL’S LUCK  FAUST, JOE CLIFFORD
2  ANGHARA  ALEXANDER, ALMA
18  ANITA BLAKE  HAMILTON, LAURELL K.
2  ANNA STRONG  STEIN, JEANNE C.
2  ANNO DRACULA  NEWMAN, KIM
3  ANTARES  MCCOLLUM, MICHAEL
2  ANTEROS  authorship varies
2  ANTHI  BLAKENEY, JAY D.
3  ANTHONY VILLIERS  PANSHIN, ALEXEI
1  ANTRIAN MESSENGER  WISLER, G. CLIFTON
1  APOTHEOSIS  HILL, DOUGLAS
4  APPLESEED  SHIROW, MASAMUNE

ENTRIES  SERIES  AUTHOR
8  APPRENTICE ADEPT  ANTHONY, PIERS
1  AQUANAUTS  STANTON, KEN
2  AQUASILVA  AUDLEY, ANSELM
3  AQUILIAD  SOMTOW, S. P.
3  ARAFEL  CHERRYH, C. J.
1  ARAFEL (REVISED)  CHERRYH, C. J.
1  ARATHELLEAN WARS  MORRIS, TEE
3  ARBITER  DOUGLAS, L. WARREN
2  ARC ONE  HUGHES, MONICA
2  ARC RIDERS  DRAKE, DAVID/MORRIS, JANET
2  ARCADY  WILLIAMS, MICHAEL
2  ARCANA  LLYWELYN, MORGAN/SCOTT, MICHAEL
1  ARCANE SOCIETY  KRENTZ, JAYNE ANN
4  ARCHANGEL  MOREHOUSE, LYDA
2  ARCHFORM BEAUTY  MODESITT, L. E. JR.
6  ARCHONATE  HUGHES, MATTHEW
10  AREA 51  DOHERTY, ROBERT
3  AREIA DARENGA  MEZO, FRANCINE
5  ARGONIA  SCARBOROUGH, ELIZABETH ANN
1  ARHAMMER BRUNNER  WERNER, C. L.
3  ARMAGEDDON  RANKIN, ROBERT
3  ARNA  ROWLEY, CHRISTOPHER
2  ARRANDIN  HERNIMAN, MARCUS
3  ARRIS  COOKE, CATHERINE
2  ART  BARKER, CLIVE
5  ARTEMIS FOWL  COLFER, EOIN
2  ARTHUR PENN  DAVID, PETER
2  ARTHUR WAR LORD  AB HUGH, DAFYDD
4  ARTHURIAN  STEWART, MARY
4  ARWIN  SMITH, GEORGE H.
3  AS IT IS ON MARS  CRONIN, THOMAS W.
2  ASAHIEL  THOMPSON, ELDON
2  ASGARD  STABLEFORD, BRIAN M.
4  ASH  GENTLE, MARY
3  ASH STAFF  FISHER, PAUL R.
9  ASHES  JOHNSTONE, WILLIAM W.
2  ASHLU  BAKER, SCOTT
1  ASHTON FORD  PENDLETON, DON
4  ASIMOV’S ROBOT MYSTERY  authorship varies
4  ASTEROID WARS  BOVA, BEN
2  ASTROLOGER  DEMARTINO, DENNY
9  ATEVI  CHERRYH, C. J.
2  ATHANOR  LINDSKOLD, JANE
2  ATHENA  HSIEH, DEAN
11  ATLAN  authorship varies
4  ATLANTIS  DONEGAN, GREG
2  ATROCITY ARCHIVES  STROSS, CHARLES
2  ATTAR  GRAHAM, ROBERT
3  AUBRY KNIGHT  BARNES, STEVEN
3  AURELIAN EMPIRE  BRENNAN, CAITLIN
4  AURIAN  FUREY, MAGGIE
5  AUTHOR’S CHOICE MONTHLY  authorship varies
2  AUTUMN WORLD MOSAIC  GRAVEL, GEARY
6  AVARYAN RISING  TARR, JUDITH
1  AVATARS  SUTHERLAND, TUI T.
36  AVENGER  ROBESON, KENNETH
7  AVENGERS  authorship varies
3  AVENGERS (MARVEL COMICS CHARACTERS)  authorship varies
2  AVERALAAN  WEST, MICHELLE
4  AVERIDAN  CLOUGH, BRENDA W.
2  AWAKENED KING  CHARRETTE, ROBERT N.
3  AWAKENERS  TEPPER, SHERI S.
3  AXIS OF TIME  BIRMINGHAM, JOHN
4  AYESHA  HAGGARD, H. RIDER
4  AZZIE  ZELAZNY, ROGER/SHECKLEY, ROBERT
12  BABYLON 5  authorship varies
3  BABYLON 5 LEGIONS OF FIRE  DAVID, PETER
3  BABYLON 5 PASSING OF THE TECHNO-MAGES  CAVELOS, JEANNE
3  BABYLON 5 PSI CORPS  KEYES, J. GREGORY
3  BACK TO THE FUTURE  authorship varies
1  BAGEL STREET DOZEN  FISH, ROBERT L.
3  BALZAN  MOORE, WALLACE
4  BAND OF FOUR  GREENWOOD, ED
5  BANNED AND THE BANISHED  CLEMENS, JAMES

ENTRIES  SERIES  AUTHOR
6  BANTAM CONAN  authorship varies
2  BARAKHAI  REICHERT, MICKEY ZUCKER
4  BARD  TAYLOR, KEITH
8  BARD’S TALE  authorship varies
1  BARDIC CHOICES  LACKEY, MERCEDES/SHERMAN, JOSEPHA
5  BARDIC VOICES  LACKEY, MERCEDES
2  BARION AND COYUL  GODWIN, PARKE
5  BARNABY  JOHNSON, CROCKETT
1  BARONESS  KENYON, PAUL
4  BAROQUE CYCLE  STEPHENSON, NEAL
2  BARRA THE PICT  STOVER, MATTHEW WOODRING
2  BARRAZAN V  GOTLIEB, PHYLLIS
26  BARSOOM  BURROUGHS, EDGAR RICE
3  BARTIMAEUS  STROUD, JONATHAN
2  BASIL CASTELLAN  WHITE, STEVE
1  BASTARD PRINCESS  EDWARDS, CLAUDIA J.
4  BATMAN  authorship varies
2  BATMAN ANIMATED SERIES  GRAVEL, GEARY
3  BATTACHARIYA  SHEFFIELD, CHARLES
3  BATTLE CIRCLE  ANTHONY, PIERS
22  BATTLESTAR GALACTICA 1978  authorship varies
4  BATTLESTAR GALACTICA 2003  authorship varies
2  BATTLESTATION  DRAKE, DAVID/FAWCETT, BILL
84  BATTLETECH  authorship varies
7  BAZIL BROKETAIL  ROWLEY, CHRISTOPHER
3  BEAUTY  ROQUELAURE, A. N.
3  BEAUTY AND THE BEAST  authorship varies
2  BECCA  FRIESNER, ESTHER M.
3  BECOMING ALIEN  ORE, REBECCA
3  BEGGARS  KRESS, NANCY
1  BEKA COOPER  PIERCE, TAMORA
10  BELGARIAD  authorship varies
7  BELISARIUS  authorship varies
1  BENEATH STRANGE SKIES  LEE, SHARON/MILLER, STEVE
2  BENTON HAWKES  SHATNER, WILLIAM
2  BERBORA  SIROTA, MIKE
3  BERKLEY CONAN  HOWARD, ROBERT E.
2  BERKLEY SIX MILLION DOLLAR MAN  JAHN, MICHAEL
11  BERNICE SUMMERFIELD  authorship varies
19  BERSERKER  SABERHAGEN, FRED
3  BERSERKER SWIFTAXE  CARLSEN, CHRIS
2  BESTIARY  FORRESTER, JOHN
5  BETSY TAYLOR  DAVIDSON, MARYJANICE
2  BEWITCHING MYSTERY  ALT, MADELYN
2  BICYCLING THROUGH SPACE AND TIME  SIROTA, MIKE
5  BIFROST GUARDIANS  REICHERT, MICKEY ZUCKER
3  BIG BRAIN  BRANDNER, GARY
7  BILL  authorship varies
4  BINDING OF THE BLADE  GRAHAM, L. B.
6  BIO OF A SPACE TYRANT  ANTHONY, PIERS
2  BIO RESCUE  VIEHL, S. L.
4  BIOFAB  BERRY, STEPHEN AMES
2  BIONIC WOMAN  LOTTMAN, EILEEN
3  BIOWARRIORS  VARDEMAN, ROBERT E.
3  BIRD OF KINSHIP  COWPER, RICHARD
4  BIRTHRIGHT  authorship varies
2  BIRTHRIGHT PROJECT  MACKEL, KATHRYN
4  BITTERBYNDE  DART-THORNTON, CECILIA
3  BJ VAUGHAN  O’DONOHOE, NICK
12  BLACK COMPANY  COOK, GLEN
6  BLACK COMPANY GLITTERING STONE  COOK, GLEN
3  BLACK COMPANY SOUTH  COOK, GLEN
4  BLACK HOLE TRAVEL AGENCY  MCKINNEY, JACK
5  BLACK JEWELS  BISHOP, ANNE
3  BLACK MAGICIAN  CANAVAN, TRUDI
2  BLACK OAK  GRANT, CHARLES L.
2  BLACK THRONE  RUSCH, KRISTINE KATHRYN
1  BLACK VULMEA  SMITH, DAVID C.
3  BLACKCOLLAR  ZAHN, TIMOTHY
6  BLACKSTONE  SAUL, JOHN
2  BLACKWATER  MCDOWELL, MICHAEL
37  BLADE  LORD, JEFFREY
2  BLADE RUNNER  JETER, K. W.
4  BLAKE’S 7  authorship varies

ENTRIES  SERIES  AUTHOR
5  BLENDING  GREEN, SHARON
3  BLENDING ENTHRONED  GREEN, SHARON
4  BLOOD AND STONE  NEWCOMB, ROBERT
3  BLOOD BOWL  FORBECK, MATT
3  BLOOD OF KERENSKY  STACKPOLE, MICHAEL A.
5  BLOOD OF TEN CHIEFS  authorship varies
3  BLOOD OF THE GODDESS  DALKEY, KARA
3  BLOOD OF THE LAMB  ROGERS, MARK E.
2  BLOOD OPERA  LEE, TANITH
3  BLOODFANG  WEAVER, MICHAEL D.
2  BLOODSHADOWS  authorship varies
4  BLOSSOM CULP  PECK, RICHARD
16  BOLOS  authorship varies
2  BOOK OF ALL HOURS  DUNCAN, HAL
2  BOOK OF MAN  WELCH, JANE
4  BOOK OF THE SHORT SUN  WOLFE, GENE
3  BOOK OF WORDS  JONES, J. V.
4  BOOK OF YEARS  MORWOOD, PETER
3  BOOKS OF MAGIC  JABLONSKI, CARLA
5  BOOKS OF THE ISLE  SPRINGER, NANCY
1  BORDERLAND  WINDLING, TERRI/SHERMAN, DELIA
6  BORDERLANDS  authorship varies
2  BORDERLANDS (IVERSTERRE)  FREEMAN, LORNA
2  BOREAL MOON  MAY, JULIAN
3  BORRIBLES  DE LARRABEITI, MICHAEL
5  BOTANY  MCCAFFREY, ANNE
1  BRAIDED PATH  WOODING, CHRIS
6  BRAK  JAKES, JOHN
4  BRAN MAK MORN  authorship varies
5  BRENTFORD  RANKIN, ROBERT
3  BRIDEI  MARILLIER, JULIET
4  BRIDGE OF D’ARNATH  BERG, CAROL
2  BRIERLY  MARCELLAS, DIANA
1  BRION BAYARD  LAUMER, KEITH
3  BROKEN CITADEL  GREGORIAN, JOYCE BALLOU
1  BROMELIAD  PRATCHETT, TERRY
3  BRONWYN  MILLER, RON
3  BRONZE CANTICLES  HICKMAN, TRACY/HICKMAN, LAURA
3  BROTHERS OF THE DRAGON  BAILEY, ROBIN WAYNE
7  BUCK ROGERS  authorship varies
5  BUCK ROGERS (NIVEN & POURNELLE’S)  authorship varies
3  BUCK ROGERS INNER PLANETS  authorship varies
2  BUCK ROGERS INVADERS OF CHARON  authorship varies
3  BUCK ROGERS MARTIAN WARS  MURDOCK, M. S.
62  BUFFY  authorship varies
3  BUFFY ANGEL UNSEEN  HOLDER, NANCY/MARIOTTE, JEFF
3  BUFFY GATEKEEPER  GOLDEN, CHRISTOPHER/HOLDER, NANCY
5  BUFFY LOST SLAYER  GOLDEN, CHRISTOPHER
1  BUFFY SPIKE  DAVID, PETER
1  BUFFY SPIKE & DRU  GOLDEN, CHRISTOPHER
3  BUFFY WICKED WILLOW  NAVARRO, YVONNE
3  BUNNICULA  authorship varies
1  BUREAU OF SPECIAL INVESTIGATIONS  ALLEN, ROGER MACBRIDE
5  BUREAU 13  POLLOTTA, NICK
1  BUTTERFLY ST. CYR  BES SHAHAR, ELUKI
6  C.A.D.S.  SIEVERT, JOHN
3  CABAL  DUNN, PHILIP
3  CADRE  O’RIORDAN, ROBERT
6  CADWAL  VANCE, JACK
3  CAGEWORLD  KAPP, COLIN
4  CAITHAN CRUSADE  SMITH, JULIE DEAN
2  CAL AND NIKO  THURMAN, ROB
1  CAL MCDONALD  NILES, STEVE
2  CALEDON  HARRIS, DEBORAH TURNER
3  CALIBAN  ALLEN, ROGER MACBRIDE
3  CALIFORNIA DRAGONS  GUNNARSSON, THORARINN
2  CALINDOR  KILIAN, CRAWFORD
10  CALLAHAN’S CROSSTIME SALOON  ROBINSON, SPIDER
4  CAMELOT  LANDIS, ARTHUR H.
1  CAMP HAUNTED HILL  COVILLE, BRUCE
8  CAMULOD  WHYTE, JACK
5  CANOPUS IN ARGOS  LESSING, DORIS
3  CANTRIX  MARLEY, LOUISE
16  CAP KENNEDY  KERN, GREGORY

ENTRIES  SERIES  AUTHOR
2  CAPELLAN SOLUTION  COLEMAN, LOREN L.
2  CAPTAINS OUTRAGEOUS  YOUNG, ROY V.
3  CAR WARRIORS  authorship varies
2  CARGO AND GHOSTER  LEWITT, SHARIANN
2  CAROLUS REX  NORTON, ANDRE/EDGHILL, ROSEMARY
7  CARPATHIAN  FEEHAN, CHRISTINE
24  CASCA  authorship varies
2  CASSANDRA KRESNOV  SHEPHERD, JOEL
2  CASSANDRA PALMER  CHANCE, KAREN
6  CASTAWAYS IN TIME  ADAMS, ROBERT
1  CASTAWAYS OF THE FLYING DUTCHMAN  JACQUES, BRIAN
10  CASTLE  authorship varies
4  CASTLE BRASS  MOORCOCK, MICHAEL
2  CASTLE FALKENSTEIN  DE CHANCIE, JOHN
4  CAT  VINGE, JOAN D.
3  CATACLYSM  WILLIAMS, SEAN
3  CATS OF MARS  SARRANTONIO, AL
4  CECILIA & KATHERINE  WREDE, PATRICIA C./STEVERMER, CAROLINE
2  CELAENO  FLETCHER, JANE
3  CELTIC CRUSADES  LAWHEAD, STEPHEN R.
3  CELYDONN GREEN LION  EDGERTON, TERESA
3  CELYDONN GWENLLIANT  EDGERTON, TERESA
6  CENOTAPH ROAD  VARDEMAN, ROBERT E.
3  CHAGA  MCDONALD, IAN
2  CHAINS  ROSEN, SELINA
3  CHALION  BUJOLD, LOIS MCMASTER
1  CHALLENGE CHAOS  STABLEFORD, BRIAN M.
3  CHANGELING  GARRISON, PETER
2  CHANGELING DETECTIVE  CUNNINGHAM, ELAINE
2  CHANGELING STAR  CARVER, JEFFREY A.
4  CHANGES  DICKINSON, PETER
4  CHANGEWINDS  CHALKER, JACK L.
5  CHANUR  CHERRYH, C. J.
3  CHAOS CHRONICLES  CARVER, JEFFREY A.
3  CHAOS GATE  COOPER, LOUISE
4  CHARLIE MOON  YARBRO, CHELSEA QUINN
38  CHARMED  authorship varies
2  CHARON  ASARO, CATHERINE
12  CHEYSULI  ROBERSON, JENNIFER
1  CHIA BLACK DRAGON  MARLEY, STEPHEN
5  CHICKS IN CHAINMAIL  authorship varies
2  CHILD QUEEN  MCKENZIE, NANCY
15  CHILDE CYCLE  authorship varies
1  CHILDERIC THE SHATTERER  EDGHILL, ROSEMARY
2  CHILDREN OF CHAOS  DUNCAN, DAVE
2  CHILDREN OF EARTH  WELLS, CATHERINE
2  CHILDREN OF THE CHANGELING  KEYES, J. GREGORY
2  CHILDREN OF THE NIGHT  HODGMAN, ANN
4  CHILDREN OF THE STARS  COULSON, JUANITA
7  CHILL  SHERMAN, JORY
2  CHOSEN  WATT-EVANS, LAWRENCE
9  CHRESTOMANCI  JONES, DIANA WYNNE
4  CHRIST CLONE  BEAUSEIGNEUR, JAMES
4  CHROMAGIC  ANTHONY, PIERS
4  CHRONICLES OF THE RAVEN  BARCLAY, JAMES
2  CHRONICLES UNBINDING  STONE, ROBERT S.
1  CHRONO CRUSADE  MORIYAMA, DAISUKE
3  CHRONOPLANE WARS  KILIAN, CRAWFORD
2  CHTHON  ANTHONY, PIERS
8  CHUNG KUO  WINGROVE, DAVID
4  CINEVERSE  GARDNER, CRAIG SHAW
2  CINGULUM  ROBERTS, JOHN MADDOX
3  CIRCLE  ROBERTS, NORA
4  CIRCLE OF LIGHT  HANCOCK, NIEL
5  CIRCLE OF MAGIC  PIERCE, TAMORA
3  CIRCUIT  SNODGRASS, MELINDA M.
2  CIRCUIT OF HEAVEN  DANVERS, DENNIS
3  CIRCUS WORLD  LONGYEAR, BARRY B.
3  CITIES IN FLIGHT  BLISH, JAMES
1  CLAIDI  LEE, TANITH
4  CLAYTON DREW  FEARN, JOHN RUSSELL
2  CLEVELAND UNDEAD  SWINIARSKI, S. A.
3  CLINGFIRE  BRADLEY, MARION ZIMMER/ROSS, DEBORAH J.
3  CLOAK AND DAGGER  GROELL, ANNE LESLEY
2  CLONE REPUBLIC  KENT, STEVEN L.

ENTRIES  SERIES  AUTHOR
4  CLOUDMAGES  FARRELL, S. L.
3  CLOUDSHIPS OF ORION  MCALLISTER, P. K.
5  CLUSTER  ANTHONY, PIERS
1  COBB ANDERSON  RUCKER, RUDY
5  COBRA  ZAHN, TIMOTHY
2  COCOON  SAPERSTEIN, DAVID
3  CODEX ALERA  BUTCHER, JIM
2  COLD CASH WAR  authorship varies
3  COLDFIRE  FRIEDMAN, C. S.
4  COLIN MACINTYRE  WEBER, DAVID
5  COLLECTED STORIES JACK WILLIAMSON  WILLIAMSON, JACK
1  COLLECTED STORIES OF PHILIP K. DICK  DICK, PHILIP K.
2  COLLEGE OF MAGICS  STEVERMER, CAROLINE
3  COLOSSUS  JONES, D. F.
3  COLSEC  HILL, DOUGLAS
1  COMBAT COMMAND CHILDE CYCLE  DICKSON, GORDON R./DENNING, TROY/GLABERSON, CORY
2  COMMANDER HALLEY  MCARTHUR, MAXINE
2  COMMONWEALTH  HAMILTON, PETER F.
13  COMPANY  BAKER, KAGE
6  COMPLEMENTARY FAIRY COLORS  DATLOW, ELLEN/WINDLING, TERRI
7  COMPLETE ELFQUEST  PINI, WENDY/PINI, RICHARD
10  COMPLETE STORIES OF THEODORE STURGEON  STURGEON, THEODORE
2  COMPLETE TENN  TENN, WILLIAM
62  CONAN  authorship varies
4  CONCLAVE OF SHADOWS  FEIST, RAYMOND E.
3  CONCORDAT  SWYCAFFER, JEFFERSON P.
2  CONEK HAYDEN  MCKEONE, LEE
1  CONFEDERATE STATES  TURTLEDOVE, HARRY
3  CONFEDERATE STATES AMERICAN EMPIRE  TURTLEDOVE, HARRY
3  CONFEDERATE STATES: GREAT WAR  TURTLEDOVE, HARRY
4  CONFLUENCE  MCAULEY, PAUL J.
1  CONNOR GREY  DEL FRANCO, MARK
3  CONQUERORS  ZAHN, TIMOTHY
7  CONRAD STARGARD  FRANKOWSKI, LEO
3  CONTINUING TIME  MORAN, DANIEL KEYS
2  CONTRARYWISE  GREENHALGH, ZOHRA
2  CORAMONDE  DALEY, BRIAN
3  CORDELIA NAISMITH  BUJOLD, LOIS MCMASTER
7  COREAN  MODESITT, L. E. JR.
5  CORMAC  authorship varies
9  CORNELIUS  authorship varies
3  CORNELIUS MURPHY  RANKIN, ROBERT
3  CORPUS DEI  MORROW, JAMES
7  CORUM  MOORCOCK, MICHAEL
2  CORWYN THE ALCHEMIST  CLARK, DOUGLAS W.
3  COYOTE  STEELE, ALLEN
6  COYOTE JONES  ELGIN, SUZETTE HADEN
4  CRABS  SMITH, GUY N.
2  CRADLE OF SATURN  HOGAN, JAMES P.
1  CRAFTERS  STASHEFF, CHRISTOPHER/FAWCETT, BILL
3  CRAIG KREIDENT  ANDERSON, KEVIN J./BEASON, DOUG
2  CRASHCOURSE  BAIRD, WILHELMINA
3  CRAY ORMORU  EISENSTEIN, PHYLLIS
2  CRETACEOUS SEA  HUBBELL, WILL
1  CRIMSON CRYSTAL  LAWSON, SUSAN
4  CRIMSON SHADOW  SALVATORE, R. A.
4  CRISIS OF EMPIRE  authorship varies
2  CROSSFIRE  KRESS, NANCY
1  CROSSROADS  ELLIOTT, KATE
2  CROSSTIME  NORTON, ANDRE
3  CROSSTIME TRAFFIC  TURTLEDOVE, HARRY
4  CROW  authorship varies
7  CROWN OF STARS  ELLIOTT, KATE
1  CROWN OF THE ISLES  DRAKE, DAVID
1  CROWTHISTLE  DART-THORNTON, CECILIA
2  CRUCIBLE  DOUGLASS, SARA
3  CRYSTAL SINGER  MCCAFFREY, ANNE
1  CRYSTAL WALLS  DRAKE, DAVID
2  CRYSTAL WARRIORS  FORSTCHEN, WILLIAM R./MORRISON, GREG
2  CULTURAL SURVEY  BIGGLE, LLOYD JR.
6  CULTURE  BANKS, IAIN M.
2  CUPS AND SORCERY  COSTIKYAN, GREG
3  CURTIS MACURDY  DALMAS, JOHN
3  CV  KNIGHT, DAMON
1  CYBERNARC  CAIN, ROBERT

ENTRIES  SERIES  AUTHOR
2  CYBERNETIC SAMURAI  MILAN, VICTOR
2  CYBERPUNK 2.0.2.0.  BILLIAS, STEPHEN
2  CYBORG  CAIDIN, MARTIN
3  CYBORG COMMANDO  O’NEILL, PAMELA/MOHAN, KIM
4  CYCLE OF FIRE  WURTS, JANNY
2  CYGNET  MCKILLIP, PATRICIA A.
2  D’SHAI  ROSENBERG, JOEL
6  DAEDALUS  STABLEFORD, BRIAN M.
4  DALEMARK  JONES, DIANA WYNNE
2  DALRIADA  WATSON, JULES
3  DAMASTES  BUNCH, CHRIS
4  DAMIANO  MACAVOY, R. A.
3  DAMNED  FOSTER, ALAN DEAN
3  DANANS  TAYLOR, KEITH
4  DANCE OF GODS  BRENNER, MAYER ALAN
3  DANCE OF THE RINGS  FANCHER, JANE S.
3  DANCER  CLAYTON, JO
5  DANCERS AT THE END OF TIME  MOORCOCK, MICHAEL
5  DANCING GODS  CHALKER, JACK L.
3  DANGEROUS JOURNEYS  GYGAX, GARY
2  DANIELLE WALTHERS  CIENCIN, SCOTT
5  DANNUS  SIROTA, MIKE
15  DANNY DUNN  WILLIAMS, JAY/ABRASHKIN, RAYMOND
2  DANTE VALENTINE  SAINTCROW, LILITH
1  DAREDEVIL  GOLDEN, CHRISTOPHER
3  DARIAN  LACKEY, MERCEDES/DIXON, LARRY
3  DARK ANGEL  COLLINS, MAX ALLAN
4  DARK BORDER  ZIMMER, PAUL EDWIN
1  DARK CHRONICLES  SOROKA, CYNTHIA
3  DARK CONSPIRACY  STACKPOLE, MICHAEL A.
2  DARK DESCENT  HARTWELL, DAVID G.
1  DARK ENCHANTMENT  COOPER, LOUISE
8  DARK FORCES  authorship varies
3  DARK FUTURE  authorship varies
1  DARK FUTURE AMERICAN MEAT  MOORE, STUART
2  DARK FUTURE DEMON DOWNLOAD  YEOVIL, JACK
1  DARK FUTURE EDDIE KALISH  STONE, DAVE
7  DARK HORSE  HERBERT, MARY H.
5  DARK IS RISING  COOPER, SUSAN
4  DARK MATTER  authorship varies
1  DARK MOON  BLAIR, CYNTHIA
5  DARK SUN ATHAS  authorship varies
5  DARK SUN PRISM  DENNING, TROY
3  DARK SUN TRIBE OF ONE  HAWKE, SIMON
9  DARK TOWER  authorship varies
3  DARK-HUNTER  KENYON, SHERRILYN
4  DARKANGEL  PIERCE, MEREDITH ANN
2  DARKGLASS MOUNTAIN  DOUGLASS, SARA
3  DARKMAN  BOYLL, RANDALL
43  DARKOVER  authorship varies
4  DARKSWORD  WEIS, MARGARET/HICKMAN, TRACY
5  DARKWAR  authorship varies
5  DARWATH  HAMBLY, BARBARA
1  DAUGHTER OF THE LIONESS  PIERCE, TAMORA
3  DAUGHTERS OF BAST  ISIDORE, SARAH
3  DAVID LYDYARD  STABLEFORD, BRIAN M.
3  DAYWORLD  FARMER, PHILIP JOSE
12  DC COMICS  authorship varies
2  DEAD  CALDER, RICHARD
2  DEAD RIVERS  KRITZER, NAOMI
4  DEATH DEALER  SILKE, JAMES
8  DEATH GATE  WEIS, MARGARET/HICKMAN, TRACY
2  DEATHDAY  DIETZ, WILLIAM C.
39  DEATHLANDS  authorship varies
1  DEATHLESS  MITCHELL, SYNE
9  DEATHSTALKER  GREEN, SIMON R.
1  DEATHWALKER  CASTEVANO, ROMAN
2  DECOY PRINCESS  COOK, DAWN
2  DEEP  VINGE, VERNOR
1  DEEPGATE  CAMPBELL, ALAN
1  DELTA GREEN  DETWILLER, DENNIS
2  DEMON BLADE  GARLAND, MARK A./MCGRAW, CHARLES G.
3  DEMON CROWN  VARDEMAN, ROBERT E.
11  DEMON PRINCES  VANCE, JACK

ENTRIES  SERIES  AUTHOR
3  DEMON WARS  MCCONNELL, ASHLEY
1  DEMON-BOUND  UNDERWOOD, LAURA J.
3  DEMONTECH  SHERMAN, DAVID
4  DEMONWARS  SALVATORE, R. A.
3  DEMONWARS II  SALVATORE, R. A.
3  DEMU  BUSBY, F. M.
3  DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS  RYAN, C. J.
2  DEPTFORD HISTORIES  JARVIS, ROBIN
1  DEPTFORD MICE  JARVIS, ROBIN
2  DERKHOLM  JONES, DIANA WYNNE
7  DERYNI  authorship varies
3  DERYNI CAMBER  KURTZ, KATHERINE
4  DERYNI CHRONICLES  KURTZ, KATHERINE
3  DERYNI HEIRS  KURTZ, KATHERINE
4  DERYNI KELSON  KURTZ, KATHERINE
3  DESCENT  TELEP, PETER
2  DESTINIES OF BLOOD AND STONE  NEWCOMB, ROBERT
5  DESTINY MAKERS  SHUPP, MIKE
3  DESTINY’S CHILDREN  BAXTER, STEPHEN
1  DESTROYER  SAPIR, RICHARD BEN/MURPHY, WARREN
2  DETECTIVE INSPECTOR CHEN  WILLIAMS, LIZ
15  DEVERRY  KERR, KATHARINE
6  DEVERRY PROPER  KERR, KATHARINE
1  DEVIL DANCES  JUBERT, HERVE
2  DEVIL’S WEST  SUMNER, MARK
4  DIABLO  authorship varies
2  DIABLO SIN WAR  KNAAK, RICHARD A.
12  DIADEM  CLAYTON, JO
2  DIANA SANTEE  GREEN, SHARON
2  DIANA TREGARDE  LACKEY, MERCEDES
2  DIEGO  FENN, LIONEL
3  DIES IRAE  STABLEFORD, BRIAN M.
5  DIG ALLEN  GREENE, JOSEPH
3  DINGILLIAD  GERROLD, DAVID
3  DINOSAUR EMPIRE  SAWYER, ROBERT J.
4  DINOSAUR PLANET  MCCAFFREY, ANNE
2  DINOTOPIA  FOSTER, ALAN DEAN
2  DION HOLYHANDS  ROUTLEY, JANE
5  DIPPLE  NORTON, ANDRE
6  DIRIGENT MERCENARY CORPS  SHELLEY, RICK
2  DIRK GENTLY  ADAMS, DOUGLAS
1  DIRK PITT  CUSSLER, CLIVE
4  DIRSHAN  LANCOUR, GENE
40  DISCWORLD  authorship varies
4  DISCWORLD DEATH  PRATCHETT, TERRY
7  DISCWORLD RINCEWIND  PRATCHETT, TERRY
3  DISCWORLD TIFFANY ACHING  PRATCHETT, TERRY
7  DISCWORLD WATCHMEN  PRATCHETT, TERRY
7  DISCWORLD WITCHES  PRATCHETT, TERRY
2  DISINHERITED  WHITE, STEVE
2  DIVINE ENDURANCE  JONES, GWYNETH
148  DOC SAVAGE  authorship varies
2  DOC SIDHE  ALLSTON, AARON
2  DOCTOR DIMENSION  DE CHANCIE, JOHN/BISCHOFF, DAVID F.
6  DOCTOR DOLITTLE  LOFTING, HUGH
1  DOCTOR STRANGE  ROTSLER, WILLIAM
254  DOCTOR WHO  authorship varies
71  DOCTOR WHO EIGHTH DOCTOR  authorship varies
33  DOCTOR WHO MISSING ADVENTURES  authorship varies
61  DOCTOR WHO NEW ADVENTURES  authorship varies
24  DOCTOR WHO NEW ADVENTURES (DOCTORLESS)  authorship varies
1  DOCTOR WHO NINTH (ANIMATED) DOCTOR  CORNELL, PAUL
6  DOCTOR WHO NINTH DOCTOR  authorship varies
74  DOCTOR WHO PAST DOCTORS  authorship varies
4  DOCTOR WHO TENTH DOCTOR  authorship varies
3  DOLPHIN  MEYERS, ROY
12  DOMINIC FLANDRY  ANDERSON, POUL
3  DOMINIONS OF IRTH  LEE, ADAM
5  DON SEBASTIAN  DANIELS, LES
4  DOOM  AB HUGH, DAFYDD/LINAWEAVER, BRAD
3  DOOM STAR  MEYERS, RICHARD S.
2  DOOM’S DAY  authorship varies
2  DOOM-QUEST OF ARA-KARN  CORBY, ADAM
11  DOOMSDAY WARRIOR  STACY, RYDER

ENTRIES  SERIES  AUTHOR
3  DOONA  authorship varies
3  DOUBLE SPIRAL WAR  NORWOOD, WARREN
2  DOWNS-LORD  WHITBOURN, JOHN
6  DR. BONES  authorship varies
18  DRACULA  authorship varies
3  DRAGAERA ADRILANKHA  BRUST, STEVEN
1  DRAGAERA FENARIO  BRUST, STEVEN
2  DRAGAERA KHAAVREN  BRUST, STEVEN
14  DRAGAERA VLAD TALTOS  authorship varies
3  DRAGON CIRCLE  GARDNER, CRAIG SHAW
2  DRAGON COMPANION  CALLANDER, DON
5  DRAGON DELASANGRE  TROOP, ALAN F.
1  DRAGON DICE  VARNEY, ALLEN
2  DRAGON HEART  KOKE, JAK
4  DRAGON JOUSTERS  LACKEY, MERCEDES
3  DRAGON KING  LAWHEAD, STEPHEN R.
9  DRAGON KNIGHT  DICKSON, GORDON R.
1  DRAGON KNIGHTS  D’ARC, BIANCA
3  DRAGON MAGE  KERR, KATHARINE
4  DRAGON NIMBUS  RADFORD, IRENE
3  DRAGON NIMBUS HISTORY  RADFORD, IRENE
1  DRAGON OF THE LOST SEA  YEP, LAURENCE
3  DRAGON PRINCE  RAWN, MELANIE
4  DRAGON QUARTET  KELLOGG, MARJORIE BRADLEY
3  DRAGON STAR  RAWN, MELANIE
3  DRAGON TEMPLE  CROSS, JANINE
3  DRAGON THRONE  BAIRD, ALISON
3  DRAGON WARS  GENTILE, GARY
2  DRAGON’S COUNTY  LYNN, ELIZABETH A.
1  DRAGON’S HEIRS  JONES, COURTWAY
3  DRAGONBACK  ZAHN, TIMOTHY
5  DRAGONCROWN WAR  STACKPOLE, MICHAEL A.
14  DRAGONLANCE  authorship varies
6  DRAGONLANCE AGE OF MORTALS  authorship varies
3  DRAGONLANCE BARBARIANS  THOMPSON, PAUL B./COOK, TONYA C.
2  DRAGONLANCE BRIDGES OF TIME  authorship varies
3  DRAGONLANCE CHAMPIONS  authorship varies
5  DRAGONLANCE CHAOS WAR  authorship varies
6  DRAGONLANCE CHRONICLES  authorship varies
4  DRAGONLANCE CLASSICS  authorship varies
4  DRAGONLANCE CROSSROADS  authorship varies
2  DRAGONLANCE DARK DISCIPLE  WEIS, MARGARET
3  DRAGONLANCE DEFENDERS OF MAGIC  KIRCHOFF, MARY
3  DRAGONLANCE DHAMON  RABE, JEAN
4  DRAGONLANCE DRAGON QUARTET  authorship varies
3  DRAGONLANCE DWARVEN NATIONS  PARKINSON, DAN
2  DRAGONLANCE ELVEN EXILES  THOMPSON, PAUL B./COOK, TONYA C.
3  DRAGONLANCE ELVEN NATIONS  authorship varies
3  DRAGONLANCE ERGOTH  THOMPSON, PAUL B./COOK, TONYA C.
4  DRAGONLANCE FIFTH AGE  authorship varies
1  DRAGONLANCE GOODLUND  SULLIVAN, STEPHEN D.
3  DRAGONLANCE HEROES  authorship varies
3  DRAGONLANCE HEROES II  authorship varies
3  DRAGONLANCE ICEWALL  NILES, DOUGLAS
1  DRAGONLANCE KANG’S REGIMENT  PERRIN, DON/WEIS, MARGARET
3  DRAGONLANCE KINGPRIEST  PIERSON, CHRIS
3  DRAGONLANCE LEGENDS  WEIS, MARGARET/HICKMAN, TRACY
3  DRAGONLANCE LINSHA  HERBERT, MARY H.
2  DRAGONLANCE LOST CHRONICLES  WEIS, MARGARET/HICKMAN, TRACY
6  DRAGONLANCE LOST HISTORIES  authorship varies
2  DRAGONLANCE LOST LEGENDS  authorship varies
6  DRAGONLANCE MEETINGS  authorship varies
3  DRAGONLANCE MINOTAUR WARS  KNAAK, RICHARD A.
8  DRAGONLANCE NEW ADVENTURES  authorship varies
1  DRAGONLANCE NEW ADVENTURES ELEMENTS  SOESBEE, REE
3  DRAGONLANCE NEW ADVENTURES ELIDOR  SOESBEE, REE
2  DRAGONLANCE NEW ADVENTURES SUNCATCHER  SAMPSON, JEFF
3  DRAGONLANCE NEW ADVENTURES TRINISTYR  authorship varies
3  DRAGONLANCE PRELUDES  authorship varies
3  DRAGONLANCE PRELUDES II  authorship varies
2  DRAGONLANCE RAISTLIN  authorship varies
3  DRAGONLANCE RISE OF SOLAMNIA  NILES, DOUGLAS
4  DRAGONLANCE SPELLBINDER QUARTET  authorship varies

ENTRIES  SERIES  AUTHOR
3  DRAGONLANCE TALADAS  PIERSON, CHRIS
3  DRAGONLANCE TALES  WEIS, MARGARET/HICKMAN, TRACY
3  DRAGONLANCE TALES II  TSR
5  DRAGONLANCE VILLAINS  authorship varies
1  DRAGONLANCE VILLIANS  WILLIAMS, MICHAEL/WILLIAMS, TERI
5  DRAGONLANCE WAR OF SOULS  authorship varies
7  DRAGONLANCE WARRIORS  authorship varies
3  DRAGONLORD LINDEN  BERTIN, JOANNE
3  DRAGONMASTER  BUNCH, CHRIS
6  DRAGONREALM  KNAAK, RICHARD A.
3  DRAGONREALM ORIGINS  KNAAK, RICHARD A.
5  DRAGONS  RINGO, JOHN
2  DRAGONS OF THE CUYAHOGA  SWANN, S. ANDREW
3  DRAGONSPELL  PAUL, DONITA K.
3  DRAGONSTAR  BISCHOFF, DAVID F./MONTELEONE, THOMAS F.
3  DRAGONSWORD  BAUDINO, GAEL
1  DRAGONTALES  VILOTT, RHONDI
2  DRAGONVARLD  WEIS, MARGARET
5  DRAKA  STIRLING, S. M.
1  DRAKA (REVISED)  STIRLING, S. M.
3  DRAKE MAIJSTRAL  WILLIAMS, WALTER JON
37  DRAY PRESCOT  AKERS, ALAN BURT
5  DREAD EMPIRE  COOK, GLEN
4  DREAD EMPIRE’S FALL  WILLIAMS, WALTER JON
3  DREAM LORDS  COLE, ADRIAN
3  DREAM PARK  NIVEN, LARRY/BARNES, STEVEN
5  DREAMERS  EDDINGS, DAVID/EDDINGS, LEIGH
2  DREAMERS OF THE DAY  ROSS, DAVID D.
2  DREAMHUNTER  KNOX, ELIZABETH
4  DREAMLAND  authorship varies
2  DREAMQUEST  SAINT ALCORN, LLOYD
1  DREAMTIME  NICHOLLS, STAN
2  DREAMWORLD  LUMLEY, BRIAN
6  DRENAI  GEMMELL, DAVID
1  DRENAI 1 (ORIGINAL)  GEMMELL, DAVID
1  DRENAI 1 (REVISED)  GEMMELL, DAVID
2  DRENAI WAYLANDER  GEMMELL, DAVID
9  DRESDEN FILES  BUTCHER, JIM
2  DREWCILA QWAH  ROSEN, SELINA
3  DRIFTER  DIETZ, WILLIAM C.
2  DRIVE-IN  LANSDALE, JOE R.
3  DRUMS OF CHAOS  authorship varies
5  DRYCO  WOMACK, JACK
1  DUBRIC BYERLY  JONES, TAMARA SILER
3  DUEL OF SORCERY  CLAYTON, JO
33  DUMAREST  TUBB, E. C.
2  DUN COW  WANGERIN, WALTER JR.
4  DUN LADY’S JESS  DURGIN, DORANNA
13  DUNE  authorship varies
6  DUNE JIHAD  HERBERT, BRIAN/ANDERSON, KEVIN J.
6  DUNGEON  authorship varies
11  DUNGEONS & DRAGONS  authorship varies
3  DUNGEONS & DRAGONS PENHALIGON  HEINRICH, D. J.
2  DUNWORTHY & FINCH  WILLIS, CONNIE
1  DURHAM RED  EVANS, PETER J.
2  DURHYS FAMILY  ISAAK, ELAINE
3  DUSHAU  LICHTENBERG, JACQUELINE
8  DYING EARTH  authorship varies
3  DYLAN MATHESON  LEE, J. ARDIAN
2  E-BRANCH  LUMLEY, BRIAN
2  EAGLEHEART  WESTCOTT, C. T.
1  EARTH  VORNHOLT, JOHN
5  EARTH FINAL CONFLICT  authorship varies
3  EARTH SONG  WEBB, SHARON
4  EARTH’S CHILDREN  AUEL, JEAN M.
3  EARTHBLOOD  AXLER, JAMES
1  EARTHBORN WARS  COLLINS, PAUL
5  EARTHDAWN  authorship varies
7  EARTHSEA  LE GUIN, URSULA K.
2  EARTH 2  authorship varies
4  EBENEZUM  GARDNER, CRAIG SHAW
1  EBERRON  SEHESTEDT, MARK
2  EBERRON BLADE OF THE FLAME  WAGGONER, TIM
3  EBERRON DRAGON BELOW  BASSINGTHWAITE, DON

ENTRIES  SERIES  AUTHOR
3  EBERRON DREAMING DARK  BAKER, KEITH
2  EBERRON HEIRS OF ASH  WULF, RICH
3  EBERRON INQUISITIVES  authorship varies
1  EBERRON LANTERNLIGHT  DEWOLF, PARKER
3  EBERRON LOST MARK  FORBECK, MATT
4  EBERRON WAR-TORN  authorship varies
3  ECHOES OF EARTH  WILLIAMS, SEAN/DIX, SHANE
6  ECOLITAN  MODESITT, L. E. JR.
2  EDDIE BEAR  RANKIN, ROBERT
3  EDEN  HARRISON, HARRY
2  EDEN MOORE  PRIEST, CHERIE
5  EINARINN  MCKENNA, JULIET E.
3  EL ISKANDRYIA  GRIMWOOD, JON COURTENAY
4  ELAKI  HIGHTOWER, LYNN S.
3  ELANTRA  SAGARA, MICHELLE
3  ELEMENTAL LOGIC  MARKS, LAURIE J.
4  ELEMENTAL MASTERS  LACKEY, MERCEDES
3  ELENIUM  EDDINGS, DAVID
1  ELFWOOD  STEIN, KEVIN
2  ELIANA  KRITZER, NAOMI
4  ELITA  JACOBY, KATE
3  ELIXIR OF LIFE  FRANKOS, STEVEN
1  ELM STREET  COOPER, JEFFREY
19  ELRIC  authorship varies
5  ELUNDIUM  JEFFERIES, MIKE
3  ELVEN KINGDOMS  COOKE, CATHERINE
4  ELVEN WAYS  authorship varies
4  ELYSIUM  SLONCZEWSKI, JOAN
3  EMANCIPATOR  ALDRIDGE, RAY
3  EMBERVERSE  STIRLING, S. M.
4  EMMA MERRIGAN  ABBEY, LYNN
6  EMORTALITY  STABLEFORD, BRIAN M.
1  EMPIRE OF ELWOLF  CARRADINE, F. J.
5  EMPIRE OF THE EAST  SABERHAGEN, FRED
5  EMPIRE PRINCESS  DIAMOND, GRAHAM
2  EMPYRION  LAWHEAD, STEPHEN R.
4  ENCHANTED FOREST  WREDE, PATRICIA C.
12  ENDER  CARD, ORSON SCOTT
1  ENDLESS FRONTIER  POURNELLE, JERRY/CARR, JOHN F.
2  ENDLESS QUEST  authorship varies
2  ENDWORLD  ROBBINS, DAVID
4  ENEMIES OF SOCIETY  DAVID, JOHN
2  ENEMY GLORY  MICHALSON, KAREN
6  ENFORCER  SUGAR, ANDREW
3  ENGINES OF LIGHT  MACLEOD, KEN
1  ENTIRE AND ROSE  KENYON, KAY
1  ENVIADO SECRETO  CAUDETT, FRANK
3  EON  BEAR, GREG
2  EPHEMERA  BISHOP, ANNE
3  ER-DA  REDMOND, JEFFREY
3  EREKOSE  MOORCOCK, MICHAEL
8  ERIC BANYON  authorship varies
2  ERIC BRIGHTEYES  authorship varies
5  ERIC CARSTAIRS  CARTER, LIN
3  ERIC MATHERS  MEREDITH, RICHARD C.
3  ERIS  SMITH, L. NEIL
4  ERON  DEITZ, TOM
4  ERTHRING  DREW, WAYLAND
3  ESCHATON  POHL, FREDERIK
1  ESTHER DIAMOND  RESNICK, LAURA
1  ETERNAL ARCHIVES  WEIS, MARGARET
6  ETERNAL CHAMPION  MOORCOCK, MICHAEL
3  ETERNAL GUARDIANS  CROSS, RONALD ANTHONY
1  ETERNAL WARRIORS  BEALE, THEODORE
3  ETHAN STARK  HEMRY, JOHN G.
3  ETHICAL SOCIETY  TURNER, GEORGE
9  ETHSHAR  WATT-EVANS, LAWRENCE
10  EVE DALLAS  authorship varies
3  EVENMERE  STODDARD, JAMES
1  EVER  SAVAGE, FELICITY
4  EVERGENCE  WILLIAMS, SEAN/DIX, SHANE
3  EVERIEN  LEITH, VALERY
1  EVERWORLD  APPLEGATE, K. A.
4  EXALTED  authorship varies

ENTRIES  SERIES  AUTHOR
3  EXALTED SECOND AGE  DANSKY, RICHARD E.
3  EXCHAMELEON  GOULART, RON
4  EXILES  authorship varies
1  EXILES OF BOQ’URAIN  CRISPIN, A. C.
2  EXORCIST  BLATTY, WILLIAM PETER
5  EXORDIUM  SMITH, SHERWOOD/TROWBRIDGE, DAVE
4  EXPENDABLES  AVERY, RICHARD
4  EXPLORATION  RANDLE, KEVIN
3  EYES OF GOD  MARCO, JOHN
2  F.R.E.E.LANCERS  ODOM, MEL
6  FACTION PARADOX  authorship varies
4  FADED SUN  CHERRYH, C. J.
2  FADING WORLDS  GRAVEL, GEARY
14  FAFHRD  authorship varies
3  FAIA RISSEDOTTE  LISLE, HOLLY
1  FAIRY REALM  RODDA, EMILY
1  FALCON GOLD  WOLFF, ROBERT SHERMAN
4  FALKENBERG’S LEGION  authorship varies
3  FALL OF ILE-RIEN  WELLS, MARTHA
1  FALL OF NIGHT  BENFORD, GREGORY
3  FALL OF THE FIRST WORLD  SMITH, DAVID C.
4  FALL OF THE TOWERS  DELANY, SAMUEL R.
5  FALL REVOLUTION  MACLEOD, KEN
4  FALLEN CLOUD  GIAMBASTIANI, KURT R. A.
3  FALLERS  KRESS, NANCY
2  FAMILY  MCCARTHY, KEVIN/SILVA, DAVID
10  FAMILY D’ALEMBERT  authorship varies
2  FANTASTIC FOUR  authorship varies
1  FANTASY ADVENTURES  EMERY, CLAYTON
6  FANUILH  HOOD, DANIEL
2  FAR SIDE OF FOREVER  GREEN, SHARON
3  FAR STARS AND FUTURE TIMES  MCENROE, RICHARD S.
2  FARADAWN  FORD, RICHARD
2  FARSALA  BELL, HILARI
4  FARSCAPE  authorship varies
3  FARSEER  HOBB, ROBIN
2  FARSTAR & SON  STARR, BILL
2  FAT WHITE VAMPIRE (JULES DUCHON)  FOX, ANDREW
3  FAULT LINES  POWERS, TIM
3  FEAR  CHEPAITIS, B. A.
2  FELINE WIZARDS  DUANE, DIANE
3  FENCER  PARKER, K. J.
3  FERN CHAPEL  SIEGEL, JAN
5  FEY  RUSCH, KRISTINE KATHRYN
3  FIFTH FOREIGN LEGION  authorship varies
8  FIFTH MILLENNIUM  authorship varies
2  FINN MACCUMHAL  FLINT, KENNETH C.
3  FINNBRANCH  HAZEL, PAUL
3  FIONAVAR TAPESTRY  KAY, GUY GAVRIEL
3  FIONN MAC CUMHAL  PAXSON, DIANA L./MARTINE-BARNES, ADRIENNE
1  FIRE-US  ARMSTRONG, JENNIFER/BUTCHER, NANCY
4  FIREBIRD  TYERS, KATHY
4  FIREBRATS  SIEGEL, BARBARA/SIEGEL, SCOTT
2  FIREBRINGER  PIERCE, MEREDITH ANN
6  FIREKEEPER  LINDSKOLD, JANE
1  FIRST AMERICANS  SARABONDE, WILLIAM
4  FIRST TRUTH  COOK, DAWN
3  FIVE COUNTRIES  WELCH, MICHELLE M.
3  FIVE HUNDRED KINGDOMS  LACKEY, MERCEDES
7  FIVE RINGS CLAN WAR  authorship varies
5  FIVE RINGS FOUR WINDS  authorship varies
3  FIVE WORLDS  SARRANTONIO, AL
12  FLASH GORDON  authorship varies
2  FLASHMAN  FRASER, GEORGE MACDONALD
7  FLAT EARTH  LEE, TANITH
2  FLATLAND  authorship varies
6  FLEET  DRAKE, DAVID/FAWCETT, BILL
2  FLEETS OF THE GREAT BATTLE  ETHERIDGE, RUTLEDGE
3  FLESH AND GOLD  GOTLIEB, PHYLLIS
3  FLIGHT ENGINEER  DOOHAN, JAMES/STIRLING, S. M.
5  FLYING NUN  JOHNSTON, WILLIAM
3  FOOL’S GOLD  FISHER, JUDE
2  FOOTPRINTS OF THUNDER  DAVID, JAMES F.
3  FORBIDDEN BORDERS  GEAR, W. MICHAEL

ENTRIES  SERIES  AUTHOR
5  FORERUNNER  NORTON, ANDRE
4  FOREVER HERO  MODESITT, L. E. JR.
3  FOREVER KING  authorship varies
2  FOREVER KNIGHT  authorship varies
2  FORGE OF GOD  BEAR, GREG
23  FORGOTTEN REALMS  authorship varies
3  FORGOTTEN REALMS ARCANE AGE NETHERIL  EMERY, CLAYTON
8  FORGOTTEN REALMS AVATAR  authorship varies
1  FORGOTTEN REALMS BALDUR’S GATE I  ATHANS, PHILIP
2  FORGOTTEN REALMS BALDUR’S GATE II  authorship varies
4  FORGOTTEN REALMS CITIES  authorship varies
5  FORGOTTEN REALMS CLERIC  SALVATORE, R. A.
3  FORGOTTEN REALMS CORMYR  authorship varies
3  FORGOTTEN REALMS COUNSELORS & KINGS  CUNNINGHAM, ELAINE
4  FORGOTTEN REALMS DARK ELF  SALVATORE, R. A.
4  FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE  authorship varies
3  FORGOTTEN REALMS DRUIDHOME  NILES, DOUGLAS
2  FORGOTTEN REALMS DUNGEONS  authorship varies
4  FORGOTTEN REALMS ELMINSTER  GREENWOOD, ED
3  FORGOTTEN REALMS EMPIRES  authorship varies
1  FORGOTTEN REALMS EMPYREAN ODYSSEY  REID, THOMAS M.
3  FORGOTTEN REALMS EREVIS CALE  KEMP, PAUL S.
4  FORGOTTEN REALMS FIGHTERS  authorship varies
3  FORGOTTEN REALMS FINDER’S STONE  NOVAK, KATE/GRUBB, JEFF
16  FORGOTTEN REALMS HARPERS  authorship varies
1  FORGOTTEN REALMS HAUNTED LANDS  BYERS, RICHARD LEE
3  FORGOTTEN REALMS HOUSE OF SERPENTS  SMEDMAN, LISA
3  FORGOTTEN REALMS HUNTER’S BLADES  SALVATORE, R. A.
3  FORGOTTEN REALMS ICEWIND DALE  SALVATORE, R. A.
1  FORGOTTEN REALMS KNIGHTS OF MYTH DRANNOR  GREENWOOD, ED
1  FORGOTTEN REALMS LADY PENITENT  SMEDMAN, LISA
4  FORGOTTEN REALMS LAST MYTHAL  authorship varies
4  FORGOTTEN REALMS LEGACY OF THE DROW  SALVATORE, R. A.
4  FORGOTTEN REALMS LOST EMPIRES  authorship varies
1  FORGOTTEN REALMS LOST GODS  NOVAK, KATE/GRUBB, JEFF
3  FORGOTTEN REALMS MAZTICA  NILES, DOUGLAS
3  FORGOTTEN REALMS MOONSHAE  NILES, DOUGLAS
6  FORGOTTEN REALMS NOBLES  authorship varies
1  FORGOTTEN REALMS PATHS OF DARKNESS  SALVATORE, R. A.
3  FORGOTTEN REALMS PRIESTS  authorship varies
4  FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS  authorship varies
4  FORGOTTEN REALMS ROGUES  authorship varies
1  FORGOTTEN REALMS RUINS OF MYTH DRANNOR  BEBRIS, CARRIE A.
3  FORGOTTEN REALMS SCIONS OF ARRABAR  REID, THOMAS M.
3  FORGOTTEN REALMS SELLSWORDS  SALVATORE, R. A.
7  FORGOTTEN REALMS SEMBIA  authorship varies
1  FORGOTTEN REALMS SEMBIA GATEWAY TO THE REALMS  ATHANS, PHILIP
3  FORGOTTEN REALMS SHADOW OF AVATAR  GREENWOOD, ED
2  FORGOTTEN REALMS SHANDRIL  GREENWOOD, ED
3  FORGOTTEN REALMS STARLIGHT AND SHADOWS  CUNNINGHAM, ELAINE
4  FORGOTTEN REALMS THREAT FROM THE SEA  authorship varies
3  FORGOTTEN REALMS TWILIGHT GIANTS  DENNING, TROY
1  FORGOTTEN REALMS TWILIGHT WAR  KEMP, PAUL
1  FORGOTTEN REALMS UNDERDARK  KING, J. ROBERT
6  FORGOTTEN REALMS WAR OF THE SPIDER QUEEN  authorship varies
2  FORGOTTEN REALMS WATERCOURSE  ATHANS, PHILIP
4  FORGOTTEN REALMS WIZARDS  authorship varies
5  FORGOTTEN REALMS YEAR OF ROGUE DRAGONS  authorship varies
2  FORTUNE  EDWARDS, GRAHAM
3  FORTY SIGNS OF RAIN  ROBINSON, KIM STANLEY
12  FOUNDATION  authorship varies
1  FOUNDATIONS OF HOPE  GRODY, LEATHEL
5  FOUR LORDS OF THE DIAMOND  CHALKER, JACK L.
1  FOURLANDS  SWAINSTON, STEPH
2  FOURTH REALM  TWELVE HAWKS, JOHN
2  FOX  CHASE, ROBERT R.
3  FRANK CARLUCCI  RUSSO, RICHARD PAUL
6  FREEDOM’S RANGERS  ANDREWS, KEITH WILLIAM
2  FREEHOLD WAR  WILLIAMSON, MICHAEL Z.
2  FRONTIER EARTH  BOXLEITNER, BRUCE
2  FRONTIERS  ELWOOD, ROGER
3  FROST  BAILEY, ROBIN WAYNE
2  FROSTFLOWER  KARR, PHYLLIS ANN

ENTRIES  SERIES  AUTHOR
3  FUTURE AT WAR  BRETNOR, REGINALD
2  FUTURE EARTHS  RESNICK, MIKE/DOZOIS, GARDNER
7  FUZZIES  authorship varies
8  G-8  HOGAN, ROBERT J.
3  G.O.D. INC.  CHALKER, JACK L.
1  GABRIEL KNIGHT  JENSEN, JANE
3  GAEA  VARLEY, JOHN
7  GALACTIC CENTER  BENFORD, GREGORY
4  GALACTIC EMPIRE  ASIMOV, ISAAC
3  GALACTIC MI  RANDLE, KEVIN
4  GALACTIC MIDWAY  RESNICK, MIKE
3  GALACTIC MILIEU  MAY, JULIAN
1  GALAHAD 1  TESTA, DOM
2  GALEN SWORD  REEVES-STEVENS, JUDITH/REEVES-STEVENS, GARFIELD
3  GAMANT  authorship varies
3  GAMBLER’S STAR  HOLDER, NANCY
2  GAMEARTH  ANDERSON, KEVIN J.
3  GAMESTER WARS  FORSTCHEN, WILLIAM R.
3  GAMING MAGI  BISCHOFF, DAVID F.
1  GAMMA QUEST  COX, GREG
4  GAMMALAW  DALEY, BRIAN
9  GANDALARA  GARRETT, RANDALL/HEYDRON, VICKI ANN
5  GAP  DONALDSON, STEPHEN R.
2  GARDNER  MANN, PHILLIP
3  GARRETH MIKAELIAN  KILLOUGH, LEE
15  GARRETT  COOK, GLEN
2  GARRISON  authorship varies
3  GATES OF TIME  PARKINSON, DAN
1  GAUNTLET DARK LEGACY  WHITE, RICHARD C.
2  GENE WARS  CHERRYH, C. J.
3  GENELLAN  GIER, SCOTT G.
9  GENERAL  authorship varies
2  GENESIS  MOFFITT, DONALD
2  GEODESICA  WILLIAMS, SEAN/DIX, SHANE
1  GEODESSEY  ANTHONY, PIERS
3  GEODYSSEY  ANTHONY, PIERS
1  GEORGIE AND BUG  RUBY, LAURA
5  GERALD KNAVE  JANIFER, LAURENCE M.
3  GERIN THE FOX  TURTLEDOVE, HARRY
1  GERIN THE FOX (REVISED)  TURTLEDOVE, HARRY
9  GET SMART  JOHNSTON, WILLIAM
3  GETTYSBURG  GINGRICH, NEWT/FORSTCHEN, WILLIAM R.
2  GHATTENS’ GAMBIT  GREENO, GAYLE
3  GHATTI’S TALE  GREENO, GAYLE
2  GHOST HOUSE  MCNALLY, CLAIRE
1  GHOST HUNTER  LAURIA, VICTORIA
1  GHOST IN THE SHELL 2  YAMADA, MASAKI
2  GHOSTBUSTERS  authorship varies
3  GHOSTS OF ALBION  BENSON, AMBER/GOLDEN, CHRISTOPHER
2  GHOSTWALKERS  FEEHAN, CHRISTINE
7  GIANTS  HOGAN, JAMES P.
2  GILGAMESH  SILVERBERG, ROBERT
2  GIRD  MOON, ELIZABETH
3  GIRL FACTORY  MURPHY, ROBERT FRANKLIN
4  GIRL FROM UNCLE  authorship varies
5  GIRL GENIUS  FOGLIO, PHIL/FOGLIO, KAJA
3  GLASSWRIGHT  KLASKY, MINDY L.
2  GLENRAVEN  BRADLEY, MARION ZIMMER/LISLE, HOLLY
3  GLOBAL 2000  BASILE, GLORIA VITANZA
3  GNOME MAN’S LAND  FRIESNER, ESTHER M.
2  GNOMEWRENCH  O’DONOHOE, NICK
2  GOBLIN MOON  EDGERTON, TERESA
3  GODMOTHER  SCARBOROUGH, ELIZABETH ANN
6  GODS  SABERHAGEN, FRED
2  GODS OF IRELAND  FLYNN, CASEY
2  GODSLAYER  CLEMENS, JAMES
3  GODWARS  WELLS, ANGUS
1  GOLDBANE  WILSON, JOHN
3  GOLDEN AGE  WRIGHT, JOHN C.
4  GOLDEN QUEEN  authorship varies
3  GOLDENWING  COPPEL, ALFRED
2  GONARD  CYMRI, CHRYS
6  GONDWANE  CARTER, LIN
1  GONJI  RYPEL, T. C.

ENTRIES  SERIES  AUTHOR
2  GOOSEBUMPS  STINE, R. L.
26  GOR  NORMAN, JOHN
5  GORD THE ROGUE  GYGAX, GARY
3  GORMENGHAST  PEAKE, MERVYN
7  GRAINGER  STABLEFORD, BRIAN M.
3  GREAT GAME  DUNCAN, DAVE
3  GREAT IMPERIUM  CARTER, LIN
2  GREAT MACHINE  MCMULLEN, SEAN
3  GREAT TREE OF AVALON  BARRON, T. A.
5  GREEN KNOWE  BOSTON, L. M.
1  GREEN MILE  KING, STEPHEN
2  GREEN PLANET  PETAJA, EMIL
2  GREEN RIDER  BRITAIN, KRISTEN
5  GREEN STAR  CARTER, LIN
3  GREEN-SKY  SNYDER, ZILPHA KEATLEY
3  GREG MANDEL  HAMILTON, PETER F.
2  GREMLINS  authorship varies
15  GREYHAWK  authorship varies
3  GREYSTONE BAY  GRANT, CHARLES L.
38  GRIMES  CHANDLER, A. BERTRAM
1  GROUNDTIES  FANCHER, JANE S.
3  GUARDIAN  GRAY, JULIA
4  GUARDIAN-KING  HANCOCK, KAREN
9  GUARDIANS  authorship varies
5  GUARDIANS (SAXON)  SAXON, PETER
1  GUARDIANS OF GADIEL  MORGAN, KATHLEEN
9  GUARDIANS OF THE FLAME  ROSENBERG, JOEL
3  GUARDIANS OF THE FLAME CASTLE CULLINANE  ROSENBERG, JOEL
4  GUARDIANS OF THE THREE  authorship varies
3  GUINEVERE  NEWMAN, SHARAN
3  GUNDAM MOBILE SUIT  TOMINO, YOSHIYUKI
3  GWALCHMAI  BRADSHAW, GILLIAN
2  GYPSY  GOULART, RON
1  HAGWOOD  JARVIS, ROBIN
7  HAINISH  LE GUIN, URSULA K.
2  HAL SPACEJOCK  HAYNES, SIMON
1  HALCYON BLITHE  WARD, JAMES M.
3  HALFBLOOD  NORTON, ANDRE/LACKEY, MERCEDES
3  HALFLIFE  SIMMONS, WM. MARK
5  HALLOWED ISLE  PAXSON, DIANA L.
4  HALO  authorship varies
3  HAMMER AND THE CROSS  authorship varies
15  HAMMER’S SLAMMERS  authorship varies
3  HANAN REBELLION  CHERRYH, C. J.
4  HANDFUL OF MEN  DUNCAN, DAVE
2  HARAN OF OAS  TAFF, JOHN F. D.
2  HARI MICHAELSON  STOVER, MATTHEW WOODRING
2  HARMONY OF WORLDS  FREIREICH, VALERIE J.
2  HARRIERS  DICKSON, GORDON R.
4  HARRY BORG  HAWKINS, WARD
10  HARRY POTTER  ROWLING, J. K.
4  HARVEST OF STARS  ANDERSON, POUL
2  HATHER LAMBISSON  KING, BERNARD
1  HATTI EMPIRE  MORRIS, JANET
7  HAWK & FISHER  GREEN, SIMON R.
2  HAWK QUEEN  GEMMELL, DAVID
3  HEAVEN  VINGE, JOAN D.
6  HEECHEE  POHL, FREDERIK
2  HEIRS TO BYZANTIUM  SHWARTZ, SUSAN
1  HEIRS TO GNARLSMYRE  JEFFERIES, MIKE
3  HEL  DRAKE, ASA
2  HEL’S CRUCIBLE  MCKIERNAN, DENNIS L.
3  HELLFLOWER  BES SHAHAR, ELUKI
1  HELLGATE LONDON  ODOM, MEL
3  HELLICONIA  ALDISS, BRIAN W.
7  HELMSMAN  BALDWIN, BILL
2  HENGHIS HAPTHORN  HUGHES, MATTHEW
2  HENRY MARTYN  SMITH, L. NEIL
2  HEOROT  NIVEN, LARRY/POURNELLE, JERRY/BARNES, STEVEN
3  HERALD CHILDE  FARMER, PHILIP JOSE
4  HERALDS OF VALDEMAR  LACKEY, MERCEDES
2  HERCULES  BETANCOURT, JOHN GREGORY
2  HERCULES THE LEGENDARY JOURNEYS  authorship varies
3  HERE BE DEMONS  FRIESNER, ESTHER M.

ENTRIES  SERIES  AUTHOR
8  HERIS SERRANO  MOON, ELIZABETH
3  HERITAGE  DOUGLAS, IAN
6  HERITAGE UNIVERSE  SHEFFIELD, CHARLES
2  HERITAGE UNIVERSE (REVISED)  SHEFFIELD, CHARLES
12  HEROES IN HELL  authorship varies
3  HIDDEN REALMS  GREEN, SHARON
4  HIGH DRUID OF SHANNARA  BROOKS, TERRY
1  HIGH PILGRIMAGE  KOMAN, VICTOR
10  HIGHLANDER  authorship varies
3  HIGHROAD  RASMUSSEN, ALIS A.
4  HIS DARK MATERIALS  PULLMAN, PHILIP
3  HISTORICAL ILLUMINATUS  WILSON, ROBERT ANTON
2  HISTORY  WELLS, H. G.
12  HISTORY OF MIDDLE-EARTH  TOLKIEN, J. R. R.
14  HITCHHIKERS  authorship varies
3  HOBART FLOYT  DALEY, BRIAN
4  HOKA  ANDERSON, POUL/DICKSON, GORDON R.
5  HOLDERLY HALL  CAMERON, KATE
5  HOLDFAST  CHARNAS, SUZY MCKEE
7  HOMECOMING  CARD, ORSON SCOTT
19  HONOR HARRINGTON  authorship varies
1  HONOR HARRINGTON SAGANAMI  WEBER, DAVID
4  HOOK  ZETFORD, TULLY
3  HORIZON WAR  WEINBERG, ROBERT
3  HORN  SLOANE, BEN
4  HORRORSCOPE  LORY, ROBERT
3  HORSE GODDESS PEOPLE  TARR, JUDITH
18  HORSECLANS  ADAMS, ROBERT
2  HORSEGIRL  ASH, CONSTANCE
6  HOSTEEM STORM  authorship varies
3  HOSTILE TAKEOVER  SWANN, S. ANDREW
1  HOT BLOOD  GELB, JEFF/GARRETT, MICHAEL
4  HOUND AND FALCON  TARR, JUDITH
2  HOUSE OF EIGIN  LEIBER, JUSTIN
2  HOUSE OF MOONS  WENTWORTH, K. D.
1  HOWARD LIBRARY  HOWARD, ROBERT E.
3  HOWLING  BRANDNER, GARY
4  HRADANI  authorship varies
3  HUFF  HUFF, TANYA
3  HULK  authorship varies
9  HULZEIN DYNASTY  BUSBY, F. M.
8  HUMANX  FOSTER, ALAN DEAN
13  HUMANX FLINX  FOSTER, ALAN DEAN
3  HUMANX FOUNDING COMMONWEALTH  FOSTER, ALAN DEAN
5  HUMANX ICERIGGER  FOSTER, ALAN DEAN
3  HUMILITY GARDEN  SAVAGE, FELICITY
7  HUNGRY CITY  REEVE, PHILIP
2  HUNTED EARTH  ALLEN, ROGER MACBRIDE
4  HUNTER  authorship varies
1  HUNTER THE RECKONING  WIECK, STEWART
3  HUNTSMAN  HILL, DOUGLAS
5  HYPERION  SIMMONS, DAN
3  HYTHRUN  FALLON, JENNIFER
2  HYTHRUN WOLFBLADE  FALLON, JENNIFER
1  I SPY  TIGER, JOHN
3  ICE  FORSTCHEN, WILLIAM R.
3  ID4  authorship varies
3  IDLEWILD  SAGAN, NICK
1  ILARIO  GENTLE, MARY
2  ILIUM  SIMMONS, DAN
3  ILLUMINATION  MCGARRY, TERRY
4  ILLUMINATUS!  SHEA, ROBERT/WILSON, ROBERT ANTON
3  IMARO  SAUNDERS, CHARLES R.
2  IMMORTAL EYES  CASSADA, JACKIE
3  IMMORTALS  authorship varies
5  IMMORTALS (PIERCE)  PIERCE, TAMORA
2  IMMORTALS AFTER DARK  COLE, KRESLEY
3  IMPERIAL STARS  POURNELLE, JERRY
6  IMPERIUM  LAUMER, KEITH
2  IN FIRE FORGED  RESNICK, LAURA
7  INCARNATIONS OF IMMORTALITY  ANTHONY, PIERS
8  INCOMPLETE ENCHANTER  authorship varies
3  INCONSTANT MOON  ASKEGREN, PIERCE
1  INDIAN IN THE CUPBOARD  BANKS, LYNNE REID

ENTRIES  SERIES  AUTHOR
8  INDIANA JONES  authorship varies
8  INDIGO  COOPER, LOUISE
1  INDIW  KERNS, DANIEL R.
2  INGERSOL  CARROLL, JERRY JAY
2  INHERITANCE  PAOLINI, CHRISTOPHER
2  INHERITOR  BRADLEY, MARION ZIMMER
2  INITIATE BROTHER  RUSSELL, SEAN
6  INQUESTOR  SUCHARITKUL, SOMTOW
2  INSTRUMENTALITIES OF THE NIGHT  COOK, GLEN
6  INSTRUMENTALITY OF MANKIND  SMITH, CORDWAINER
2  INTEGRAL TREES  NIVEN, LARRY
3  INTERFACE  ADLARD, MARK
1  INTERPLANETARY SPY  MCEVOY, SETH
2  INTERSTELLAR PATROL  ANVIL, CHRISTOPHER
3  INTERVENTION  MAY, JULIAN
2  INTERZONE  authorship varies
6  INTO THE DARKNESS  TURTLEDOVE, HARRY
4  INVADERS  authorship varies
2  INVASION AMERICA  GOLDEN, CHRISTIE
1  INVASION OF THE BODY SNATCHERS  MCCARTHY, KEVIN/GORMAN, ED
1  INVASION: EARTH  CAVE, PETER
2  IRA VEDDER  BOND, EVELYN
3  IRON ANGEL  MORRESSY, JOHN
1  IRON DRAGONS  ESTES, ROSE
1  IRON MAN  ROTSLER, WILLIAM
3  IRON TOWER  MCKIERNAN, DENNIS L.
3  IROSHI  OSBORNE, CARY
4  ISAAC’S UNIVERSE  authorship varies
4  ISAVALTA  ZETTEL, SARAH
5  ISIS  authorship varies
2  ISLAND WARRIOR  authorship varies
4  ISLAND WORLDS  KOTANI, ERIC/ROBERTS, JOHN MADDOX
2  ISLANDIA  SAXTON, MARK
3  ISLES OF GLORY  LARKE, GLENDA
4  IVORY  EGAN, DORIS
3  JACK HONG  WU, WILLIAM F.
1  JACK RYAN  CLANCY, TOM
4  JACKY ROWAN  DE LINT, CHARLES
2  JADE DARCY  GOLDIN, STEPHEN/MASON, MARY
4  JADE DEMONS  VARDEMAN, ROBERT E.
3  JADE PHOENIX  THURSTON, ROBERT
3  JAK JINNAKA  BARNES, JOHN
2  JAK REBO  DIETZ, WILLIAM C.
5  JALAV  GREEN, SHARON
23  JAMES BOND  authorship varies
2  JAMES UNDERHILL  JACCOMA, RICHARD
5  JAMETHIEL DREAM-WEAVER  HODGELL, P. C.
2  JAMIE WATERMAN  BOVA, BEN
5  JAN DARZEK  BIGGLE, LLOYD JR.
8  JANDAR  CARTER, LIN
6  JANI KILIAN  SMITH, KRISTINE
5  JANISSARIES  authorship varies
4  JANNA BRILL  KILLOUGH, LEE
2  JANUS  NORTON, ANDRE
2  JARAN  ELLIOTT, KATE
1  JARVED 9  O’SHEA, PATTI
2  JASON COSMO  MCGIRT, DAN
2  JASON WANDER  BUETTNER, ROBERT
6  JEFFERSON’S WAR  RANDLE, KEVIN
1  JENNA  YOLEN, JANE
3  JENNY CASEY  BEAR, ELIZABETH
2  JENNY SIXA  FULLILOVE, ERIC JAMES
2  JERN MURDOCK  NORTON, ANDRE
3  JEWELFIRE  WARRINGTON, FREDA
2  JEWELS OF ELVISH  BERBERICK, NANCY VARIAN
2  JIANA  AB HUGH, DAFYDD
2  JIG THE GOBLIN  HINES, JIM C.
1  JOAN OF ARC  CHAMBERLIN, ANN
3  JOE GREY  MURPHY, SHIRLEY ROUSSEAU
2  JOE’S WORLD  authorship varies
3  JOHAN ESBACH  MODESITT, L. E. JR.
1  JOHN EAGLE EXPEDITER  EDWARDS, PAUL
7  JOHN THE BALLADEER  WELLMAN, MANLY WADE
2  JOHN THUNSTONE  WELLMAN, MANLY WADE

ENTRIES  SERIES  AUTHOR
3  JOHNNY MAXWELL  PRATCHETT, TERRY
4  JONATHAN BARRETT  ELROD, P. N.
3  JORKENS  DUNSANY, LORD
4  JOURNEYS OF MCGILL FEIGHAN  O’DONNELL, KEVIN JR.
3  JOURNEYS OF THE CATECHIST  FOSTER, ALAN DEAN
7  JUDGE DREDD  authorship varies
5  JUDICIARY  MATTHEWS, SUSAN R.
2  JUMPER  GOULD, STEVEN C.
1  JUMP 225  EDELMAN, DAVID LOUIS
1  JUNO  HAMMOND, WARREN
2  JURASSIC PARK  CRICHTON, MICHAEL
3  JUSTICE  HAMILTON, VIRGINIA
1  JUSTICE LEAGUE OF AMERICA GREEN LANTERN  O’NEIL, DENNIS
3  JUSTIN BELL  FORSTCHEN, WILLIAM R.
3  K-9 CORPS  VON GUNDEN, KENNETH
2  K-PAX  BREWER, GENE
3  KAI LUNG  BRAMAH, ERNEST
4  KALEVALA  PETAJA, EMIL
1  KAMUS OF KADIZHAR  SHIRLEY, JOHN
3  KANTMORIE  DIBELL, ANSEN
3  KAPHTU  PURTILL, RICHARD
2  KARAVANS  ROBERSON, JENNIFER
2  KARL GLOGAUER  MOORCOCK, MICHAEL
3  KATIE CHANDLER  SWENDSON, SHANNA
5  KEDRIGERN  MORRESSY, JOHN
1  KEDRIGERN CHRONICLES  MORRESSY, JOHN
8  KEEPER MARTIN  STANEK, ROBERT
3  KEEPER’S CHRONICLES  HUFF, TANYA
3  KEEPERS OF THE HIDDEN WAYS  ROSENBERG, JOEL
4  KELTIAD  KENNEALY-MORRISON, PATRICIA
3  KELTIAD ARTHUR  KENNEALY-MORRISON, PATRICIA
1  KELTIAD MISC  KENNEALY-MORRISON, PATRICIA
5  KELVIN HACKLEBERRY  ANTHONY, PIERS/MARGROFF, ROBERT E.
3  KENDAR GANGLE  FRIESNER, ESTHER M.
5  KENT MONTANA  FENN, LIONEL
1  KEROWYN’S TALE  LACKEY, MERCEDES
3  KERRION EMPIRE  MORRIS, JANET
3  KEVIN BRADLEY AND EVAN DAVIS  BACON-SMITH, CAMILLE
2  KEVLA OF ARUKAN  GOLDEN, CHRISTIE
3  KEYS OF POWER  BROWN, SIMON
3  KEYS TO PARADISE  authorship varies
8  KEYS TO THE DIMENSIONS  BULMER, KENNETH
4  KEYS TO THE KINGDOM  NIX, GARTH
1  KHYREN  BOUCHER-KAPLAN, ALINE
4  KI AND VANDIEN  LINDHOLM, MEGAN
2  KILLSTAR  KELLY, DAVID J.
6  KING OF YS  ANDERSON, POUL/ANDERSON, KAREN
1  KING RAVEN  LAWHEAD, STEPHEN R.
2  KING WITHOUT A THRONE  COOK, GLEN
6  KING’S BLADES  DUNCAN, DAVE
1  KING’S DAGGERS  DUNCAN, DAVE
3  KING’S QUEST  authorship varies
4  KING’S TRAMP  DE HAVEN, TOM
3  KINGDOMS  WELLS, ANGUS
3  KINGDOMS OF THORN AND BONE  KEYES, J. GREGORY
1  KINGKILLER  ROTHFUSS, PATRICK
4  KIOGA  CHESTER, WILLIAM L.
3  KITTY NORVILLE  VAUGHN, CARRIE
2  KNIGHTS OF THE BLOOD  MACMILLAN, SCOTT
2  KNIGHTS OF THE DRAGONS REVELATIONS  FORBECK, MATT
12  KNIGHTS OF THE SILVER DRAGON  authorship varies
12  KNOWN SPACE  authorship varies
5  KNOWN SPACE RINGWORLD  NIVEN, LARRY
2  KONSTANTIN  CADIGAN, PAT
1  KORRE  LISLE, HOLLY
5  KOTHAR  FOX, GARDNER F.
2  KREN  LICHTENBERG, JACQUELINE
4  KRIS LONGKNIFE  SHEPHERD, MIKE
9  KRISHNA  authorship varies
3  KRISPOS  TURTLEDOVE, HARRY
3  KULL  authorship varies
3  KURTIS:P385XL47:RUTH  STAMEY, SARA
5  KUSHIEL  CAREY, JACQUELINE
3  KYDAN  BROWN, SIMON

ENTRIES  SERIES  AUTHOR
4  KYRIK  FOX, GARDNER F.
2  LADY BRANWEN  MEANEY, DEE MORRISON
3  LADY IN GIL  BRADLEY, REBECCA
4  LAFAYETTE O’LEARY  LAUMER, KEITH
2  LAMB AMONG THE STARS  WALLEY, CHRIS
2  LAND BETWEEN THE RIVERS  ACRES, MARK
4  LAND BEYOND THE GATE  ESHBACH, LLOYD ARTHUR
4  LAND OF THE GIANTS  authorship varies
1  LAND OF THE LOST  GRAHAM, KENNON
4  LAND THAT TIME FORGOT  BURROUGHS, EDGAR RICE
5  LANDOVER  BROOKS, TERRY
4  LANDUC  WILLEY, ELIZABETH
2  LANEN KAELAR  KERNER, ELIZABETH
2  LANNAT  MCQUINN, DONALD E.
1  LARA CROFT TOMB RAIDER  GARDNER, JAMES ALAN
2  LAST CLANSMAN  LAU, MILLER
4  LAST HERALD MAGE  LACKEY, MERCEDES
4  LAST LEGION  BUNCH, CHRIS
5  LAST LEGIONARY  HILL, DOUGLAS
3  LAST ON EARTH  KAYE, MARILYN
3  LAST RANGER  SARGENT, CRAIG
7  LAST RUNE  ANTHONY, MARK
1  LAST THOMAS COVENANT  DONALDSON, STEPHEN R.
1  LAST VAMPIRE  PIKE, CHRISTOPHER
3  LAST VIKING  ANDERSON, POUL
5  LAWS OF THE BLOOD  SIZEMORE, SUSAN
1  LCSI  HARPER, STEVEN
6  LEAGUE OF PEOPLES  GARDNER, JAMES ALAN
2  LEAR’S DAUGHTERS  KELLOGG, MARJORIE BRADLEY/ROSSOW, WILLIAM B.
4  LEFT BEHIND  LAHAYE, TIM/JENKINS, JERRY B.
3  LEGACY  DOUGLAS, IAN
3  LEGENDS OF LONE WOLF  DEVER, JOE/GRANT, JOHN
3  LEGENDS OF THE RAVEN  BARCLAY, JAMES
1  LEGENDS OF THE RIFTWAR  FEIST, RAYMOND E./FORSTCHEN, WILLIAM R.
6  LEGION OF THE DAMNED  DIETZ, WILLIAM C.
2  LEIBOWITZ  MILLER, WALTER M. JR.
3  LEMMUS  SAVARIN, JULIAN JAY
3  LENS OF THE WORLD  MACAVOY, R. A.
9  LENSMEN  authorship varies
3  LER  FOSTER, M. A.
11  LESTAT  authorship varies
5  LIADEN ADVENTURES  LEE, SHARON/MILLER, STEVE
17  LIADEN UNIVERSE  LEE, SHARON/MILLER, STEVE
2  LIADEN UNIVERSE GREAT MIGRATION  LEE, SHARON/MILLER, STEVE
6  LIAVEK  authorship varies
2  LIFE PROBE  MCCOLLUM, MICHAEL
2  LIGHT  HARRISON, M. JOHN
3  LIGHTBRINGER  JOHNSON, OLIVER
2  LIGHTSTONE  ZINDELL, DAVID
3  LIJE BALEY  ASIMOV, ISAAC
2  LILY YU  WILKS, EILEEN
4  LINCOLN BLACKTHORNE  MARSH, GEOFFREY
1  LINDORMYN  BILSBOROUGH, DAVID
2  LION’S BLOOD  BARNES, STEVEN
1  LITTLE GODDESS  LANE, AMY
3  LIVESHIP TRADERS  HOBB, ROBIN
3  LIVING FLAME  PATTON, FIONA
4  LIVING PLANET  COON, SUSAN
2  LIVING WORLD  POLLACK, RACHEL
2  LIZARD WAR  DALMAS, JOHN
3  LLADRANA  OWENS, ROBIN D.
2  LLARN  FOX, GARDNER F.
4  LOGAN  authorship varies
2  LOGICAL MAGICIAN  WEINBERG, ROBERT
1  LOIS & CLARK  CHERRYH, C. J.
3  LON TOBYN  COE, DAVID B.
2  LONDON VAMPIRES  HAMBLY, BARBARA
21  LONE WOLF  authorship varies
1  LONG PRICE  ABRAHAM, DANIEL
6  LONG SUN  WOLFE, GENE
3  LONGDIRK  HOOD, KEN
1  LOOKING GLASS  RINGO, JOHN
7  LORD DARCY  authorship varies
6  LORD OF THE ISLES  DRAKE, DAVID

ENTRIES  SERIES  AUTHOR
16  LORD OF THE RINGS  authorship varies
4  LORD TEDRIC  authorship varies
2  LORDS OF AVALON  MACGREGOR, KINLEY
5  LORDS OF DUS  WATT-EVANS, LAWRENCE
1  LORDS OF THE SKY  WELLS, ANGUS
1  LOREGIVER  CHRISTIAN, DEBORAH
2  LOST CONTINENT  ASARO, CATHERINE
2  LOST FLEET  CAMPBELL, JACK
6  LOST IN SPACE  authorship varies
1  LOST IN SPACE NEW JOURNEYS  KRULIK, NANCY
3  LOST LANDS  ESTES, ROSE
3  LOST MILLENNIUM  MOSCOE, MIKE
9  LOST REGIMENT  FORSTCHEN, WILLIAM R.
11  LOST SWORDS  SABERHAGEN, FRED
1  LOST WORLD  SOUZA, MARCIO
2  LOST YEARS  LUMLEY, BRIAN
4  LOST YEARS OF MERLIN  BARRON, T. A.
6  LUCIFER COVE  COFFMAN, VIRGINIA
2  LUCIFER JONES  RESNICK, MIKE
3  LUCKY 13TH  SHELLEY, RICK
12  LUCKY STARR  ASIMOV, ISAAC
3  LUKE  CHRISTOPHER, JOHN
6  LYONESSE  VANCE, JACK
6  LYRA  WREDE, PATRICIA C.
3  LYREMOUTH  FLETCHER, JANE
4  MABINOGI  WALTON, EVANGELINE
2  MAC MCLANAHAN  BROWN, DALE
2  MACEDON  GEMMELL, DAVID
2  MACNAIR OF MACNAIR  PEIRCE, HAYFORD
4  MADAGASCAR MANIFESTO  BERLINER, JANET/GUTHRIDGE, GEORGE
2  MAGE AND TAKUMO  DEDMAN, STEPHEN
1  MAGE ASCENSION  STEELE, JOHN H.
3  MAGE STORMS  LACKEY, MERCEDES
3  MAGE WARS  LACKEY, MERCEDES/DIXON, LARRY
3  MAGE WINDS  LACKEY, MERCEDES
3  MAGE: THE ASCENSION  WEINBERG, ROBERT
2  MAGELORD  MARTIN, THOMAS K.
3  MAGES OF GARILLON  HARRIS, DEBORAH TURNER
1  MAGEVERSE MASTERS  KNIGHT, ANGELA
7  MAGEWORLDS  DOYLE, DEBRA/MACDONALD, JAMES D.
1  MAGI  CURTIS, GENE
5  MAGIC  authorship varies
1  MAGIC LANDS  STANEK, ROBERT
1  MAGIC OF THE PLAINS  COSTIKYAN, GREG
2  MAGIC OR NOT?  EAGER, EDWARD
3  MAGIC SHOP  COVILLE, BRUCE
2  MAGIC TIME  authorship varies
2  MAGIC’S RETURN  SERVICE, PAMELA F.
19  MAGIC: THE GATHERING  authorship varies
4  MAGIC: THE GATHERING ARTIFACTS  authorship varies
3  MAGIC: THE GATHERING ICE AGE  GRUBB, JEFF
3  MAGIC: THE GATHERING INVASION  KING, J. ROBERT
3  MAGIC: THE GATHERING KAMIGAWA  MCGOUGH, SCOTT
3  MAGIC: THE GATHERING LEGENDS  EMERY, CLAYTON
3  MAGIC: THE GATHERING LEGENDS TWO  MCGOUGH, SCOTT
2  MAGIC: THE GATHERING MASQUERADE  authorship varies
3  MAGIC: THE GATHERING MIRRODIN  authorship varies
3  MAGIC: THE GATHERING ODYSSEY  authorship varies
3  MAGIC: THE GATHERING ONSLAUGHT  KING, J. ROBERT
3  MAGIC: THE GATHERING RAVNICA  HERNDON, CORY J.
3  MAGIC: THE GATHERING TIME SPIRAL  authorship varies
1  MAGICIAN’S HOUSE  CORLETT, WILLIAM
3  MAGICKERS  DRAKE, EMILY
3  MAGIRA  WALKER, HUGH
1  MAGISTER  FRIEDMAN, C. S.
3  MAGRAVANDIAS  CONSTANTINE, STORM
3  MAIRELON AND KIM  WREDE, PATRICIA C.
4  MAJIPOOR  SILVERBERG, ROBERT
3  MAJIPOOR PRESTIMION  SILVERBERG, ROBERT
7  MALAZAN  ERIKSON, STEVEN
2  MALKAR  TUBB, E. C.
8  MALLOREON  authorship varies
2  MALLWORLD  authorship varies
3  MAMMOTH SF  WAUGH, CHARLES G./GREENBERG, MARTIN H.

ENTRIES  SERIES  AUTHOR
3  MAMMOTHS  BAXTER, STEPHEN
4  MAN FROM ATLANTIS  WOODLEY, RICHARD
2  MAN FROM ORGY  MARK, TED
2  MAN FROM PLANET X  ADAMS, HUNTER
1  MAN FROM TOMCAT  KNIGHT, MALLORY T.
28  MAN FROM UNCLE  authorship varies
4  MAN OF HIS WORD  DUNCAN, DAVE
2  MAN PLUS  authorship varies
4  MAN WHO LOVED MARS  CARTER, LIN
17  MAN-KZIN WARS  authorship varies
4  MANCER  CALLANDER, DON
2  MANSE  CANTRELL, LISA W.
3  MARA  FEIST, RAYMOND E./WURTS, JANNY
2  MARATHON  SMITH, DAVID ALEXANDER
4  MARCH UPCOUNTRY  WEBER, DAVID/RINGO, JOHN
1  MARCO  SLEATOR, WILLIAM
3  MARIANNE  TEPPER, SHERI S.
4  MARID AUDRAN  EFFINGER, GEORGE ALEC
4  MARK HOOD  DARK, JAMES
2  MARKED MAN  INGRID, CHARLES
2  MARLEY FAMILY WEREWOLF  SCOTCH, CHERI
3  MARQ’SSAN  DUCHAMP, L. TIMMEL
2  MARROW  REED, ROBERT
5  MARS  ROBINSON, KIM STANLEY
3  MARS ATTACKS  authorship varies
9  MARVEL COMICS  authorship varies
1  MARVEL’S FINEST  SHOOTER, JIM
2  MARVIN BREWSTER  HAWKE, SIMON
1  MASS EFFECT  KARPYSHYN, DREW
3  MASTER LI  HUGHART, BARRY
3  MASTERS OF SPACE  VARDEMAN, ROBERT E.
8  MATADOR  authorship varies
3  MATERNAS  FORREST, KATHERINE V.
3  MATHEW LEWIN  GRAHAM, MITCHELL
3  MATHEW SWAIN  MCQUAY, MIKE
2  MATILDA BLACK  HOFFMAN, NINA KIRIKI
3  MATT CRUSE  authorship varies
2  MATTHEW DILKE  GUTTERIDGE, LINDSAY
5  MATTHEWS  BURGAUER, STEVEN
2  MAURA WOODBURY  HALE, DEBORAH
2  MAUREEN PIERCE  HETLEY, JAMES A.
3  MAVIN MANYSHAPED  TEPPER, SHERI S.
3  MAX MERRIWELL  MURPHY, PAT
2  MAXIMUM ICE  KENYON, KAY
5  MAYFAIR WITCHES  authorship varies
1  MAYFLOWER  CARD, ORSON SCOTT/KIDD, KATHRYN H.
2  MECHANICAL SKY  MOFFITT, DONALD
5  MECHWARRIOR  authorship varies
24  MECHWARRIOR DARK AGE  authorship varies
3  MECHWARRIOR DARK AGE PROVING GROUNDS  DELRIO, MARTIN
6  MED  LEINSTER, MURRAY
4  MEG MURRY  L’ENGLE, MADELEINE
1  MEL DAWNFIELD  ASARO, CATHERINE
3  MELVINGE  authorship varies
5  MEMORY, SORROW AND THORN  WILLIAMS, TAD
4  MEN IN BLACK  authorship varies
2  MENAGERIE  GOLDEN, CHRISTOPHER/SNIEGOSKI, THOMAS E.
3  MENNENKALTENEI  TURNER, DELIA MARSHALL
1  MENNYMS  WAUGH, SYLVIA
3  MERCHANT PRINCE  authorship varies
3  MERCHANT PRINCES  STROSS, CHARLES
2  MERCY THOMPSON  BRIGGS, PATRICIA
5  MEREDITH GENTRY  HAMILTON, LAURELL K.
3  MERI  BOHNHOFF, MAYA KAATHRYN
3  MERLIN CODEX  HOLDSTOCK, ROBERT P.
4  MERLIN’S DESCENDANTS  RADFORD, IRENE
7  MEROVINGEN NIGHTS  CHERRYH, C. J.
2  MESSENGER  KELLY, JAMES PATRICK
2  MESSIAH STONE  CAIDIN, MARTIN
2  METAPLANETARY  DANIEL, TONY
9  MICHAEL KANE  authorship varies
2  MICHAEL PERRIN  BEAR, GREG
1  MICHAEL PERRIN (REVISED)  BEAR, GREG
3  MICRONAUTS  WILLIAMS, GORDON

ENTRIES  SERIES  AUTHOR
2  MIDNIGHTERS  WESTERFELD, SCOTT
3  MIGHT AND MAGIC  authorship varies
8  MIKE MARS  WOLLHEIM, DONALD A.
15  MILES VORKOSIGAN  BUJOLD, LOIS MCMASTER
1  MILLENIUM  HAND, ELIZABETH
2  MILLENNIUM QUARTET  GRANT, CHARLES L.
3  MIND BROTHERS  HEATH, PETER
4  MIND MASTERS  authorship varies
3  MINDKILLER  ROBINSON, SPIDER
2  MIREI  BRENNAN, MARIE
1  MISS FROM SIS  TRALINS, ROBERT
5  MISS PICKERELL  authorship varies
1  MISSION CRITICAL  CHAFE, PAUL C.
11  MISSION EARTH  HUBBARD, L. RON
2  MISSION IMPOSSIBLE  WALKER, MAX
1  MISTBORN  SANDERSON, BRANDON
1  MISTMANTLE  MCALLISTER, M. I.
3  MISTS OF AVALON  authorship varies
15  MITHGAR  MCKIERNAN, DENNIS L.
4  MODE  ANTHONY, PIERS
4  MODESTY BLAISE  O’DONNELL, PETER
5  MONARCHIES OF GOD  KEARNEY, PAUL
1  MONERE  SUNNY
4  MONGO FREDERICKSON  CHESBRO, GEORGE C.
2  MONTE COOK’S ARCANA EVOLVED  COOK, SUE WEINLEIN
5  MOON SINGER  NORTON, ANDRE
2  MOONBASE  BOVA, BEN
2  MOONRUNNER  TOOMBS, JANE
2  MOONTIDE AND MAGIC RISE  RUSSELL, SEAN
3  MOONWORLDS  MCMULLEN, SEAN
2  MORDANT’S NEED  DONALDSON, STEPHEN R.
2  MORDRED’S CURSE  MCDOWELL, IAN
4  MOREAU  SWANN, S. ANDREW
6  MORGAINE  CHERRYH, C. J.
2  MORGAN OF ALB’NY  MCQUAY, MIKE
2  MORGANVILLE VAMPIRES  CAINE, RACHEL
2  MORIGU  PERRY, MARK C.
1  MORK AND MINDY  WAGNER, ROBIN S.
3  MORPHODITE  FOSTER, M. A.
3  MORTAL KOMBAT  authorship varies
2  MOTE IN GOD’S EYE  NIVEN, LARRY/POURNELLE, JERRY
2  MOUNTAIN MADE OF LIGHT  MYERS, EDWARD
1  MULTIVERSE  WEBER, DAVID/EVANS, LINDA
3  MUTANT CHRONICLES APOSTLE OF INSANITY  authorship varies
4  MUTANT SEASON  authorship varies
4  MUTANTS AMOK  GRANT, MARK
4  MY TEACHERS  COVILLE, BRUCE
3  MYRON TANY  VANCE, JACK
3  MYST  authorship varies
2  MYSTARA  authorship varies
3  MYSTARA DRAGONLORD  GUNNARSSON, THORARINN
1  MYSTIC REBEL  SYVERTSEN, RYDER
18  MYTH ADVENTURES  authorship varies
3  MYTH INC  ASPRIN, ROBERT LYNN
3  MYTHOLOGY  NYE, JODY LYNN
2  NAGA TEOT  GLADNEY, HEATHER
2  NANNY AND THE PROFESSOR  JOHNSTON, WILLIAM
2  NANOTECH  GOONAN, KATHLEEN ANN
2  NANOTECH VERITY  GOONAN, KATHLEEN ANN
3  NANTUCKET  STIRLING, S. M.
3  NAPHAR  BAKER, SHARON
7  NARNIA  LEWIS, C. S.
3  NATIVE TONGUE  ELGIN, SUZETTE HADEN
4  NAUSICAA PERFECT  MIYAZAKI, HAYAO
3  NEANDERTHAL PARALLAX  SAWYER, ROBERT J.
1  NEAR SPACE  STEELE, ALLEN
1  NECROMANCER  MARTIN, GAIL Z.
9  NECROMUNDA  authorship varies
11  NECROSCOPE  LUMLEY, BRIAN
1  NEDAO  EMERSON, RU
4  NETWORK-CONSORTIUM  FRANKLIN, CHERYL J.
3  NEUROMANCER  GIBSON, WILLIAM
3  NEUSTRIAN CYCLE  BARRINGER, LESLIE
1  NEW CROBUZON  MIEVILLE, CHINA

ENTRIES  SERIES  AUTHOR
2  NEW ICE AGE  MORAN, RICHARD
6  NEW SUN  WOLFE, GENE
2  NEW TALES  RICE, ANNE
11  NEW TOM SWIFT  APPLETON, VICTOR
9  NEWFORD  DE LINT, CHARLES
3  NEWS FROM THE EDGE  SUMNER, MARK
3  NEXT WAVE  authorship varies
8  NICHOLAS SEAFORT  FEINTUCH, DAVID
2  NICOLAS RATHE  SCOTT, MELISSA/BARNETT, LISA A.
3  NICOLE SHEA  CLAREMONT, CHRIS
2  NIFFT  SHEA, MICHAEL
1  NIGHT OF THE SOLSTICE  SMITH, L. J.
2  NIGHT WHISTLERS  TREVOR, DAN
8  NIGHT’S DAWN  HAMILTON, PETER F.
6  NIGHT-THREADS  EMERSON, RU
1  NIGHTFALL  REICHERT, MICKEY ZUCKER
3  NIGHTRUNNER  FLEWELLING, LYNN
8  NIGHTSIDE  GREEN, SIMON R.
2  NIGHTWALKERS  FRANK, JACQUELYN
3  NIGHTWATCH  LUKYANENKO, SERGEI
1  NIKOLAI DANTE  BISHOP, DAVID
2  NIMH  authorship varies
3  NIMNESTL  CRAWFORD, DAN
1  NINE KINGDOMS  KURLAND, LYNN
5  NOBLE DEAD SERIES 1  HENDEE, BARB/HENDEE, J. C.
3  NOEL KEDRAN  DALTON, SEAN
7  NORBY  ASIMOV, JANET/ASIMOV, ISAAC
3  NOREN  ENGDAHL, SYLVIA LOUISE
8  NORTH AMERICAN CONFEDERACY  SMITH, L. NEIL
4  NORTHWORLD  DRAKE, DAVID
2  NOT QUITE HUMAN  MCEVOY, SETH
3  NUALA  KIMBRIEL, KATHERINE ELISKA
3  NULAPEIRON  MEANEY, JOHN
4  OAK, YEW, ASH, AND ROWAN  NORTON, ANDRE/MILLER, SASHA
5  OATH OF EMPIRE  HARLAN, THOMAS
4  OBERNEWTYN  CARMODY, ISOBELLE
3  OBSIDIAN  LACKEY, MERCEDES/MALLORY, JAMES
3  OBSIDIAN CHRONICLES  WATT-EVANS, LAWRENCE
3  OF MAN AND MANTA  ANTHONY, PIERS
2  OHNDARIEN  CARWYN, GILES/FAHNESTOCK, TODD
4  OLD MAN’S WAR  SCALZI, JOHN
1  OMAR  DUNCAN, DAVE
4  OMARAN  COLE, ADRIAN
6  OMEGA SUB  CAMERON, J. D.
5  OMEN  authorship varies
1  ONE LAND  STOUT, AMY
2  ONE ROSE  DAYTON, GAIL
1  OPENING OF THE WORLD  TURTLEDOVE, HARRY
2  OPERATION CHAOS  ANDERSON, POUL
1  OPERATION PLEIADES  SCHARTZ, VIJAYA
3  OPERATOR 5  STEELE, CURTIS
1  ORACLE  FISHER, CATHERINE
3  ORCS FIRST BLOOD  NICHOLLS, STAN
6  ORGANIC FUTURE  EASTON, THOMAS A.
2  ORIANI  GADALLAH, LESLIE
2  ORIANUS CREATION  BURTON, D. SHANE
5  ORION  BOVA, BEN
1  ORLANDO FURIOSO  ARIOSTO, LUDOVICO
5  OROKON  ARDEN, TOM
5  ORON  SMITH, DAVID C.
1  ORPHAN’S TALES  VALENTE, CATHERYNNE M
3  ORPHANS OF CHAOS  WRIGHT, JOHN C.
2  ORTHE  GENTLE, MARY
1  OSSERIAN  FORBES, DAVID
3  OSWALD BASTABLE  MOORCOCK, MICHAEL
3  OTHER SINBAD  GARDNER, CRAIG SHAW
4  OTHERLAND  WILLIAMS, TAD
3  OTHERS (BONANNO)  BONANNO, MARGARET WANDER
3  OTHERS (WARREN)  WARREN, CHRISTINE
10  OTHERWORLD  authorship varies
5  OTORI  HEARN, LIAN
1  OUT OF THIS WORLD  KAYE, MARILYN
3  OUT OF TIME  authorship varies
5  OUTER LIMITS  PEEL, JOHN
4  OUTLANDER (GABALDON)  GABALDON, DIANA

ENTRIES  SERIES  AUTHOR
1  OUTLANDER (SALTERBERG)  SALTERBERG, B. J.
6  OUTLANDERS  AXLER, JAMES
2  OUTLANDERS HEART OF THE WORLD  AXLER, JAMES
1  OUTREACH INVESTIGATIONS  FOOTE, KEITH D.
8  OUTREMER  BRENCHLEY, CHAZ
4  OUTRIDER  HARDING, RICHARD
17  OZ  authorship varies
2  OZ CHRONICLES  RIDLEY, R. W.
4  OZARK  ELGIN, SUZETTE HADEN
1  PAINTER  GILLULY, SHEILA
4  PAKSENARRION  MOON, ELIZABETH
2  PALADINS  ROSENBERG, JOEL
2  PANDORA  LANCE, KATHRYN
2  PANDRAGORE  JONES, DENNIS
2  PANGAEA  MASON, LISA
6  PARADYS  LEE, TANITH
5  PARATIME POLICE  authorship varies
3  PARATWA  HINZ, CHRISTOPHER
2  PARRISH PLESSIS  DE PIERRES, MARIANNE
3  PARSINA  GOLDIN, STEPHEN
4  PARSIVAL  MONACO, RICHARD
6  PATTERNS OF CHAOS  INGRID, CHARLES
2  PAUL CARPENTER  HOLT, TOM
2  PAUL HOLTON  HUNT, CHARLOTTE
4  PAUL SINCLAIR  HEMRY, JOHN G.
2  PAX  NORTON, ANDRE
3  PEACE COMPANY  GREEN, ROLAND J.
4  PEARL  LUSTBADER, ERIC VAN
2  PEARL OF IMMORTALITY  CLOUGH, BRENDA W.
4  PEGASUS  MCCAFFREY, ANNE
7  PELBAR  WILLIAMS, PAUL O.
9  PELLUCIDAR  BURROUGHS, EDGAR RICE
3  PELMEN THE POWERSHAPER  HUGHES, ROBERT DON
3  PEMBROKE HALL  FAUST, JOE CLIFFORD
5  PENDRAGON  LAWHEAD, STEPHEN R.
2  PEOPLE OF THE WOLF  GEAR, W. MICHAEL/GEAR, KATHLEEN O’NEAL
1  PEPPER MARTIN  DANIELS, CASEY
1  PERCHERON  MCINTOSH, FIONA
1  PERCY JACKSON & THE OLYMPIANS  RIORDAN, RICK
4  PERELANDRA  LEWIS, C. S.
19  PERN  authorship varies
4  PERN DRAGONRIDERS  MCCAFFREY, ANNE
4  PERN HARPER HALL  MCCAFFREY, ANNE
174  PERRY RHODAN  authorship varies
1  PERRY RHODAN ANDROMEDA  ANTON, UWE
1  PERRY RHODAN LEMURA  BORSCH, FRANK
15  PERRY RHODAN SILBERBANDE  PABEL MOEWIG KG
3  PETAYBEE  MCCAFFREY, ANNE/SCARBOROUGH, ELIZABETH ANN
2  PETER  BARRY, DAVE/PEARSON, RIDLEY
2  PHAESTOR  GERROLD, DAVID
2  PHAID  FARREN, MICK
6  PHANTOM  authorship varies
35  PHENWICK  KIMBROUGH, KATHERYN
3  PHIL D’AMATO  LEVINSON, PAUL
2  PHOENIX  ALEXANDER, DAVID
3  PHOENIX LEGACY  WREN, M. K.
7  PHOTON  authorship varies
6  PHULE  authorship varies
3  PILGRIM  SABERHAGEN, FRED
2  PINCH  authorship varies
4  PIT DRAGONS  YOLEN, JANE
1  PLANE SCAPE  VALLESE, RAY/VALLESE, VALERIE
2  PLANESCAPE BLOOD WARS  KING, J. ROBERT
6  PLANET BUILDERS  TALLIS, ROBYN
1  PLANET DWELLER  PALMER, JANE
5  PLANET OF THE APES  authorship varies
4  PLANET PIRATES  authorship varies
2  PLANETFALL  COVER, ARTHUR BYRON
2  PLAYERS IN THE CONTEST OF WORLDS  BRODERICK, DAMIEN
3  PLENTY  GREENLAND, COLIN
6  PLIOCENE EXILE  MAY, JULIAN
3  POL DETSON  ZELAZNY, ROGER
2  POLAR CITY  authorship varies
2  POLITY  ASHER, NEAL

ENTRIES  SERIES  AUTHOR
4  POLITY GRIDLINKED  ASHER, NEAL
2  POLITY SKINNER  ASHER, NEAL
1  POLTERGEIST THE LEGACY  HAUTALA, RICK
8  POSLEEN  authorship varies
2  POWER AND THE PATTERN  BUSH, ANNE KELLEHER
1  POWER OF THE BOOD  KILPATRICK, NANCY
6  PREDATOR  authorship varies
3  PRETERNATURAL  BONANNO, MARGARET WANDER
3  PRIMAL LAND  LUMLEY, BRIAN
1  PRIMAL RAGE  VORNHOLT, JOHN
1  PRIMORTALS  PERRY, STEVE
3  PRINCE OF NOTHING  BAKKER, R. SCOTT
2  PRINCE OF THE SIDHE  SHERMAN, JOSEPHA
4  PRISCILLA HUTCHINS  MCDEVITT, JACK
4  PRISONER  authorship varies
6  PRIVATE SCHOOL  CHARLES, STEVEN
2  PRIVATEERS  BOVA, BEN
2  PROBE  DOUGLAS, CAROLE NELSON
3  PROCURATOR  MITCHELL, KIRK
2  PROF. CHALLENGER  DOYLE, ARTHUR CONAN
5  PROFESSOR JAMESON  JONES, NEIL R.
1  PROFESSOR SHANDY  MACLEOD, CHARLOTTE
2  PROMETHEAN AGE  BEAR, ELIZABETH
4  PROTECTOR OF THE SMALL  PIERCE, TAMORA
5  PROTEUS  SHEFFIELD, CHARLES
2  PROWLERS  GOLDEN, CHRISTOPHER
7  PRYDAIN  ALEXANDER, LLOYD
3  PSI PATROL  authorship varies
6  PSI-MAN  DAVID, PETER
1  PSYCHIC WARRIOR  DOHERTY, ROBERT
3  PSYCHO  BLOCH, ROBERT
2  PSYCHODROME  HAWKE, SIMON
3  PSYCHOMECH  LUMLEY, BRIAN
1  PTOLIA  BAKER, FRED
2  PUNKTOWN  THOMAS, JEFFREY

1 PURRFECT SERIES WAUGH, CAROL-LYNN ROSSEL/GREENBERG, MARTIN/

ASIMOV, ISAAC

1  PYAT  MOORCOCK, MICHAEL
2  Q  HOYLE, TREVOR
2  QUAG KEEP  authorship varies
1  QUANTUM GRAVITY  ROBSON, JUSTINA
17  QUANTUM LEAP  authorship varies
4  QUARTER SINGERS  HUFF, TANYA
4  QUEEN OF ANGELS  BEAR, GREG
3  QUEEN’S QUARTER  SNYDER, MIDORI
5  QUEENDOM OF SOL  MCCARTHY, WIL
3  QUEST FOR THE WHITE DUCK  FENN, LIONEL
5  QUEST FOR TOMORROW  SHATNER, WILLIAM
2  QUEST TRIAD  NILES, DOUGLAS
3  QUESTIONER  SCHMIDT, DENNIS
3  QUICKENING  MCINTOSH, FIONA
3  QUINTARA  CHALKER, JACK L.
3  RABELAIS  CARR, JAYGE
5  RACHEL MORGAN  HARRISON, KIM
5  RACK  JAMES, LAURENCE
2  RADIANT DAWN  GOODFELLOW, CODY
4  RADIX  ATTANASIO, A. A.
1  RAFE AND ASKANA  LEE, LISA/MORRIS, TEE
7  RAGNARSON  COOK, GLEN
2  RAHXEPHON  IZUBUCHI, YUTAKA/BONES
6  RAMA  authorship varies
2  RAMAYANA  BANKER, ASHOK K.
3  RAMPART WORLDS  MAY, JULIAN
3  RANKS OF BRONZE  authorship varies
3  RATHA  BELL, CLARE
2  RATS & BATS  authorship varies
4  RAVEN  KIRK, RICHARD
19  RAVENLOFT  authorship varies
1  RAVENSCLIFF  HUNTINGTON, GEOFFREY
6  RAY BRADBURY’S DINOSAURS  authorship varies
6  RCN  DRAKE, DAVID
3  REACHES  DRAKE, DAVID
2  REALMS OF ARKANIA THALIONMEL  KRAMER, INA
2  REALMS OF CHAOS  STACKPOLE, MICHAEL A.

ENTRIES  SERIES  AUTHOR
2  REANDN  DURGIN, DORANNA
13  RECLUCE  MODESITT, L. E. JR.
3  RECURSION  BALLANTYNE, TONY
5  RED DWARF  authorship varies
2  RED PAVILIONS  HUNTER, KIM
6  RED SONJA  SMITH, DAVID C./TIERNEY, RICHARD L.
2  RED THUNDER  VARLEY, JOHN
16  REDWALL  JACQUES, BRIAN
5  REGIMENT  DALMAS, JOHN
4  REID MALENFANT  BAXTER, STEPHEN
3  REIGN OF SHADOWS  CHESTER, DEBORAH
3  REISIL  FRANCIS, DIANA PHARAOH
1  RELAMS OF ARKANIA  KIESOW, ULRICH
2  RELIC  PRESTON, DOUGLAS/CHILD, LINCOLN
5  RELUCTANT KING  DE CAMP, L. SPRAGUE
3  REMEMBER!  RANDLE, KEVIN/CORNETT, ROBERT
3  RENSHAI  REICHERT, MICKEY ZUCKER
3  RENSHAI CHRONICLES  REICHERT, MICKEY ZUCKER
11  REPAIRMAN JACK  WILSON, F. PAUL
4  REQUIEM FOR HOMO SAPIENS  ZINDELL, DAVID
7  RESIDENT EVIL  PERRY, S. D.
3  RESURRECTION MAN  STEWART, SEAN
22  RETIEF  authorship varies
5  RETRIEVAL ARTIST  RUSCH, KRISTINE KATHRYN
4  RETRIEVERS  GILMAN, LAURA ANNE
3  RETURN TO THE PLANET OF THE APES  ARROW, WILLIAM
6  REVELATION SPACE  REYNOLDS, ALASTAIR
1  REWIND TO YESTERDAY  PFEFFER, SUSAN BETH
4  RHADA  GILMAN, ROBERT CHAM
3  RHIANNON’S RIDE  FORSYTH, KATE
23  RICK BRANT  BLAINE, JOHN
4  RIFTERS  WATTS, PETER
9  RIFTWAR  authorship varies
3  RIFTWAR LEGACY  FEIST, RAYMOND E.
2  RIFTWORLD  MCCAY, BILL
5  RIGANTE  GEMMELL, DAVID
4  RILEY JENSON  ARTHUR, KERI
2  RILSIN  BRENNAN, NOEL-ANNE
4  RINGS OF THE MASTER  CHALKER, JACK L.
4  RISING STARS  authorship varies
2  RIVER INTO DARKNESS  RUSSELL, SEAN
3  RIVERRUN  SOMTOW, S. P.
1  RIVERS OF WAR  FLINT, ERIC
9  RIVERWORLD  FARMER, PHILIP JOSE
4  RO-LAN  SIROTA, MIKE
6  ROBOT CITY  authorship varies
1  ROBOT ODYSSEY  D’IGNAZIO, FRED
22  ROBOTECH  MCKINNEY, JACK
6  ROBOTS AND ALIENS  authorship varies
6  ROBOTS IN TIME  WU, WILLIAM F.
5  ROCHEWORLD  authorship varies
2  ROCK AND ROLL REICH  JONES, GWYNETH
2  ROD ALLBRIGHT  COVILLE, BRUCE
2  RODERICK THE ROBOT  SLADEK, JOHN
2  RODRIGO OF CALEDON  FEINTUCH, DAVID
2  ROGER RABBIT  WOLF, GARY K.
4  ROGER ZELAZNY’S DAWN OF AMBER  BETANCOURT, JOHN GREGORY
2  ROGUE TROOPER  authorship varies
9  ROGUE WIZARD  STASHEFF, CHRISTOPHER
2  ROLAND  VAN ASTEN, GAIL
3  ROSE OF THE PROPHET  WEIS, MARGARET/HICKMAN, TRACY
3  ROSINANTE  GILLILAND, ALEXIS A.
2  ROUMANIA  PARK, PAUL
4  ROVER  ODOM, MEL
4  ROWAN  MCCAFFREY, ANNE
6  RUIN MIST  STANEK, ROBERT
1  RUIN MIST DARK PATH  STANEK, ROBERT
3  RULERS OF HYLOR  WILDER, CHERRY
3  RUNE BLADE  MARSTON, ANN
1  RUNE OF UNMAKING  HOWARD, MADELINE
5  RUNELORDS  FARLAND, DAVID
3  RUNESPELL  WELCH, JANE
7  RUNESTAFF  MOORCOCK, MICHAEL
6  RUNESWORD  authorship varies

ENTRIES  SERIES  AUTHOR
2  RUSTUM  ANDERSON, POUL
3  RYHOPE WOOD  HOLDSTOCK, ROBERT P.
26  SABRINA THE TEENAGE WITCH  authorship varies
7  SABRINA THE TEENAGE WITCH SALEM  authorship varies
3  SACRED STONES  CALDECOTT, MOYRA
2  SAGAMORE  DALKEY, KARA
17  SAINT-GERMAIN  YARBRO, CHELSEA QUINN
3  SAINT-GERMAIN OLIVIA  YARBRO, CHELSEA QUINN
3  SALLY FORTH  LEIBER, JUSTIN
3  SAM GUNN  BOVA, BEN
5  SAM MCCADE  DIETZ, WILLIAM C.
2  SAM SPACE  NOLAN, WILLIAM F.
4  SAMANTHA SLADE  SMITH, SUSAN
5  SAMARIA  SHINN, SHARON
2  SAMUEL CLEMENS FOWLER  BROCK, DARRYL
5  SAMURAI CAT  ROGERS, MARK E.
6  SAND WARS  INGRID, CHARLES
13  SANDMAN  authorship varies
1  SANDWRITER  HUGHES, MONICA
1  SANGREAL  HEMINGWAY, AMANDA
2  SANTIAGO  RESNICK, MIKE
3  SARANTINE MOSAIC  KAY, GUY GAVRIEL
1  SATAN SLEUTH  AVALLONE, MICHAEL
3  SATELLITE NIGHT  HOPKINS, JACK
6  SAVAGE EMPIRE  authorship varies
3  SCAR  SARTI, RON
3  SCAVENGER  PARKER, K. J.
3  SCEPTER OF MERCY  CHERNENKO, DAN
3  SCEPTER’D ISLE  LACKEY, MERCEDES/GELLIS, ROBERTA
4  SCHRODINGER’S CAT  WILSON, ROBERT ANTON
3  SCIENCE OF DISCWORLD  PRATCHETT, TERRY/STEWART, IAN/COHEN, JACK
6  SCORPIO  MCDONOUGH, ALEX
3  SCOTT ST. ANDREW  WEAVER, BEN
2  SEA BEGGARS  KEARNEY, PAUL
3  SEA KING  SPRINGER, NANCY
1  SEA OF GRASS  GERROLD, DAVID
3  SEAQUEST DSV  authorship varies
3  SEARCHERS  WILLIAMSON, CHET
1  SEAS OF KILMOYN  GREEN, ROLAND J.
4  SECANTIS  TIEDEMANN, MARK W.
3  SECOND COMMONWEALTH  SHELLEY, RICK
3  SECOND ETHER  MOORCOCK, MICHAEL
3  SECOND FOUNDATION  authorship varies
3  SECOND SONS  FALLON, JENNIFER
1  SECRET BOOKS OF VENUS  LEE, TANITH
1  SECRET CIRCLE  SMITH, L. J.
3  SECRET COUNTRY  DEAN, PAMELA C.
3  SECRET TEXTS  LISLE, HOLLY
3  SECRETS OF POWER  CHARRETTE, ROBERT N.
13  SECTOR GENERAL  WHITE, JAMES
3  SEEDS OF WAR  RANDLE, KEVIN/CORNETT, ROBERT
1  SEIKAI CREST OF THE STARS  MORIOKA, HIROYUKI
2  SENTIENCE  ADAMS, TERRY A.
5  SENTINELS  MCKINNEY, JACK
1  SEPTIMUS HEAP  SAGE, ANGIE
2  SERAPH  BRIGGS, PATRICIA
5  SERENITY  authorship varies
2  SERPENT’S EGG  MCCURDY, J. FITSGERALD
5  SERPENTWAR  FEIST, RAYMOND E.
5  SERRATED EDGE  authorship varies
3  SERVANTS OF ARK  WYLIE, JONATHAN
3  SETTLING ACCOUNTS  TURTLEDOVE, HARRY
3  SEVEN BROTHERS  BENJAMIN, CURT
3  SEVEN CIRCLES  NILES, DOUGLAS
4  SEVEN CITADELS  HARRIS, GERALDINE
6  SEVEN SUNS  authorship varies
2  SEVEN TOWERS  SHELLEY, RICK
1  SEVENTH CARRIER  ALBANO, PETER
3  SEVENTH SWORD  DUNCAN, DAVE
3  SEVENWATERS  MARILLIER, JULIET
3  SEYONNE  BERG, CAROL
3  SHADITH’S QUEST  CLAYTON, JO
5  SHADOW  LOGSTON, ANNE
4  SHADOW OF THE LION  authorship varies

ENTRIES  SERIES  AUTHOR
1  SHADOW SAGA  GOLDEN, CHRISTOPHER
3  SHADOW WAR  authorship varies
3  SHADOW WARRIOR  BUNCH, CHRIS
4  SHADOW WORLD  HAMMELL, IAN
1  SHADOWKINGS  COBLEY, MICHAEL
3  SHADOWLEAGUE  FUREY, MAGGIE
4  SHADOWMARCH  WILLIAMS, TAD
43  SHADOWRUN  authorship varies
1  SHADOWRUN DRAGON HEART  KOKE, JAK
2  SHADOWRUN KELLAN COLT  KENSON, STEPHEN
3  SHADOWS  GOLDEN, CHRISTOPHER
3  SHADOWSONG  CLAYTON, JO
1  SHALRODAN SAGA  HOTCHKISS, ROBERT R.
1  SHAMER  KAABERSOL, LENE
1  SHANDRIL  GREENWOOD, ED
2  SHANJI INTERSTELLAR EMPIRE  GLASS, JAMES C.
4  SHANNARA  BROOKS, TERRY
4  SHANNARA HERITAGE  BROOKS, TERRY
3  SHANNARA JERLE  BROOKS, TERRY
3  SHAPER EXILE  FINCH, SHEILA
2  SHARING KNIFE  BUJOLD, LOIS MCMASTER
3  SHATTERED LIGHT  authorship varies
4  SHATTERZONE  authorship varies
8  SHELLPEOPLE  authorship varies
2  SHERWOOD  GODWIN, PARKE
2  SHEVA  BEAR, GREG
2  SHIELD AND SOURCE  MOORE, MOIRA J.
4  SHIP  authorship varies
3  SHIPS OF MERIOR  WURTS, JANNY
8  SHOLAN ALLIANCE  NORMAN, LISANNE
3  SHUNLAR  HELLER, CAROL
3  SIDHE  FLINT, KENNETH C.
1  SIDHE AND SHADOWLANDS  KELLEHER, ANNE
2  SIGNAL TO NOISE  NYLUND, ERIC S.
4  SILENCE LEIGH  SCOTT, MELISSA
4  SILENT EMPIRE  HARPER, STEVEN
4  SILISTRA  MORRIS, JANET
1  SILISTRA (REVISED)  MORRIS, JANET
2  SILKEN MAGIC  GILLIGAN, ELIZABETH
2  SILVER AGE SENTINELS  LOWDER, JAMES
2  SILVER CALL  MCKIERNAN, DENNIS L.
3  SILVER METAL LOVER  LEE, TANITH
2  SILVER WYRM  KERR, KATHARINE
4  SILVERGLASS  RIVKIN, J. F.
8  SIME/GEN  authorship varies
1  SIMS  WILSON, F. PAUL
3  SINGULARITY SKY  STROSS, CHARLES
2  SIR APROPOS OF NOTHING  DAVID, PETER
1  SISTERS OF THE NIGHT  YARBRO, CHELSEA QUINN
2  SISTERS OF THE RAVEN  HAMBLY, BARBARA
2  SIX KINGDOMS  FERGUSSON, BRUCE
6  SIX MILLION DOLLAR MAN  authorship varies
9  SIXTEEN THIRTY-TWO  authorship varies
1  SIXTH SENSE  BLOCH, JOHN W.
4  SKAITH  BRACKETT, LEIGH
3  SKEEN  CLAYTON, JO
1  SKILGANNON THE DAMNED  GEMMELL, DAVID
9  SKOLIAN EMPIRE  ASARO, CATHERINE
1  SKOLIAN TRIAD  ASARO, CATHERINE
2  SKY LORDS  BROSNAN, JOHN
3  SKYLA  BOYER, ELIZABETH H.
4  SKYLARK  SMITH, E. E.
5  SKYRIDER  MICHAELS, MELISA C.
3  SKYWAY  DE CHANCIE, JOHN
3  SLAVE TRADE  WRIGHT, SUSAN
3  SLOW WORLD  RIPLEY, KAREN
2  SLUGGY FREELANCE  ABRAMS, PETE
9  SMALLVILLE  authorship varies
4  SNOW QUEEN  VINGE, JOAN D.
2  SNOW-EYES  SMITH, STEPHANIE A.
3  SNOWFALL  SMITH, MITCHELL
3  SOCIETY OF HUMANITY/UNITY PARTY  MOSCOE, MIKE
3  SOLACE  ALLEN, ROGER MACBRIDE
7  SOLAR QUEEN  authorship varies

ENTRIES  SERIES  AUTHOR
2  SOLDIER OF THE MIST  WOLFE, GENE
2  SOLDIER SON  HOBB, ROBIN
5  SOLOMON KANE  HOWARD, ROBERT E.
3  SONG CALLED YOUTH  SHIRLEY, JOHN
3  SONG OF ALBION  LAWHEAD, STEPHEN R.
3  SONG OF EARTH  CONEY, MICHAEL G.
8  SONG OF ICE AND FIRE  authorship varies
2  SONG OF THE DWARVES  GUNNARSSON, THORARINN
4  SONG OF THE LIONESS  PIERCE, TAMORA
3  SONGKILLER  SCARBOROUGH, ELIZABETH ANN
1  SONGS OF THE STELLAR WIND  HICKMAN, TRACY
5  SONJA BLUE  COLLINS, NANCY A.
1  SOONE  SOMERS, JAMES
2  SORCERER’S TALE  DOUGLAS, L. WARREN
5  SORCERY!  JACKSON, STEVE
4  SOUL DRINKER  CLAYTON, JO
5  SOUL RIDER  CHALKER, JACK L.
3  SOULSMITH  DEITZ, TOM
9  SOUTHERN VAMPIRES  HARRIS, CHARLAINE
4  SOVEREIGN STONE  WEIS, MARGARET/HICKMAN, TRACY
2  SPACE ANGEL  ROBERTS, JOHN MADDOX
3  SPACE COPS  DUANE, DIANE/MORWOOD, PETER
2  SPACE HAWKS  PACKARD, EDWARD
2  SPACE MAVERICKS  KRING, MICHAEL
2  SPACE MERCHANTS  authorship varies
8  SPACE ODYSSEY  CLARKE, ARTHUR C.
3  SPACE PRECINCT  BISCHOFF, DAVID F.
2  SPACE: ABOVE AND BEYOND  TELEP, PETER
14  SPACE: 1999  authorship varies
7  SPACE: 1999 YEAR 2  BUTTERWORTH, MICHAEL
19  SPACEWAYS  CLEVE, JOHN
3  SPEARWIELDER  SALVATORE, R. A.
3  SPEC OPS SQUAD  SHELLEY, RICK
3  SPECIES IMPERATIVE  CZERNEDA, JULIE E.
4  SPECTROS  WINTERS, LOGAN
2  SPELL WARS  HALE, F. J.
6  SPELLJAMMER CLOAKMASTER  authorship varies
1  SPELLKEY  DOWNER, ANN
1  SPELLKEY (REVISED)  DOWNER, ANN
10  SPELLSINGER  FOSTER, ALAN DEAN
5  SPELLSONG  MODESITT, L. E. JR.
4  SPIDER  STOCKBRIDGE, GRANT
4  SPIDER WORLD  WILSON, COLIN
6  SPIDER-MAN  authorship varies
2  SPIDERMAN  authorship varies
1  SPIRITS OF ANCIENT SANDS  KASHINA, ANNA
3  SQUIRE  TELEP, PETER
1  ST. JOHN THE PURSUER  HENRICK, RICHARD
2  STACY GORDON  SHIPLETT, JUNE LUND
14  STAINLESS STEEL RAT  HARRISON, HARRY
2  STAIRWAY TO FOREVER  ADAMS, ROBERT
9  STAR COMMANDOS  GRIFFIN, P. M.
1  STAR CONTROL  QUICK, W. T.
2  STAR DRAGONS  CARVER, JEFFREY A.
1  STAR FRONTIER  HUDSON, EDWARD S.
2  STAR HAWKS  GOULART, RON
4  STAR OF THE GUARDIANS  WEIS, MARGARET
3  STAR OF THE GUARDIANS MAG FORCE 7  WEIS, MARGARET/PERRIN, DON
3  STAR PRECINCT  authorship varies
4  STAR REQUIEM  COLE, ADRIAN
5  STAR RISK, LTD.  authorship varies
3  STAR SHARDS  SHUSTERMAN, NEAL
4  STAR SOLDIERS  NORTON, ANDRE
2  STAR STONE  STASHEFF, CHRISTOPHER
245  STAR TREK  authorship varies
6  STAR TREK CAPTAIN’S TABLE  authorship varies
1  STAR TREK CHALLENGER  CAREY, DIANE
3  STAR TREK CRUCIBLE  GEORGE, DAVID R. III
4  STAR TREK DAY OF HONOR  authorship varies
4  STAR TREK DOMINION WAR  authorship varies
56  STAR TREK DS9  authorship varies
3  STAR TREK DS9 MILLENIUM  REEVES-STEVENS, JUDITH/REEVES-STEVENS, GARFIELD
4  STAR TREK DS9 MISSION GAMMA  authorship varies
3  STAR TREK DS9 REBELS  AB HUGH, DAFYDD

ENTRIES  SERIES  AUTHOR
1  STAR TREK DS9 SECTION 31  WEDDLE, DAVID/LANG, JEFFREY
3  STAR TREK DS9 WORLDS  authorship varies
11  STAR TREK ENTERPRISE  authorship varies
2  STAR TREK ERRAND OF FURY  RYAN, KEVIN
3  STAR TREK ERRAND OF VENGEANCE  RYAN, KEVIN
7  STAR TREK GATEWAYS  authorship varies
3  STAR TREK I.K.S. GORKON  DECANDIDO, KEITH R. A.
4  STAR TREK INVASION!  authorship varies
4  STAR TREK JANUS GATE  GRAF, L. A.
5  STAR TREK KEY COLLECTION  authorship varies
1  STAR TREK LOST ERA (2298 AD)  MARTIN, MICHAEL A./MANGELS, ANDY
1  STAR TREK LOST ERA (2311 AD)  GEORGE, DAVID R. III
1  STAR TREK LOST ERA (2328-46 AD)  DECANDIDO, KEITH R. A.
1  STAR TREK LOST ERA (2336 AD)  BICK, ILSA J.
1  STAR TREK LOST ERA (2355-57 AD)  MARIOTTE, JEFF
1  STAR TREK LOST ERA (2360 AD)  BONANNO, MARGARET WANDER
5  STAR TREK MIRROR  authorship varies
3  STAR TREK MY BROTHER’S KEEPER  FRIEDMAN, MICHAEL JAN
6  STAR TREK NEW EARTH  authorship varies
18  STAR TREK NEW FRONTIER  DAVID, PETER
3  STAR TREK ODYSSEY  authorship varies
6  STAR TREK RIHANNSU  authorship varies
8  STAR TREK S.C.E.  authorship varies
1  STAR TREK SAMUEL COGLEY  INGERSOLL, BOB/ISABELLA, TONY
1  STAR TREK SECTION 31  PERRY, S. D.
3  STAR TREK STARFLEET ACADEMY  authorship varies
3  STAR TREK TITAN  authorship varies
102  STAR TREK TNG  authorship varies
6  STAR TREK TNG DOUBLE HELIX  authorship varies
2  STAR TREK TNG MAXIMUM WARP  GALANTER, DAVE/BRODEUR, GREG
1  STAR TREK TNG SECTION 31  MANGELS, ANDY/MARTIN, MICHAEL A.
11  STAR TREK TNG STARFLEET ACADEMY  authorship varies
7  STAR TREK TNG STARGAZER  FRIEDMAN, MICHAEL JAN
9  STAR TREK TNG TIME TO  authorship varies
3  STAR TREK VANGUARD  authorship varies
28  STAR TREK VOYAGER  authorship varies
3  STAR TREK VOYAGER DARK MATTERS  GOLDEN, CHRISTIE
2  STAR TREK VOYAGER HOMECOMING  GOLDEN, CHRISTIE
1  STAR TREK VOYAGER SECTION 31  SMITH, DEAN WESLEY/RUSCH, KRISTINE KATHRYN
2  STAR TREK VOYAGER SPIRIT WALK  GOLDEN, CHRISTIE
1  STAR TREK VOYAGER STARFLEET ACADEMY  BARNES-SVARNEY, PATRICIA
3  STAR TREK VOYAGER STRING THEORY  authorship varies
3  STAR TREK VULCAN’S SOUL  SHERMAN, JOSEPHA/SHWARTZ, SUSAN
1  STAR TREK, THE NEXT GENERATION  FRIEDMAN, MICHAEL JAN/KRAUSE, PETER/MARCO, PABLO
52  STAR WARS  authorship varies
3  STAR WARS BLACK FLEET CRISIS  KUBE-MCDOWELL, MICHAEL P.
4  STAR WARS BOUNTY HUNTER WARS  JETER, K. W.
3  STAR WARS CLONE WARS  authorship varies
2  STAR WARS CLONE WARS MEDSTAR  REAVES, MICHAEL/PERRY, STEVE
1  STAR WARS CLONE WARS YODA  STEWART, SEAN
3  STAR WARS CORELLIAN  ALLEN, ROGER MACBRIDE
2  STAR WARS DARK FORCES  DIETZ, WILLIAM C.
1  STAR WARS DARK NEST  DENNING, TROY
1  STAR WARS DARTH BANE  KARPYSHYN, DREW
1  STAR WARS GALAXIES  WHITNEY-ROBINSON, VORONICA/BLACKMAN, HADEN
5  STAR WARS GALAXY OF FEAR  WHITMAN, JOHN
7  STAR WARS HAN SOLO  authorship varies
3  STAR WARS JEDI ACADEMY  ANDERSON, KEVIN J.
1  STAR WARS LAST OF THE JEDI  WATSON, JUDE
5  STAR WARS LEGACY OF THE FORCE  authorship varies
6  STAR WARS MOVIE  authorship varies
9  STAR WARS NEW JEDI ORDER  authorship varies
3  STAR WARS NEW JEDI ORDER AGENTS OF CHAOS  LUCENO, JAMES
3  STAR WARS NEW JEDI ORDER DARK TIDE  STACKPOLE, MICHAEL A.
3  STAR WARS NEW JEDI ORDER EDGE OF VICTORY  KEYES, J. GREGORY
3  STAR WARS NEW JEDI ORDER ENEMY LINES  ALLSTON, AARON
4  STAR WARS NEW JEDI ORDER FORCE HERETIC  WILLIAMS, SEAN/DIX, SHANE
2  STAR WARS REPUBLIC COMMANDO  TRAVISS, KAREN
9  STAR WARS X-WING  authorship varies
14  STAR WARS YOUNG JEDI KNIGHTS  ANDERSON, KEVIN J./MOESTA, REBECCA
7  STAR WOLF  GERROLD, DAVID
4  STAR WRECK  REWOLINSKI, LEAH
5  STAR*DRIVE  authorship varies

ENTRIES  SERIES  AUTHOR
3  STAR*DRIVE HARBINGER  DUANE, DIANE
7  STARBRIDGE  authorship varies
4  STARBRIDGE CHRONICLES  PARK, PAUL
4  STARCRAFT  authorship varies
1  STARCRAFT DARK TEMPLAR  GOLDEN, CHRISTIE
1  STARCRAFT GHOST  DECANDIDO, KEITH R. A.
6  STARCRUISER SHENANDOAH  GREEN, ROLAND J.
6  STARDANCE  ROBINSON, SPIDER/ROBINSON, JEANNE
7  STARDOC  VIEHL, S. L.
1  STARDOC JORY RASK  VIEHL, S. L.
4  STARFARERS  MCINTYRE, VONDA N.
5  STARFIRE  authorship varies
4  STARFISHERS  COOK, GLEN
13  STARFIST  SHERMAN, DAVID/CRAGG, DAN
2  STARFIST FORCE RECON  SHERMAN, DAVID/CRAGG, DAN
7  STARGATE  authorship varies
5  STARGATE ATLANTIS  authorship varies
9  STARGATE SG-1  authorship varies
4  STARGATE SG-1 (MCCONNELL)  MCCONNELL, ASHLEY
1  STARGAZER  HAUSMAN, GERALD
1  STARGODS  RADFORD, IRENE
5  STARHAWK  MALONEY, MACK
6  STARHAWKS  DALTON, SEAN
3  STARHOUNDS  BISCHOFF, DAVID F.
3  STARHUNTERS  DRAKE, DAVID
4  STARS  FLYNN, MICHAEL
3  STARS & STRIPES  HARRISON, HARRY
3  STARSEA INVADERS  STINE, G. HARRY
3  STARSHIELD  WEIS, MARGARET/HICKMAN, TRACY
1  STARSHIP  RESNICK, MIKE
3  STARSHIP ORPHEUS  JADE, SYMON
1  STARSHIP THEODORE ROOSEVELT  RESNICK, MIKE
3  STARSHIP TROUPERS  STASHEFF, CHRISTOPHER
1  STARSTRIKE  DOCKERY, KEVIN/NILES, DOUGLAS
4  STARWOLF  HAMILTON, EDMOND
4  STARWOLVES  GUNNARSSON, THORARINN
2  STEELE  MASTERS, J. D.
2  STEELEYE  DUNN, SAUL
2  STEERSWOMAN  KIRSTEIN, ROSEMARY
2  STELLAR RANGER  PERRY, STEVE
8  STEN  COLE, ALLAN/BUNCH, CHRIS
3  STEPHEN FISHER  ROHAN, MICHAEL SCOTT
2  STINKER  SERVICE, PAMELA F.
3  STONE  EDWARDS, GRAHAM
2  STONE DANCE OF THE CHAMELEON  PINTO, RICARDO
2  STONEFORT, MAINE  HETLEY, JAMES A.
6  STONES OF POWER  GEMMELL, DAVID
5  STORMLANDS  ROBERTS, JOHN MADDOX
5  STRAND  LEE, JOHN
5  STRANDS OF STARLIGHT  BAUDINO, GAEL
1  STRONTIUM DOG  LEVENE, REBECCA
2  STRUCTURE  SKETCHLEY, MARTIN
3  SUCCESSION  WESTERFELD, SCOTT
2  SUMMERLANDS  FOXXE, ELLEN
6  SUN SWORD  WEST, MICHELLE
3  SUN-CROSS  HAMBLY, BARBARA
6  SUNDERED  authorship varies
2  SUNDERING  CAREY, JACQUELINE
4  SUNSET WARRIOR  LUSTBADER, ERIC VAN
2  SUNSPACE  ZEBROWSKI, GEORGE
4  SUNWOLF AND STARHAWK  HAMBLY, BARBARA
2  SUPREME CONSTELLATIONS  BROOKE, GUN
18  SURVIVALIST  AHERN, JERRY
2  SURVIVAL 2000  MCPHEE, JAMES
3  SWANS’ WAR  RUSSELL, SEAN
3  SWORD & SORCERY DEAD GOD  BYERS, RICHARD LEE
3  SWORD AND CIRCLET  DOUGLAS, CAROLE NELSON
3  SWORD IN EXILE  MARSTON, ANN
4  SWORD OF CHANGE  BRAY, PATRICIA
2  SWORD OF HEAVEN  ELLIOTT, KATE
4  SWORD OF KNOWLEDGE  authorship varies
2  SWORD OF SHADOWS  JONES, J. V.
11  SWORD OF TRUTH  GOODKIND, TERRY
3  SWORD, RING AND CHALICE  CHESTER, DEBORAH

ENTRIES  SERIES  AUTHOR
4  SWORD, RING AND CHALICE PHERESA  CHESTER, DEBORAH
6  SWORDS  authorship varies
6  SWORDS OF RAEMLLYN  VARDEMAN, ROBERT E./PROCTOR, GEO. W.
3  SWORDSPOINT  authorship varies
3  SYMPATHY FOR THE DEVIL  authorship varies
6  SYMPHONY OF AGES  HAYDON, ELIZABETH
3  SYSTEMIC SHOCK  ING, DEAN
1  TAKEN  FOSTER, ALAN DEAN
3  TAKESHI KOVACS  MORGAN, RICHARD
4  TALE OF THE FIVE  DUANE, DIANE
1  TALES FROM THE ONE-EYED CROW  MCKIERNAN, DENNIS L.
2  TALES OF GOM  CHETWIN, GRACE
2  TALES OF GUINEVERE  BORCHARDT, ALICE
2  TALES OF THE BARD  SCOTT, MICHAEL
2  TALES OF THE WEREWOLF CLAN  MUNN, H. WARNER
4  TALES OF TOMORROW  HOUSTON, DAVID
2  TALISWOMAN  DOUGLAS, CAROLE NELSON
2  TAMSON HOUSE  DE LINT, CHARLES
3  TAMULI  EDDINGS, DAVID
2  TANAGARAN  SMEDS, DAVE
3  TANAQUIL  LEE, TANITH
2  TAORMIN  FRANKLIN, CHERYL J.
4  TAROT  ANTHONY, PIERS
1  TAROT DECK  PELZ, BRUCE
31  TARZAN  authorship varies
1  TAVAN  BURTON, JOSH MICHAEL
3  TAWNY MAN  HOBB, ROBIN
3  TAYLOR’S ARK  NYE, JODY LYNN
4  TEARS OF ARTAMON  ASH, SARAH
1  TEDDY LONDON, P.I.  MORGAN, ROBERT
1  TEETH  HUCKELBERY, TIMOTHY LEE
9  TEK  SHATNER, WILLIAM
2  TEKUMEL  BARKER, M. A. R.
3  TELNARIAN  NORMAN, JOHN
7  TELZEY  authorship varies
3  TEMERAIRE  authorship varies
1  TEMPS  GAIMAN, NEIL/STEWART, ALEX
3  TEN THOUSAND WILLOWS  GRANT, KATHRYN
3  TENEBREA  DAWSON, ROXANN/GRAHAM, DANIEL
3  TENTH PLANET  SMITH, DEAN WESLEY/RUSCH, KRISTINE KATHRYN
4  TERMINATOR  authorship varies
3  TERMINATOR T 2  STIRLING, S. M.
1  TERMINATOR 2  TIEDEMANN, MARK W.
3  TERMINATOR 2 NEW JOHN CONNOR  BLACKFORD, RUSSELL
1  TERMINATOR 3  ALLSTON, AARON
4  TERRA MAGICA  CARTER, LIN
1  TERRA NOVA  OHLANDER, BEN/DRAKE, DAVID
5  TERRILIAN  GREEN, SHARON
1  TESSERACTS  HOPKINSON, NALO/RYMAN, GEOFF
2  TEX MURPHY  CONNERS, AARON
3  TEXAS  DA CRUZ, DANIEL
8  THERE WILL BE WAR  POURNELLE, JERRY
2  THESE LAWLESS WORLDS  COMSTOCK, JARROD
20  THIEVES’ WORLD  authorship varies
2  THIEVES’ WORLD SHADOWSPAWN  OFFUTT, ANDREW J.
7  THIEVES’ WORLD TEMPUS  authorship varies
5  THLASSA MEY  MCCARTY, DENNIS
6  THOMAS COVENANT  DONALDSON, STEPHEN R.
2  THORGRIM  BAILEY, GERALD EARL
4  THOUSAND CULTURES  BARNES, JOHN
2  THRAXAS  SCOTT, MARTIN
3  THREE CALIFORNIAS  ROBINSON, KIM STANLEY
3  THREE DAMOSELS  CHAPMAN, VERA
4  THREE KINGS  CHALKER, JACK L.
3  THREE LANDS  EVANS, D. R.
4  THREE WORLDS  WATT-EVANS, LAWRENCE
1  THRILLING TALES  ANDERSON, M. T.
1  THRONE OF SHERRAN  GREEN, ROLAND J./MURRAY, FRIEDA A.
4  THURSDAY NEXT  FFORDE, JASPER
3  TIELMARAN  REIMANN, KATYA
6  TIGER  ROBERSON, JENNIFER
2  TIGUAFAYA  GRAY, JULIA
2  TIME AND AGAIN  FINNEY, JACK
2  TIME GATE  SILVERBERG, ROBERT/FAWCETT, BILL

ENTRIES  SERIES  AUTHOR
10  TIME HUNTER  authorship varies
3  TIME KEEPER  BARTHOLOMEW, BARBARA
2  TIME LOTTERY  MOSER, NANCY
3  TIME MACHINE  authorship varies
4  TIME MASTER  COOPER, LOUISE
4  TIME ODYSSEY  CLARKE, ARTHUR C./BAXTER, STEPHEN
2  TIME OF THE SIXTH SUN  HARLAN, THOMAS
3  TIME OF TROUBLES  TURTLEDOVE, HARRY
3  TIME POLICE  authorship varies
5  TIME SCOUT  ASPRIN, ROBERT LYNN/EVANS, LINDA
3  TIME STATION  EVANS, DAVID
3  TIME STREAM  EDMONDSON, G. C.
2  TIME TOURS  authorship varies
10  TIME TRADERS  authorship varies
3  TIME TRAP  DALTON, SEAN
1  TIME TRAVELER  LYONS, STEVE
2  TIME TUNNEL  LEINSTER, MURRAY
2  TIME WARRIOR  COSTELLO, MATTHEW J.
2  TIME WARRIORS  NORTH, DAVID
1  TIME’S TAPESTRY  BAXTER, STEPHEN
3  TIMECOP  PARKINSON, DAN
3  TIMEDIVERS  MODESITT, L. E. JR.
1  TIMELAPSE  NIGHBERT, DAVID E.
3  TIMELINE WARS  BARNES, JOHN
3  TIMEQUEST  TEDFORD, WILLIAM
2  TIMERAIDER  BARNES, JOHN
3  TIMESHARE  DANN, JOSHUA
12  TIMEWARS  HAWKE, SIMON
2  TIMOTHY VALENTINE  SOMTOW, S. P.
3  TIMURA  COLE, ALLAN
2  TINKER  SPENCER, WEN
3  TIR ALAINN  BISHOP, ANNE
4  TIR TANAGIRI  WALTON, JO
2  TITAN A.E.  ANDERSON, KEVIN J./MOESTA, REBECCA
3  TITANS  ADKINS, PATRICK H.
4  TITUS CROW  LUMLEY, BRIAN
4  TO THE STARS  HARRISON, HARRY
3  TOBIN OF SKALA  FLEWELLING, LYNN
3  TOLLER MARAQUINE  SHAW, BOB
9  TOM CORBETT  authorship varies
38  TOM SWIFT  APPLETON, VICTOR
12  TOM SWIFT (ARCHWAY)  APPLETON, VICTOR
33  TOM SWIFT JR.  APPLETON, VICTOR II
2  TOMMASO  ROESSNER, MICHAELA
3  TOMOE GOZEN  SALMONSON, JESSICA AMANDA
3  TONY FOSTER  HUFF, TANYA
2  TORCH OF HONOR  ALLEN, ROGER MACBRIDE
5  TORG POSSIBILITY WARS  authorship varies
3  TORIN  WILDER, CHERRY
3  TORIN KERR  HUFF, TANYA
3  TORNOR  LYNN, ELIZABETH A.
2  TOUR OF THE MERRIMACK  MELUCH, R. M.
3  TRADE PACT  CZERNEDA, JULIE E.
13  TRAVELER  DRUMM, D. B.
1  TRAVELLER  ASKEGREN, PIERCE
1  TRIAD  ASARO, CATHERINE
1  TRICKSTER’S GAME  CAMPBELL, BARBARA
3  TRIFFIDS  authorship varies
3  TRIGON DISUNITY  KUBE-MCDOWELL, MICHAEL P.
5  TRILLIUM  authorship varies
3  TRILOGY THAT DARE NOT SPEAK ITS NAME  RANKIN, ROBERT
1  TRIO  MONTGOMERY, R. A.
4  TRIPODS  CHRISTOPHER, JOHN
6  TRISTEN  CHERRYH, C. J.
1  TROPHY CHASE  POLIVKA, GEORGE BRYAN
4  TROS OF SAMOTHRACE  MUNDY, TALBOT
2  TROY  GEMMELL, DAVID
4  TROY GAME  DOUGLASS, SARA
8  TRUE GAME  TEPPER, SHERI S.
3  TURING HOPPER  ANDREWS, DONNA
3  TWELVE HOUSES  SHINN, SHARON
4  TWELVE KINGDOMS  FRIESNER, ESTHER M.
3  TWELVE TREASURES  EDGHILL, ROSEMARY
1  TWENTY NINETY-NINE  PEEL, JOHN

ENTRIES  SERIES  AUTHOR
17  TWILIGHT  authorship varies
8  TWILIGHT OF THE CLANS  authorship varies
3  TWILIGHT OF THE GODS  SCHMIDT, DENNIS
1  TWILIGHT ZONE  MILLER, JOHN J.
2  TWINS OF PETAYBEE  MCCAFFREY, ANNE/SCARBOROUGH, ELIZABETH ANN
3  TWO-EDGED SWORD  MARTIN, THOMAS K.
5  TY MERRICK  VITOLA, DENISE
2  TYPHON  ANZETTI, TONI
3  TYRANTS AND KINGS  MARCO, JOHN
1  U.S.S. REPULSE  DE CHANCIE, JOHN
2  UFO  MIALL, ROBERT
4  UKIAH OREGON  SPENCER, WEN
2  ULTIMA  ABBEY, LYNN
3  ULTIMA TECHNOCRAT WAR  ANDREWS, AUSTEN
3  UNBALANCED EARTH  WYLIE, JONATHAN
4  UNDERSEA  POHL, FREDERIK/WILLIAMSON, JACK
3  UNDERWORLD  COX, GREG
5  UNEXPECTED DRAGON  BROWN, MARY
2  UNICORN AND DRAGON  ABBEY, LYNN
3  UNIFIED WORLDS  THORNLEY, DIANN
6  UNION/ALLIANCE  CHERRYH, C. J.
4  UNION/ALLIANCE CYTEEN  CHERRYH, C. J.
2  UNREAL  authorship varies
8  UPLIFT  BRIN, DAVID
3  USS SKIPJACK  LANG, SIMON
16  V  authorship varies
9  VALDEMAR  authorship varies
3  VALENTINE MARSH  CHARNAS, SUZY MCKEE
5  VAMPIRE EARTH  KNIGHT, E. E.
10  VAMPIRE FILES  ELROD, P. N.
3  VAMPIRE HUNTRESS  BANKS, L. A.
1  VAMPIRE KISSES  SHCREIBER, ELLEN
3  VAMPIRE LEGACY  TAYLOR, KAREN E.
1  VAMPIRE WORLD  LUMLEY, BRIAN
5  VAMPIRELLA  GOULART, RON
1  VAMPYRICON  CLEGG, DOUGLAS
44  VANCE INTEGRAL EDITION  VANCE, JACK
2  VANG  ROWLEY, CHRISTOPHER
3  VARAYAN  SHELLEY, RICK
1  VARIS  PALMATIER, JOSHUA
3  VARNEY  RYMER, JAMES
1  VARNEY THE VAMPIRE  RYMER, JAMES
4  VATTA’S WAR  MOON, ELIZABETH
1  VEIL  GOLDEN, CHRISTOPHER
4  VELVET COMET  RESNICK, MIKE
8  VENUS  authorship varies
3  VENUS OF DREAMS  SARGENT, PAMELA
6  VENUS PRIME  PREUSS, PAUL
3  VERESHCHAGIN  FREZZA, ROBERT
2  VICTOR RENQUIST  FARREN, MICK
5  VICTORY NELSON  HUFF, TANYA
6  VIDESSOS  TURTLEDOVE, HARRY
4  VIEW FROM THE MIRROR  IRVINE, IAN
2  VINTAGE SEASON  authorship varies
2  VIOLETS  WOODWORTH, STEPHEN
1  VIRGA  SCHROEDER, KARL
6  VIRICONIUM  HARRISON, M. JOHN
3  VIRTUAL REALITY  BESHER, ALEXANDER
1  VIRTUAL WORLD  FINDLEY, NIGEL D.
3  VIS  LEE, TANITH
3  VISION OF BEASTS  LOVEJOY, JACK
3  VOID & WORD  BROOKS, TERRY
6  VON BEK  MOORCOCK, MICHAEL
2  VOR THE MAELSTROM  authorship varies
3  VOWS AND HONOR  LACKEY, MERCEDES
3  VOYAGE OF THE JERLE SHANNARA  BROOKS, TERRY
3  VOYAGERS  BOVA, BEN
3  WALKER  MURPHY, C. E.
4  WANDOR  GREEN, ROLAND J.
4  WAR AGAINST THE CHTORR  GERROLD, DAVID
3  WAR AGAINST THE CHTORR (ABRIDGED)  GERROLD, DAVID
3  WAR BETWEEN THE PROVINCES  TURTLEDOVE, HARRY
6  WAR OF POWERS  VARDEMAN, ROBERT E./MILAN, VICTOR
1  WAR OF THE DREAMING  WRIGHT, JOHN C.

ENTRIES  SERIES  AUTHOR
3  WAR OF THE GODS ON EARTH  OFFUTT, ANDREW J.
1  WAR OF THE ROSE  BRYAN, KATHLEEN
3  WAR OF THE WIZARDS  OFFUTT, ANDREW J./LYON, RICHARD K.
1  WAR OF THE WORLDS  DILLARD, J. M.
2  WAR SURPLUS  WATT-EVANS, LAWRENCE
2  WAR WITH THE MEIN  DURHAM, DAVID ANTHONY
6  WAR WORLD  authorship varies
3  WAR YEARS  FAWCETT, BILL
6  WARBOTS  STINE, G. HARRY
2  WARCHILD  BOWES, RICHARD
4  WARCRAFT  authorship varies
2  WARCRAFT SUNWELL  KNAAK, RICHARD A.
3  WARCRAFT WAR OF THE ANCIENTS  KNAAK, RICHARD A.
2  WARD OF HUROG  BRIGGS, PATRICIA
5  WARE  RUCKER, RUDY
36  WARHAMMER  authorship varies
3  WARHAMMER ANGELIKA FLEISCHER  LAWS, ROBIN D.
2  WARHAMMER BLACK HEARTS  LONG, NATHAN
2  WARHAMMER BLOOD ON THE REIK  MITCHELL, SANDY
1  WARHAMMER BLOODQUEST  RENNIE, GORDON/MACNEIL, COLIN
2  WARHAMMER BRUNNER  WERNER, C. L.
1  WARHAMMER DARKBLADE  ABNETT, DAN
4  WARHAMMER GENEVIEVE  YEOVIL, JACK
1  WARHAMMER GILEAD  ABNETT, DAN/VINCENT, NIK
9  WARHAMMER GOTREK & FELIX  KING, WILLIAM
4  WARHAMMER INQUISITION WAR  WATSON, IAN
1  WARHAMMER INQUISITOR ASCENDANT  ABNETT, DAN
4  WARHAMMER KONRAD  FERRING, DAVID
4  WARHAMMER MALUS DARKBLADE  ABNETT, DAN/LEE, MIKE
2  WARHAMMER MARK OF CHAOS  authorship varies
3  WARHAMMER ORFEO  CRAIG, BRIAN
3  WARHAMMER SLAVES TO DARKNESS  THORPE, GAV
2  WARHAMMER STEFAN KUMANSKY  MCINTOSH, NEIL
1  WARHAMMER VON CARSTEIN  SAVILE, STEVEN
2  WARHAMMER VON VELTEN  MCNEILL, GRAHAM
1  WARHAMMER WITCH HUNTER  WERNER, C. L.
1  WARHAMMER ZAVANT KONNIGER  RENNIE, GORDON
32  WARHAMMER 40K  authorship varies
2  WARHAMMER 40K BLACK TEMPLARS  GREEN, JONATHAN
2  WARHAMMER 40K BLOOD ANGELS  SWALLOW, JAMES
5  WARHAMMER 40K CIAPHAS CAIN  MITCHELL, SANDY
1  WARHAMMER 40K DAEMONIFUGE  CAMPBELL, JIM/SPURRIER, SIMON/RENNIE, GORDON/WALKER, KEV
2  WARHAMMER 40K DAWN OF WAR  GOTO, C. S.
2  WARHAMMER 40K DEATHWATCH  GOTO, C. S.
4  WARHAMMER 40K EISENHORN  ABNETT, DAN
1  WARHAMMER 40K ELDAR  KING, WILLIAM
10  WARHAMMER 40K GAUNT’S GHOSTS  ABNETT, DAN
2  WARHAMMER 40K GREY KNIGHTS  COUNTER, BEN
4  WARHAMMER 40K HORUS HERESY  authorship varies
3  WARHAMMER 40K LAST CHANCERS  THORPE, GAV
3  WARHAMMER 40K RAVENOR  ABNETT, DAN
2  WARHAMMER 40K SHIRA CALPURNIA  FARRER, MATTHEW
1  WARHAMMER 40K SISTERS OF BATTLE  SWALLOW, JAMES
3  WARHAMMER 40K SOUL DRINKERS  COUNTER, BEN
5  WARHAMMER 40K SPACE WOLF  KING, WILLIAM
4  WARHAMMER 40K ULTRAMARINES  MCNEILL, GRAHAM
3  WARHORSE OF ESDRAGON  DEXTER, SUSAN
1  WARKEEP 2030  KASNER, MICHAEL
18  WARLOCK  STASHEFF, CHRISTOPHER
5  WARLOCK’S HEIRS  STASHEFF, CHRISTOPHER
6  WARLORD  FROST, JASON
6  WARRIOR  authorship varies
2  WARRIOR OF VENGEANCE  COE, ROSS ANTON
1  WARRIORS NEW PROPHECY  HUNTER, ERIN
1  WARRIORS OF ESTAVIA  PATTON, FIONA
3  WARRIORS OF SPIDER  GEAR, W. MICHAEL
7  WARS OF LIGHT AND SHADOW  WURTS, JANNY
6  WARSTRIDER  KEITH, WILLIAM H. JR.
1  WARTORN  ASPRIN, ROBERT LYNN/DEL CARLO, ERIC
2  WASTEWORLD  BARTON, JAMES
4  WATCHERS  authorship varies
4  WATCHERS (CIRCLE LIGHTFALL)  SAINTCROW, LILITH
2  WATER  DALKEY, KARA
3  WATER!  BAUDINO, GAEL

ENTRIES  SERIES  AUTHOR
3  WATERSHED  NILES, DOUGLAS
2  WATERSHIP DOWN  ADAMS, RICHARD
6  WAYFARER REDEMPTION  DOUGLASS, SARA
3  WEAPONS OF CHAOS  VARDEMAN, ROBERT E.
5  WEATHER WARDEN  CAINE, RACHEL
3  WEB  BRADLEY, MARION ZIMMER
3  WEB SHIFTERS  CZERNEDA, JULIE E.
2  WEB WARRIORS  LUCENO, JAMES
5  WEB, THE  authorship varies
3  WEIRD HEROES  authorship varies
1  WEIRD ZONE  ABBOTT, TONY
8  WELL WORLD  CHALKER, JACK L.
3  WELL WORLD WATCHERS AT THE WELL  CHALKER, JACK L.
3  WELL-BUILT CITY  FORD, JEFFREY
6  WESS’HAR WARS  TRAVISS, KAREN
3  WESTERN LIGHTS  BARLOUGH, JEFFREY E.
1  WESTERN SHORE  LE GUIN, URSULA K.
5  WESTLANDS  KERR, KATHARINE
3  WESTMARK  ALEXANDER, LLOYD
8  WESTRIA  PAXSON, DIANA L.
5  WHAT MIGHT HAVE BEEN  BENFORD, GREGORY/GREENBERG, MARTIN H.
3  WHEEL  FRANKOS, STEVEN
16  WHEEL OF TIME  authorship varies
3  WHEN WORLDS COLLIDE  authorship varies
4  WHISPERING HILLS  CAMERON, KATE
1  WICKED DEAD  PETRUCHA, STEFAN/PENDLETON, THOMAS
5  WIDOWMAKER  RESNICK, MIKE
13  WILD CARDS  authorship varies
3  WILD CARDS NEW CYCLE  MARTIN, GEORGE R. R.
2  WILD HUNT  FOXE, JOCELIN
3  WILD MAGIC  CLAYTON, JO
2  WILD ROAD  KING, GABRIEL
4  WILDERNESS OF FOUR  HANCOCK, NIEL
2  WILLIAM SHAKESPEARE  HOYT, SARAH A.
1  WIND ON FIRE  NICHOLSON, WILLIAM
4  WINDAMEIR CIRCLE  HANCOCK, NIEL
4  WINDHOVER TAPES  NORWOOD, WARREN
4  WINDROSE  HAMBLY, BARBARA
5  WINDS OF THE FORELANDS  COE, DAVID B.
9  WING COMMANDER  authorship varies
6  WINGMAN  MALONEY, MACK
3  WINTER  SILVERBERG, ROBERT
3  WINTER KING’S WAR  DEXTER, SUSAN
4  WINTER OF THE WORLD  ROHAN, MICHAEL SCOTT
4  WINTER WORLD  MILLS, C. J.
4  WINTERLANDS  HAMBLY, BARBARA
2  WINTERLONG  HAND, ELIZABETH
1  WISHBONE  STEELE, ALEXANDER
6  WITCH WORLD  authorship varies
9  WITCH WORLD ESTCARP  NORTON, ANDRE
6  WITCH WORLD HIGH HALLACK  authorship varies
3  WITCH WORLD HIGH HALLACK GRYPHON  authorship varies
3  WITCH WORLD SECRETS  authorship varies
3  WITCH WORLD TURNING  authorship varies
4  WITCHBLADE  authorship varies
2  WITCHES OF CHISWICK  RANKIN, ROBERT
6  WITCHES OF EILEANAN  FORSYTH, KATE
2  WITCHES OF KARRES  authorship varies
6  WIZ ZUMWALT  COOK, RICK
2  WIZARD AND DRAGON  authorship varies
8  WIZARD IN RHYME  STASHEFF, CHRISTOPHER
2  WIZARD KNIGHT  WOLFE, GENE
5  WIZARD OF YURT  BRITTAIN, C. DALE
5  WIZARD WAR  COOK, HUGH
4  WIZARD’S WAR  BOYER, ELIZABETH H.
1  WIZARDRY  REAGEN, JAMES E.
2  WIZENBEAK  GILLILAND, ALEXIS A.
3  WODAN’S CHILDREN  PAXSON, DIANA L.
3  WOLF RICE  BORCHARDT, ALICE
2  WOLFKINGS  WOOD, BRIDGET
7  WOLFWALKER  HARPER, TARA K.
3  WOMAN OF THE IRON PEOPLE  ARNASON, ELEANOR
1  WOMEN OF GENESIS  CARD, ORSON SCOTT
5  WOMEN OF THE OTHERWORLD  ARMSTRONG, KELLEY

ENTRIES  SERIES  AUTHOR
3  WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE  ARMSTRONG, KELLEY
3  WONDERLAND GAMBIT  CHALKER, JACK L.
2  WOODEN SWORD  ABBEY, LYNN
1  WORKS OF KIPLING  KIPLING, RUDYARD
3  WORLD GATES  LISLE, HOLLY
2  WORLD IN AMBER  ORR, A.
3  WORLD OF ADEN THUNDERSCAPE  authorship varies
6  WORLD OF DARKNESS  authorship varies
2  WORLD OF DARKNESS DARK AGES CLAN  authorship varies
2  WORLD OF DARKNESS DEMON THE FALLEN  STOLZE, GREG
3  WORLD OF DARKNESS GRAILS COVENANT  WILSON, DAVID NIALL
1  WORLD OF DARKNESS MAGE  BASSINGTHWAITE, DON
1  WORLD OF DARKNESS ORPHEUS  BOULLE, PHILIPPE
6  WORLD OF DARKNESS PREDATOR & PREY  authorship varies
1  WORLD OF DARKNESS RAGE  BRIDGES, WILLIAM
1  WORLD OF DARKNESS RAGE JAY NO-NAME  MURRAY, DOUG
6  WORLD OF DARKNESS VAMPIRE  authorship varies
1  WORLD OF DARKNESS VAMPIRE DIARY  WEINBERG, ROBERT/REIN-HAGEN, MARK

3 WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD

CURSE

FLEMING, GHERBOD

9  WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN  authorship varies
4  WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA  WIECK, STEWART/FLEMING, GHERBOD/GRIFFIN, ERIC
1  WORLD OF DARKNESS VAMPIRE MASQUERADE RED DEATH  WEINBERG, ROBERT

2 WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN

AGE

BOULLE, PHILIPPE

3  WORLD OF DARKNESS WEREOLF  authorship varies
5  WORLD OF DARKNESS WEREWOLF  authorship varies
6  WORLD OF DARKNESS WRAITH  authorship varies
3  WORLD OF DARKNESS YEAR OF THE SCARAB  BATES, ANDREW
8  WORLD OF TIERS  FARMER, PHILIP JOSE
2  WORLD OF WARCRAFT  authorship varies
3  WORLDS  HALDEMAN, JOE
1  WORLDS OF CHTHON  PLATT, CHARLES
4  WORLDS OF HONOR  WEBER, DAVID
1  WORLDS OF POWER  MILES, ELLEN
7  WORLDWAR  TURTLEDOVE, HARRY
3  WORTHING  CARD, ORSON SCOTT
4  WRAETHTHU  CONSTANTINE, STORM
1  WRAETHTHU HISTORIES  CONSTANTINE, STORM
4  WUNTVOR  GARDNER, CRAIG SHAW
1  WYNDHAM WEREWOLF  DAVIDSON, MARYJANICE
10  WYRDRUNE  HAWKE, SIMON
1  X-COM  DUANE, DIANE
11  X-FILES  authorship varies
6  X-MEN  authorship varies
33  XANTH  authorship varies
5  XEELEE  BAXTER, STEPHEN
7  XENA  authorship varies
5  XENOGENESIS  BUTLER, OCTAVIA E.
3  XOR  HUNT, WALTER H.
4  YAMATO  KATO, KEN
2  YELINA  SNYDER, MARIA V.
5  YNELL  MURPHY, SHIRLEY ROUSSEAU
4  YNGLING  DALMAS, JOHN
3  YNIS AIELLE  SALVATORE, R. A.
2  YOLANDA  VERSEAU, DOMINIQUE
1  YOUNG MERLIN  YOLEN, JANE
8  YOUNG WIZARDS  DUANE, DIANE
4  ZACHARY NIXON JOHNSON  authorship varies
4  ZARKON  CARTER, LIN
2  ZENDA  HOPE, ANTHONY
5  ZIMIAMVIA  EDDISON, E. R.
2  ZODIAC  PETTERSON, VICKI
1  ZOE MARTINIQUE  WELDON, PHAEDRA
1  ZONE  ROUCH, JAMES
2  ZORACHUS  ROGERS, MARK E.
6  ZORK  authorship varies
2  ZOTHIQUE  SMITH, CLARK ASHTON

The MITSFS Seriesdex  September 24, 2007  Page 1
TITLE  AUTHOR  CODE
ABACAR THE WIZAR [ABACAR 1]  ERENBERGER, TIMOTHY  LP
ABACAR THE WIZARD AND THE DARK LORD OF DRAGONS [ABACAR 2]  ERENBERGER, TIMOTHY  LP
DAYS OF MAGIC, NIGHTS OF WAR [ABARAT 2]  BARKER, CLIVE  H
MEDUSA WAR, THE [@ABC WARRIORS]  MILLS, PAT/MITCHELL, ALAN  P
ABHORSEN TRILOGY, THE [ABHORSEN 1,2,3]  NIX, GARTH  H
ACROSS THE WALL [ABHORSEN]  NIX, GARTH  H
VELOCITY GOSPEL, THE [ACCOMPLICE 2]  AYLETT, STEVE  LP
ACORNA [ACORNA 1]  MCCAFFREY, ANNE/BALL, MARGARET  H

ACORNA’S PEOPLE [ACORNA 3] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H,P

ACORNA’S QUEST [ACORNA 2]  MCCAFFREY, ANNE/BALL, MARGARET  C/P,P,H

ACORNA’S REBELS [ACORNA 6] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

ACORNA’S REBELS (PROOF) [ACORNA 6] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

SR-LP

ACORNA’S SEARCH [ACORNA 5]  MCCAFFREY, ANNE  H

ACORNA’S TRIUMPH [ACORNA 7] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

P

ACORNA’S TRIUMPH (PROOF) [ACORNA 7] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

SR-LP

ACORNA’S WORLD [ACORNA 4] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

FIRST WARNING [ACORNA’S CHILDREN 1] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

SECOND WAVE [ACORNA’S CHILDREN 2] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

ADDAMS FAMILY, THE [ADDAMS FAMILY]  SHARKEY, JACK  C/P:2,P
ADDAMS FAMILY VALUES [ADDAMS FAMILY]  HODGMAN, ANN  P

DAGGER MAGIC [ADEPT 4] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,P,H

DEATH OF AN ADEPT [ADEPT 5] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

H,P

ADEPT, THE [@ADEPT #1] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,SFWA-P:2,P

LODGE OF THE LYNX, THE [@ADEPT #2] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,SFWA-P,P

TEMPLAR TREASURE, THE [@ADEPT #3] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,SFWA-P,P

KEEP, THE [ADVERSARY 1]  WILSON, F. PAUL  H,P
NIGHTWORLD [ADVERSARY 6]  WILSON, F. PAUL  P
REBORN [ADVERSARY 4]  WILSON, F. PAUL  P
REPRISAL [ADVERSARY 5]  WILSON, F. PAUL  P
TOMB, THE [ADVERSARY 2/REPAIRMAN JACK]  WILSON, F. PAUL  C/P:2,P
TOUCH, THE [ADVERSARY 3]  WILSON, F. PAUL  P
MASTER WALK [ADVOCACY 1]  LEE, SHARON/MILLER, STEVE  SR-LP
AEGYPT [AEGYPT 1]  CROWLEY, JOHN  H,LP
ENDLESS THINGS [AEGYPT]  CROWLEY, JOHN  H
LOVE AND SLEEP [AEGYPT 2]  CROWLEY, JOHN  LP
EYE OF THE SERPENT [AELWYN 2]  CHARRETTE, ROBERT N.  C/P,P
TIMESPELL [AELWYN 1]  CHARRETTE, ROBERT N.  P
WIZARD OF BONES [AELWYN 3]  CHARRETTE, ROBERT N.  P
AFRIKORPS [@AFRIKORPS #1]  DOLAN, BILL  P
IRON HORSE [@AFRIKORPS #2]  DOLAN, BILL  P
SEA STALLION [@AFRIKORPS #4]  DOLAN, BILL  P
LION MOUNTAIN [@AFRIKORPS #5]  DOLAN, BILL  P
SCION OF THE SERPENT [@AGE OF CONAN ANOK #1]  YORK, J. STEVEN  P
BLOOD OF WOLVES [@AGE OF CONAN KERN #1]  COLEMAN, LOREN L.  P
CIMMERIAN RAGE [@AGE OF CONAN KERN #2]  COLEMAN, LOREN L.  P
SONGS OF VICTORY [@AGE OF CONAN KERN #3]  COLEMAN, LOREN L.  P
GHOST OF THE WALL [@AGE OF CONAN MARAUDERS #1]  MARIOTTE, JEFF  P
WINDS OF THE WILD SEA [@AGE OF CONAN MARAUDERS #2]  MARIOTTE, JEFF  P
DAWN OF THE ICE BEAR [@AGE OF CONAN MARAUDERS #3]  MARIOTTE, JEFF  P
GOD IN THE MOON, THE [@AGE OF CONAN SOLDIER’S QUEST #1]  KNAAK, RICHARD A.  P
EYE OF CHARON, THE [@AGE OF CONAN SOLDIER’S QUEST #2]  KNAAK, RICHARD A.  P
SILENT ENEMY, THE [@AGE OF CONAN SOLDIER’S QUEST #3]  KNAAK, RICHARD A.  P
TYRANNOSAURUS REX [AGE OF DINOSAURS 1]  RIVKIN, J. F.  P,SFWA-P
CARTOMANCY [AGE OF DISCOVERY 2]  STACKPOLE, MICHAEL A.  LP
SECRET ATLAS, A [AGE OF DISCOVERY 1]  STACKPOLE, MICHAEL A.  H
DRAGON CHAMPION [@AGE OF FIRE #1]  KNIGHT, E. E.  LP
DRAGON AVENGER [@AGE OF FIRE #2]  KNIGHT, E. E.  LP
ALWAYS FOREVER [AGE OF MISRULE 3]  CHADBOURN, MARK  LP
DARKEST HOUR [AGE OF MISRULE 2]  CHADBOURN, MARK  P
WORLD’S END [AGE OF MISRULE 1]  CHADBOURN, MARK  P
LAST OF THE WILDS [AGE OF THE FIVE 2]  CANAVAN, TRUDI  P

The MITSFS Seriesdex  September 24, 2007  Page 2
TITLE  AUTHOR  CODE
PRIESTESS OF THE WHITE [AGE OF THE FIVE 1]  CANAVAN, TRUDI  P
VOICE OF THE GODS [AGE OF THE FIVE 3]  CANAVAN, TRUDI  P
CALCULUS OF ANGELS, A [AGE OF UNREASON 2]  KEYES, J. GREGORY  H:2,LP
EMPIRE OF UNREASON [AGE OF UNREASON 3]  KEYES, J. GREGORY  LP
NEWTON’S CANNON [AGE OF UNREASON 1]  KEYES, J. GREGORY  H:2,LP
SHADOWS OF GOD, THE [AGE OF UNREASON 4]  KEYES, J. GREGORY  H,LP
FLYING SAUCER GAMBIT, THE [@AGENT OF TERRA #1]  MADDOCK, LARRY  C/P:2,P
GOLDEN GODDESS GAMBIT, THE [@AGENT OF TERRA #2]  MADDOCK, LARRY  C/P:2,P
EMERALD ELEPHANT GAMBIT, THE [@AGENT OF TERRA #3]  MADDOCK, LARRY  C/P:2,P
TIME TRAP GAMBIT, THE [@AGENT OF TERRA #4]  MADDOCK, LARRY  C/P:2,P
FIRES OF AGGAR [AGGAR 2]  WOLFE, CHRIS ANNE  LP
SHADOWS OF AGGAR [AGGAR 1]  WOLFE, CHRIS ANNE  LP
AGGRESSOR SIX [AGGRESSOR SIX 1]  MCCARTHY, WIL  C/P,P
FALL OF SIRIUS, THE [AGGRESSOR SIX 2]  MCCARTHY, WIL  P
THIRTEENTH SCROLL, THE [AGHAMORE 2]  NEASON, REBECCA  P
TRUEST POWER, THE [AGHAMORE 1]  NEASON, REBECCA  P
CITY ON FIRE [AIAH 2]  WILLIAMS, WALTER JON  H,P
METROPOLITAN [AIAH 1]  WILLIAMS, WALTER JON  H,P
FIRE ME UP [AISLING GREY 2]  MACALISTER, KATIE  C/P,P
LIGHT MY FIRE [AISLING GREY 3]  MACALISTER, KATIE  C/P,P
YOU SLAY ME [AISLING GREY 4]  MACALISTER, KATIE  C/P,P
ALAMUT [ALAMUT 1]  TARR, JUDITH  C/P,LP
DAGGER AND THE CROSS, THE [ALAMUT 2]  TARR, JUDITH  H,P
ALASTOR [ALASTOR]  VANCE, JACK  H
MARUNE: ALASTOR 933 [ALASTOR 2]  VANCE, JACK  C/P:2,P
TRULLION: ALASTOR 2262 [ALASTOR 1]  VANCE, JACK  C/P:2,P
WYST: ALASTOR 1716 [ALASTOR 3]  VANCE, JACK  C/P:2,P
SOUTHERN FIRE [ALDABRESHIN COMPASS 1]  MCKENNA, JULIET E.  H
NORTHERN STORM [@ALDABRESHIN COMPASS #2]  MCKENNA, JULIET E.  P
WESTERN SHORE [@ALDABRESHIN COMPASS #3]  MCKENNA, JULIET E.  P
ALDAIR, ACROSS THE MISTY SEA [ALDAIR 3]  BARRETT, NEAL JR.  P
ALDAIR IN ALBION [ALDAIR 1]  BARRETT, NEAL JR.  P
ALDAIR, MASTER OF SHIPS [ALDAIR 2]  BARRETT, NEAL JR.  C/P,P
ALDAIR: THE LEGION OF BEASTS [ALDAIR 4]  BARRETT, NEAL JR.  P
MOON OF GOMRATH, THE [ALDERLY 2]  GARNER, ALAN  C/P:2,P
WEIRDSTONE OF BRISINGAMEN, THE [ALDERLY 1]  GARNER, ALAN  C/P:2,P
NORTH WIND [ALEUTIAN 2]  JONES, GWYNETH  H,LP
PHOENIX CAFE [ALEUTIAN 3]  JONES, GWYNETH  LP
WHITE QUEEN [ALEUTIAN 1]  JONES, GWYNETH  H
POLARIS [ALEX BENEDICT 2]  MCDEVITT, JACK  H
SEEKER [ALEX BENEDICT 3]  MCDEVITT, JACK  H
TALENT FOR WAR, A [ALEX BENEDICT 1]  MCDEVITT, JACK  H,SFWA-P,P
DARK MIRROR, DARK DREAMS [ALEXIA AND TIRAN 2/HIDDEN REALMS 2] GREEN, SHARON  P
SILVER PRINCESS, GOLDEN KNIGHT [ALEXIA AND TIRAN 1]  GREEN, SHARON  P,SFWA-P
WIND WHISPERS, SHADOW SHOUTS [ALEXIA AND TIRAN 3]  GREEN, SHARON  C/P,P
ELVES AND THE OTTERSKIN, THE [ALFAR 2]  BOYER, ELIZABETH H.  C/P:2,P
SWORD AND THE SATCHEL, THE [ALFAR 1]  BOYER, ELIZABETH H.  C/P:2,P
THRALL AND THE DRAGON’S HEART, THE [ALFAR 3]  BOYER, ELIZABETH H.  C/P:2,P
WIZARD AND THE WARLORD, THE [ALFAR 4]  BOYER, ELIZABETH H.  C/P:2,P
GOLDEN ONE, THE [@ALIEN CHRONICLES #1]  CHESTER, DEBORAH  C/P,P
CRIMSON CLAW, THE [@ALIEN CHRONICLES #2]  CHESTER, DEBORAH  P
CRYSTAL EYE, THE [@ALIEN CHRONICLES #3]  CHESTER, DEBORAH  P
AWAKEN ME DARKLY [ALIEN HUNTRESS 1]  SHOWALTER, GENA  LP
ALIEN NATION [ALIEN NATION]  FOSTER, ALAN DEAN  C/P:2,P
BODY AND SOUL [ALIEN NATION 3]  DAVID, PETER  C/P:2,P
PASSING FANCY [ALIEN NATION 6]  SPENCER, DAVID  C/P,P

DAY OF DESCENT, THE [@ALIEN NATION #1] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

DARK HORIZON [@ALIEN NATION #2]  JETER, K. W.  P
CHANGE, THE [@ALIEN NATION #4]  LONGYEAR, BARRY B.  C/P:2,P
SLAG LIKE ME [@ALIEN NATION #5]  LONGYEAR, BARRY B.  C/P,P
EXTREME PREJUDICE [@ALIEN NATION #7]  GRAF, L. A.  C/P,P
CROSS OF BLOOD [@ALIEN NATION #8]  JETER, K. W.  P
CLYPSIS [@ALIEN SPEEDWAY 1]  CARVER, JEFFREY A.  C/P:2,P
PITFALL [ALIEN SPEEDWAY 2]  WYLDE, THOMAS  C/P:2,P
PITFALL (PROOF) [ALIEN SPEEDWAY 2]  WYLDE, THOMAS  SR-LP
WEB, THE [ALIEN SPEEDWAY 3]  WYLDE, THOMAS  C/P:2,P
AFTER THE FLAMES [ALIEN STARS 2]  MITCHELL, ELIZABETH  PA
ALIEN STARS [ALIEN STARS 1]  MITCHELL, ELIZABETH  PA
FREE LANCERS [ALIEN STARS 4]  MITCHELL, ELIZABETH  C/PA,PA
UNDER THE WHEEL [ALIEN STARS 3]  MITCHELL, ELIZABETH  PA
ALIEN TRACE, THE [@ALIEN TRACE #1]  MAJOR, H. M.  P
TIME TWISTER [@ALIEN TRACE #2]  MAJOR, H. M.  P

The MITSFS Seriesdex  September 24, 2007  Page 3
TITLE  AUTHOR  CODE
ALIEN [ALIENS]  FOSTER, ALAN DEAN  H:2,P
ALIEN 3 [ALIENS]  FOSTER, ALAN DEAN  C/P,P
ALIEN HARVEST [ALIENS 5]  SHECKLEY, ROBERT  C/P,P
ALIEN RESURRECTION [ALIENS]  CRISPIN, A. C./O’MALLEY, KATHLEEN  P
ALIENS [ALIENS]  FOSTER, ALAN DEAN  P
ALIENS COLONIAL MARINES TECHNICAL MANUAL [@ALIENS]  BRIMMICOMBE-WOOD, LEE  VLP
AVP: ALIEN VS. PREDATOR [ALIENS/PREDATOR]  CERASINI, MARC  P
BERSERKER [@ALIENS 9]  PERRY, S. D.  P
GENOCIDE [@ALIENS 4]  BISCHOFF, DAVID F.  P
LABYRINTH [@ALIENS 7]  PERRY, S. D.  P
MUSIC OF THE SPEARS [@ALIENS 8]  NAVARRO, YVONNE  P
ROGUE [@ALIENS 6]  SCHOFIELD, SANDY  C/P,P
EARTH HIVE [@ALIENS #1]  PERRY, STEVE  C/P,P
NIGHTMARE ASYLUM [@ALIENS #2]  PERRY, STEVE  P
FEMALE WAR, THE [@ALIENS #3]  PERRY, STEVE/PERRY, STEPHANI  C/P,P
HUNTER’S PLANET [@ALIENS VS. PREDATOR 2]  BISCHOFF, DAVID F.  C/P,P
PREY [@ALIENS VS. PREDATOR 1]  PERRY, STEVE/PERRY, STEPHANI  P
WAR [@ALIENS VS. PREDATOR 3]  PERRY, S. D.  C/P,P
ALLAN’S WIFE [ALLAN QUATERMAIN]  HAGGARD, H. RIDER  LP
KING SOLOMON’S MINES [ALLAN QUATERMAIN]  HAGGARD, H. RIDER  C/P,P,H
RIDE THE LIGHTNING [ALO NUDGER]  LUTZ, JOHN  SFWA-P
CENTAURI DAWN [@ALPHA CENTAURI #1]  ELY, MICHAEL  P
DRAGON SUN [@ALPHA CENTAURI #2]  ELY, MICHAEL  P
TWILIGHT OF THE MIND [@ALPHA CENTAURI #3]  ELY, MICHAEL  P
ONE-ARMED QUEEN, THE [ALTA 3]  YOLEN, JANE  C/P,P
SISTER LIGHT, SISTER DARK [ALTA 1]  YOLEN, JANE  P,SFWA-P
WHITE JENNA [ALTA 2]  YOLEN, JANE  C/P,P,H
ALVIN JOURNEYMAN [ALVIN MAKER 4]  CARD, ORSON SCOTT  C/P:2,P
ALVIN WANDERING [ALVIN MAKER 4,5]  CARD, ORSON SCOTT  H
CRYSTAL CITY, THE [ALVIN MAKER 6]  CARD, ORSON SCOTT  H
HATRACK RIVER [ALVIN MAKER 1,2,3]  CARD, ORSON SCOTT  H
HEARTFIRE [ALVIN MAKER 5]  CARD, ORSON SCOTT  H,P
PRENTICE ALVIN [ALVIN MAKER 3]  CARD, ORSON SCOTT  C/P,SFWA-P,P,H
RED PROPHET [ALVIN MAKER 2]  CARD, ORSON SCOTT  H:2,SFWA-P,P
SEVENTH SON [ALVIN MAKER 1]  CARD, ORSON SCOTT  SFWA-P,H:2,P
BLOOD OF AMBER [AMBER 7]  ZELAZNY, ROGER  C/P,H,@P,SFWA
CHRONICLES OF AMBER 1, THE [AMBER 1,2,3]  ZELAZNY, ROGER  H:2
CHRONICLES OF AMBER 2, THE [AMBER 4,5]  ZELAZNY, ROGER  H:2
COMPLETE AMBER SOURCEBOOK, THE [AMBER]  KRULIK, THEODORE  R/REF-H
COURTS OF CHAOS, THE [AMBER 5]  ZELAZNY, ROGER  C/P,@P,H,SFWA
GREAT BOOK OF AMBER, THE [AMBER 1,2,3,4,5,6,7,8,9,10]  ZELAZNY, ROGER  LP
GUNS OF AVALON, THE [AMBER 2]  ZELAZNY, ROGER  C/P:3,@P
HAND OF OBERON, THE [AMBER 4]  ZELAZNY, ROGER  C/P,@P,H,SFWA
KNIGHT OF SHADOWS [AMBER 9]  ZELAZNY, ROGER  C/P,H,@P,SFWA
NINE PRINCES IN AMBER [AMBER 1]  ZELAZNY, ROGER  C/P:2,@P
PRINCE OF CHAOS [AMBER 10]  ZELAZNY, ROGER  H:2,@P
SEVEN NO-TRUMP [AMBER]  RANDALL, NEIL  SFWA-P
SIGN OF CHAOS [AMBER 8]  ZELAZNY, ROGER  C/P,H,@P
SIGN OF THE UNICORN [AMBER 3]  ZELAZNY, ROGER  C/P:2,@P
TRUMPS OF DOOM [AMBER 6]  ZELAZNY, ROGER  C/P,H,@P,SFWA
CITY OF SAINTS AND MADMEN [AMBERGRIS]  VANDERMEER, JEFF  LP
SHRIEK: AN AFTERWORD [AMBERGRIS]  VANDERMEER, JEFF  LP
DOOR TO AMBERMERE, THE [AMBERMERE 1]  PIERCE, J. CALVIN  C/P,P
SORCERESS OF AMBERMERE, THE [AMBERMERE 2]  PIERCE, J. CALVIN  P,SFWA-P
WIZARD OF AMBERMERE, THE [AMBERMERE 3]  PIERCE, J. CALVIN  P
AMERICA 2040 [AMERICA 2040 1]  INNES, EVAN  P
CITY IN THE MIST [AMERICA 2040 3]  INNES, EVAN  P
GOLDEN WORLD, THE [AMERICA 2040 2]  INNES, EVAN  C/P,P
RETURN, THE [AMERICA 2040 4]  INNES, EVAN  P
CLOUD WARRIOR [AMTRAK WARS 1]  TILLEY, PATRICK  C/P:2,P
FIRST FAMILY, THE [AMTRAK WARS 2]  TILLEY, PATRICK  C/P:2,P
IRON MASTER [AMTRAK WARS 3]  TILLEY, PATRICK  C/P,P
DEATH-BRINGER [@AMTRAK WARS #5]  TILLEY, PATRICK  P
WOLF BROTHER [ANCIENT DARKNESS 1]  PAVER, MICHELLE  H
"A" COMME ANDROMEDE [ANDROMEDA 1]  HOYLE, FRED/ELLIOT, JOHN  FP
A FOR ANDROMEDA [ANDROMEDA 1]  HOYLE, FRED/ELLIOT, JOHN  H:2,P
AMEACA DE ANDROMEDA [ANDROMEDA 1]  HOYLE, FRED/ELLIOT, JOHN  FP
ANDROMEDA BREAKTHROUGH [ANDROMEDA 2]  HOYLE, FRED/ELLIOT, JOHN  C/P,P,H
ATTITUDE OF SILENCE, THE [ANDROMEDA (GENE RODDENBERRY’S) 5]  MARIOTTE, JEFF  H
BROKEN PLACES, THE [ANDROMEDA (GENE RODDENBERRY’S)]  VARE, ETHLIE ANN/MORRIS, DANIEL  H
PARADISE DRIFT [ANDROMEDA (GENE RODDENBERRY’S)]  SMITH, SHERWOOD  H
DESTRUCTION OF ILLUSIONS [@ANDROMEDA (GENE RODDENBERRY’S) #1]  DECANDIDO, KEITH R. A.  P

TITLE  AUTHOR  CODE

THROUGH THE LOOKING GLASS [@ANDROMEDA (GENE RODDENBERRY’S) #4] SHERMAN, JOSEPHA

P

AVATAR [@ANGEL]  PASSARELLA, JOHN  P
BOOK OF THE DEAD [@ANGEL]  MCCONNELL, ASHLEY  P
BRUJA [@ANGEL]  ODOM, MEL  P
CITY OF [@ANGEL]  HOLDER, NANCY  P
CLOSE TO THE GROUND [@ANGEL]  MARIOTTE, JEFF  C/P,P
DARK MIRROR [@ANGEL]  GARDNER, CRAIG SHAW  C/P,P
ENDANGERED SPECIES [@ANGEL]  HOLDER, NANCY/MARIOTTE, JEFF  P
FEARLESS [@ANGEL]  DURGIN, DORANNA  P
HAUNTED [@ANGEL]  MARIOTTE, JEFF  P
HOLLYWOOD NOIR [@ANGEL]  MARIOTTE, JEFF  P
IMAGE [@ANGEL]  ODOM, MEL  P
IMPRESSIONS [@ANGEL]  DURGIN, DORANNA  P
LONGEST NIGHT 1, THE [@ANGEL]  TWENTIETH CENTURY FOX  PA
LOVE AND DEATH [@ANGEL]  MARIOTTE, JEFF  P
MONOLITH [@ANGEL]  PASSARELLA, JOHN  C/P,P
NEMESIS [@ANGEL]  CIENCIN, SCOTT/CIENCIN, DENISE  P
NOT FORGOTTEN [@ANGEL]  HOLDER, NANCY  P
REDEMPTION [@ANGEL]  ODOM, MEL  P
SANCTUARY [@ANGEL]  MARIOTTE, JEFF  P
SHAKEDOWN [@ANGEL]  DEBRANDT, DON  P
SOLITARY MAN [@ANGEL]  MARIOTTE, JEFF  P
SOUL TRADE [@ANGEL]  SNIEGOSKI, THOMAS E.  P
STRANGER TO THE SUN [@ANGEL]  MARIOTTE, JEFF  P
SUMMONED, THE [@ANGEL]  DOKEY, CAMERON  P
VENGEANCE [@ANGEL]  CIENCIN, SCOTT/JOLLEY, DAN  P
DESPERATE MEASURES [ANGEL’S LUCK 1]  FAUST, JOE CLIFFORD  C/P,P,H
ESSENCE OF EVIL, THE [ANGEL’S LUCK 3]  FAUST, JOE CLIFFORD  H,P
PRECIOUS CARGO [ANGEL’S LUCK 2]  FAUST, JOE CLIFFORD  H,P
CHANGER OF DAYS [ANGHARA 2]  ALEXANDER, ALMA  P
HIDDEN QUEEN, THE [ANGHARA 1]  ALEXANDER, ALMA  P
BLACK MOON INN [ANITA BLAKE 7,8]  HAMILTON, LAURELL K.  H
BLOODY BONES [ANITA BLAKE 5]  HAMILTON, LAURELL K.  C/P:2,P
BLUE MOON [ANITA BLAKE 8]  HAMILTON, LAURELL K.  C/P:2,P
BURNT OFFERINGS [ANITA BLAKE 7]  HAMILTON, LAURELL K.  C/P:2,P
CERULEAN SINS [ANITA BLAKE 11]  HAMILTON, LAURELL K.  C/P,H
CIRCUS OF THE DAMNED [ANITA BLAKE 3]  HAMILTON, LAURELL K.  C/P:2,P
CLUB VAMPYRE [ANITA BLAKE 1,2,3]  HAMILTON, LAURELL K.  H
DANSE MACABRE [ANITA BLAKE 14]  HAMILTON, LAURELL K.  H:2
GUILTY PLEASURES [ANITA BLAKE 1]  HAMILTON, LAURELL K.  C/P:2,P
HARLEQUIN, THE [ANITA BLAKE 15]  HAMILTON, LAURELL K.  H
INCUBUS DREAMS [ANITA BLAKE 12]  HAMILTON, LAURELL K.  H
KILLING DANCE, THE [ANITA BLAKE 6]  HAMILTON, LAURELL K.  C/P:2,P
LAUGHING CORPSE, THE [ANITA BLAKE 2]  HAMILTON, LAURELL K.  C/P:2,P
LUNATIC CAFE, THE [ANITA BLAKE 4]  HAMILTON, LAURELL K.  C/P:2,P
MACAH [ANITA BLAKE 13]  HAMILTON, LAURELL K.  P
MIDNIGHT CAFE, THE [ANITA BLAKE 4,5,6]  HAMILTON, LAURELL K.  H
NARCISSUS IN CHAINS [ANITA BLAKE 10]  HAMILTON, LAURELL K.  C/P,P,H
OBSIDIAN BUTTERFLY [ANITA BLAKE 9]  HAMILTON, LAURELL K.  C/P,P,H
BECOMING, THE [ANNA STRONG 1]  STEIN, JEANNE C.  P
BLOOD DRIVE [ANNA STRONG 2]  STEIN, JEANNE C.  P
ANNO DRACULA [ANNO DRACULA 1]  NEWMAN, KIM  H,P
BLOODY RED BARON, THE [ANNO DRACULA 2]  NEWMAN, KIM  P
ANTARES DAWN [ANTARES 1]  MCCOLLUM, MICHAEL  C/P:2,P
ANTARES PASSAGE [ANTARES 2]  MCCOLLUM, MICHAEL  C/P,P
ANTARES VICTORY [ANTARES 3]  MCCOLLUM, MICHAEL  LP
KINGDOMS OF THE NIGHT [ANTEROS]  COLE, ALLAN/BUNCH, CHRIS  C/P,P
WARRIOR RETURNS, THE [@ANTEROS]  COLE, ALLAN  H,P
CHILDREN OF ANTHI, THE [ANTHI]  BLAKENEY, JAY D.  P
REQUIEM FOR ANTHI [ANTHI]  BLAKENEY, JAY D.  P
MASQUE WORLD [ANTHONY VILLIERS 3]  PANSHIN, ALEXEI  C/P:2,P
STAR WELL [ANTHONY VILLIERS 1]  PANSHIN, ALEXEI  SR-P,C/P:2,P
THURB REVOLUTION, THE [ANTHONY VILLIERS 2]  PANSHIN, ALEXEI  SR-P,C/P:2
SEER, THE [ANTRIAN MESSENGER 2]  WISLER, G. CLIFTON  H
LIGHTLESS DOME, THE [APOTHEOSIS 1]  HILL, DOUGLAS  P
PROMETHEAN CHALLENGE, THE [@APPLESEED #1]  SHIROW, MASAMUNE  CX
PROMETHEUS UNBOUND [@APPLESEED #2]  SHIROW, MASAMUNE  CX
SCALES OF PROMETHEUS, THE [@APPLESEED #3]  SHIROW, MASAMUNE  CX
PROMETHEAN BALANCE, THE [@APPLESEED #4]  SHIROW, MASAMUNE  CX
DOUBLE EXPOSURE [APPRENTICE ADEPT 1,2,3]  ANTHONY, PIERS  H:2
SPLIT INFINITY [APPRENTICE ADEPT #1]  ANTHONY, PIERS  C/P:2,@P
BLUE ADEPT [APPRENTICE ADEPT #2]  ANTHONY, PIERS  C/P,@P,H

TITLE  AUTHOR  CODE
JUXTAPOSITION [APPRENTICE ADEPT #3]  ANTHONY, PIERS  C/P,@P,H
OUT OF PHAZE [APPRENTICE ADEPT #4]  ANTHONY, PIERS  C/P:2,@P
ROBOT ADEPT [APPRENTICE ADEPT #5]  ANTHONY, PIERS  C/P:2,@P,SFWA-P
UNICORN POINT [APPRENTICE ADEPT #6]  ANTHONY, PIERS  C/P,SFWA-P,@P,H
PHAZE DOUBT [APPRENTICE ADEPT #7]  ANTHONY, PIERS  C/P,@P,SFWA-P:2,H
WHIRLWIND BENEATH THE SEA [AQUANAUTS 6]  STANTON, KEN  P
HERESY [AQUASILVA 1]  AUDLEY, ANSELM  H
INQUISITION [@AQUASILVA #2]  AUDLEY, ANSELM  P
AQUILA IN THE NEW WORLD [@AQUILIAD #1]  SOMTOW, S. P.  C/P,P
AQUILA AND THE IRON HORSE [@AQUILIAD #2]  SOMTOW, S. P.  P
AQUILA AND THE SPHINX [@AQUILIAD #3]  SOMTOW, S. P.  P
ARAFEL’S SAGA [ARAFEL 1,2]  CHERRYH, C. J.  H:2
DREAMSTONE, THE [ARAFEL 1]  CHERRYH, C. J.  C/P:2,P
TREE OF SWORDS AND JEWELS, THE [ARAFEL 2]  CHERRYH, C. J.  C/P:3,P
DREAMING TREE, THE [ARAFEL (REVISED) 1,2]  CHERRYH, C. J.  P
LEGACY OF MOREVI [@ARATHELLEAN WARS #1]  MORRIS, TEE  LP
GLAICE [ARBITER 1]  DOUGLAS, L. WARREN  P
STEPWATER [ARBITER]  DOUGLAS, L. WARREN  P
WHEELS OF PHYRE, THE [ARBITER 2]  DOUGLAS, L. WARREN  P
DEVIL ON MY BACK [ARC ONE 1]  HUGHES, MONICA  P
DREAM CATCHER, THE [ARC ONE 2]  HUGHES, MONICA  SFWA
ARC RIDERS [ARC RIDERS]  DRAKE, DAVID/MORRIS, JANET  C/P:2,P
FOURTH ROME, THE [ARC RIDERS]  DRAKE, DAVID/MORRIS, JANET  C/P,P
ALLAMANDA [ARCADY 2]  WILLIAMS, MICHAEL  H,LP
ARCADY [ARCADY 1]  WILLIAMS, MICHAEL  C/P,P
SILVERHAND [@ARCANA #1]  LLYWELYN, MORGAN/SCOTT, MICHAEL  P
SILVERLIGHT [@ARCANA #2]  LLYWELYN, MORGAN/SCOTT, MICHAEL  P
WHITE LIES [ARCANE SOCIETY]  KRENTZ, JAYNE ANN  H
APOCALYPSE ARRAY [ARCHANGEL 4]  MOREHOUSE, LYDA  P
ARCHANGEL PROTOCOL [ARCHANGEL 1]  MOREHOUSE, LYDA  P
FALLEN HOST [ARCHANGEL 2]  MOREHOUSE, LYDA  P
MESSIAH NODE [ARCHANGEL 3]  MOREHOUSE, LYDA  P
ARCHFORM: BEAUTY [ARCHFORM BEAUTY 1]  MODESITT, L. E. JR.  H
FLASH [ARCHFORM BEAUTY 2]  MODESITT, L. E. JR.  H
BLACK BRILLION [ARCHONATE 3]  HUGHES, MATTHEW  H
FOOL ME TWICE [ARCHONATE 2]  HUGHES, MATTHEW  C/P,P
FOOLS ERRANT [ARCHONATE 1]  HUGHES, MATTHEW  P
GIST HUNTER, THE [ARCHONATE/HENGHIS HAPTHORN]  HUGHES, MATTHEW  H
GULLIBLE’S TRAVELS [ARCHONATE 1,2]  HUGHES, MATTHEW  H
MAJESTRUM [ARCHONATE/HENGHIS HAPTHORN]  HUGHES, MATTHEW  H
AREA 51 [AREA 51 1]  DOHERTY, ROBERT  P
AREA 51: EXCALIBUR [AREA 51 6]  DOHERTY, ROBERT  P
AREA 51: LEGEND [AREA 51 9]  DOHERTY, ROBERT  C/P,P
AREA 51: NOSFERATU [@AREA 51 8]  DOHERTY, ROBERT  P
AREA 51: NOSFERATU (PROOF) [AREA 51 8]  DOHERTY, ROBERT  SR-P
AREA 51: THE GRAIL [AREA 51 5]  DOHERTY, ROBERT  P
AREA 51: THE MISSION [AREA 51 3]  DOHERTY, ROBERT  P
AREA 51: THE REPLY [AREA 51 2]  DOHERTY, ROBERT  P
AREA 51: THE SPHINX [AREA 51 4]  DOHERTY, ROBERT  P
AREA 51: THE TRUTH [@AREA 51 7]  DOHERTY, ROBERT  P
FALL OF WORLDS, THE [@AREIA DARENGA #1]  MEZO, FRANCINE  C/P:2,P
UNLESS SHE BURN [@AREIA DARENGA #2]  MEZO, FRANCINE  P
NO EARTHLY SHORE [@AREIA DARENGA #3]  MEZO, FRANCINE  C/P,P
BRONWYN’S BANE [ARGONIA 3]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
CHRISTENING QUEST, THE [ARGONIA 4]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
SONG OF SORCERY [ARGONIA 1]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
SONGS FROM THE SEASHELL ARCHIVES [ARGONIA 1,2]  SCARBOROUGH, ELIZABETH ANN  P
UNICORN CREED, THE [ARGONIA 2]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
BLOOD MONEY [@ARHAMMER BRUNNER 1]  WERNER, C. L.  P
ARMAGEDDON: THE MUSICAL [ARMAGEDDON 1]  RANKIN, ROBERT  LP
ARMAGEDDON: THE MUSICAL (PROOF) [ARMAGEDDON 1]  RANKIN, ROBERT  SR-LP
SUBURBAN BOOK OF THE DEAD, THE [ARMAGEDDON 3]  RANKIN, ROBERT  P
ANCIENT ENEMY, THE [ARNA 1]  ROWLEY, CHRISTOPHER  P
SHAST WAR, THE [@ARNA 2]  ROWLEY, CHRISTOPHER  P
DOOM’S BREAK [@ARNA #3]  ROWLEY, CHRISTOPHER  P
SIEGE OF ARRANDIN, THE [@ARRANDIN #1]  HERNIMAN, MARCUS  P
TREASON OF DORTREAN, THE [@ARRANDIN #2]  HERNIMAN, MARCUS  P
CRIMSON GODDESS, THE [ARRIS 3]  COOKE, CATHERINE  C/P:2,P
REALM OF THE GODS [ARRIS 2]  COOKE, CATHERINE  P
WINGED ASSASSIN, THE [ARRIS 1]  COOKE, CATHERINE  P
EVERVILLE [ART 2]  BARKER, CLIVE  H,LP
GREAT AND SECRET SHOW, THE [ART 1]  BARKER, CLIVE  H,P
ARTEMIS FOWL [ARTEMIS FOWL 1]  COLFER, EOIN  C/P,LP

The MITSFS Seriesdex  September 24, 2007  Page 6
TITLE  AUTHOR  CODE
ARTEMIS FOWL: THE ARCTIC INCIDENT [ARTEMIS FOWL 2]  COLFER, EOIN  LP
ARTEMIS FOWL: THE ETERNITY CODE [ARTEMIS FOWL 3]  COLFER, EOIN  H
ARTEMIS FOWL: THE LOST COLONY [ARTEMIS FOWL 5]  COLFER, EOIN  H
ARTEMIS FOWL: THE OPAL DECEPTION [ARTEMIS FOWL 4]  COLFER, EOIN  H
KNIGHT LIFE [ARTHUR PENN 1]  DAVID, PETER  P,SFWA-P
ONE KNIGHT ONLY [ARTHUR PENN 2]  DAVID, PETER  P
ARTHUR WAR LORD [ARTHUR WAR LORD 1]  AB HUGH, DAFYDD  P
FAR BEYOND THE WAVE [@ARTHUR WAR LORD #2]  AB HUGH, DAFYDD  P
CRYSTAL CAVE, THE [ARTHURIAN 1]  STEWART, MARY  C/P:2,P,LP
HOLLOW HILLS, THE [ARTHURIAN 2]  STEWART, MARY  C/P,P,LP,H
LAST ENCHANTMENT, THE [ARTHURIAN 3]  STEWART, MARY  H:2,LP,P
WICKED DAY, THE [ARTHURIAN 4]  STEWART, MARY  H:2,LP,P
ISLAND SNATCHERS, THE [ARWIN 5]  SMITH, GEORGE H.  C/P,P
KAR KABALLA [ARWIN 3]  SMITH, GEORGE H.  D42900
SECOND WAR OF THE WORLDS, THE [ARWIN 4]  SMITH, GEORGE H.  P
WITCH QUEEN OF LOCHLANN [ARWIN 2]  SMITH, GEORGE H.  C/P,P
AS IT IS ON MARS (REVISED SECOND EDITION) [AS IT IS ON MARS 1]  CRONIN, THOMAS W.  H
GIVE US THIS MARS [AS IT IS ON MARS 2]  CRONIN, THOMAS W.  H
GLORY BE TO MARS [AS IT IS ON MARS 3]  CRONIN, THOMAS W.  H
CRIMSON SWORD, THE [ASAHIEL 1]  THOMPSON, ELDON  P
OBSIDIAN KEY, THE [ASAHIEL 2]  THOMPSON, ELDON  H
INVADERS FROM THE CENTRE [ASGARD 2]  STABLEFORD, BRIAN M.  P
JOURNEY TO THE CENTER [ASGARD 1]  STABLEFORD, BRIAN M.  H,P
SECRET HISTORY, A [@ASH #1]  GENTLE, MARY  C/P:2,P
CARTHAGE ASCENDANT [@ASH #2]  GENTLE, MARY  C/P,P
WILD MACHINES, THE [@ASH #3]  GENTLE, MARY  C/P,P
LOST BURGUNDY [@ASH #4]  GENTLE, MARY  C/P,P
ASH STAFF, THE [ASH STAFF 1]  FISHER, PAUL R.  C/P,P
HAWKS OF FELLHEATH, THE [ASH STAFF 2]  FISHER, PAUL R.  C/P,P
PRINCESS AND THE THORN, THE [ASH STAFF 3]  FISHER, PAUL R.  P
ALONE IN THE ASHES [ASHES 5]  JOHNSTONE, WILLIAM W.  P
ANARCHY IN THE ASHES [ASHES 3]  JOHNSTONE, WILLIAM W.  P
BLOOD IN THE ASHES [ASHES 4]  JOHNSTONE, WILLIAM W.  P
DANGER IN THE ASHES [ASHES 8]  JOHNSTONE, WILLIAM W.  P
FIRE IN THE ASHES [ASHES 2]  JOHNSTONE, WILLIAM W.  P
OUT OF THE ASHES [ASHES 1]  JOHNSTONE, WILLIAM W.  C/P,P
SMOKE FROM THE ASHES [ASHES 7]  JOHNSTONE, WILLIAM W.  P
TRAPPED IN THE ASHES [ASHES 9]  JOHNSTONE, WILLIAM W.  P
WIND IN THE ASHES [ASHES 6]  JOHNSTONE, WILLIAM W.  C/P,P
DRINK THE FIRE FROM THE FLAMES [ASHLU]  BAKER, SCOTT  P,SFWA-P
FIREDANCE [ASHLU]  BAKER, SCOTT  C/P,P
TIME TO TIME [ASHTON FORD 6]  PENDLETON, DON  P
AURORA [ASIMOV’S ROBOT MYSTERY]  TIEDEMANN, MARK W.  P
CHIMERA [@ASIMOV’S ROBOT MYSTERY]  TIEDEMANN, MARK W.  LP
HAVE ROBOT, WILL TRAVEL [@ASIMOV’S ROBOT MYSTERY]  IRVINE, ALEXANDER C.  H
MIRAGE [@ASIMOV’S ROBOT MYSTERY]  TIEDEMANN, MARK W.  LP
ASTEROID WARS, THE [ASTEROID WARS 1,2,3]  BOVA, BEN  H
PRECIPICE, THE [ASTEROID WARS 1]  BOVA, BEN  H,P
ROCK RATS, THE [ASTEROID WARS 2]  BOVA, BEN  P
SILENT WAR, THE [ASTEROID WARS 3]  BOVA, BEN  P
HEART OF STONE [@ASTROLOGER 1]  DEMARTINO, DENNY  P
WAYWARD MOON [@ASTROLOGER 2]  DEMARTINO, DENNY  P
DEFENDER [ATEVI 5]  CHERRYH, C. J.  H:2,P
DELIVERER [ATEVI 9]  CHERRYH, C. J.  H
DESTROYER [ATEVI 7]  CHERRYH, C. J.  H:2
EXPLORER [ATEVI 6]  CHERRYH, C. J.  H:2
FOREIGNER [ATEVI 1]  CHERRYH, C. J.  H:2,P
INHERITOR [ATEVI 3]  CHERRYH, C. J.  H:2,P
INVADER [ATEVI 2]  CHERRYH, C. J.  H:2
PRECURSOR [ATEVI 4]  CHERRYH, C. J.  H:2,P
PRETENDER [ATEVI 8]  CHERRYH, C. J.  H
CHANGER [ATHANOR 1]  LINDSKOLD, JANE  P
LEGENDS WALKING [ATHANOR 2]  LINDSKOLD, JANE  P
ATHENA, ACT 1 [ATHENA 1]  HSIEH, DEAN  SCX
ATHENA, ACT 2 [ATHENA 2]  HSIEH, DEAN  SCX
ATLAN [ATLAN]  GASKELL, JANE  C/P,P,H
CITY, THE [ATLAN]  GASKELL, JANE  C/P,P,H
DRAGON, THE [ATLAN]  GASKELL, JANE  C/P:2,P
KING’S DAUGHTER [ATLAN]  GASKELL, JANE  C/P,P
SERPENT, THE [ATLAN]  GASKELL, JANE  C/P:2,P
SOME SUMMER LANDS [ATLAN]  GASKELL, JANE  C/P,P,H

SPIDER DESERT [@ATLAN #1]/ WASP MEN ATTACK, THE (SPECIAL) [PERRY RHODAN] ACE PA

The MITSFS Seriesdex  September 24, 2007  Page 7
TITLE  AUTHOR  CODE

FLIGHT FROM TARKIHL [@ATLAN #2]/ MENACE OF ATOMIGEDDON (SPECIAL) [PERRY RHODAN] ACE

PA

PALE COUNTRY PURSUIT [@ATLAN #3]/ ROBOT THREAT: NEW YORK (SPECIAL) [PERRY RHODAN] ACE

PA

CRYSTAL PRINCE, THE [@ATLAN #4]/WAR OF THE GHOSTS [ATLAN 5]  ACE  PA
ASSAULT ON ATLANTIS [ATLANTIS 5]  DONEGAN, GREG  P
ATLANTIS [ATLANTIS 1]  DONEGAN, GREG  C/P,P
ATLANTIS: DEVIL’S SEA [ATLANTIS 3]  DONEGAN, GREG  P
BATTLE FOR ATLANTIS [ATLANTIS 6]  DONEGAN, GREG  P
ATROCITY ARCHIVES, THE [ATROCITY ARCHIVES 1]  STROSS, CHARLES  H
JENNIFER MORGUE, THE [ATROCITY ARCHIVES 2]  STROSS, CHARLES  H:2
ATTAR’S REVENGE [@ATTAR #1]  GRAHAM, ROBERT  C/P,P
WAR OF NERVES [@ATTAR #2]  GRAHAM, ROBERT  P
FIREDANCE [AUBRY KNIGHT]  BARNES, STEVEN  H
GORGON CHILD [AUBRY KNIGHT 2]  BARNES, STEVEN  C/P:2,SFWA-P,P
STREETLETHAL [AUBRY KNIGHT 1]  BARNES, STEVEN  C/P:2,SFWA-P,P
MOUNTAIN’S CALL, THE [AURELIAN EMPIRE 1]  BRENNAN, CAITLIN  LP
SHATTERED DANCE [AURELIAN EMPIRE 3]  BRENNAN, CAITLIN  LP
SONG OF UNMAKING [AURELIAN EMPIRE 2]  BRENNAN, CAITLIN  LP
AURIAN [AURIAN 1]  FUREY, MAGGIE  C/P,P
DHIAMMARA [AURIAN 4]  FUREY, MAGGIE  P
HARP OF WINDS [AURIAN 2]  FUREY, MAGGIE  C/P:2,P
SWORD OF FLAME [AURIAN 3]  FUREY, MAGGIE  C/P:2,P
OLD FUNNY STUFF, THE [@AUTHOR’S CHOICE MONTHLY #1]  EFFINGER, GEORGE ALEC  SR-H

UNTHREATENED BY THE MORNING LIGHT [@AUTHOR’S CHOICE MONTHLY #2] WAGNER, KARL EDWARD

SR-H

SKYROCKET STEELE CONQUERS THE UNIVERSE [@AUTHOR’S CHOICE MONTHLY #11] GOULART, RON

SR-H

STORIES BY MAMA LANSDALE’S YOUNGEST BOY [@AUTHOR’S CHOICE MONTHLY #18] LANSDALE, JOE R.

SR-H

SENSITIVE DEPENDENCE ON INITIAL CONDITIONS, A [@AUTHOR’S CHOICE MONTHLY #20] ROBINSON, KIM STANLEY

SR-LP

ALCHEMISTS, THE [AUTUMN WORLD MOSAIC 1]  GRAVEL, GEARY  C/P:2,P
PATHFINDERS, THE [AUTUMN WORLD MOSAIC 2]  GRAVEL, GEARY  P
AVARYAN RESPLENDENT [AVARYAN RISING 4,5,6]  TARR, JUDITH  LP
HALL OF THE MOUNTAIN KING, THE [@AVARYAN RISING #1]  TARR, JUDITH  H,SFWA,SFWA-P,P
LADY OF HAN-GILEN, THE [@AVARYAN RISING #2]  TARR, JUDITH  H,SFWA-P,P
FALL OF PRINCES, A [@AVARYAN RISING #3]  TARR, JUDITH  C/P,SFWA-P,P,H
ARROWS OF THE SUN [@AVARYAN RISING #4]  TARR, JUDITH  H:2,P
SPEAR OF HEAVEN [@AVARYAN RISING #5]  TARR, JUDITH  P
SO THIS IS HOW IT ENDS [@AVATARS #1]  SUTHERLAND, TUI T.  H
JUSTICE INC. [@AVENGER #1]  ROBESON, KENNETH  C/P:2,P
YELLOW HOARD, THE [@AVENGER #2]  ROBESON, KENNETH  C/P:2,P
SKY WALKER, THE [@AVENGER #3]  ROBESON, KENNETH  C/P:2,P
DEVIL’S HORNS, THE [@AVENGER #4]  ROBESON, KENNETH  C/P:2,P
FROSTED DEATH, THE [@AVENGER #5]  ROBESON, KENNETH  C/P:2,P
BLOOD RING, THE [@AVENGER #6]  ROBESON, KENNETH  C/P:2,P
STOCKHOLDERS IN DEATH [@AVENGER #7]  ROBESON, KENNETH  C/P:2,P
GLASS MOUNTAIN, THE [@AVENGER #8]  ROBESON, KENNETH  C/P:2,P
TUNED FOR MURDER [@AVENGER #9]  ROBESON, KENNETH  C/P:2,P
SMILING DOGS, THE [@AVENGER #10]  ROBESON, KENNETH  C/P:2,P
RIVER OF ICE [@AVENGER #11]  ROBESON, KENNETH  C/P:2,P
FLAME BREATHERS, THE [@AVENGER #12]  ROBESON, KENNETH  C/P:2,P
MURDER ON WHEELS [@AVENGER #13]  ROBESON, KENNETH  C/P:2,P
THREE GOLD CROWNS [@AVENGER #14]  ROBESON, KENNETH  C/P:2,P
HOUSE OF DEATH [@AVENGER #15]  ROBESON, KENNETH  C/P,P
HATE MASTER, THE [@AVENGER #16]  ROBESON, KENNETH  C/P:2,P
NEVLO [@AVENGER #17]  ROBESON, KENNETH  C/P,P
DEATH IN SLOW MOTION [@AVENGER #18]  ROBESON, KENNETH  C/P:2,P
PICTURES OF DEATH [@AVENGER #19]  ROBESON, KENNETH  C/P:2,P
GREEN KILLER, THE [@AVENGER #20]  ROBESON, KENNETH  C/P,P
HAPPY KILLERS, THE [@AVENGER #21]  ROBESON, KENNETH  C/P:2,P
BLACK DEATH, THE [@AVENGER #22]  ROBESON, KENNETH  C/P,P
WILDER CURSE, THE [@AVENGER #23]  ROBESON, KENNETH  C/P,P
MIDNIGHT MURDER [@AVENGER #24]  ROBESON, KENNETH  C/P:2,P
MAN FROM ATLANTIS, THE [@AVENGER #25]  ROBESON, KENNETH  C/P:2,P
RED MOON [@AVENGER #26]  ROBESON, KENNETH  C/P:2,P
PURPLE ZOMBIE, THE [@AVENGER #27]  ROBESON, KENNETH  C/P:2,P
DOCTOR TIME [@AVENGER #28]  ROBESON, KENNETH  C/P,P
NIGHTWITCH DEVIL, THE [@AVENGER #29]  ROBESON, KENNETH  C/P:2,P
BLACK CHARIOTS [@AVENGER #30]  ROBESON, KENNETH  C/P:2,P
CARTOON CRIMES, THE [@AVENGER #31]  ROBESON, KENNETH  C/P,P
DEATH MACHINE, THE [@AVENGER #32]  ROBESON, KENNETH  P

The MITSFS Seriesdex  September 24, 2007  Page 8
TITLE  AUTHOR  CODE
BLOOD COUNTESS, THE [@AVENGER #33]  ROBESON, KENNETH  C/P:2,P
GLASS MAN, THE [@AVENGER #34]  ROBESON, KENNETH  C/P,P
IRON SKULL, THE [@AVENGER #35]  ROBESON, KENNETH  P
DEMON ISLAND [@AVENGER #36]  ROBESON, KENNETH  C/P:2,P
DROWNED QUEEN, THE [AVENGERS 6]  LAUMER, KEITH  P
FLOATING GAME, THE [@AVENGERS 1]  GARFORTH, JOHN  C/P:2,P
GOLD BOMB, THE [AVENGERS 7]  LAUMER, KEITH  C/P,P
LAUGH WAS ON LAZARUS, THE [@AVENGERS 2]  GARFORTH, JOHN  P
PASSING OF GLORIA MUNDAY, THE [@AVENGERS 3]  GARFORTH, JOHN  C/P,P
MAGNETIC MAN, THE [AVENGERS #8]  DANIELS, NORMAN  C/P:2,P
MOON EXPRESS [AVENGERS #9]  DANIELS, NORMAN  C/P,P

AVENGERS AND THE THUNDERBOLTS, THE [AVENGERS (MARVEL COMICS CHARACTERS)] ASKEGREN, PIERCE

C/P,P

AVENGERS BATTLE THE EARTH-WRECKER, THE [AVENGERS (MARVEL COMICS CHARACTERS)] BINDER, OTTO

C/P,P

MAN WHO STOLE TOMORROW, THE [AVENGERS (MARVEL COMICS CHARACTERS)] MICHELINIE, DAVID

C/P,P

HUNTER’S DEATH [AVERALAAN 2]  WEST, MICHELLE  C/P,P
HUNTER’S OATH [AVERALAAN 1]  WEST, MICHELLE  C/P,P
CRYSTAL CROWN, THE [AVERIDAN 1]  CLOUGH, BRENDA W.  C/P:2,P
DRAGON OF MISHBIL, THE [AVERIDAN 4]  CLOUGH, BRENDA W.  C/P:2,P
NAME OF THE SUN, THE [AVERIDAN 3]  CLOUGH, BRENDA W.  C/P:2,P
REALM BENEATH, THE [AVERIDAN 2]  CLOUGH, BRENDA W.  C/P:2,P
KING BENEATH THE MOUNTAIN, A [AWAKENED KING 2]  CHARRETTE, ROBERT N.  P
KNIGHT AMONG KNAVES, A [AWAKENED KING 3]  CHARRETTE, ROBERT N.  P
AWAKENERS, THE [AWAKENERS 1,2]  TEPPER, SHERI S.  H:2,LP
NORTHSHORE [@AWAKENERS #1]  TEPPER, SHERI S.  C/P,SFWA,SFWA-P,P
SOUTHSHORE [@AWAKENERS #2]  TEPPER, SHERI S.  C/P:2,SFWA-P,P,H
DESIGNATED TARGETS [AXIS OF TIME 2]  BIRMINGHAM, JOHN  LP
FINAL IMPACT [AXIS OF TIME 3]  BIRMINGHAM, JOHN  LP
WEAPONS OF CHOICE [AXIS OF TIME 1]  BIRMINGHAM, JOHN  LP
AYESHA: THE RETURN OF SHE [AYESHA 2]  HAGGARD, H. RIDER  C/P,P
SHE [AYESHA 1]  HAGGARD, H. RIDER  H:2,P
SHE AND ALLAN [AYESHA 4/ALLAN QUATERMAIN]  HAGGARD, H. RIDER  LP
WISDOM’S DAUGHTER [AYESHA 3]  HAGGARD, H. RIDER  P
BRING ME THE HEAD OF PRINCE CHARMING [AZZIE 1]  ZELAZNY, ROGER/SHECKLEY, ROBERT  H:2,LP,P
FARCE TO BE RECKONED WITH, A [AZZIE 3]  ZELAZNY, ROGER/SHECKLEY, ROBERT  H,LP
IF AT FAUST YOU DON’T SUCCEED [AZZIE 2]  ZELAZNY, ROGER/SHECKLEY, ROBERT  P,LP
IF AT FAUST YOU DON’T SUCCEED (PROOF) [AZZIE 2]  ZELAZNY, ROGER/SHECKLEY, ROBERT  SR-LP

The MITSFS Seriesdex  September 24, 2007  Page 9
TITLE  AUTHOR  CODE
CALL TO ARMS, A [@BABYLON 5]  SHECKLEY, ROBERT  C/P,P
IN THE BEGINNING [@BABYLON 5]  DAVID, PETER  C/P,P
THIRDSPACE [@BABYLON 5]  DAVID, PETER  P
VOICES [@BABYLON 5 #1]  VORNHOLT, JOHN  C/P:2,P
ACCUSATIONS [@BABYLON 5 #2]  TILTON, LOIS  C/P,P
BLOOD OATH [@BABYLON 5 #3]  VORNHOLT, JOHN  C/P,P
CLARK’S LAW [@BABYLON 5 #4]  MORTIMORE, JIM  C/P:2,P

TOUCH OF YOUR SHADOW, THE WHISPER OF YOUR NAME, THE [@BABYLON 5 #5] BARRETT, NEAL JR.

P

BETRAYALS [@BABYLON 5 #6]  STIRLING, S. M.  C/P,P
SHADOW WITHIN, THE [@BABYLON 5 #7]  CAVELOS, JEANNE  C/P,P
PERSONAL AGENDAS [@BABYLON 5 #8]  SARRANTONIO, AL  C/P,P
TO DREAM IN THE CITY OF SORROWS [@BABYLON 5 #9]  DRENNAN, KATHRYN M.  C/P,P

LONG NIGHT OF CENTAURI PRIME, THE [@BABYLON 5 LEGIONS OF FIRE #1] DAVID, PETER

P

ARMIES OF LIGHT AND DARK [@BABYLON 5 LEGIONS OF FIRE #2]  DAVID, PETER  C/P,P
OUT OF THE DARKNESS [@BABYLON 5 LEGIONS OF FIRE #3]  DAVID, PETER  C/P,P
CASTING SHADOWS [@BABYLON 5 PASSING OF THE TECHNO-MAGES #1]  CAVELOS, JEANNE  C/P,P
SUMMONING LIGHT [@BABYLON 5 PASSING OF THE TECHNO-MAGES #2]  CAVELOS, JEANNE  C/P,P
INVOKING DARKNESS [@BABYLON 5 PASSING OF THE TECHNO-MAGES #3]  CAVELOS, JEANNE  C/P,P
DARK GENESIS [@BABYLON 5 PSI CORPS 1]  KEYES, J. GREGORY  C/P,P
DEADLY RELATIONS [@BABYLON 5 PSI CORPS 2]  KEYES, J. GREGORY  P
FINAL RECKONING [@BABYLON 5 PSI CORPS 3]  KEYES, J. GREGORY  P
BACK TO THE FUTURE [BACK TO THE FUTURE 1]  GIPE, GEORGE  C/P,P
BACK TO THE FUTURE PART 2 [BACK TO THE FUTURE 2]  GARDNER, CRAIG SHAW  P
BACK TO THE FUTURE PART 3 [BACK TO THE FUTURE 3]  GARDNER, CRAIG SHAW  P
MEMOIRS OF SCHLOCK HOMES, THE [BAGEL STREET DOZEN 2]  FISH, ROBERT L.  H
BLOOD STONES, THE [BALZAN 1]  MOORE, WALLACE  P
CAVES OF MADNESS, THE [BALZAN 2]  MOORE, WALLACE  P
LIGHTS OF ZETAR, THE [BALZAN 3]  MOORE, WALLACE  P
DRAGON’S ASCENSION, A [BAND OF FOUR]  GREENWOOD, ED  C/P,P
DRAGON’S DOOM [BAND OF FOUR]  GREENWOOD, ED  H
KINGLESS LAND, THE [BAND OF FOUR]  GREENWOOD, ED  P
VACANT THRONE, THE [BAND OF FOUR]  GREENWOOD, ED  H,P
WIT’CH FIRE [BANNED AND THE BANISHED 1]  CLEMENS, JAMES  C/P:2,LP
WIT’CH GATE [BANNED AND THE BANISHED 4]  CLEMENS, JAMES  P
WIT’CH STAR [BANNED AND THE BANISHED 5]  CLEMENS, JAMES  P
WIT’CH STORM [BANNED AND THE BANISHED 2]  CLEMENS, JAMES  C/P,LP
WIT’CH WAR [BANNED AND THE BANISHED 3]  CLEMENS, JAMES  H,LP
CONAN AND THE SPIDER GOD [BANTAM CONAN 5]  DE CAMP, L. SPRAGUE  C/P,P
CONAN THE LIBERATOR [BANTAM CONAN 2]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN THE REBEL [BANTAM CONAN 6]  ANDERSON, POUL  C/P,P

CONAN THE SWORDSMAN [BANTAM CONAN 1] DE CAMP, L. SPRAGUE/CARTER, LIN/

NYBERG, BJORN

C/P,P

ROAD OF KINGS, THE [BANTAM CONAN 4]  WAGNER, KARL EDWARD  P
SWORD OF SKELOS, THE [BANTAM CONAN 3]  OFFUTT, ANDREW J.  C/P,P
BEASTS OF BARAKHAI [@BARAKHAI #1]  REICHERT, MICKEY ZUCKER  H,P
LOST DRAGONS OF BARAKHAI, THE [@BARAKHAI #2]  REICHERT, MICKEY ZUCKER  H,P
BARD [@BARD #1]  TAYLOR, KEITH  C/P:2,P
BARD II [@BARD #2]  TAYLOR, KEITH  C/P,P
WILD SEA, THE [@BARD #3]  TAYLOR, KEITH  P
RAVENS’ GATHERING [@BARD #4]  TAYLOR, KEITH  P
CAST OF CORBIES, A [BARDIC CHOICES]  LACKEY, MERCEDES/SHERMAN, JOSEPHA  C/P:2,P
FOUR AND TWENTY BLACKBIRDS [BARDIC VOICES 4]  LACKEY, MERCEDES  H:2
LARK AND THE WREN, THE [BARDIC VOICES 1]  LACKEY, MERCEDES  C/P:2,P,H
LARK AND THE WREN, THE (EXCERPT) [BARDIC VOICES 1]  LACKEY, MERCEDES  SR-LP
ROBIN AND THE KESTREL, THE [BARDIC VOICES 2]  LACKEY, MERCEDES  C/P,@H,P
EAGLE & THE NIGHTINGALES, THE [@BARDIC VOICES #3]  LACKEY, MERCEDES  H,P
CASTLE OF DECEPTION [BARD’S TALE 1]  LACKEY, MERCEDES/SHERMAN, JOSEPHA  C/P:2,SFWA-P,P
CHAOS GATE, THE [BARD’S TALE 5]  SHERMAN, JOSEPHA  P
CURSE OF THE BLACK HERON [BARD’S TALE 8]  LISLE, HOLLY  C/P:2,P
ESCAPE FROM ROKSAMUR [BARD’S TALE 4]  SHEPHERD, MARK  P
FORTRESS OF FROST AND FIRE [BARD’S TALE 2]  LACKEY, MERCEDES/EMERSON, RU  C/P:2,SFWA-P,P
PRISON OF SOULS [BARD’S TALE 3]  LACKEY, MERCEDES/SHEPHERD, MARK  C/P,SFWA-P,P
THUNDER OF THE CAPTAINS [BARD’S TALE 6]  LISLE, HOLLY/ALLSTON, AARON  P
WRATH OF THE PRINCES [BARD’S TALE 7]  LISLE, HOLLY/ALLSTON, AARON  P
SNAKE OIL WARS, THE [BARION AND COYUL 2]  GODWIN, PARKE  H
WAITING FOR THE GALACTIC BUS [BARION AND COYUL 1]  GODWIN, PARKE  H:2,P
WANTED: A FAIRY GODFATHER [@BARNABY #1]  JOHNSON, CROCKETT  SCX
MR. O’MALLEY AND THE HAUNTED HOUSE [@BARNABY #2]  JOHNSON, CROCKETT  SCX
JACKEEN J. O’MALLEY FOR CONGRESS [@BARNABY #3]  JOHNSON, CROCKETT  SCX
MR. O’MALLEY GOES FOR THE GOLD [@BARNABY #4]  JOHNSON, CROCKETT  SCX
MR. O’MALLEY, WIZARD OF WALL STREET [@BARNABY #5]  JOHNSON, CROCKETT  SCX

The MITSFS Seriesdex  September 24, 2007  Page 10
TITLE  AUTHOR  CODE
OPERATION DOOMSDAY [BARONESS 5]  KENYON, PAUL  P
CONFUSION, THE [BAROQUE CYCLE 2]  STEPHENSON, NEAL  H,LP
QUICKSILVER [BAROQUE CYCLE 1]  STEPHENSON, NEAL  H:3,LP
SYSTEM OF THE WORLD, THE [BAROQUE CYCLE 3]  STEPHENSON, NEAL  H
KING OF THE VAGABONDS [BAROQUE CYCLE 1B]  STEPHENSON, NEAL  P
IRON DAWN [BARRA THE PICT 1]  STOVER, MATTHEW WOODRING  C/P,P
JERICHO MOON [BARRA THE PICT 2]  STOVER, MATTHEW WOODRING  LP
HEART OF RED IRON [BARRAZAN V 2]  GOTLIEB, PHYLLIS  H
O MASTER CALIBAN! [BARRAZAN V 1]  GOTLIEB, PHYLLIS  C/P,SFWA,P,H
GODS OF MARS, THE [BARSOOM 2]/WARLORD OF MARS, THE [BARSOOM 3] BURROUGHS, EDGAR RICE  H:2
JOHN CARTER OF MARS (ORIGINAL) [BARSOOM 11] BURROUGHS, EDGAR RICE SR-L
LLANA OF GATHOL [BARSOOM 10]/JOHN CARTER OF MARS [BARSOOM 11] BURROUGHS, EDGAR RICE  H:2

MASTERMIND OF MARS, THE [BARSOOM 6]/ FIGHTING MAN OF MARS, A [BARSOOM 7] BURROUGHS, EDGAR RICE

H:2

PRINCESS OF MARS, A [BARSOOM 1]/ FIGHTING MAN OF MARS, A [BARSOOM 7] BURROUGHS, EDGAR RICE

LP

SCIENCE FICTION CLASSICS [BARSOOM 4,5,6/PELLUCIDAR 2,3]  BURROUGHS, EDGAR RICE  H
SWORDS OF MARS [BARSOOM 8]/SYNTHETIC MEN OF MARS [BARSOOM 9] BURROUGHS, EDGAR RICE  H:2

THUVIA, MAID OF MARS [BARSOOM 4]/ CHESSMEN OF MARS, THE [BARSOOM 5] BURROUGHS, EDGAR RICE

H:2

PRINCESS OF MARS, A [BARSOOM #1]  BURROUGHS, EDGAR RICE  H:3,@P,SR-H
GODS OF MARS, THE [BARSOOM #2]  BURROUGHS, EDGAR RICE  C/P,@P,H
WARLORD OF MARS, THE [BARSOOM #3]  BURROUGHS, EDGAR RICE  C/P,@P,H:2
THUVIA, MAID OF MARS [BARSOOM #4]  BURROUGHS, EDGAR RICE  H:2,@P
3 MARTIAN NOVELS [BARSOOM #4,5,6]  BURROUGHS, EDGAR RICE  H:2,LP
CHESSMEN OF MARS, THE [BARSOOM #5]  BURROUGHS, EDGAR RICE  C/P:2,@P
MASTERMIND OF MARS, THE [BARSOOM #6]  BURROUGHS, EDGAR RICE  C/P,@P,H
FIGHTING MAN OF MARS, A [BARSOOM #7]  BURROUGHS, EDGAR RICE  C/P,@P,H
SWORDS OF MARS [BARSOOM #8]  BURROUGHS, EDGAR RICE  C/P:2,@P
SYNTHETIC MEN OF MARS [BARSOOM #9]  BURROUGHS, EDGAR RICE  C/P:2,@P
LLANA OF GATHOL [BARSOOM #10]  BURROUGHS, EDGAR RICE  C/P:2,@P
JOHN CARTER OF MARS [BARSOOM #11]  BURROUGHS, EDGAR RICE  C/P:2,@P
AMULET OF SAMARKAND, THE [@BARTIMAEUS #1]  STROUD, JONATHAN  H
GOLEM’S EYE, THE [@BARTIMAEUS #2]  STROUD, JONATHAN  H
PTOLEMY’S GATE [@BARTIMAEUS #3]  STROUD, JONATHAN  H
EMPEROR OF DAWN [BASIL CASTELLAN 2]  WHITE, STEVE  P
PRINCE OF SUNSET [BASIL CASTELLAN 1]  WHITE, STEVE  C/P,P
ELDRIE THE HEALER [@BASTARD PRINCESS #1]  EDWARDS, CLAUDIA J.  P
BATMAN [BATMAN 1]  GARDNER, CRAIG SHAW  C/P:2,P
BATMAN MURDERS, THE [BATMAN 2]  GARDNER, CRAIG SHAW  P
MASK OF THE PHANTASM [@BATMAN]  GRAVEL, GEARY  P
RED RAIN [BATMAN]  MOENCH, DOUG  CX
DUAL TO THE DEATH [@BATMAN ANIMATED SERIES 2]  GRAVEL, GEARY  P
SHADOWS OF THE PAST [@BATMAN ANIMATED SERIES 1]  GRAVEL, GEARY  P
COLD AS ICE [BATTACHARIYA 1]  SHEFFIELD, CHARLES  C/P:2,P,H
DARK AS DAY [BATTACHARIYA 3]  SHEFFIELD, CHARLES  H
GANYMEDE CLUB, THE [BATTACHARIYA 2]  SHEFFIELD, CHARLES  H,P
BATTLE CIRCLE [BATTLE CIRCLE 1,2,3]  ANTHONY, PIERS  C/P:2,P
SOS THE ROPE [BATTLE CIRCLE 1]  ANTHONY, PIERS  C/P,@P,H
VAR THE STICK [BATTLE CIRCLE #2]  ANTHONY, PIERS  C/P:2,@P
ARMAGEDDON [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/GOLDEN, CHRISTOPHER  P
DESTINY [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/LINAWEAVER, BRAD  P

OFFICIAL BATTLESTAR GALACTICA SCRAPBOOK, THE [BATTLESTAR GALACTICA 1978] NEYLAND, JAMES

R/REF-H

PARADIS [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/LINAWEAVER, BRAD  P
REBELLION [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/RODGERS, ALAN  H
REDEMPTION [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/LINAWEAVER, BRAD  H
SAGA OF A STAR WORLD [@BATTLESTAR GALACTICA 1978]  LARSON, GLEN A./THURSTON, ROBERT  P
WARHAWK [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/GOLDEN, CHRISTOPHER  H
BATTLESTAR GALACTICA [BATTLESTAR GALACTICA 1978 #1]  LARSON, GLEN A./THURSTON, ROBERT  H:2,@P
CYLON DEATH MACHINE, THE [@BATTLESTAR GALACTICA 1978 #2]  LARSON, GLEN A./THURSTON, ROBERT  C/P,P
TOMBS OF KOBOL, THE [@BATTLESTAR GALACTICA 1978 #3]  LARSON, GLEN A./THURSTON, ROBERT  P
YOUNG WARRIORS, THE [@BATTLESTAR GALACTICA 1978 #4]  LARSON, GLEN A./THURSTON, ROBERT  P
GALACTICA DISCOVERS EARTH [@BATTLESTAR GALACTICA 1978 #5]  LARSON, GLEN A./RESNICK, MIKE  P
LIVING LEGEND, THE [@BATTLESTAR GALACTICA 1978 #6]  LARSON, GLEN A./YERMAKOV, NICHOLAS  P
WAR OF THE GODS [@BATTLESTAR GALACTICA 1978 #7]  LARSON, GLEN A./YERMAKOV, NICHOLAS  P
GREETINGS FROM EARTH [@BATTLESTAR GALACTICA 1978 #8]  LARSON, GLEN A./GOULART, RON  P
EXPERIMENT IN TERRA [@BATTLESTAR GALACTICA 1978 #9]  LARSON, GLEN A./GOULART, RON  C/P,P
LONG PATROL, THE [@BATTLESTAR GALACTICA 1978 #10]  LARSON, GLEN A./GOULART, RON  P
NIGHTMARE MACHINE, THE [@BATTLESTAR GALACTICA 1978 #11]  LARSON, GLEN A./THURSTON, ROBERT  P
"DIE, CHAMELEON!" [@BATTLESTAR GALACTICA 1978 #12]  LARSON, GLEN A./THURSTON, ROBERT  P
APOLLO’S WAR [@BATTLESTAR GALACTICA 1978 #13]  LARSON, GLEN A./THURSTON, ROBERT  P,SFWA-P
SURRENDER THE GALACTICA! [@BATTLESTAR GALACTICA 1978 #14]  LARSON, GLEN A./THURSTON, ROBERT  P

TITLE  AUTHOR  CODE
CYLONS’ SECRET, THE [@BATTLESTAR GALACTICA 2003]  GARDNER, CRAIG SHAW  LP
CYLON’S SECRET, THE (PROOF) [@BATTLESTAR GALACTICA 2003]  GARDNER, CRAIG SHAW  SR-LP
SAGITTARIUS IS BLEEDING [@BATTLESTAR GALACTICA 2003]  DAVID, PETER  LP
UNITY [@BATTLESTAR GALACTICA 2003]  HARPER, STEVEN  LP
BATTLESTATION [@BATTLESTATION #1]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P:2
VANGUARD [@BATTLESTATION #2]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P:2
ASSUMPTION OF RISK [@BATTLETECH 12]  STACKPOLE, MICHAEL A.  C/P,P
BINDING FORCE [@BATTLETECH 32]  COLEMAN, LOREN L.  P
BLACK DRAGON [@BATTLETECH 29]  MILAN, VICTOR  C/P,P
BLOOD OF HEROES [@BATTLETECH 11]  KEITH, ANDREW  C/P,P
BRED FOR WAR [@BATTLETECH 16]  STACKPOLE, MICHAEL A.  P
CALL OF DUTY [BATTLETECH]  PARDOE, BLAINE LEE  P
CLOSE QUARTERS [@BATTLETECH 15]  MILAN, VICTOR  P
DAGGER POINT [@BATTLETECH]  GRESSMAN, THOMAS S.  P
DECISION AT THUNDER RIFT [@BATTLETECH 6]  KEITH, WILLIAM H. JR.  P
DOUBLE-BLIND [@BATTLETECH 31]  COLEMAN, LOREN L.  P
D.R.T. [@BATTLETECH 14]  LONG, JAMES D.  P
ENDGAME [@BATTLETECH]  COLEMAN, LOREN L.  P
FAR COUNTRY [@BATTLETECH 13]  RICE, PETER L.  C/P,P
FLASHPOINT [@BATTLETECH 30]  COLEMAN, LOREN L.  P
HEARTS OF CHAOS [@BATTLETECH 26]  MILAN, VICTOR  P
HEIR TO THE DRAGON [@BATTLETECH 28]  CHARRETTE, ROBERT N.  P
HIGHLANDER GAMBIT [@BATTLETECH 18]  PARDOE, BLAINE LEE  P
I AM JADE FALCON [@BATTLETECH 17]  THURSTON, ROBERT  P
IDEAL WAR [@BATTLETECH 9]  KUBASIK, CHRIS  C/P,P
ILLUSIONS OF VICTORY [@BATTLETECH 47]  COLEMAN, LOREN L.  P
IMPETUS OF WAR [@BATTLETECH 30]  PARDOE, BLAINE LEE  P
MAIN EVENT [@BATTLETECH 10]  LONG, JIM  C/P,P
MALICIOUS INTENT [@BATTLETECH 24]  STACKPOLE, MICHAEL A.  P
MEASURE OF A HERO [@BATTLETECH 48]  PARDOE, BLAINE LEE  P
MERCENARY’S STAR [@BATTLETECH 7]  KEITH, WILLIAM H. JR.  C/P:2,P
OPERATION AUDACITY [BATTLETECH]  PARDOE, BLAINE LEE  P
OPERATION EXCALIBUR [@BATTLETECH 27]  KEITH, WILLIAM H. JR.  P
PATH OF GLORY [@BATTLETECH 49]  BILLS, RANDALL N.  P
PATRIOTS AND TYRANTS [@BATTLETECH 52]  COLEMAN, LOREN L.  P
PRICE OF GLORY, THE [@BATTLETECH 8]  KEITH, WILLIAM H. JR.  C/P:2,P
STAR LORD [@BATTLETECH 23]  PHILLIPS, DONALD G.  P
STORMS OF FATE [@BATTLETECH]  COLEMAN, LOREN L.  C/P,P
SWORD AND THE DAGGER, THE [@BATTLETECH]  MAYHAR, ARDATH  P
TACTICS OF DUTY [@BATTLETECH 19]  KEITH, WILLIAM H. JR.  P
TEST OF VENGEANCE [@BATTLETECH 51]  NYSTUL, BRYAN  P
WOLF PACK [BATTLETECH 4]  CHARRETTE, ROBERT N.  C/P,P
WOLVES ON THE BORDER [BATTLETECH 25]  CHARRETTE, ROBERT N.  P
WAY OF THE CLANS [@BATTLETECH #1/JADE PHOENIX 1]  THURSTON, ROBERT  C/P:2,SFWA-P:2,P
BLOODNAME [@BATTLETECH #2/JADE PHOENIX 2]  THURSTON, ROBERT  C/P,P
FALCON GUARD [@BATTLETECH #3/JADE PHOENIX 3]  THURSTON, ROBERT  C/P:2,P
NATURAL SELECTION [@BATTLETECH #5]  STACKPOLE, MICHAEL A.  C/P,P
LETHAL HERITAGE [@BATTLETECH #20/BLOOD OF KERENSKY 1]  STACKPOLE, MICHAEL A.  P,SFWA-P
BLOOD LEGACY [@BATTLETECH #21/BLOOD OF KERENSKY 2]  STACKPOLE, MICHAEL A.  P
LOST DESTINY [@BATTLETECH #22/BLOOD OF KERENSKY 3]  STACKPOLE, MICHAEL A.  P
EXODUS ROAD [@BATTLETECH #33/TWILIGHT OF THE CLANS 1]  PARDOE, BLAINE LEE  C/P,P
GRAVE COVENANT [@BATTLETECH #34/TWILIGHT OF THE CLANS 2]  STACKPOLE, MICHAEL A.  C/P,P
HUNTERS, THE [@BATTLETECH #35/TWILIGHT OF THE CLANS 3]  GRESSMAN, THOMAS S.  C/P,P
FREEBIRTH [@BATTLETECH #36/TWILIGHT OF THE CLANS 4]  THURSTON, ROBERT  C/P,P
SWORD AND FIRE [@BATTLETECH #37/TWILIGHT OF THE CLANS 5]  GRESSMAN, THOMAS S.  C/P,P
SHADOWS OF WAR [@BATTLETECH #38/TWILIGHT OF THE CLANS 6]  GRESSMAN, THOMAS S.  P
PRINCE OF HAVOC [@BATTLETECH #39/TWILIGHT OF THE CLANS 7]  STACKPOLE, MICHAEL A.  P
FALCON RISING [@BATTLETECH #40/TWILIGHT OF THE CLANS 8]  THURSTON, ROBERT  P
WARRIOR: EN GARDE [@BATTLETECH #41/WARRIOR 1]  STACKPOLE, MICHAEL A.  C/P,P
WARRIOR: RIPOSTE [@BATTLETECH #42/WARRIOR 2]  STACKPOLE, MICHAEL A.  C/P:2,P
WARRIOR: COUPE [@BATTLETECH #43/WARRIOR 3]  STACKPOLE, MICHAEL A.  C/P:2,P
THREADS OF AMBITION [@BATTLETECH #44/CAPELLAN SOLUTION 1]  COLEMAN, LOREN L.  P
KILLING FIELDS, THE [@BATTLETECH #45/CAPELLAN SOLUTION 2]  COLEMAN, LOREN L.  P
BATTLEDRAGON [BAZIL BROKETAIL 4]  ROWLEY, CHRISTOPHER  P
BAZIL BROKETAIL [BAZIL BROKETAIL 1]  ROWLEY, CHRISTOPHER  C/P,SFWA-P,P
DRAGON AT WORLD’S END, A [BAZIL BROKETAIL 5]  ROWLEY, CHRISTOPHER  P
DRAGON ULTIMATE [BAZIL BROKETAIL 7]  ROWLEY, CHRISTOPHER  P
DRAGONS OF ARGONATH [BAZIL BROKETAIL 6]  ROWLEY, CHRISTOPHER  P
DRAGONS OF WAR [BAZIL BROKETAIL 3]  ROWLEY, CHRISTOPHER  C/P,P
SWORD FOR A DRAGON, A [BAZIL BROKETAIL 2]  ROWLEY, CHRISTOPHER  C/P:2,SFWA-P,P
BEAUTY’S PUNISHMENT [BEAUTY 2]  ROQUELAURE, A. N.  SR-L
BEAUTY’S RELEASE [BEAUTY 3]  ROQUELAURE, A. N.  H,SR-L
CLAIMING OF SLEEPING BEAUTY, THE [BEAUTY 1]  ROQUELAURE, A. N.  H,SR-L

TITLE  AUTHOR  CODE
BEAUTY AND THE BEAST [BEAUTY AND THE BEAST]  HAMBLY, BARBARA  C/P,P
MASQUES [@BEAUTY AND THE BEAST]  EMERSON, RU  P
SONG OF ORPHEUS [BEAUTY AND THE BEAST]  HAMBLY, BARBARA  C/P:3,P
PSALMS OF HEROD, THE [BECCA 1]  FRIESNER, ESTHER M.  P
SWORD OF MARY, THE [BECCA 2]  FRIESNER, ESTHER M.  P
BECOMING ALIEN [BECOMING ALIEN 1]  ORE, REBECCA  H:2,P
BEING ALIEN [BECOMING ALIEN 2]  ORE, REBECCA  H:2,SFWA-P
HUMAN TO HUMAN [BECOMING ALIEN 3]  ORE, REBECCA  H:2,SFWA-P,P
BEGGARS AND CHOOSERS [BEGGARS 2]  KRESS, NANCY  H:2,P
BEGGARS IN SPAIN [BEGGARS 1]  KRESS, NANCY  H:2,P
BEGGARS RIDE [BEGGARS 3]  KRESS, NANCY  C/P:2,P
TERRIER [@BEKA COOPER #1]  PIERCE, TAMORA  H
BELGARATH THE SORCERER [BELGARIAD/MALLOREON]  EDDINGS, DAVID/EDDINGS, LEIGH  H:2,P
BELGARIAD 1, THE [BELGARIAD 1,2,3]  EDDINGS, DAVID  H:2
BELGARIAD 2, THE [BELGARIAD 4,5]  EDDINGS, DAVID  H:2
CASTLE OF WIZARDRY [BELGARIAD 4]  EDDINGS, DAVID  C/P:3,P
ENCHANTERS’ END GAME [BELGARIAD 5]  EDDINGS, DAVID  C/P:2,P
MAGICIAN’S GAMBIT [BELGARIAD 3]  EDDINGS, DAVID  C/P:2,P
PAWN OF PROPHECY [BELGARIAD 1]  EDDINGS, DAVID  C/P:2,P
POLGARA THE SORCERESS [BELGARIAD/MALLOREON]  EDDINGS, DAVID/EDDINGS, LEIGH  H:3
QUEEN OF SORCERY [BELGARIAD 2]  EDDINGS, DAVID  C/P:2,P
RIVAN CODEX, THE [BELGARIAD/MALLOREON]  EDDINGS, DAVID/EDDINGS, LEIGH  P
DANCE OF TIME, THE [BELISARIUS 6]  FLINT, ERIC/DRAKE, DAVID  H
DESTINY’S SHIELD [BELISARIUS 3]  FLINT, ERIC/DRAKE, DAVID  C/P:2,P
FORTUNE’S STROKE [BELISARIUS 4]  FLINT, ERIC/DRAKE, DAVID  H:2
IN THE HEART OF DARKNESS [BELISARIUS 2]  FLINT, ERIC/DRAKE, DAVID  C/P,P
OBLIQUE APPROACH, AN [BELISARIUS 1]  FLINT, ERIC/DRAKE, DAVID  C/P,P
TIDE OF VICTORY, THE [BELISARIUS 5]  FLINT, ERIC/DRAKE, DAVID  H
SWORD OF ORION [@BENEATH STRANGE SKIES #1]  LEE, SHARON/MILLER, STEVE  LP
LAW OF WAR, THE [BENTON HAWKES 2]  SHATNER, WILLIAM  H,P
MAN O’ WAR [BENTON HAWKES 1]  SHATNER, WILLIAM  H,P
BERBORA [@BERBORA #1]  SIROTA, MIKE  C/P,P
FLIGHT FROM BERBORA [@BERBORA #2]  SIROTA, MIKE  P
HOUR OF THE DRAGON, THE [BERKLEY CONAN 1]  HOWARD, ROBERT E.  H,P
PEOPLE OF THE BLACK CIRCLE, THE [BERKLEY CONAN 3]  HOWARD, ROBERT E.  P
RED NAILS [BERKLEY CONAN 2]  HOWARD, ROBERT E.  P
INTERNATIONAL INCIDENTS [BERKLEY SIX MILLION DOLLAR MAN 2]  JAHN, MICHAEL  P
SECRET OF BIGFOOT PASS, THE [BERKLEY SIX MILLION DOLLAR MAN 1]  JAHN, MICHAEL  P
BIG HUNT, THE [BERNICE SUMMERFIELD]  PARKIN, LANCE  H
LIFE DURING WARTIME [BERNICE SUMMERFIELD]  CORNELL, PAUL  HA

LIFE IN PIECES, A [BERNICE SUMMERFIELD] STONE, DAVE/SUTTON, PAUL/

LIDSTER, JOSEPH

H

LIFE OF SURPRISES, A [BERNICE SUMMERFIELD]  CORNELL, PAUL  HA
LIFE WORTH LIVING, A [BERNICE SUMMERFIELD]  GUERRIER, SIMON  HA

PROFESOR BERNICE SUMMERFIELD AND THE DEAD MEN DIARIES [BERNICE SUMMERFIELD] CORNELL, PAUL

PA

PROFESOR BERNICE SUMMERFIELD AND THE GODS OF THE UNDERWORLD [BERNICE SUMMERFIELD 2] COLE, STEPHEN

C/P,P

PROFESSOR BERNICE SUMMERFIELD AND THE DOOMSDAY MANUSCRIPT [BERNICE SUMMERFIELD 1] RICHARDS, JUSTIN

P

PROFESSOR BERNICE SUMMERFIELD AND THE GLASS PRISON [BERNICE SUMMERFIELD 5] RAYNER, JACQUELINE

P

PROFESSOR BERNICE SUMMERFIELD AND THE INFERNAL NEXUS [BERNICE SUMMERFIELD 4] STONE, DAVE

P

PROFESSOR BERNICE SUMMERFIELD AND THE SQUIRE’S CRYSTAL [BERNICE SUMMERFIELD 3] RAYNER, JACQUELINE

P

BERSERKER [BERSERKER 1]  SABERHAGEN, FRED  C/P:2,P

BERSERKER BASE [BERSERKER 2] SABERHAGEN, FRED SFWA,SFWA-P,HA:2,

LPA

BERSERKER: BLUE DEATH [BERSERKER 3]  SABERHAGEN, FRED  C/P:2,SFWA,SFWA-P,LP
BERSERKER FURY [BERSERKER 7]  SABERHAGEN, FRED  H,P
BERSERKER KILL [BERSERKER 6]  SABERHAGEN, FRED  H:2
BERSERKER LIES [BERSERKER]  SABERHAGEN, FRED  C/P,P
BERSERKER MAN [BERSERKER]  SABERHAGEN, FRED  C/P:2,P
BERSERKER PRIME [BERSERKER]  SABERHAGEN, FRED  P
BERSERKER THRONE, THE [BERSERKER 4]  SABERHAGEN, FRED  H:2,SFWA-P,LP
BERSERKER WARS, THE [BERSERKER 1]  SABERHAGEN, FRED  C/P,SFWA-P,P
BERSERKER’S PLANET [BERSERKER 5]  SABERHAGEN, FRED  C/P:2,P
BERSERKER’S STAR [BERSERKER]  SABERHAGEN, FRED  P
BERSERKERS: THE BEGINNING [BERSERKER]  SABERHAGEN, FRED  C/P,P
BERSERKERS: THE ULTIMATE ENEMY [BERSERKER 4]  SABERHAGEN, FRED  C/P,P
BROTHER ASSASSIN [BERSERKER 2]  SABERHAGEN, FRED  C/P:2,P
BROTHER BERSERKER [BERSERKER 2]  SABERHAGEN, FRED  H
ROGUE BERSERKER [BERSERKER]  SABERHAGEN, FRED  H

The MITSFS Seriesdex  September 24, 2007  Page 13
TITLE  AUTHOR  CODE
SHIVA IN STEEL [BERSERKER 8]  SABERHAGEN, FRED  C/P,P,H
ULTIMATE ENEMY, THE [BERSERKER 4]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
BULL CHIEF, THE [@BERSERKER SWIFTAXE 2]  CARLSEN, CHRIS  P
HORNED WARRIOR, THE [@BERSERKER SWIFTAXE 3]  CARLSEN, CHRIS  P
SHADOW OF THE WOLF [@BERSERKER SWIFTAXE 1]  CARLSEN, CHRIS  P
BESTIARY MOUNTAIN [BESTIARY 1]  FORRESTER, JOHN  H,P
SECRET OF THE ROUND BEAST, THE [BESTIARY 2]  FORRESTER, JOHN  P
UNDEAD AND UNAPPRECIATED [BETSY TAYLOR 3]  DAVIDSON, MARYJANICE  H,P
UNDEAD AND UNEASY [BETSY TAYLOR 6]  DAVIDSON, MARYJANICE  H
UNDEAD AND UNEMPLOYED [BETSY TAYLOR 2]  DAVIDSON, MARYJANICE  P
UNDEAD AND UNPOPULAR [BETSY TAYLOR 5]  DAVIDSON, MARYJANICE  H:2
UNDEAD AND UNRETURNABLE [BETSY TAYLOR 4]  DAVIDSON, MARYJANICE  H
CHARMED DEATH, A [@BEWITCHING MYSTERY 2]  ALT, MADELYN  P
TROUBLE WITH MAGIC, THE [@BEWITCHING MYSTERY 1]  ALT, MADELYN  P

BICYCLING THROUGH SPACE AND TIME [BICYCLING THROUGH SPACE AND TIME 1] SIROTA, MIKE

C/P,SFWA-P:2,P

ULTIMATE BIKE PATH, THE [BICYCLING THROUGH SPACE AND TIME 2]  SIROTA, MIKE  C/P,SFWA-P,P
BY CHAOS CURSED [BIFROST GUARDIANS 5]  REICHERT, MICKEY ZUCKER  P
DRAGONRANK MASTER [BIFROST GUARDIANS 3]  REICHERT, MICKEY ZUCKER  C/P,P
GODSLAYER [BIFROST GUARDIANS 1]  REICHERT, MICKEY ZUCKER  C/P:2,P
SHADOW CLIMBER [BIFROST GUARDIANS 2]  REICHERT, MICKEY ZUCKER  C/P,P
SHADOW’S REALM [BIFROST GUARDIANS 4]  REICHERT, MICKEY ZUCKER  P
AARDVARK AFFAIR, THE [@BIG BRAIN #1]  BRANDNER, GARY  C/P:2,P
BEELZEBUB BUSINESS, THE [@BIG BRAIN #2]  BRANDNER, GARY  P
ENERGY ZERO [@BIG BRAIN #3]  BRANDNER, GARY  P
BILL, THE GALACTIC HERO [BILL 0]  HARRISON, HARRY  C/P,P,H
BILL, THE GALACTIC HERO ON THE PLANET OF BOTTLED BRAINS [BILL 2] HARRISON, HARRY/SHECKLEY, ROBERT  C/P:2,P

BILL, THE GALACTIC HERO ON THE PLANET OF TASTELESS PLEASURE [BILL 3] HARRISON, HARRY/BISCHOFF, DAVID F.

C/P:2,P

BILL, THE GALACTIC HERO ON THE PLANET OF TEN THOUSAND BARS [BILL 5] HARRISON, HARRY/BISCHOFF, DAVID F.

C/P,P

BILL, THE GALACTIC HERO ON THE PLANET OF ZOMBIE VAMPIRES [BILL 4] HARRISON, HARRY/HALDEMAN, JACK C. II

C/P:2,P

BILL, THE GALACTIC HERO: THE FINAL INCOHERENT ADVENTURE [BILL 6] HARRISON, HARRY/HARRIS, DAVID  C/P,SFWA-P,P
BILL THE GALACTIC HERO: THE PLANET OF THE ROBOT SLAVES [BILL 1]  HARRISON, HARRY  C/P,P
BEYOND THE SUMMERLAND [BINDING OF THE BLADE 1]  GRAHAM, L. B.  LP
BRINGER OF STORMS [BINDING OF THE BLADE 2]  GRAHAM, L. B.  LP
SHADOW IN THE DEEP [@BINDING OF THE BLADE #3]  GRAHAM, L. B.  LP
FATHER OF DRAGONS [@BINDING OF THE BLADE #4]  GRAHAM, L. B.  LP
IRON MAIDEN, THE [BIO OF A SPACE TYRANT 6]  ANTHONY, PIERS  LP
REFUGEE [@BIO OF A SPACE TYRANT #1]  ANTHONY, PIERS  C/P:2,P
MERCENARY [@BIO OF A SPACE TYRANT #2]  ANTHONY, PIERS  C/P:2,P
POLITICIAN [@BIO OF A SPACE TYRANT #3]  ANTHONY, PIERS  C/P:2,P
EXECUTIVE [@BIO OF A SPACE TYRANT #4]  ANTHONY, PIERS  C/P:2,P
STATESMAN [@BIO OF A SPACE TYRANT #5]  ANTHONY, PIERS  C/P:2,P
AFTERBURN [BIO RESCUE 2]  VIEHL, S. L.  H
BIO RESCUE [BIO RESCUE 1]  VIEHL, S. L.  H
AI WAR, THE [BIOFAB 3]  BERRY, STEPHEN AMES  SFWA-P
BATTLE FOR TERRA TWO, THE [BIOFAB 2]  BERRY, STEPHEN AMES  C/P,P
BIOFAB WAR, THE [BIOFAB 1]  BERRY, STEPHEN AMES  P
FINAL ASSAULT [BIOFAB 4]  BERRY, STEPHEN AMES  P,SFWA-P
EXTRACURRICULAR ACTIVITIES [BIONIC WOMAN 2]  LOTTMAN, EILEEN  C/P:2,P
WELCOME HOME, JAIME [BIONIC WOMAN 1]  LOTTMAN, EILEEN  P
INFINITY PLAGUE, THE [@BIOWARRIORS 1]  VARDEMAN, ROBERT E.  P
CRISIS AT STARLIGHT [@BIOWARRIORS #2]  VARDEMAN, ROBERT E.  C/P,P
SPACE VECTORS [@BIOWARRIORS #3]  VARDEMAN, ROBERT E.  C/P,P
DREAM OF KINSHIP, A [BIRD OF KINSHIP 2]  COWPER, RICHARD  C/P,P
ROAD TO CORLAY, THE [BIRD OF KINSHIP 1]  COWPER, RICHARD  C/P,P,H
TAPESTRY OF TIME, A [BIRD OF KINSHIP 3]  COWPER, RICHARD  P
GREATHEART [BIRTHRIGHT 2]  MCKEONE, DIXIE LEE  P
HAG’S CONTRACT [BIRTHRIGHT 4]  BETANCOURT, JOHN GREGORY  P
IRON THRONE, THE [BIRTHRIGHT 1]  HAWKE, SIMON  P
SPIDER’S TEST, THE [@BIRTHRIGHT 5]  MCKEONE, DIXIE LEE  P
OUTRIDERS [BIRTHRIGHT PROJECT 1]  MACKEL, KATHRYN  LP
TRACKERS [BIRTHRIGHT PROJECT 2]  MACKEL, KATHRYN  LP
BATTLE OF EVERNIGHT, THE [BITTERBYNDE 3]  DART-THORNTON, CECILIA  C/P,P
BATTLE OF EVERNIGHT, THE (PROOF) [BITTERBYNDE 3]  DART-THORNTON, CECILIA  SR-LP
ILL-MADE MUTE, THE [BITTERBYNDE 1]  DART-THORNTON, CECILIA  C/P:2,P
LADY OF THE SORROWS, THE [BITTERBYNDE 2]  DART-THORNTON, CECILIA  C/P,P
HEALING OF CROSSROADS, THE [BJ VAUGHAN 3]  O’DONOHOE, NICK  P
MAGIC AND THE HEALING, THE [BJ VAUGHAN 1]  O’DONOHOE, NICK  P,SFWA-P
UNDER THE HEALING SIGN [BJ VAUGHAN 2]  O’DONOHOE, NICK  P
ANNALS OF THE BLACK COMPANY [BLACK COMPANY 1,2,3]  COOK, GLEN  H:2

The MITSFS Seriesdex  September 24, 2007  Page 14
TITLE  AUTHOR  CODE
BLACK COMPANY, THE [BLACK COMPANY 1]  COOK, GLEN  C/P:2,SFWA-P,P

BLACK COMPANY GOES SOUTH, THE [BLACK COMPANY 4,5/ BLACK COMPANY SOUTH 1,2] COOK, GLEN

H

DREAMS OF STEEL [BLACK COMPANY 5/BLACK COMPANY SOUTH 2]  COOK, GLEN  C/P:2,SFWA-P,P
SHADOW GAMES [BLACK COMPANY 4/BLACK COMPANY SOUTH 1]  COOK, GLEN  C/P,SFWA-P,P
SHADOWS LINGER [BLACK COMPANY 2]  COOK, GLEN  SR-P,C/P:2
SILVER SPIKE, THE [BLACK COMPANY]  COOK, GLEN  C/P:2,SFWA-P,P
WHITE ROSE, THE [BLACK COMPANY 3]  COOK, GLEN  C/P:2,SFWA-P,P

BLACK COMPANY, THE: GLITTERING STONE I [BLACK COMPANY GLITTERING STONE 1,2] COOK, GLEN

H

BLACK COMPANY, THE: GLITTERING STONE II [BLACK COMPANY GLITTERING STONE 3,4] COOK, GLEN

H

BLEAK SEASONS [@BLACK COMPANY GLITTERING STONE 1/ BLACK COMPANY 6] COOK, GLEN

H,P

SHE IS THE DARKNESS [@BLACK COMPANY GLITTERING STONE 2/ BLACK COMPANY 7] COOK, GLEN

H,P

SOLDIERS LIVE [@BLACK COMPANY GLITTERING STONE 4/ BLACK COMPANY 9] COOK, GLEN

H:2

WATER SLEEPS [@BLACK COMPANY GLITTERING STONE 3/ BLACK COMPANY 8] COOK, GLEN

H

EVENT HORIZON [@BLACK HOLE TRAVEL AGENCY #1]  MCKINNEY, JACK  C/P:2,P
ARTIFACT OF THE SYSTEM [@BLACK HOLE TRAVEL AGENCY #2]  MCKINNEY, JACK  C/P:2,P
FREE RADICALS [@BLACK HOLE TRAVEL AGENCY #3]  MCKINNEY, JACK  C/P:2,P
HOSTILE TAKEOVER [@BLACK HOLE TRAVEL AGENCY #4]  MCKINNEY, JACK  C/P:2,P
BLACK JEWELS TRILOGY, THE [BLACK JEWELS 1,2,3]  BISHOP, ANNE  LP
DAUGHTER OF THE BLOOD [BLACK JEWELS 1]  BISHOP, ANNE  C/P,P
HEIR TO THE SHADOWS [BLACK JEWELS 2]  BISHOP, ANNE  C/P,P
INVISIBLE RING, THE [BLACK JEWELS]  BISHOP, ANNE  C/P,P
QUEEN OF THE DARKNESS [BLACK JEWELS 3]  BISHOP, ANNE  C/P,P
HIGH LORD, THE [BLACK MAGICIAN 3]  CANAVAN, TRUDI  C/P,P
MAGICIANS’ GUILD, THE [BLACK MAGICIAN 1]  CANAVAN, TRUDI  C/P,P
NOVICE, THE [BLACK MAGICIAN 2]  CANAVAN, TRUDI  C/P,P
HUNTING GROUND [@BLACK OAK 4]  GRANT, CHARLES L.  P
WINTER KNIGHT [@BLACK OAK 3]  GRANT, CHARLES L.  P
BLACK KING, THE [BLACK THRONE 2]  RUSCH, KRISTINE KATHRYN  P
BLACK QUEEN, THE [BLACK THRONE 1]  RUSCH, KRISTINE KATHRYN  P
WITCH OF THE INDIES, THE [BLACK VULMEA]  SMITH, DAVID C.  P
BACKLASH MISSION, THE [BLACKCOLLAR 2]  ZAHN, TIMOTHY  C/P,P
BLACKCOLLAR, THE [BLACKCOLLAR 1]  ZAHN, TIMOTHY  C/P:2,P
BLACKCOLLAR: THE JUDAS SOLUTION [BLACKCOLLAR 2]  ZAHN, TIMOTHY  H
EYE FOR AN EYE: THE DOLL, AN [@BLACKSTONE #1]  SAUL, JOHN  P
TWIST OF FATE: THE LOCKET [@BLACKSTONE #2]  SAUL, JOHN  P
ASHES TO ASHES: THE DRAGON’S FLAME [@BLACKSTONE #3]  SAUL, JOHN  P
IN THE SHADOW OF EVIL: THE HANDKERCHIEF [@BLACKSTONE #4]  SAUL, JOHN  C/P,P
DAY OF RECKONING: THE STEREOSCOPE [@BLACKSTONE #5]  SAUL, JOHN  C/P,P
ASYLUM [@BLACKSTONE #6]  SAUL, JOHN  P
FORTUNE, THE [@BLACKWATER #5]  MCDOWELL, MICHAEL  P
WAR, THE [@BLACKWATER #6]  MCDOWELL, MICHAEL  P
BRONZE AXE, THE [@BLADE #1]  LORD, JEFFREY  C/P:2,P
JADE WARRIOR, THE [@BLADE #2]  LORD, JEFFREY  C/P,P
JEWEL OF THARN [@BLADE #3]  LORD, JEFFREY  C/P:2,P
SLAVE OF SARMA [@BLADE #4]  LORD, JEFFREY  C/P:2,P
LIBERATOR OF JEDD [@BLADE #5]  LORD, JEFFREY  C/P:2,P
MONSTER OF THE MAZE [@BLADE #6]  LORD, JEFFREY  C/P,P
PEARL OF PATMOS [@BLADE #7]  LORD, JEFFREY  C/P,P
UNDYING WORLD [@BLADE #8]  LORD, JEFFREY  C/P,P
KINGDOM OF ROYTH [@BLADE #9]  LORD, JEFFREY  P
ICE DRAGON [@BLADE #10]  LORD, JEFFREY  C/P,P
DIMENSION OF DREAMS [@BLADE #11]  LORD, JEFFREY  C/P,P
KING OF ZUNGA [@BLADE #12]  LORD, JEFFREY  P
GOLDEN STEED, THE [@BLADE #13]  LORD, JEFFREY  P
TEMPLES OF AYOCAN, THE [@BLADE #14]  LORD, JEFFREY  C/P,P
TOWERS OF MELNON, THE [@BLADE #15]  LORD, JEFFREY  P
CRYSTAL SEAS, THE [@BLADE #16]  LORD, JEFFREY  P
MOUNTAINS OF BREGA, THE [@BLADE #17]  LORD, JEFFREY  P
WARLORDS OF GAIKON [@BLADE #18]  LORD, JEFFREY  P
LOOTERS OF THARN [@BLADE #19]  LORD, JEFFREY  P
GUARDIANS OF THE CORAL THRONE [@BLADE #20]  LORD, JEFFREY  C/P,P
CHAMPION OF THE GODS [@BLADE #21]  LORD, JEFFREY  P
FORESTS OF GLEOR, THE [@BLADE #22]  LORD, JEFFREY  P
EMPIRE OF BLOOD [@BLADE #23]  LORD, JEFFREY  P
DRAGONS OF ENGLOR, THE [@BLADE #24]  LORD, JEFFREY  P
TORIAN PEARLS, THE [@BLADE #25]  LORD, JEFFREY  P

The MITSFS Seriesdex  September 24, 2007  Page 15
TITLE  AUTHOR  CODE
CITY OF THE LIVING DEAD [@BLADE #26]  LORD, JEFFREY  P
MASTER OF THE HASHOMI [@BLADE #27]  LORD, JEFFREY  P
WIZARD OF RENTORO [@BLADE #28]  LORD, JEFFREY  P
TREASURE OF THE STARS [@BLADE #29]  LORD, JEFFREY  P
DIMENSION OF HORROR [@BLADE #30]  LORD, JEFFREY  P
GLADIATORS OF HAPANU [@BLADE #31]  LORD, JEFFREY  P
PIRATES OF GOHAR [@BLADE #32]  LORD, JEFFREY  P
KILLER PLANTS OF BINAARK [@BLADE #33]  LORD, JEFFREY  P
RUINS OF KALDAK, THE [@BLADE #34]  LORD, JEFFREY  P
LORDS OF THE CRIMSON RIVER, THE [@BLADE #35]  LORD, JEFFREY  P
RETURN TO KALDAK [@BLADE #36]  LORD, JEFFREY  P
WARRIORS OF LATAN [@BLADE #37]  LORD, JEFFREY  P
REPLICANT NIGHT [@BLADE RUNNER]  JETER, K. W.  P
EDGE OF HUMAN, THE [@BLADE RUNNER #2]  JETER, K. W.  P
BLAKE’S SEVEN [BLAKE’S 7]  HOYLE, TREVOR  C/P,P
PROJECT AVALON [BLAKE’S 7]  HOYLE, TREVOR  C/P,P
SCORPIO ATTACK [BLAKE’S 7]  HOYLE, TREVOR  P
TERRY NATION’S BLAKE’S 7 THE PROGRAMME GUIDE [BLAKE’S 7]  ATTWOOD, TONY  R/REF-P
CONVERGENCE [@BLENDING #1]  GREEN, SHARON  C/P,P
COMPETITIONS [@BLENDING #2]  GREEN, SHARON  C/P,P
CHALLENGES [@BLENDING #3]  GREEN, SHARON  C/P,P
BETRAYALS [@BLENDING #4]  GREEN, SHARON  C/P,P
PROPHECY [@BLENDING #5]  GREEN, SHARON  C/P,P
INTRIGUES [@BLENDING ENTHRONED #1]  GREEN, SHARON  C/P,P
DECEPTIONS [@BLENDING ENTHRONED #2]  GREEN, SHARON  P
DESTINY [@BLENDING ENTHRONED #3]  GREEN, SHARON  C/P,P
FIFTH SORCERESS, THE [BLOOD AND STONE 1]  NEWCOMB, ROBERT  H
GATES OF DAWN, THE (PROOF) [BLOOD AND STONE 2]  NEWCOMB, ROBERT  H,SR-LP
SCROLLS OF THE ANCIENTS, THE [BLOOD AND STONE 3]  NEWCOMB, ROBERT  H,P
GATES OF DAWN, THE [@BLOOD AND STONE #2]  NEWCOMB, ROBERT  C/P,P,H
BLOOD BOWL [@BLOOD BOWL 1]  FORBECK, MATT  P
DEAD BALL [@BLOOD BOWL 2]  FORBECK, MATT  P
DEATH MATCH [@BLOOD BOWL 3]  FORBECK, MATT  P

BLOOD OF TEN CHIEFS, THE [@BLOOD OF TEN CHIEFS #1] PINI, RICHARD/ASPRIN, ROBERT LYNN/

ABBEY, LYNN

SFWA-P,C/PA,PA

WOLFSONG [@BLOOD OF TEN CHIEFS #2] PINI, RICHARD/ASPRIN, ROBERT LYNN/

ABBEY, LYNN

SFWA-P,C/PA,LPA

WINDS OF CHANGE [@BLOOD OF TEN CHIEFS #3]  PINI, RICHARD  PA,C/PA,SFWA-P,SFWA
AGAINST THE WIND [@BLOOD OF TEN CHIEFS #4]  PINI, RICHARD  C/PA,LPA,SFWA
DARK HOURS [@BLOOD OF TEN CHIEFS #5]  PINI, RICHARD  LPA
BHAGAVATI [BLOOD OF THE GODDESS 3]  DALKEY, KARA  H
BIJAPUR [BLOOD OF THE GODDESS 2]  DALKEY, KARA  P
GOA [BLOOD OF THE GODDESS 1]  DALKEY, KARA  P
EXPECTED ONE, THE [BLOOD OF THE LAMB 1]  ROGERS, MARK E.  P
DEVOURING VOID, THE [@BLOOD OF THE LAMB #2]  ROGERS, MARK E.  P
RIDDLED MAN, THE [@BLOOD OF THE LAMB #3]  ROGERS, MARK E.  P
DARK DANCE [BLOOD OPERA 1]  LEE, TANITH  H
PERSONAL DARKNESS [BLOOD OPERA 2]  LEE, TANITH  H
BLOODFANG [BLOODFANG 3]  WEAVER, MICHAEL D.  P
NIGHTREAVER [BLOODFANG 2]  WEAVER, MICHAEL D.  P
WOLF-DREAMS [BLOODFANG 1]  WEAVER, MICHAEL D.  P
FIFTH HORSEMAN, THE [BLOODSHADOWS]  STARK, ED  LP
HELL’S FEAST [BLOODSHADOWS]  FARSHTEY, GREG  LPA
BLOSSOM CULP AND THE SLEEP OF DEATH [BLOSSOM CULP 4]  PECK, RICHARD  P
DREADFUL FUTURE OF BLOSSOM CULP, THE [BLOSSOM CULP 3]  PECK, RICHARD  C/P,P
GHOST BELONGED TO ME, THE [BLOSSOM CULP 1]  PECK, RICHARD  P
GHOSTS I HAVE BEEN [BLOSSOM CULP 2]  PECK, RICHARD  C/P:2,P
BOLO [BOLOS]  LAUMER, KEITH  C/P:2,P
BOLO BRIGADE [BOLOS]  KEITH, WILLIAM H. JR.  C/P,P
BOLO RISING [BOLOS]  KEITH, WILLIAM H. JR.  C/P,P
BOLO STRIKE [BOLOS]  KEITH, WILLIAM H. JR.  H
COMPLEAT BOLO, THE [BOLOS]  LAUMER, KEITH  C/P,SFWA-P,P
OLD SOLDIERS [BOLOS]  WEBER, DAVID  H
ROAD TO DAMASCUS, THE [BOLOS]  RINGO, JOHN/EVANS, LINDA  H
ROAD TO DAMASCUS, THE (PROOF) [BOLOS]  RINGO, JOHN/EVANS, LINDA  SR-LP
ROGUE BOLO [BOLOS]  LAUMER, KEITH  C/P:2,P
STARS MUST WAIT, THE [BOLOS]  LAUMER, KEITH  C/P:2,SFWA-P,P
HONOR OF THE REGIMENT [@BOLOS #1]  FAWCETT, BILL  SFWA-P,C/PA,PA
UNCONQUERABLE, THE [@BOLOS #2]  FAWCETT, BILL  PA,C/PA
TRIUMPHANT, THE [@BOLOS #3]  WEBER, DAVID/EVANS, LINDA  P
LAST STAND [@BOLOS #4]  FAWCETT, BILL  C/PA,PA
OLD GUARD [@BOLOS #5]  FAWCETT, BILL  PA
COLD STEEL [@BOLOS #6]  FAWCETT, BILL  PA

TITLE  AUTHOR  CODE
INK [BOOK OF ALL HOURS 2]  DUNCAN, HAL  LP
VELLUM [BOOK OF ALL HOURS 1]  DUNCAN, HAL  LP
BROKEN CHALICE, THE [BOOK OF MAN 2]  WELCH, JANE  P
DAWN OF A DARK AGE [BOOK OF MAN 1]  WELCH, JANE  P
BOOK OF THE SHORT SUN, THE [BOOK OF THE SHORT SUN 1,2,3]  WOLFE, GENE  H:2
IN GREEN’S JUNGLES [BOOK OF THE SHORT SUN 2]  WOLFE, GENE  LP
ON BLUE’S WATERS [BOOK OF THE SHORT SUN 1]  WOLFE, GENE  H,LP
RETURN TO THE WHORL [BOOK OF THE SHORT SUN 3]  WOLFE, GENE  LP
BAKER’S BOY, THE [BOOK OF WORDS 1]  JONES, J. V.  C/P,LP
MAN BETRAYED, A [BOOK OF WORDS 2]  JONES, J. V.  C/P,LP,H
MASTER AND FOOL [BOOK OF WORDS 3]  JONES, J. V.  C/P:2,LP
DEMON LORD, THE [BOOK OF YEARS 2]  MORWOOD, PETER  C/P,P
DRAGON LORD [BOOK OF YEARS 3]  MORWOOD, PETER  P
HORSE LORD, THE [BOOK OF YEARS 1]  MORWOOD, PETER  P
WARLORD’S DOMAIN, THE [BOOK OF YEARS 4]  MORWOOD, PETER  P
INVITATION, THE [@BOOKS OF MAGIC 1]  JABLONSKI, CARLA  P
BINDINGS [@BOOKS OF MAGIC #2]  JABLONSKI, CARLA  P
RECONINGS [@BOOKS OF MAGIC #6]  JABLONSKI, CARLA  P
BLACK BEAST, THE [BOOKS OF THE ISLE 4]  SPRINGER, NANCY  C/P,P
GOLDEN SWAN, THE [BOOKS OF THE ISLE 5]  SPRINGER, NANCY  C/P:2,P
SABLE MOON, THE [BOOKS OF THE ISLE 3]  SPRINGER, NANCY  C/P:2,P
SILVER SUN, THE [BOOKS OF THE ISLE 2]  SPRINGER, NANCY  C/P:2,P
WHITE HART, THE [BOOKS OF THE ISLE 1]  SPRINGER, NANCY  H:2,P
ESSENTIAL BORDERTOWN, THE [BORDERLAND]  WINDLING, TERRI/SHERMAN, DELIA  LP
BORDERLAND [BORDERLANDS 1]  WINDLING, TERRI/ARNOLD, MARK ALAN  PA,C/PA
BORDERTOWN [BORDERLANDS 2]  WINDLING, TERRI/ARNOLD, MARK ALAN  C/PA,PA
ELSEWHERE [BORDERLANDS 5]  SHETTERLY, WILL  H:2,P
FINDER [BORDERLANDS 4]  BULL, EMMA  C/P:2,P
LIFE ON THE BORDER [BORDERLANDS 3]  WINDLING, TERRI  PA,C/PA
NEVERNEVER [BORDERLANDS 6]  SHETTERLY, WILL  C/P,P,H
COVENANTS [BORDERLANDS (IVERSTERRE)]  FREEMAN, LORNA  P
KING’S OWN, THE [@BORDERLANDS (IVERSTERRE)]  FREEMAN, LORNA  P
CONQUEROR’S MOON [BOREAL MOON 1]  MAY, JULIAN  H
IRONCROWN MOON [BOREAL MOON 2]  MAY, JULIAN  H
ACROSS THE DARK METROPOLIS [BORRIBLES 3]  DE LARRABEITI, MICHAEL  P
BORRIBLES, THE [BORRIBLES 1]  DE LARRABEITI, MICHAEL  C/P,P
BORRIBLES GO FOR BROKE, THE [BORRIBLES 2]  DE LARRABEITI, MICHAEL  P
BOOK OF FREEDOMS, THE [BOTANY 1,2,3]  MCCAFFREY, ANNE  H
FREEDOM’S CHALLENGE [BOTANY 3]  MCCAFFREY, ANNE  H:2,P
FREEDOM’S CHOICE [BOTANY 2]  MCCAFFREY, ANNE  C/P:2,P,H
FREEDOM’S LANDING [BOTANY 1]  MCCAFFREY, ANNE  H:3,P
FREEDOM’S RANSOM [BOTANY 4]  MCCAFFREY, ANNE  H:2
WEAVERS OF SARAMYR [BRAIDED PATH 1]  WOODING, CHRIS  LP
BRAK THE BARBARIAN [BRAK 1]  JAKES, JOHN  P
BRAK THE BARBARIAN (JAPANESE) [BRAK 1]  JAKES, JOHN  FP
BRAK THE BARBARIAN VERSUS THE MARK OF THE DEMONS [BRAK 2]  JAKES, JOHN  C/P:2,P
BRAK THE BARBARIAN VERSUS THE SORCERESS [BRAK 3]  JAKES, JOHN  C/P,P
FORTUNES OF BRAK, THE [BRAK 5]  JAKES, JOHN  P
WHEN THE IDOLS WALKED [BRAK 4]  JAKES, JOHN  P
BRAN MAK MORN [BRAN MAK MORN 1]  HOWARD, ROBERT E.  SR-P,C/P
BRAN MAK MORN: THE LAST KING [BRAN MAK MORN]  HOWARD, ROBERT E.  H
FOR THE WITCH OF THE MISTS [BRAN MAK MORN 3]  SMITH, DAVID C./TIERNEY, RICHARD L.  C/P,P
LEGION FROM THE SHADOWS [BRAN MAK MORN 2]  WAGNER, KARL EDWARD  C/P:2,SFWA-P,P
ANTIPOPE, THE [BRENTFORD 1]  RANKIN, ROBERT  P
BRENTFORD TRIANGLE, THE [BRENTFORD 2]  RANKIN, ROBERT  P
EAST OF EALING [BRENTFORD 3]  RANKIN, ROBERT  P
KNEES UP MOTHER EARTH [BRENTFORD 7/WITCHES OF CHISWICK 2]  RANKIN, ROBERT  H
SPROUTS OF WRATH, THE [BRENTFORD 4]  RANKIN, ROBERT  P
WELL OF SHADES, THE [BRIDEI 3]  MARILLIER, JULIET  H
DARK MIRROR, THE [@BRIDEI #1]  MARILLIER, JULIET  H
BLADE OF FORTRIU [@BRIDEI #2]  MARILLIER, JULIET  H
DAUGHTER OF ANCIENTS [BRIDGE OF D’ARNATH 4]  BERG, CAROL  P
GUARDIANS OF THE KEEP [BRIDGE OF D’ARNATH 2]  BERG, CAROL  P
SON OF AVONAR [BRIDGE OF D’ARNATH 1]  BERG, CAROL  P
SOUL WEAVER, THE [BRIDGE OF D’ARNATH 3]  BERG, CAROL  P
MOTHER OCEAN, DAUGHTER SEA [BRIERLY 1]  MARCELLAS, DIANA  P
SEA LARK’S SONG, THE [BRIERLY 2]  MARCELLAS, DIANA  P
EMBASSY [BRION BAYARD 1]  LAUMER, KEITH  P
BROKEN CITADEL, THE [BROKEN CITADEL 1]  GREGORIAN, JOYCE BALLOU  C/P,P
CASTLEDOWN [BROKEN CITADEL 2]  GREGORIAN, JOYCE BALLOU  P
GREAT WHEEL, THE [BROKEN CITADEL 3]  GREGORIAN, JOYCE BALLOU  SFWA-P
BROMELIAD TRILOGY, THE [BROMELIAD 1,2,3]  PRATCHETT, TERRY  H
HEARTS AND ARMOR [BRONWYN 3]  MILLER, RON  P

The MITSFS Seriesdex  September 24, 2007  Page 17
TITLE  AUTHOR  CODE
PALACES AND PRISONS (REVISED) [BRONWYN 1]  MILLER, RON  LP
SILK AND STEEL [BRONWYN 2]  MILLER, RON  C/P,P
MYSTIC WARRIOR [@BRONZE CANTICLES #1]  HICKMAN, TRACY/HICKMAN, LAURA  P
MYSTIC QUEST [@BRONZE CANTICLES #2]  HICKMAN, TRACY/HICKMAN, LAURA  H:2
MYSTIC EMPIRE [@BRONZE CANTICLES #3]  HICKMAN, TRACY/HICKMAN, LAURA  H
BROTHERS OF THE DRAGON [BROTHERS OF THE DRAGON 1]  BAILEY, ROBIN WAYNE  C/P:2,P
FLAMES OF THE DRAGON [BROTHERS OF THE DRAGON 2]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
TRIUMPH OF THE DRAGON [BROTHERS OF THE DRAGON 3]  BAILEY, ROBIN WAYNE  C/P,P
ARMAGEDDON 2419 A.D. [BUCK ROGERS]  NOWLAN, PHILIP FRANCIS  C/P,P,H

ARMAGEDDON 2419 A.D. (REVISED) [BUCK ROGERS/ BUCK ROGERS (NIVEN & POURNELLE’S) 1] NOWLAN, PHILIP FRANCIS

C/P,P

ARRIVAL [BUCK ROGERS]  TSR  PA

COLLECTED WORKS OF BUCK ROGERS IN THE 25TH CENTURY, THE [BUCK ROGERS] DILLE, ROBERT C.

CX

LIFE IN THE FUTURE, A [BUCK ROGERS]  CAIDIN, MARTIN  H
BUCK ROGERS IN THE 25TH CENTURY [@BUCK ROGERS #1]  STEELE, ADDISON E.  P
THAT MAN ON BETA [@BUCK ROGERS #2]  STEELE, ADDISON E.  C/P,P

MORDRED [BUCK ROGERS (NIVEN & POURNELLE’S) 2] HOLMES, JOHN ERIC/NIVEN, LARRY/

POURNELLE, JERRY

C/P,P

ROGERS’ RANGERS [BUCK ROGERS (NIVEN & POURNELLE’S) 5]  SILBERSACK, JOHN  P
WARRIOR’S BLOOD [BUCK ROGERS (NIVEN & POURNELLE’S) 3]  MCENROE, RICHARD S.  P
WARRIOR’S WORLD [BUCK ROGERS (NIVEN & POURNELLE’S) 4]  MCENROE, RICHARD S.  P
FIRST POWER PLAY [BUCK ROGERS INNER PLANETS 1]  MILLER, JOHN  C/P:2,P
MATRIX CUBED [BUCK ROGERS INNER PLANETS 3]  BLOOM, BRITTON  C/P,P
PRIME SQUARED [BUCK ROGERS INNER PLANETS 2]  MURDOCK, M. S.  C/P:2,P
GENESIS WEB, THE [BUCK ROGERS INVADERS OF CHARON 1]  BRENNAN, C. M.  P
NOMADS OF THE SKY [BUCK ROGERS INVADERS OF CHARON 2]  KEITH, WILLIAM H. JR.  P
ARMAGEDDON OFF VESTA [BUCK ROGERS MARTIAN WARS 3]  MURDOCK, M. S.  C/P,P
HAMMER OF MARS [BUCK ROGERS MARTIAN WARS 2]  MURDOCK, M. S.  C/P,P
REBELLION 2456 [BUCK ROGERS MARTIAN WARS 1]  MURDOCK, M. S.  C/P,P
AFTERIMAGE [@BUFFY]  ASKEGREN, PIERCE  P
ANGEL CHRONICLES, THE, VOL. 1 [BUFFY]  HOLDER, NANCY  C/P,P
ANGEL CHRONICLES, THE, VOL. 2 [BUFFY]  TANKERSLEY, RICHIE  C/P,P
ANGEL CHRONICLES, THE, VOL. 3 [BUFFY]  HOLDER, NANCY  C/P,P
APOCALYPSE MEMORIES [@BUFFY]  BURNS, LAURA J./METZ, MELINDA  C/P,P
BAD BARGAIN [@BUFFY]  GALLAGHER, DIANA G.  P
BLACKOUT [@BUFFY]  DECANDIDO, KEITH R. A.  P
BLOOD AND FOG [@BUFFY]  HOLDER, NANCY  P
BLOODED [BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
BOOK OF FOURS, THE [@BUFFY]  HOLDER, NANCY  H:2,P
BUFFY THE VAMPIRE SLAYER [BUFFY]  CUSICK, RICHIE TANKERSLEY  P
CARNIVAL OF SOULS [@BUFFY]  HOLDER, NANCY  P
CHAOS BLEEDS [@BUFFY]  MOORE, JAMES A.  P
CHILD OF THE HUNT [@BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
CHOSEN [@BUFFY]  HOLDER, NANCY  P
CORDELIA COLLECTION, THE, VOL. 1 [@BUFFY]  KRULIK, NANCY  P
COYOTE MOON [BUFFY]  VORNHOLT, JOHN  P
CROSSINGS [@BUFFY]  ODOM, MEL  P
CURSED [@BUFFY/ANGEL]  ODOM, MEL  C/P,P
DEATHLESS, THE [@BUFFY]  DECANDIDO, KEITH R. A.  P
DEEP WATER [BUFFY]  GILMAN, LAURA ANNE/SHERMAN, JOSEPHA P
DOOMSDAY DECK [BUFFY]  GALLAGHER, DIANA G.  P
EVIL THAT MEN DO, THE [@BUFFY]  HOLDER, NANCY  C/P,P
FAITH TRIALS, THE, VOL. 1 [BUFFY]  LAURENCE, JAMES  P
GHOUL TROUBLE [BUFFY]  PASSARELLA, JOHN  P
GO ASK MALICE [@BUFFY]  LEVY, JOSEPH  H
HALLOWEEN RAIN [BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  P
HARVEST, THE [BUFFY]  CUSICK, RICHIE TANKERSLEY  P
HEAT [@BUFFY/ANGEL]  HOLDER, NANCY  P
HERE BE MONSTERS [BUFFY]  DOKEY, CAMERON  P
HOW I SURVIVED MY SUMMER VACATION, VOL. 1 [BUFFY]  TWENTIETH CENTURY FOX  C/PA,PA
IMMORTAL [BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
LITTLE THINGS [@BUFFY]  MOESTA, REBECCA  P

MONSTER ISLAND [@BUFFY/ANGEL] GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

P

MORTAL FEAR [@BUFFY]  CIENCIN, SCOTT/CIENCIN, DENISE  LP
NIGHT OF THE LIVING RERUN [BUFFY]  COVER, ARTHUR BYRON  P
OBSIDIAN FATE [@BUFFY]  GALLAGHER, DIANA G.  C/P,P
OZ: INTO THE WILD [@BUFFY]  GOLDEN, CHRISTOPHER  P
PALEO [@BUFFY]  NAVARRO, YVONNE  C/P,P
PORTAL THROUGH TIME [@BUFFY]  HENDERSON, ALICE  P
POWER OF PERSUASION [BUFFY]  MASSIE, ELIZABETH  P
PRIME EVIL [@BUFFY]  GALLAGHER, DIANA G.  C/P,P

TITLE  AUTHOR  CODE
QUEEN OF THE SLAYERS [@BUFFY]  HOLDER, NANCY  H,LP
RESURRECTING RAVANA [BUFFY]  GARTON, RAY  C/P,P
RETURN TO CHAOS [@BUFFY]  GARDNER, CRAIG SHAW  C/P:2,P
REVENANT [BUFFY]  ODOM, MEL  C/P,P
SEVEN CROWN [@BUFFY/ANGEL]  VORNHOLT, JOHN  C/P,P
SINS OF THE FATHER [BUFFY]  GOLDEN, CHRISTOPHER  C/P,P
SPARK AND BURN [@BUFFY]  GALLAGHER, DIANA G.  P
SWEET SIXTEEN [@BUFFY]  CIENCIN, SCOTT  P
TALES OF THE SLAYER, VOL. 1 [BUFFY]  POCKET PULSE  LP
TALES OF THE SLAYER, VOL. 4 [@BUFFY]  SIMON SPOTLIGHT ENTERTAINMENT  LPA
TALES OF THE SLAYER: VOLUMES 2 & 3 [BUFFY]  SFBC  HA
TEMPTED CHAMPIONS [@BUFFY]  NAVARRO, YVONNE  C/P,P
THESE OUR ACTORS [@BUFFY]  MCCONNELL, ASHLEY/KOOGLER, DORI  C/P,P
UNNATURAL SELECTION [BUFFY]  ODOM, MEL  P
VISITORS [BUFFY]  GILMAN, LAURA ANNE/SHERMAN, JOSEPHA P
WILLOW FILES, THE, VOL. 1 [@BUFFY]  NAVARRO, YVONNE  P
WILLOW FILES, THE, VOL. 2 [@BUFFY]  NAVARRO, YVONNE  P
WISDOM OF WAR, THE [@BUFFY]  GOLDEN, CHRISTOPHER  C/P,P
XANDER YEARS, THE, VOL. 1 [BUFFY]  DECANDIDO, KEITH R. A.  P
XANDER YEARS, THE, VOL. 2 [BUFFY]  MARIOTTE, JEFF  P
BURNING, THE [@BUFFY ANGEL UNSEEN 1]  HOLDER, NANCY/MARIOTTE, JEFF  P
DOOR TO ALTERNITY [@BUFFY ANGEL UNSEEN 2]  HOLDER, NANCY/MARIOTTE, JEFF  C/P,P
LONG WAY HOME [@BUFFY ANGEL UNSEEN 3]  HOLDER, NANCY/MARIOTTE, JEFF  P
GHOST ROADS [@BUFFY GATEKEEPER 2]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P:2,P
OUT OF THE MADHOUSE [@BUFFY GATEKEEPER 1]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
SONS OF ENTROPY [@BUFFY GATEKEEPER 3]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  P
LOST SLAYER, THE [@BUFFY LOST SLAYER 1,2,3,4]  GOLDEN, CHRISTOPHER  P
PROPHECIES [@BUFFY LOST SLAYER #1]  GOLDEN, CHRISTOPHER  C/P
DARK TIMES [@BUFFY LOST SLAYER #2]  GOLDEN, CHRISTOPHER  C/P
KING OF THE DEAD [@BUFFY LOST SLAYER #3]  GOLDEN, CHRISTOPHER  C/P:2
ORIGINAL SINS [@BUFFY LOST SLAYER #4]  GOLDEN, CHRISTOPHER  C/P
SPIKE VS. DRACULA [BUFFY SPIKE]  DAVID, PETER  CX
PRETTY MAIDS ALL IN A ROW [BUFFY SPIKE & DRU]  GOLDEN, CHRISTOPHER  H,P
DARKENING, THE [@BUFFY WICKED WILLOW #1]  NAVARRO, YVONNE  C/P,P
SHATTERED TWILIGHT [@BUFFY WICKED WILLOW #2]  NAVARRO, YVONNE  P
BROKEN SUNRISE [@BUFFY WICKED WILLOW #3]  NAVARRO, YVONNE  C/P,P
BUNNICULA [BUNNICULA 1]  HOWE, DEBORAH/HOWE, JAMES  LP
CELERY STALKS AT MIDNIGHT, THE [BUNNICULA 3]  HOWE, JAMES  LP
HOWLIDAY INN [BUNNICULA 2]  HOWE, JAMES  LP
CAUSE OF DEATH, THE [BUREAU OF SPECIAL INVESTIGATIONS]  ALLEN, ROGER MACBRIDE  P
BUREAU 13 [BUREAU 13 1]  POLLOTTA, NICK  P
DAMNED NATION [@BUREAU 13]  POLLOTTA, NICK  LP
DOOMSDAY EXAM [BUREAU 13 2]  POLLOTTA, NICK  C/P,P
FULL MOONSTER [BUREAU 13 3]  POLLOTTA, NICK  P
JUDGMENT NIGHT [@BUREAU 13]  POLLOTTA, NICK  LP
BUTTERFLY AND HELLFLOWER [BUTTERFLY ST. CYR 1,2,3]  BES SHAHAR, ELUKI  H

The MITSFS Seriesdex  September 24, 2007  Page 19
TITLE  AUTHOR  CODE
BLACK MOON, THE [CABAL 2]  DUNN, PHILIP  P
CABAL, THE [CABAL 1]  DUNN, PHILIP  P
EVANGELIST, THE [CABAL 3]  DUNN, PHILIP  P
CADRE LUCIFER [CADRE 2]  O’RIORDAN, ROBERT  C/P,SFWA-P,P
CADRE MESSIAH [CADRE 3]  O’RIORDAN, ROBERT  C/P:2,P
CADRE ONE [CADRE 1]  O’RIORDAN, ROBERT  C/P:2,P
C.A.D.S. [@C.A.D.S. #1]  SIEVERT, JOHN  P
TECH BATTLEGROUND [@C.A.D.S. #2]  SIEVERT, JOHN  P
TECH COMMANDO [@C.A.D.S. #3]  SIEVERT, JOHN  P
TECH STRIKE FORCE [@C.A.D.S. #4]  SIEVERT, JOHN  P
TECH SATAN [@C.A.D.S. #5]  SIEVERT, JOHN  P
DOOM COMMANDER [@C.A.D.S. #7]  SIEVERT, JOHN  P
ARAMINTA STATION [CADWAL 1]  VANCE, JACK  C/P,SFWA-P,P,H
ECCE AND OLD EARTH [CADWAL 2]  VANCE, JACK  C/P,P,H
THROY [CADWAL 3]  VANCE, JACK  H
LOST WORLDS OF CRONUS, THE [CAGEWORLD 2]  KAPP, COLIN  C/P,P
SEARCH FOR THE SUN [CAGEWORLD 1]  KAPP, COLIN  C/P:2,P
TYRANT OF HADES, THE [CAGEWORLD 3]  KAPP, COLIN  C/P,P
CALL OF MADNESS [CAITHAN CRUSADE 1]  SMITH, JULIE DEAN  C/P,P
MISSION OF MAGIC [CAITHAN CRUSADE 2]  SMITH, JULIE DEAN  P
SAGE OF SARE [CAITHAN CRUSADE 3]  SMITH, JULIE DEAN  C/P,P
WIZARD KING, THE [CAITHAN CRUSADE 4]  SMITH, JULIE DEAN  P
MOONSHINE [CAL AND NIKO 2]  THURMAN, ROB  P
NIGHTLIFE [CAL AND NIKO 1]  THURMAN, ROB  P
GUNS, DRUGS, AND MONSTERS [CAL MCDONALD 2]  NILES, STEVE  P
CALEDON OF THE MISTS [CALEDON 1]  HARRIS, DEBORAH TURNER  C/P,SFWA-P,P
QUEEN OF ASHES, THE [CALEDON 2]  HARRIS, DEBORAH TURNER  C/P,P
CALIBAN [CALIBAN 1]  ALLEN, ROGER MACBRIDE  H,SFWA:2,P
INFERNO [CALIBAN 2]  ALLEN, ROGER MACBRIDE  P
UTOPIA [CALIBAN 3]  ALLEN, ROGER MACBRIDE  H,P
DRAGONS ON THE TOWN [CALIFORNIA DRAGONS 3]  GUNNARSSON, THORARINN  P
HUMAN, BEWARE! [CALIFORNIA DRAGONS 2]  GUNNARSSON, THORARINN  P
MAKE WAY FOR DRAGONS! [CALIFORNIA DRAGONS 1]  GUNNARSSON, THORARINN  C/P,P
GREENMAGIC [CALINDOR 1]  KILIAN, CRAWFORD  P
REDMAGIC [CALINDOR 2]  KILIAN, CRAWFORD  P
CALLAHAN TOUCH, THE [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P,P,H
CALLAHAN’S CON [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  H
CALLAHAN’S CROSSTIME SALOON [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P:2,P
CALLAHAN’S KEY [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  H,P
CALLAHAN’S LADY [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P,SFWA-P:2,P,H,SFWA
CALLAHAN’S LEGACY [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  H,P
CALLAHAN’S SECRET [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P:2,P
LADY SLINGS THE BOOZE [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P,P,H
OFF THE WALL AT CALLAHAN’S [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  LP
TIME TRAVELERS STRICTLY CASH [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P:2,P
CAMELOT IN ORBIT [CAMELOT 2]  LANDIS, ARTHUR H.  C/P,P
HOME — TO AVALON [CAMELOT 4]  LANDIS, ARTHUR H.  C/P,P
MAGICK OF CAMELOT, THE [CAMELOT 3]  LANDIS, ARTHUR H.  C/P:2,P
WORLD CALLED CAMELOT, A [CAMELOT 1]  LANDIS, ARTHUR H.  C/P,P
DINOSAUR THAT FOLLOWED ME HOME, THE [CAMP HAUNTED HILL 3]  COVILLE, BRUCE  LP
EAGLE, THE [CAMULOD 9]  WHYTE, JACK  H:2
EAGLES’ BROOD, THE [CAMULOD 3]  WHYTE, JACK  P
FORT AT RIVER’S BEND, THE [CAMULOD 5]  WHYTE, JACK  LP
LANCE THROWER, THE [CAMULOD]  WHYTE, JACK  P
SAXON SHORE, THE [CAMULOD 4]  WHYTE, JACK  P
SKYSTONE, THE [CAMULOD 1]  WHYTE, JACK  P
SORCERER: METAMORPHOSIS, THE [CAMULOD 6]  WHYTE, JACK  P
UTHER [CAMULOD]  WHYTE, JACK  P

DOCUMENTS RELATING TO THE SENTIMENTAL AGENTS IN THE VOLYEN EMPIRE [CANOPUS IN ARGOS 5] LESSING, DORIS

H,LP

MAKING OF THE REPRESENTATIVE FOR PLANET 8, THE [CANOPUS IN ARGOS 4] LESSING, DORIS

H,LP

MARRIAGES BETWEEN ZONES THREE, FOUR, AND FIVE, THE [CANOPUS IN ARGOS 2] LESSING, DORIS

H:2,LP

RE: COLONISED PLANET 5, SHIKASTA [CANOPUS IN ARGOS 1]  LESSING, DORIS  H,LP
SIRIAN EXPERIMENTS, THE [CANOPUS IN ARGOS 3]  LESSING, DORIS  H:2,LP
RECEIVE THE GIFT [CANTRIX 3]  MARLEY, LOUISE  P
SING THE LIGHT [CANTRIX 1]  MARLEY, LOUISE  P
SING THE WARMTH [CANTRIX 2]  MARLEY, LOUISE  P
GALAXY OF THE LOST [@CAP KENNEDY #1]  KERN, GREGORY  C/P:2,P
SLAVE SHIP FROM SERGAN [@CAP KENNEDY #2]  KERN, GREGORY  C/P,P
MONSTER OF METELAZE [@CAP KENNEDY #3]  KERN, GREGORY  P
ENEMY WITHIN THE SKULL [@CAP KENNEDY #4]  KERN, GREGORY  P

TITLE  AUTHOR  CODE
JEWEL OF JARHEN [@CAP KENNEDY #5]  KERN, GREGORY  P
SEETEE ALERT! [@CAP KENNEDY #6]  KERN, GREGORY  C/P,P
GHOLAN GATE, THE [@CAP KENNEDY #7]  KERN, GREGORY  P
EATER OF WORLDS, THE [@CAP KENNEDY #8]  KERN, GREGORY  C/P,P
EARTH ENSLAVED [@CAP KENNEDY #9]  KERN, GREGORY  P
PLANET OF DREAD [@CAP KENNEDY #10]  KERN, GREGORY  C/P,P
SPAWN OF LABAN [@CAP KENNEDY #11]  KERN, GREGORY  C/P,P
GENETIC BUCCANEER, THE [@CAP KENNEDY #12]  KERN, GREGORY  C/P,P
WORLD AFLAME, A [@CAP KENNEDY #13]  KERN, GREGORY  P
GHOSTS OF EPIDORIS, THE [@CAP KENNEDY #14]  KERN, GREGORY  C/P,P
MIMICS OF DEPHENE [@CAP KENNEDY #15]  KERN, GREGORY  C/P,P
BEYOND THE GALACTIC LENS [@CAP KENNEDY #16]  KERN, GREGORY  P
CAPTAINS OUTRAGEOUS [CAPTAINS OUTRAGEOUS 1]  YOUNG, ROY V.  P
YOR’S REVENGE [CAPTAINS OUTRAGEOUS 2]  YOUNG, ROY V.  P
BACK FROM HELL [CAR WARRIORS 3]  FARREN, MICK  P
DOUBLE JEOPARDY [CAR WARRIORS 2]  ALLSTON, AARON  P
SQUARE DEAL, THE [@CAR WARRIORS #1]  DRAKE, DAVID  LP
CYBERSTEALTH [CARGO AND GHOSTER 1]  LEWITT, SHARIANN  P
DANCING VAC [CARGO AND GHOSTER 2]  LEWITT, SHARIANN  P
LEOPARD IN EXILE [CAROLUS REX 2]  NORTON, ANDRE/EDGHILL, ROSEMARY  H,P
SHADOW OF ALBION, THE [CAROLUS REX 1]  NORTON, ANDRE/EDGHILL, ROSEMARY  H:2,P
DARK CELEBRATION [CARPATHIAN 17]  FEEHAN, CHRISTINE  H
DARK DEMON [CARPATHIAN 16]  FEEHAN, CHRISTINE  P
DARK DESIRE [CARPATHIAN 2]  FEEHAN, CHRISTINE  P
DARK DESTINY [CARPATHIAN 13]  FEEHAN, CHRISTINE  P
DARK MAGIC [CARPATHIAN 4]  FEEHAN, CHRISTINE  P
DARK SECRET [CARPATHIAN 15]  FEEHAN, CHRISTINE  H
DARK SYMPHONY [CARPATHIAN 10]  FEEHAN, CHRISTINE  P
CASCA THE DEFIANT [CASCA]  DENGELEGI, PAUL  P
CASCA THE LIBERATOR [CASCA]  DENGELEGI, PAUL  P
ETERNAL MERCENARY, THE [@CASCA #1]  SADLER, BARRY  C/P,P
GOD OF DEATH [@CASCA #2]  SADLER, BARRY  P
WARLORD, THE [@CASCA #3]  SADLER, BARRY  C/P,P
PANZER SOLDIER [@CASCA #4]  SADLER, BARRY  C/P,P
BARBARIAN, THE [@CASCA #5]  SADLER, BARRY  P
PERSIAN, THE [@CASCA #6]  SADLER, BARRY  P
DAMNED, THE [@CASCA #7]  SADLER, BARRY  P
SOLDIER OF FORTUNE [@CASCA #8]  SADLER, BARRY  P
SENTINEL, THE [@CASCA #9]  SADLER, BARRY  P
CONQUISTADOR, THE [@CASCA #10]  SADLER, BARRY  P
LEGIONNAIRE, THE [@CASCA #11]  SADLER, BARRY  P
AFRICAN MERCENARY, THE [@CASCA #12]  SADLER, BARRY  P
ASSASSIN, THE [@CASCA #13]  SADLER, BARRY  P
PHOENIX, THE [@CASCA #14]  SADLER, BARRY  P
PIRATE, THE [@CASCA #15]  SADLER, BARRY  P
DESERT MERCENARY [@CASCA #16]  SADLER, BARRY  P
WARRIOR, THE [@CASCA #17]  SADLER, BARRY  P
CURSED, THE [@CASCA #18]  SADLER, BARRY  P
SAMURAI, THE [@CASCA #19]  SADLER, BARRY  P
SOLDIER OF GIDEON [@CASCA #20]  SADLER, BARRY  P
TRENCH SOLDIER, THE [@CASCA #21]  SADLER, BARRY  P
MONGOL, THE [@CASCA #22]  SADLER, BARRY  P
BREAKAWAY [CASSANDRA KRESNOV 2]  SHEPHERD, JOEL  LP
CROSSOVER [CASSANDRA KRESNOV 1]  SHEPHERD, JOEL  LP
CLAIMED BY SHADOW [CASSANDRA PALMER]  CHANCE, KAREN  P
TOUCH THE DARK [CASSANDRA PALMER]  CHANCE, KAREN  P
CASTAWAYS IN TIME [@CASTAWAYS IN TIME #1]  ADAMS, ROBERT  C/P,LP
SEVEN MAGICAL JEWELS OF IRELAND, THE [@CASTAWAYS IN TIME #2]  ADAMS, ROBERT  C/P,P
OF QUESTS AND KINGS [@CASTAWAYS IN TIME #3]  ADAMS, ROBERT  C/P:2,P
OF CHIEFS AND CHAMPIONS [@CASTAWAYS IN TIME #4]  ADAMS, ROBERT  P
OF MYTHS AND MONSTERS [@CASTAWAYS IN TIME #5]  ADAMS, ROBERT  P
OF BEGINNINGS AND ENDINGS [@CASTAWAYS IN TIME #6]  ADAMS, ROBERT  P
ANGEL’S COMMAND, THE [CASTAWAYS OF THE FLYING DUTCHMAN]  JACQUES, BRIAN  P
BRIDE OF THE CASTLE [CASTLE 8]  DE CHANCIE, JOHN  C/P,P
CASTLE DREAMS [CASTLE 6]  DE CHANCIE, JOHN  C/P:2,P
CASTLE FOR RENT [CASTLE 2]  DE CHANCIE, JOHN  C/P:2,SFWA-P,P
CASTLE IN THE AIR [CASTLE 2]  JONES, DIANA WYNNE  H
CASTLE KIDNAPPED [CASTLE 3]  DE CHANCIE, JOHN  C/P:2,SFWA-P,P
CASTLE MURDERS [CASTLE 5]  DE CHANCIE, JOHN  C/P:2,P
CASTLE PERILOUS [CASTLE 1]  DE CHANCIE, JOHN  C/P:2,P
CASTLE SPELLBOUND [CASTLE 7]  DE CHANCIE, JOHN  C/P,P
CASTLE WAR! [CASTLE 4]  DE CHANCIE, JOHN  C/P:2,SFWA-P,P
HOWL’S MOVING CASTLE [CASTLE 1]  JONES, DIANA WYNNE  P

TITLE  AUTHOR  CODE
CHAMPION OF GARATHORM, THE [CASTLE BRASS 2]  MOORCOCK, MICHAEL  C/P,P
COUNT BRASS [CASTLE BRASS 1]  MOORCOCK, MICHAEL  C/P,P
QUEST FOR TANELORN, THE [CASTLE BRASS 3]  MOORCOCK, MICHAEL  C/P:2,P
QUEST FOR TANELORN, THE (BRITISH) [CASTLE BRASS 3]  MOORCOCK, MICHAEL  SR-P
FROM PRUSSIA WITH LOVE [CASTLE FALKENSTEIN]  DE CHANCIE, JOHN  P
MASTERMINDS OF FALKENSTEIN [@CASTLE FALKENSTEIN]  DE CHANCIE, JOHN  P
ALIEN BLOOD [CAT 1,2]  VINGE, JOAN D.  H:2
CATSPAW [CAT 2]  VINGE, JOAN D.  C/P:2,P
DREAMFALL [CAT 3]  VINGE, JOAN D.  H:2
PSION [CAT 1]  VINGE, JOAN D.  C/P:2,P
CROOKED LETTER, THE [CATACLYSM 1]  WILLIAMS, SEAN  H
BLOOD DEBT, THE [@CATACLYSM #2]  WILLIAMS, SEAN  H
HANGING MOUNTAINS [@CATACLYSM #3]  WILLIAMS, SEAN  H
HAYDN OF MARS [CATS OF MARS 1]  SARRANTONIO, AL  P
QUEEN OF MARS [CATS OF MARS 3]  SARRANTONIO, AL  P
SEBASTIAN OF MARS [CATS OF MARS 2]  SARRANTONIO, AL  P
GRAND TOUR, THE [CECILIA & KATHERINE 2]  WREDE, PATRICIA C./STEVERMER, CAROLINE H
MAGICIANS OF QUALITY [CECILIA & KATHERINE 1,2]  WREDE, PATRICIA C./STEVERMER, CAROLINE H
MISLAID MAGICIAN, THE [CECILIA & KATHERINE 3]  WREDE, PATRICIA C./STEVERMER, CAROLINE H:2
SORCERY AND CECELIA [CECILIA & KATHERINE 1]  WREDE, PATRICIA C./STEVERMER, CAROLINE C/P,P
TEMPLE AT LANDFALL, THE [CELAENO]  FLETCHER, JANE  LP
WALLS OF WESTERNFORT, THE [CELAENO]  FLETCHER, JANE  LP
BLACK ROOD, THE [CELTIC CRUSADES 2]  LAWHEAD, STEPHEN R.  P
IRON LANCE, THE [CELTIC CRUSADES 1]  LAWHEAD, STEPHEN R.  P
MYSTIC ROSE, THE [@CELTIC CRUSADES 3]  LAWHEAD, STEPHEN R.  P
CHILD OF SATURN [CELYDONN GREEN LION 1]  EDGERTON, TERESA  C/P:2,SFWA-P,P
MOON IN HIDING, THE [CELYDONN GREEN LION 2]  EDGERTON, TERESA  C/P:2,SFWA-P,P
WORK OF THE SUN, THE [CELYDONN GREEN LION 3]  EDGERTON, TERESA  P,SFWA-P:2
CASTLE OF THE SILVER WHEEL, THE [CELYDONN GWENLLIANT 1]  EDGERTON, TERESA  SFWA-P:2
GRAIL AND THE RING, THE [CELYDONN GWENLLIANT 2]  EDGERTON, TERESA  C/P,SFWA-P,P
MOON AND THE THORN, THE [CELYDONN GWENLLIANT 3]  EDGERTON, TERESA  C/P,P
CENOTAPH ROAD [CENOTAPH ROAD 1]  VARDEMAN, ROBERT E.  P
FIRE AND FOG [CENOTAPH ROAD 5]  VARDEMAN, ROBERT E.  C/P,P
IRON TONGUE [CENOTAPH ROAD 4]  VARDEMAN, ROBERT E.  C/P,P
PILLAR OF NIGHT [CENOTAPH ROAD 6]  VARDEMAN, ROBERT E.  C/P:2,P
SORCERER’S SKULL, THE [CENOTAPH ROAD 2]  VARDEMAN, ROBERT E.  C/P:2,P
WORLD OF MAZES [CENOTAPH ROAD 3]  VARDEMAN, ROBERT E.  C/P,P
EVOLUTION’S SHORE [CHAGA 1]  MCDONALD, IAN  C/P,P,H
KIRINYA [CHAGA 2]  MCDONALD, IAN  P
TENDELEO’S STORY [CHAGA]  MCDONALD, IAN  GZD07305,SR-LP
CHAINS OF DESTRUCTION [CHAINS 2]  ROSEN, SELINA  LP
CHAINS OF FREEDOM [CHAINS 1]  ROSEN, SELINA  LP
CURSE OF CHALION, THE [CHALION 1]  BUJOLD, LOIS MCMASTER  H:2
HALLOWED HUNT, THE [CHALION 3]  BUJOLD, LOIS MCMASTER  H:2
PALADIN OF SOULS [CHALION 2]  BUJOLD, LOIS MCMASTER  H
TO CHALLENGE CHAOS [CHALLENGE CHAOS 1]  STABLEFORD, BRIAN M.  C/P:2,P
CHANGELING WAR, THE [@CHANGELING #1]  GARRISON, PETER  C/P,LP
SORCERER’S GUN, THE [@CHANGELING #2]  GARRISON, PETER  C/P,LP
MAGIC DEAD, THE [@CHANGELING #3]  GARRISON, PETER  P
SHADOWS IN THE DARKNESS [CHANGELING DETECTIVE 1]  CUNNINGHAM, ELAINE  H
SHADOWS IN THE STARLIGHT [CHANGELING DETECTIVE 2]  CUNNINGHAM, ELAINE  H
DOWN THE STREAM OF STARS [CHANGELING STAR 2]  CARVER, JEFFREY A.  C/P:2,P
FROM A CHANGELING STAR [CHANGELING STAR 1]  CARVER, JEFFREY A.  H:2,P
CHANGES, THE [CHANGES 1,2,3]  DICKINSON, PETER  H,LP
DEVIL’S CHILDREN, THE [CHANGES 1]  DICKINSON, PETER  H
HEARTSEASE [CHANGES 2]  DICKINSON, PETER  H:2
WEATHERMONGER, THE [CHANGES 3]  DICKINSON, PETER  C/P,P,H
CHANGEWINDS, THE [CHANGEWINDS 1,2,3]  CHALKER, JACK L.  LP
RIDERS OF THE WINDS [CHANGEWINDS 2]  CHALKER, JACK L.  C/P:2,SFWA-P,P
WAR OF THE MAELSTROM [CHANGEWINDS 3]  CHALKER, JACK L.  C/P:3,SFWA-P,P
WHEN THE CHANGEWINDS BLOW [CHANGEWINDS 1]  CHALKER, JACK L.  C/P:2,P
CHANUR’S HOMECOMING [CHANUR 4]  CHERRYH, C. J.  H:2,P
CHANUR’S LEGACY [CHANUR 5]  CHERRYH, C. J.  C/P,SFWA,P,H:2
CHANUR’S VENTURE [CHANUR 2]  CHERRYH, C. J.  H:2,P
KIF STRIKE BACK, THE [CHANUR 3]  CHERRYH, C. J.  C/P,P,H
PRIDE OF CHANUR, THE [CHANUR 1]  CHERRYH, C. J.  C/P,P,H
INFINITE SEA, THE [CHAOS CHRONICLES 3]  CARVER, JEFFREY A.  H,P
NEPTUNE CROSSING [CHAOS CHRONICLES 1]  CARVER, JEFFREY A.  C/P:2,P,H
STRANGE ATTRACTORS [CHAOS CHRONICLES 2]  CARVER, JEFFREY A.  C/P,P,H
AVENGER, THE [CHAOS GATE 3]  COOPER, LOUISE  C/P:2,P
DECEIVER, THE [CHAOS GATE 1]  COOPER, LOUISE  C/P,P
PRETENDER, THE [CHAOS GATE 2]  COOPER, LOUISE  P
BAD MEDICINE [CHARLIE MOON 1]  YARBRO, CHELSEA QUINN  P

TITLE  AUTHOR  CODE
CAT’S CLAW [CHARLIE MOON 4]  YARBRO, CHELSEA QUINN  P
FALSE NOTES [CHARLIE MOON 2]  YARBRO, CHELSEA QUINN  P
POISON FRUIT [CHARLIE MOON 3]  YARBRO, CHELSEA QUINN  P
AS PUCK WOULD HAVE IT [@CHARMED]  RUDITIS, PAUL  P
BEWARE WHAT YOU WISH [@CHARMED]  GALLAGHER, DIANA G.  P
BREWING STORM, THE [@CHARMED]  RUDITIS, PAUL  P
CHANGELING PLACES [@CHARMED]  OSTOW, MICOL  P
CHARMED AGAIN [@CHARMED]  LENHARD, ELIZABETH  P
CRIMSON SPELL, THE [@CHARMED]  GOLDSBOROUGH, F.  P
DARK VENGEANCE [@CHARMED]  GALLAGHER, DIANA G.  P
DATE WITH DEATH [@CHARMED]  LENHARD, ELIZABETH  P
DEMON DOPPELGANGERS [@CHARMED]  ELLIOT, GREG  P
GARDEN OF EVIL [@CHARMED]  HARRISON, EMMA  P
GYPSY ENCHANTMENT, THE [@CHARMED]  JABLONSKI, CARLA  P
HAUNTED BY DESIRE [@CHARMED]  DOKEY, CAMERON  P
HIGH SPIRTS [CHARMED]  CIENCIN, SCOTT  P
HOUSE OF SHARDS [@CHARMED]  OSTOW, MICOL  P
HURRICANE HEX [@CHARMED]  GALLAGHER, DIANA G.  P
INHERIT THE WITCH [@CHARMED]  BURNS, LAURA J.  P
KISS OF DARKNESS [@CHARMED]  ALEXANDER, BRANDON  P
LIGHT OF THE WORLD [@CHARMED]  CIENCIN, SCOTT  P
LUCK BE A LADY [@CHARMED]  CIENCIN, SCOTT  P
MIRROR IMAGE [@CHARMED]  MARIOTTE, JEFF  P
MIST AND STONE [@CHARMED]  GALLAGHER, DIANA G.  P
MYSTIC KNOLL [@CHARMED]  GALLAGHER, DIANA G.  P
PHOEBE WHO? [@CHARMED]  HARRISON, EMMA  P
PICTURE PERFECT [@CHARMED]  DOKEY, CAMERON  P
PIED PIPER [@CHARMED]  VIGUIE, DEBBIE  P
POWER OF THREE, THE [@CHARMED]  WILLARD, ELIZA  P
QUEEN’S CURSE, THE [@CHARMED]  HARRISON, EMMA  P
SHADOW OF THE SPHINX [@CHARMED]  JABLONSKI, CARLA  P
SOMETHING WICCAN THIS WAY COMES [@CHARMED]  HARRISON, EMMA  P
SOUL OF THE BRIDE [@CHARMED]  LENHARD, ELIZABETH  P
SPIRIT OF THE WOLF [@CHARMED]  GALLAGHER, DIANA G.  P
SURVIVAL OF THE FITTEST [@CHARMED]  MARIOTTE, JEFF  P
SWEET TALKIN’ DEMON [@CHARMED]  BURNS, LAURA J.  P
TALE OF TWO PIPERS, A [@CHARMED]  HARRISON, EMMA  P
TRUTH AND CONSEQUENCES [@CHARMED]  DOKEY, CAMERON  P
VOODOO MOON [@CHARMED]  STAUB, WENDY CORSI  P
WARREN WITCHES, THE [CHARMED]  SIMON SPOTLIGHT ENTERTAINMENT  LPA
WHISPERS FROM THE PAST [@CHARMED]  NOONAN, ROSILAND  P
ALPHA [CHARON 2]  ASARO, CATHERINE  H
SUNRISE ALLEY [CHARON 1]  ASARO, CATHERINE  H
DAUGHTER OF THE LION [CHEYSULI 6]  ROBERSON, JENNIFER  C/P,P
FLIGHT OF THE RAVEN [CHEYSULI 7]  ROBERSON, JENNIFER  C/P:2,P
LEGACY OF THE SWORD [CHEYSULI 3]  ROBERSON, JENNIFER  C/P:2,P
PRIDE OF PRINCES, A [CHEYSULI 5]  ROBERSON, JENNIFER  C/P:2,P
SHAPECHANGERS [CHEYSULI 1]  ROBERSON, JENNIFER  C/P:2,P
SONG OF HOMANA, THE [CHEYSULI 2]  ROBERSON, JENNIFER  C/P:2,P
TAPESTRY OF LIONS, A [CHEYSULI 8]  ROBERSON, JENNIFER  C/P,P
TRACK OF THE WHITE WOLF [CHEYSULI 4]  ROBERSON, JENNIFER  C/P,P
SHAPECHANGER’S SONG (INCLUDES "KINSPIRIT") [@CHEYSULI #1,2,3]  ROBERSON, JENNIFER  P
LEGACY OF THE WOLF [@CHEYSULI #3,4]  ROBERSON, JENNIFER  P
CHILDREN OF THE LION [@CHEYSULI #5,6]  ROBERSON, JENNIFER  P
LION THRONE, THE [@CHEYSULI #7,8]  ROBERSON, JENNIFER  P
MORTAL MASK [CHIA BLACK DRAGON 2]  MARLEY, STEPHEN  LP
CHICK IS IN THE MAIL, THE [CHICKS IN CHAINMAIL 5]  FRIESNER, ESTHER M.  PA,C/PA
CHICKS IN CHAINMAIL [CHICKS IN CHAINMAIL 1]  FRIESNER, ESTHER M.  PA
CHICKS ’N CHAINED MALES [CHICKS IN CHAINMAIL 3]  FRIESNER, ESTHER M.  HA,PA
DID YOU SAY CHICKS?! [CHICKS IN CHAINMAIL 2]  FRIESNER, ESTHER M.  C/PA,PA
MATHEMAGICS [CHICKS IN CHAINMAIL 4]  BALL, MARGARET  C/P,P
CHILD QUEEN, THE [CHILD QUEEN 1]  MCKENZIE, NANCY  P
HIGH QUEEN, THE [CHILD QUEEN 2]  MCKENZIE, NANCY  P
ANTAGONIST [CHILDE CYCLE]  DICKSON, GORDON R./WIXON, DAVID W.  H
CHANTRY GUILD, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,SFWA-P,P,H
DORSAI! [CHILDE CYCLE]  DICKSON, GORDON R.  C/P:2,P
DORSAI COMPANION, THE [CHILDE CYCLE]  DICKSON, GORDON R.  H,LP
FINAL ENCYCLOPEDIA, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H
GENETIC GENERAL, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,DD449,P
LOST DORSAI [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,LP,H
NECROMANCER [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H
NO ROOM FOR MAN [CHILDE CYCLE]  DICKSON, GORDON R.  C/P:2,P
OTHER [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H

TITLE  AUTHOR  CODE
SOLDIER, ASK NOT [CHILDE CYCLE]  DICKSON, GORDON R.  C/P:2,P
SPIRIT OF DORSAI, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,LP,H
TACTICS OF MISTAKE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H,SFWA
THREE TO DORSAI [CHILDE CYCLE]  DICKSON, GORDON R.  H:2
YOUNG BLEYS [CHILDE CYCLE]  DICKSON, GORDON R.  H:2,P
VENGEANCE OF MASKS [CHILDERIC THE SHATTERER 1]  EDGHILL, ROSEMARY  H
CHILDREN OF CHAOS [CHILDREN OF CHAOS 1]  DUNCAN, DAVE  H
MOTHER OF LIES [CHILDREN OF CHAOS 2]  DUNCAN, DAVE  H
EARTH IS ALL THAT LASTS, THE [CHILDREN OF EARTH 1]  WELLS, CATHERINE  P
EARTH SAVER, THE [CHILDREN OF EARTH 3]  WELLS, CATHERINE  P
BLACKGOD, THE [CHILDREN OF THE CHANGELING 2]  KEYES, J. GREGORY  H,P
WATERBORN, THE [CHILDREN OF THE CHANGELING 1]  KEYES, J. GREGORY  H,P
DARK DREAMS [CHILDREN OF THE NIGHT 1]  HODGMAN, ANN  P
DARK MUSIC [CHILDREN OF THE NIGHT 2]  HODGMAN, ANN  P
TOMORROW’S HERITAGE [@CHILDREN OF THE STARS #1]  COULSON, JUANITA  C/P,P,H
OUTWARD BOUND [@CHILDREN OF THE STARS #2]  COULSON, JUANITA  C/P,P,H
LEGACY OF EARTH [@CHILDREN OF THE STARS #3]  COULSON, JUANITA  C/P,P
PAST OF FOREVER, THE [@CHILDREN OF THE STARS #4]  COULSON, JUANITA  C/P,P
SATAN’S SEED [@CHILL #1]  SHERMAN, JORY  P
CHILL [@CHILL #2]  SHERMAN, JORY  P
BAMBOO DEMONS, THE [@CHILL #3]  SHERMAN, JORY  P
VEGAS VAMPIRE [@CHILL #4]  SHERMAN, JORY  P
PHOENIX MAN, THE [@CHILL #5]  SHERMAN, JORY  P
HOUSE OF SCORPIONS [@CHILL #6]  SHERMAN, JORY  P
SHADOWS [@CHILL #7]  SHERMAN, JORY  C/P,P
NINTH TALISMAN, THE [CHOSEN 2]  WATT-EVANS, LAWRENCE  H
WIZARD LORD, THE [CHOSEN 1]  WATT-EVANS, LAWRENCE  H
CHARMED LIFE [CHRESTOMANCI 1]  JONES, DIANA WYNNE  C/P:2,P
CHRESTOMANCI QUARTET, THE [CHRESTOMANCI 1,2,3,4]  JONES, DIANA WYNNE  H
CHRONICLES OF CHRESTOMANCI, VOLUME II [CHRESTOMANCI 3,2]  JONES, DIANA WYNNE  P
CONRAD’S FATE [CHRESTOMANCI 5]  JONES, DIANA WYNNE  H
LIVES OF CHRISTOPHER CHANT, THE [CHRESTOMANCI 4]  JONES, DIANA WYNNE  LP
MAGICIANS OF CAPRONA, THE [CHRESTOMANCI 3]  JONES, DIANA WYNNE  C/P,P
MIXED MAGICS [CHRESTOMANCI]  JONES, DIANA WYNNE  LP
PINHOE EGG, THE [CHRESTOMANCI]  JONES, DIANA WYNNE  H
WITCH WEEK [CHRESTOMANCI 2]  JONES, DIANA WYNNE  H,LP
BIRTH OF AN AGE (PROOF) [CHRIST CLONE 2]  BEAUSEIGNEUR, JAMES  SR-LP
IN HIS IMAGE [@CHRIST CLONE #1]  BEAUSEIGNEUR, JAMES  P
BIRTH OF AN AGE [@CHRIST CLONE #2]  BEAUSEIGNEUR, JAMES  P
ACTS OF GOD [@CHRIST CLONE #3]  BEAUSEIGNEUR, JAMES  H
KEY TO CHROMA [CHROMAGIC 2]  ANTHONY, PIERS  H
KEY TO DESTINY [CHROMAGIC 3]  ANTHONY, PIERS  H
KEY TO HAVOC [CHROMAGIC 1]  ANTHONY, PIERS  LP
KEY TO LIBERTY [CHROMAGIC 4]  ANTHONY, PIERS  H
CHRONICLES OF THE RAVEN 1 [CHRONICLES OF THE RAVEN 1,2]  BARCLAY, JAMES  H
DAWNTHIEF [CHRONICLES OF THE RAVEN 1]  BARCLAY, JAMES  C/P:2,P
NIGHTCHILD [CHRONICLES OF THE RAVEN 3]  BARCLAY, JAMES  LP
NOONSHADE [CHRONICLES OF THE RAVEN 2]  BARCLAY, JAMES  LP
DARK WATERS [CHRONICLES UNBINDING 2]  STONE, ROBERT S.  P
HAZARD’S PRICE [CHRONICLES UNBINDING 1]  STONE, ROBERT S.  P
CHRONO CRUSADE, VOL. 1 [CHRONO CRUSADE 1]  MORIYAMA, DAISUKE  SCX
EMPIRE OF TIME, THE [CHRONOPLANE WARS 3]  KILIAN, CRAWFORD  C/P,P
FALL OF THE REPUBLIC, THE [CHRONOPLANE WARS 1]  KILIAN, CRAWFORD  C/P,P
ROGUE EMPEROR [CHRONOPLANE WARS 2]  KILIAN, CRAWFORD  C/P,P
CHTHON [CHTHON #1]  ANTHONY, PIERS  C/P:2,@P
PHTHOR [CHTHON #2]  ANTHONY, PIERS  C/P:2,@P
MARRIAGE OF THE LIVING DARK, THE [@CHUNG KUO 8]  WINGROVE, DAVID  P
STONE WITHIN, THE [CHUNG KUO 4]  WINGROVE, DAVID  C/P:2,LP
MIDDLE KINGDOM, THE [@CHUNG KUO #1]  WINGROVE, DAVID  C/P,P,H
BROKEN WHEEL, THE [@CHUNG KUO #2]  WINGROVE, DAVID  C/P:2,P
WHITE MOUNTAIN, THE [@CHUNG KUO #3]  WINGROVE, DAVID  C/P,P,H
BENEATH THE TREE OF HEAVEN [@CHUNG KUO #5]  WINGROVE, DAVID  C/P:2,LP
WHITE MOON, RED DRAGON [@CHUNG KUO #6]  WINGROVE, DAVID  C/P:2,LP
DAYS OF BITTER STRENGTH [@CHUNG KUO #7]  WINGROVE, DAVID  C/P,LP,H
BRIDE OF THE SLIME MONSTER [CINEVERSE 2]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P:2,P
CINEVERSE CYCLE, THE [CINEVERSE 1,2,3]  GARDNER, CRAIG SHAW  H
REVENGE OF THE FLUFFY BUNNIES [CINEVERSE 3]  GARDNER, CRAIG SHAW  C/P,SFWA-P:2,P
SLAVES OF THE VOLCANO GOD [CINEVERSE 1]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P,P
CLOAK OF ILLUSION [CINGULUM 2]  ROBERTS, JOHN MADDOX  P,SFWA-P
SWORD, THE JEWEL AND THE MIRROR, THE [CINGULUM 3]  ROBERTS, JOHN MADDOX  P,SFWA-P
DANCE OF THE GODS [CIRCLE 2]  ROBERTS, NORA  H
MORRIGAN’S CROSS [CIRCLE 1]  ROBERTS, NORA  H
VALLEY OF SILENCE [CIRCLE 3]  ROBERTS, NORA  H

The MITSFS Seriesdex  September 24, 2007  Page 24
TITLE  AUTHOR  CODE
GREYFAX GRIMWALD [@CIRCLE OF LIGHT #1]  HANCOCK, NIEL  C/P:2,P
FARAGON FAIRINGAY [@CIRCLE OF LIGHT #2]  HANCOCK, NIEL  C/P:2,P
CALIX STAY [@CIRCLE OF LIGHT #3]  HANCOCK, NIEL  C/P,P
SQUARING THE CIRCLE [@CIRCLE OF LIGHT #4]  HANCOCK, NIEL  C/P:2,P
WILL OF THE EMPRESS, THE [CIRCLE OF MAGIC]  PIERCE, TAMORA  H
SANDRY’S BOOK [@CIRCLE OF MAGIC #1]  PIERCE, TAMORA  P
TRIS’S BOOK [@CIRCLE OF MAGIC #2]  PIERCE, TAMORA  P
DAJA’S BOOK [@CIRCLE OF MAGIC #3]  PIERCE, TAMORA  P
BRIAR’S BOOK [@CIRCLE OF MAGIC #4]  PIERCE, TAMORA  C/P,P
CIRCUIT [CIRCUIT 1]  SNODGRASS, MELINDA M.  C/P:2,P
CIRCUIT BREAKER [CIRCUIT 2]  SNODGRASS, MELINDA M.  P,SFWA-P
FINAL CIRCUIT [CIRCUIT 3]  SNODGRASS, MELINDA M.  P
CIRCUIT OF HEAVEN [CIRCUIT OF HEAVEN 1]  DANVERS, DENNIS  H,P
END OF DAYS [CIRCUIT OF HEAVEN 2]  DANVERS, DENNIS  C/P,P
CIRCUS WORLD [CIRCUS WORLD 3]  LONGYEAR, BARRY B.  C/P,P,H
CITY OF BARABOO [CIRCUS WORLD 1]  LONGYEAR, BARRY B.  C/P,SFWA,P,H
ELEPHANT SONG [CIRCUS WORLD 2]  LONGYEAR, BARRY B.  C/P:2,P
CITIES IN FLIGHT [CITIES IN FLIGHT 1,2,3,4]  BLISH, JAMES  C/P,H:2,P
CITIES IN FLIGHT, VOL. 1 [CITIES IN FLIGHT 1,2]  BLISH, JAMES  C/P:2,SFWA-P,P
CITIES IN FLIGHT, VOL. 2 [CITIES IN FLIGHT 3,4]  BLISH, JAMES  P
CLAIDI JOURNALS, THE [CLAIDI 1,2,3]  LEE, TANITH  H
EMPEROR OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
GODDESS OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
RED MEN OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
WARRIOR OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
BLOOD & RUST [CLEVELAND UNDEAD 1,2]  SWINIARSKI, S. A.  P
RAVEN [CLEVELAND UNDEAD 1]  SWINIARSKI, S. A.  P

FALL OF NESKAYA, THE [@CLINGFIRE #1/DARKOVER] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H,P

ZANDRU’S FORGE [@CLINGFIRE #2] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H

FLAME IN HALI, A [@CLINGFIRE #3] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H

ANVIL OF THE SUN [@CLOAK AND DAGGER #1]  GROELL, ANNE LESLEY  P
BRIDGE OF VALOR [@CLOAK AND DAGGER #2]  GROELL, ANNE LESLEY  P
CAULDRON OF INIQUITY [@CLOAK AND DAGGER #3]  GROELL, ANNE LESLEY  P
CLONE REPUBLIC, THE [CLONE REPUBLIC 1]  KENT, STEVEN L.  P
ROGUE CLONE [CLONE REPUBLIC 2]  KENT, STEVEN L.  P
HOLDER OF LIGHTNING [@CLOUDMAGES #1]  FARRELL, S. L.  H
HOLDER OF LIGHTNINGS (PROOF) [@CLOUDMAGES #1]  FARRELL, S. L.  SR-LP
MAGE OF CLOUDS [@CLOUDMAGES #2]  FARRELL, S. L.  P
HEIR OF STONE [@CLOUDMAGES #3]  FARRELL, S. L.  H
MAIA’S VEIL [CLOUDSHIPS OF ORION 2]  MCALLISTER, P. K.  P
ORION’S DAGGER [CLOUDSHIPS OF ORION 3]  MCALLISTER, P. K.  P
SIDURI’S NET [CLOUDSHIPS OF ORION 1]  MCALLISTER, P. K.  P
CLUSTER [CLUSTER #1]  ANTHONY, PIERS  C/P:2,@P
CHAINING THE LADY [CLUSTER #2]  ANTHONY, PIERS  C/P:2,@P
KIRLIAN QUEST [CLUSTER #3]  ANTHONY, PIERS  C/P:2,@P
THOUSANDSTAR [CLUSTER #4]  ANTHONY, PIERS  C/P:2,@P
VISCOUS CIRCLE [CLUSTER #5]  ANTHONY, PIERS  C/P:2,@P
LIVE ROBOTS [COBB ANDERSON 1,2]  RUCKER, RUDY  P
COBRA [COBRA 1]  ZAHN, TIMOTHY  C/P:2,P
COBRA BARGAIN [COBRA 3]  ZAHN, TIMOTHY  C/P:3,P
COBRA STRIKE [COBRA 2]  ZAHN, TIMOTHY  C/P:2,P
COBRA TRILOGY, THE [COBRA 1,2,3]  ZAHN, TIMOTHY  H
COBRAS TWO [COBRA 1,2]  ZAHN, TIMOTHY  C/P:2,SFWA-P,P
COCOON [COCOON 1]  SAPERSTEIN, DAVID  C/P,P
METAMORPHOSIS [COCOON 2]  SAPERSTEIN, DAVID  P
FURIES OF CALDERON [CODEX ALERA 1]  BUTCHER, JIM  H
ACADEM’S FURY [@CODEX ALERA #2]  BUTCHER, JIM  H
CURSOR’S FURY [@CODEX ALERA #3]  BUTCHER, JIM  H:2
COLD CASH WAR, THE [COLD CASH WAR 1]  ASPRIN, ROBERT LYNN  C/P,SFWA-P:2,P,H
COLD CASH WARRIOR [COLD CASH WAR 2]  ASPRIN, ROBERT LYNN/FAWCETT, BILL  P
BLACK SUN RISING [COLDFIRE 1]  FRIEDMAN, C. S.  H:2,P
CROWN OF SHADOWS [COLDFIRE 3]  FRIEDMAN, C. S.  H,P
WHEN TRUE NIGHT FALLS [COLDFIRE 2]  FRIEDMAN, C. S.  H:2,P
ARMAGEDDON INHERITANCE, THE [COLIN MACINTYRE 2]  WEBER, DAVID  C/P,SFWA-P,P
EMPIRE FROM THE ASHES [COLIN MACINTYRE 1,2,3]  WEBER, DAVID  H,SR-LP
HEIRS OF EMPIRE [COLIN MACINTYRE 3]  WEBER, DAVID  C/P:2,P
MUTINEERS’ MOON [COLIN MACINTYRE 1]  WEBER, DAVID  C/P:2,SFWA-P,P

METAL MAN AND OTHERS, THE [@COLLECTED STORIES JACK WILLIAMSON #1] WILLIAMSON, JACK

H

WOLVES OF DARKNESS [@COLLECTED STORIES JACK WILLIAMSON #2]  WILLIAMSON, JACK  H

The MITSFS Seriesdex  September 24, 2007  Page 25
TITLE  AUTHOR  CODE
WIZARD’S ISLE [@COLLECTED STORIES JACK WILLIAMSON #3]  WILLIAMSON, JACK  H
SPIDER ISLAND [@COLLECTED STORIES JACK WILLIAMSON #4]  WILLIAMSON, JACK  H
CRUCIBLE OF POWER, THE [@COLLECTED STORIES JACK WILLIAMSON #5]  WILLIAMSON, JACK  H
MINORITY REPORT, THE [COLLECTED STORIES OF PHILIP K. DICK 4]  DICK, PHILIP K.  LP
COLLEGE OF MAGICS, A [COLLEGE OF MAGICS 1]  STEVERMER, CAROLINE  H
SCHOLAR OF MAGICS, A [COLLEGE OF MAGICS 2]  STEVERMER, CAROLINE  H
COLOSSUS [COLOSSUS 1]  JONES, D. F.  H:2,P
COLOSSUS AND THE CRAB [COLOSSUS 3]  JONES, D. F.  C/P,P
FALL OF COLOSSUS, THE [COLOSSUS 2]  JONES, D. F.  H:2,P
CAVES OF KLYDOR, THE [COLSEC 2]  HILL, DOUGLAS  C/P,P,H
COLSEC REBELLION [COLSEC 3]  HILL, DOUGLAS  H,P
EXILES OF COLSEC [COLSEC 1]  HILL, DOUGLAS  C/P,P,H

DORSAI’S COMMAND [@COMBAT COMMAND CHILDE CYCLE] DICKSON, GORDON R./DENNING, TROY/

GLABERSON, CORY

P

TIME FUTURE [COMMANDER HALLEY 1]  MCARTHUR, MAXINE  P
TIME PAST [COMMANDER HALLEY 2]  MCARTHUR, MAXINE  P
JUDAS UNCHAINED [COMMONWEALTH 2]  HAMILTON, PETER F.  H
PANDORA’S STAR [COMMONWEALTH 1]  HAMILTON, PETER F.  H
BLACK PROJECTS, WHITE KNIGHTS [COMPANY]  BAKER, KAGE  H
CHILDREN OF THE COMPANY, THE [COMPANY 6]  BAKER, KAGE  H:2
GODS AND PAWNS [COMPANY 8]  BAKER, KAGE  H
GRAVEYARD GAME, THE [COMPANY 4]  BAKER, KAGE  LP
IN BAD COMPANY [COMPANY 3,4]  BAKER, KAGE  H
IN THE GARDEN OF IDEN [COMPANY 1]  BAKER, KAGE  C/P:2,P
LIFE OF THE WORLD TO COME, THE [COMPANY 5]  BAKER, KAGE  H
MACHINE’S CHILD, THE [COMPANY 7]  BAKER, KAGE  H
MENDOZA IN HOLLYWOOD [COMPANY 3]  BAKER, KAGE  P
ON COMPANY TIME [COMPANY 1,2]  BAKER, KAGE  H
RUDE MECHANICALS [COMPANY]  BAKER, KAGE  SR-H
SKY COYOTE [COMPANY 2]  BAKER, KAGE  P
SONS OF HEAVEN, THE [COMPANY 9]  BAKER, KAGE  H
BLACK HEART, IVORY BONES [COMPLEMENTARY FAIRY COLORS 6]  DATLOW, ELLEN/WINDLING, TERRI  LPA
BLACK SWAN, WHITE RAVEN [COMPLEMENTARY FAIRY COLORS 4]  DATLOW, ELLEN/WINDLING, TERRI  HA
BLACK THORN, WHITE ROSE [COMPLEMENTARY FAIRY COLORS 2]  DATLOW, ELLEN/WINDLING, TERRI  C/PA,PA
RUBY SLIPPERS, GOLDEN TEARS [COMPLEMENTARY FAIRY COLORS 3]  DATLOW, ELLEN/WINDLING, TERRI  HA,PA
SILVER BIRCH, BLOOD MOON [COMPLEMENTARY FAIRY COLORS 5]  DATLOW, ELLEN/WINDLING, TERRI  HA,LPA
SNOW WHITE, BLOOD RED [COMPLEMENTARY FAIRY COLORS 1]  DATLOW, ELLEN/WINDLING, TERRI  SR-HA,PA,C/PA,SFWA
QUEST BEGINS, THE [COMPLETE ELFQUEST]  PINI, WENDY/PINI, RICHARD  LP
FIRE AND FLIGHT [@COMPLETE ELFQUEST #1]  PINI, WENDY/PINI, RICHARD  CX
FORBIDDEN GROVE, THE [@COMPLETE ELFQUEST #2]  PINI, WENDY/PINI, RICHARD  CX
CAPTIVES OF BLUE MOUNTAIN [@COMPLETE ELFQUEST #3]  PINI, WENDY/PINI, RICHARD  CX
QUEST’S END [@COMPLETE ELFQUEST #4]  PINI, WENDY/PINI, RICHARD  CX
SIEGE AT BLUE MOUNTAIN [@COMPLETE ELFQUEST #5]  PINI, WENDY/PINI, RICHARD  CX
SECRET OF TWO-EDGE, THE [@COMPLETE ELFQUEST #6]  PINI, WENDY/PINI, RICHARD  CX
AND NOW THE NEWS... [COMPLETE STORIES OF THEODORE STURGEON 9] STURGEON, THEODORE  H
BABY IS THREE [COMPLETE STORIES OF THEODORE STURGEON 6]  STURGEON, THEODORE  H
KILLDOZER! [COMPLETE STORIES OF THEODORE STURGEON 3]  STURGEON, THEODORE  H

MAN WHO LOST THE SEA, THE [COMPLETE STORIES OF THEODORE STURGEON 10] STURGEON, THEODORE

H

MICROCOSMIC GOD [@COMPLETE STORIES OF THEODORE STURGEON 2]  STURGEON, THEODORE  LP

NAIL AND THE ORACLE, THE [COMPLETE STORIES OF THEODORE STURGEON 11] STURGEON, THEODORE

H

PERFECT HOST, THE [COMPLETE STORIES OF THEODORE STURGEON 5]  STURGEON, THEODORE  H

SAUCER OF LONELINESS, A [COMPLETE STORIES OF THEODORE STURGEON 7] STURGEON, THEODORE

H

THUNDER AND ROSES [COMPLETE STORIES OF THEODORE STURGEON 4]  STURGEON, THEODORE  H

ULTIMATE EGOIST, THE [@COMPLETE STORIES OF THEODORE STURGEON 1] STURGEON, THEODORE

LP

HERE COMES CIVILIZATION [COMPLETE TENN 2]  TENN, WILLIAM  H
IMMODEST PROPOSALS [COMPLETE TENN 1]  TENN, WILLIAM  H
BLADE OF CONAN, THE [CONAN]  DE CAMP, L. SPRAGUE  R/REF-P,C/REF-P
BLOODY CROWN OF CONAN, THE [CONAN]  HOWARD, ROBERT E.  H
COMING OF CONAN THE CIMMERIAN, THE [CONAN]  HOWARD, ROBERT E.  H,LP

CONAN [CONAN 1] HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/

CARTER, LIN

C/P:2,P

CONAN AND THE AMAZON [CONAN]  ROBERTS, JOHN MADDOX  P
CONAN AND THE DEATH LORD OF THANZA [CONAN]  GREEN, ROLAND J.  P
CONAN AND THE EMERALD LOTUS [CONAN 38]  HOCKING, JOHN C.  LP
CONAN AND THE GODS OF THE MOUNTAIN [CONAN]  GREEN, ROLAND J.  P
CONAN AND THE GRIM GREY GOD [CONAN]  MOORE, SEAN A.  LP
CONAN AND THE MANHUNTERS [CONAN]  ROBERTS, JOHN MADDOX  P
CONAN AND THE MISTS OF DOOM [@CONAN]  GREEN, ROLAND J.  P,SFWA-P
CONAN AND THE SHAMAN’S CURSE [CONAN]  MOORE, SEAN A.  P

The MITSFS Seriesdex  September 24, 2007  Page 26
TITLE  AUTHOR  CODE
CONAN AND THE SORCERER [CONAN]  OFFUTT, ANDREW J.  C/P,LP
CONAN AND THE TREASURE OF PYTHON [CONAN]  ROBERTS, JOHN MADDOX  LP
CONAN AT THE DEMON’S GATE [CONAN]  GREEN, ROLAND J.  LP
CONAN OF AQUILONIA [CONAN 11]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P

CONAN OF CIMMERIA [CONAN 2] HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/

CARTER, LIN

C/P:2,P

CONAN OF THE ISLES [CONAN 12]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN OF THE RED BROTHERHOOD [CONAN]  CARPENTER, LEONARD  P
CONAN OF VENARIUM [CONAN]  TURTLEDOVE, HARRY  P
CONAN READER, THE [CONAN]  DE CAMP, L. SPRAGUE  R/REF-H
CONAN, SCOURGE OF THE BLOODY COAST [CONAN]  CARPENTER, LEONARD  P
CONAN THE ADVENTURER [CONAN 5]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P

CONAN THE AVENGER [CONAN 10] HOWARD, ROBERT E./NYBERG, BJORN/

DE CAMP, L. SPRAGUE

C/P:2,P

CONAN THE BARBARIAN [CONAN]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN THE BOLD [CONAN]  ROBERTS, JOHN MADDOX  P,SFWA-P
CONAN THE BUCCANEER [CONAN 6]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN THE CHAMPION [CONAN]  ROBERTS, JOHN MADDOX  P
CONAN THE CONQUEROR [CONAN 9]  HOWARD, ROBERT E.  C/P:2,SR-DD036,P
CONAN THE DEFENDER [CONAN]  JORDAN, ROBERT  C/P,SFWA-P,LP
CONAN THE DEFIANT [CONAN]  PERRY, STEVE  SFWA-P,LP
CONAN THE DESTROYER [CONAN]  JORDAN, ROBERT  C/P:2,P
CONAN THE FEARLESS [CONAN]  PERRY, STEVE  H,SFWA-P
CONAN: THE FLAME KNIFE [CONAN]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE P
CONAN THE FORMIDABLE [CONAN] PERRY, STEVE C/P,SFWA-P,LP,P
CONAN THE FREEBOOTER [CONAN 3]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P
CONAN THE FREELANCE [CONAN]  PERRY, STEVE  P,SFWA-P
CONAN THE GLADIATOR [CONAN]  CARPENTER, LEONARD  P
CONAN THE GREAT [CONAN]  CARPENTER, LEONARD  P,SFWA-P
CONAN THE GUARDIAN [CONAN]  GREEN, ROLAND J.  P,SFWA-P
CONAN THE HERO [CONAN]  CARPENTER, LEONARD  C/P,P
CONAN THE INDOMITABLE [CONAN]  PERRY, STEVE  C/P,SFWA,SFWA-P,LP
CONAN THE INVINCIBLE [CONAN]  JORDAN, ROBERT  C/P:2,P
CONAN THE MAGNIFICENT [CONAN]  JORDAN, ROBERT  C/P:2,P
CONAN THE MARAUDER [CONAN]  ROBERTS, JOHN MADDOX  P,SFWA-P
CONAN THE MERCENARY [CONAN]  OFFUTT, ANDREW J.  P
CONAN THE OUTCAST [CONAN]  CARPENTER, LEONARD  P
CONAN THE RAIDER [CONAN]  CARPENTER, LEONARD  P,SFWA,SFWA-P,LP
CONAN THE RENEGADE [CONAN]  CARPENTER, LEONARD  P,SFWA-P
CONAN THE ROGUE [CONAN]  ROBERTS, JOHN MADDOX  H,LP
CONAN THE SAVAGE [CONAN]  CARPENTER, LEONARD  P
CONAN THE TRIUMPHANT [CONAN]  JORDAN, ROBERT  C/P:2,SFWA-P,P
CONAN THE UNCONQUERED [CONAN]  JORDAN, ROBERT  C/P:2,SFWA-P,P
CONAN THE USURPER [CONAN 8]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P
CONAN THE VALIANT [CONAN]  GREEN, ROLAND J.  P,SFWA-P:2
CONAN THE VALOROUS [CONAN]  ROBERTS, JOHN MADDOX  SFWA,LP
CONAN THE VICTORIOUS [CONAN]  JORDAN, ROBERT  C/P,SFWA,SFWA-P:2,P,H

CONAN THE WANDERER [CONAN 4] HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/

CARTER, LIN

C/P,P

CONAN THE WARLORD [CONAN]  CARPENTER, LEONARD  P,SFWA-P
CONAN THE WARRIOR [CONAN 7]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P
CONQUERING SWORD OF CONAN, THE [CONAN]  HOWARD, ROBERT E.  LP
LORD OF THE BLACK RIVER [CONAN]  CARPENTER, LEONARD  P
EXILE’S RETURN [CONCLAVE OF SHADOWS 3]  FEIST, RAYMOND E.  H
KING OF FOXES [CONCLAVE OF SHADOWS 2]  FEIST, RAYMOND E.  H
KING OF FOXES (PROOF) [CONCLAVE OF SHADOWS 2]  FEIST, RAYMOND E.  SR-LP
TALON OF THE SILVER HAWK [CONCLAVE OF SHADOWS 1]  FEIST, RAYMOND E.  P
EMPIRE’S LEGACY, THE [@CONCORDAT #1]  SWYCAFFER, JEFFERSON P.  C/P,P
VOYAGE OF THE PLANETSLAYER [@CONCORDAT #2]  SWYCAFFER, JEFFERSON P.  C/P:2,P
REVOLT AND REBIRTH [@CONCORDAT #3]  SWYCAFFER, JEFFERSON P.  P
BACKBLAST [CONEK HAYDEN]  MCKEONE, LEE  P
STARFIRE DOWN [CONEK HAYDEN]  MCKEONE, LEE  P
HOW FEW REMAIN [CONFEDERATE STATES]  TURTLEDOVE, HARRY  C/P:2,P
BLOOD AND IRON [CONFEDERATE STATES AMERICAN EMPIRE 1]  TURTLEDOVE, HARRY  C/P,P,H

CENTER CANNOT HOLD, THE [CONFEDERATE STATES AMERICAN EMPIRE 2] TURTLEDOVE, HARRY

H

VICTORIOUS OPPOSITION, THE [CONFEDERATE STATES AMERICAN EMPIRE 3] TURTLEDOVE, HARRY

H

AMERICAN FRONT [CONFEDERATE STATES: GREAT WAR 1]  TURTLEDOVE, HARRY  C/P,P
BREAKTHROUGHS [CONFEDERATE STATES: GREAT WAR 3]  TURTLEDOVE, HARRY  H,P
WALK IN HELL [CONFEDERATE STATES: GREAT WAR 2]  TURTLEDOVE, HARRY  C/P,P,H
CONFLUENCE [CONFLUENCE 1,2,3]  MCAULEY, PAUL J.  H
CHILD OF THE RIVER [@CONFLUENCE #1]  MCAULEY, PAUL J.  H,P

The MITSFS Seriesdex  September 24, 2007  Page 27
TITLE  AUTHOR  CODE
ANCIENT OF DAYS [@CONFLUENCE #2]  MCAULEY, PAUL J.  C/P,P,H
SHRINE OF STARS [@CONFLUENCE #3]  MCAULEY, PAUL J.  H
UNSHAPELY THINGS [CONNOR GREY 1]  DEL FRANCO, MARK  P
CONQUERORS’ HERITAGE [CONQUERORS 2]  ZAHN, TIMOTHY  C/P,P,H
CONQUERORS’ LEGACY [CONQUERORS 3]  ZAHN, TIMOTHY  H,P
CONQUERORS’ PRIDE [CONQUERORS 1]  ZAHN, TIMOTHY  C/P,P,H
CONRAD’S QUEST FOR RUBBER [CONRAD STARGARD 6]  FRANKOWSKI, LEO  C/P,P
CONRAD’S TIME MACHINE [CONRAD STARGARD 0]  FRANKOWSKI, LEO  H
CROSS-TIME ENGINEER, THE [CONRAD STARGARD 1]  FRANKOWSKI, LEO  C/P:2,P
FLYING WARLORD, THE [CONRAD STARGARD 4]  FRANKOWSKI, LEO  C/P:2,P
HIGH-TECH KNIGHT, THE [CONRAD STARGARD 2]  FRANKOWSKI, LEO  C/P:2,P
LORD CONRAD’S LADY [CONRAD STARGARD 5]  FRANKOWSKI, LEO  C/P:2,P
RADIANT WARRIOR, THE [CONRAD STARGARD 3]  FRANKOWSKI, LEO  C/P:2,P
EMERALD EYES [CONTINUING TIME 1]  MORAN, DANIEL KEYS  P
LAST DANCER, THE [CONTINUING TIME 3]  MORAN, DANIEL KEYS  C/P,P
LONG RUN, THE [CONTINUING TIME 2]  MORAN, DANIEL KEYS  C/P,P
CONTRARYWISE [CONTRARYWISE 1]  GREENHALGH, ZOHRA  P
TRICKSTER’S TOUCH [CONTRARYWISE 2]  GREENHALGH, ZOHRA  C/P,P
DOOMFARERS OF CORAMONDE, THE [CORAMONDE 1]  DALEY, BRIAN  C/P:2,P
STARFOLLOWERS OF CORAMONDE, THE [CORAMONDE 2]  DALEY, BRIAN  C/P:2,P
ALECTOR’S CHOICE [COREAN 4]  MODESITT, L. E. JR.  H
CADMIAN’S CHOICE [COREAN 5]  MODESITT, L. E. JR.  H
DARKNESSES [COREAN 2]  MODESITT, L. E. JR.  P
LEGACIES [COREAN 1]  MODESITT, L. E. JR.  H
LEGACIES (PROOF) [COREAN 1]  MODESITT, L. E. JR.  SR-LP
SCEPTERS [COREAN 3]  MODESITT, L. E. JR.  H
SOARER’S CHOICE [COREAN 6]  MODESITT, L. E. JR.  H
MISTS OF DOOM, THE [CORMAC 1]  OFFUTT, ANDREW J.  C/P,P
SIGN OF THE MOONBOW, THE [CORMAC 6]  OFFUTT, ANDREW J.  P
SWORD OF THE GAEL [CORMAC 4]  OFFUTT, ANDREW J.  C/P:2,P
TOWER OF DEATH, THE [CORMAC 2]  OFFUTT, ANDREW J./TAYLOR, KEITH  P
WHEN DEATH BIRDS FLY [CORMAC 3]  OFFUTT, ANDREW J./TAYLOR, KEITH  C/P,P
CORNELIUS CHRONICLES, THE [CORNELIUS 1,2,3,4]  MOORCOCK, MICHAEL  C/P,P
CORNELIUS CHRONICLES II, THE [CORNELIUS 5,6]  MOORCOCK, MICHAEL  P
CORNELIUS CHRONICLES III, THE [CORNELIUS 7,8]  MOORCOCK, MICHAEL  P
CURE FOR CANCER, A [CORNELIUS 2]  MOORCOCK, MICHAEL  H
ENGLISH ASSASSIN, THE [CORNELIUS 3]  MOORCOCK, MICHAEL  H
FINAL PROGRAMME, THE [CORNELIUS 1]  MOORCOCK, MICHAEL  C/P:2,P

LIVES AND TIMES OF JERRY CORNELIUS, THE (1979 EDITION)

[CORNELIUS 5] MOORCOCK, MICHAEL

P

LIVES AND TIMES OF JERRY CORNELIUS, THE (2003 EDITION)

[CORNELIUS 5] MOORCOCK, MICHAEL

LP

NEW NATURE OF THE CATASTROPHE, THE (OMNIBUS EDITION)

[CORNELIUS] JONES, LANGDON/MOORCOCK, MICHAEL

PA

BOOK OF ULTIMATE TRUTHS, THE [CORNELIUS MURPHY 1]  RANKIN, ROBERT  P
MOST AMAZING MAN WHO EVER LIVED, THE [CORNELIUS MURPHY 3]  RANKIN, ROBERT  P
RAIDERS OF THE LOST CAR PARK [CORNELIUS MURPHY 2]  RANKIN, ROBERT  P
BLAMELESS IN ABADDON [CORPUS DEI 2]  MORROW, JAMES  H
ETERNAL FOOTMAN, THE [CORPUS DEI 3]  MORROW, JAMES  H
TOWING JEHOVAH [CORPUS DEI 1]  MORROW, JAMES  H:2,LP
BULL AND THE SPEAR, THE [CORUM 4]  MOORCOCK, MICHAEL  C/P:2,P
CHRONICLES OF CORUM, THE [CORUM 1,2,3]  MOORCOCK, MICHAEL  P
KING OF THE SWORDS, THE [CORUM 3]  MOORCOCK, MICHAEL  C/P:2,P
KNIGHT OF THE SWORDS, THE [CORUM 1]  MOORCOCK, MICHAEL  C/P:2,P
OAK AND THE RAM, THE [CORUM 5]  MOORCOCK, MICHAEL  C/P,P,H
QUEEN OF THE SWORDS, THE [CORUM 2]  MOORCOCK, MICHAEL  C/P:2,P
SWORD AND THE STALLION, THE [CORUM 6]  MOORCOCK, MICHAEL  C/P,P,H
ALCHEMY UNLIMITED [CORWYN THE ALCHEMIST 1]  CLARK, DOUGLAS W.  C/P,P
WHIRLWIND ALCHEMY [CORWYN THE ALCHEMIST 2]  CLARK, DOUGLAS W.  C/P,SFWA-P,P
COYOTE [COYOTE 1]  STEELE, ALLEN  H
COYOTE FRONTIER [COYOTE 3]  STEELE, ALLEN  H
COYOTE RISING [COYOTE 2]  STEELE, ALLEN  H
AT THE SEVENTH LEVEL [COYOTE JONES 3]  ELGIN, SUZETTE HADEN  C/P,P
COMMUNIPATH WORLDS [COYOTE JONES 1,2,3]  ELGIN, SUZETTE HADEN  C/P:2,P
COMMUNIPATHS, THE [COYOTE JONES 1]  ELGIN, SUZETTE HADEN  D11560
FURTHEST [COYOTE JONES 2]  ELGIN, SUZETTE HADEN  C/P:2,P
STAR-ANCHORED, STAR-ANGERED [COYOTE JONES 4]  ELGIN, SUZETTE HADEN  H:2,SFWA,P
YONDER COMES THE OTHER END OF TIME [COYOTE JONES 5]  ELGIN, SUZETTE HADEN  C/P:2,P
CRABS’ MOON [CRABS 2]  SMITH, GUY N.  P
CRABS ON THE RAMPAGE [CRABS 3]  SMITH, GUY N.  C/P,P
NIGHT OF THE CRABS [CRABS]  SMITH, GUY N.  P
ORIGIN OF THE CRABS, THE [CRABS 1]  SMITH, GUY N.  P
ANGUISHED DAWN, THE [CRADLE OF SATURN 2]  HOGAN, JAMES P.  P

TITLE  AUTHOR  CODE
CRADLE OF SATURN [CRADLE OF SATURN 1]  HOGAN, JAMES P.  H:2,P
BLESSINGS AND CURSES [@CRAFTERS #2]  STASHEFF, CHRISTOPHER/FAWCETT, BILL  C/PA,PA,SFWA-P:2
FALLOUT [CRAIG KREIDENT 2]  ANDERSON, KEVIN J./BEASON, DOUG  P
LETHAL EXPOSURE [CRAIG KREIDENT 3]  ANDERSON, KEVIN J./BEASON, DOUG  C/P,P
VIRTUAL DESTRUCTION [CRAIG KREIDENT 1]  ANDERSON, KEVIN J./BEASON, DOUG  C/P,P
CLIPJOINT [CRASHCOURSE 2]  BAIRD, WILHELMINA  P
CRASHCOURSE [CRASHCOURSE 1]  BAIRD, WILHELMINA  C/P:2,SFWA-P,P
CRYSTAL PALACE, THE [CRAY ORMORU 2]  EISENSTEIN, PHYLLIS  C/P:2,P
SORCERER’S SON [CRAY ORMORU 1]  EISENSTEIN, PHYLLIS  C/P:2,P
SORCERER’S SON (PROOF) [CRAY ORMORU 1]  EISENSTEIN, PHYLLIS  SR-LP
CRETACEOUS SEA [CRETACEOUS SEA 1]  HUBBELL, WILL  P
SEA OF TIME [CRETACEOUS SEA 2]  HUBBELL, WILL  P
RIDDLE OF THE GRIFFON [@CRIMSON CRYSTAL 1]  LAWSON, SUSAN  P
CRIMSON SHADOW, THE [CRIMSON SHADOW 1,2,3]  SALVATORE, R. A.  LP
DRAGON KING, THE [CRIMSON SHADOW 3]  SALVATORE, R. A.  P
LUTHIEN’S GAMBLE [CRIMSON SHADOW 2]  SALVATORE, R. A.  P
SWORD OF BEDWYR, THE [CRIMSON SHADOW 1]  SALVATORE, R. A.  P
CLUSTER COMMAND [CRISIS OF EMPIRE 2]  DIETZ, WILLIAM C./DRAKE, DAVID  P,SFWA-P
CROWN OF EMPIRE [CRISIS OF EMPIRE 4]  YARBRO, CHELSEA QUINN  P
HONORABLE DEFENSE, AN [CRISIS OF EMPIRE 1]  DRAKE, DAVID/THOMAS, THOMAS T.  C/P:2,SFWA-P,P
WAR MACHINE, THE [CRISIS OF EMPIRE 3]  ALLEN, ROGER MACBRIDE/DRAKE, DAVID  P,SFWA-P:2
CROSSFIRE [CROSSFIRE 1]  KRESS, NANCY  H
CRUCIBLE [CROSSFIRE 2]  KRESS, NANCY  H
SPIRIT GATE [CROSSROADS 1]  ELLIOTT, KATE  H
CROSSROADS OF TIME, THE [CROSSTIME]  NORTON, ANDRE  C/P:2,DD164,P
QUEST CROSSTIME [CROSSTIME]  NORTON, ANDRE  H:2,P
GUNPOWDER EMPIRE [CROSSTIME TRAFFIC 1]  TURTLEDOVE, HARRY  H
IN HIGH PLACES [CROSSTIME TRAFFIC 3]  TURTLEDOVE, HARRY  H
DISUNITED STATES OF AMERICA, THE [@CROSSTIME TRAFFIC #4]  TURTLEDOVE, HARRY  H
CLASH BY NIGHT [CROW]  WILLIAMSON, CHET  LP
HELLBOUND [@CROW]  ATTANASIO, A. A.  H,LP
LAZARUS HEART, THE [@CROW]  BRITE, POPPY Z.  LP
TEMPLE OF NIGHT [@CROW]  SOMTOW, S. P.  P
BURNING STONE, THE [CROWN OF STARS 3]  ELLIOTT, KATE  C/P,P,H
CHILD OF FLAME [CROWN OF STARS 4]  ELLIOTT, KATE  H:2
CROWN OF STARS [CROWN OF STARS 7]  ELLIOTT, KATE  H
GATHERING STORM, THE [CROWN OF STARS 5]  ELLIOTT, KATE  H
IN THE RUINS [CROWN OF STARS 6]  ELLIOTT, KATE  H
KING’S DRAGON [CROWN OF STARS 1]  ELLIOTT, KATE  C/P,P,H
PRINCE OF DOGS [CROWN OF STARS 2]  ELLIOTT, KATE  C/P,P,H
FORTRESS OF GLASS, THE [CROWN OF THE ISLES 1]  DRAKE, DAVID  H
IRON TREE, THE [CROWTHISTLE 1]  DART-THORNTON, CECILIA  H,P
NAMELESS DAY, THE [CRUCIBLE 1]  DOUGLASS, SARA  P
WOUNDED HAWK, THE [@CRUCIBLE 2]  DOUGLASS, SARA  H
CRYSTAL LINE [CRYSTAL SINGER 3]  MCCAFFREY, ANNE  H:2,P
CRYSTAL SINGER [CRYSTAL SINGER 1]  MCCAFFREY, ANNE  C/P:2,P,H:3
KILLASHANDRA [CRYSTAL SINGER 2]  MCCAFFREY, ANNE  C/P,P,H,SFWA
SEA HAG, THE [CRYSTAL WALLS 1]  DRAKE, DAVID  C/P,SFWA-P,P
CRYSTAL SORCERORS, THE [CRYSTAL WARRIORS 2]  FORSTCHEN, WILLIAM R./MORRISON, GREG P
CRYSTAL WARRIORS, THE [CRYSTAL WARRIORS 1]  FORSTCHEN, WILLIAM R./MORRISON, GREG C/P:2,P
STILL, SMALL VOICE OF TRUMPETS, THE [CULTURAL SURVEY 1]  BIGGLE, LLOYD JR.  C/P,P,H:2
WORLD MENDERS, THE [CULTURAL SURVEY 2]  BIGGLE, LLOYD JR.  C/P:2,SFWA,P
CONSIDER PHLEBAS [CULTURE]  BANKS, IAIN M.  H,P
EXCESSION [CULTURE]  BANKS, IAIN M.  C/P,P,H
LOOK TO WINDWARD [CULTURE]  BANKS, IAIN M.  H
PLAYER OF GAMES, THE [CULTURE]  BANKS, IAIN M.  C/P:3,P,H
STATE OF THE ART, THE [CULTURE]  BANKS, IAIN M.  H,P
USE OF WEAPONS [CULTURE]  BANKS, IAIN M.  H:2,P
ANOTHER DAY, ANOTHER DUNGEON [CUPS AND SORCERY 1]  COSTIKYAN, GREG  P,SFWA-P
ONE QUEST, HOLD THE DRAGONS [CUPS AND SORCERY 2]  COSTIKYAN, GREG  P
BAVARIAN GATE, THE [CURTIS MACURDY 2]  DALMAS, JOHN  P
LION OF FARSIDE, THE [CURTIS MACURDY 1]  DALMAS, JOHN  C/P,P
LION RETURNS, THE [CURTIS MACURDY 3]  DALMAS, JOHN  C/P,P
CV [CV 1]  KNIGHT, DAMON  C/P,SFWA-P,P,H
OBSERVERS, THE [CV 2]  KNIGHT, DAMON  H,SFWA-P
REASONABLE WORLD, A [CV 3]  KNIGHT, DAMON  H:2,SFWA,P
CAPO’S REVENGE [@CYBERNARC #4]  CAIN, ROBERT  P
CYBERNETIC SAMURAI, THE [CYBERNETIC SAMURAI 1]  MILAN, VICTOR  H:2,SFWA,P
CYBERNETIC SHOGUN, THE [CYBERNETIC SAMURAI 2]  MILAN, VICTOR  H
HOLO MEN [CYBERPUNK 2.0.2.0.]  BILLIAS, STEPHEN  P
RAVENGERS, THE [CYBERPUNK 2.0.2.0.]  BILLIAS, STEPHEN  P
CYBORG [@CYBORG #1]  CAIDIN, MARTIN  C/P:2,P
OPERATION NUKE [@CYBORG #2]  CAIDIN, MARTIN  P

TITLE  AUTHOR  CODE
PLANET IN PERIL [@CYBORG COMMANDO #1]  O’NEILL, PAMELA/MOHAN, KIM  C/P:2,P
CHASE INTO SPACE [@CYBORG COMMANDO #2]  O’NEILL, PAMELA/MOHAN, KIM  P
ULTIMATE PRIZE, THE [@CYBORG COMMANDO #3]  O’NEILL, PAMELA/MOHAN, KIM  P
CYCLE OF FIRE, THE [CYCLE OF FIRE 1,2,3]  WURTS, JANNY  LP
KEEPER OF THE KEYS [CYCLE OF FIRE 2]  WURTS, JANNY  C/P:2,P
SHADOWFANE [CYCLE OF FIRE 3]  WURTS, JANNY  C/P:2,P
STORMWARDEN [CYCLE OF FIRE 1]  WURTS, JANNY  C/P:2,SFWA-P,P
CYGNET AND THE FIREBIRD, THE [CYGNET 2]  MCKILLIP, PATRICIA A.  SFWA
SORCERESS AND THE CYGNET, THE [CYGNET 1]  MCKILLIP, PATRICIA A.  C/P:2,SFWA,SFWA-P:2,P

TITLE  AUTHOR  CODE
BALANCE OF POWER [DAEDALUS 5]  STABLEFORD, BRIAN M.  C/P,P
CITY OF THE SUN [DAEDALUS 4]  STABLEFORD, BRIAN M.  C/P,P
CRITICAL THRESHOLD [DAEDALUS 2]  STABLEFORD, BRIAN M.  C/P:2,P
FLORIANS, THE [DAEDALUS 1]  STABLEFORD, BRIAN M.  C/P,P
PARADOX OF THE SETS, THE [DAEDALUS 6]  STABLEFORD, BRIAN M.  C/P:2,P
WILDEBLOOD’S EMPIRE [DAEDALUS 3]  STABLEFORD, BRIAN M.  C/P,P
CART AND CWIDDER [DALEMARK 1]  JONES, DIANA WYNNE  C/P:2,LP
CROWN OF DALEMARK, THE [DALEMARK 4]  JONES, DIANA WYNNE  C/P,P
DROWNED AMMET [DALEMARK 2]  JONES, DIANA WYNNE  C/P,P
SPELLCOATS, THE [DALEMARK 3]  JONES, DIANA WYNNE  C/P:2,P
DAWN STAG, THE [DALRIADA 2]  WATSON, JULES  H
WHITE MARE, THE [DALRIADA 1]  WATSON, JULES  LP
DEMON KING, THE [DAMASTES 2]  BUNCH, CHRIS  LP
SEER KING, THE [DAMASTES 1]  BUNCH, CHRIS  C/P,P
WARRIOR KING, THE [DAMASTES 3]  BUNCH, CHRIS  P
DAMIANO [DAMIANO 1]  MACAVOY, R. A.  C/P:2,P
DAMIANO’S LUTE [DAMIANO 2]  MACAVOY, R. A.  C/P:2,P
RAPHAEL [DAMIANO 3]  MACAVOY, R. A.  C/P:2,P
TRIO FOR LUTE, A [DAMIANO 1,2,3]  MACAVOY, R. A.  H:2,P
CALL TO ARMS, A [@DAMNED #1]  FOSTER, ALAN DEAN  C/P,P,H
FALSE MIRROR, THE [@DAMNED #2]  FOSTER, ALAN DEAN  C/P,P,H
SPOILS OF WAR, THE [@DAMNED #3]  FOSTER, ALAN DEAN  C/P:2,P
CAULDRON OF PLENTY, THE [DANANS 2]  TAYLOR, KEITH  C/P,P
SEARCH FOR THE STARBLADE [DANANS 3]  TAYLOR, KEITH  P
SORCERER’S SACRED ISLE, THE [DANANS 1]  TAYLOR, KEITH  C/P,P
CATASTROPHE’S SPELL [DANCE OF GODS 1]  BRENNER, MAYER ALAN  C/P,P
SPELL OF APOCALYPSE [DANCE OF GODS 4]  BRENNER, MAYER ALAN  P
SPELL OF FATE [DANCE OF GODS 3]  BRENNER, MAYER ALAN  P
SPELL OF INTRIGUE [DANCE OF GODS 2]  BRENNER, MAYER ALAN  P
RING OF DESTINY [DANCE OF THE RINGS 3]  FANCHER, JANE S.  P
RING OF INTRIGUE [DANCE OF THE RINGS 2]  FANCHER, JANE S.  C/P,P
RING OF LIGHTNING [DANCE OF THE RINGS 1]  FANCHER, JANE S.  P
DANCER’S RISE [@DANCER #1]  CLAYTON, JO  C/P:2,P
SERPENT WALTZ [@DANCER #2]  CLAYTON, JO  C/P:2,P
DANCE DOWN THE STARS [@DANCER #3]  CLAYTON, JO  C/P:2,P
ALIEN HEAT, AN [DANCERS AT THE END OF TIME 1]  MOORCOCK, MICHAEL  H:2,P,SFWA
END OF ALL SONGS, THE [DANCERS AT THE END OF TIME 3]  MOORCOCK, MICHAEL  H:2,P,SFWA
HOLLOW LANDS, THE [DANCERS AT THE END OF TIME 2]  MOORCOCK, MICHAEL  H:2,SFWA,P
LEGENDS FROM THE END OF TIME [DANCERS AT THE END OF TIME 4]  MOORCOCK, MICHAEL  H,SFWA,P
MESSIAH AT THE END OF TIME, A [DANCERS AT THE END OF TIME 5]  MOORCOCK, MICHAEL  C/P:2,P
DEMONS OF THE DANCING GODS [DANCING GODS 2]  CHALKER, JACK L.  C/P:2,P
HORRORS OF THE DANCING GODS [DANCING GODS 5]  CHALKER, JACK L.  C/P:2,P
RIVER OF DANCING GODS, THE [DANCING GODS 1]  CHALKER, JACK L.  C/P:2,P
SONGS OF THE DANCING GODS [DANCING GODS 4]  CHALKER, JACK L.  C/P:2,P
VENGEANCE OF THE DANCING GODS [DANCING GODS 3]  CHALKER, JACK L.  C/P:2,P
ANUBIS MURDERS, THE [@DANGEROUS JOURNEYS 1]  GYGAX, GARY  C/P,SFWA-P,P
DEATH IN DELHI [@DANGEROUS JOURNEYS 3]  GYGAX, GARY  P
SAMARKAND SOLUTION, THE [@DANGEROUS JOURNEYS 2]  GYGAX, GARY  C/P,P
VAMPRIRE ODYSSEY, THE [DANIELLE WALTHERS 1]  CIENCIN, SCOTT  P
WILDINGS, THE [DANIELLE WALTHERS 2]  CIENCIN, SCOTT  P
PRISONER OF REGLATHIUM, THE [@DANNUS #1]  SIROTA, MIKE  P
CONQUERORS OF REGLATHIUM, THE [@DANNUS #2]  SIROTA, MIKE  P
CAVES OF REGLATHIUM, THE [@DANNUS #3]  SIROTA, MIKE  C/P,P
DARK STRAITS OF REGLATHIUM, THE [@DANNUS #4]  SIROTA, MIKE  P
SLAVES OF REGLATHIUM, THE [@DANNUS #5]  SIROTA, MIKE  P
DANNY DUNN AND THE ANTI-GRAVITY PAINT [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P,H
DANNY DUNN AND THE AUTOMATIC HOUSE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN AND THE FOSSIL CAVE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P,H
DANNY DUNN AND THE HEAT RAY [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE HOMEWORK MACHINE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE SMALLIFYING MACHINE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE SWAMP MONSTER [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN AND THE UNIVERSAL GLUE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN AND THE VOICE FROM SPACE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE WEATHER MACHINE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P,H
DANNY DUNN, INVISIBLE BOY [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN ON A DESERT ISLAND [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN ON THE OCEAN FLOOR [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN, SCIENTIFIC DETECTIVE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN, TIME TRAVELER [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DEAD MAN RISING [DANTE VALENTINE 2]  SAINTCROW, LILITH  P
WORKING FOR THE DEVIL [DANTE VALENTINE 1]  SAINTCROW, LILITH  P
PREDATOR’S SMILE [@DAREDEVIL]  GOLDEN, CHRISTOPHER  P

TITLE  AUTHOR  CODE
OWLFLIGHT [DARIAN 1]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
OWLKNIGHT [DARIAN 3]  LACKEY, MERCEDES/DIXON, LARRY  C/P:2,P,H
OWLSIGHT [DARIAN 2]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
AFTER THE DARK [@DARK ANGEL]  COLLINS, MAX ALLAN  P
BEFORE THE DAWN [@DARK ANGEL]  COLLINS, MAX ALLAN  C/P,P
SKIN GAME [@DARK ANGEL]  COLLINS, MAX ALLAN  C/P,P
GATHERING OF HEROES, A [@DARK BORDER]  ZIMMER, PAUL EDWIN  C/P,P
INGULF THE MAD [DARK BORDER]  ZIMMER, PAUL EDWIN  C/P,SFWA-P,P
LOST PRINCE, THE [@DARK BORDER #1]  ZIMMER, PAUL EDWIN  C/P:2,P
KING CHONDOS’ RIDE [@DARK BORDER #2]  ZIMMER, PAUL EDWIN  C/P:2,P
BEGINNING, THE [@DARK CHRONICLES #1]  SOROKA, CYNTHIA  P
EVIL ASCENDING [DARK CONSPIRACY 2]  STACKPOLE, MICHAEL A.  P
EVIL TRIUMPHANT [DARK CONSPIRACY 3]  STACKPOLE, MICHAEL A.  P
GATHERING EVIL, A [DARK CONSPIRACY 1]  STACKPOLE, MICHAEL A.  P
COLOR OF EVIL, THE [DARK DESCENT 1]  HARTWELL, DAVID G.  PA
MEDUSA IN THE SHIELD, THE [DARK DESCENT 2]  HARTWELL, DAVID G.  PA
HEART OF GLASS [DARK ENCHANTMENT]  COOPER, LOUISE  P
ASHTON HORROR, THE [@DARK FORCES 12]  BRIDGES, LAURIE  P
BARGAIN, THE [DARK FORCES 5]  SPARGER, REX  C/P,P
DEVIL WIND [@DARK FORCES 4]  BRIDGES, LAURIE/ALEXANDER, PAUL  P
DOLL, THE [DARK FORCES 3]  SPARGER, REX  P
EYES OF THE TAROT [DARK FORCES 9]  COVILLE, BRUCE  P
GAME, THE [DARK FORCES 1]  LOGAN, LES  P
MAGIC SHOW [@DARK FORCES 2]  BRIDGES, LAURIE/ALEXANDER, PAUL  P
UNNATURAL TALENT [DARK FORCES 7]  LOGAN, LES  P
COMEBACK TOUR [DARK FUTURE]  YEOVIL, JACK  LP
KROKODIL TEARS [DARK FUTURE]  YEOVIL, JACK  LP
ROUTE 666 [DARK FUTURE]  PRINGLE, DAVID  LP
AMERICAN MEAT [@DARK FUTURE AMERICAN MEAT 1]  MOORE, STUART  P
DEMON DOWNLOAD [DARK FUTURE DEMON DOWNLOAD 1]  YEOVIL, JACK  LP
ROUTE 666 [DARK FUTURE DEMON DOWNLOAD 2]  YEOVIL, JACK  P
GOLGOTHA RUN [@DARK FUTURE EDDIE KALISH 1]  STONE, DAVE  P
CITY OF THE SORCERERS [DARK HORSE 3]  HERBERT, MARY H.  P
DARK HORSE [DARK HORSE 1]  HERBERT, MARY H.  P
LIGHTNING’S DAUGHTER [DARK HORSE 2]  HERBERT, MARY H.  P
VALORIAN [DARK HORSE 0]  HERBERT, MARY H.  P
VALORIAN’S CHILDREN [DARK HORSE 1,2]  HERBERT, MARY H.  H
VALORIAN’S LEGACY [DARK HORSE 3,4]  HERBERT, MARY H.  H
WINGED MAGIC [DARK HORSE 4]  HERBERT, MARY H.  P
DANCE WITH THE DEVIL [DARK-HUNTER]  KENYON, SHERRILYN  P
SINS OF THE NIGHT [@DARK-HUNTER]  KENYON, SHERRILYN  P
UNLEASH THE NIGHT [@DARK-HUNTER]  KENYON, SHERRILYN  P
DARK IS RISING, THE [DARK IS RISING 2]  COOPER, SUSAN  C/P,LP,H:2
GREENWITCH [DARK IS RISING 3]  COOPER, SUSAN  H:3,LP
GREY KING, THE [DARK IS RISING 4]  COOPER, SUSAN  C/P,LP,H:2
OVER SEA, UNDER STONE [DARK IS RISING 1]  COOPER, SUSAN  C/P,SR-P,H:2
SILVER ON THE TREE [DARK IS RISING 5]  COOPER, SUSAN  C/P,P,H:2
IF WHISPERS CALL [DARK MATTER 2]  BASSINGTHWAITE, DON  P
IN FLUID SILENCE [DARK MATTER 3]  TIRPA, G. W.  P
IN HOLLOW HOUSES [DARK MATTER 1]  BRAUNBECK, GARY A.  P
OF AGED ANGELS [DARK MATTER 4]  COOK, MONTE  P
CURSE, THE [@DARK MOON #1]  BLAIR, CYNTHIA  P
BRAZEN GAMBIT, THE [@DARK SUN ATHAS #1]  ABBEY, LYNN  P
DARKNESS BEFORE THE DAWN, THE [DARK SUN ATHAS #2]  HUGHES, RYAN  P
BROKEN BLADE, THE [@DARK SUN ATHAS #3]  HAWKE, SIMON  P
CINNABAR SHADOWS [@DARK SUN ATHAS #4]  ABBEY, LYNN  P
RISE AND FALL OF A DRAGON KING, THE [@DARK SUN ATHAS #5]  ABBEY, LYNN  P
VERDANT PASSAGE, THE [@DARK SUN PRISM #1]  DENNING, TROY  C/P,P
CRIMSON LEGION, THE [@DARK SUN PRISM #2]  DENNING, TROY  C/P,P
AMBER ENCHANTRESS, THE [@DARK SUN PRISM #3]  DENNING, TROY  C/P,P
OBSIDIAN ORACLE, THE [@DARK SUN PRISM #4]  DENNING, TROY  P
CERULEAN STORM, THE [@DARK SUN PRISM #5]  DENNING, TROY  C/P,P
OUTCAST, THE [@DARK SUN TRIBE OF ONE #1]  HAWKE, SIMON  C/P,P
SEEKER, THE [@DARK SUN TRIBE OF ONE #2]  HAWKE, SIMON  P
NOMAD, THE [@DARK SUN TRIBE OF ONE #3]  HAWKE, SIMON  P
DRAWING OF THE THREE, THE [DARK TOWER 2]  KING, STEPHEN  C/P:2,@H:2,P
GUNSLINGER, THE [DARK TOWER 1]  KING, STEPHEN  C/P:2,SR-L,LP
GUNSLINGER, THE (REVISED) [DARK TOWER 1]  KING, STEPHEN  H

STEPHEN KING’S THE DARK TOWER A CONCORDANCE VOLUME I [DARK TOWER] FURTH, ROBIN

R/REF-P

WASTE LANDS, THE [@DARK TOWER #3]  KING, STEPHEN  C/P,SFWA,LP,H
WIZARD AND GLASS [@DARK TOWER #4]  KING, STEPHEN  C/P,LP,H
WOLVES OF THE CALLA [@DARK TOWER #5]  KING, STEPHEN  H:2

The MITSFS Seriesdex  September 24, 2007  Page 33
TITLE  AUTHOR  CODE
SONG OF SUSANNAH [@DARK TOWER #6]  KING, STEPHEN  H
DARK TOWER, THE [@DARK TOWER #7]  KING, STEPHEN  H:2
DARKANGEL, THE [DARKANGEL 1]  PIERCE, MEREDITH ANN  C/P,P
DARKANGEL TRILOGY, THE [DARKANGEL 1,2,3]  PIERCE, MEREDITH ANN  H
GATHERING OF GARGOYLES, A [DARKANGEL 2]  PIERCE, MEREDITH ANN  P,SFWA-P
PEARL OF THE SOUL OF THE WORLD, THE [DARKANGEL 3]  PIERCE, MEREDITH ANN  P
SERPENT BRIDE, THE [DARKGLASS MOUNTAIN 1]  DOUGLASS, SARA  H
SERPENT BRIDE, THE (PROOF) [DARKGLASS MOUNTAIN 1]  DOUGLASS, SARA  SR-LP
HANGMAN, THE [@DARKMAN #1]  BOYLL, RANDALL  P
PRICE OF FEAR, THE [@DARKMAN #2]  BOYLL, RANDALL  P
IN THE FACE OF DEATH [@DARKMAN #4]  BOYLL, RANDALL  P

ALTON GIFT, THE [DARKOVER] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H

BLOODY SUN, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
BLOODY SUN, THE (EXPANDED) [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
CHILDREN OF HASTUR [DARKOVER]  BRADLEY, MARION ZIMMER  H:2
CITY OF SORCERY [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
DARKOVER LANDFALL [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
DOMAINS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
EXILE’S SONG [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
FORBIDDEN TOWER, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
FOUR MOONS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
FREE AMAZONS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  PA,C/PA
HAWKMISTRESS! [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
HEIRS OF HAMMERFELL, THE [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
HERITAGE OF HASTUR, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
KEEPER’S PRICE, THE [DARKOVER]  BRADLEY, MARION ZIMMER  PA,C/PA
LERONI OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
MARION ZIMMER BRADLEY’S DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
OATH OF THE RENUNCIATES [DARKOVER]  BRADLEY, MARION ZIMMER  H
OTHER SIDE OF THE MIRROR, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
PLANET SAVERS, THE [DARKOVER]  BRADLEY, MARION ZIMMER  DF153

PLANET SAVERS, THE [DARKOVER]/ SWORD OF ALDONES, THE [DARKOVER] BRADLEY, MARION ZIMMER

C/P,P

PLANET SAVERS, THE (INCLUDES THE WATERFALL) [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P
RED SUN OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA

REDISCOVERY [DARKOVER] BRADLEY, MARION ZIMMER/ LACKEY, MERCEDES

H:2,P

RENUNCIATES OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
SAGA OF THE RENUNCIATES, THE [DARKOVER]  BRADLEY, MARION ZIMMER  P
SHADOW MATRIX, THE [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
SHARRA’S EXILE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
SHATTERED CHAIN, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
SNOWS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
SPELL SWORD, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
STAR OF DANGER [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
STORMQUEEN! [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
SWORD OF ALDONES, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,DF153,P
SWORD OF CHAOS [DARKOVER]  BRADLEY, MARION ZIMMER  PA,C/PA
THENDARA HOUSE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
TO SAVE A WORLD [DARKOVER]  BRADLEY, MARION ZIMMER  P
TOWERS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
TRAITOR’S SUN [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
TWO TO CONQUER [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
WINDS OF DARKOVER, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H,D89250
WORLD WRECKERS, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
LEGACY OF THE DARK SWORD [DARKSWORD]  WEIS, MARGARET/HICKMAN, TRACY  H
FORGING THE DARKSWORD [@DARKSWORD #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DOOM OF THE DARKSWORD [@DARKSWORD #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
TRIUMPH OF THE DARKSWORD [@DARKSWORD #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
FLIGHT OF THE NIGHTHAWKS [DARKWAR 1]  FEIST, RAYMOND E.  H
INTO A DARK REALM [DARKWAR 2]  FEIST, RAYMOND E.  H
DOOMSTALKER [@DARKWAR #1]  COOK, GLEN  P
WARLOCK [@DARKWAR #2]  COOK, GLEN  C/P,P
CEREMONY [@DARKWAR #3]  COOK, GLEN  P
ARMIES OF DAYLIGHT, THE [DARWATH 3]  HAMBLY, BARBARA  C/P:2,P
ICEFALCON’S QUEST [DARWATH 5]  HAMBLY, BARBARA  C/P:2,P,H
MOTHER OF WINTER [DARWATH 4]  HAMBLY, BARBARA  H:2,P
TIME OF THE DARK, THE [DARWATH 1]  HAMBLY, BARBARA  C/P:2,P
WALLS OF AIR, THE [DARWATH 2]  HAMBLY, BARBARA  C/P:2,P
TRICKSTERS [DAUGHTER OF THE LIONESS 1,2]  PIERCE, TAMORA  H
HIDDEN LAND, THE [DAUGHTERS OF BAST 1]  ISIDORE, SARAH  P
SHRINE OF LIGHT [DAUGHTERS OF BAST 2]  ISIDORE, SARAH  P

TITLE  AUTHOR  CODE
WORLD TREE, THE [DAUGHTERS OF BAST 3]  ISIDORE, SARAH  P
ANGEL OF PAIN, THE [DAVID LYDYARD 2]  STABLEFORD, BRIAN M.  H
CARNIVAL OF DESTRUCTION, THE [DAVID LYDYARD 3]  STABLEFORD, BRIAN M.  P
WEREWOLVES OF LONDON, THE [DAVID LYDYARD 1]  STABLEFORD, BRIAN M.  C/P,P,H
DAYWORLD [DAYWORLD 1]  FARMER, PHILIP JOSE  C/P,P,H
DAYWORLD BREAKUP [DAYWORLD 3]  FARMER, PHILIP JOSE  H:2,SFWA-P,P
DAYWORLD REBEL [DAYWORLD 2]  FARMER, PHILIP JOSE  C/P,SFWA-P,P,H
DEATH AND LIFE OF SUPERMAN, THE [DC COMICS]  STERN, ROGER  C/P,P,H
FURTHER ADVENTURES OF BATMAN, THE [DC COMICS]  GREENBERG, MARTIN H.  C/PA,PA
FURTHER ADVENTURES OF BATMAN 2, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF BATMAN 3, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF SUPERMAN, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF THE JOKER, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF WONDER WOMAN, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
SUPERMAN: DOOMSDAY & BEYOND [DC COMICS]  SIMONSON, LOUISE  LP
SUPERMAN III [DC COMICS]  KOTZWINKLE, WILLIAM  C/P,P
SUPERMAN: LAST SON OF KRYPTON [DC COMICS]  MAGGIN, ELLIOT S.  C/P:2,P
SUPERMAN: MIRACLE MONDAY [DC COMICS]  MAGGIN, ELLIOT S.  P
SUPERMAN: THE MAN OF STEEL [DC COMICS]  HELFER, ANDREW  P
DEAD GIRLS [DEAD 1]  CALDER, RICHARD  P
DEAD GIRLS DEAD BOYS DEAD THINGS [DEAD 1,2,3]  CALDER, RICHARD  LP
FREEDOM’S APPRENTICE [DEAD RIVERS 2]  KRITZER, NAOMI  P
FREEDOM’S GATE [DEAD RIVERS 1]  KRITZER, NAOMI  C/P,P
LORDS OF DESTRUCTION [DEATH DEALER 2]  SILKE, JAMES  P,SFWA-P
PLAGUE OF KNIVES [DEATH DEALER 4]  SILKE, JAMES  C/P,SFWA-P,P
PRISONER OF THE HORNED HELMET [DEATH DEALER 1]  SILKE, JAMES  C/P:2,SFWA-P,P
TOOTH AND CLAW [DEATH DEALER 3]  SILKE, JAMES  C/P,SFWA-P,P
HAND OF CHAOS, THE (PROOF) [DEATH GATE 5]  WEIS, MARGARET/HICKMAN, TRACY  SR-LP
DRAGON WING [@DEATH GATE #1]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
ELVEN STAR [@DEATH GATE #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P,P,H
FIRE SEA [@DEATH GATE #3]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
SERPENT MAGE [@DEATH GATE #4]  WEIS, MARGARET/HICKMAN, TRACY  C/P,P,H
HAND OF CHAOS, THE [@DEATH GATE #5]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
INTO THE LABYRINTH [@DEATH GATE #6]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
SEVENTH GATE, THE [@DEATH GATE #7]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
DEATHDAY [DEATHDAY 1]  DIETZ, WILLIAM C.  C/P,P
EARTHRISE [DEATHDAY 2]  DIETZ, WILLIAM C.  P
BITTER FRUIT [@DEATHLANDS]  AXLER, JAMES  P
BLACK HARVEST [@DEATHLANDS]  AXLER, JAMES  P
COLD ASYLUM [@DEATHLANDS]  AXLER, JAMES  P
CRATER LAKE [@DEATHLANDS]  AXLER, JAMES  P
CROSSWAYS [@DEATHLANDS]  AXLER, JAMES  P
CRUCIBLE OF TIME [@DEATHLANDS]  AXLER, JAMES  C/P,P
DARK EMBLEM [@DEATHLANDS]  AXLER, JAMES  P
DECTRA CHAIN [@DEATHLANDS]  AXLER, JAMES  P
DEMONS OF EDEN [@DEATHLANDS]  AXLER, JAMES  P
ECLIPSE AT NOON [@DEATHLANDS]  AXLER, JAMES  P
EMERALD FIRE [@DEATHLANDS]  AXLER, JAMES  P
ENCOUNTER [@DEATHLANDS]  AXLER, JAMES  P
FREEDOM LOST [@DEATHLANDS]  AXLER, JAMES  P
GROUND ZERO [@DEATHLANDS]  AXLER, JAMES  P
HOMEWARD BOUND [@DEATHLANDS]  AXLER, JAMES  P
ICE AND FIRE [@DEATHLANDS]  AXLER, JAMES  P
KEEPERS OF THE SUN [@DEATHLANDS]  AXLER, JAMES  P
LATITUDE ZERO [@DEATHLANDS]  AXLER, JAMES  P
MARS ARENA, THE [@DEATHLANDS]  AXLER, JAMES  C/P,P
MOON FATE [@DEATHLANDS]  AXLER, JAMES  P
NEUTRON SOLSTICE [@DEATHLANDS]  AXLER, JAMES  P
NIGHTMARE PASSAGE [@DEATHLANDS]  AXLER, JAMES  P
NORTHSTAR RISING [DEATHLANDS]  AXLER, JAMES  P
PILGRIMAGE TO HELL [@DEATHLANDS]  ADRIAN, JACK  C/P,P
PONY SOLDIERS [@DEATHLANDS]  AXLER, JAMES  P
RAT KING [@DEATHLANDS]  AXLER, JAMES  P
RED HOLOCAUST [@DEATHLANDS]  AXLER, JAMES  P
ROAD WARS [@DEATHLANDS]  AXLER, JAMES  P
SHADOW WORLD [@DEATHLANDS]  AXLER, JAMES  P
SHADOWFALL [@DEATHLANDS]  AXLER, JAMES  P
SHAKING EARTH [@DEATHLANDS]  AXLER, JAMES  P
SKYDARK SPAWN [@DEATHLANDS]  AXLER, JAMES  P
STONEFACE [@DEATHLANDS]  AXLER, JAMES  P
TIME NOMADS [@DEATHLANDS]  AXLER, JAMES  P
TRADER REDUX [@DEATHLANDS]  AXLER, JAMES  P
TWILIGHT CHILDREN [@DEATHLANDS]  AXLER, JAMES  P

The MITSFS Seriesdex  September 24, 2007  Page 35
TITLE  AUTHOR  CODE
WATERSLEEP [@DEATHLANDS]  AXLER, JAMES  P
WAY OF THE WOLF [@DEATHLANDS]  AXLER, JAMES  P
ZERO CITY [@DEATHLANDS]  AXLER, JAMES  P
LAST MORTAL MAN, THE [DEATHLESS 1]  MITCHELL, SYNE  P
DEATHSTALKER [DEATHSTALKER 1]  GREEN, SIMON R.  C/P:2,P
DEATHSTALKER CODA [DEATHSTALKER]  GREEN, SIMON R.  H
DEATHSTALKER DESTINY [DEATHSTALKER 5]  GREEN, SIMON R.  C/P,P
DEATHSTALKER HONOR [DEATHSTALKER 4]  GREEN, SIMON R.  P
DEATHSTALKER LEGACY [DEATHSTALKER]  GREEN, SIMON R.  H
DEATHSTALKER REBELLION [DEATHSTALKER 2]  GREEN, SIMON R.  C/P,P
DEATHSTALKER RETURN [DEATHSTALKER]  GREEN, SIMON R.  P
DEATHSTALKER WAR [DEATHSTALKER 3]  GREEN, SIMON R.  C/P,P
TWILIGHT OF THE EMPIRE [DEATHSTALKER]  GREEN, SIMON R.  C/P,P
RITES OF THE DEMON [@DEATHWALKER #1]  CASTEVANO, ROMAN  P
DECOY PRINCESS, THE [DECOY PRINCESS 1]  COOK, DAWN  C/P,P
PRINCESS AT SEA [DECOY PRINCESS 2]  COOK, DAWN  C/P,P
DEEPNESS IN THE SKY, A [DEEP 2]  VINGE, VERNOR  C/P,P,H
FIRE UPON THE DEEP, A [DEEP 1]  VINGE, VERNOR  H:2,P
SCAR NIGHT [DEEPGATE 1]  CAMPBELL, ALAN  H
DELTA GREEN: DENIED TO THE ENEMY [DELTA GREEN]  DETWILLER, DENNIS  LP
DEMON BLADE [DEMON BLADE 1]  GARLAND, MARK A./MCGRAW, CHARLES G. C/P,P
FROST [DEMON BLADE 2]  GARLAND, MARK A./MCGRAW, CHARLES G. P
DRAGON’S TONGUE [DEMON-BOUND 1]  UNDERWOOD, LAURA J.  LP
GLASS WARRIOR, THE [@DEMON CROWN #1]  VARDEMAN, ROBERT E.  P,SFWA-P
PHANTOMS ON THE WIND [@DEMON CROWN #2]  VARDEMAN, ROBERT E.  P,SFWA-P
SYMPHONY OF STORMS, A [@DEMON CROWN #3]  VARDEMAN, ROBERT E.  C/P:2,SFWA-P,P
BOOK OF DREAMS, THE [DEMON PRINCES 5]  VANCE, JACK  C/P:2,P
DEMON PRINCES, THE [DEMON PRINCES 1,2,3,4,5]  VANCE, JACK  H
FACE, THE [DEMON PRINCES 4]  VANCE, JACK  C/P:2,P
KILLING MACHINE, THE [DEMON PRINCES 2]  VANCE, JACK  C/P:2,P
PALACE OF LOVE, THE [DEMON PRINCES 3]  VANCE, JACK  C/P:2,P
STAR KING, THE [DEMON PRINCES 1]  VANCE, JACK  C/P:2,P
COURTS OF SORCERY, THE [DEMON WARS 3]  MCCONNELL, ASHLEY  P
FOUNTAINS OF MIRLACCA, THE [DEMON WARS 1]  MCCONNELL, ASHLEY  P
ITINERANT EXORCIST, THE [DEMON WARS 2]  MCCONNELL, ASHLEY  P
GULF RUN [@DEMONTECH 3]  SHERMAN, DAVID  P
ONSLAUGHT [@DEMONTECH 1]  SHERMAN, DAVID  P
RALLY POINT [@DEMONTECH 2]  SHERMAN, DAVID  P
DEMON APOSTLE, THE [DEMONWARS 3]  SALVATORE, R. A.  C/P,P,H
DEMON AWAKENS, THE [DEMONWARS 1]  SALVATORE, R. A.  C/P,SR-LP,H
DEMON SPIRIT, THE [DEMONWARS 2]  SALVATORE, R. A.  C/P,P
MORTALIS [DEMONWARS 4]  SALVATORE, R. A.  H
ASCENDANCE [DEMONWARS II 1]  SALVATORE, R. A.  H,P
IMMORTALIS [DEMONWARS II 3]  SALVATORE, R. A.  H
TRANSCENDENCE [DEMONWARS II 2]  SALVATORE, R. A.  H
CAGE A MAN [DEMU 1]  BUSBY, F. M.  H:2,P
DEMU TRILOGY, THE [DEMU 1,2,3]  BUSBY, F. M.  C/P:2,P
PROUD ENEMY, THE [DEMU 2]  BUSBY, F. M.  C/P:2,P
DEXTA [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS 1]  RYAN, C. J.  P
GLORIOUS TREASON [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS 2]  RYAN, C. J.  P

FIFTH QUADRANT, THE [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS #2] RYAN, C. J.

P

ALCHYMIST’S CAT, THE [@DEPTFORD HISTORIES #2]  JARVIS, ROBIN  LP
OAKEN THRONE, THE [@DEPTFORD HISTORIES #2]  JARVIS, ROBIN  LP
CRYSTAL PRISON, THE [DEPTFORD MICE 2]  JARVIS, ROBIN  LP
DARK LORD OF DERKHOLM [DERKHOLM 1]  JONES, DIANA WYNNE  H
YEAR OF THE GRIFFIN [DERKHOLM 2]  JONES, DIANA WYNNE  P
CHILDE MORGAN [DERYNI]  KURTZ, KATHERINE  H
CODEX DERYNIANUS II [DERYNI]  KURTZ, KATHERINE/REGINALD, ROBERT  R/REF-H
DERYNI ARCHIVES, THE [DERYNI]  KURTZ, KATHERINE  C/P,P,H
DERYNI CHALLENGE [DERYNI]  BILLIAS, STEPHEN  P,SFWA-P
DERYNI MAGIC [DERYNI]  KURTZ, KATHERINE  C/P:2,P
DERYNI TALES [DERYNI]  KURTZ, KATHERINE  C/PA,PA
IN THE KING’S SERVICE [@DERYNI]  KURTZ, KATHERINE  P
CAMBER OF CULDI [DERYNI CAMBER 1]  KURTZ, KATHERINE  C/P,P,H
CAMBER THE HERETIC [DERYNI CAMBER 3]  KURTZ, KATHERINE  C/P:2,P,H
SAINT CAMBER [DERYNI CAMBER 2]  KURTZ, KATHERINE  C/P,P,H
CHRONICLES OF THE DERYNI, THE [DERYNI CHRONICLES 1,2,3]  KURTZ, KATHERINE  H
DERYNI CHECKMATE [DERYNI CHRONICLES 2]  KURTZ, KATHERINE  C/P:3,P
DERYNI RISING [DERYNI CHRONICLES 1]  KURTZ, KATHERINE  C/P:3,P
HIGH DERYNI [DERYNI CHRONICLES 3]  KURTZ, KATHERINE  C/P:2,SR-P,P
BASTARD PRINCE, THE [DERYNI HEIRS 3]  KURTZ, KATHERINE  H:2,P
HARROWING OF GWYNEDD, THE [DERYNI HEIRS 1]  KURTZ, KATHERINE  C/P,P,H

TITLE  AUTHOR  CODE
KING JAVAN’S YEAR [DERYNI HEIRS 2]  KURTZ, KATHERINE  H
BISHOP’S HEIR, THE [DERYNI KELSON 1]  KURTZ, KATHERINE  C/P,P,H,SFWA
KING KELSON’S BRIDE [DERYNI KELSON 4]  KURTZ, KATHERINE  H
KING’S JUSTICE, THE [DERYNI KELSON 2]  KURTZ, KATHERINE  H,P,SFWA
QUEST FOR SAINT CAMBER, THE [DERYNI KELSON 3]  KURTZ, KATHERINE  C/P,H,P
DESCENT [@DESCENT 1]  TELEP, PETER  P
EQUINOX [@DESCENT 3]  TELEP, PETER  P
STEALING THUNDER [@DESCENT 2]  TELEP, PETER  P
MARCH INTO DARKNESS, A [DESTINIES OF BLOOD AND STONE 2]  NEWCOMB, ROBERT  H
SAVAGE MESSIAH [DESTINIES OF BLOOD AND STONE 1]  NEWCOMB, ROBERT  H
WITH FATE CONSPIRE [@DESTINY MAKERS 1]  SHUPP, MIKE  C/P:2,P
MORNING OF CREATION [@DESTINY MAKERS #2]  SHUPP, MIKE  C/P:2,P
SOLDIER OF ANOTHER FORTUNE [@DESTINY MAKERS #3]  SHUPP, MIKE  P
DEATH’S GRAY LAND [@DESTINY MAKERS #4]  SHUPP, MIKE  C/P,P
LAST RECKONING, THE [@DESTINY MAKERS #5]  SHUPP, MIKE  C/P,P
COALESCENT [DESTINY’S CHILDREN 1]  BAXTER, STEPHEN  H,P
EXULTANT [DESTINY’S CHILDREN 2]  BAXTER, STEPHEN  H
TRANSCENDENT [DESTINY’S CHILDREN 3]  BAXTER, STEPHEN  H
DR. QUAKE [DESTROYER 5]  SAPIR, RICHARD BEN/MURPHY, WARREN  P
DEMON AND THE CITY, THE [DETECTIVE INSPECTOR CHEN 2]  WILLIAMS, LIZ  LP
PRECIOUS DRAGON [DETECTIVE INSPECTOR CHEN]  WILLIAMS, LIZ  H
BLACK RAVEN, THE [DEVERRY 10/DRAGON MAGE 2]  KERR, KATHARINE  H,LP
BRISTLING WOOD, THE [DEVERRY 3/DEVERRY PROPER 3]  KERR, KATHARINE  C/P:2,P
DAGGERSPELL [DEVERRY 1/DEVERRY PROPER 1]  KERR, KATHARINE  H:2,SFWA,P
DAGGERSPELL (REVISED) [DEVERRY 1/DEVERRY PROPER 1]  KERR, KATHARINE  C/P:2,P
DARKSPELL [DEVERRY 2/DEVERRY PROPER 2]  KERR, KATHARINE  C/P,SFWA,P,H
DARKSPELL (REVISED) [DEVERRY 2/DEVERRY PROPER 2]  KERR, KATHARINE  C/P,P
DAYS OF AIR AND DARKNESS [DEVERRY 8/WESTLANDS 4]  KERR, KATHARINE  C/P:2,LP
DAYS OF BLOOD AND FIRE [DEVERRY 7/WESTLANDS 3]  KERR, KATHARINE  C/P:2,LP
DRAGON REVENANT, THE [DEVERRY 4/DEVERRY PROPER 4]  KERR, KATHARINE  C/P,LP,H
FIRE DRAGON, THE [DEVERRY 11/DRAGON MAGE 3]  KERR, KATHARINE  P
RED WYVERN, THE [DEVERRY 9/DRAGON MAGE 1]  KERR, KATHARINE  C/P,LP
TIME OF EXILE, A [DEVERRY 5/WESTLANDS 1]  KERR, KATHARINE  C/P,P,H
TIME OF JUSTICE, A [DEVERRY 7/WESTLANDS 3]  KERR, KATHARINE  P
TIME OF OMENS, A [DEVERRY 6/WESTLANDS 2]  KERR, KATHARINE  H:2,P
TIME OF WAR, A [DEVERRY]  KERR, KATHARINE  P
DEVIL’S TANGO [DEVIL DANCES 2]  JUBERT, HERVE  H
DEVIL’S ENGINE [DEVIL’S WEST 2]  SUMNER, MARK  P
DEVIL’S TOWER [DEVIL’S WEST 1]  SUMNER, MARK  C/P,P
MOON OF THE SPIDER [@DIABLO]  KNAAK, RICHARD A.  P
LEGACY OF BLOOD [@DIABLO #1]  KNAAK, RICHARD A.  P
BLACK ROAD, THE [@DIABLO #2]  ODOM, MEL  C/P,P
KINGDOM OF SHADOW, THE [@DIABLO #3]  KNAAK, RICHARD A.  C/P,P
BIRTHRIGHT [@DIABLO SIN WAR #1]  KNAAK, RICHARD A.  P
SCALES OF THE SERPENT [@DIABLO SIN WAR #2]  KNAAK, RICHARD A.  P
DIADEM FROM THE STARS [DIADEM 1]  CLAYTON, JO  C/P:3,P
GHOSTHUNT [DIADEM 7]  CLAYTON, JO  C/P:2,P
IRSUD [DIADEM 3]  CLAYTON, JO  C/P:2,P
LAMARCHOS [DIADEM 2]  CLAYTON, JO  C/P:2,P
MAEVE [DIADEM 4]  CLAYTON, JO  C/P:2,P
NOWHERE HUNT, THE [DIADEM 6]  CLAYTON, JO  C/P:2,P
QUESTER’S ENDGAME [DIADEM 9]  CLAYTON, JO  C/P:2,P
SNARES OF IBEX, THE [DIADEM 8]  CLAYTON, JO  C/P:2,P
STAR HUNTERS [DIADEM 5]  CLAYTON, JO  C/P:2,P
GATEWAY TO XANADU [DIANA SANTEE 2]  GREEN, SHARON  P
MIND GUEST [DIANA SANTEE 1]  GREEN, SHARON  P
CHILDREN OF THE NIGHT [DIANA TREGARDE]  LACKEY, MERCEDES  C/P:2,P
JINX HIGH [DIANA TREGARDE]  LACKEY, MERCEDES  C/P:2,P
BY THE TIME I GET TO NASHVILLE [DIEGO 2]  FENN, LIONEL  P
ONCE UPON A TIME IN THE EAST [DIEGO 1]  FENN, LIONEL  P
DAY OF WRATH [DIES IRAE 3]  STABLEFORD, BRIAN M.  C/P:2,P
DAYS OF GLORY, THE [DIES IRAE 1]  STABLEFORD, BRIAN M.  C/P:2,P
IN THE KINGDOM OF THE BEASTS [DIES IRAE 2]  STABLEFORD, BRIAN M.  C/P:2,P
FORGOTTEN STAR, THE [DIG ALLEN #1]  GREENE, JOSEPH  P,@H
CAPTIVES IN SPACE [@DIG ALLEN #2]  GREENE, JOSEPH  P,@H
JOURNEY TO JUPITER [@DIG ALLEN #3]  GREENE, JOSEPH  H
TRAPPERS OF VENUS [@DIG ALLEN #4]  GREENE, JOSEPH  H
ROBOTS OF SATURN [@DIG ALLEN #5]  GREENE, JOSEPH  H
BOUNCING OFF THE MOON [DINGILLIAD 2]  GERROLD, DAVID  H
JUMPING OFF THE PLANET [DINGILLIAD 1]  GERROLD, DAVID  H,P
LEAPING TO THE STARS [DINGILLIAD 3]  GERROLD, DAVID  H
FAR-SEER [DINOSAUR EMPIRE 1]  SAWYER, ROBERT J.  C/P:2,SFWA-P:2,P
FOREIGNER [DINOSAUR EMPIRE 3]  SAWYER, ROBERT J.  P

The MITSFS Seriesdex  September 24, 2007  Page 37
TITLE  AUTHOR  CODE
FOSSIL HUNTER [DINOSAUR EMPIRE 2]  SAWYER, ROBERT J.  P,SFWA-P
DINOSAUR PLANET [DINOSAUR PLANET 1]  MCCAFFREY, ANNE  C/P:3,P
DINOSAUR PLANET SURVIVORS [DINOSAUR PLANET 2]  MCCAFFREY, ANNE  C/P:2,P
IRETA ADVENTURE, THE [DINOSAUR PLANET 1,2]  MCCAFFREY, ANNE  H
SURVIVORS, THE [DINOSAUR PLANET 2]  MCCAFFREY, ANNE  P
DINOTOPIA LOST [DINOTOPIA]  FOSTER, ALAN DEAN  H
HAND OF DINOTOPIA, THE [DINOTOPIA]  FOSTER, ALAN DEAN  H
FIRE ANGELS [DION HOLYHANDS 1]  ROUTLEY, JANE  H,LP
MAGE HEART [DION HOLYHANDS 2]  ROUTLEY, JANE  C/P,P
CATSEYE [DIPPLE 1]  NORTON, ANDRE  C/P,P,H
MASKS OF THE OUTCASTS [DIPPLE 1,3]  NORTON, ANDRE  H
NIGHT OF MASKS [DIPPLE 3]  NORTON, ANDRE  C/P,P,H
CAPTAIN [DIRIGENT MERCENARY CORPS 3]  SHELLEY, RICK  C/P,P
COLONEL [DIRIGENT MERCENARY CORPS 6]  SHELLEY, RICK  C/P,P
LIEUTENANT [DIRIGENT MERCENARY CORPS 2]  SHELLEY, RICK  C/P,P
LIEUTENANT COLONEL [DIRIGENT MERCENARY CORPS 5]  SHELLEY, RICK  C/P,P
MAJOR [DIRIGENT MERCENARY CORPS 4]  SHELLEY, RICK  C/P,P
OFFICER-CADET [DIRIGENT MERCENARY CORPS 1]  SHELLEY, RICK  C/P,P
DIRK GENTLY’S HOLISTIC DETECTIVE AGENCY [DIRK GENTLY 1]  ADAMS, DOUGLAS  H:2,P
LONG DARK TEA-TIME OF THE SOUL, THE [DIRK GENTLY 2]  ADAMS, DOUGLAS  C/P,P,H
TREASURE [DIRK PITT]  CUSSLER, CLIVE  P
LERIOS MECCA, THE [DIRSHAN 1]  LANCOUR, GENE  H,SFWA
MAN-EATERS OF CASCALON, THE [DIRSHAN 4]  LANCOUR, GENE  SFWA
SWORD FOR THE EMPIRE [DIRSHAN 3]  LANCOUR, GENE  H,SFWA
WAR MACHINES OF KALINTH, THE [DIRSHAN 2]  LANCOUR, GENE  H:2,SFWA
AMAZING MAURICE AND HIS EDUCATED RODENTS, THE [DISCWORLD 27]  PRATCHETT, TERRY  H
CARPE JUGULUM [DISCWORLD 23/DISCWORLD WITCHES 6]  PRATCHETT, TERRY  H:2,P
COLOUR OF MAGIC, THE [DISCWORLD 1/DISCWORLD RINCEWIND 1]  PRATCHETT, TERRY  C/P,P,H:2
DISCWORLD COMPANION, THE [DISCWORLD]  PRATCHETT, TERRY/BRIGGS, STEPHEN  R/REF-P
EQUAL RITES [DISCWORLD 3/DISCWORLD WITCHES 1]  PRATCHETT, TERRY  C/P:2,P,H
ERIC [DISCWORLD 9/DISCWORLD RINCEWIND 4]  PRATCHETT, TERRY  C/P:3,SFWA-P,VLP
FEET OF CLAY [DISCWORLD 19/DISCWORLD WATCHMEN 3]  PRATCHETT, TERRY  H:2,P
FIFTH ELEPHANT, THE [DISCWORLD 24/DISCWORLD WATCHMEN 5]  PRATCHETT, TERRY  H:2,P
GOING POSTAL [DISCWORLD 32]  PRATCHETT, TERRY  C/P,P,H
GUARDS! GUARDS! [DISCWORLD 8/DISCWORLD WATCHMEN 1]  PRATCHETT, TERRY  H:2,P
HAT FULL OF SKY, A [DISCWORLD 31/DISCWORLD TIFFANY ACHING 2]  PRATCHETT, TERRY  H
HOGFATHER [DISCWORLD 20/DISCWORLD DEATH 4]  PRATCHETT, TERRY  H,P
INTERESTING TIMES [DISCWORLD 17/DISCWORLD RINCEWIND 5]  PRATCHETT, TERRY  C/P,P,H
JINGO [DISCWORLD 21/DISCWORLD WATCHMEN 4]  PRATCHETT, TERRY  H:2,P
LAST CONTINENT, THE [DISCWORLD 22/DISCWORLD RINCEWIND 6]  PRATCHETT, TERRY  C/P,P
LAST HERO, THE [DISCWORLD]  PRATCHETT, TERRY  VLP
LIGHT FANTASTIC, THE [DISCWORLD 2/DISCWORLD RINCEWIND 2]  PRATCHETT, TERRY  C/P:2,P,H
LORDS AND LADIES [DISCWORLD 14/DISCWORLD WITCHES 4]  PRATCHETT, TERRY  C/P:2,LP,P
LORDS AND LADIES (PROOF) [DISCWORLD 14/DISCWORLD WITCHES 4]  PRATCHETT, TERRY  SR-LP
MASKERADE [DISCWORLD 18/DISCWORLD WITCHES 5]  PRATCHETT, TERRY  C/P,P,H
MEN AT ARMS [DISCWORLD 15/DISCWORLD WATCHMEN 2]  PRATCHETT, TERRY  C/P,P,H
MONSTROUS REGIMENT [DISCWORLD 30]  PRATCHETT, TERRY  H:2,P
MORT [DISCWORLD 4/DISCWORLD DEATH 1]  PRATCHETT, TERRY  C/P,H,P
MOVING PICTURES [DISCWORLD 10]  PRATCHETT, TERRY  C/P:3,P,H
NIGHT WATCH [DISCWORLD 28/DISCWORLD WATCHMEN 6]  PRATCHETT, TERRY  H:2
PYRAMIDS [DISCWORLD 7]  PRATCHETT, TERRY  H:2,P
REAPER MAN [DISCWORLD 11/DISCWORLD DEATH 2]  PRATCHETT, TERRY  H:2,SFWA-P,P

RINCEWIND THE WIZARD [DISCWORLD 1,2,5,9/ DISCWORLD RINCEWIND 1,2,3,4] PRATCHETT, TERRY

H

SMALL GODS [DISCWORLD 13]  PRATCHETT, TERRY  H:2,P
SOUL MUSIC [DISCWORLD 16/DISCWORLD DEATH 3]  PRATCHETT, TERRY  C/P:2,P
SOURCERY [DISCWORLD 5/DISCWORLD RINCEWIND 3]  PRATCHETT, TERRY  C/P,P,H
THIEF OF TIME [DISCWORLD 26]  PRATCHETT, TERRY  H:2,P
THUD! [DISCWORLD 33/DISCWORLD WATCHMEN 7]  PRATCHETT, TERRY  H,P
TRUTH, THE [DISCWORLD 25]  PRATCHETT, TERRY  H:2,P
WEE FREE MEN, THE [DISCWORLD 29/DISCWORLD TIFFANY ACHING 1]  PRATCHETT, TERRY  H,P
WHERE’S MY COW? [DISCWORLD]  PRATCHETT, TERRY  H
WITCHES ABROAD [DISCWORLD 12/DISCWORLD WITCHES 3]  PRATCHETT, TERRY  C/P,SFWA-P,P
WYRD SISTERS [DISCWORLD 6/DISCWORLD WITCHES 2]  PRATCHETT, TERRY  C/P,P,H
WINTERSMITH [DISCWORLD TIFFANY ACHING 3]  PRATCHETT, TERRY  H:2
DISINHERITED, THE [DISINHERITED 1]  WHITE, STEVE  P,SFWA-P
LEGACY [DISINHERITED 2]  WHITE, STEVE  P
DIVINE ENDURANCE [DIVINE ENDURANCE 1]  JONES, GWYNETH  H,SFWA-P,P
FLOWERDUST [DIVINE ENDURANCE 2]  JONES, GWYNETH  LP
DOC SAVAGE FILES, THE [DOC SAVAGE]  DENT, LESTER  R/REF-P
DOC SAVAGE OMNIBUS 1 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 3 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 4 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P

TITLE  AUTHOR  CODE
DOC SAVAGE OMNIBUS 5 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 6 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 7 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 8 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 9 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 10 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 11 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 12 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 13 [DOC SAVAGE]  ROBESON, KENNETH  P
ESCAPE FROM LOKI [DOC SAVAGE]  FARMER, PHILIP JOSE  P
FLIGHT INTO FEAR [DOC SAVAGE]  ROBESON, KENNETH  P
FORGOTTEN REALM, THE [DOC SAVAGE]  ROBESON, KENNETH  P
FRIGHTENED FISH, THE [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
GHOST LEGION, THE [DOC SAVAGE 3]  ROBESON, KENNETH  H
JADE OGRE, THE [DOC SAVAGE]  ROBESON, KENNETH  C/P:2,P
PYTHON ISLE [DOC SAVAGE]  ROBESON, KENNETH  C/P:2,P
WHISTLING WRAITH, THE [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
WHITE EYES [DOC SAVAGE]  ROBESON, KENNETH  P
MAN OF BRONZE, THE [DOC SAVAGE #1]  ROBESON, KENNETH  C/P:2,@P
THOUSAND-HEADED MAN, THE [DOC SAVAGE #2]  ROBESON, KENNETH  C/P:2,@P
METEOR MENACE [DOC SAVAGE #3]  ROBESON, KENNETH  C/P:2,@P
POLAR TREASURE, THE [DOC SAVAGE #4]  ROBESON, KENNETH  C/P,@P
BRAND OF THE WEREWOLF [DOC SAVAGE #5]  ROBESON, KENNETH  C/P:2,@P
LOST OASIS, THE [DOC SAVAGE #6]  ROBESON, KENNETH  C/P:2,@P
MONSTERS, THE [DOC SAVAGE #7]  ROBESON, KENNETH  C/P:2,@P
LAND OF TERROR, THE [DOC SAVAGE #8]  ROBESON, KENNETH  C/P,@P
MYSTIC MULLAH, THE [DOC SAVAGE #9]  ROBESON, KENNETH  @P
PHANTOM CITY, THE [DOC SAVAGE #10]  ROBESON, KENNETH  C/P,@P
FEAR CAY [DOC SAVAGE #11]  ROBESON, KENNETH  C/P:2,@P
QUEST OF QUI [DOC SAVAGE #12]  ROBESON, KENNETH  C/P:2,@P
LAND OF ALWAYS-NIGHT [DOC SAVAGE #13]  ROBESON, KENNETH  C/P:2,@P
FANTASTIC ISLAND, THE [DOC SAVAGE #14]  ROBESON, KENNETH  C/P,@P
MURDER MELODY [DOC SAVAGE #15]  ROBESON, KENNETH  C/P:2,@P
SPOOK LEGION, THE [DOC SAVAGE #16]  ROBESON, KENNETH  C/P:2,@P
RED SKULL, THE [DOC SAVAGE #17]  ROBESON, KENNETH  C/P:2,@P
SARGASSO OGRE, THE [DOC SAVAGE #18]  ROBESON, KENNETH  H,@P
PIRATE OF THE PACIFIC [DOC SAVAGE #19]  ROBESON, KENNETH  C/P,@P
SECRET IN THE SKY, THE [DOC SAVAGE #20]  ROBESON, KENNETH  C/P:2,@P
COLD DEATH [DOC SAVAGE #21]  ROBESON, KENNETH  C/P,@P
CZAR OF FEAR, THE [DOC SAVAGE #22]  ROBESON, KENNETH  C/P:2,@P
FORTRESS OF SOLITUDE [DOC SAVAGE #23]  ROBESON, KENNETH  @P
GREEN EAGLE, THE [DOC SAVAGE #24]  ROBESON, KENNETH  C/P,@P
DEVIL’S PLAYGROUND, THE [DOC SAVAGE #25]  ROBESON, KENNETH  C/P:2,@P
DEATH IN SILVER [DOC SAVAGE #26]  ROBESON, KENNETH  C/P:2,@P
MYSTERY UNDER THE SEA [DOC SAVAGE #27]  ROBESON, KENNETH  C/P,@P
DEADLY DWARF, THE [DOC SAVAGE #28]  ROBESON, KENNETH  C/P:2,@P
OTHER WORLD, THE [DOC SAVAGE #29]  ROBESON, KENNETH  @P
FLAMING FALCONS, THE [DOC SAVAGE #30]  ROBESON, KENNETH  C/P:2,@P
ANNIHILIST [DOC SAVAGE #31]  ROBESON, KENNETH  C/P:2,@P
DUST OF DEATH [DOC SAVAGE #32]  ROBESON, KENNETH  C/P:2,@P
TERROR IN THE NAVY, THE [DOC SAVAGE #33]  ROBESON, KENNETH  C/P:2,@P
MAD EYES [DOC SAVAGE #34]  ROBESON, KENNETH  C/P,@P
SQUEAKING GOBLIN, THE [DOC SAVAGE #35]  ROBESON, KENNETH  C/P:2,@P
RESURRECTION DAY [DOC SAVAGE #36]  ROBESON, KENNETH  C/P:2,@P
HEX [DOC SAVAGE #37]  ROBESON, KENNETH  C/P:2,@P
RED SNOW [DOC SAVAGE #38]  ROBESON, KENNETH  C/P,@P
WORLD’S FAIR GOBLIN [DOC SAVAGE #39]  ROBESON, KENNETH  C/P,@P
DAGGER IN THE SKY, THE [DOC SAVAGE #40]  ROBESON, KENNETH  C/P:2,@P
MERCHANTS OF DISASTER [DOC SAVAGE #41]  ROBESON, KENNETH  C/P:2,@P
GOLD OGRE, THE [DOC SAVAGE #42]  ROBESON, KENNETH  C/P,@P
MAN WHO SHOOK THE EARTH, THE [DOC SAVAGE #43]  ROBESON, KENNETH  C/P:2,@P
SEA MAGICIAN, THE [DOC SAVAGE #44]  ROBESON, KENNETH  C/P:2,@P
MEN WHO SMILED NO MORE, THE [DOC SAVAGE #45]  ROBESON, KENNETH  C/P:2,@P
MIDAS MAN, THE [DOC SAVAGE #46]  ROBESON, KENNETH  C/P:2,@P
LAND OF LONG JUJU [DOC SAVAGE #47]  ROBESON, KENNETH  C/P,@P
FEATHERED OCTOPUS, THE [DOC SAVAGE #48]  ROBESON, KENNETH  C/P:2,@P
SEA ANGEL, THE [DOC SAVAGE #49]  ROBESON, KENNETH  C/P:2,@P
DEVIL ON THE MOON [DOC SAVAGE #50]  ROBESON, KENNETH  C/P:2,@P
HAUNTED OCEAN [DOC SAVAGE #51]  ROBESON, KENNETH  C/P,@P
VANISHER, THE [DOC SAVAGE #52]  ROBESON, KENNETH  C/P:2,@P
MENTAL WIZARD, THE [DOC SAVAGE #53]  ROBESON, KENNETH  C/P,@P
HE COULD STOP THE WORLD [DOC SAVAGE #54]  ROBESON, KENNETH  C/P,@P
GOLDEN PERIL, THE [DOC SAVAGE #55]  ROBESON, KENNETH  C/P,@P

The MITSFS Seriesdex  September 24, 2007  Page 39
TITLE  AUTHOR  CODE
GIGGLING GHOSTS, THE [DOC SAVAGE #56]  ROBESON, KENNETH  C/P:2,@P
POISON ISLAND [DOC SAVAGE #57]  ROBESON, KENNETH  C/P,@P
MUNITIONS MASTER, THE [DOC SAVAGE #58]  ROBESON, KENNETH  C/P:2,@P
YELLOW CLOUD, THE [DOC SAVAGE #59]  ROBESON, KENNETH  C/P:2,@P
MAJII, THE [DOC SAVAGE #60]  ROBESON, KENNETH  C/P,@P
LIVING FIRE MENACE, THE [DOC SAVAGE #61]  ROBESON, KENNETH  C/P:2,@P
PIRATE’S GHOST, THE [DOC SAVAGE #62]  ROBESON, KENNETH  C/P:2,@P
SUBMARINE MYSTERY, THE [DOC SAVAGE #63]  ROBESON, KENNETH  C/P:2,@P
MOTION MENACE [DOC SAVAGE #64]  ROBESON, KENNETH  C/P,@P
GREEN DEATH, THE [DOC SAVAGE #65]  ROBESON, KENNETH  C/P:2,@P
MAD MESA [DOC SAVAGE #66]  ROBESON, KENNETH  C/P:2,@P
FRECKLED SHARK, THE [DOC SAVAGE #67]  ROBESON, KENNETH  C/P:2,@P
QUEST OF THE SPIDER [DOC SAVAGE #68]  ROBESON, KENNETH  C/P:2,@P
MYSTERY ON THE SNOW, THE [DOC SAVAGE #69]  ROBESON, KENNETH  C/P:2,@P
SPOOK HOLE [DOC SAVAGE #70]  ROBESON, KENNETH  C/P:2,@P
MURDER MIRAGE [DOC SAVAGE #71]  ROBESON, KENNETH  C/P:2,@P
METAL MASTER, THE [DOC SAVAGE #72]  ROBESON, KENNETH  C/P:2,@P
SEVEN AGATE DEVILS, THE [DOC SAVAGE #73]  ROBESON, KENNETH  C/P:2,@P
DERRICK DEVIL, THE [DOC SAVAGE #74]  ROBESON, KENNETH  C/P:2,@P
LAND OF FEAR, THE [DOC SAVAGE #75]  ROBESON, KENNETH  C/P:2,@P
BLACK SPOT, THE [DOC SAVAGE #76]  ROBESON, KENNETH  C/P:2,@P
SOUTH POLE TERROR, THE [DOC SAVAGE #77]  ROBESON, KENNETH  C/P,@P
CRIMSON SERPENT, THE [DOC SAVAGE #78]  ROBESON, KENNETH  C/P:2,@P
DEVIL GENGHIS, THE [DOC SAVAGE #79]  ROBESON, KENNETH  C/P:2,@P
KING MAKER, THE [DOC SAVAGE #80]  ROBESON, KENNETH  C/P:2,@P
STONE MAN, THE [DOC SAVAGE #81]  ROBESON, KENNETH  @P
EVIL GNOME, THE [DOC SAVAGE #82]  ROBESON, KENNETH  C/P,@P
RED TERRORS, THE [DOC SAVAGE #83]  ROBESON, KENNETH  C/P,@P
MOUNTAIN MONSTER, THE [DOC SAVAGE #84]  ROBESON, KENNETH  C/P:2,@P
BOSS OF TERROR, THE [DOC SAVAGE #85]  ROBESON, KENNETH  @P
ANGRY GHOST, THE [DOC SAVAGE #86]  ROBESON, KENNETH  @P
SPOTTED MEN, THE [DOC SAVAGE #87]  ROBESON, KENNETH  @P
ROAR DEVIL, THE [DOC SAVAGE #88]  ROBESON, KENNETH  C/P,@P
MAGIC ISLAND, THE [DOC SAVAGE #89]  ROBESON, KENNETH  C/P,@P
FLYING GOBLIN, THE [DOC SAVAGE #90]  ROBESON, KENNETH  C/P,@P
PURPLE DRAGON, THE [DOC SAVAGE #91]  ROBESON, KENNETH  @P
AWFUL EGG, THE [DOC SAVAGE #92]  ROBESON, KENNETH  @P
TUNNEL TERROR [DOC SAVAGE #93]  ROBESON, KENNETH  @P
HATE GENIUS, THE [DOC SAVAGE #94]  ROBESON, KENNETH  @P
RED SPIDER, THE [DOC SAVAGE #95]  ROBESON, KENNETH  @P
MYSTERY ON HAPPY BONES, THE [DOC SAVAGE #96]  ROBESON, KENNETH  @P
SATAN BLACK [DOC SAVAGE #97]/CARGO UNKNOWN [DOC SAVAGE 98]  ROBESON, KENNETH  @P
HELL BELOW [DOC SAVAGE #99]/LOST GIANT, THE [DOC SAVAGE 100]  ROBESON, KENNETH  C/P,@P

PHAROAH’S GHOST, THE [DOC SAVAGE #101]/ TIME TERROR, THE [DOC SAVAGE 102] ROBESON, KENNETH

@P

WHISKER OF HERCULES, THE [DOC SAVAGE #103]/ MAN WHO WAS SCARED, THE [DOC SAVAGE 104] ROBESON, KENNETH

@P

THEY DIED TWICE [DOC SAVAGE #105]/ SCREAMING MAN, THE [DOC SAVAGE 106] ROBESON, KENNETH

@P

JIU SAN [DOC SAVAGE #107]/BLACK, BLACK WITCH, THE [DOC SAVAGE 108] ROBESON, KENNETH  @P

SHAPE OF TERROR, THE [DOC SAVAGE #109]/ DEATH HAD YELLOW EYES [DOC SAVAGE 110] ROBESON, KENNETH

@P

ONE-EYED MYSTIC [DOC SAVAGE #111]/ MAN WHO FELL UP, THE [DOC SAVAGE 112] ROBESON, KENNETH

@P

TALKING DEVIL, THE [DOC SAVAGE #113]/ TEN TON SNAKES, THE [DOC SAVAGE 114] ROBESON, KENNETH

@P

PIRATE ISLE [DOC SAVAGE #115]/SPEAKING STONE, THE [DOC SAVAGE 116] ROBESON, KENNETH  @P

GOLDEN MAN, THE [DOC SAVAGE #117]/ PERIL IN THE NORTH [DOC SAVAGE 118] ROBESON, KENNETH

@P

LAUGH OF DEATH, THE [DOC SAVAGE #119]/ KING OF TERROR, THE [DOC SAVAGE 120] ROBESON, KENNETH

@P

THREE WILD MEN, THE [DOC SAVAGE #121]/ FIERY MENACE, THE [DOC SAVAGE 122] ROBESON, KENNETH

@P

DEVILS OF THE DEEP [DOC SAVAGE #123]/ HEADLESS MEN, THE [DOC SAVAGE 124] ROBESON, KENNETH

@P

GOBLINS, THE [DOC SAVAGE #125]/ SECRET OF THE SU, THE [DOC SAVAGE 126] ROBESON, KENNETH

@P

DOC SIDHE [DOC SIDHE 1]  ALLSTON, AARON  P
SIDHE DEVIL [DOC SIDHE 2]  ALLSTON, AARON  P
DR. DIMENSION [DOCTOR DIMENSION]  DE CHANCIE, JOHN/BISCHOFF, DAVID F.  C/P:2,P
DR. DIMENSION: MASTERS OF SPACETIME [DOCTOR DIMENSION]  DE CHANCIE, JOHN/BISCHOFF, DAVID F.  P
DOCTOR DOLITTLE’S CIRCUS [DOCTOR DOLITTLE]  LOFTING, HUGH  H,P
DOCTOR DOLITTLE’S GARDEN [DOCTOR DOLITTLE]  LOFTING, HUGH  H

The MITSFS Seriesdex  September 24, 2007  Page 40
TITLE  AUTHOR  CODE
DOCTOR DOLITTLE’S POST OFFICE [DOCTOR DOLITTLE]  LOFTING, HUGH  H,P
DOCTOR DOLITTLE’S RETURN [DOCTOR DOLITTLE]  LOFTING, HUGH  H
DOCTOR DOLITTLE’S ZOO [DOCTOR DOLITTLE]  LOFTING, HUGH  P
VOYAGES OF DOCTOR DOLITTLE, THE [DOCTOR DOLITTLE]  LOFTING, HUGH  C/P,P,H
NIGHTMARE [DOCTOR STRANGE]  ROTSLER, WILLIAM  C/P,P
ACE! [DOCTOR WHO]  ALDRED, SOPHIE/TUCKER, MIKE  R/REF-H
ADVENTURES OF DOCTOR WHO, THE [DOCTOR WHO]  DICKS, TERRANCE  H

ADVENTURES OF K9 AND OTHER MECHANICAL CREATURES [DOCTOR WHO] DICKS, TERRANCE

P

ADVENTURESS OF HENRIETTA STREET, THE [DOCTOR WHO]  MILES, LAWRENCE  P
ALIENS AND ENEMIES [DOCTOR WHO]  RICHARDS, JUSTIN  R/REF-H
AMBASSADORS OF DEATH, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P
ARC OF INFINITY [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
ARK, THE [DOCTOR WHO]  ERICKSON, PAUL  SR-P,C/P
ART OF DESTRUCTION, THE [@DOCTOR WHO]  COLE, STEPHEN  H
AWAKENING, THE [DOCTOR WHO]  PRINGLE, ERIC  SR-P,C/P
AZTECS, THE [DOCTOR WHO]  LUCAROTTI, JOHN  SR-P,C/P:2
BATTLEFIELD [DOCTOR WHO]  PLATT, MARC  C/P,P
BLACK ORCHID [DOCTOR WHO]  DUDLEY, TERRENCE  SR-P,C/P
BLOOD AND HOPE [DOCTOR WHO]  MCLAUGHLIN, IAIN  SR-H
CABINET OF LIGHT, THE [DOCTOR WHO/TIME HUNTER 0]  O’MAHONY, DANIEL  SR-H
CASTROVALVA [DOCTOR WHO]  BIDMEAD, CHRISTOPHER H.  SR-P,C/P
CAT WHO WALKED THROUGH TIME, THE [DOCTOR WHO]  KELLY, TOM/KELLY, ALRYSSA  SR-VLP
CAVES OF ANDROZANI, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
CELESTIAL TOYMAKER, THE [DOCTOR WHO]  DAVIS, GERRY/BINGEMAN, ALISON  SR-P
CITADEL OF DREAMS [DOCTOR WHO]  STONE, DAVE  SR-H
COMPANION PIECE [DOCTOR WHO]  PERRY, ROBERT/TUCKER, MIKE  SR-H
COMPANIONS [DOCTOR WHO]  HOWE, DAVID J./STAMMERS, MARK  R/REF-H
CREATURES AND DEMONS [DOCTOR WHO]  RICHARDS, JUSTIN  R/REF-H
DALEK FACTOR, THE [DOCTOR WHO]  CLARK, SIMON  SR-H
DEATH TO THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P

DECALOG [DOCTOR WHO] STAMMERS, MARK/

WALKER, STEPHEN JAMES

C/PA,PA

DECALOG 2: LOST PROPERTY [DOCTOR WHO] STAMMERS, MARK/

WALKER, STEPHEN JAMES

C/PA:2,PA

DECALOG 3: CONSEQUENCES [DOCTOR WHO]  LANE, ANDY/RICHARDS, JUSTIN  C/PA,PA
DECALOG 4 — RE:GENERATIONS [DOCTOR WHO]  LANE, ANDY/RICHARDS, JUSTIN  PA
DECALOG 5 — WONDERS [DOCTOR WHO]  LEONARD, PAUL/MORTIMORE, JIM  PA
DELTA AND THE BANNERMEN (MISSPELLED ON SPINE) [DOCTOR WHO]  KOHLL, MALCOLM  SR-P,C/P
DOCTOR AND THE ENTERPRISE, THE [DOCTOR WHO/STAR TREK]  AIREY, JEAN  SR-LP
DOCTOR WHO [DOCTOR WHO]  RUSSELL, GARY  C/P,P
DOCTOR WHO 1985 SUMMER SPECIAL CLASSIC [DOCTOR WHO]  MARVEL COMICS  CX
DOCTOR WHO AND AN UNEARTHLY CHILD [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE ABOMINABLE SNOWMEN [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE ANDROID INVASION [DOCTOR WHO]  DICKS, TERRANCE  C/P,P
DOCTOR WHO AND THE ANDROIDS OF TARA [DOCTOR WHO]  DICKS, TERRANCE  SR-P,P
DOCTOR WHO AND THE ARK IN SPACE [DOCTOR WHO]  MARTER, IAN  SR-P
DOCTOR WHO AND THE ARMAGEDDON FACTOR [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE AUTON INVASION [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE BRAIN OF MORBIUS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,H,P
DOCTOR WHO AND THE CARNIVAL OF MONSTERS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE CAVE-MONSTERS [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE CLAWS OF AXOS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE CREATURE FROM THE PIT [DOCTOR WHO]  FISHER, DAVID  SR-P,C/P
DOCTOR WHO AND THE CRUSADERS [DOCTOR WHO]  WHITAKER, DAVID  SR-P,C/P:2
DOCTOR WHO AND THE CURSE OF PELADON [DOCTOR WHO]  HAYLES, BRIAN  SR-P,C/P:2
DOCTOR WHO AND THE CYBERMEN [DOCTOR WHO]  DAVIS, GERRY  SR-P,C/P:2
DOCTOR WHO AND THE DAEMONS [DOCTOR WHO]  LETTS, BARRY  SR-P,C/P:2
DOCTOR WHO AND THE DALEK INVASION OF EARTH [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:3
DOCTOR WHO AND THE DALEKS [DOCTOR WHO]  WHITAKER, DAVID  SR-P,P
DOCTOR WHO AND THE DAY OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE DEADLY ASSASSIN [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE DESTINY OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE DINOSAUR INVASION [DOCTOR WHO]  HULKE, MALCOLM  P
DOCTOR WHO AND THE DOOMSDAY WEAPON [DOCTOR WHO]  HULKE, MALCOLM  C/P:2,P
DOCTOR WHO AND THE ENEMY OF THE WORLD [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
DOCTOR WHO AND THE FACE OF EVIL [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE GENESIS OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE GIANT ROBOT [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE GREEN DEATH [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE HAND OF FEAR [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE HORNS OF NIMON [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE HORROR OF FANG ROCK [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P

TITLE  AUTHOR  CODE
DOCTOR WHO AND THE ICE WARRIORS [DOCTOR WHO]  HAYLES, BRIAN  SR-P,C/P
DOCTOR WHO AND THE IMAGE OF THE FENDAHL [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE INVASION OF TIME [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE INVISIBLE ENEMY [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE KEEPER OF TRAKEN [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE KEYS OF MARINUS [DOCTOR WHO]  HINCHCLIFFE, PHILIP  SR-P,C/P:2
DOCTOR WHO AND THE LEISURE HIVE [DOCTOR WHO]  FISHER, DAVID  SR-P
DOCTOR WHO AND THE LOCH NESS MONSTER [DOCTOR WHO]  DICKS, TERRANCE  C/P,P
DOCTOR WHO AND THE MASQUE OF MANDRAGORA [DOCTOR WHO]  HINCHCLIFFE, PHILIP  C/P,P
DOCTOR WHO AND THE MONSTER OF PELADON [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE MUTANTS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE NIGHTMARE OF EDEN [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE PLANET OF EVIL [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE PLANET OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE PLANET OF THE SPIDERS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE POWER OF KROLL [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE PYRAMIDS OF MARS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE REVENGE OF THE CYBERMEN [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE RIBOS OPERATION [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
DOCTOR WHO AND THE ROBOTS OF DEATH [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE SEA-DEVILS [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE SEEDS OF DOOM [DOCTOR WHO]  HINCHCLIFFE, PHILIP  C/P:2,P
DOCTOR WHO AND THE SONTARAN EXPERIMENT [DOCTOR WHO]  MARTER, IAN  SR-P
DOCTOR WHO AND THE SPACE WAR [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P
DOCTOR WHO AND THE STATE OF DECAY [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE STONES OF BLOOD [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE SUNMAKERS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE TALONS OF WENG-CHIANG [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE TENTH PLANET [DOCTOR WHO]  DAVIS, GERRY  SR-P,C/P:2
DOCTOR WHO AND THE TERROR OF THE AUTONS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE THREE DOCTORS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE TIME MONSTER, THE [DOCTOR WHO]  DICKS, TERRANCE  P
DOCTOR WHO AND THE TIME WARRIOR [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE TOMB OF THE CYBERMEN [DOCTOR WHO]  DAVIS, GERRY  SR-P,C/P:2
DOCTOR WHO AND THE UNDERWORLD [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE VISITATION [DOCTOR WHO]  SAWARD, ERIC  SR-P
DOCTOR WHO AND THE WAR GAMES [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE WEB OF FEAR [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE ZARBI [DOCTOR WHO]  STRUTTON, BILL  SR-P,C/P:2
DOCTOR WHO AND WARRIORS’ GATE [DOCTOR WHO]  LYDECKER, JOHN  SR-P,C/P:2
DOCTOR WHO ANNUAL 1985 [DOCTOR WHO]  BBC TV  HA
DOCTOR WHO ANNUAL 1986 [DOCTOR WHO]  BBC TV  HA

DOCTOR WHO ANNUAL (YEAR UNKNOWN, COPYRIGHT 1981)

[DOCTOR WHO] BBC TV

HA

DOCTOR WHO: ATTACK OF THE CYBERMEN [DOCTOR WHO]  SAWARD, ERIC  P
DOCTOR WHO CROSSWORD BOOK, THE [DOCTOR WHO]  ROBINSON, NIGEL  [C/RAN-P]
DOCTOR WHO CYBERMEN [DOCTOR WHO]  BANKS, DAVID  R/REF-H

DOCTOR WHO IN AN EXCITING ADVENTURE WITH THE DALEKS [DOCTOR WHO] WHITAKER, DAVID

C/P:2,P

DOCTOR WHO: JOURNEY THROUGH TIME [DOCTOR WHO]  BBC TV  H

DOCTOR WHO PROGRAMME GUIDE 1: THE PROGRAMMES, THE [DOCTOR WHO] LOFFICIER, JEAN MARC

R/REF-P,C/REF-P:2

DOCTOR WHO PROGRAMME GUIDE 2:

WHAT’S WHAT AND WHO’S WHO, THE [DOCTOR WHO] LOFFICIER, JEAN MARC

R/REF-P,C/REF-P:2

DOCTOR WHO PROGRAMME GUIDE, THE (2ND EDITION) [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P,C/REF-P

DOCTOR WHO TECHNICAL MANUAL, THE (OFFICIAL ANNIVERSARY VOLUME) [DOCTOR WHO] HARRIS, MARK

R/REF-H,C/REF-H

DOCTOR WHO: THE AUDIO SCRIPTS, VOLUME TWO [DOCTOR WHO]  BIG FINISH  HA
DOCTOR WHO: THE AUDIO SCRIPTS, VOLUME THREE [DOCTOR WHO]  BIG FINISH  HA
DOCTOR WHO: THE CHASE [DOCTOR WHO]  PEEL, JOHN  P

DOCTOR WHO: THE DALEK’S MASTERPLAN - PART I:

MISSION TO THE UNKNOWN [DOCTOR WHO] PEEL, JOHN

P

DOCTOR WHO: THE INSIDE STORY [DOCTOR WHO]  RUSSELL, GARY  R/REF-H
DOCTOR WHO: THE KEY TO TIME [DOCTOR WHO]  HAINING, PETER  R/REF-H
DOCTOR WHO: THE LEGEND CONTINUES [DOCTOR WHO]  RICHARDS, JUSTIN  R/REF-H
DOCTOR WHO: THE SCRIPT OF THE FILM [DOCTOR WHO]  JACOBS, MATTHEW  LP
DOCTOR WHO: THE UNFOLDING TEXT [DOCTOR WHO]  TULLOCH, JOHN/ALVARADO, MANUEL  R/REF-P,C/REF-H
DOCTOR WHO: THE WAR MACHINES [DOCTOR WHO]  BLACK, IAN STUART  P
DOCTOR WHO: THE WHEEL IN SPACE [DOCTOR WHO]  DICKS, TERRANCE  P
DOCTOR WHO YEARBOOK (1991) [DOCTOR WHO]  MARVEL COMICS  HA
DOMINATORS, THE [DOCTOR WHO]  MARTER, IAN  SR-P,C/P:2
EARTHSHOCK [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
EDGE OF DESTRUCTION, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P,C/P

The MITSFS Seriesdex  September 24, 2007  Page 42
TITLE  AUTHOR  CODE

EIGHTIES, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-H

ENLIGHTENMENT [DOCTOR WHO]  CLEGG, BARBARA  SR-P,C/P:2
EVIL OF THE DALEKS [DOCTOR WHO]  PEEL, JOHN  P
EYE OF THE TYGER, THE [DOCTOR WHO]  MCAULEY, PAUL J.  SR-H
FACELESS ONES, THE [@DOCTOR WHO]  DICKS, TERRANCE  P
FALLEN GODS [DOCTOR WHO]  BLUM, JONATHAN/ORMAN, KATE  SR-H
FIFTH DOCTOR, THE [DOCTOR WHO]  HOWE, DAVID J./WALKER, STEPHEN JAMES  R/REF-P,C/REF-P

FIRST DOCTOR, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-P,C/REF-P

FIVE DOCTORS, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
FOREIGN DEVILS [DOCTOR WHO]  CARTMEL, ANDREW  SR-H,H
FOUR TO DOOMSDAY [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2

FOURTH DOCTOR, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-P

FRAYED [DOCTOR WHO]  SAMMS, TARA  SR-H
FRONTIOS [DOCTOR WHO]  BIDMEAD, CHRISTOPHER H.  SR-P,P
FULL CIRCLE [DOCTOR WHO]  SMITH, ANDREW  SR-P,C/P
FURTHER ADVENTURES OF DOCTOR WHO, THE [DOCTOR WHO]  DICKS, TERRANCE  H
FURY FROM THE DEEP [DOCTOR WHO]  PEMBERTON, VICTOR  SR-P
GALAXY FOUR [DOCTOR WHO]  EMMS, WILLIAM  SR-P,C/P
GHOST SHIP [DOCTOR WHO]  TOPPING, KEITH  SR-H
GREATEST SHOW IN THE GALAXY, THE [DOCTOR WHO]  WYATT, STEPHEN  P
GUNFIGHTERS, THE [DOCTOR WHO]  COTTON, DONALD  SR-P
HARRY SULLIVAN’S WAR [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
HISTORY OF THE UNIVERSE, A [DOCTOR WHO]  PARKIN, LANCE  R/REF-P,C/REF-P
I AM A DALEK [DOCTOR WHO]  ROBERTS, GARETH  LP
INFERNO [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
INSTRUMENTS OF DARKNESS [DOCTOR WHO]  RUSSELL, GARY  P
INVASION, THE [DOCTOR WHO]  MARTER, IAN  SR-P,P
INVASION OF THE ORMAZOIDS [DOCTOR WHO]  MARTIN, PHILIP  [C/RAN-P]
K-9 AND COMPANY [DOCTOR WHO]  DUDLEY, TERRENCE  SR-P,C/P
KINDA [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
KING’S DEMONS, THE [DOCTOR WHO]  DUDLEY, TERRENCE  SR-P
KROTONS, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
LICENCE DENIED [DOCTOR WHO]  CORNELL, PAUL  R/REF-P
LIFEDEATH [DOCTOR WHO]  COWE-SPIGAI, HERETH/NEIGHLY, PATRICK  SR-PA
LOGOPOLIS [DOCTOR WHO]  BIDMEAD, CHRISTOPHER H.  SR-P,C/P:2
MACRA TERROR, THE [DOCTOR WHO]  BLACK, IAN STUART  SR-P
MARCO POLO [DOCTOR WHO]  LUCAROTTI, JOHN  SR-P
MARK OF THE MANDRAGORA, THE [DOCTOR WHO]  MARVEL COMICS  CX
MARK OF THE RANI, THE [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  P
MASSACRE, THE [DOCTOR WHO]  LUCAROTTI, JOHN  SR-P
MAWDRYN UNDEAD [DOCTOR WHO]  GRIMWADE, PETER  SR-P,C/P:2
MEGLOS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
MIND OF EVIL, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,P
MINDWARP [DOCTOR WHO]  MARTIN, PHILIP  P
MISSING PIECES [DOCTOR WHO]  PHIPPEN, MARK/LYON, SHAUN  VLPA
MISSION TO MAGNUS [DOCTOR WHO]  MARTIN, PHILIP  P
MISSION TO VENUS [DOCTOR WHO]  EMMS, WILLIAM  [C/RAN-P]
MORE SHORT TRIPS [DOCTOR WHO]  COLE, STEPHEN  PA
MYSTERIOUS PLANET, THE [DOCTOR WHO]  DICKS, TERRANCE  P
MYTH MAKERS, THE [DOCTOR WHO]  COTTON, DONALD  C/P,P
NIGHTDREAMERS [DOCTOR WHO]  ARDEN, TOM  SR-H
NIGHTMARE FAIR, THE [DOCTOR WHO]  WILLIAMS, GRAHAM  SR-P,P
NIGHTMARE OF BLACK ISLAND, THE [@DOCTOR WHO]  TUCKER, MIKE  H
NTH DOCTOR, THE [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P
PARADISE TOWERS [DOCTOR WHO]  WYATT, STEPHEN  P
PERFECT TIMING [DOCTOR WHO]  PHIPPEN, MARK/FAYLE, HELEN  SR-LPA
PERFECT TIMING 2 [DOCTOR WHO]  FAYLE, HELEN/EALES, JULIAN  SR-LPA
PESCATONS, THE [@DOCTOR WHO]  PEMBERTON, VICTOR  P
PLANET OF FIRE [DOCTOR WHO]  GRIMWADE, PETER  SR-P,C/P
POWER OF THE DALEKS, THE [DOCTOR WHO]  PEEL, JOHN  C/P,P
PSI-ENCE FICTION [DOCTOR WHO]  BOUCHER, CHRIS  P
RACE AGAINST TIME [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  [C/RAN-P]
REIGN OF TERROR, THE [DOCTOR WHO]  MARTER, IAN  SR-P
REMEMBRANCE OF THE DALEKS [DOCTOR WHO]  AARONOVITCH, BEN  C/P,P
RESCUE, THE [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
RIP TIDE [DOCTOR WHO]  COOPER, LOUISE  SR-H
SAVAGES, THE [DOCTOR WHO]  BLACK, IAN STUART  SR-P
SEARCH FOR THE DOCTOR [DOCTOR WHO]  MARTIN, DAVID  [C/RAN-P]

SECOND DOCTOR, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES R/REF-P

TITLE  AUTHOR  CODE
SEEDS OF DEATH, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P
SENSORITES, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P
SHADA [DOCTOR WHO]  ADAMS, DOUGLAS  SR-P
SHADOWS OF AVALON, THE [DOCTOR WHO]  CORNELL, PAUL  P
SHELL SHOCK [DOCTOR WHO]  FORWARD, SIMON A.  SR-H
SHORT TRIPS [DOCTOR WHO]  COLE, STEPHEN  C/PA,PA
SHORT TRIPS: A CHRISTMAS TREASURY [DOCTOR WHO]  CORNELL, PAUL  HA
SHORT TRIPS: A DAY IN THE LIFE [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: A UNIVERSE OF TERRORS [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS AND SIDE STEPS [DOCTOR WHO]  COLE, STEPHEN/RAYNER, JACQUELINE  PA
SHORT TRIPS: COMPANIONS [DOCTOR WHO]  RAYNER, JACQUELINE  HA
SHORT TRIPS: DALEK EMPIRE [DOCTOR WHO]  BRIGGS, NICHOLAS/GUERRIER, SIMON  HA
SHORT TRIPS: LIFE SCIENCES [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS: MONTSERS [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: PAST TENSE [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: REPERCUSSIONS [DOCTOR WHO]  RUSSELL, GARY  HA
SHORT TRIPS: SEVEN DEADLY SINS [DOCTOR WHO]  BAILEY, DAVID  HA
SHORT TRIPS: STEEL SKIES [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS: THE CENTENARIAN [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: THE MUSES [DOCTOR WHO]  RAYNER, JACQUELINE  HA
SHORT TRIPS: THE SOLAR SYSTEM [DOCTOR WHO]  RUSSELL, GARY  HA
SHORT TRIPS: 2040 [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS: ZODIAC [DOCTOR WHO]  RAYNER, JACQUELINE  HA
SLIPBACK [DOCTOR WHO]  SAWARD, ERIC  SR-P,C/P
SMUGGLERS, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
SNAKEDANCE [DOCTOR WHO]  DICKS, TERRANCE  SR-P
SPACE PIRATES, THE [DOCTOR WHO]  DICKS, TERRANCE  P
STONE ROSE, THE [@DOCTOR WHO]  RAYNER, JACQUELINE  H
SURVIVAL [DOCTOR WHO]  MUNRO, RONA  SR-P
TERMINUS [DOCTOR WHO]  LYDECKER, JOHN  SR-P,C/P:2
TERRESTRIAL INDEX, THE [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P
TERROR OF THE VERVOIDS [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  SR-P,C/P
THIRD DOCTOR, THE [DOCTOR WHO]  HOWE, DAVID J./WALKER, STEPHEN JAMES  R/REF-P
TIME AND RELATIVE [DOCTOR WHO]  NEWMAN, KIM  SR-H
TIME AND THE RANI [DOCTOR WHO]  BAKER, JANE/BAKER, PIP  P
TIME-FLIGHT [DOCTOR WHO]  GRIMWADE, PETER  SR-P,C/P
TIME MEDDLER, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P,P
TIMELASH [DOCTOR WHO]  MCCOY, GLEN  SR-P,C/P
TIMEVIEW [DOCTOR WHO]  BELLAMY, FRANK/BELLAMY, DAVID  A
TRAVEL WITHOUT THE TARDIS [DOCTOR WHO]  AIREY, JEAN/HALDEMAN, LINDA  R/REF-P
TURLOUGH AND THE EARTHLINK DILEMMA [DOCTOR WHO]  ATTWOOD, TONY  SR-P
TWIN DILEMMA, THE [DOCTOR WHO]  SAWARD, ERIC  SR-P,C/P
TWO DOCTORS, THE [DOCTOR WHO]  HOLMES, ROBERT  SR-P,C/P
ULTIMATE FOE, THE [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  P
UNDERWATER MENACE, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P,C/P
UNIVERSAL DATABANK, THE [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P
VENGEANCE ON VAROS [DOCTOR WHO]  MARTIN, PHILIP  SR-P,C/P
WALKING IN ETERNITY [DOCTOR WHO]  EALES, JULIAN  SR-LPA
WARRIORS OF THE DEEP [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
WEB OF FEAR, THE [DOCTOR WHO]  DICKS, TERRANCE  P
WHO KILLED KENNEDY [DOCTOR WHO]  STEVENS, JAMES/BISHOP, DAVID  C/P,P
WONDERLAND [DOCTOR WHO]  CHADBOURN, MARK  SR-H
WOODEN HEART [@DOCTOR WHO]  DAY, MARTIN  H
ALIEN BODIES [DOCTOR WHO EIGHTH DOCTOR]  MILES, LAWRENCE  C/P:2,P
ANACHROPHOBIA [DOCTOR WHO EIGHTH DOCTOR]  MORRIS, JONATHAN  P
ANCESTOR CELL, THE [DOCTOR WHO EIGHTH DOCTOR]  ANGHELIDES, PETER/COLE, STEPHEN  P
AUTUMN MIST [DOCTOR WHO EIGHTH DOCTOR]  MCINTEE, DAVID A.  P
BANQUO LEGACY, THE [DOCTOR WHO EIGHTH DOCTOR]  LANE, ANDY/RICHARDS, JUSTIN  P
BELTEMPEST [DOCTOR WHO EIGHTH DOCTOR]  MORTIMORE, JIM  P
BLUE ANGEL, THE [DOCTOR WHO EIGHTH DOCTOR]  MAGRS, PAUL/HOAD, JEREMY  P
BODYSNATCHERS, THE [DOCTOR WHO EIGHTH DOCTOR]  MORRIS, MARK  C/P:2,P
BOOK OF THE STILL, THE [DOCTOR WHO EIGHTH DOCTOR]  EBBS, PAUL  P
BURNING, THE [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  P
CAMERA OBSCURA [DOCTOR WHO EIGHTH DOCTOR]  ROSE, LLOYD  P
CASUALTIES OF WAR [DOCTOR WHO EIGHTH DOCTOR]  EMMERSON, STEVE  C/P,P
CITY OF THE DEAD, THE [DOCTOR WHO EIGHTH DOCTOR]  ROSE, LLOYD  P
COLDHEART [DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
CROOKED WORLD, THE [DOCTOR WHO EIGHTH DOCTOR]  LYONS, STEVE  P
DARK PROGENY [DOCTOR WHO EIGHTH DOCTOR]  EMMERSON, STEVE  P
DEADSTONE MEMORIAL, THE [@DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
DEMONTAGE [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  C/P,P
DOMINION [DOCTOR WHO EIGHTH DOCTOR]  WALTERS, NICK  C/P,P
DOMINO EFFECT, THE [DOCTOR WHO EIGHTH DOCTOR]  BISHOP, DAVID  P

TITLE  AUTHOR  CODE
DREAMSTONE MOON [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
EARTHWORLD [DOCTOR WHO EIGHTH DOCTOR]  RAYNER, JACQUELINE  P
EATER OF WASPS [DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
EIGHT DOCTORS, THE [DOCTOR WHO EIGHTH DOCTOR]  DICKS, TERRANCE  C/P:2,P
EMOTIONAL CHEMISTRY [DOCTOR WHO EIGHTH DOCTOR]  FORWARD, SIMON A.  P
ENDGAME [DOCTOR WHO EIGHTH DOCTOR]  DICKS, TERRANCE  P
ESCAPE VELOCITY [DOCTOR WHO EIGHTH DOCTOR]  BRAKE, COLIN  P
FACE-EATER, THE [DOCTOR WHO EIGHTH DOCTOR]  MESSINGHAM, SIMON  P
FALL OF YQUATINE, THE [DOCTOR WHO EIGHTH DOCTOR]  WALTERS, NICK  P
FATHER TIME [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
FRONTIER WORLDS [DOCTOR WHO EIGHTH DOCTOR]  ANGHELIDES, PETER  P
GALLIFREY CHRONICLES, THE [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
GENOCIDE [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
GRIMM REALITY [DOCTOR WHO EIGHTH DOCTOR]  BUCHER-JONES, SIMON/HALE, KELLY  P
HALFLIFE [DOCTOR WHO EIGHTH DOCTOR]  MICHALOWSKI, MARK  P
HISTORY 101 [DOCTOR WHO EIGHTH DOCTOR]  HALLIDAY, MAGS L.  P
HOPE [DOCTOR WHO EIGHTH DOCTOR]  CLAPHAM, MARK  P
INFINITY DOCTORS, THE [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
INFINITY RACE, THE [DOCTOR WHO EIGHTH DOCTOR]  MESSINGHAM, SIMON  P

INTERFERENCE BOOK TWO: THE HOUR OF THE GEEK [DOCTOR WHO EIGHTH DOCTOR] MILES, LAWRENCE

P

JANUS CONJUNCTION, THE [DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
KURSAAL [DOCTOR WHO EIGHTH DOCTOR]  ANGHELIDES, PETER  C/P,P
LAST RESORT, THE [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  P
LEGACY OF THE DALEKS [DOCTOR WHO EIGHTH DOCTOR]  PEEL, JOHN  C/P,P
LONGEST DAY [DOCTOR WHO EIGHTH DOCTOR]  COLLIER, MICHAEL  C/P,P
MAD DOGS AND ENGLISHMEN [DOCTOR WHO EIGHTH DOCTOR]  MAGRS, PAUL  P
OPTION LOCK [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  C/P:2,P
PARALLEL 59 [DOCTOR WHO EIGHTH DOCTOR]  DALLAIRE, NATALIE/COLE, STEPHEN  P
PLACEBO EFFECT [DOCTOR WHO EIGHTH DOCTOR]  RUSSELL, GARY  C/P:2,P
RECKLESS ENGINEERING [DOCTOR WHO EIGHTH DOCTOR]  WALTERS, NICK  P
REVOLUTION MAN [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
SCARLET EMPRESS, THE [DOCTOR WHO EIGHTH DOCTOR]  MAGRS, PAUL  C/P,P
SEEING I [DOCTOR WHO EIGHTH DOCTOR]  BLUM, JONATHAN/ORMAN, KATE  C/P:2,P
SLEEP OF REASON, THE [@DOCTOR WHO EIGHTH DOCTOR]  DAY, MARTIN  P
SLOW EMPIRE, THE [DOCTOR WHO EIGHTH DOCTOR]  STONE, DAVE  P
SOMETIME NEVER... [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  P
SPACE AGE, THE [DOCTOR WHO EIGHTH DOCTOR]  LYONS, STEVE  P
TAINT, THE [DOCTOR WHO EIGHTH DOCTOR]  COLLIER, MICHAEL  C/P,P
TAKING OF PLANET 5, THE [DOCTOR WHO EIGHTH DOCTOR]  BUCHER-JONES, SIMON/CLAPHAM, MARK  P
TIME ZERO [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  P
TIMELESS [DOCTOR WHO EIGHTH DOCTOR]  COLE, STEPHEN  P
TO THE SLAUGHTER [DOCTOR WHO EIGHTH DOCTOR]  COLE, STEPHEN  P
TOMORROW WINDOWS, THE [DOCTOR WHO EIGHTH DOCTOR]  MORRIS, JONATHAN  P
TRADING FUTURES [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
TURING TEST, THE [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
UNNATURAL HISTORY [DOCTOR WHO EIGHTH DOCTOR]  BLUM, JONATHAN/ORMAN, KATE  C/P,P
VAMPIRE SCIENCE [DOCTOR WHO EIGHTH DOCTOR]  BLUM, JONATHAN/ORMAN, KATE  C/P:2,P
VANDERDEKEN’S CHILDREN [DOCTOR WHO EIGHTH DOCTOR]  BULIS, CHRISTOPHER  C/P,P
VANISHING POINT [DOCTOR WHO EIGHTH DOCTOR]  COLE, STEPHEN  P
WAR OF THE DALEKS [DOCTOR WHO EIGHTH DOCTOR]  PEEL, JOHN  C/P:2,P
YEAR OF INTELLIGENT TIGERS, THE [DOCTOR WHO EIGHTH DOCTOR]  ORMAN, KATE  P
BURNING HEART [DOCTOR WHO MISSING ADVENTURES]  STONE, DAVE  C/P,P
COLD FUSION [DOCTOR WHO MISSING ADVENTURES]  PARKIN, LANCE  C/P,P
CRYSTAL BUCEPHALUS, THE [DOCTOR WHO MISSING ADVENTURES]  HINTON, CRAIG  C/P:2,P
DANCING THE CODE [DOCTOR WHO MISSING ADVENTURES]  LEONARD, PAUL  C/P:2,P
DARK PATH, THE [DOCTOR WHO MISSING ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
DEVICE OF DEATH, A [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P,P
DOWNTIME [DOCTOR WHO MISSING ADVENTURES]  PLATT, MARC  C/P:2,P
EMPIRE OF GLASS, THE [DOCTOR WHO MISSING ADVENTURES]  LANE, ANDY  C/P:2,P
ENGLISH WAY OF DEATH, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
EVOLUTION [DOCTOR WHO MISSING ADVENTURES]  PEEL, JOHN  C/P:2,P
EYE OF THE GIANT, THE [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P,P
GHOSTS OF N-SPACE, THE [DOCTOR WHO MISSING ADVENTURES]  LETTS, BARRY  C/P,P
GOTH OPERA [DOCTOR WHO MISSING ADVENTURES]  CORNELL, PAUL  C/P:2,P
INVASION OF THE CAT-PEOPLE [DOCTOR WHO MISSING ADVENTURES]  RUSSELL, GARY  C/P:2,P
KILLING GROUND [DOCTOR WHO MISSING ADVENTURES]  LYONS, STEVE  C/P:2,P
LORDS OF THE STORM [DOCTOR WHO MISSING ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
MAN IN THE VELVET MASK, THE [DOCTOR WHO MISSING ADVENTURES]  O’MAHONY, DANIEL  C/P,P
MANAGRA [DOCTOR WHO MISSING ADVENTURES]  MARLEY, STEPHEN  C/P:2,P
MENAGERIE, THE [DOCTOR WHO MISSING ADVENTURES]  DAY, MARTIN  C/P:2,P
MILLENNIAL RITES [DOCTOR WHO MISSING ADVENTURES]  HINTON, CRAIG  C/P:2,P
PLOTTERS, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P

TITLE  AUTHOR  CODE
ROMANCE OF CRIME, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
SANDS OF TIME, THE [DOCTOR WHO MISSING ADVENTURES]  RICHARDS, JUSTIN  C/P,P
SCALES OF INJUSTICE, THE [DOCTOR WHO MISSING ADVENTURES]  RUSSELL, GARY  C/P:2,P
SHADOW OF WENG-CHIANG, THE [DOCTOR WHO MISSING ADVENTURES]  MCINTEE, DAVID A.  C/P,P
SORCERER’S APPRENTICE, THE [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P:2,P
SPEED OF FLIGHT [DOCTOR WHO MISSING ADVENTURES]  LEONARD, PAUL  C/P,P
STATE OF CHANGE [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P:2,P
SYSTEM SHOCK [DOCTOR WHO MISSING ADVENTURES]  RICHARDS, JUSTIN  C/P:2,P
TIME OF YOUR LIFE [DOCTOR WHO MISSING ADVENTURES]  LYONS, STEVE  C/P:2,P
TWILIGHT OF THE GODS [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P,P
VENUSIAN LULLABY [DOCTOR WHO MISSING ADVENTURES]  LEONARD, PAUL  C/P,P
WELL-MANNERED WAR, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
ALL-CONSUMING FIRE [DOCTOR WHO NEW ADVENTURES]  LANE, ANDY  C/P:2,P
ALSO PEOPLE, THE [DOCTOR WHO NEW ADVENTURES]  AARONOVITCH, BEN  C/P:2,P
BAD THERAPY [DOCTOR WHO NEW ADVENTURES]  JONES, MATTHEW  C/P,P
BIRTHRIGHT [DOCTOR WHO NEW ADVENTURES]  ROBINSON, NIGEL  C/P:2,P
BLOOD HARVEST [DOCTOR WHO NEW ADVENTURES]  DICKS, TERRANCE  C/P:2,P
BLOOD HEAT [DOCTOR WHO NEW ADVENTURES]  MORTIMORE, JIM  C/P,P
CAT’S CRADLE: TIME’S CRUCIBLE [DOCTOR WHO NEW ADVENTURES]  PLATT, MARC  C/P:2,P
CAT’S CRADLE: WARHEAD [DOCTOR WHO NEW ADVENTURES]  CARTMEL, ANDREW  C/P:2,P
CAT’S CRADLE: WITCH MARK [DOCTOR WHO NEW ADVENTURES]  HUNT, ANDREW  C/P:2,P
CHRISTMAS ON A RATIONAL PLANET [DOCTOR WHO NEW ADVENTURES]  MILES, LAWRENCE  P
CONUNDRUM [DOCTOR WHO NEW ADVENTURES]  LYONS, STEVE  C/P:2,P
DAMAGED GOODS [DOCTOR WHO NEW ADVENTURES]  DAVIES, RUSSELL T.  C/P,P
DEATH AND DIPLOMACY [DOCTOR WHO NEW ADVENTURES]  STONE, DAVE  C/P,P
DEATH OF ART, THE [DOCTOR WHO NEW ADVENTURES]  BUCHER-JONES, SIMON  C/P,P
DECEIT [DOCTOR WHO NEW ADVENTURES]  DARVILL-EVANS, PETER  C/P:2,P
DIMENSION RIDERS, THE [DOCTOR WHO NEW ADVENTURES]  BLYTHE, DANIEL  C/P:2,P
DYING DAYS, THE [DOCTOR WHO NEW ADVENTURES]  PARKIN, LANCE  C/P,P
ETERNITY WEEPS [DOCTOR WHO NEW ADVENTURES]  MORTIMORE, JIM  C/P,P
FALLS THE SHADOW [DOCTOR WHO NEW ADVENTURES]  O’MAHONY, DANIEL  C/P:2,P
FIRST FRONTIER [DOCTOR WHO NEW ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
GODENGINE [DOCTOR WHO NEW ADVENTURES]  HINTON, CRAIG  P
HAPPY ENDINGS [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  P
HEAD GAMES [DOCTOR WHO NEW ADVENTURES]  LYONS, STEVE  C/P:2,P
HIGHEST SCIENCE, THE [DOCTOR WHO NEW ADVENTURES]  ROBERTS, GARETH  C/P:2,P
HUMAN NATURE [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
ICEBERG [DOCTOR WHO NEW ADVENTURES]  BANKS, DAVID  C/P,P
INFINITE REQUIEM [DOCTOR WHO NEW ADVENTURES]  BLYTHE, DANIEL  C/P:2,P
JUST WAR [DOCTOR WHO NEW ADVENTURES]  PARKIN, LANCE  C/P:2,P
LEFT-HANDED HUMMINGBIRD, THE [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P:2,P
LEGACY [DOCTOR WHO NEW ADVENTURES]  RUSSELL, GARY  C/P:2,P
LOVE AND WAR [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
LUCIFER RISING [DOCTOR WHO NEW ADVENTURES]  LANE, ANDY/MORTIMORE, JIM  C/P:2,P
LUNGBARROW [DOCTOR WHO NEW ADVENTURES]  PLATT, MARC  P
NIGHTSHADE [DOCTOR WHO NEW ADVENTURES]  GATISS, MARK  C/P:2,P
NO FUTURE [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
ORIGINAL SIN [DOCTOR WHO NEW ADVENTURES]  LANE, ANDY  C/P:2,P
PARASITE [DOCTOR WHO NEW ADVENTURES]  MORTIMORE, JIM  C/P:2,P
PIT, THE [DOCTOR WHO NEW ADVENTURES]  PENSWICK, NEIL  C/P:2,P
RETURN OF THE LIVING DAD [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P,P
ROOM WITH NO DOORS, THE [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P,P
SANCTUARY [DOCTOR WHO NEW ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
SET PIECE [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P,P
SHADOWMIND [DOCTOR WHO NEW ADVENTURES]  BULIS, CHRISTOPHER  C/P:2,P
SHAKEDOWN [DOCTOR WHO NEW ADVENTURES]  DICKS, TERRANCE  C/P:2,P
SKY PIRATES! [DOCTOR WHO NEW ADVENTURES]  STONE, DAVE  C/P:2,P
SLEEPY [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P:2,P
SO VILE A SIN [DOCTOR WHO NEW ADVENTURES]  AARONOVITCH, BEN/ORMAN, KATE  C/P,P
ST ANTHONY’S FIRE [DOCTOR WHO NEW ADVENTURES]  GATISS, MARK  C/P:2,P
STRANGE ENGLAND [DOCTOR WHO NEW ADVENTURES]  MESSINGHAM, SIMON  C/P:2,P
THEATRE OF WAR [DOCTOR WHO NEW ADVENTURES]  RICHARDS, JUSTIN  C/P:2,P
TIMEWYRM: APOCALYPSE [DOCTOR WHO NEW ADVENTURES]  ROBINSON, NIGEL  C/P,P
TIMEWYRM: EXODUS [DOCTOR WHO NEW ADVENTURES]  DICKS, TERRANCE  C/P,P
TIMEWYRM: GENESYS [DOCTOR WHO NEW ADVENTURES]  PEEL, JOHN  C/P:2,P
TIMEWYRM: REVELATION [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
TOY SOLDIERS [DOCTOR WHO NEW ADVENTURES]  LEONARD, PAUL  C/P:2,P
TRAGEDY DAY [DOCTOR WHO NEW ADVENTURES]  ROBERTS, GARETH  C/P:2,P
TRANSIT [DOCTOR WHO NEW ADVENTURES]  AARONOVITCH, BEN  C/P,P
WARCHILD [DOCTOR WHO NEW ADVENTURES]  CARTMEL, ANDREW  C/P:2,P
WARLOCK [DOCTOR WHO NEW ADVENTURES]  CARTMEL, ANDREW  C/P:2,P
WHITE DARKNESS [DOCTOR WHO NEW ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
ZAMPER [DOCTOR WHO NEW ADVENTURES]  ROBERTS, GARETH  C/P:2,P

TITLE  AUTHOR  CODE

ANOTHER GIRL, ANOTHER PLANET [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] DAY, MARTIN/BEECH, LEN

C/P,P

BEIGE PLANET MARS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  PARKIN, LANCE/CLAPHAM, MARK  C/P,P
BEYOND THE SUN [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  JONES, MATTHEW  C/P,P
DEAD ROMANCE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  MILES, LAWRENCE  P

DEAD ROMANCE (WITH ADDITIONAL STORIES)

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] MILES, LAWRENCE

LP

DEADFALL [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RUSSELL, GARY  C/P,P
DOWN [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  MILES, LAWRENCE  C/P,P
DRAGONS’ WRATH [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RICHARDS, JUSTIN  C/P,P
DRY PILGRIMAGE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  LEONARD, PAUL/WALTERS, NICK  C/P,P
GHOST DEVICES [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  BUCHER-JONES, SIMON  C/P,P
JOY DEVICE, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RICHARDS, JUSTIN  P

MARY-SUE EXTRUSION, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] STONE, DAVE

P

MEAN STREETS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  DICKS, TERRANCE  P
MEDUSA EFFECT, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RICHARDS, JUSTIN  P,PA
OBLIVION [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  STONE, DAVE  C/P,P
OH NO IT ISN’T! [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  CORNELL, PAUL  C/P,P

RETURN TO THE FRACTURED PLANET [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] STONE, DAVE

P

SHIP OF FOOLS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  STONE, DAVE  C/P,P

SWORD OF FOREVER, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] MORTIMORE, JIM

C/P,P

TEARS OF THE ORACLE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] RICHARDS, JUSTIN

P

TEMPEST [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  BULIS, CHRISTOPHER  C/P,P

TWILIGHT OF THE GODS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] CLAPHAM, MARK/MILLER, JON DE BURGH

P

WALKING TO BABYLON [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] ORMAN, KATE

P,PA

WHERE ANGELS FEAR [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  LEVENE, REBECCA/WINSTONE, SIMON  C/P,P
SCREAM OF THE SHALKA [@DOCTOR WHO NINTH (ANIMATED) DOCTOR]  CORNELL, PAUL  P
CLOCKWISE MAN, THE [DOCTOR WHO NINTH DOCTOR]  RICHARDS, JUSTIN  H
DEVIANT STRAIN, THE [DOCTOR WHO NINTH DOCTOR]  RICHARDS, JUSTIN  H
MONSTERS INSIDE, THE [DOCTOR WHO NINTH DOCTOR]  COLE, STEPHEN  H
ONLY HUMAN [DOCTOR WHO NINTH DOCTOR]  ROBERTS, GARETH  H
STEALER OF DREAMS, THE [DOCTOR WHO NINTH DOCTOR]  LYONS, STEVE  H
WINNER TAKES ALL [DOCTOR WHO NINTH DOCTOR]  RAYNER, JACQUELINE  H
ALGEBRA OF ICE, THE [DOCTOR WHO PAST DOCTORS]  ROSE, LLOYD  P
AMORALITY TALE [DOCTOR WHO PAST DOCTORS]  BISHOP, DAVID  P
ASYLUM [DOCTOR WHO PAST DOCTORS]  DARVILL-EVANS, PETER  P
ATOM BOMB BLUES [DOCTOR WHO PAST DOCTORS]  CARTMEL, ANDREW  P
BLUE BOX [DOCTOR WHO PAST DOCTORS]  ORMAN, KATE  P
BULLET TIME [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  P
BUNKER SOLDIERS [DOCTOR WHO PAST DOCTORS]  DAY, MARTIN  P
BUSINESS UNUSUAL [DOCTOR WHO PAST DOCTORS]  RUSSELL, GARY  C/P:2,P
BYZANTIUM! [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH  P
CAMPAIGN [DOCTOR WHO PAST DOCTORS]  MORTIMORE, JIM  SR-P
CATASTROPHEA [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  C/P,P
CITY AT WORLD’S END [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  P
COLONY OF LIES, THE [DOCTOR WHO PAST DOCTORS]  BRAKE, COLIN  P
COMBAT ROCK [DOCTOR WHO PAST DOCTORS]  LEWIS, MICK  P
CORPSE MARKER [DOCTOR WHO PAST DOCTORS]  BOUCHER, CHRIS  P
DEADLY REUNION [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE/LETTS, BARRY  P
DEEP BLUE [DOCTOR WHO PAST DOCTORS]  MORRIS, MARK  C/P,P
DEVIL GOBLINS FROM NEPTUNE, THE [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH/DAY, MARTIN  C/P:2,P
DIVIDED LOYALTIES [DOCTOR WHO PAST DOCTORS]  RUSSELL, GARY  P
DREAMS OF EMPIRE [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN  P
DRIFT [DOCTOR WHO PAST DOCTORS]  FORWARD, SIMON A.  P
DYING IN THE SUN [DOCTOR WHO PAST DOCTORS]  MILLER, JON DE BURGH  P
ELEVENTH TIGER, THE [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  P
EMPIRE OF DEATH [DOCTOR WHO PAST DOCTORS]  BISHOP, DAVID  P
EYE OF HEAVEN [DOCTOR WHO PAST DOCTORS]  MORTIMORE, JIM  C/P,P
FACE OF THE ENEMY, THE [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  C/P:2,P
FEAR ITSELF [DOCTOR WHO PAST DOCTORS]  WALLACE, NICK  P
FEAR OF THE DARK [DOCTOR WHO PAST DOCTORS]  BAXENDALE, TREVOR  P
FESTIVAL OF DEATH [DOCTOR WHO PAST DOCTORS]  MORRIS, JONATHAN  C/P,P
FINAL SANCTION, THE [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  P
GRAVE MATTER [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN  P
HEART OF TARDIS [DOCTOR WHO PAST DOCTORS]  STONE, DAVE  P
HERITAGE [DOCTOR WHO PAST DOCTORS]  SMITH, DALE  P
HOLLOW MEN, THE [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH/DAY, MARTIN  C/P,P
ILLEGAL ALIEN [DOCTOR WHO PAST DOCTORS]  TUCKER, MIKE/PERRY, ROBERT  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 47
TITLE  AUTHOR  CODE
IMPERIAL MOON [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  C/P,P
INDEPENDENCE DAY [DOCTOR WHO PAST DOCTORS]  DARVILL-EVANS, PETER  C/P,P
INDESTRUCTIBLE MAN, THE [@DOCTOR WHO PAST DOCTORS]  MESSINGHAM, SIMON  P

INTERFERENCE BOOK ONE: SHOCK TACTIC [DOCTOR WHO PAST DOCTORS] MILES, LAWRENCE

P

ISLAND OF DEATH [DOCTOR WHO PAST DOCTORS]  LETTS, BARRY  P
KING OF TERROR, THE [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH  C/P,P
LAST MAN RUNNING [DOCTOR WHO PAST DOCTORS]  BOUCHER, CHRIS  C/P:2,P
LAST OF THE GADERENE [DOCTOR WHO PAST DOCTORS]  GATISS, MARK  C/P,P
LOVING THE ALIEN [DOCTOR WHO PAST DOCTORS]  TUCKER, MIKE/PERRY, ROBERT  P
MATCH OF THE DAY [@DOCTOR WHO PAST DOCTORS]  BOUCHER, CHRIS  P
MATRIX [DOCTOR WHO PAST DOCTORS]  PERRY, ROBERT/TUCKER, MIKE  C/P,P
MILLENNIUM SHOCK [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN  P
MISSION: IMPRACTICAL [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  C/P,P
MURDER GAME, THE [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  C/P:2,P
PALACE OF THE RED SUN [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  P
PLAYERS [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  C/P,P
PRIME TIME [DOCTOR WHO PAST DOCTORS]  TUCKER, MIKE  P
QUANTUM ARCHANGEL, THE [DOCTOR WHO PAST DOCTORS]  HINTON, CRAIG  P
RAGS [DOCTOR WHO PAST DOCTORS]  LEWIS, MICK  P
RELATIVE DEMENTIAS [DOCTOR WHO PAST DOCTORS]  MICHALOWSKI, MARK  P
ROUNDHEADS, THE [DOCTOR WHO PAST DOCTORS]  GATISS, MARK  C/P:2,P
SALVATION [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  C/P:2,P
SHADOW IN THE GLASS, THE [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN/COLE, STEPHEN  P
SPIRAL SCRATCH [DOCTOR WHO PAST DOCTORS]  RUSSELL, GARY  P
STORM HARVEST [DOCTOR WHO PAST DOCTORS]  PERRY, ROBERT/TUCKER, MIKE  C/P,P
SUNS OF CARESH, THE [DOCTOR WHO PAST DOCTORS]  SAINT, PAUL  P
SUPERIOR BEINGS [DOCTOR WHO PAST DOCTORS]  WALTERS, NICK  P
SYNTHESPIANS[TM] [@DOCTOR WHO PAST DOCTORS]  HINTON, CRAIG  P
TEN LITTLE ALIENS [DOCTOR WHO PAST DOCTORS]  COLE, STEPHEN  P
TIME TRAVELLERS, THE [DOCTOR WHO PAST DOCTORS]  GUERRIER, SIMON  P
TOMB OF VALDEMAR [DOCTOR WHO PAST DOCTORS]  MESSINGHAM, SIMON  P
ULTIMATE TREASURE, THE [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  C/P:2,P
VERDIGRIS [DOCTOR WHO PAST DOCTORS]  MAGRS, PAUL  C/P,P
WAGES OF SIN, THE [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  C/P:2,P
WARMONGER [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  P
WITCH HUNTERS, THE [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  C/P:2,P
WOLFSBANE [DOCTOR WHO PAST DOCTORS]  RAYNER, JACQUELINE  P
WORLD GAME [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  P
ZETA MAJOR [DOCTOR WHO PAST DOCTORS]  MESSINGHAM, SIMON  C/P,P
FEAST OF THE DROWNED, THE [DOCTOR WHO TENTH DOCTOR]  COLE, STEPHEN  H
PRICE OF PARADISE, THE [DOCTOR WHO TENTH DOCTOR]  BRAKE, COLIN  H
RESURRECTION CASKET, THE [DOCTOR WHO TENTH DOCTOR]  RICHARDS, JUSTIN  H
SHOOTING SCRIPTS, THE [DOCTOR WHO TENTH DOCTOR]  BBC BOOKS  R/REF-H
DAUGHTERS OF THE DOLPHIN [DOLPHIN 2]  MEYERS, ROY  C/P:2,P
DESTINY AND THE DOLPHINS [DOLPHIN 3]  MEYERS, ROY  C/P:2,P
DOLPHIN BOY [DOLPHIN 1]  MEYERS, ROY  C/P:2,P

AGENT OF THE TERRAN EMPIRE (ORIGINAL) (BRITISH)

[DOMINIC FLANDRY] ANDERSON, POUL

SR-P

AGENT OF THE TERRAN EMPIRE (REVISED) [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P
CIRCUS OF HELLS, A [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P
ENSIGN FLANDRY [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P
FLANDRY [DOMINIC FLANDRY]  ANDERSON, POUL  P,SFWA-P
FLANDRY OF TERRA [DOMINIC FLANDRY]  ANDERSON, POUL  C/P,P,H
GAME OF EMPIRE, THE [DOMINIC FLANDRY]  ANDERSON, POUL  C/P,P,H
IMPERIAL STARS, THE [DOMINIC FLANDRY]  ANDERSON, POUL  H
KNIGHT OF GHOSTS AND SHADOWS, A [DOMINIC FLANDRY]  ANDERSON, POUL  H:2,P
REBEL WORLDS, THE [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P

REBEL WORLDS, THE [DOMINIC FLANDRY]/ KNIGHT OF GHOSTS AND SHADOWS, A [DOMINIC FLANDRY] ANDERSON, POUL

P

STONE IN HEAVEN, A [DOMINIC FLANDRY]  ANDERSON, POUL  H:2,P
DARK SHORE, THE [@DOMINIONS OF IRTH #1]  LEE, ADAM  C/P,P
SHADOW EATER, THE [@DOMINIONS OF IRTH #2]  LEE, ADAM  P
OCTOBERLAND [@DOMINIONS OF IRTH #3]  LEE, ADAM  H,P
BLACK CASTLE, THE [DON SEBASTIAN 1]  DANIELS, LES  H,P
CITIZEN VAMPIRE [DON SEBASTIAN 3]  DANIELS, LES  P
NO BLOOD SPILLED [DON SEBASTIAN 5]  DANIELS, LES  P,SFWA-P
SILVER SKULL, THE [DON SEBASTIAN 2]  DANIELS, LES  P
YELLOW FOG [DON SEBASTIAN 4]  DANIELS, LES  SR-H,SFWA-P
KNEE-DEEP IN THE DEAD [@DOOM #1]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P,P
HELL ON EARTH [@DOOM #2]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P,P
INFERNAL SKY [@DOOM #3]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P:2,P
ENDGAME [@DOOM #4]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P:2,P

TITLE  AUTHOR  CODE
DIVINE QUEEN, THE [DOOM-QUEST OF ARA-KARN 2]  CORBY, ADAM  P
FORMER KING, THE [DOOM-QUEST OF ARA-KARN 1]  CORBY, ADAM  C/P,P
DOOM STAR [DOOM STAR 1]  MEYERS, RICHARD S.  SR-P,C/P:2
DOOM STAR NUMBER TWO [DOOM STAR 2]  MEYERS, RICHARD S.  SR-P
RETURN TO DOOMSTAR [DOOM STAR]  MEYERS, RICHARD S.  C/P,P
RAMPAGE [@DOOM’S DAY 1]  FINGEROTH, DANNY/FEIN ERIC  P
SABOTAGE [@DOOM’S DAY #2]  ASKEGREN, PIERCE/FINGEROTH, DANNY  C/P,P
DOOMSDAY WARRIOR [@DOOMSDAY WARRIOR #1]  STACY, RYDER  P
RED AMERICA [@DOOMSDAY WARRIOR #2]  STACY, RYDER  P
LAST AMERICAN, THE [@DOOMSDAY WARRIOR #3]  STACY, RYDER  P
BLOODY AMERICA [@DOOMSDAY WARRIOR #4]  STACY, RYDER  P
AMERICA’S LAST DECLARATION [@DOOMSDAY WARRIOR #5]  STACY, RYDER  P
AMERICAN REBELLION [@DOOMSDAY WARRIOR #6]  STACY, RYDER  P
AMERICAN DEFIANCE [@DOOMSDAY WARRIOR #7]  STACY, RYDER  P
AMERICAN GLORY [@DOOMSDAY WARRIOR #8]  STACY, RYDER  P
AMERICA’S ZERO HOUR [@DOOMSDAY WARRIOR #9]  STACY, RYDER  P
AMERICAN EDEN [@DOOMSDAY WARRIOR #11]  STACY, RYDER  P
DEATH AMERICAN STYLE [@DOOMSDAY WARRIOR #12]  STACY, RYDER  P
CRISIS ON DOONA [DOONA]  MCCAFFREY, ANNE/NYE, JODY LYNN  C/P:2,SFWA-P,P
DECISION AT DOONA [DOONA]  MCCAFFREY, ANNE  C/P:2,P
TREATY AT DOONA [DOONA]  MCCAFFREY, ANNE/NYE, JODY LYNN  P
FINAL COMMAND [DOUBLE SPIRAL WAR 3]  NORWOOD, WARREN  P
MIDWAY BETWEEN [DOUBLE SPIRAL WAR 1]  NORWOOD, WARREN  C/P:2,P
POLAR FLEET [DOUBLE SPIRAL WAR 2]  NORWOOD, WARREN  C/P,P
DOWNS-LORD DAWN [DOWNS-LORD 1]  WHITBOURN, JOHN  P
DOWNS-LORD DAY [DOWNS-LORD 2]  WHITBOURN, JOHN  P
COSMIC BOMBER, THE [DR. BONES]  WU, WILLIAM F.  P
DRAGONS OF KOMAKO, THE [DR. BONES]  BETANCOURT, JOHN GREGORY  P
GARUKAN BLOOD [DR. BONES]  WYLDE, THOMAS  P
JOURNEY TO RILLA [DR. BONES]  WYLDE, THOMAS  P
NIGHTMARE WORLD [DR. BONES]  STERN, DAVID  P
SECRET OF THE LONA, THE [DR. BONES 1]  LEIGH, STEPHEN  C/P,P
COLDNESS IN THE BLOOD, A [DRACULA]  SABERHAGEN, FRED  H
DOMINION [DRACULA]  SABERHAGEN, FRED  C/P:2,P
DRACULA TAPE, THE [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
HOLMES-DRACULA FILE, THE [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
MATTER OF TASTE, A [DRACULA]  SABERHAGEN, FRED  C/P,P,H
OLD FRIEND OF THE FAMILY, AN [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
QUESTION OF TIME, A [DRACULA]  SABERHAGEN, FRED  C/P,P,H
SEANCE FOR A VAMPIRE [DRACULA]  SABERHAGEN, FRED  H,P
SHARPNESS ON THE NECK, A [DRACULA]  SABERHAGEN, FRED  H,P
THORN [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
DRACULA RETURNS [@DRACULA #1]  LORY, ROBERT  P
HAND OF DRACULA, THE [@DRACULA #2]  LORY, ROBERT  C/P,P
DRACULA’S BROTHERS [@DRACULA #3]  LORY, ROBERT  C/P,P
DRACULA’S GOLD [@DRACULA #4]  LORY, ROBERT  P
DRUMS OF DRACULA [@DRACULA #5]  LORY, ROBERT  P
WITCHING OF DRACULA, THE [@DRACULA #6]  LORY, ROBERT  P
DRACULA’S DISCIPLE [@DRACULA #8]  LORY, ROBERT  C/P,P
CHALLENGE TO DRACULA [@DRACULA #9]  LORY, ROBERT  P
LORD OF CASTLE BLACK, THE [DRAGAERA ADRILANKHA 2]  BRUST, STEVEN  H
PATHS OF THE DEAD, THE [DRAGAERA ADRILANKHA 1]  BRUST, STEVEN  H:2,P
SETHRA LAVODE [DRAGAERA ADRILANKHA 3]  BRUST, STEVEN  H
BROKEDOWN PALACE [DRAGAERA FENARIO]  BRUST, STEVEN  C/P:3,P
FIVE HUNDRED YEARS AFTER [DRAGAERA KHAAVREN 2]  BRUST, STEVEN  C/P,P,H
PHOENIX GUARDS, THE [DRAGAERA KHAAVREN 1]  BRUST, STEVEN  C/P,P,H
ATHYRA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,SFWA-P:2,P
BOOK OF ATHYRA, THE [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  LP
BOOK OF JHEREG, THE [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  H
BOOK OF TALTOS, THE [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  H:2
DRAGON [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:2,P,H
DZUR [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  H
ISSOLA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P,P,H
JHEREG [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:2,P
ORCA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
PHOENIX [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:2,SFWA-P:2,P
STEVEN BRUST’S JHEREG [DRAGAERA VLAD TALTOS]  ZELENETZ, ALAN/PIERARD, JOHN  CX
TALTOS [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
TECKLA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
YENDI [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
DRAGON BURNING [DRAGON CIRCLE 3]  GARDNER, CRAIG SHAW  H,@P
DRAGON SLEEPING [DRAGON CIRCLE 1]  GARDNER, CRAIG SHAW  H:2,SFWA,P
DRAGON WAKING [DRAGON CIRCLE 2]  GARDNER, CRAIG SHAW  C/P,P,H

TITLE  AUTHOR  CODE
DRAGON COMPANION [DRAGON COMPANION 1]  CALLANDER, DON  C/P:2,P
DRAGON RESCUE [DRAGON COMPANION 2]  CALLANDER, DON  C/P,P
DRAGON DELASANGRE, THE [DRAGON DELASANGRE 1]  TROOP, ALAN F.  P
DRAGON DELASANGRE, THE (PROOF) [DRAGON DELASANGRE 1]  TROOP, ALAN F.  SR-LP
DRAGON MOON [DRAGON DELASANGRE 2]  TROOP, ALAN F.  P
HOST OF DRAGONS, A [DRAGON DELASANGRE 4]  TROOP, ALAN F.  P
SEADRAGON’S DAUGHTER, THE [DRAGON DELASANGRE 3]  TROOP, ALAN F.  P
CAST OF FATE [DRAGON DICE]  VARNEY, ALLEN  SR-P
AERIE [DRAGON JOUSTERS 4]  LACKEY, MERCEDES  H
ALTA [DRAGON JOUSTERS 2]  LACKEY, MERCEDES  H,P
JOUST [DRAGON JOUSTERS 1]  LACKEY, MERCEDES  H:2
SANCTUARY [DRAGON JOUSTERS 3]  LACKEY, MERCEDES  H:2
IN THE HALL OF THE DRAGON KING [@DRAGON KING #1]  LAWHEAD, STEPHEN R.  C/P,P
WARLORDS OF NIN, THE [@DRAGON KING #2]  LAWHEAD, STEPHEN R.  C/P,P
SWORD AND THE FLAME, THE [@DRAGON KING #3]  LAWHEAD, STEPHEN R.  C/P:2,SFWA-P,P
DRAGON AND THE DJINN, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  C/P,P
DRAGON AND THE FAIR MAID OF KENT, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  H
DRAGON AND THE GEORGE, THE [DRAGON KNIGHT 1]  DICKSON, GORDON R.  C/P,P,H:2
DRAGON AND THE GNARLY KING, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  C/P,P,H
DRAGON AT WAR, THE [DRAGON KNIGHT 4]  DICKSON, GORDON R.  H,SFWA,SFWA-P:2,P
DRAGON IN LYONESSE, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  H
DRAGON KNIGHT, THE [DRAGON KNIGHT 2]  DICKSON, GORDON R.  H:2,SFWA,P
DRAGON ON THE BORDER, THE [DRAGON KNIGHT 3]  DICKSON, GORDON R.  H:2,SFWA-P:2,P
DRAGON, THE EARL, AND THE TROLL, THE [DRAGON KNIGHT 5]  DICKSON, GORDON R.  H,P
ICE DRAGON, THE [@DRAGON KNIGHTS #3]  D’ARC, BIANCA  LP
GLASS DRAGON, THE [DRAGON NIMBUS 1]  RADFORD, IRENE  C/P:2,P
LONELIEST MAGICIAN, THE [DRAGON NIMBUS 3]  RADFORD, IRENE  P
PERFECT PRINCESS, THE [DRAGON NIMBUS 2]  RADFORD, IRENE  C/P:2,P
WIZARD’S TREASURE, THE [DRAGON NIMBUS 4]  RADFORD, IRENE  P
DRAGON’S TOUCHSTONE, THE [DRAGON NIMBUS HISTORY 1]  RADFORD, IRENE  C/P,P
LAST BATTLEMAGE, THE [DRAGON NIMBUS HISTORY 2]  RADFORD, IRENE  P
RENEGADE DRAGON, THE [DRAGON NIMBUS HISTORY 3]  RADFORD, IRENE  P
DRAGON CAULDRON [DRAGON OF THE LOST SEA 3]  YEP, LAURENCE  SFWA
DRAGON PRINCE [DRAGON PRINCE 1]  RAWN, MELANIE  C/P,P
STAR SCROLL, THE [DRAGON PRINCE 2]  RAWN, MELANIE  C/P:2,P
SUN-RUNNER’S FIRE [DRAGON PRINCE 3]  RAWN, MELANIE  C/P:2,P
BOOK OF EARTH, THE [@DRAGON QUARTET #1]  KELLOGG, MARJORIE BRADLEY  C/P,P
BOOK OF WATER, THE [@DRAGON QUARTET #2]  KELLOGG, MARJORIE BRADLEY  C/P,P
BOOK OF FIRE, THE [@DRAGON QUARTET #3]  KELLOGG, MARJORIE BRADLEY  C/P:2,P
BOOK OF AIR, THE [@DRAGON QUARTET #4]  KELLOGG, MARJORIE BRADLEY  C/P,P
DRAGON TOKEN, THE [DRAGON STAR 2]  RAWN, MELANIE  H:2,P
SKYBOWL [DRAGON STAR 3]  RAWN, MELANIE  C/P,SFWA-P,H,P,SFWA
STRONGHOLD [DRAGON STAR 1]  RAWN, MELANIE  C/P,P,H
TOUCHED BY VENOM [@DRAGON TEMPLE #1]  CROSS, JANINE  LP
SHADOWED BY WINGS [@DRAGON TEMPLE #2]  CROSS, JANINE  LP
FORGED BY FIRE [@DRAGON TEMPLE #3]  CROSS, JANINE  LP
STONE OF STARS, THE [@DRAGON THRONE #1]  BAIRD, ALISON  LP
EMPIRE OF THE STARS, THE [@DRAGON THRONE #2]  BAIRD, ALISON  LP
ARCHONS OF THE STARS, THE [@DRAGON THRONE #3]  BAIRD, ALISON  LP
DRAGONS PAST [DRAGON WARS 2]  GENTILE, GARY  C/P,P
NO FUTURE FOR DRAGONS [DRAGON WARS 3]  GENTILE, GARY  P
TIME FOR DRAGONS, A [DRAGON WARS 1]  GENTILE, GARY  C/P:2,P
DRAGON AND HERDSMAN [DRAGONBACK 4]  ZAHN, TIMOTHY  H
DRAGON AND THIEF [DRAGONBACK 1]  ZAHN, TIMOTHY  H
DRAGONBACK BARGAIN [DRAGONBACK 1,2,3]  ZAHN, TIMOTHY  H
DARK GLORY WAR, THE [DRAGONCROWN WAR]  STACKPOLE, MICHAEL A.  P
FORTRESS DRACONIS [DRAGONCROWN WAR 1]  STACKPOLE, MICHAEL A.  LP
GRAND CRUSADE, THE [DRAGONCROWN WAR 3]  STACKPOLE, MICHAEL A.  LP
GRAND CRUSADE, THE (PROOF) [DRAGONCROWN WAR 3]  STACKPOLE, MICHAEL A.  SR-LP
WHEN DRAGONS RAGE [DRAGONCROWN WAR 2]  STACKPOLE, MICHAEL A.  C/P,LP
BEST OF TALES, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
BEST OF TALES II, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  C/PA,PA
DRAGONLANCE: THE ANNOTATED CHRONICLES [DRAGONLANCE 1,2,3]  WEIS, MARGARET/HICKMAN, TRACY  LP
DRAGONS AT WAR, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS IN THE ARCHIVES [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS OF CHAOS, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS OF KRYNN, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS OF TIME [DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
PLAYERS OF GILEAN, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
REBELS AND TYRANTS [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
RELICS AND OMENS [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
SEARCH FOR POWER, THE [@DRAGONLANCE]  WEIS, MARGARET  PA
SECOND GENERATION, THE [DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  C/P,P,H

TITLE  AUTHOR  CODE
TALES OF UNCLE TRAPSPRINGER [@DRAGONLANCE]  MCKEONE, DIXIE LEE  P
CONUNDRUM [@DRAGONLANCE AGE OF MORTALS]  CROOK, JEFF  P
DARK THANE [@DRAGONLANCE AGE OF MORTALS]  CROOK, JEFF  P
LAKE OF DEATH, THE [@DRAGONLANCE AGE OF MORTALS]  RABE, JEAN  P
LIONESS, THE [@DRAGONLANCE AGE OF MORTALS]  BERBERICK, NANCY VARIAN  P
PRISONER OF HAVEN [@DRAGONLANCE AGE OF MORTALS]  BERBERICK, NANCY VARIAN  P
WIZARDS’ CONCLAVE [@DRAGONLANCE AGE OF MORTALS]  NILES, DOUGLAS  P
CHILDREN OF THE PLAINS [@DRAGONLANCE BARBARIANS #1]  THOMPSON, PAUL B./COOK, TONYA C.  P
BROTHER OF THE DRAGON [@DRAGONLANCE BARBARIANS #2]  THOMPSON, PAUL B./COOK, TONYA C.  P
SISTER OF THE SWORD [@DRAGONLANCE BARBARIANS #3]  THOMPSON, PAUL B./COOK, TONYA C.  P
DEZRA’S QUEST [@DRAGONLANCE BRIDGES OF TIME]  PIERSON, CHRIS  P
SILVER STAIR, THE [@DRAGONLANCE BRIDGES OF TIME]  RABE, JEAN  P
ALIEN SEA, THE [@DRAGONLANCE CHAMPIONS]  SOULBAN, LUCIEN  P
GREAT WHITE WYRM, THE [@DRAGONLANCE CHAMPIONS]  ARCHER, PETER  P
SAVING SOLACE [@DRAGONLANCE CHAMPIONS]  CLARK, DOUGLAS W.  P
DOOM BRIGADE, THE [@DRAGONLANCE CHAOS WAR 1]  WEIS, MARGARET/PERRIN, DON  P
LAST THANE, THE [@DRAGONLANCE CHAOS WAR 2]  NILES, DOUGLAS  P
PUPPET KING, THE [@DRAGONLANCE CHAOS WAR 4]  NILES, DOUGLAS  P
REAVERS OF THE BLOOD SEA [@DRAGONLANCE CHAOS WAR 5]  KNAAK, RICHARD A.  P
SIEGE OF MT. NEVERMIND, THE [@DRAGONLANCE CHAOS WAR 6]  RYAN, FERGUS  P
CLANDESTINE CIRCLE, THE [@DRAGONLANCE CHRONICLES]  HERBERT, MARY H.  P

DRAGONLANCE: THE ANNOTATED LEGENDS [DRAGONLANCE CHRONICLES 1,2,3] WEIS, MARGARET/HICKMAN, TRACY

LP

DRAGONS OF SUMMER FLAME [DRAGONLANCE CHRONICLES 4]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DRAGONS OF AUTUMN TWILIGHT [@DRAGONLANCE CHRONICLES #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DRAGONS OF WINTER NIGHT [@DRAGONLANCE CHRONICLES #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DRAGONS OF SPRING DAWNING [@DRAGONLANCE CHRONICLES #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
CITADEL, THE [@DRAGONLANCE CLASSICS]  KNAAK, RICHARD A.  P
DALAMAR THE DARK [@DRAGONLANCE CLASSICS]  BERBERICK, NANCY VARIAN  P
INHERITANCE, THE [@DRAGONLANCE CLASSICS]  BERBERICK, NANCY VARIAN  P
MURDER IN TARSIS [@DRAGONLANCE CLASSICS]  ROBERTS, JOHN MADDOX  P
DRAGON ISLES, THE [@DRAGONLANCE CROSSROADS]  SULLIVAN, STEPHEN D.  P
DRAGON’S BLUFF [@DRAGONLANCE CROSSROADS]  HERBERT, MARY H.  P
MIDDLE OF NOWHERE, THE [@DRAGONLANCE CROSSROADS]  THOMPSON, PAUL B.  P
THIEVES’ GUILD, THE [@DRAGONLANCE CROSSROADS]  CROOK, JEFF  P
AMBER AND ASHES [@DRAGONLANCE DARK DISCIPLE #1]  WEIS, MARGARET  H
AMBER AND IRON [@DRAGONLANCE DARK DISCIPLE #2]  WEIS, MARGARET  H
NIGHT OF THE EYE [@DRAGONLANCE DEFENDERS OF MAGIC #1]  KIRCHOFF, MARY  C/P,P
MEDUSA PLAGUE, THE [@DRAGONLANCE DEFENDERS OF MAGIC #2]  KIRCHOFF, MARY  P
SEVENTH SENTINEL, THE [@DRAGONLANCE DEFENDERS OF MAGIC #3]  KIRCHOFF, MARY  P
DOWNFALL [@DRAGONLANCE DHAMON #1]  RABE, JEAN  H,P
BETRAYAL [@DRAGONLANCE DHAMON #2]  RABE, JEAN  C/P,P
REDEMPTION [@DRAGONLANCE DHAMON #3]  RABE, JEAN  H
COVENANT OF THE FORGE, THE [@DRAGONLANCE DWARVEN NATIONS #1]  PARKINSON, DAN  C/P,P
HAMMER AND AXE [@DRAGONLANCE DWARVEN NATIONS #2]  PARKINSON, DAN  P
SWORDSHEATH SCROLL, THE [@DRAGONLANCE DWARVEN NATIONS #3]  PARKINSON, DAN  P
SANCTUARY [@DRAGONLANCE ELVEN EXILES #1]  THOMPSON, PAUL B./COOK, TONYA C.  P
ALLIANCES [@DRAGONLANCE ELVEN EXILES #2]  THOMPSON, PAUL B./COOK, TONYA C.  P
FIRSTBORN [@DRAGONLANCE ELVEN NATIONS #1]  THOMPSON, PAUL B./CARTER, TONYA R.  P
KINSLAYER WARS, THE [@DRAGONLANCE ELVEN NATIONS #2]  NILES, DOUGLAS  P
QUALINESTI, THE [@DRAGONLANCE ELVEN NATIONS #3]  THOMPSON, PAUL B./CARTER, TONYA R.  P
WARRIOR’S JOURNEY, A [@DRAGONLANCE ERGOTH #1]  THOMPSON, PAUL B./COOK, TONYA C.  C/P,P
WIZARD’S FATE, THE [@DRAGONLANCE ERGOTH #2]  THOMPSON, PAUL B./COOK, TONYA C.  P
HERO’S JUSTICE, A [@DRAGONLANCE ERGOTH #3]  THOMPSON, PAUL B./COOK, TONYA C.  P
DAWNING OF A NEW AGE, THE [@DRAGONLANCE FIFTH AGE 1]  RABE, JEAN  P
DAY OF THE TEMPEST, THE [@DRAGONLANCE FIFTH AGE]  RABE, JEAN  P
EVE OF THE MAELSTROM, THE [@DRAGONLANCE FIFTH AGE]  RABE, JEAN  P
HEROES AND FOOLS [@DRAGONLANCE FIFTH AGE]  WEIS, MARGARET/HICKMAN, TRACY  C/PA,PA
WARRIOR’S HEART [@DRAGONLANCE GOODLUND #1]  SULLIVAN, STEPHEN D.  LP
STORMBLADE [DRAGONLANCE HEROES 2]  BERBERICK, NANCY VARIAN  C/P:2,P
WEASEL’S LUCK [@DRAGONLANCE HEROES 3]  WILLIAMS, MICHAEL  C/P:2,P
LEGEND OF HUMA, THE [@DRAGONLANCE HEROES #1]  KNAAK, RICHARD A.  C/P:2,P
KAZ THE MINOTAUR [@DRAGONLANCE HEROES II #1]  KNAAK, RICHARD A.  C/P:2,P
GATES OF THORBARDIN, THE [@DRAGONLANCE HEROES II #2]  PARKINSON, DAN  C/P:2,P
GALEN BEKNIGHTED [@DRAGONLANCE HEROES II #3]  WILLIAMS, MICHAEL  C/P,P
MESSENGER, THE [@DRAGONLANCE ICEWALL #1]  NILES, DOUGLAS  P
GOLDEN ORB, THE [@DRAGONLANCE ICEWALL #2]  NILES, DOUGLAS  P
WINTERHEIM [@DRAGONLANCE ICEWALL #3]  NILES, DOUGLAS  P
DRACONIAN MEASURES [@DRAGONLANCE KANG’S REGIMENT #2]  PERRIN, DON/WEIS, MARGARET  P
CHOSEN OF THE GODS [@DRAGONLANCE KINGPRIEST #1]  PIERSON, CHRIS  P
DIVINE HAMMER [@DRAGONLANCE KINGPRIEST #2]  PIERSON, CHRIS  C/P,P
SACRED FIRE [@DRAGONLANCE KINGPRIEST #3]  PIERSON, CHRIS  P
TIME OF THE TWINS [@DRAGONLANCE LEGENDS #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 51
TITLE  AUTHOR  CODE
WAR OF THE TWINS [@DRAGONLANCE LEGENDS #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
TEST OF THE TWINS [@DRAGONLANCE LEGENDS #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
CITY OF THE LOST [@DRAGONLANCE LINSHA #1]  HERBERT, MARY H.  P
FLIGHT OF THE FALLEN [@DRAGONLANCE LINSHA #2]  HERBERT, MARY H.  P
RETURN OF THE EXILE [@DRAGONLANCE LINSHA #3]  HERBERT, MARY H.  P

DRAGONS OF THE DWARVEN DEPTHS [@DRAGONLANCE LOST CHRONICLES #1] WEIS, MARGARET/HICKMAN, TRACY

H

DRAGONS OF THE HIGHLORD SKIES [@DRAGONLANCE LOST CHRONICLES #2] WEIS, MARGARET/HICKMAN, TRACY

H

IRDA, THE [DRAGONLANCE LOST HISTORIES 2]  BAKER, LINDA P.  C/P,P
KAGONESTI, THE [@DRAGONLANCE LOST HISTORIES #1]  NILES, DOUGLAS  P
DARGONESTI, THE [@DRAGONLANCE LOST HISTORIES #3]  THOMPSON, PAUL B./COOK, TONYA C.  P
LAND OF THE MINOTAURS [@DRAGONLANCE LOST HISTORIES #4]  KNAAK, RICHARD A.  P
GULLY DWARVES, THE [@DRAGONLANCE LOST HISTORIES #5]  PARKINSON, DAN  P
DRAGONS, THE [@DRAGONLANCE LOST HISTORIES #6]  NILES, DOUGLAS  P
VINAS SOLAMNUS [@DRAGONLANCE LOST LEGENDS #1]  KING, J. ROBERT  P
FISTANDANTILUS REBORN [@DRAGONLANCE LOST LEGENDS #2]  NILES, DOUGLAS  P
KINDRED SPIRITS [DRAGONLANCE MEETINGS 1]  ANTHONY, MARK/PORATH, ELLEN  C/P:2,P
OATH AND THE MEASURE, THE [@DRAGONLANCE MEETINGS 4]  WILLIAMS, MICHAEL  P
STEEL AND STONE [@DRAGONLANCE MEETINGS 5]  PORATH, ELLEN  C/P,P
WANDERLUST [@DRAGONLANCE MEETINGS 2]  KIRCHOFF, MARY/WINTER, STEVE  C/P:2,P
DARK HEART [@DRAGONLANCE MEETINGS #3]  DANIELL, TINA  C/P:2,P
COMPANIONS, THE [@DRAGONLANCE MEETINGS #6]  DANIELL, TINA  C/P,P
NIGHT OF BLOOD [@DRAGONLANCE MINOTAUR WARS #1]  KNAAK, RICHARD A.  P
TIDES OF BLOOD [@DRAGONLANCE MINOTAUR WARS #2]  KNAAK, RICHARD A.  P
EMPIRE OF BLOOD [@DRAGONLANCE MINOTAUR WARS #3]  KNAAK, RICHARD A.  H

TEMPLE OF THE DRAGONSLAYER [@DRAGONLANCE NEW ADVENTURES #1/ DRAGONLANCE SPELLBINDER QUARTET 1] WAGGONER, TIM

LP

DYING KINGDOM, THE [@DRAGONLANCE NEW ADVENTURES #2/ DRAGONLANCE SPELLBINDER QUARTET 2] SULLIVAN, STEPHEN D.

LP

DRAGON WELL, THE [@DRAGONLANCE NEW ADVENTURES #3/ DRAGONLANCE SPELLBINDER QUARTET 3] WILLIS, DAN

LP

RETURN OF THE SORCERESS [@DRAGONLANCE NEW ADVENTURES #4/ DRAGONLANCE SPELLBINDER QUARTET 4] WAGGONER, TIM

LP

DRAGON SWORD [@DRAGONLANCE NEW ADVENTURES #5/ DRAGONLANCE DRAGON QUARTET 1] SOESBEE, REE

LP

DRAGON DAY [@DRAGONLANCE NEW ADVENTURES #6/ DRAGONLANCE DRAGON QUARTET 2] BROWN, STAN

LP

DRAGON KNIGHT [@DRAGONLANCE NEW ADVENTURES #7/ DRAGONLANCE DRAGON QUARTET 3] WILLIS, DAN

LP

DRAGON SPELL [@DRAGONLANCE NEW ADVENTURES #8/ DRAGONLANCE DRAGON QUARTET 4] SAMPSON, JEFF

LP

PILLAR OF FLAME, THE [@DRAGONLANCE NEW ADVENTURES ELEMENTS #1] SOESBEE, REE

LP

CROWN OF THIEVES [@DRAGONLANCE NEW ADVENTURES ELIDOR #1]  SOESBEE, REE  LP
CRYSTAL CHALICE, THE [@DRAGONLANCE NEW ADVENTURES ELIDOR #2]  SOESBEE, REE  LP
CITY OF FORTUNE [@DRAGONLANCE NEW ADVENTURES ELIDOR #3]  SOESBEE, REE  LP

WAYWARD WIZARD, THE [@DRAGONLANCE NEW ADVENTURES SUNCATCHER #1] SAMPSON, JEFF

LP

EBONY EYE, THE [@DRAGONLANCE NEW ADVENTURES SUNCATCHER #2]  SAMPSON, JEFF  LP
WIZARD’S CURSE [@DRAGONLANCE NEW ADVENTURES TRINISTYR #1]  WOODS, CHRISTINA  LP

WIZARD’S BETRAYAL [@DRAGONLANCE NEW ADVENTURES TRINISTYR #2] SAMPSON, JEFF

LP

WIZARD’S RETURN [@DRAGONLANCE NEW ADVENTURES TRINISTYR #3]  WILLIS, DAN  LP
BROTHERS MAJERE [DRAGONLANCE PRELUDES 3]  STEIN, KEVIN  C/P:2,P
DARKNESS & LIGHT [DRAGONLANCE PRELUDES 1]  THOMPSON, PAUL B./CARTER, TONYA R.  C/P:2,P
KENDERMORE [@DRAGONLANCE PRELUDES 2]  KIRCHOFF, MARY  C/P:2,P
FLINT THE KING [@DRAGONLANCE PRELUDES II 2]  KIRCHOFF, MARY/NILES, DOUGLAS  C/P,P
RIVERWIND THE PLAINSMAN [DRAGONLANCE PRELUDES II 1]  THOMPSON, PAUL B./CARTER, TONYA R.  C/P,P
TANIS, THE SHADOW YEARS [DRAGONLANCE PRELUDES II 3]  SIEGEL, BARBARA/SIEGEL, SCOTT  C/P:2,P
BROTHERS IN ARMS [@DRAGONLANCE RAISTLIN 2]  WEIS, MARGARET/PERRIN, DON  P
SOULFORGE, THE [@DRAGONLANCE RAISTLIN 1]  WEIS, MARGARET  P
LORD OF THE ROSE [@DRAGONLANCE RISE OF SOLAMNIA #1]  NILES, DOUGLAS  P
CROWN AND THE SWORD, THE [@DRAGONLANCE RISE OF SOLAMNIA #2]  NILES, DOUGLAS  P

MEASURE AND THE TRUTH, THE [@DRAGONLANCE RISE OF SOLAMNIA #3] NILES, DOUGLAS

P

BLADES OF THE TIGER [@DRAGONLANCE TALADAS #1]  PIERSON, CHRIS  P
TRAIL OF THE BLACK WYRM [@DRAGONLANCE TALADAS #2]  PIERSON, CHRIS  P
SHADOW OF THE FLAME [@DRAGONLANCE TALADAS #3]  PIERSON, CHRIS  P
MAGIC OF KRYNN, THE [@DRAGONLANCE TALES #1]  WEIS, MARGARET/HICKMAN, TRACY  PA,C/PA
KENDER, GULLY DWARVES, AND GNOMES [@DRAGONLANCE TALES #2]  WEIS, MARGARET/HICKMAN, TRACY  PA,C/PA
LOVE AND WAR [@DRAGONLANCE TALES #3]  WEIS, MARGARET/HICKMAN, TRACY  PA,C/PA
REIGN OF ISTAR, THE [@DRAGONLANCE TALES II #1]  TSR  C/PA,PA

The MITSFS Seriesdex  September 24, 2007  Page 52
TITLE  AUTHOR  CODE
CATACLYSM, THE [@DRAGONLANCE TALES II #2]  TSR  C/PA,PA
WAR OF THE LANCE, THE [@DRAGONLANCE TALES II #3]  TSR  C/PA,PA
LORD TOEDE [DRAGONLANCE VILLAINS 5]  GRUBB, JEFF  P
BLACK WING, THE [@DRAGONLANCE VILLAINS #2]  KIRCHOFF, MARY  C/P,P
EMPEROR OF ANSALON [@DRAGONLANCE VILLAINS #3]  NILES, DOUGLAS  P
HEDERICK THE THEOCRAT [@DRAGONLANCE VILLAINS #4]  SEVERSON, ELLEN DODGE  C/P,P
DARK QUEEN, THE [@DRAGONLANCE VILLAINS #6]  WILLIAMS, MICHAEL/WILLIAMS, TERI  P
BEFORE THE MASK [@DRAGONLANCE VILLIANS #1]  WILLIAMS, MICHAEL/WILLIAMS, TERI  C/P,P

BERTREM’S GUIDE TO THE WAR OF SOULS, VOL. ONE [@DRAGONLANCE WAR OF SOULS] CROOK, JEFF/GRUBBER, JOHN/

HERBERT, MARY H./

BERBERICK, NANCY VARIAN

P

DRAGONS OF A FALLEN SUN [DRAGONLANCE WAR OF SOULS 1]  WEIS, MARGARET/HICKMAN, TRACY  H
DRAGONS OF A LOST STAR [DRAGONLANCE WAR OF SOULS 2]  WEIS, MARGARET/HICKMAN, TRACY  H
DRAGONS OF A VANISHD MOON [DRAGONLANCE WAR OF SOULS 3]  WEIS, MARGARET/HICKMAN, TRACY  H
SEARCH FOR MAGIC, THE [DRAGONLANCE WAR OF SOULS]  WEIS, MARGARET/HICKMAN, TRACY  PA
KNIGHTS OF THE CROWN [@DRAGONLANCE WARRIORS #1]  GREEN, ROLAND J.  P
MAQUESTA KAR-THON [@DRAGONLANCE WARRIORS #2]  DANIELL, TINA  P
KNIGHTS OF THE SWORD [@DRAGONLANCE WARRIORS #3]  GREEN, ROLAND J.  P
THEROS IRONFELD [@DRAGONLANCE WARRIORS #4]  PERRIN, DON  P
KNIGHTS OF THE ROSE [@DRAGONLANCE WARRIORS #5]  GREEN, ROLAND J.  P
LORD SOTH [@DRAGONLANCE WARRIORS #6]  VAN BELKOM, EDO  P
WAYWARD KNIGHTS, THE [@DRAGONLANCE WARRIORS #7]  GREEN, ROLAND J.  P
DRAGON AND PHOENIX [DRAGONLORD LINDEN 2]  BERTIN, JOANNE  H:2
DRAGON AND PHOENIX (PROOF) [DRAGONLORD LINDEN 2]  BERTIN, JOANNE  SR-LP
LAST DRAGONLORD, THE [DRAGONLORD LINDEN 1]  BERTIN, JOANNE  H,P
STORM OF WINGS [@DRAGONMASTER #1]  BUNCH, CHRIS  LP
KNIGHTHOOD OF THE DRAGON [@DRAGONMASTER #2]  BUNCH, CHRIS  LP
LAST BATTLE, THE [@DRAGONMASTER #3]  BUNCH, CHRIS  LP
CRYSTAL DRAGON, THE [DRAGONREALM 5]  KNAAK, RICHARD A.  C/P,P
FIREDRAKE [DRAGONREALM 1]  KNAAK, RICHARD A.  C/P:2,P
HORSE KING, THE [DRAGONREALM 7]  KNAAK, RICHARD A.  P
ICEDRAGON [DRAGONREALM 2]  KNAAK, RICHARD A.  C/P,P
SHADOW STEED [DRAGONREALM 4]  KNAAK, RICHARD A.  P
WOLFHELM [DRAGONREALM 3]  KNAAK, RICHARD A.  P
CHILDREN OF THE DRAKE [DRAGONREALM ORIGINS 2]  KNAAK, RICHARD A.  P
DRAGON TOME [DRAGONREALM ORIGINS 3]  KNAAK, RICHARD A.  P
SHROUDED REALM, THE [DRAGONREALM ORIGINS 1]  KNAAK, RICHARD A.  P
AGAINST THE TIDE [DRAGONS 3]  RINGO, JOHN  H
EAST OF THE SUN, WEST OF THE MOON [DRAGONS 4]  RINGO, JOHN  H
EMERALD SEA [DRAGONS 2]  RINGO, JOHN  H
THERE WILL BE DRAGONS (HARDCOVER AND CD) [DRAGONS 1]  RINGO, JOHN  H:2
THERE WILL BE DRAGONS (PROOF) [DRAGONS 1]  RINGO, JOHN  SR-LP
DRAGON’S TREASURE [DRAGON’S COUNTY 2]  LYNN, ELIZABETH A.  H
DRAGON’S WINTER [DRAGON’S COUNTY 1]  LYNN, ELIZABETH A.  C/P,P
PRINCE IN CAMELOT, A [DRAGON’S HEIRS 3]  JONES, COURTWAY  P
DRAGONS OF THE CUYAHOGA, THE [DRAGONS OF THE CUYAHOGA 1]  SWANN, S. ANDREW  C/P,P
DWARVES OF WHISKEY ISLAND [DRAGONS OF THE CUYAHOGA 2]  SWANN, S. ANDREW  P
DRAGONKNIGHT [DRAGONSPELL 3]  PAUL, DONITA K.  LP
DRAGONQUEST [DRAGONSPELL 2]  PAUL, DONITA K.  LP
DRAGONSPELL [DRAGONSPELL 1]  PAUL, DONITA K.  LP

DAY OF THE DRAGONSTAR [DRAGONSTAR 1] BISCHOFF, DAVID F./

MONTELEONE, THOMAS F.

C/P:2,P

DRAGONSTAR DESTINY [DRAGONSTAR 3] BISCHOFF, DAVID F./

MONTELEONE, THOMAS F.

C/P,P

NIGHT OF THE DRAGONSTAR [DRAGONSTAR 2] BISCHOFF, DAVID F./

MONTELEONE, THOMAS F.

C/P:2,P

DRAGON DEATH [DRAGONSWORD 3]  BAUDINO, GAEL  C/P,SFWA-P,P
DRAGONSWORD [DRAGONSWORD 1]  BAUDINO, GAEL  C/P:2,P
DUEL OF DRAGONS [DRAGONSWORD 2]  BAUDINO, GAEL  C/P,SFWA-P:2,P
TOWERS OF REXOR, THE [@DRAGONTALES #4]  VILOTT, RHONDI  P
MASTER OF DRAGONS [DRAGONVARLD 3]  WEIS, MARGARET  H
MISTRESS OF DRAGONS [DRAGONVARLD 1]  WEIS, MARGARET  H
DRAKAS! [DRAKA]  STIRLING, S. M.  C/PA,PA
DRAKON [DRAKA 4]  STIRLING, S. M.  C/P:2,P
MARCHING THROUGH GEORGIA [DRAKA 1]  STIRLING, S. M.  C/P:2,SFWA-P,P
STONE DOGS, THE [DRAKA 3]  STIRLING, S. M.  C/P,P
UNDER THE YOKE [DRAKA 2]  STIRLING, S. M.  C/P,SFWA-P,P
DOMINATION, THE [DRAKA (REVISED) 1,2,3]  STIRLING, S. M.  H
CROWN JEWELS, THE [DRAKE MAIJSTRAL 1]  WILLIAMS, WALTER JON  C/P:2,P
HOUSE OF SHARDS [DRAKE MAIJSTRAL 2]  WILLIAMS, WALTER JON  P,SFWA-P
ROCK OF AGES [DRAKE MAIJSTRAL 3]  WILLIAMS, WALTER JON  H,SFWA
ALLIES OF ANTARES [DRAY PRESCOT 26]  AKERS, ALAN BURT  P
ARENA OF ANTARES [DRAY PRESCOT 7]  AKERS, ALAN BURT  P

TITLE  AUTHOR  CODE
ARMADA OF ANTARES [DRAY PRESCOT 11]  AKERS, ALAN BURT  C/P,P
AVENGER OF ANTARES [DRAY PRESCOT 10]  AKERS, ALAN BURT  P
BEASTS OF ANTARES [DRAY PRESCOT 23]  AKERS, ALAN BURT  P
BLADESMAN OF ANTARES [DRAY PRESCOT 9]  AKERS, ALAN BURT  P
CAPTIVE SCORPIO [DRAY PRESCOT 17]  AKERS, ALAN BURT  P
DELIA OF VALLIA [DRAY PRESCOT 28]  AKERS, ALAN BURT  P
FIRES OF SCORPIO [DRAY PRESCOT 29]  AKERS, ALAN BURT  P
FLIERS OF ANTARES [DRAY PRESCOT 8]  AKERS, ALAN BURT  P
FORTUNE FOR KREGEN, A [DRAY PRESCOT 21]  AKERS, ALAN BURT  P
GOLDEN SCORPIO [DRAY PRESCOT 18]  AKERS, ALAN BURT  P
KROZAIR OF KREGEN [DRAY PRESCOT 14]  AKERS, ALAN BURT  P
LEGIONS OF ANTARES [DRAY PRESCOT 25]  AKERS, ALAN BURT  P
LIFE FOR KREGEN, A [DRAY PRESCOT 19]  AKERS, ALAN BURT  P
MANHOUNDS OF ANTARES [DRAY PRESCOT 6]  AKERS, ALAN BURT  C/P,P
MASKS OF SCORPIO [DRAY PRESCOT 31]  AKERS, ALAN BURT  C/P,P
MAZES OF SCORPIO [DRAY PRESCOT 27]  AKERS, ALAN BURT  P
OMENS OF KREGEN [DRAY PRESCOT 36]  AKERS, ALAN BURT  P
PRINCE OF SCORPIO [DRAY PRESCOT 5]  AKERS, ALAN BURT  C/P,P
REBEL OF ANTARES [DRAY PRESCOT 24]  AKERS, ALAN BURT  P
RENEGADE OF KREGEN [DRAY PRESCOT 13]  AKERS, ALAN BURT  P
SAVAGE SCORPIO [DRAY PRESCOT 16]  AKERS, ALAN BURT  C/P:2,P
SECRET SCORPIO [DRAY PRESCOT 15]  AKERS, ALAN BURT  P
SEG THE BOWMAN [DRAY PRESCOT 32]  AKERS, ALAN BURT  C/P,P
STORM OVER VALLIA [DRAY PRESCOT 35]  AKERS, ALAN BURT  P
SUNS OF SCORPIO, THE [DRAY PRESCOT 2]  AKERS, ALAN BURT  C/P,P
SWORD FOR KREGEN, A [DRAY PRESCOT 20]  AKERS, ALAN BURT  P
SWORDSHIPS OF SCORPIO [DRAY PRESCOT 4]  AKERS, ALAN BURT  C/P,P
TALONS OF SCORPIO [DRAY PRESCOT 30]  AKERS, ALAN BURT  P
TIDES OF KREGEN, THE [DRAY PRESCOT 12]  AKERS, ALAN BURT  P
TRANSIT TO SCORPIO [DRAY PRESCOT 1]  AKERS, ALAN BURT  C/P:2,P
VICTORY FOR KREGEN, A [DRAY PRESCOT 22]  AKERS, ALAN BURT  C/P,P
WARLORD OF ANTARES [DRAY PRESCOT 37]  AKERS, ALAN BURT  P
WARRIOR OF SCORPIO [DRAY PRESCOT 3]  AKERS, ALAN BURT  C/P,P
WEREWOLVES OF KREGEN [DRAY PRESCOT 33]  AKERS, ALAN BURT  P
WITCHES OF KREGEN [DRAY PRESCOT 34]  AKERS, ALAN BURT  P
ALL DARKNESS MET [DREAD EMPIRE 3/RAGNARSON 5]  COOK, GLEN  C/P:2,P
ILL FATE MARSHALLING, AN [DREAD EMPIRE 4/RAGNARSON 6]  COOK, GLEN  C/P,SFWA-P,P
OCTOBER’S BABY [DREAD EMPIRE 2/RAGNARSON 4]  COOK, GLEN  C/P:2,P
REAP THE EAST WIND [DREAD EMPIRE 5/RAGNARSON 7]  COOK, GLEN  C/P:2,SFWA-P,P
SHADOW OF ALL NIGHT FALLING, A [DREAD EMPIRE 1/RAGNARSON 3]  COOK, GLEN  C/P:2,P
CONVENTIONS OF WAR [@DREAD EMPIRE’S FALL 3]  WILLIAMS, WALTER JON  P
PRAXIS, THE [@DREAD EMPIRE’S FALL 1]  WILLIAMS, WALTER JON  P
PRAXIS, THE (PROOF) [@DREAD EMPIRE’S FALL 1]  WILLIAMS, WALTER JON  SR-P
SUNDERING, THE [@DREAD EMPIRE’S FALL 2]  WILLIAMS, WALTER JON  P
PLAGUE OF NIGHTMARES, A [@DREAM LORDS #1]  COLE, ADRIAN  C/P:2,P
LORD OF NIGHTMARES [@DREAM LORDS #2]  COLE, ADRIAN  P
BANE OF NIGHTMARES [@DREAM LORDS #3]  COLE, ADRIAN  P
BARSOOM PROJECT, THE [DREAM PARK]  NIVEN, LARRY/BARNES, STEVEN  C/P,SFWA-P,P,H
CALIFORNIA VOODOO GAME, THE [DREAM PARK]  NIVEN, LARRY/BARNES, STEVEN  C/P,P,H
DREAM PARK [DREAM PARK]  NIVEN, LARRY/BARNES, STEVEN  H:3,P
ELDER GODS, THE (PROOF) [DREAMERS 1]  EDDINGS, DAVID/EDDINGS, LEIGH  SR-LP
ELDER GODS, THE [@DREAMERS #1]  EDDINGS, DAVID/EDDINGS, LEIGH  H:2
TREASURED ONE, THE [@DREAMERS #2]  EDDINGS, DAVID/EDDINGS, LEIGH  H
CRYSTAL GORGE [@DREAMERS #3]  EDDINGS, DAVID/EDDINGS, LEIGH  H
YOUNGER GODS, THE [@DREAMERS #4]  EDDINGS, DAVID/EDDINGS, LEIGH  H
ARGUS GAMBIT, THE [DREAMERS OF THE DAY 1]  ROSS, DAVID D.  H,SFWA
EIGHTH RANK, THE [DREAMERS OF THE DAY 2]  ROSS, DAVID D.  H,SFWA
DREAMHUNTER [DREAMHUNTER 1]  KNOX, ELIZABETH  H
DREAMQUAKE [DREAMHUNTER 2]  KNOX, ELIZABETH  H
DREAMLAND CHRONICLES [DREAMLAND 1,2,3]  SIMMONS, WM. MARK  LP
ICED ON ARAN [DREAMLAND 4]  LUMLEY, BRIAN  P
IN THE NET OF DREAMS [DREAMLAND 1]  SIMMONS, WM. MARK  P
MAD MOON OF DREAMS [DREAMLAND 3]  LUMLEY, BRIAN  P
HALBERD, DREAM WARRIOR [DREAMQUEST 1]  SAINT ALCORN, LLOYD  P
ON THE SHOULDERS OF GIANTS [DREAMQUEST 2]  SAINT ALCORN, LLOYD  P
COVENANT RISING, THE [DREAMTIME 1]  NICHOLLS, STAN  H,LP
HERO OF DREAMS [DREAMWORLD 1]  LUMLEY, BRIAN  LP
SHIP OF DREAMS [DREAMWORLD 2]  LUMLEY, BRIAN  H
FIRST CHRONICLES OF DRUSS THE LEGEND, THE [DRENAI]  GEMMELL, DAVID  C/P,P
HERO IN THE SHADOWS [DRENAI]  GEMMELL, DAVID  P
KING BEYOND THE GATE, THE [DRENAI 2]  GEMMELL, DAVID  P
QUEST FOR LOST HEROES [DRENAI 3]  GEMMELL, DAVID  P
WHITE WOLF [DRENAI]  GEMMELL, DAVID  P

TITLE  AUTHOR  CODE
WINTER WARRIORS [DRENAI]  GEMMELL, DAVID  P
AGAINST THE HORDE [DRENAI 1 (ORIGINAL)]  GEMMELL, DAVID  P
LEGEND [DRENAI 1 (REVISED)]  GEMMELL, DAVID  C/P:2,P
IN THE REALM OF THE WOLF [DRENAI WAYLANDER 2]  GEMMELL, DAVID  P
WAYLANDER [DRENAI WAYLANDER 1]  GEMMELL, DAVID  C/P,P
DEAD BEAT [DRESDEN FILES 7]  BUTCHER, JIM  H
PROVEN GUILTY [DRESDEN FILES 8]  BUTCHER, JIM  H
WHITE NIGHT [DRESDEN FILES 9]  BUTCHER, JIM  H:2
STORM FRONT [@DRESDEN FILES #1]  BUTCHER, JIM  C/P,P
FOOL MOON [@DRESDEN FILES #2]  BUTCHER, JIM  C/P:2,P
GRAVE PERIL [@DRESDEN FILES #3]  BUTCHER, JIM  C/P,P
SUMMER KNIGHT [@DRESDEN FILES #4]  BUTCHER, JIM  C/P,P
DEATH MASKS [@DRESDEN FILES #5]  BUTCHER, JIM  C/P,P
BLOOD RITES [@DRESDEN FILES #6]  BUTCHER, JIM  C/P,P
QUEEN OF DENIAL [DREWCILA QWAH 1]  ROSEN, SELINA  LP
RECYCLED [DREWCILA QWAH 2]  ROSEN, SELINA  LP
DRIFTER [DRIFTER 1]  DIETZ, WILLIAM C.  C/P:2,SFWA-P:2,P
DRIFTER’S RUN [DRIFTER 2]  DIETZ, WILLIAM C.  C/P,P
DRIFTER’S WAR [DRIFTER 3]  DIETZ, WILLIAM C.  C/P:2,SFWA-P,P
DRIVE-IN, THE [@DRIVE-IN #1]  LANSDALE, JOE R.  C/P,P
DRIVE-IN 2, THE [@DRIVE-IN #2]  LANSDALE, JOE R.  C/P,P
DRUM CALLS [DRUMS OF CHAOS 2]  CLAYTON, JO  P
DRUM INTO SILENCE [DRUMS OF CHAOS 3]  CLAYTON, JO/MURPHY, KEVIN ANDREW  H,P
DRUM WARNING [DRUMS OF CHAOS 1]  CLAYTON, JO  C/P,P
AMBIENT [DRYCO 1]  WOMACK, JACK  H,P
ELVISSEY [DRYCO 4]  WOMACK, JACK  H:2,LP
HEATHERN [DRYCO 3]  WOMACK, JACK  P,SFWA
RANDOM ACTS OF SENSELESS VIOLENCE [DRYCO 5]  WOMACK, JACK  H,LP
TERRAPLANE [DRYCO 2]  WOMACK, JACK  C/P,SFWA-P,P,H
D’SHAI [D’SHAI 1]  ROSENBERG, JOEL  C/P,SFWA-P:2,P,H
HOUR OF THE OCTOPUS [D’SHAI 2]  ROSENBERG, JOEL  P
GHOSTS IN THE SNOW [DUBRIC BYERLY 1]  JONES, TAMARA SILER  C/P,P
CHANGER’S MOON [DUEL OF SORCERY 3]  CLAYTON, JO  C/P:2,P
MOONGATHER [DUEL OF SORCERY 1]  CLAYTON, JO  C/P:2,P
MOONSCATTER [DUEL OF SORCERY 2]  CLAYTON, JO  C/P:2,P
ANGADO [DUMAREST 29]  TUBB, E. C.  C/P:2,P
COMING EVENT, THE [DUMAREST 26]  TUBB, E. C.  C/P,P
DERAI [DUMAREST 2]  TUBB, E. C.  C/P,DH077,D89301
EARTH IS HEAVEN [DUMAREST 27]  TUBB, E. C.  C/P,P
ELOISE [DUMAREST 12]  TUBB, E. C.  C/P:2,P
EYE OF THE ZODIAC [DUMAREST 13]  TUBB, E. C.  C/P,P
HAVEN OF DARKNESS [DUMAREST 16]  TUBB, E. C.  C/P:2,P
IDUNA’S UNIVERSE [DUMAREST 21]  TUBB, E. C.  C/P,P
INCIDENT ON ATH [DUMAREST 18]  TUBB, E. C.  C/P:2,P
JACK OF SWORDS [DUMAREST 14]  TUBB, E. C.  C/P:2,P
JESTER AT SCAR, THE [DUMAREST 5]  TUBB, E. C.  C/P,D81610
JONDELLE [DUMAREST 10]  TUBB, E. C.  C/P,P
KALIN [DUMAREST 4]  TUBB, E. C.  C/P,D42800,P
LALLIA [DUMAREST 6]  TUBB, E. C.  D71082
MAYENNE [DUMAREST 9]  TUBB, E. C.  C/P:2,P
MELOME [DUMAREST 28]  TUBB, E. C.  C/P,P
NECTAR OF HEAVEN [DUMAREST 24]  TUBB, E. C.  C/P:2,P
PRISON OF NIGHT [DUMAREST 17]  TUBB, E. C.  C/P:2,P
QUILLIAN SECTOR, THE [DUMAREST 19]  TUBB, E. C.  C/P,P
SPECTRUM OF A FORGOTTEN SUN [DUMAREST 15]  TUBB, E. C.  C/P:2,P
SYMBOL OF TERRA [DUMAREST 30]  TUBB, E. C.  C/P,P
TECHNOS [DUMAREST 7]  TUBB, E. C.  C/P,D79975
TEMPLE OF TRUTH, THE [DUMAREST 31]  TUBB, E. C.  C/P,P
TERRA DATA, THE [DUMAREST 22]  TUBB, E. C.  C/P,P
TERRIDAE, THE [DUMAREST 25]  TUBB, E. C.  C/P,P
TOYMAN [DUMAREST 3]  TUBB, E. C.  D23140
VERUCHIA [DUMAREST 8]  TUBB, E. C.  C/P:2,P
WEB OF SAND [DUMAREST 20]  TUBB, E. C.  C/P:2,P
WINDS OF GATH, THE [DUMAREST 1]  TUBB, E. C.  C/P,DH027,D89301
WINDS OF GATH, THE (WITH "FRESH GUY") [DUMAREST 1]  TUBB, E. C.  LP
WORLD OF PROMISE [DUMAREST 23]  TUBB, E. C.  C/P,P
ZENYA [DUMAREST 11]  TUBB, E. C.  C/P:2,P
RETURN, THE [@DUMAREST #32]  TUBB, E. C.  LP
BOOK OF SORROWS, THE [DUN COW 2]  WANGERIN, WALTER JR.  H
BOOK OF THE DUN COW, THE [DUN COW 1]  WANGERIN, WALTER JR.  C/P:2,P
BARRENLANDS [DUN LADY’S JESS 4]  DURGIN, DORANNA  P
CHANGESPELL [DUN LADY’S JESS 2]  DURGIN, DORANNA  P
CHANGESPELL LEGACY [DUN LADY’S JESS 3]  DURGIN, DORANNA  P

The MITSFS Seriesdex  September 24, 2007  Page 55
TITLE  AUTHOR  CODE
DUN LADY’S JESS [DUN LADY’S JESS 1]  DURGIN, DORANNA  P
CHAPTERHOUSE: DUNE [DUNE 6]  HERBERT, FRANK  C/P,SFWA-P,P,H
CHILDREN OF DUNE [DUNE 3]  HERBERT, FRANK  C/P:2,P,H
DUNE [DUNE 1]  HERBERT, FRANK  C/P:3,SFWA-P,P,H
DUNE MESSIAH [DUNE 2]  HERBERT, FRANK  C/P,P,H
GOD EMPEROR OF DUNE [DUNE 4]  HERBERT, FRANK  C/P,P,H,SFWA
HERETICS OF DUNE [DUNE 5]  HERBERT, FRANK  C/P,LP,H
HUNTERS OF DUNE [DUNE 7]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H
ILLUSTRATED DUNE, THE [DUNE 1]  HERBERT, FRANK  LP

ROAD TO DUNE, THE [DUNE] HERBERT, FRANK/HERBERT, BRIAN/

ANDERSON, KEVIN J.

R/REF-H

DUNE: HOUSE ATREIDES [DUNE 0.1]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2,P
DUNE: HOUSE HARKONNEN [DUNE 0.2]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2,P
DUNE: HOUSE CORRINO [DUNE 0.3]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2,P
DUNE: HOUSE CORRINO (6 CASSETTES, BOXED) [DUNE 0.3]  HERBERT, BRIAN/ANDERSON, KEVIN J.  SR-MM
BATTLE OF CORRIN, THE [DUNE JIHAD 3]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H
DUNE: THE BUTLERIAN JIHAD [DUNE JIHAD 1]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2
DUNE: THE MACHINE CRUSADE [DUNE JIHAD 2]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2
DUNE: THE MACHINE CRUSADE (PROOF) [DUNE JIHAD 2]  HERBERT, BRIAN/ANDERSON, KEVIN J.  SR-LP
DUNE: WHIPPING MEK [DUNE JIHAD]  HERBERT, BRIAN/ANDERSON, KEVIN J.  B
DUNE: WHIPPING MEK (BOOKLET AND CD) [DUNE JIHAD]  HERBERT, BRIAN/ANDERSON, KEVIN J.  SR-MM:2
DARK ABYSS, THE [DUNGEON 2]  COVILLE, BRUCE  C/P:2,P
HIDDEN CITY, THE [DUNGEON 5]  DE LINT, CHARLES  C/P:2,P
LAKE OF FIRE, THE [DUNGEON 4]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
VALLEY OF THUNDER, THE [DUNGEON 3]  DE LINT, CHARLES  C/P:2,P
BLACK TOWER, THE [DUNGEON #1]  LUPOFF, RICHARD A.  C/P:2,P
FINAL BATTLE, THE [DUNGEON #6]  LUPOFF, RICHARD A.  C/P:2,P
BLOODY EYE, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
CITY OF FIRE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
DEATH RAY, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
DUNGEONS & DRAGONS THE MOVIE [@DUNGEONS & DRAGONS]  BARRETT, NEAL JR.  P
LIVING DEAD, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
OATH OF NERULL [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
PLAGUE OF ICE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
RETURN OF THE DAMNED [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
SAVAGE CAVES, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
SUNDERED ARMS, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
TREACHERY’S WAKE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
TAINTED SWORD, THE [@DUNGEONS & DRAGONS PENHALIGON #1]  HEINRICH, D. J.  P
DRAGON’S TOMB, THE [@DUNGEONS & DRAGONS PENHALIGON #2]  HEINRICH, D. J.  P
FALL OF MAGIC, THE [@DUNGEONS & DRAGONS PENHALIGON #3]  HEINRICH, D. J.  C/P,P
DOOMSDAY BOOK [DUNWORTHY & FINCH 1]  WILLIS, CONNIE  C/P,LP,H:2
TO SAY NOTHING OF THE DOG [DUNWORTHY & FINCH 2]  WILLIS, CONNIE  H:2,P
UNQUIET GRAVE, THE [@DURHAM RED]  EVANS, PETER J.  P
EUNUCH’S HEIR, THE [DURHYS FAMILY 2]  ISAAK, ELAINE  LP
SINGER’S CROWN, THE [DURHYS FAMILY 1]  ISAAK, ELAINE  LP
DUSHAU [DUSHAU 1]  LICHTENBERG, JACQUELINE  C/P,P
FARFETCH [DUSHAU 2]  LICHTENBERG, JACQUELINE  C/P,P
OUTREACH [DUSHAU 3]  LICHTENBERG, JACQUELINE  C/P,P
COMPLEAT DYING EARTH, THE [DYING EARTH 1,2,3,4]  VANCE, JACK  H:2
CUGEL’S SAGA [DYING EARTH 3]  VANCE, JACK  C/P,P,H
DYING EARTH, THE [DYING EARTH 1]  VANCE, JACK  C/P:2,P
EYES OF THE OVERWORLD, THE [DYING EARTH 2]  VANCE, JACK  SR-H,C/P:2,P
QUEST FOR SIMBILIS, A [DYING EARTH]  SHEA, MICHAEL  C/P,P
RHIALTO THE MARVELLOUS [DYING EARTH 4]  VANCE, JACK  H,P
OUTLAW SWORDS [DYLAN MATHESON 2]  LEE, J. ARDIAN  LP
SON OF THE SWORD [DYLAN MATHESON 1]  LEE, J. ARDIAN  LP
SWORD OF KING JAMES [DYLAN MATHESON 3]  LEE, J. ARDIAN  LP

The MITSFS Seriesdex  September 24, 2007  Page 57
TITLE  AUTHOR  CODE
BLOOD AND BONE [EAGLEHEART 3]  WESTCOTT, C. T.  C/P:2,P
SILVER WINGS AND LEATHER JACKETS [EAGLEHEART 1]  WESTCOTT, C. T.  P
LEATHER WINGS [EARTH 2]  VORNHOLT, JOHN  P
AUGUR’S TEACHER [EARTH FINAL CONFLICT]  SMITH, SHERWOOD  H,LP
FIRST PROTECTOR, THE [EARTH FINAL CONFLICT]  WHITE, JAMES  H,P
HERITAGE [EARTH FINAL CONFLICT]  DURGIN, DORANNA  H
LEGACY [EARTH FINAL CONFLICT]  SIXBURY, GLENN R.  LP
REQUIEM FOR BOONE [@EARTH FINAL CONFLICT]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P,LP
EARTH SONG [EARTH SONG 2]  WEBB, SHARON  C/P,P,H
EARTHCHILD [EARTH SONG 1]  WEBB, SHARON  C/P,P,H
RAM SONG [EARTH SONG 3]  WEBB, SHARON  C/P,SFWA,P,H
EARTH 2 [EARTH 2]  CRANDALL, MELISSA  C/P,P
PUZZLE [EARTH 2]  DALTON, SEAN  C/P,P
AURORA QUEST [@EARTHBLOOD]  AXLER, JAMES  P
DEEP TREK [@EARTHBLOOD]  AXLER, JAMES  P
EARTHBLOOD [@EARTHBLOOD]  AXLER, JAMES  P
EARTHBORN, THE [EARTHBORN WARS 1]  COLLINS, PAUL  H
LONGING RING, THE [@EARTHDAWN]  KUBASIK, CHRISTOPHER  P
LOST KAER [@EARTHDAWN]  FINDLEY, NIGEL D.  P
POISONED MEMORIES [@EARTHDAWN]  KUBASIK, CHRISTOPHER  P
PROPHECY [@EARTHDAWN]  GORDEN, GREG  P
TALISMAN [@EARTHDAWN]  LEWIS, SAM  PA
CLAN OF THE CAVE BEAR, THE [EARTH’S CHILDREN 1]  AUEL, JEAN M.  C/P:2,P
MAMMOTH HUNTERS, THE [EARTH’S CHILDREN 3]  AUEL, JEAN M.  C/P,P,H
PLAINS OF PASSAGE, THE [EARTH’S CHILDREN 4]  AUEL, JEAN M.  P
VALLEY OF HORSES, THE [EARTH’S CHILDREN 2]  AUEL, JEAN M.  C/P,P,H
FARTHEST SHORE, THE [EARTHSEA 3]  LE GUIN, URSULA K.  C/P,P,H,SFWA
OTHER WIND, THE [EARTHSEA]  LE GUIN, URSULA K.  LP
TALES FROM EARTHSEA [EARTHSEA]  LE GUIN, URSULA K.  H
TALES FROM EARTHSEA [EARTHSEA]/OTHER WIND, THE [EARTHSEA]  LE GUIN, URSULA K.  H
TEHANU [EARTHSEA 4]  LE GUIN, URSULA K.  H:2,P
TOMBS OF ATUAN, THE [EARTHSEA 2]  LE GUIN, URSULA K.  C/P,P,H
WIZARD OF EARTHSEA, A [EARTHSEA 1]  LE GUIN, URSULA K.  C/P,P,H
EXPLOITS OF EBENEZUM, THE [EBENEZUM 1,2,3]  GARDNER, CRAIG SHAW  H
MALADY OF MAGICKS, A [EBENEZUM 1]  GARDNER, CRAIG SHAW  C/P:2,P
MULTITUDE OF MONSTERS, A [EBENEZUM 2]  GARDNER, CRAIG SHAW  C/P:2,P
NIGHT IN THE NETHERHELLS, A [EBENEZUM 3]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P,P
TALES OF THE LAST WAR, THE [@EBERRON]  SEHESTEDT, MARK  PA
THIEVES OF BLOOD [@EBERRON BLADE OF THE FLAME #1]  WAGGONER, TIM  P
FORGE OF THE MINDSLAYERS [@EBERRON BLADE OF THE FLAME #2]  WAGGONER, TIM  P
BINDING STONE, THE [@EBERRON DRAGON BELOW #1]  BASSINGTHWAITE, DON  P
GRIEVING TREE, THE [@EBERRON DRAGON BELOW #2]  BASSINGTHWAITE, DON  P
KILLING SONG, THE [@EBERRON DRAGON BELOW #3]  BASSINGTHWAITE, DON  P
CITY OF TOWERS, THE [@EBERRON DREAMING DARK #1]  BAKER, KEITH  P
SHATTERED LAND, THE [@EBERRON DREAMING DARK #2]  BAKER, KEITH  P
GATES OF NIGHT, THE [@EBERRON DREAMING DARK #3]  BAKER, KEITH  P
VOYAGE OF THE MOURNING DAWN [@EBERRON HEIRS OF ASH #1]  WULF, RICH  P
FLIGHT OF THE DYING SUN [@EBERRON HEIRS OF ASH #2]  WULF, RICH  P
BOUND BY IRON [@EBERRON INQUISITIVES]  BOLME, EDWARD  P
LEGACY OF WOLVES [@EBERRON INQUISITIVES]  ROCKWELL, MARSHEILA  P
NIGHT OF THE LONG SHADOWS [@EBERRON INQUISITIVES]  CRILLEY, PAUL  P
LEFT HAND OF DEATH, THE [@EBERRON LANTERNLIGHT #1]  DEWOLF, PARKER  P
MARKED FOR DEATH [@EBERRON LOST MARK #1]  FORBECK, MATT  P
ROAD TO DEATH, THE [@EBERRON LOST MARK #2]  FORBECK, MATT  P
QUEEN OF DEATH, THE [@EBERRON LOST MARK #3]  FORBECK, MATT  P
CRIMSON TALISMAN, THE [@EBERRON WAR-TORN #1]  COLE, ADRIAN  P
ORB OF XORIAT, THE [@EBERRON WAR-TORN #2]  BOLME, EDWARD  P
IN THE CLAWS OF THE TIGER [@EBERRON WAR-TORN #3]  WYATT, JAMES  P
BLOOD AND HONOR [@EBERRON WAR-TORN #4]  DAVIS, GRAEME  P
ECHOES OF EARTH [ECHOES OF EARTH 1]  WILLIAMS, SEAN/DIX, SHANE  C/P,P
HEIRS OF EARTH [ECHOES OF EARTH 3]  WILLIAMS, SEAN/DIX, SHANE  P
ORPHANS OF EARTH [ECHOES OF EARTH 2]  WILLIAMS, SEAN/DIX, SHANE  P
ECOLITAN ENIGMA, THE [ECOLITAN 4]  MODESITT, L. E. JR.  H,P
ECOLITAN OPERATION, THE [ECOLITAN 2]  MODESITT, L. E. JR.  SFWA-P
ECOLITAN PRIME [ECOLITAN 1,4]  MODESITT, L. E. JR.  LP
ECOLOGIC ENVOY, THE [ECOLITAN 1]  MODESITT, L. E. JR.  C/P,P
ECOLOGIC SECESSION, THE [ECOLITAN 3]  MODESITT, L. E. JR.  C/P:2,P
EMPIRE & ECOLITAN [ECOLITAN 2,3]  MODESITT, L. E. JR.  LP
EDDIE BEAR, PRIVATE DETECTIVE [EDDIE BEAR 1,2]  RANKIN, ROBERT  H

HOLLOW CHOCOLATE BUNNIES OF THE APOCALYPSE, THE [EDDIE BEAR 1] RANKIN, ROBERT

H

RETURN TO EDEN [EDEN 3]  HARRISON, HARRY  H:2,P
WEST OF EDEN [EDEN 1]  HARRISON, HARRY  C/P,P,H

TITLE  AUTHOR  CODE
WINTER IN EDEN [EDEN 2]  HARRISON, HARRY  H:2,P
FOUR AND TWENTY BLACKBIRDS [EDEN MOORE]  PRIEST, CHERIE  LP
WINGS TO THE KINGDOM [EDEN MOORE]  PRIEST, CHERIE  LP
ASSASSIN’S EDGE, THE [EINARINN 5]  MCKENNA, JULIET E.  P
GAMBLER’S FORTUNE, THE [EINARINN 3]  MCKENNA, JULIET E.  P
SWORDSMAN’S OATH, THE [EINARINN 2]  MCKENNA, JULIET E.  H,P
THIEF’S GAMBLE, THE [EINARINN 1]  MCKENNA, JULIET E.  C/P:2,P
WARRIOR’S BLOOD, THE [EINARINN 4]  MCKENNA, JULIET E.  P
EFFENDI [EL ISKANDRYIA 2]  GRIMWOOD, JON COURTENAY  LP
FELAHEEN [EL ISKANDRYIA 3]  GRIMWOOD, JON COURTENAY  H
PASHAZADE [EL ISKANDRYIA 1]  GRIMWOOD, JON COURTENAY  H
ALIEN BLUES [ELAKI 1]  HIGHTOWER, LYNN S.  P,SFWA-P
ALIEN EYES [ELAKI 2]  HIGHTOWER, LYNN S.  C/P,P
ALIEN HEAT [ELAKI 3]  HIGHTOWER, LYNN S.  P
ALIEN RITES [ELAKI 4]  HIGHTOWER, LYNN S.  P
CAST IN COURTLIGHT [ELANTRA 2]  SAGARA, MICHELLE  LP
CAST IN SECRET [ELANTRA 3]  SAGARA, MICHELLE  LP
CAST IN SHADOW [ELANTRA 1]  SAGARA, MICHELLE  P
EARTH LOGIC [ELEMENTAL LOGIC 2]  MARKS, LAURIE J.  H
FIRE LOGIC [ELEMENTAL LOGIC 1]  MARKS, LAURIE J.  H
WATER LOGIC [ELEMENTAL LOGIC 3]  MARKS, LAURIE J.  LP
GATES OF SLEEP, THE [ELEMENTAL MASTERS 2]  LACKEY, MERCEDES  H
PHOENIX AND ASHES [ELEMENTAL MASTERS 3]  LACKEY, MERCEDES  H:2
SERPENT’S SHADOW, THE [ELEMENTAL MASTERS 1]  LACKEY, MERCEDES  H
WIZARD OF LONDON, THE [ELEMENTAL MASTERS 4]  LACKEY, MERCEDES  H:2
DIAMOND THRONE, THE [ELENIUM 1]  EDDINGS, DAVID  H:2,P
RUBY KNIGHT, THE [ELENIUM 2]  EDDINGS, DAVID  H:2,P
SAPPHIRE ROSE, THE [ELENIUM 3]  EDDINGS, DAVID  C/P,P,H
TWISTED DRAGON [ELFWOOD]  STEIN, KEVIN  P
FIRES OF THE FAITHFUL [ELIANA 1]  KRITZER, NAOMI  P
TURNING THE STORM [ELIANA 2]  KRITZER, NAOMI  P
EXILE’S RETURN [@ELITA 1]  JACOBY, KATE  P
VOICE OF THE DEMON [@ELITA #2]  JACOBY, KATE  P
BLACK EAGLE RISING [@ELITA #3]  JACOBY, KATE  P
REBEL’S CAGE [@ELITA #4]  JACOBY, KATE  LP
BEYOND LICH GATE [ELIXIR OF LIFE 1]  FRANKOS, STEVEN  P
CATHEDRAL OF THORNS [ELIXIR OF LIFE 2]  FRANKOS, STEVEN  P
LEGEND REBORN, A [ELIXIR OF LIFE 3]  FRANKOS, STEVEN  P
NIGHTMARES ON ELM STREET, THE [ELM STREET 1,2,3]  COOPER, JEFFREY  P
BANE OF THE BLACK SWORD, THE [ELRIC 5]  MOORCOCK, MICHAEL  C/P:2,P
DREAMING CITY, THE [ELRIC]  MOORCOCK, MICHAEL  C/P:2,P
ELRIC AT THE END OF TIME [ELRIC]  MOORCOCK, MICHAEL  C/P:2,P
ELRIC OF MELNIBONE [ELRIC 1]  MOORCOCK, MICHAEL  P
ELRIC OF MELNIBONE (OMNIBUS EDITION) [ELRIC]  MOORCOCK, MICHAEL  P
ELRIC SAGA PART I, THE [ELRIC 1,2,3]  MOORCOCK, MICHAEL  H:2
ELRIC SAGA PART II, THE [ELRIC 4,5,6]  MOORCOCK, MICHAEL  H:2
ELRIC SAGA PART III, THE [ELRIC]  MOORCOCK, MICHAEL  H
FORTRESS OF THE PEARL, THE [ELRIC]  MOORCOCK, MICHAEL  H:2,SFWA,SFWA-P:2,P
REVENGE OF THE ROSE, THE [ELRIC]  MOORCOCK, MICHAEL  H:2,SFWA:2,P
SAILOR ON THE SEAS OF FATE, THE [ELRIC 2]  MOORCOCK, MICHAEL  C/P:2,P
SKRAYLING TREE, THE [ELRIC/VON BEK]  MOORCOCK, MICHAEL  P
SLEEPING SORCERESS, THE [ELRIC]  MOORCOCK, MICHAEL  C/P:2,P
STEALER OF SOULS, THE [ELRIC]  MOORCOCK, MICHAEL  SR-H,C/P:2
STORMBRINGER [ELRIC]  MOORCOCK, MICHAEL  SR-H,C/P:2,P
STORMBRINGER (REVISED) [ELRIC 6]  MOORCOCK, MICHAEL  C/P,P
TALES OF THE WHITE WOLF [ELRIC]  KRAMER, EDWARD E.  PA
VANISHING TOWER, THE [ELRIC 4]  MOORCOCK, MICHAEL  C/P,P
WEIRD OF THE WHITE WOLF, THE [ELRIC 3]  MOORCOCK, MICHAEL  C/P:2,P
CITADEL OF SHADOWS [ELUNDIUM 4]  JEFFERIES, MIKE  C/P,P
KNIGHTS OF CAWDOR, THE [ELUNDIUM 5]  JEFFERIES, MIKE  P
PALACE OF KINGS [ELUNDIUM 2]  JEFFERIES, MIKE  P
ROAD TO UNDERFALL, THE [ELUNDIUM 1]  JEFFERIES, MIKE  C/P:2,P
SHADOWLIGHT [ELUNDIUM 3]  JEFFERIES, MIKE  P
HIDDEN TEMPLE, THE [ELVEN KINGDOMS 3]  COOKE, CATHERINE  C/P,SFWA-P,P
MASK OF THE WIZARD [ELVEN KINGDOMS 1]  COOKE, CATHERINE  C/P,SFWA-P,P
VEIL OF SHADOW [ELVEN KINGDOMS 2]  COOKE, CATHERINE  P,SFWA-P
FOUR FORGES, THE [ELVEN WAYS 1]  RHODES, JENNA  H
WAYS OF MAGIC, THE [@ELVEN WAYS #1]  CIENCIN, SCOTT  P
ANCIENT GAMES [@ELVEN WAYS #2]  CIENCIN, SCOTT  P
NIGHT OF GLORY [@ELVEN WAYS #3]  CIENCIN, SCOTT  P
BRAIN PLAGUE [ELYSIUM]  SLONCZEWSKI, JOAN  H
CHILDREN STAR, THE [ELYSIUM]  SLONCZEWSKI, JOAN  H:2,P
DAUGHTER OF ELYSIUM [ELYSIUM]  SLONCZEWSKI, JOAN  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 59
TITLE  AUTHOR  CODE
DOOR INTO OCEAN, A [ELYSIUM]  SLONCZEWSKI, JOAN  H:2,P,SFWA
PHARAOH CONTRACT, THE [@EMANCIPATOR #1]  ALDRIDGE, RAY  C/P:2,P
EMPEROR OF EVERYTHING, THE [@EMANCIPATOR #2]  ALDRIDGE, RAY  C/P,P
ORPHEUS MACHINE, THE [@EMANCIPATOR #3]  ALDRIDGE, RAY  C/P,P
DIES THE FIRE [EMBERVERSE 1]  STIRLING, S. M.  H
MEETING AT CORVALLIS, A [EMBERVERSE 3]  STIRLING, S. M.  H
PROTECTOR’S WAR, THE [EMBERVERSE 2]  STIRLING, S. M.  H
BEHIND TIME [EMMA MERRIGAN 2]  ABBEY, LYNN  C/P,P
DOWN TIME [EMMA MERRIGAN 4]  ABBEY, LYNN  P
OUT OF TIME [EMMA MERRIGAN 1]  ABBEY, LYNN  C/P,P
TAKING TIME [EMMA MERRIGAN 3]  ABBEY, LYNN  P
ARCHITECTS OF EMORTALITY [EMORTALITY]  STABLEFORD, BRIAN M.  H
CASSANDRA COMPLEX, THE [EMORTALITY]  STABLEFORD, BRIAN M.  H
DARK ARARAT [EMORTALITY]  STABLEFORD, BRIAN M.  H
FOUNTAINS OF YOUTH, THE [EMORTALITY]  STABLEFORD, BRIAN M.  H,LP
INHERIT THE EARTH [EMORTALITY]  STABLEFORD, BRIAN M.  H
OMEGA EXPEDITION, THE [EMORTALITY]  STABLEFORD, BRIAN M.  H
RUNES & RIDDLES [EMPIRE OF ELWOLF]  CARRADINE, F. J.  LP
ARDNEH’S WORLD [EMPIRE OF THE EAST 3]  SABERHAGEN, FRED  P,SFWA-P
BLACK MOUNTAINS, THE [EMPIRE OF THE EAST 2]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
BROKEN LANDS, THE [EMPIRE OF THE EAST 1]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
CHANGELING EARTH [EMPIRE OF THE EAST 3]  SABERHAGEN, FRED  C/P:2,P
EMPIRE OF THE EAST [EMPIRE OF THE EAST 1,2,3]  SABERHAGEN, FRED  H:2,SFWA-P,P
BEASTS OF HADES, THE [EMPIRE PRINCESS 4]  DIAMOND, GRAHAM  P
DUNGEONS OF KUBA [EMPIRE PRINCESS 2]  DIAMOND, GRAHAM  P
FALCON OF EDEN, THE [EMPIRE PRINCESS 3]  DIAMOND, GRAHAM  P
HAVEN, THE [EMPIRE PRINCESS 0]  DIAMOND, GRAHAM  P
LADY OF THE HAVEN [EMPIRE PRINCESS 1]  DIAMOND, GRAHAM  P
SEARCH FOR FIERRA, THE [@EMPYRION #1]  LAWHEAD, STEPHEN R.  LP
SIEGE OF DOME, THE [@EMPYRION #2]  LAWHEAD, STEPHEN R.  LP
CALLING ON DRAGONS [ENCHANTED FOREST 3]  WREDE, PATRICIA C.  P
DEALING WITH DRAGONS [ENCHANTED FOREST 1]  WREDE, PATRICIA C.  P
SEARCHING FOR DRAGONS [ENCHANTED FOREST 2]  WREDE, PATRICIA C.  P
TALKING TO DRAGONS [ENCHANTED FOREST 4]  WREDE, PATRICIA C.  P
CHILDREN OF THE MIND [ENDER 4]  CARD, ORSON SCOTT  H:3,P
ENDER’S GAME (ORIGINAL) [ENDER 1]  CARD, ORSON SCOTT  P
ENDER’S GAME (REVISED 1991 EDITION) [ENDER 1]  CARD, ORSON SCOTT  C/P:3,SFWA-P,P

ENDER’S WAR (CONTAINS ORIGINAL NOVEL VERSION OF ENDER’S GAME)

[ENDER 1,2] CARD, ORSON SCOTT

H:2

FIRST MEETINGS IN THE ENDERVERSE [ENDER]  CARD, ORSON SCOTT  H
SPEAKER FOR THE DEAD [ENDER 2]  CARD, ORSON SCOTT  C/P:2,SFWA,SFWA-P,P
XENOCIDE [ENDER 3]  CARD, ORSON SCOTT  H:3,P
XENOCIDE (PROOF) [ENDER 3]  CARD, ORSON SCOTT  SR-LP
ENDER’S SHADOW [ENDER 1.1]  CARD, ORSON SCOTT  H:2,P
SHADOW OF THE HEGEMON [ENDER 1.2]  CARD, ORSON SCOTT  C/P,P,H
SHADOW PUPPETS [ENDER 1.3]  CARD, ORSON SCOTT  H:2,P
SHADOW OF THE GIANT [ENDER 1.4]  CARD, ORSON SCOTT  H:2,P
CITIES IN SPACE [ENDLESS FRONTIER 3]  POURNELLE, JERRY/CARR, JOHN F.  PA
KNIGHT OF ILLUSION [ENDLESS QUEST 33]  KIRCHOFF, MARY  C/P,P
VILLAINS OF VOLTURNUS [ENDLESS QUEST 8]  BLASHFIELD, JEAN  P
FOX RUN, THE [@ENDWORLD 1]  ROBBINS, DAVID  P
TWIN CITIES RUN [@ENDWORLD #3]  ROBBINS, DAVID  P
PASSION FOR DEAD LEAVES [ENEMIES OF SOCIETY 3]  DAVID, JOHN  LP
SEND IN THE SWORDS [ENEMIES OF SOCIETY 4]  DAVID, JOHN  LP
BECOMING DEATH [@ENEMIES OF SOCIETY #1]  DAVID, JOHN  LP
DESTROYING WORLDS [@ENEMIES OF SOCIETY #2]  DAVID, JOHN  LP
ENEMY GLORY [ENEMY GLORY 1]  MICHALSON, KAREN  H,P
HECATE’S GLORY [ENEMY GLORY 2]  MICHALSON, KAREN  H
ENFORCER, THE [@ENFORCER #1]  SUGAR, ANDREW  P
CALLING DOCTOR KILL! [@ENFORCER #2]  SUGAR, ANDREW  P
KILL CITY [@ENFORCER #3]  SUGAR, ANDREW  P
KILL DEADLINE [@ENFORCER #4]  SUGAR, ANDREW  P
BIO BLITZ [@ENFORCER #5]  SUGAR, ANDREW  P
STEEL TRAP [@ENFORCER #6]  SUGAR, ANDREW  P
COSMONAUT KEEP [ENGINES OF LIGHT 1]  MACLEOD, KEN  H
DARK LIGHT [ENGINES OF LIGHT 2]  MACLEOD, KEN  H
ENGINE CITY [ENGINES OF LIGHT 3]  MACLEOD, KEN  H
BRIGHT OF THE SKY [ENTIRE AND ROSE 1]  KENYON, KAY  H
A LA LUNA EN UN SEGUNDO [ENVIADO SECRETO 81]  CAUDETT, FRANK  FP
EON [EON 1]  BEAR, GREG  C/P,SFWA-P:2,H:2,P
ETERNITY [EON 2]  BEAR, GREG  SFWA-P,H:2,P
LEGACY [EON 0]  BEAR, GREG  H:2,P
BELLADONNA [EPHEMERA]  BISHOP, ANNE  H

The MITSFS Seriesdex  September 24, 2007  Page 60
TITLE  AUTHOR  CODE
SEBASTIAN [EPHEMERA]  BISHOP, ANNE  H
CHRONICLES OF ER-DA: BOOK ONE [ER-DA 1]  REDMOND, JEFFREY  LP
CHRONICLES OF ER-DA: BOOK TWO [ER-DA 2]  REDMOND, JEFFREY  LP
ER-DAN STORIES, THE [ER-DA]  REDMOND, JEFFREY  LP
DRAGON IN THE SWORD, THE [EREKOSE 3/ETERNAL CHAMPION 3]  MOORCOCK, MICHAEL  C/P,SFWA-P,P,H
ETERNAL CHAMPION, THE [EREKOSE 1/ETERNAL CHAMPION 1]  MOORCOCK, MICHAEL  C/P,SFWA,SFWA-P,P
SILVER WARRIORS, THE [EREKOSE 2]  MOORCOCK, MICHAEL  C/P:2,P
BEDLAM BOYZ [ERIC BANYON]  GUON, ELLEN  C/P:2,SFWA-P,P
BEDLAM’S BARD [ERIC BANYON 1,2]  LACKEY, MERCEDES/GUON, ELLEN  H,P
BEYOND WORLD’S END [ERIC BANYON 3]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H:2
KNIGHT OF GHOSTS AND SHADOWS [ERIC BANYON 1]  LACKEY, MERCEDES/GUON, ELLEN  C/P:2,SFWA-P,P
MAD MAUDLIN [ERIC BANYON 5]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H:2
MUSIC TO MY SORROW [ERIC BANYON 6]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H
SPIRITS WHITE AS LIGHTNING [ERIC BANYON 4]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H
SUMMONED TO TOURNEY [ERIC BANYON 2]  LACKEY, MERCEDES/GUON, ELLEN  C/P:2,SFWA-P,P
ERIC BRIGHTEYES [ERIC BRIGHTEYES 1]  HAGGARD, H. RIDER  C/P,LP
WITCH’S WELCOME, A [ERIC BRIGHTEYES 2]  SKALDASPILLIR, SIGFRIOUR  P
DARYA OF THE BRONZE AGE [ERIC CARSTAIRS 4]  CARTER, LIN  P
ERIC OF ZANTHODON [ERIC CARSTAIRS 5]  CARTER, LIN  P
HUROK OF THE STONE AGE [ERIC CARSTAIRS 3]  CARTER, LIN  P
JOURNEY TO THE UNDERGROUND WORLD [ERIC CARSTAIRS 1]  CARTER, LIN  P
ZANTHODON [ERIC CARSTAIRS 2]  CARTER, LIN  P
AT THE NARROW PASSAGE [ERIC MATHERS 1]  MEREDITH, RICHARD C.  C/P,P,H
NO BROTHER, NO FRIEND [ERIC MATHERS 2]  MEREDITH, RICHARD C.  C/P,P,H
VESTIGES OF TIME [ERIC MATHERS 3]  MEREDITH, RICHARD C.  H,SFWA,P
CONTACT AND COMMUNE [ERIS 1]  SMITH, L. NEIL  P
CONVERSE AND CONFLICT [ERIS 2]  SMITH, L. NEIL  P
FORGE OF THE ELDERS [ERIS 1,2]  SMITH, L. NEIL  P
BLOODWINTER [ERON 1]  DEITZ, TOM  LP
SPRINGWAR [ERON 2]  DEITZ, TOM  LP
SUMMERBLOOD [ERON 3]  DEITZ, TOM  P
WARAUTUMN [ERON 4]  DEITZ, TOM  P
ERTHRING CYCLE, THE [ERTHRING 1,2,3]  DREW, WAYLAND  H:2
GAIAN EXPEDIENT [ERTHRING 2]  DREW, WAYLAND  C/P:2,P
MASTER OF NORRIYA, THE [ERTHRING 3]  DREW, WAYLAND  P
MEMOIRS OF ALCHERINGIA, THE [ERTHRING 1]  DREW, WAYLAND  C/P:2,P
FAR SHORE OF TIME, THE [ESCHATON 3]  POHL, FREDERIK  C/P,P,H
OTHER END OF TIME, THE [ESCHATON 1]  POHL, FREDERIK  C/P:2,P
SIEGE OF ETERNITY, THE [ESCHATON 2]  POHL, FREDERIK  C/P,P,H
DISAPPEARING NIGHTLY [ESTHER DIAMOND 1]  RESNICK, LAURA  LP
LEGENDS [ETERNAL ARCHIVES #1]  WEIS, MARGARET  PA
EARL AUBEC AND OTHER STORIES [ETERNAL CHAMPION]  MOORCOCK, MICHAEL  P
ETERNAL CHAMPION, THE (OMNIBUS) [ETERNAL CHAMPION]  MOORCOCK, MICHAEL  LP
PAWN OF CHAOS [ETERNAL CHAMPION]  MOORCOCK, MICHAEL  LPA
PHOENIX IN OBSIDIAN [ETERNAL CHAMPION 2]  MOORCOCK, MICHAEL  SR-P
WHITE GUARDIAN, THE [ETERNAL GUARDIANS 3]  CROSS, RONALD ANTHONY  H,LP
FOURTH GUARDIAN, THE [@ETERNAL GUARDIANS #1]  CROSS, RONALD ANTHONY  LP
LOST GUARDIAN, THE [@ETERNAL GUARDIANS #2]  CROSS, RONALD ANTHONY  H,P
WAR IN HEAVEN, THE [ETERNAL WARRIORS 1]  BEALE, THEODORE  P
STARK’S COMMAND [ETHAN STARK 2]  HEMRY, JOHN G.  P
STARK’S CRUSADE [ETHAN STARK 3]  HEMRY, JOHN G.  P
STARK’S WAR [ETHAN STARK 1]  HEMRY, JOHN G.  C/P,P
BELOVED SON [ETHICAL SOCIETY 1]  TURNER, GEORGE  C/P,P
VANEGLORY [ETHICAL SOCIETY 2]  TURNER, GEORGE  P
YESTERDAY’S MEN [ETHICAL SOCIETY 3]  TURNER, GEORGE  P
BLOOD OF A DRAGON, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
ITHANALIN’S RESTORATION [ETHSHAR]  WATT-EVANS, LAWRENCE  H
MISENCHANTED SWORD, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
NIGHT OF MADNESS [ETHSHAR]  WATT-EVANS, LAWRENCE  H
SPELL OF THE BLACK DAGGER, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
SPRIGGAN MIRROR, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  LP
TAKING FLIGHT [ETHSHAR]  WATT-EVANS, LAWRENCE  P
UNWILLING WARLORD, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  P
WITH A SINGLE SPELL [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
BORN IN DEATH [EVE DALLAS]  ROBB, J. D.  H
DIVIDED IN DEATH [EVE DALLAS]  ROBB, J. D.  H
IMMORTAL IN DEATH [EVE DALLAS]  ROBB, J. D.  P
JUDGMENT IN DEATH [EVE DALLAS]  ROBB, J. D.  P
ORIGIN IN DEATH [EVE DALLAS]  ROBB, J. D.  H

OUT OF THIS WORLD [EVE DALLAS] ROBB, J. D./HAMILTON, LAURELL K./ KRINARD, SUSAN/SHAYNE, MAGGIE

C/P,P

PURITY IN DEATH [EVE DALLAS]  ROBB, J. D.  P
RAPTURE IN DEATH [EVE DALLAS]  ROBB, J. D.  P

TITLE  AUTHOR  CODE
SURVIVOR IN DEATH [EVE DALLAS]  ROBB, J. D.  H
VISIONS IN DEATH [EVE DALLAS]  ROBB, J. D.  H
FALSE HOUSE, THE [EVENMERE 2]  STODDARD, JAMES  P
HIGH HOUSE, THE [EVENMERE 1]  STODDARD, JAMES  H,P
HIGH HOUSE, THE (PROOF) [EVENMERE 1]  STODDARD, JAMES  SR-P
EVER: THE WAR IN THE WASTE [EVER 1]  SAVAGE, FELICITY  LP
EVERGENCE [EVERGENCE 1,2,3]  WILLIAMS, SEAN/DIX, SHANE  H
EVERGENCE: A DARK IMBALANCE [EVERGENCE 3]  WILLIAMS, SEAN/DIX, SHANE  P
EVERGENCE: THE DYING LIGHT [EVERGENCE 2]  WILLIAMS, SEAN/DIX, SHANE  P
EVERGENCE: THE PRODIGAL SUN [EVERGENCE 1]  WILLIAMS, SEAN/DIX, SHANE  P
COMPANY OF GLASS, THE [EVERIEN 1]  LEITH, VALERY  C/P,LP
RIDDLED NIGHT, THE [EVERIEN 2]  LEITH, VALERY  C/P,LP
WAY OF THE ROSE, THE [EVERIEN 3]  LEITH, VALERY  LP
ENTER THE ENCHANTED [@EVERWORLD #3]  APPLEGATE, K. A.  P
DAY DARK AS NIGHT, A [@EXALTED #1]  BOWEN, CARL  P
RELIC OF THE DAWN [@EXALTED #2]  WILSON, DAVID NIALL  P
IN NORTHERN TWILIGHT [@EXALTED #3]  HARTLEY, JESS  P
PILLAR OF THE SUN [@EXALTED #4]  BOWEN, CARL  P
CHOSEN THE SUN [@EXALTED SECOND AGE #1]  DANSKY, RICHARD E.  P
BELOVED OF THE DEAD [@EXALTED SECOND AGE #2]  DANSKY, RICHARD E.  P
CHILDREN OF THE DRAGON [@EXALTED SECOND AGE #3]  DANSKY, RICHARD E.  P
DAREDEVILS, LTD. [EXCHAMELEON 1]  GOULART, RON  C/P:2,P
EVERYBODY COMES TO COSMO’S [EXCHAMELEON 3]  GOULART, RON  C/P,P
STARPIRATE’S BRAIN [EXCHAMELEON 2]  GOULART, RON  C/P:2,P
EXILE’S CHALLENGE [EXILES 2]  WELLS, ANGUS  LP
EXILE’S CHILDREN [EXILES 1]  WELLS, ANGUS  LP
RUINS OF AMBRAI, THE [@EXILES #1]  RAWN, MELANIE  C/P,P,H
MAGEBORN TRAITOR, THE [@EXILES #2]  RAWN, MELANIE  C/P,P,H
STORMS OF DESTINY [EXILES OF BOQ’URAIN 1]  CRISPIN, A. C.  P
EXORCIST, THE [EXORCIST 1]  BLATTY, WILLIAM PETER  C/P:2,P
LEGION [EXORCIST 2]  BLATTY, WILLIAM PETER  H:2
PHOENIX IN FLIGHT, THE [EXORDIUM 1]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
PRISON UNSOUGHT, A [EXORDIUM 3]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
RIFTER’S COVENANT, THE [EXORDIUM 4]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
RULER OF NAUGHT [EXORDIUM 2]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
THRONES OF KRONOS, THE [EXORDIUM 5]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
DEATHWORMS OF KRATOS, THE [EXPENDABLES 1]  AVERY, RICHARD  C/P:2,P
RINGS OF TANTALUS, THE [EXPENDABLES 2]  AVERY, RICHARD  C/P:2,P
VENOM OF ARGUS, THE [EXPENDABLES 4]  AVERY, RICHARD  C/P:2,P
WAR GAMES OF ZELOS, THE [EXPENDABLES 3]  AVERY, RICHARD  C/P:2,P
SIGNALS [@EXPLORATION #1]  RANDLE, KEVIN  C/P,P
STARSHIP [@EXPLORATION #2]  RANDLE, KEVIN  C/P,P
F.T.L. [@EXPLORATION #3]  RANDLE, KEVIN  P
GATE, THE [@EXPLORATION #4]  RANDLE, KEVIN  P
DEVIL’S ARMOR [EYES OF GOD 2]  MARCO, JOHN  P
EYES OF GOD [EYES OF GOD 1]  MARCO, JOHN  H,P
SWORD OF ANGELS, THE [EYES OF GOD 3]  MARCO, JOHN  H

TITLE  AUTHOR  CODE
BOOK OF THE WAR, THE [@FACTION PARADOX 0]  MILES, LAWRENCE  LPA
ERASING SHERLOCK [FACTION PARADOX 5]  HALE, KELLY  LP
OF THE CITY OF THE SAVED... [@FACTION PARADOX 2]  PURSER-HALLARD, PHILIP  LP
THIS TOWN WILL NEVER LET US GO [@FACTION PARADOX 1]  MILES, LAWRENCE  LP
WARLORDS OF UTOPIA [@FACTION PARADOX 3]  PARKIN, LANCE  LP
WARRING STATES [@FACTION PARADOX 4]  HALLIDAY, MAGS L.  LP
FADED SUN, THE [FADED SUN 1,2,3]  CHERRYH, C. J.  C/P:2,P
FADED SUN: KESRITH, THE [FADED SUN 1]  CHERRYH, C. J.  H:2,P
FADED SUN: KUTATH, THE [FADED SUN 3]  CHERRYH, C. J.  H:2,P
FADED SUN: SHON’JIR, THE [FADED SUN 2]  CHERRYH, C. J.  H:2,P
KEY FOR THE NONESUCH, A [FADING WORLDS 1]  GRAVEL, GEARY  C/P:2,P
RETURN OF THE BREAKNECK BOYS [FADING WORLDS 2]  GRAVEL, GEARY  P
ILL MET IN LANKHMAR [FAFHRD 1,2]  LEIBER, FRITZ  H,SFWA-TD19,TD19
ILL MET IN LANKHMAR (AND OTHER STORIES) [FAFHRD]  LEIBER, FRITZ  P
ILL MET IN LANKHMAR (SHORT STORY) [FAFHRD]  LEIBER, FRITZ  H
KNIGHT AND KNAVE OF SWORDS, THE [FAFHRD]  LEIBER, FRITZ  C/P,SFWA-P:2,P,H
LANKHMAR [FAFHRD]  LEIBER, FRITZ  P
SWORDS AGAINST DEATH [FAFHRD 2]  LEIBER, FRITZ  C/P:2,P
SWORDS AGAINST THE SHADOWLAND [FAFHRD]  BAILEY, ROBIN WAYNE  H,LP
SWORDS AGAINST WIZARDRY [FAFHRD 4]  LEIBER, FRITZ  C/P:2,P
SWORDS AND DEVILTRY [FAFHRD 1]  LEIBER, FRITZ  C/P:2,P
SWORDS AND ICE MAGIC [FAFHRD 6]  LEIBER, FRITZ  C/P:2,P
SWORDS IN THE MIST [FAFHRD 3]  LEIBER, FRITZ  C/P:2,P
SWORDS’ MASTERS [FAFHRD 4,5,6]  LEIBER, FRITZ  H:2
SWORDS OF LANKHMAR, THE [FAFHRD 5]  LEIBER, FRITZ  SR-P,C/P:2
THREE OF SWORDS, THE [FAFHRD 1,2,3]  LEIBER, FRITZ  H
BONES OF THE PAST [FAIA RISSEDOTTE 2]  LISLE, HOLLY  C/P:2,SFWA-P,P
FIRE IN THE MIST [FAIA RISSEDOTTE 1]  LISLE, HOLLY  C/P:2,SFWA-P,P
MIND OF THE MAGIC [FAIA RISSEDOTTE 3]  LISLE, HOLLY  C/P:2,P
WATER SPRITES, THE [FAIRY REALM 8]  RODDA, EMILY  H
CAVES OF MARS (PROOF) [FALCON GOLD 1]  WOLFF, ROBERT SHERMAN  SR-LP
FALKENBERG’S LEGION [FALKENBERG’S LEGION]  POURNELLE, JERRY  C/P,P
GO TELL THE SPARTANS [FALKENBERG’S LEGION]  POURNELLE, JERRY/STIRLING, S. M.  C/P:2,SFWA-P,P
PRINCE OF MERCENARIES [FALKENBERG’S LEGION]  POURNELLE, JERRY  C/P:2,SFWA-P,P
PRINCE OF SPARTA [FALKENBERG’S LEGION]  POURNELLE, JERRY/STIRLING, S. M.  C/P:2,SFWA-P,P
GATE OF GODS, THE [FALL OF ILE-RIEN 3]  WELLS, MARTHA  H
SHIPS OF AIR, THE [FALL OF ILE-RIEN 2]  WELLS, MARTHA  H
WIZARD HUNTERS, THE [FALL OF ILE-RIEN 1]  WELLS, MARTHA  H:2
BEYOND INFINITY [FALL OF NIGHT]  BENFORD, GREGORY  H
MASTER OF EVIL, THE [FALL OF THE FIRST WORLD 1]  SMITH, DAVID C.  P
PASSING OF THE GODS, THE [FALL OF THE FIRST WORLD 3]  SMITH, DAVID C.  P
SORROWING VENGEANCE [FALL OF THE FIRST WORLD 2]  SMITH, DAVID C.  P
CITY OF A THOUSAND SUNS [FALL OF THE TOWERS 3]  DELANY, SAMUEL R.  C/P:2,P
FALL OF THE TOWERS, THE [FALL OF THE TOWERS 1,2,3]  DELANY, SAMUEL R.  C/P:2,P
OUT OF THE DEAD CITY [FALL OF THE TOWERS 1]  DELANY, SAMUEL R.  P
TOWERS OF TORON, THE [FALL OF THE TOWERS 2]  DELANY, SAMUEL R.  C/P,DF261,P
CASSINI DIVISION, THE [FALL REVOLUTION]  MACLEOD, KEN  H,P
FALL REVOLUTION, THE [FALL REVOLUTION]  MACLEOD, KEN  H
SKY ROAD, THE [FALL REVOLUTION]  MACLEOD, KEN  H
STAR FRACTION, THE [FALL REVOLUTION]  MACLEOD, KEN  H,LP
STONE CANAL, THE [FALL REVOLUTION]  MACLEOD, KEN  C/P,LP
FROM THE HEART OF THE STORM [FALLEN CLOUD 4]  GIAMBASTIANI, KURT R. A.  P
SHADOW OF THE STORM [FALLEN CLOUD 3]  GIAMBASTIANI, KURT R. A.  P
SPIRIT OF THUNDER, THE [FALLEN CLOUD 2]  GIAMBASTIANI, KURT R. A.  H,P
YEAR THE CLOUD FELL, THE [FALLEN CLOUD 1]  GIAMBASTIANI, KURT R. A.  P
PROBABILITY MOON [FALLERS 1]  KRESS, NANCY  H:2,P
PROBABILITY SPACE [FALLERS 3]  KRESS, NANCY  H
PROBABILITY SUN [FALLERS 2]  KRESS, NANCY  H
SPECIAL EFFECTS [@FAMILY #1]  MCCARTHY, KEVIN/SILVA, DAVID  P
INTO THE DARKNESS [@FAMILY #2]  MCCARTHY, KEVIN/SILVA, DAVID  P
IMPERIAL STARS [FAMILY D’ALEMBERT #1]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
STRANGLER’S MOON [FAMILY D’ALEMBERT #2]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
CLOCKWORK TRAITOR, THE [FAMILY D’ALEMBERT #3]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
GETAWAY WORLD [FAMILY D’ALEMBERT #4]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
APPOINTMENT AT BLOODSTAR [FAMILY D’ALEMBERT #5]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
PURITY PLOT, THE [FAMILY D’ALEMBERT #6]  SMITH, E. E./EKLUND, GORDON  C/P,@P
PLANET OF TREACHERY [FAMILY D’ALEMBERT #7]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
ECLIPSING BINARIES [FAMILY D’ALEMBERT #8]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
OMICRON INVASION, THE [FAMILY D’ALEMBERT #9]  SMITH, E. E./GOLDIN, STEPHEN  @P
REVOLT OF THE GALAXY [FAMILY D’ALEMBERT #10]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
DOOMSDAY [FANTASTIC FOUR]  WOLFMAN, MARV  C/P,P
TO FREE ATLANTIS [FANTASTIC FOUR]  COLLINS, NANCY A.  P
CARD MASTER [FANTASY ADVENTURES]  EMERY, CLAYTON  P

The MITSFS Seriesdex  September 24, 2007  Page 64
TITLE  AUTHOR  CODE
BEGGAR’S BANQUET [FANUILH 3]  HOOD, DANIEL  P
FAMILIAR DRAGON, THE [FANUILH 1,2,3]  HOOD, DANIEL  H
FANUILH [FANUILH 1]  HOOD, DANIEL  C/P,P
KING’S CURE [FANUILH 5]  HOOD, DANIEL  P
SCALES OF JUSTICE [FANUILH 4]  HOOD, DANIEL  P
WIZARD’S HEIR [FANUILH 2]  HOOD, DANIEL  P
FAR SIDE OF FOREVER, THE [@FAR SIDE OF FOREVER #1]  GREEN, SHARON  C/P,P
HELLHOUND MAGIC [@FAR SIDE OF FOREVER #2]  GREEN, SHARON  P
FLIGHT OF HONOR [FAR STARS AND FUTURE TIMES]  MCENROE, RICHARD S.  P
SHATTERED STARS, THE [FAR STARS AND FUTURE TIMES]  MCENROE, RICHARD S.  C/P:2,P
SKINNER [FAR STARS AND FUTURE TIMES]  MCENROE, RICHARD S.  C/P:2,P
CHILDREN OF ASHGAROTH, THE [FARADAWN 3]  FORD, RICHARD  P
QUEST FOR THE FARADAWN [FARADAWN 1]  FORD, RICHARD  H,P
FALL OF A KINGDOM [@FARSALA #1]  BELL, HILARI  P
RISE OF A HERO [@FARSALA #2]  BELL, HILARI  H
DARK SIDE OF THE SUN [@FARSCAPE]  DYMOND, ANDREW  C/P,P
FARSCAPE: THE ILLUSTRATED COMPANION [FARSCAPE]  SIMPSON, PAUL/HUGHES, DAVID  R/REF-P
HOUSE OF CARDS [@FARSCAPE]  DECANDIDO, KEITH R. A.  C/P,P
SHIP OF GHOSTS [@FARSCAPE]  BISCHOFF, DAVID F.  C/P,P
ASSASSIN’S APPRENTICE [FARSEER 1]  HOBB, ROBIN  C/P:3,P
ASSASSIN’S QUEST [FARSEER 3]  HOBB, ROBIN  H,P
ROYAL ASSASSIN [FARSEER 2]  HOBB, ROBIN  C/P:2,P
TREASURE OF WONDERWHAT, THE [FARSTAR & SON 2]  STARR, BILL  C/P:2,P
WAY TO DAWNWORLD, THE [FARSTAR & SON 1]  STARR, BILL  C/P:2,P

BRIDE OF THE FAT WHITE VAMPIRE [FAT WHITE VAMPIRE (JULES DUCHON) 2] FOX, ANDREW

LP

FAT WHITE VAMPIRE BLUES [FAT WHITE VAMPIRE (JULES DUCHON) 1]  FOX, ANDREW  LP
EARTHQUAKE WEATHER [FAULT LINES 2]  POWERS, TIM  P
EXPIRATION DATE [FAULT LINES 1]  POWERS, TIM  H,P
FAULT LINES [FAULT LINES 1,2]  POWERS, TIM  H
FEAR OF GOD, THE [FEAR 2]  CHEPAITIS, B. A.  C/P,P
FEAR PRINCIPLE, THE [FEAR 1]  CHEPAITIS, B. A.  P
LEARNING FEAR [FEAR 3]  CHEPAITIS, B. A.  P
BOOK OF NIGHT WITH MOON, THE [FELINE WIZARDS 1]  DUANE, DIANE  C/P,LP
TO VISIT THE QUEEN [FELINE WIZARDS 2]  DUANE, DIANE  C/P,LP
BELLY OF THE BOW, THE [FENCER 2]  PARKER, K. J.  LP
COLOURS IN THE STEEL [FENCER 1]  PARKER, K. J.  LP
PROOF HOUSE, THE [FENCER 3]  PARKER, K. J.  LP
DRAGON CHARMER, THE [FERN CHAPEL 2]  SIEGEL, JAN  P
PROSPERO’S CHILDREN [FERN CHAPEL 1]  SIEGEL, JAN  P
WITCH QUEEN, THE [FERN CHAPEL 3]  SIEGEL, JAN  H
CHANGELING, THE [@FEY 2]  RUSCH, KRISTINE KATHRYN  C/P,P
RESISTANCE, THE [@FEY 4]  RUSCH, KRISTINE KATHRYN  C/P,P
RIVAL, THE [@FEY 3]  RUSCH, KRISTINE KATHRYN  P
SACRIFICE, THE [@FEY 1]  RUSCH, KRISTINE KATHRYN  C/P,P
VICTORY [@FEY 5]  RUSCH, KRISTINE KATHRYN  P
MARCH OR DIE [@FIFTH FOREIGN LEGION #1]  KEITH, ANDREW/KEITH, WILLIAM H. JR.  P
HONOR AND FIDELITY [@FIFTH FOREIGN LEGION #2]  KEITH, ANDREW/KEITH, WILLIAM H. JR.  P
COHORT OF THE DAMNED [@FIFTH FOREIGN LEGION #3]  KEITH, ANDREW  C/P,P
CAGE, THE [FIFTH MILLENNIUM]  STIRLING, S. M./MEIER, SHIRLEY  C/P,SFWA-P:2,P
LION’S HEART [FIFTH MILLENNIUM]  WEHRSTEIN, KAREN  C/P,P
LION’S SOUL [FIFTH MILLENNIUM]  WEHRSTEIN, KAREN  C/P:2,P
SABER AND SHADOW [FIFTH MILLENNIUM]  STIRLING, S. M./MEIER, SHIRLEY  C/P,P
SHADOW’S DAUGHTER [FIFTH MILLENNIUM]  MEIER, SHIRLEY  C/P:2,P

SHADOW’S SON [FIFTH MILLENNIUM] MEIER, SHIRLEY/STIRLING, S. M./

WEHRSTEIN, KAREN

C/P,SFWA-P,P

SHARPEST EDGE, THE [FIFTH MILLENNIUM]  STIRLING, S. M./MEIER, SHIRLEY  C/P,P
SNOWBROTHER [FIFTH MILLENNIUM]  STIRLING, S. M.  C/P:2,P
CHALLENGE OF THE CLANS [FINN MACCUMHAL 1]  FLINT, KENNETH C.  C/P,P
DARK DRUID, THE [FINN MACCUMHAL 3]  FLINT, KENNETH C.  P
UNDERSEA [FINNBRANCH 2]  HAZEL, PAUL  P
WINTERKING [FINNBRANCH 3]  HAZEL, PAUL  H,P
YEARWOOD [FINNBRANCH 1]  HAZEL, PAUL  C/P:2,P
DARKEST ROAD, THE [FIONAVAR TAPESTRY 3]  KAY, GUY GAVRIEL  H:2,SFWA,P
SUMMER TREE, THE [FIONAVAR TAPESTRY 1]  KAY, GUY GAVRIEL  H:2,SFWA,P
WANDERING FIRE, THE [FIONAVAR TAPESTRY 2]  KAY, GUY GAVRIEL  H:2,SFWA,SFWA-P,P

MASTER OF EARTH AND WATER [FIONN MAC CUMHAL 1] PAXSON, DIANA L./

MARTINE-BARNES, ADRIENNE

H,P

SHIELD BETWEEN THE WORLDS, THE [FIONN MAC CUMHAL 2] PAXSON, DIANA L./

MARTINE-BARNES, ADRIENNE

H,P

SWORD OF FIRE AND SHADOW [FIONN MAC CUMHAL 3] PAXSON, DIANA L./

MARTINE-BARNES, ADRIENNE

P

KILN, THE [FIRE-US]  ARMSTRONG, JENNIFER/BUTCHER, NANCY  H

TITLE  AUTHOR  CODE
CROWN OF FIRE [FIREBIRD 3]  TYERS, KATHY  LP
FIREBIRD [FIREBIRD 1]  TYERS, KATHY  C/P,P
FIREBIRD TRILOGY, THE [FIREBIRD 1,2,3]  TYERS, KATHY  H
FUSION FIRE [FIREBIRD 2]  TYERS, KATHY  C/P,P
BURNING LAND, THE [@FIREBRATS #1]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
SURVIVORS [@FIREBRATS #2]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
THUNDER MOUNTAIN [@FIREBRATS #3]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
SHOCKWAVE [@FIREBRATS #4]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
FIREBRINGER TRILOGY, THE [FIREBRINGER 1,2,3]  PIERCE, MEREDITH ANN  H
SON OF SUMMER STARS, THE [FIREBRINGER 3]  PIERCE, MEREDITH ANN  P
DRAGON OF DESPAIR, THE [FIREKEEPER 3]  LINDSKOLD, JANE  P
THROUGH WOLF’S EYES [FIREKEEPER 1]  LINDSKOLD, JANE  P
WOLF CAPTURED [FIREKEEPER 4]  LINDSKOLD, JANE  H
WOLF HUNTING [FIREKEEPER 5]  LINDSKOLD, JANE  H
WOLF’S BLOOD [FIREKEEPER 6]  LINDSKOLD, JANE  H
WOLF’S HEAD, WOLF’S HEART [FIREKEEPER 2]  LINDSKOLD, JANE  P
EDGE OF THE WORLD, THE [FIRST AMERICANS]  SARABONDE, WILLIAM  P
FIRST TRUTH [FIRST TRUTH 1]  COOK, DAWN  P
FORGOTTEN TRUTH [FIRST TRUTH 3]  COOK, DAWN  P
HIDDEN TRUTH [FIRST TRUTH 2]  COOK, DAWN  P
LOST TRUTH [FIRST TRUTH 4]  COOK, DAWN  P
BRIGHT AND THE DARK, THE [FIVE COUNTRIES]  WELCH, MICHELLE M.  C/P,P
CONFIDENCE GAME [FIVE COUNTRIES]  WELCH, MICHELLE M.  P
CONFIDENCE GAME (PROOF) [FIVE COUNTRIES]  WELCH, MICHELLE M.  SR-P
FAIRY GODMOTHER, THE [FIVE HUNDRED KINGDOMS]  LACKEY, MERCEDES  P
FORTUNE’S FOOL [FIVE HUNDRED KINGDOMS]  LACKEY, MERCEDES  H
ONE GOOD KNIGHT [FIVE HUNDRED KINGDOMS]  LACKEY, MERCEDES  H
CRAB, THE [FIVE RINGS CLAN WAR 5]  BROWN, STAN  P
CRANE, THE [FIVE RINGS CLAN WAR 3]  SOESBEE, REE  P
DRAGON, THE [FIVE RINGS CLAN WAR 6]  SOESBEE, REE  P
LION, THE [FIVE RINGS CLAN WAR 7]  SULLIVAN, STEPHEN D.  P
PHOENIX, THE [FIVE RINGS CLAN WAR 4]  SULLIVAN, STEPHEN D.  P
SCORPION, THE [FIVE RINGS CLAN WAR 1]  SULLIVAN, STEPHEN D.  P
UNICORN, THE [FIVE RINGS CLAN WAR 2]  LASSIEUR, A. L.  P
STEEL THRONE, THE [FIVE RINGS FOUR WINDS 0]  BOLME, EDWARD  P
WIND OF HONOR [FIVE RINGS FOUR WINDS #1]  SOESBEE, REE  P
WIND OF WAR [FIVE RINGS FOUR WINDS #2]  LEBOW, JESS  P
WIND OF JUSTICE [@FIVE RINGS FOUR WINDS #3]  WULFF, RICH  P
WIND OF TRUTH [@FIVE RINGS FOUR WINDS #4]  SOESBEE, REE  P
EXILE [@FIVE WORLDS #1]  SARRANTONIO, AL  P
JOURNEY [@FIVE WORLDS #2]  SARRANTONIO, AL  P
RETURN [@FIVE WORLDS #3]  SARRANTONIO, AL  P
LION MEN OF MONGO [FLASH GORDON 1]  STEFFANSON, CON  P
ON THE LOST CONTINENT OF MONGO [FLASH GORDON]  WILLIAMSON, AL  SCX
PLAGUE OF SOUND, THE [FLASH GORDON 2]  STEFFANSON, CON  P
SPACE CIRCUS, THE [FLASH GORDON 3]  STEFFANSON, CON  C/P,P
TIME TRAP OF MING XIII, THE [FLASH GORDON 4]  STEFFANSON, CON  C/P,P
WAR OF THE CYBERNAUTS, THE [FLASH GORDON 6]  BINGHAM, CARSON  C/P:2,P
WITCH QUEEN OF MONGO, THE [FLASH GORDON 5]  BINGHAM, CARSON  C/P:2,P
MASSACRE IN THE 22ND CENTURY [@FLASH GORDON #1]  ANONYMOUS  C/P:2,P
CRISIS ON CITADEL II [@FLASH GORDON #3]  ANONYMOUS  P
FORCES FROM THE FEDERATION [@FLASH GORDON #4]  ANONYMOUS  P
CITADELS UNDER ATTACK [@FLASH GORDON #5]  ANONYMOUS  P
CITADELS ON EARTH [@FLASH GORDON #6]  ANONYMOUS  C/P,P
FLASHMAN [FLASHMAN]  FRASER, GEORGE MACDONALD  P
FLASHMAN AT THE CHARGE [FLASHMAN]  FRASER, GEORGE MACDONALD  P
LORDS OF DARKNESS, THE [@FLAT EARTH 3,4,5]  LEE, TANITH  H:2
NIGHT’S DAUGHTER [@FLAT EARTH 1,2]  LEE, TANITH  H:2
NIGHT’S MASTER [@FLAT EARTH #1]  LEE, TANITH  C/P,P
DEATH’S MASTER [FLAT EARTH #2]  LEE, TANITH  C/P,@P
DELUSION’S MASTER [@FLAT EARTH #3]  LEE, TANITH  C/P:2,P
DELIRIUM’S MISTRESS [@FLAT EARTH #4]  LEE, TANITH  P
NIGHT’S SORCERIES [@FLAT EARTH #5]  LEE, TANITH  P
ESCAPE FROM FLATLAND [FLATLAND]  HAVERSINE, THEDA  SR-P
FLATLAND [FLATLAND]  ABBOTT, EDWIN A.  H:2,LP
FLEET, THE [@FLEET #1]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P
COUNTERATTACK [@FLEET #2]  DRAKE, DAVID/FAWCETT, BILL  PA,C/PA,SFWA-P
BREAKTHROUGH [@FLEET #3]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P
SWORN ALLIES [@FLEET #4]  DRAKE, DAVID/FAWCETT, BILL  PA,SFWA-P:2
TOTAL WAR [@FLEET #5]  DRAKE, DAVID/FAWCETT, BILL  PA,SFWA-P:2
CRISIS [@FLEET #6]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P:2
AGENT OF CHAOS [FLEETS OF THE GREAT BATTLE 2]  ETHERIDGE, RUTLEDGE  P
AGENT OF DESTRUCTION [FLEETS OF THE GREAT BATTLE 1]  ETHERIDGE, RUTLEDGE  P

The MITSFS Seriesdex  September 24, 2007  Page 66
TITLE  AUTHOR  CODE
FLESH AND GOLD [FLESH AND GOLD 1]  GOTLIEB, PHYLLIS  H,LP
MINDWORLDS [FLESH AND GOLD 3]  GOTLIEB, PHYLLIS  H
VIOLENT STARS [FLESH AND GOLD 2]  GOTLIEB, PHYLLIS  H
INDEPENDENT COMMAND, THE [FLIGHT ENGINEER 3]  DOOHAN, JAMES/STIRLING, S. M.  H,P
PRIVATEER, THE [FLIGHT ENGINEER 2]  DOOHAN, JAMES/STIRLING, S. M.  C/P,P
RISING, THE [FLIGHT ENGINEER 1]  DOOHAN, JAMES/STIRLING, S. M.  P
MIRACLE AT SAN TANCO [@FLYING NUN #1]  JOHNSTON, WILLIAM  P
LITTLEST REBELS, THE [@FLYING NUN #2]  JOHNSTON, WILLIAM  C/P,P
MOTHER OF INVENTION [@FLYING NUN #3]  JOHNSTON, WILLIAM  P
LITTLE GREEN MEN, THE [@FLYING NUN #4]  JOHNSTON, WILLIAM  P
UNDERGROUND PICNIC, THE [@FLYING NUN #5]  JOHNSTON, WILLIAM  P
ROSE OF THE WORLD, THE [FOOL’S GOLD 3]  FISHER, JUDE  H
SORCERY RISING [FOOL’S GOLD 1]  FISHER, JUDE  C/P,P
WILD MAGIC [FOOL’S GOLD 2]  FISHER, JUDE  H
FOOTPRINTS OF THUNDER [FOOTPRINTS OF THUNDER 1]  DAVID, JAMES F.  P
THUNDER OF TIME [FOOTPRINTS OF THUNDER 2]  DAVID, JAMES F.  H
REQUIEM FOR THE CONQUEROR [@FORBIDDEN BORDERS #1]  GEAR, W. MICHAEL  C/P:2,SFWA-P:2,P
RELIC OF EMPIRE [@FORBIDDEN BORDERS #2]  GEAR, W. MICHAEL  C/P:2,SFWA-P,P
COUNTERMEASURES [@FORBIDDEN BORDERS #3]  GEAR, W. MICHAEL  C/P:2,P
FORERUNNER [FORERUNNER 4/DIPPLE 4]  NORTON, ANDRE  C/P:2,SFWA-P,P
FORERUNNER FORAY [FORERUNNER 3]  NORTON, ANDRE  H:2,P
FORERUNNER: THE SECOND VENTURE [FORERUNNER 5]  NORTON, ANDRE  H:2,SFWA,SFWA-P,P
ORDEAL IN OTHERWHERE [FORERUNNER 2]  NORTON, ANDRE  C/P:2,P
STORM OVER WARLOCK [FORERUNNER 1]  NORTON, ANDRE  C/P,P,H
DAWN FOR A DISTANT EARTH [FOREVER HERO 1]  MODESITT, L. E. JR.  C/P:2,SFWA-P,P
FOREVER HERO, THE [FOREVER HERO 1,2,3]  MODESITT, L. E. JR.  H
IN ENDLESS TWILIGHT [FOREVER HERO 3]  MODESITT, L. E. JR.  C/P,P
SILENT WARRIOR, THE [FOREVER HERO 2]  MODESITT, L. E. JR.  C/P,P
BROKEN SWORD, THE [FOREVER KING 2]  COCHRAN, MOLLY/MURPHY, WARREN  C/P,P,H
FOREVER KING, THE [FOREVER KING 1]  COCHRAN, MOLLY/MURPHY, WARREN  H:2,P
THIRD MAGIC, THE [FOREVER KING 3]  COCHRAN, MOLLY  P
INTIMATIONS OF MORTALITY [FOREVER KNIGHT]  GARRETT, SUSAN M.  P
THESE OUR REVELS [@FOREVER KNIGHT]  HATHAWAY-NAYNE, ANNE  P
ANVIL OF STARS [FORGE OF GOD 2]  BEAR, GREG  H:2,P
FORGE OF GOD, THE [FORGE OF GOD 1]  BEAR, GREG  H:2,SFWA-P,P
BEST OF THE REALMS, BOOK II, THE [@FORGOTTEN REALMS]  GREENWOOD, ED  P
BEST OF THE REALMS BOOK III, THE [@FORGOTTEN REALMS]  CUNNINGHAM, ELAINE  P
BEST OF THE REALMS BOOK I, THE [@FORGOTTEN REALMS]  WIZARDS OF THE COAST  PA
DREAM SPHERES [@FORGOTTEN REALMS]  CUNNINGHAM, ELAINE  P
EVERMEET: ISLAND OF ELVES [FORGOTTEN REALMS]  CUNNINGHAM, ELAINE  P
FORGOTTEN REALMS ATLAS, THE [FORGOTTEN REALMS]  FONSTAD, KAREN WYNN  A
GLASS PRISON, THE [@FORGOTTEN REALMS]  COOK, MONTE  P
LADY OF POISON [@FORGOTTEN REALMS]  CORDELL, BRUCE R.  P
MURDER IN CORMYR [@FORGOTTEN REALMS]  WILLIAMSON, CHET  P
ONCE AROUND THE REALMS [@FORGOTTEN REALMS]  THOMSEN, BRIAN M.  P
POOL OF RADIANCE [@FORGOTTEN REALMS]  WARD, JAMES M./HONG, JANE COOPER  C/P:2,P
POOL OF TWILIGHT [@FORGOTTEN REALMS]  WARD, JAMES M./BROWN, ANNE K.  C/P,P
POOLS OF DARKNESS [@FORGOTTEN REALMS]  WARD, JAMES M./BROWN, ANNE K.  C/P,P
REALMS OF INFAMY [FORGOTTEN REALMS]  LOWDER, JAMES  PA
REALMS OF MAGIC [@FORGOTTEN REALMS]  THOMSEN, BRIAN M./KING, J. ROBERT  PA
REALMS OF MYSTERY [@FORGOTTEN REALMS]  ATHANS, PHILIP  PA
REALMS OF THE ARCANE [@FORGOTTEN REALMS]  THOMSEN, BRIAN M.  PA
REALMS OF VALOR [FORGOTTEN REALMS]  LOWDER, JAMES  PA
SEA OF SWORDS [@FORGOTTEN REALMS]  SALVATORE, R. A.  P
SHADOW STONE [@FORGOTTEN REALMS]  BAKER, RICHARD  P
SILVERFALL [@FORGOTTEN REALMS]  GREENWOOD, ED  P
SPINE OF THE WORLD, THE [@FORGOTTEN REALMS]  SALVATORE, R. A.  P
TEMPTATION OF ELMINSTER, THE [@FORGOTTEN REALMS]  GREENWOOD, ED  P
SWORD PLAY [@FORGOTTEN REALMS ARCANE AGE NETHERIL #1]  EMERY, CLAYTON  P
DANGEROUS GAMES [@FORGOTTEN REALMS ARCANE AGE NETHERIL #2]  EMERY, CLAYTON  P

MORTAL CONSEQUENCES [@FORGOTTEN REALMS ARCANE AGE NETHERIL #3] EMERY, CLAYTON

P

CRUCIBLE: THE TRIAL OF CYRIC THE MAD [@FORGOTTEN REALMS AVATAR 5] DENNING, TROY

P

PRINCE OF LIES [FORGOTTEN REALMS AVATAR 4]  LOWDER, JAMES  P
SHADOWDALE [FORGOTTEN REALMS AVATAR 1]  AWLINSON, RICHARD  C/P:2,P
TANTRAS [FORGOTTEN REALMS AVATAR 2]  AWLINSON, RICHARD  C/P:2,P
WATERDEEP [FORGOTTEN REALMS AVATAR 3]  AWLINSON, RICHARD  C/P:2,P
SHADOWDALE [@FORGOTTEN REALMS AVATAR #1]  CIENCIN, SCOTT  P
TANTRAS [@FORGOTTEN REALMS AVATAR #2]  CIENCIN, SCOTT  P
WATERDEEP [@FORGOTTEN REALMS AVATAR #3]  DENNING, TROY  P
BALDUR’S GATE [@FORGOTTEN REALMS BALDUR’S GATE I]  ATHANS, PHILIP  P
SHADOWS OF AMN [@FORGOTTEN REALMS BALDUR’S GATE II]  ATHANS, PHILIP  P

The MITSFS Seriesdex  September 24, 2007  Page 67
TITLE  AUTHOR  CODE
THRONE OF BHAAL [@FORGOTTEN REALMS BALDUR’S GATE II]  KARPYSHYN, DREW  P
CITY OF RAVENS, THE [@FORGOTTEN REALMS CITIES]  BAKER, RICHARD  P
CITY OF SPLENDORS, THE [@FORGOTTEN REALMS CITIES]  GREENWOOD, ED/CUNNINGHAM, ELAINE  P
JEWEL OF TURMISH, THE [@FORGOTTEN REALMS CITIES]  ODOM, MEL  P
TEMPLE HILL [FORGOTTEN REALMS CITIES]  KARPYSHYN, DREW  P
CANTICLE [@FORGOTTEN REALMS CLERIC #1]  SALVATORE, R. A.  P
IN SYLVAN SHADOWS [@FORGOTTEN REALMS CLERIC #2]  SALVATORE, R. A.  P
NIGHT MASKS [@FORGOTTEN REALMS CLERIC #3]  SALVATORE, R. A.  P
FALLEN FORTRESS, THE [@FORGOTTEN REALMS CLERIC #4]  SALVATORE, R. A.  P
CHAOS CURSE, THE [@FORGOTTEN REALMS CLERIC #5]  SALVATORE, R. A.  C/P,P
BEYOND THE HIGH ROAD [@FORGOTTEN REALMS CORMYR 2]  DENNING, TROY  P
CORMYR [@FORGOTTEN REALMS CORMYR #1]  GREENWOOD, ED/GRUBB, JEFF  P
DEATH OF THE DRAGON [@FORGOTTEN REALMS CORMYR #3]  GREENWOOD, ED/DENNING, TROY  P
MAGEHOUND, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #1]  CUNNINGHAM, ELAINE  P
FLOODGATE, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #2]  CUNNINGHAM, ELAINE  P
WIZARDWAR, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #3]  CUNNINGHAM, ELAINE  P
DARK ELF TRILOGY, THE [@FORGOTTEN REALMS DARK ELF]  SALVATORE, R. A.  LP
SOJOURN [@FORGOTTEN REALMS DARK ELF 3]  SALVATORE, R. A.  C/P,P
HOMELAND [@FORGOTTEN REALMS DARK ELF #1]  SALVATORE, R. A.  C/P:2,P
EXILE [@FORGOTTEN REALMS DARK ELF #2]  SALVATORE, R. A.  C/P,P

ABDUCTION, THE [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #1] KING, J. ROBERT

P

PALADINS, THE [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #2]  WARD, JAMES S./WISE, DAVID  P

ERRAND OF MERCY [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #4] MOORE, ROGER E.

P

OPPORTUNITY FOR PROFIT, AN [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #5] GROSS, DAVE

P

PROPHET OF MOONSHAE [@FORGOTTEN REALMS DRUIDHOME #1]  NILES, DOUGLAS  C/P,P
CORAL KINGDOM, THE [@FORGOTTEN REALMS DRUIDHOME #2]  NILES, DOUGLAS  C/P,P
DRUID QUEEN, THE [@FORGOTTEN REALMS DRUIDHOME #3]  NILES, DOUGLAS  C/P,P
DEPTHS OF MADNESS [@FORGOTTEN REALMS DUNGEONS]  DE BIE, ERIK SCOTT  P
HOWLING DELVE, THE [@FORGOTTEN REALMS DUNGEONS]  JOHNSON, JALEIGH  P
ELMINSTER IN HELL [@FORGOTTEN REALMS ELMINSTER]  GREENWOOD, ED  P
ELMINSTER IN MYTH DRANNOR [@FORGOTTEN REALMS ELMINSTER]  GREENWOOD, ED  H

ELMINSTER: THE MAKING OF A MAGE [@FORGOTTEN REALMS ELMINSTER] GREENWOOD, ED

C/P,P,H

ELMINSTER’S DAUGHTER [@FORGOTTEN REALMS ELMINSTER]  GREENWOOD, ED  H
CRUSADE [FORGOTTEN REALMS EMPIRES 3]  LOWDER, JAMES  C/P:2,P
DRAGONWALL [FORGOTTEN REALMS EMPIRES 2]  DENNING, TROY  C/P:2,P
HORSELORDS [FORGOTTEN REALMS EMPIRES 1]  COOK, DAVID  C/P:2,P
GOSSAMER PLAIN, THE [@FORGOTTEN REALMS EMPYREAN ODYSSEY #1]  REID, THOMAS M.  P
TWILIGHT FALLING [@FORGOTTEN REALMS EREVIS CALE #1]  KEMP, PAUL S.  P
DAWN OF NIGHT [@FORGOTTEN REALMS EREVIS CALE #2]  KEMP, PAUL S.  P
MIDNIGHT’S MASK [@FORGOTTEN REALMS EREVIS CALE #3]  KEMP, PAUL S.  P
BLADESINGER [@FORGOTTEN REALMS FIGHTERS]  STROHM, KEITH FRANCIS  P
GHOSTWALKER [@FORGOTTEN REALMS FIGHTERS]  DEBIE, ERIK SCOTT  P
MASTER OF CHAINS [@FORGOTTEN REALMS FIGHTERS]  LEBOW, JESS  P
SON OF THUNDER [@FORGOTTEN REALMS FIGHTERS]  LEEDER, MURRAY J. D.  P
AZURE BONDS [FORGOTTEN REALMS FINDER’S STONE 1]  NOVAK, KATE/GRUBB, JEFF  C/P:2,P
SONG OF THE SAURIALS [FORGOTTEN REALMS FINDER’S STONE 3]  NOVAK, KATE/GRUBB, JEFF  P
WYVERN’S SPUR, THE [FORGOTTEN REALMS FINDER’S STONE 2]  NOVAK, KATE/GRUBB, JEFF  C/P,P
FINDER’S BANE [@FORGOTTEN REALMS HARPERS]  NOVAK, KATE/GRUBB, JEFF  P
MASQUERADES [FORGOTTEN REALMS HARPERS 10]  NOVAK, KATE/GRUBB, JEFF  P
THORNHOLD [@FORGOTTEN REALMS HARPERS]  CUNNINGHAM, ELAINE  P
PARCHED SEA, THE [@FORGOTTEN REALMS HARPERS #1]  DENNING, TROY  C/P,P
ELFSHADOW [@FORGOTTEN REALMS HARPERS #2]  CUNNINGHAM, ELAINE  C/P:2,P
RED MAGIC [@FORGOTTEN REALMS HARPERS #3]  RABE, JEAN  P
NIGHT PARADE, THE [FORGOTTEN REALMS HARPERS #4]  CIENCIN, SCOTT  C/P,P
RING OF WINTER, THE [FORGOTTEN REALMS HARPERS #5]  LOWDER, JAMES  P
CRYPT OF THE SHADOWKING [@FORGOTTEN REALMS HARPERS #6]  ANTHONY, MARK  P
SOLDIERS OF ICE [@FORGOTTEN REALMS HARPERS #7]  COOK, DAVID  P
ELFSONG [@FORGOTTEN REALMS HARPERS #8]  CUNNINGHAM, ELAINE  C/P,P
CROWN OF FIRE [@FORGOTTEN REALMS HARPERS #9/SHANDRIL 2]  GREENWOOD, ED  C/P:2,P
CURSE OF THE SHADOWMAGE [@FORGOTTEN REALMS HARPERS #11]  ANTHONY, MARK  P
VEILED DRAGON, THE [@FORGOTTEN REALMS HARPERS #12]  DENNING, TROY  P
SILVER SHADOWS [@FORGOTTEN REALMS HARPERS #13]  CUNNINGHAM, ELAINE  P
STORMLIGHT [@FORGOTTEN REALMS HARPERS #14]  GREENWOOD, ED  P
UNCLEAN [@FORGOTTEN REALMS HAUNTED LANDS 1]  BYERS, RICHARD LEE  P
VENOM’S TASTE [@FORGOTTEN REALMS HOUSE OF SERPENTS #1]  SMEDMAN, LISA  P
VIPER’S KISS [@FORGOTTEN REALMS HOUSE OF SERPENTS #2]  SMEDMAN, LISA  P
VANITY’S BROOD [@FORGOTTEN REALMS HOUSE OF SERPENTS #3]  SMEDMAN, LISA  P
THOUSAND ORCS, THE [@FORGOTTEN REALMS HUNTER’S BLADES #1]  SALVATORE, R. A.  P
LONE DROW, THE [@FORGOTTEN REALMS HUNTER’S BLADES #2]  SALVATORE, R. A.  P

The MITSFS Seriesdex  September 24, 2007  Page 68
TITLE  AUTHOR  CODE
TWO SWORDS, THE [@FORGOTTEN REALMS HUNTER’S BLADES #3]  SALVATORE, R. A.  H
CRYSTAL SHARD, THE [@FORGOTTEN REALMS ICEWIND DALE #1]  SALVATORE, R. A.  C/P:2,P
STREAMS OF SILVER [@FORGOTTEN REALMS ICEWIND DALE #2]  SALVATORE, R. A.  C/P:2,P
HALFLING’S GEM, THE [@FORGOTTEN REALMS ICEWIND DALE #3]  SALVATORE, R. A.  P

SWORDS OF EVENINGSTAR [@FORGOTTEN REALMS KNIGHTS OF MYTH DRANNOR #1] GREENWOOD, ED

H

SACRIFICE OF THE WIDOW [@FORGOTTEN REALMS LADY PENITENT #1]  SMEDMAN, LISA  C/P,P
REALMS OF THE ELVES [@FORGOTTEN REALMS LAST MYTHAL]  ATHANS, PHILIP  PA
FORSAKEN HOUSE [@FORGOTTEN REALMS LAST MYTHAL #1]  BAKER, RICHARD  P
FARTHEST REACH [@FORGOTTEN REALMS LAST MYTHAL #2]  BAKER, RICHARD  P
FINAL GATE [@FORGOTTEN REALMS LAST MYTHAL #3]  BAKER, RICHARD  P
LEGACY, THE [FORGOTTEN REALMS LEGACY OF THE DROW 1]  SALVATORE, R. A.  C/P,P,H
PASSAGE TO DAWN [FORGOTTEN REALMS LEGACY OF THE DROW 4]  SALVATORE, R. A.  P
SIEGE OF DARKNESS [FORGOTTEN REALMS LEGACY OF THE DROW 3]  SALVATORE, R. A.  P
STARLESS NIGHT [FORGOTTEN REALMS LEGACY OF THE DROW 2]  SALVATORE, R. A.  H,P
FACES OF DECEPTION [@FORGOTTEN REALMS LOST EMPIRES]  DENNING, TROY  P

LOST LIBRARY OF CORMANTHYR, THE [@FORGOTTEN REALMS LOST EMPIRES] ODOM, MEL

P

NETHER SCROLL, THE [@FORGOTTEN REALMS LOST EMPIRES]  ABBEY, LYNN  P
STAR OF CURSRAH [FORGOTTEN REALMS LOST EMPIRES]  EMERY, CLAYTON  P
TYMORA’S LUCK [@FORGOTTEN REALMS LOST GODS]  NOVAK, KATE/GRUBB, JEFF  P
IRONHELM [@FORGOTTEN REALMS MAZTICA #1]  NILES, DOUGLAS  P
VIPERHAND [@FORGOTTEN REALMS MAZTICA #2]  NILES, DOUGLAS  C/P:2,P
FEATHERED DRAGON, THE [@FORGOTTEN REALMS MAZTICA #3]  NILES, DOUGLAS  P
DARKWALKER ON MOONSHAE [@FORGOTTEN REALMS MOONSHAE #1]  NILES, DOUGLAS  C/P:2,P
BLACK WIZARDS [@FORGOTTEN REALMS MOONSHAE #2]  NILES, DOUGLAS  C/P,P
DARKWELL [@FORGOTTEN REALMS MOONSHAE #3]  NILES, DOUGLAS  C/P:2,P
WAR IN TETHYR [FORGOTTEN REALMS NOBLES 2]  MILAN, VICTOR  P
KING PINCH [@FORGOTTEN REALMS NOBLES #1]  COOK, DAVID  P
ESCAPE FROM UNDERMOUNTAIN [@FORGOTTEN REALMS NOBLES #3]  ANTHONY, MARK  P
MAGE IN THE IRON MASK, THE [@FORGOTTEN REALMS NOBLES #4]  THOMSEN, BRIAN M.  P
COUNCIL OF BLADES, THE [@FORGOTTEN REALMS NOBLES #6]  KIDD, PAUL  P
SIMBUL’S GIFT, THE [@FORGOTTEN REALMS NOBLES #6]  ABBEY, LYNN  P
SILENT BLADE, THE [FORGOTTEN REALMS PATHS OF DARKNESS 1]  SALVATORE, R. A.  H
MAIDEN OF PAIN [@FORGOTTEN REALMS PRIESTS]  FRANKLIN, KAMERON M.  P
MISTRESS OF THE NIGHT [@FORGOTTEN REALMS PRIESTS]  BASSINGTHWAITE, DON/GROSS, DAVE  P
QUEEN OF THE DEPTHS [@FORGOTTEN REALMS PRIESTS]  BYERS, RICHARD LEE  P

REALMS OF SHADOW [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS] BALDWIN, LIZZ

PA

SUMMONING, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #1] DENNING, TROY

P

SIEGE, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #2]  DENNING, TROY  P

SORCERER, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #3] DENNING, TROY

P

ALABASTER STAFF, THE [@FORGOTTEN REALMS ROGUES]  BOLME, EDWARD  P
BLACK BOUQUET, THE [@FORGOTTEN REALMS ROGUES]  BYERS, RICHARD LEE  P
CRIMSON GOLD, THE [@FORGOTTEN REALMS ROGUES]  WHITNEY-ROBINSON, VORONICA  P
YELLOW SILK, THE [@FORGOTTEN REALMS ROGUES]  BASSINGTHWAITE, DON  P
POOL OF RADIANCE [@FORGOTTEN REALMS RUINS OF MYTH DRANNOR]  BEBRIS, CARRIE A.  P

SAPPHIRE CRESCENT, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #1] REID, THOMAS M.

P

RUBY GUARDIAN, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #2]  REID, THOMAS M.  P

EMERALD SCEPTER, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #3] REID, THOMAS M.

P

SERVANT OF THE SHARD [@FORGOTTEN REALMS SELLSWORDS #1]  SALVATORE, R. A.  P
PROMISE OF THE WITCHKING [@FORGOTTEN REALMS SELLSWORDS #2]  SALVATORE, R. A.  P
ROAD OF THE PATRIARCH [@FORGOTTEN REALMS SELLSWORDS #3]  SALVATORE, R. A.  H
BLACK WOLF [@FORGOTTEN REALMS SEMBIA]  GROSS, DAVE  P

HALLS OF STORMWEATHER, THE [@FORGOTTEN REALMS SEMBIA] BYERS, RICHARD LEE/EMERY, CLAYTON/ GREENWOOD, ED/GROSS, DAVE/ KEMP, PAUL S./SMEDMAN, LISA/ WHITNEY-ROBINSON, VORONICA

P

HEIRS OF PROPHECY [@FORGOTTEN REALMS SEMBIA]  SMEDMAN, LISA  P
LORD OF STORMWEATHER [@FORGOTTEN REALMS SEMBIA]  GROSS, DAVE  P
SANDS OF THE SOUL [@FORGOTTEN REALMS SEMBIA]  WHITNEY-ROBINSON, VORONICA  P
SHADOW’S WITNESS [@FORGOTTEN REALMS SEMBIA]  KEMP, PAUL S.  P
SHATTERED MASK, THE [@FORGOTTEN REALMS SEMBIA]  BYERS, RICHARD LEE  P

HALLS OF STORMWEATHER, THE [@FORGOTTEN REALMS SEMBIA GATEWAY TO THE REALMS #1] ATHANS, PHILIP

PA

SHADOWS OF DOOM [@FORGOTTEN REALMS SHADOW OF AVATAR #1]  GREENWOOD, ED  P
CLOAK OF SHADOWS [@FORGOTTEN REALMS SHADOW OF AVATAR #2]  GREENWOOD, ED  P
ALL SHADOWS FLED [@FORGOTTEN REALMS SHADOW OF AVATAR #3]  GREENWOOD, ED  P
SPELLFIRE [FORGOTTEN REALMS SHANDRIL 1]  GREENWOOD, ED  C/P:2,P
HAND OF FIRE [@FORGOTTEN REALMS SHANDRIL #3]  GREENWOOD, ED  P

TITLE  AUTHOR  CODE

DAUGHTER OF THE DROW [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 1] CUNNINGHAM, ELAINE

P

TANGLED WEBS [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 2]  CUNNINGHAM, ELAINE  P
WINDWALKER [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 3]  CUNNINGHAM, ELAINE  P
REALMS OF THE DEEP [@FORGOTTEN REALMS THREAT FROM THE SEA]  ATHANS, PHILIP  PA
RISING TIDE [@FORGOTTEN REALMS THREAT FROM THE SEA #1]  ODOM, MEL  P

UNDER FALLEN STARS [@FORGOTTEN REALMS THREAT FROM THE SEA #2] ODOM, MEL

P

SEA DEVIL’S EYE, THE [@FORGOTTEN REALMS THREAT FROM THE SEA #3] ODOM, MEL

P

OGRE’S PACT, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #1]  DENNING, TROY  P
GIANT AMONG US, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #2]  DENNING, TROY  P
TITAN OF TWILIGHT, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #3]  DENNING, TROY  P
SHADOWBRED [@FORGOTTEN REALMS TWILIGHT WAR #1]  KEMP, PAUL  P
REALMS OF THE UNDERDARK [@FORGOTTEN REALMS UNDERDARK]  KING, J. ROBERT  PA
DISSOLUTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #1]  BYERS, RICHARD LEE  P
INSURRECTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #2]  REID, THOMAS M.  H
CONDEMNATION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #3]  BAKER, RICHARD  H,P
EXTINCTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #4]  SMEDMAN, LISA  P
ANNIHILATION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #5]  ATHANS, PHILIP  P
RESURRECTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #6]  KEMP, PAUL S.  H
WHISPER OF WAVES [@FORGOTTEN REALMS WATERCOURSE #1]  ATHANS, PHILIP  P
SCREAM OF STONE [@FORGOTTEN REALMS WATERCOURSE #3]  ATHANS, PHILIP  P
BLACKSTAFF [FORGOTTEN REALMS WIZARDS]  SCHEND, STEVEN E  P
BLOODWALK [FORGOTTEN REALMS WIZARDS 5]  DAVIS, JAMES P  P
DARKVISION [@FORGOTTEN REALMS WIZARDS]  CORDELL, BRUCE R.  P
FROSTFELL [@FORGOTTEN REALMS WIZARDS]  SEHESTEDT, MARK  P

REALMS OF THE DRAGONS [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS] ATHANS, PHILIP

PA

REALMS OF THE DRAGONS II [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS] ATHANS, PHILIP

PA

RAGE, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #1]  BYERS, RICHARD LEE  P
RITE, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #2]  BYERS, RICHARD LEE  P
RUIN, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #3]  BYERS, RICHARD LEE  P
DRAGONCHARM [FORTUNE 1]  EDWARDS, GRAHAM  C/P,P
DRAGONSTORM [FORTUNE 2]  EDWARDS, GRAHAM  P
FIFTY DEGREES BELOW [FORTY SIGNS OF RAIN 2]  ROBINSON, KIM STANLEY  H:2
FORTY SIGNS OF RAIN [FORTY SIGNS OF RAIN 1]  ROBINSON, KIM STANLEY  P
SIXTY DAYS AND COUNTING [FORTY SIGNS OF RAIN 3]  ROBINSON, KIM STANLEY  H
FONDAZIONE E TERRA (ITALIAN) [FOUNDATION 5]  ASIMOV, ISAAC  FH
FORWARD THE FOUNDATION [FOUNDATION]  ASIMOV, ISAAC  C/P,P,H:2
FOUNDATION [FOUNDATION 1]  ASIMOV, ISAAC  C/P:2,P
FOUNDATION (REVISED) [FOUNDATION 1]  ASIMOV, ISAAC  P
FOUNDATION AND EARTH [FOUNDATION 5]  ASIMOV, ISAAC  H:3,SFWA,P
FOUNDATION AND EMPIRE [FOUNDATION 2]  ASIMOV, ISAAC  C/P:2,P,SR-H,H
FOUNDATION TRILOGY, THE [FOUNDATION 1,2,3]  ASIMOV, ISAAC  H:2,LP
FOUNDATION’S EDGE [FOUNDATION 4]  ASIMOV, ISAAC  C/P,P,H:2,SFWA
FOUNDATION’S FRIENDS [FOUNDATION]  GREENBERG, MARTIN H.  C/PA,SFWA-P,LPA
PRELUDE TO FOUNDATION [FOUNDATION 0]  ASIMOV, ISAAC  C/P,P,H
SECOND FOUNDATION [FOUNDATION 3]  ASIMOV, ISAAC  C/P:2,P,H:3,SR-P,SR-H
SECONDE FONDATION [FOUNDATION 3]  ASIMOV, ISAAC  FP
FAITH OF THE UNFORGOTTEN [@FOUNDATIONS OF HOPE #1]  GRODY, LEATHEL  H

FOUR LORDS OF THE DIAMOND, THE [FOUR LORDS OF THE DIAMOND 1,2,3,4] CHALKER, JACK L.

H:2

LILITH: A SNAKE IN THE GRASS [@FOUR LORDS OF THE DIAMOND #1]  CHALKER, JACK L.  C/P:2,P
CERBERUS: A WOLF IN THE FOLD [@FOUR LORDS OF THE DIAMOND #2]  CHALKER, JACK L.  C/P:2,P
CHARON: A DRAGON AT THE GATE [@FOUR LORDS OF THE DIAMOND #3]  CHALKER, JACK L.  C/P:2,P
MEDUSA: A TIGER BY THE TAIL [@FOUR LORDS OF THE DIAMOND #4]  CHALKER, JACK L.  C/P:2,P
YEAR OF OUR WAR, THE [FOURLANDS 1]  SWAINSTON, STEPH  H
TRAVELER, THE [FOURTH REALM 1]  TWELVE HAWKS, JOHN  LP
DARK RIVER, THE [@FOURTH REALM #2]  TWELVE HAWKS, JOHN  H
CRUCIBLE [FOX 2]  CHASE, ROBERT R.  P
GAME OF FOX AND LION, THE [FOX 1]  CHASE, ROBERT R.  P
CARLUCCI’S EDGE [FRANK CARLUCCI 2]  RUSSO, RICHARD PAUL  P
CARLUCCI’S HEART [FRANK CARLUCCI 3]  RUSSO, RICHARD PAUL  P
DESTROYING ANGEL [FRANK CARLUCCI 1]  RUSSO, RICHARD PAUL  C/P,P
FREEDOM’S RANGERS [@FREEDOM’S RANGERS 1]  ANDREWS, KEITH WILLIAM  C/P,P
RAIDERS OF THE REVOLUTION [@FREEDOM’S RANGERS 2]  ANDREWS, KEITH WILLIAM  P
SEARCH AND DESTROY [@FREEDOM’S RANGERS #3]  ANDREWS, KEITH WILLIAM  P
TREASON IN TIME [@FREEDOM’S RANGERS #4]  ANDREWS, KEITH WILLIAM  P
SINK THE ARMADA! [@FREEDOM’S RANGERS #5]  ANDREWS, KEITH WILLIAM  P
SNOW KILL [@FREEDOM’S RANGERS #6]  ANDREWS, KEITH WILLIAM  P
FREEHOLD [FREEHOLD WAR 1]  WILLIAMSON, MICHAEL Z.  P
WEAPON, THE [FREEHOLD WAR 2]  WILLIAMSON, MICHAEL Z.  H

TITLE  AUTHOR  CODE
F.R.E.E.FALL [F.R.E.E.LANCERS 2]  ODOM, MEL  P
F.R.E.E.LANCERS [F.R.E.E.LANCERS 1]  ODOM, MEL  P
FRONTIER EARTH [@FRONTIER EARTH 1]  BOXLEITNER, BRUCE  H
SEARCHER [@FRONTIER EARTH 2]  BOXLEITNER, BRUCE  P
NEW MIND, THE [FRONTIERS 2]  ELWOOD, ROGER  C/PA,PA
TOMORROW’S ALTERNATIVES [FRONTIERS 1]  ELWOOD, ROGER  HA,PA
BLOODSONGS [FROST 3]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
FROST [FROST 1]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
SKULL GATE [FROST 2]  BAILEY, ROBIN WAYNE  C/P,SFWA-P,P
FROSTFLOWER AND THORN [FROSTFLOWER]  KARR, PHYLLIS ANN  C/P:2,P
FROSTFLOWER AND WINDBOURNE [FROSTFLOWER]  KARR, PHYLLIS ANN  C/P:2,P
THOR’S HAMMER [@FUTURE AT WAR #1]  BRETNOR, REGINALD  SFWA-P,C/PA,PA
SPEAR OF MARS, THE [@FUTURE AT WAR #2]  BRETNOR, REGINALD  PA,C/PA,SFWA-P
ORION’S SWORD [@FUTURE AT WAR #3]  BRETNOR, REGINALD  PA,C/PA,SFWA-P
UNDER AFRICAN SKIES [@FUTURE EARTHS]  RESNICK, MIKE/DOZOIS, GARDNER  C/PA,PA
UNDER SOUTH AMERICAN SKIES [@FUTURE EARTHS]  RESNICK, MIKE/DOZOIS, GARDNER  PA
FUZZIES AND OTHER PEOPLE [FUZZIES 3]  PIPER, H. BEAM  H,P
FUZZY BONES [FUZZIES]  TUNING, WILLIAM  C/P:2,P
FUZZY PAPERS, THE [FUZZIES 1,2]  PIPER, H. BEAM  H:2,P
FUZZY SAPIENS [FUZZIES 2]  PIPER, H. BEAM  P
GOLDEN DREAM: A FUZZY ODYSSEY [FUZZIES]  MAYHAR, ARDATH  C/P,LP
LITTLE FUZZY [FUZZIES 1]  PIPER, H. BEAM  C/P:2,P
OTHER HUMAN RACE, THE [FUZZIES 2]  PIPER, H. BEAM  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 71
TITLE  AUTHOR  CODE
BAT STAFFEL, THE [@G-8 #1]  HOGAN, ROBERT J.  SR-P
SCOURGE OF THE STEEL MASK [@G-8 #1]  HOGAN, ROBERT J.  P
PURPLE ACES [@G-8 #2]  HOGAN, ROBERT J.  P
ACE OF THE WHITE DEATH [@G-8 #3]  HOGAN, ROBERT J.  SR-P
BOMBS FROM THE MURDER WOLVES [@G-8 #4]  HOGAN, ROBERT J.  SR-P
VULTURES OF THE WHITE DEATH [@G-8 #5]  HOGAN, ROBERT J.  SR-P
FLIGHT FROM THE GRAVE [@G-8 #6]  HOGAN, ROBERT J.  P
MARK OF THE VULTURE, THE [@G-8 #8]  HOGAN, ROBERT J.  SR-P
SINS OF THE FATHERS [@GABRIEL KNIGHT]  JENSEN, JANE  P
DEMON [GAEA 3]  VARLEY, JOHN  H:2,LP
TITAN [GAEA 1]  VARLEY, JOHN  H:2,P
WIZARD [GAEA 2]  VARLEY, JOHN  H:2,P,SFWA
ACROSS THE SEA OF SUNS (ORIGINAL) [GALACTIC CENTER 2]  BENFORD, GREGORY  H:2

ACROSS THE SEA OF SUNS (REVISED: EXTRA CHAPTER AT END)

[GALACTIC CENTER 2] BENFORD, GREGORY

C/P,P

FURIOUS GULF [GALACTIC CENTER 5]  BENFORD, GREGORY  C/P,P,H
GREAT SKY RIVER [GALACTIC CENTER 3]  BENFORD, GREGORY  C/P,SFWA,P,H
IN THE OCEAN OF NIGHT [GALACTIC CENTER 1]  BENFORD, GREGORY  C/P,P,H
SAILING BRIGHT ETERNITY [GALACTIC CENTER 6]  BENFORD, GREGORY  C/P:2,P
TIDES OF LIGHT [GALACTIC CENTER 4]  BENFORD, GREGORY  C/P:2,SFWA-P,P,H
CURRENTS OF SPACE, THE [GALACTIC EMPIRE 2]  ASIMOV, ISAAC  C/P:2,P,H
PEBBLE IN THE SKY [GALACTIC EMPIRE 3]  ASIMOV, ISAAC  C/P,P,H,GN3.14
STARS, LIKE DUST, THE [GALACTIC EMPIRE 1]  ASIMOV, ISAAC  C/P:2,P,H
TRIANGLE [GALACTIC EMPIRE 1,2,3]  ASIMOV, ISAAC  H
CITADEL, THE [GALACTIC MI]  RANDLE, KEVIN  P
GALACTIC MI [GALACTIC MI]  RANDLE, KEVIN  P
RAT TRAP, THE [GALACTIC MI 2]  RANDLE, KEVIN  C/P,P

BEST ROOTIN’ TOOTIN’ SHOOTIN’ GUNSLINGER IN THE WHOLE DAMNED GALAXY, THE [GALACTIC MIDWAY 4] RESNICK, MIKE

C/P:2,P

SIDESHOW [GALACTIC MIDWAY 1]  RESNICK, MIKE  C/P,P
THREE-LEGGED HOOTCH DANCER, THE [GALACTIC MIDWAY 2]  RESNICK, MIKE  C/P,P
WILD ALIEN TAMER, THE [GALACTIC MIDWAY 3]  RESNICK, MIKE  C/P,P
DIAMOND MASK [GALACTIC MILIEU 2]  MAY, JULIAN  C/P,P,H
JACK THE BODILESS [GALACTIC MILIEU 1]  MAY, JULIAN  H:2,P
MAGNIFICAT [GALACTIC MILIEU 3]  MAY, JULIAN  C/P,P,H
COMET’S CURSE, THE [GALAHAD 1]  TESTA, DOM  LP

NIGHTFEEDER [GALEN SWORD 2] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

P

SHIFTER [@GALEN SWORD #1] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P:2,P

ABYSS OF LIGHT, AN [GAMANT 1]  O’NEAL, KATHLEEN M.  P
REDEMPTION OF LIGHT [GAMANT 3]  O’NEAL, KATHLEEN M.  P
TREASURE OF LIGHT [GAMANT 2]  O’NEAL, KATHLEEN N.  P
SIX FAMILIES, THE [@GAMBLER’S STAR #1]  HOLDER, NANCY  P
LEGACIES AND LIES [@GAMBLER’S STAR #2]  HOLDER, NANCY  P
INVASIONS [@GAMBLER’S STAR #3]  HOLDER, NANCY  P
GAMEARTH [GAMEARTH 1]  ANDERSON, KEVIN J.  P
GAMEPLAY [GAMEARTH 2]  ANDERSON, KEVIN J.  P
ALEXANDRIAN RING, THE [GAMESTER WARS 1]  FORSTCHEN, WILLIAM R.  C/P:2,P
ASSASSIN GAMBIT, THE [GAMESTER WARS 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
NAPOLEON WAGER, THE [GAMESTER WARS 3]  FORSTCHEN, WILLIAM R.  P
DESTINY DICE, THE [GAMING MAGI 1]  BISCHOFF, DAVID F.  C/P:2,P
UNICORN GAMBIT, THE [GAMING MAGI 3]  BISCHOFF, DAVID F.  C/P,P
WRAITH BOARD [GAMING MAGI 2]  BISCHOFF, DAVID F.  C/P,P
LOST AND FOUND [@GAMMA QUEST #1]  COX, GREG  P
BROKEN COUNTRY, THE [@GAMMALAW 3]  DALEY, BRIAN  C/P,P
SCREAMING ACROSS THE SKY, A [@GAMMALAW 2]  DALEY, BRIAN  C/P,P
SMOKE ON THE WATER [@GAMMALAW 1]  DALEY, BRIAN  C/P,P
TO WATERS’ END [@GAMMALAW 4]  DALEY, BRIAN  C/P,P
BRONZE OF EDDARTA, THE [GANDALARA 3]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P,P
GANDALARA CYCLE I, THE [GANDALARA 1,2,3]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
GANDALARA CYCLE II, THE [GANDALARA 4,5,6]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
GLASS OF DYSKORNIS, THE [GANDALARA 2]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
RETURN TO EDDARTA [GANDALARA 6]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P,P
RIVER WALL, THE [GANDALARA 7]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
SEARCH FOR KA, THE [GANDALARA 5]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P,P
STEEL OF RAITHSKAR, THE [GANDALARA 1]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
WELL OF DARKNESS, THE [GANDALARA 4]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
CHAOS AND ORDER [GAP 4]  DONALDSON, STEPHEN R.  H:2,P
DARK AND HUNGRY GOD ARISES, A [GAP 3]  DONALDSON, STEPHEN R.  H:2,@P
FORBIDDEN KNOWLEDGE [GAP 2]  DONALDSON, STEPHEN R.  H:2,P
REAL STORY, THE [GAP 1]  DONALDSON, STEPHEN R.  H:2,P
THIS DAY ALL GODS DIE [GAP 5]  DONALDSON, STEPHEN R.  H:2,P

TITLE  AUTHOR  CODE
FALL OF THE FAMILIES, THE [GARDNER 2]  MANN, PHILLIP  H
MASTER OF PAXWAX [GARDNER 1]  MANN, PHILLIP  H
BLOOD HUNT [GARRETH MIKAELIAN 1]  KILLOUGH, LEE  C/P,SFWA-P,P
BLOODLINKS [GARRETH MIKAELIAN 2]  KILLOUGH, LEE  P,SFWA-P
BLOODWALK [GARRETH MIKAELIAN 1,2]  KILLOUGH, LEE  LP
ANGRY LEAD SKIES [GARRETT 10]  COOK, GLEN  C/P,P
BITTER GOLD HEARTS [GARRETT 2]  COOK, GLEN  C/P:2,P
COLD COPPER TEARS [GARRETT 3]  COOK, GLEN  C/P:2,P
DEADLY QUICKSILVER LIES [GARRETT 7]  COOK, GLEN  P,SFWA-P
DREAD BRASS SHADOWS [GARRETT 5]  COOK, GLEN  C/P,P
FADED STEEL HEAT [GARRETT 9]  COOK, GLEN  C/P,P
GARRETT FILES, THE [GARRETT 1,2,3]  COOK, GLEN  H
GARRETT INVESTIGATES [GARRETT 7,8,9]  COOK, GLEN  H
GARRETT ON THE CASE [GARRETT 10,11]  COOK, GLEN  H
GARRETT, P.I. [GARRETT 4,5,6]  COOK, GLEN  H
OLD TIN SORROWS [GARRETT 4]  COOK, GLEN  C/P,P
PETTY PEWTER GODS [GARRETT 8]  COOK, GLEN  C/P:2,P
RED IRON NIGHTS [GARRETT 6]  COOK, GLEN  P,SFWA-P:2
SWEET SILVER BLUES [GARRETT 1]  COOK, GLEN  C/P:2,P
WHISPERING NICKEL IDOLS [GARRETT 11]  COOK, GLEN  P
EVIDENCE [GARRISON 2]  JOHNSON, KARMELA  LP
LOCKDOWN [GARRISON 1]  SCHARTZ, VIJAYA  LP
FACES OF INFINITY [GATES OF TIME 2]  PARKINSON, DAN  P
PARADOX GATE [GATES OF TIME 3]  PARKINSON, DAN  P
WHISPERS, THE [GATES OF TIME 1]  PARKINSON, DAN  P
PATHS OF EVIL [@GAUNTLET DARK LEGACY 1]  WHITE, RICHARD C.  P
FORGE OF HEAVEN [GENE WARS 2]  CHERRYH, C. J.  H:2
HAMMERFALL [GENE WARS 1]  CHERRYH, C. J.  H:2,P
GENELLAN: PLANETFALL [@GENELLAN #1]  GIER, SCOTT G.  P
GENELLAN: IN THE SHADOW OF THE MOON [@GENELLAN #2]  GIER, SCOTT G.  P
GENELLAN: FIRST VICTORY [@GENELLAN #3]  GIER, SCOTT G.  P
ANVIL, THE [GENERAL 3]  STIRLING, S. M./DRAKE, DAVID  C/P:2,SFWA-P,P
CHOSEN, THE [GENERAL]  STIRLING, S. M./DRAKE, DAVID  C/P:2,P
CONQUEROR (PROOF) [GENERAL 3,4,5]  STIRLING, S. M./DRAKE, DAVID  SR-LP
REFORMER, THE [GENERAL]  STIRLING, S. M./DRAKE, DAVID  H,P
TYRANT, THE [GENERAL]  FLINT, ERIC/DRAKE, DAVID  P
FORGE, THE [@GENERAL #1]  STIRLING, S. M./DRAKE, DAVID  C/P:2,P
HAMMER, THE [@GENERAL #2]  STIRLING, S. M./DRAKE, DAVID  C/P:2,SFWA-P,P
STEEL, THE [@GENERAL #4]  STIRLING, S. M./DRAKE, DAVID  P,SFWA-P
SWORD, THE [@GENERAL #5]  STIRLING, S. M./DRAKE, DAVID  C/P:2,P
GENESIS QUEST, THE [GENESIS 1]  MOFFITT, DONALD  C/P:2,P
SECOND GENESIS [GENESIS 2]  MOFFITT, DONALD  C/P:2,P
GEODESICA ASCENT [GEODESICA 1]  WILLIAMS, SEAN/DIX, SHANE  P
GEODESICA DESCENT [GEODESICA 2]  WILLIAMS, SEAN/DIX, SHANE  P
SHAME OF MAN [GEODESSEY 2]  ANTHONY, PIERS  H,P
HOPE OF EARTH [GEODYSSEY 3]  ANTHONY, PIERS  H
ISLE OF WOMAN [GEODYSSEY 1]  ANTHONY, PIERS  H:2,P
MUSE OF ART [GEODYSSEY 4]  ANTHONY, PIERS  H
CHAOS KING, THE [GEORGIE AND BUG 2]  RUBY, LAURA  SR-LP
ALIENIST [GERALD KNAVE]  JANIFER, LAURENCE M.  LP
KNAVE AND THE GAME [GERALD KNAVE]  JANIFER, LAURENCE M.  H,SFWA
KNAVE IN HAND [GERALD KNAVE 2]  JANIFER, LAURENCE M.  C/P,P
SURVIVOR [GERALD KNAVE 1]  JANIFER, LAURENCE M.  C/P:2,P
TWO [GERALD KNAVE]  JANIFER, LAURENCE M.  LP
FOX AND EMPIRE [GERIN THE FOX 3]  TURTLEDOVE, HARRY  P
PRINCE OF THE NORTH [GERIN THE FOX 2]  TURTLEDOVE, HARRY  C/P,P
WERENIGHT [GERIN THE FOX 1]  TURTLEDOVE, HARRY  P,SFWA-P
WISDOM OF THE FOX [GERIN THE FOX (REVISED) 1,2]  TURTLEDOVE, HARRY  C/P,P
AND LOVING IT! [GET SMART 6]  JOHNSTON, WILLIAM  P
GET SMART! [GET SMART 1]  JOHNSTON, WILLIAM  P
GET SMART ONCE AGAIN! [GET SMART 3]  JOHNSTON, WILLIAM  C/P,P
MAX SMART AND THE GHASTLY GHOST AFFAIR [GET SMART 9]  JOHNSTON, WILLIAM  P
MAX SMART AND THE PERILOUS PELLETS [GET SMART 4]  JOHNSTON, WILLIAM  C/P,P
MAX SMART LOSES CONTROL [GET SMART 8]  JOHNSTON, WILLIAM  P
MAX SMART — THE SPY WHO WENT OUT TO THE COLD [GET SMART 7]  JOHNSTON, WILLIAM  P
MISSED IT BY THAT MUCH! [GET SMART 5]  JOHNSTON, WILLIAM  P
SORRY, CHIEF... [GET SMART 2]  JOHNSTON, WILLIAM  P
GETTYSBURG [GETTYSBURG 1]  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H
GRANT COMES EAST [GETTYSBURG 2]  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H
NEVER CALL RETREAT [GETTYSBURG 3]  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H
SUNDERLIES SEEKING [@GHATTENS’ GAMBIT #1]  GREENO, GAYLE  P
FARTHEST SEEKING, THE [@GHATTENS’ GAMBIT #2]  GREENO, GAYLE  C/P,P
FINDER’S SEEKERS [@GHATTI’S TALE #1]  GREENO, GAYLE  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 73
TITLE  AUTHOR  CODE
MIND-SPEAKERS’ CALL [@GHATTI’S TALE #2]  GREENO, GAYLE  C/P,SFWA-P,P
EXILES’ RETURN [@GHATTI’S TALE #3]  GREENO, GAYLE  C/P,P
GHOST HOUSE [GHOST HOUSE 1]  MCNALLY, CLAIRE  P
GHOST HOUSE REVENGE [GHOST HOUSE 2]  MCNALLY, CLAIRE  P
WHAT’S A GHOUL TO DO? [@GHOST HUNTER]  LAURIA, VICTORIA  P
AFTER THE LONG GOODBYE [GHOST IN THE SHELL 2]  YAMADA, MASAKI  H
GHOSTBUSTERS 2 [GHOSTBUSTERS]  NAHA, ED  P
GHOSTBUSTERS: THE RETURN [GHOSTBUSTERS]  FISCH, SHOLLY  P
ACCURSED [@GHOSTS OF ALBION]  BENSON, AMBER/GOLDEN, CHRISTOPHER  LP
INITIATION [@GHOSTS OF ALBION]  BENSON, AMBER/GOLDEN, CHRISTOPHER  SR-H
WITCHERY [@GHOSTS OF ALBION]  BENSON, AMBER/GOLDEN, CHRISTOPHER  LP
MIND GAME [GHOSTWALKERS 2]  FEEHAN, CHRISTINE  P
SHADOW GAME [GHOSTWALKERS 1]  FEEHAN, CHRISTINE  H
ENTOVERSE [GIANTS 4]  HOGAN, JAMES P.  H:2,P
GENTLE GIANTS OF GANYMEDE, THE [GIANTS 2]  HOGAN, JAMES P.  C/P:2,SR-P
GIANTS NOVELS, THE [GIANTS 1,2,3]  HOGAN, JAMES P.  C/P,P,H
GIANTS’ STAR [GIANTS 3]  HOGAN, JAMES P.  C/P:2,P
INHERIT THE STARS [GIANTS 1]  HOGAN, JAMES P.  C/P:2,SR-P
MINERVAN EXPERIMENT, THE [GIANTS 1,2,3]  HOGAN, JAMES P.  H:2
MISSION TO MINERVA [GIANTS 5]  HOGAN, JAMES P.  H
GILGAMESH THE KING [GILGAMESH]  SILVERBERG, ROBERT  C/P,P,H:2
TO THE LAND OF THE LIVING [GILGAMESH]  SILVERBERG, ROBERT  C/P,P
LIAR’S OATH [GIRD 2]  MOON, ELIZABETH  C/P:2,SFWA-P,P
SURRENDER NONE: THE LEGACY OF GIRD [GIRD 1]  MOON, ELIZABETH  C/P:2,SFWA-P,P
GIRL FACTORY, THE [GIRL FACTORY 1]  MURPHY, ROBERT FRANKLIN  C/P,P
KING’S MATE [GIRL FACTORY 2]  MURPHY, ROBERT FRANKLIN  P
MAN-MADE WOMAN, THE [GIRL FACTORY 3]  MURPHY, ROBERT FRANKLIN  P
BIRDS-OF-A-FEATHER AFFAIR, THE [@GIRL FROM UNCLE #1]  AVALLONE, MICHAEL  P
GLOBAL GLOBULES AFFAIR, THE (BRITISH) [@GIRL FROM UNCLE #1]  LATTER, SIMON  SR-P
BLAZING AFFAIR, THE [@GIRL FROM UNCLE #2]  AVALLONE, MICHAEL  P

GOLDEN BOATS OF TARADATA AFFAIR, THE (BRITISH)

[@GIRL FROM UNCLE #3] LATTER, SIMON

SR-P

AGATHA HETERODYNE AND THE BEETLEBURG CLANK [@GIRL GENIUS #1]  FOGLIO, PHIL/FOGLIO, KAJA  CX
AGATHA HETERODYNE AND THE AIRSHIP CITY [@GIRL GENIUS #2]  FOGLIO, PHIL/FOGLIO, KAJA  CX
AGATHA HETERODYNE AND THE MONSTER ENGINE [@GIRL GENIUS #3]  FOGLIO, PHIL/FOGLIO, KAJA  CX
AGATHA HETERODYNE AND THE CIRCUS OF DREAMS [@GIRL GENIUS #4]  FOGLIO, PHIL/FOGLIO, KAJA  CX

AGATHA HETERODYNE AND THE CLOCKWORK PRINCESS [@GIRL GENIUS #5] FOGLIO, PHIL/FOGLIO, KAJA

CX

GLASSWRIGHT’S APPRENTICE, THE [GLASSWRIGHT 1]  KLASKY, MINDY L.  C/P,P
GLASSWRIGHT’S JOURNEYMAN, THE [GLASSWRIGHT 3]  KLASKY, MINDY L.  P
GLASSWRIGHT’S PROGRESS, THE [GLASSWRIGHT 2]  KLASKY, MINDY L.  P
GLENRAVEN [GLENRAVEN 1]  BRADLEY, MARION ZIMMER/LISLE, HOLLY  C/P,P
IN THE RIFT [GLENRAVEN 2]  BRADLEY, MARION ZIMMER/LISLE, HOLLY  H:2,P
EYE OF THE EAGLE [GLOBAL 2000 1]  BASILE, GLORIA VITANZA  P
JACKAL HELIX, THE [GLOBAL 2000 2]  BASILE, GLORIA VITANZA  P
STING OF THE SCORPION, THE [GLOBAL 2000 3]  BASILE, GLORIA VITANZA  P
GNOME MAN’S LAND [GNOME MAN’S LAND 1]  FRIESNER, ESTHER M.  C/P:2,SFWA-P:2,P
HARPY HIGH [GNOME MAN’S LAND 2]  FRIESNER, ESTHER M.  C/P,P
UNICORN U. [GNOME MAN’S LAND 3]  FRIESNER, ESTHER M.  C/P,SFWA-P:2,P
GNOMEWRENCH IN THE DWARFWORKS, THE [GNOMEWRENCH 1]  O’DONOHOE, NICK  C/P,P
GNOMEWRENCH IN THE PEOPLEWORKS, THE [GNOMEWRENCH 2]  O’DONOHOE, NICK  P
GNOME’S ENGINE, THE [GOBLIN MOON 2]  EDGERTON, TERESA  C/P:2,SFWA-P:2,P
GOBLIN MOON [GOBLIN MOON 1]  EDGERTON, TERESA  C/P:2,SFWA-P,P
LABYRINTH OF DREAMS, THE [@G.O.D. INC. #1]  CHALKER, JACK L.  C/P:2,SFWA-P,P
SHADOW DANCERS, THE [@G.O.D. INC. #2]  CHALKER, JACK L.  C/P:2,SFWA-P,P
MAZE IN THE MIRROR, THE [@G.O.D. INC. #3]  CHALKER, JACK L.  C/P:2,SFWA-P,P
GODMOTHER, THE [GODMOTHER 1]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
GODMOTHER’S APPRENTICE, THE [GODMOTHER 2]  SCARBOROUGH, ELIZABETH ANN  C/P,P
GODMOTHER’S WEB, THE [GODMOTHER 3]  SCARBOROUGH, ELIZABETH ANN  H
ARIADNE’S WEB [GODS 2]  SABERHAGEN, FRED  P
ARMS OF HERCULES, THE [GODS 3]  SABERHAGEN, FRED  C/P,LP
BOOKS OF THE GODS, THE, PART ONE [GODS 1,2,3]  SABERHAGEN, FRED  H
FACE OF APOLLO, THE [GODS 1]  SABERHAGEN, FRED  C/P,LP
GOD OF THE GOLDEN FLEECE [GODS 4]  SABERHAGEN, FRED  H
GODS OF FIRE AND THUNDER [GODS 5]  SABERHAGEN, FRED  H
MOST ANCIENT SONG [GODS OF IRELAND 1]  FLYNN, CASEY  C/P,P
ENCHANTED ISLES, THE [@GODS OF IRELAND #2]  FLYNN, CASEY  P
HINTERLAND [GODSLAYER 2]  CLEMENS, JAMES  H
SHADOWFALL [GODSLAYER 1]  CLEMENS, JAMES  H:2
FORBIDDEN MAGIC [@GODWARS #1]  WELLS, ANGUS  P
DARK MAGIC [@GODWARS #2]  WELLS, ANGUS  C/P:2,P
WILD MAGIC [@GODWARS #3]  WELLS, ANGUS  C/P:2,P
GUARDIANS OF ALEXANDER [GOLDBANE 1]  WILSON, JOHN  LP

TITLE  AUTHOR  CODE
GOLDEN AGE, THE [GOLDEN AGE 1]  WRIGHT, JOHN C.  H:2
GOLDEN TRANSCENDENCE, THE [GOLDEN AGE 3]  WRIGHT, JOHN C.  H
PHOENIX EXULTANT, THE [GOLDEN AGE 2]  WRIGHT, JOHN C.  H
BEYOND THE GATE [GOLDEN QUEEN 2]  WOLVERTON, DAVE  H,SFWA:2,P
GOLDEN QUEEN, THE [GOLDEN QUEEN 1]  WOLVERTON, DAVE  P
LORDS OF THE SEVENTH SWARM [GOLDEN QUEEN 3]  WOLVERTON, DAVE  H,P
WORLDS OF THE GOLDEN QUEEN [GOLDEN QUEEN 1,2]  FARLAND, DAVID  LP
GLORY [GOLDENWING 1]  COPPEL, ALFRED  H
GLORY’S PEOPLE [GOLDENWING 3]  COPPEL, ALFRED  H,P
GLORY’S WAR [GOLDENWING 2]  COPPEL, ALFRED  P
DRAGON REFORGED [GONARD 2]  CYMRI, CHRYS  P
DRAGONS CAN ONLY RUST [GONARD 1]  CYMRI, CHRYS  P
BARBARIAN OF WORLD’S END, THE [GONDWANE 4]  CARTER, LIN  P
ENCHANTRESS OF WORLD’S END, THE [GONDWANE 2]  CARTER, LIN  C/P,P
GIANT OF WORLD’S END [GONDWANE]  CARTER, LIN  P
IMMORTAL OF WORLD’S END, THE [GONDWANE 3]  CARTER, LIN  C/P,P
PIRATE OF WORLD’S END, THE [GONDWANE 5]  CARTER, LIN  P
WARRIOR OF WORLD’S END, THE [GONDWANE 1]  CARTER, LIN  C/P,P
GONJI [GONJI]  RYPEL, T. C.  P
ATTACK OF THE MUTANT [GOOSEBUMPS]  STINE, R. L.  LP
STAY OUT OF THE BASEMENT [GOOSEBUMPS]  STINE, R. L.  LP
ASSASSIN OF GOR [GOR 5]  NORMAN, JOHN  C/P:2,P
BEASTS OF GOR [GOR 12]  NORMAN, JOHN  C/P:2,P
BLOOD BROTHERS OF GOR [GOR 18]  NORMAN, JOHN  C/P,P
CAPTIVE OF GOR [GOR 7]  NORMAN, JOHN  C/P:2,P
DANCER OF GOR [GOR 22]  NORMAN, JOHN  P
EXPLORERS OF GOR [GOR 13]  NORMAN, JOHN  C/P:2,P
FIGHTING SLAVE OF GOR [GOR 14]  NORMAN, JOHN  C/P:2,P
GUARDSMAN OF GOR [GOR 16]  NORMAN, JOHN  P
HUNTERS OF GOR [GOR 8]  NORMAN, JOHN  C/P:2,P
KAJIRA OF GOR [GOR 19]  NORMAN, JOHN  P
MAGICIANS OF GOR [GOR 25]  NORMAN, JOHN  C/P:2,P
MARAUDERS OF GOR [GOR 9]  NORMAN, JOHN  C/P:2,P
MERCENARIES OF GOR [GOR 21]  NORMAN, JOHN  C/P,P
NOMADS OF GOR [GOR 4]  NORMAN, JOHN  C/P:2,P
OUTLAW OF GOR [GOR 2]  NORMAN, JOHN  C/P:2,P
PLAYERS OF GOR [GOR 20]  NORMAN, JOHN  C/P:2,P
PRIEST-KINGS OF GOR [GOR 3]  NORMAN, JOHN  C/P:2,P
RAIDERS OF GOR [GOR 6]  NORMAN, JOHN  C/P:2,P
RENEGADES OF GOR [GOR 23]  NORMAN, JOHN  P
ROGUE OF GOR [GOR 15]  NORMAN, JOHN  C/P:2,P
SAVAGES OF GOR [GOR 17]  NORMAN, JOHN  C/P,P
SLAVE GIRL OF GOR [GOR 11]  NORMAN, JOHN  C/P,P
TARNSMAN OF GOR [GOR 1]  NORMAN, JOHN  C/P:2,P
TRIBESMEN OF GOR [GOR 10]  NORMAN, JOHN  C/P:2,P
VAGABONDS OF GOR [GOR 24]  NORMAN, JOHN  P
WITNESS OF GOR [GOR 26]  NORMAN, JOHN  H
CITY OF HAWKS [GORD THE ROGUE]  GYGAX, GARY  C/P:2,SFWA-P,P
COME ENDLESS DARKNESS [GORD THE ROGUE]  GYGAX, GARY  C/P:2,P
DANCE OF DEMONS [GORD THE ROGUE]  GYGAX, GARY  C/P,P
NIGHT ARRANT [GORD THE ROGUE]  GYGAX, GARY  C/P:2,P
SEA OF DEATH [GORD THE ROGUE]  GYGAX, GARY  C/P:2,P
GORMENGHAST [GORMENGHAST 2]  PEAKE, MERVYN  C/P:2,P,H
TITUS ALONE [GORMENGHAST 3]  PEAKE, MERVYN  C/P,P,H
TITUS GROAN [GORMENGHAST 1]  PEAKE, MERVYN  C/P,P,H
FENRIS DEVICE, THE [GRAINGER 5]  STABLEFORD, BRIAN M.  C/P,P
HALCYON DRIFT, THE [GRAINGER 1]  STABLEFORD, BRIAN M.  C/P:2,P
PARADISE GAME, THE [GRAINGER 4]  STABLEFORD, BRIAN M.  C/P,P
PROMISED LAND [GRAINGER 3]  STABLEFORD, BRIAN M.  H:2,P
RHAPSODY IN BLACK [GRAINGER 2]  STABLEFORD, BRIAN M.  C/P:2,P
SWAN SONG [GRAINGER 6]  STABLEFORD, BRIAN M.  P
SWAN SONGS [GRAINGER 1,2,3,4,5,6]  STABLEFORD, BRIAN M.  LP
PAST IMPERATIVE [@GREAT GAME #1]  DUNCAN, DAVE  H,P
PRESENT TENSE [@GREAT GAME #2]  DUNCAN, DAVE  H
FUTURE INDEFINITE [@GREAT GAME #3]  DUNCAN, DAVE  P
MAN WITHOUT A PLANET, THE [GREAT IMPERIUM]  CARTER, LIN  DG606
OUTWORLDER [GREAT IMPERIUM]  CARTER, LIN  C/P:2,P
STAR ROGUE [GREAT IMPERIUM]  CARTER, LIN  C/P:2,P
EYES OF THE CALCULOR [GREAT MACHINE 2]  MCMULLEN, SEAN  H
SOULS IN THE GREAT MACHINE [GREAT MACHINE 1]  MCMULLEN, SEAN  LP
CHILD OF THE DARK PROPHECY [@GREAT TREE OF AVALON 1]  BARRON, T. A.  P
SHADOWS ON THE STARS [GREAT TREE OF AVALON 2]  BARRON, T. A.  P
ETERNAL FLAME [@GREAT TREE OF AVALON #3]  BARRON, T. A.  H

The MITSFS Seriesdex  September 24, 2007  Page 75
TITLE  AUTHOR  CODE
CHILDREN OF GREEN KNOWE, THE [GREEN KNOWE 1]  BOSTON, L. M.  H,LP
ENEMY AT GREEN KNOWE, AN [GREEN KNOWE 5]  BOSTON, L. M.  H,LP
RIVER AT GREEN KNOWE, THE [GREEN KNOWE 3]  BOSTON, L. M.  H,LP
STRANGER AT GREEN KNOWE, A [GREEN KNOWE 4]  BOSTON, L. M.  H,LP
TREASURE OF GREEN KNOWE [GREEN KNOWE 2]  BOSTON, L. M.  H,LP
GREEN MILE, THE [GREEN MILE 1,2,3,4,5,6]  KING, STEPHEN  P
DOOM OF THE GREEN PLANET [GREEN PLANET 2]  PETAJA, EMIL  DH070
LORD OF THE GREEN PLANET [GREEN PLANET 1]  PETAJA, EMIL  DH022
FIRST RIDER’S CALL [GREEN RIDER 2]  BRITAIN, KRISTEN  H,P
GREEN RIDER [GREEN RIDER 1]  BRITAIN, KRISTEN  H:2,P
AND ALL BETWEEN [GREEN-SKY 2]  SNYDER, ZILPHA KEATLEY  C/P,SFWA-P,P
BELOW THE ROOT [GREEN-SKY 1]  SNYDER, ZILPHA KEATLEY  P,SFWA-P
UNTIL THE CELEBRATION [GREEN-SKY 3]  SNYDER, ZILPHA KEATLEY  P
AS THE GREEN STAR RISES [GREEN STAR 4]  CARTER, LIN  C/P,P
BY THE LIGHT OF THE GREEN STAR [GREEN STAR 3]  CARTER, LIN  C/P:2,P
IN THE GREEN STAR’S GLOW [GREEN STAR 5]  CARTER, LIN  P
UNDER THE GREEN STAR [GREEN STAR 1]  CARTER, LIN  C/P:2,P
WHEN THE GREEN STAR CALLS [GREEN STAR 2]  CARTER, LIN  C/P:2,P
MINDSTAR RISING [GREG MANDEL 1]  HAMILTON, PETER F.  C/P:2,P
NANO FLOWER, THE [GREG MANDEL 3]  HAMILTON, PETER F.  H
QUANTUM MURDER, A [GREG MANDEL 2]  HAMILTON, PETER F.  C/P,P,H
GREMLINS [GREMLINS]  GIPE, GEORGE  C/P,P
GREMLINS 2 [GREMLINS]  BISCHOFF, DAVID F.  P
AGAINST THE GIANTS [@GREYHAWK]  EMERSON, RU  P
DESCENT INTO THE DEPTHS OF THE EARTH [@GREYHAWK]  KIDD, PAUL  P
EYES HAVE IT, THE [GREYHAWK]  ESTES, ROSE  P
KEEP ON THE BORDERLANDS [@GREYHAWK]  EMERSON, RU  P
NIGHTWATCH [GREYHAWK]  BAILEY, ROBIN WAYNE  C/P,P
QUEEN OF THE DEMONWEB PITS [@GREYHAWK]  KIDD, PAUL  P
TEMPLE OF ELEMENTAL EVIL, THE [@GREYHAWK]  REID, THOMAS M.  P
TOMB OF HORRORS, THE [@GREYHAWK]  STROHM, KEITH FRANCIS  P
WHITE PLUME MOUNTAIN [@GREYHAWK]  KIDD, PAUL  P
SAGA OF OLD CITY [@GREYHAWK #1]  GYGAX, GARY  C/P:2,P
ARTIFACT OF EVIL [@GREYHAWK #2]  GYGAX, GARY  C/P:2,P
MASTER WOLF [@GREYHAWK #3]  ESTES, ROSE  C/P,P
PRICE OF POWER [@GREYHAWK #4]  ESTES, ROSE  C/P:2,P
DEMON HAND, THE [@GREYHAWK #5]  ESTES, ROSE  C/P:2,P
NAME OF THE GAME, THE [@GREYHAWK #6]  ESTES, ROSE  C/P:2,P
DOOM CITY [GREYSTONE BAY 2]  GRANT, CHARLES L.  PA,SFWA-P
GREYSTONE BAY [GREYSTONE BAY 1]  GRANT, CHARLES L.  SFWA-P
SEAHARP HOTEL, THE [GREYSTONE BAY 3]  GRANT, CHARLES L.  SFWA-P
ALTERNATE ORBITS [GRIMES]  CHANDLER, A. BERTRAM  D13783
ANARCH LORDS, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
BEYOND THE GALACTIC RIM [GRIMES]  CHANDLER, A. BERTRAM  DF237
BIG BLACK MARK, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
BRING BACK YESTERDAY [GRIMES]  CHANDLER, A. BERTRAM  DD517
BROKEN CYCLE, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
COMMODORE AT SEA, THE [GRIMES]/SPARTAN PLANET [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
CONTRABAND FROM OTHERSPACE [GRIMES]  CHANDLER, A. BERTRAM  DG609
DARK DIMENSIONS, THE [GRIMES]  CHANDLER, A. BERTRAM  D13783
DARK DIMENSIONS, THE [GRIMES]/RIM GODS, THE [GRIMES]  CHANDLER, A. BERTRAM  P
FAR TRAVELER, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
GATEWAY TO NEVER, THE [GRIMES]  CHANDLER, A. BERTRAM  D37062
HARD WAY UP, THE [GRIMES]  CHANDLER, A. BERTRAM  D31755
INHERITORS, THE [GRIMES]  CHANDLER, A. BERTRAM  D37062
INHERITORS, THE [GRIMES]/GATEWAY TO NEVER [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
INTO THE ALTERNATE UNIVERSE [GRIMES]  CHANDLER, A. BERTRAM  DM107

INTO THE ALTERNATE UNIVERSE [GRIMES]/ CONTRABAND FROM OTHERSPACE [GRIMES] CHANDLER, A. BERTRAM

C/P,P

JOHN GRIMES: LIEUTENANT OF THE SURVEY SERVICE [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: RESERVE COMMODORE [GRIMES]  CHANDLER, A. BERTRAM  H:2
JOHN GRIMES: RIM COMMANDER [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: RIM RUNNER [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: SURVEY CAPTAIN [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: TRAMP CAPTAIN [GRIMES]  CHANDLER, A. BERTRAM  H
LAST AMAZON, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
MATILDA’S STEPCHILDREN [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
RIM GODS, THE [GRIMES]  CHANDLER, A. BERTRAM  D72400
RIM OF SPACE, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,DF133,P,H
RIM OF SPACE, THE [GRIMES]/SHIP FROM OUTSIDE, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
ROAD TO THE RIM, THE [GRIMES]  CHANDLER, A. BERTRAM  DH029
ROAD TO THE RIM, THE [GRIMES]/HARD WAY UP, THE [GRIMES]  CHANDLER, A. BERTRAM  P
SHIP FROM OUTSIDE, THE [GRIMES]  CHANDLER, A. BERTRAM  DF237

TITLE  AUTHOR  CODE
SPARTAN PLANET [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
STAR COURIER [GRIMES]  CHANDLER, A. BERTRAM  P
STAR LOOT [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
TO KEEP THE SHIP [GRIMES]  CHANDLER, A. BERTRAM  P
TO PRIME THE PUMP [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
WAY BACK, THE [GRIMES]  CHANDLER, A. BERTRAM  P
WILD ONES, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
HARMONIES OF THE NET [GROUNDTIES 3]  FANCHER, JANE S.  C/P,P
DARK MOON, THE [@GUARDIAN #1]  GRAY, JULIA  P
JASPER FOREST, THE [@GUARDIAN #2]  GRAY, JULIA  P
CRYSTAL DESERT, THE [@GUARDIAN #3]  GRAY, JULIA  P
LIGHT OF EIDON [GUARDIAN-KING 1]  HANCOCK, KAREN  LP
RETURN OF THE GUARDIAN-KING [GUARDIAN-KING 4]  HANCOCK, KAREN  LP
SHADOW OVER KIRIATH [GUARDIAN-KING 3]  HANCOCK, KAREN  LP
SHADOW WITHIN, THE [GUARDIAN-KING 2]  HANCOCK, KAREN  LP
BRUTE FORCE [GUARDIANS 7]  AUSTIN, RICHARD  P
DESOLATION ROAD [GUARDIANS 8]  AUSTIN, RICHARD  P
GUARDIANS, THE [GUARDIANS 1]  AUSTIN, RICHARD  P
KRILOV CONTINUUM, THE [GUARDIANS 1]  LOVEGROVE, J. M. H.  LP
NIGHT OF THE PHOENIX [GUARDIANS 4]  AUSTIN, RICHARD  P
PLAGUE YEARS [GUARDIANS 12]  AUSTIN, RICHARD  P
THUNDER OF HELL [GUARDIANS 3]  AUSTIN, RICHARD  P
TRIAL BY FIRE [GUARDIANS 2]  AUSTIN, RICHARD  P
WAR ZONE [GUARDIANS 6]  AUSTIN, RICHARD  P
CURSE OF RATHLAW, THE [GUARDIANS (SAXON) 0]  SAXON, PETER  C/P,P
HAUNTING OF ALAN MAIS, THE [GUARDIANS (SAXON) 3]  SAXON, PETER  C/P,P
KILLING BONE, THE [GUARDIANS (SAXON) 1]  SAXON, PETER  P
THROUGH THE DARK CURTAIN [GUARDIANS (SAXON) 4]  SAXON, PETER  P
VAMPIRES OF FINISTERE, THE [GUARDIANS (SAXON) 4]  SAXON, PETER  P
GIVER OF ROSES [@GUARDIANS OF GADIEL]  MORGAN, KATHLEEN  LP
HEROES, THE [GUARDIANS OF THE FLAME 4,5]  ROSENBERG, JOEL  H:2
ROAD HOME, THE [GUARDIANS OF THE FLAME]  ROSENBERG, JOEL  P
ROAD TO EHVENOR, THE [GUARDIANS OF THE FLAME 6]  ROSENBERG, JOEL  H,P
WARRIOR LIVES, THE [GUARDIANS OF THE FLAME 5]  ROSENBERG, JOEL  C/P,P,H
WARRIORS, THE [GUARDIANS OF THE FLAME 1,2,3]  ROSENBERG, JOEL  H:2
SLEEPING DRAGON, THE [@GUARDIANS OF THE FLAME #1]  ROSENBERG, JOEL  C/P:2,P
SWORD AND THE CHAIN, THE [@GUARDIANS OF THE FLAME #2]  ROSENBERG, JOEL  C/P:2,P
SILVER CROWN, THE [@GUARDIANS OF THE FLAME #3]  ROSENBERG, JOEL  C/P:2,P
HEIR APPARENT, THE [@GUARDIANS OF THE FLAME #4]  ROSENBERG, JOEL  C/P,P

NOT EXACTLY THE THREE MUSKETEERS [GUARDIANS OF THE FLAME CASTLE CULLINANE 1] ROSENBERG, JOEL

H

NOT QUITE SCARAMOUCHE [GUARDIANS OF THE FLAME CASTLE CULLINANE 2] ROSENBERG, JOEL

H

NOT REALLY THE PRISONER OF ZENDA [GUARDIANS OF THE FLAME CASTLE CULLINANE 3] ROSENBERG, JOEL

H

DEFENDERS OF AR [GUARDIANS OF THE THREE 4]  LOVEJOY, JACK  C/P:2,P
KEEPER OF THE CITY [@GUARDIANS OF THE THREE 2]  MORWOOD, PETER/DUANE, DIANE  C/P,P
LORD OF CRAGSCLAW [GUARDIANS OF THE THREE 1]  FAWCETT, BILL/RANDALL, NEIL  C/P:2,P
WIZARD OF TIZARE [GUARDIANS OF THE THREE 3]  COSTELLO, MATTHEW J.  C/P:2,P
CHESSBOARD QUEEN, THE [GUINEVERE 2]  NEWMAN, SHARAN  LP
GUINEVERE [GUINEVERE 1]  NEWMAN, SHARAN  P
GUINEVERE EVERMORE [GUINEVERE 3]  NEWMAN, SHARAN  H
AWAKENING [@GUNDAM MOBILE SUIT #1]  TOMINO, YOSHIYUKI  C/P,P
ESCALATION [@GUNDAM MOBILE SUIT #2]  TOMINO, YOSHIYUKI  C/P:2,P
CONFRONTATION [@GUNDAM MOBILE SUIT #3]  TOMINO, YOSHIYUKI  C/P,P
HAWK OF MAY [GWALCHMAI 1]  BRADSHAW, GILLIAN  H:2,P
HAWK OF MAY (PROOF) [GWALCHMAI 1]  BRADSHAW, GILLIAN  SR-LP
KINGDOM OF SUMMER [GWALCHMAI 2]  BRADSHAW, GILLIAN  H:2,P
EYE OF THE VULTURE [GYPSY 2/WEIRD HEROES 7]  GOULART, RON  P
QUEST OF THE GYPSY [GYPSY 1]  GOULART, RON  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 77
TITLE  AUTHOR  CODE
THORN OGRES OF HAGWOOD [HAGWOOD 1]  JARVIS, ROBIN  H
CITY OF ILLUSIONS [HAINISH 3]  LE GUIN, URSULA K.  C/P:2,SFWA,P
PLANET OF EXILE [HAINISH 2]  LE GUIN, URSULA K.  C/P:2,P,SFWA,DG597
ROCANNON’S WORLD [HAINISH 1]  LE GUIN, URSULA K.  C/P:2,SFWA,DG574,P
TELLING, THE [HAINISH]  LE GUIN, URSULA K.  H:2
TELLING, THE (PROOF) [HAINISH]  LE GUIN, URSULA K.  SR-LP
THREE HAINISH NOVELS [HAINISH 1,2,3]  LE GUIN, URSULA K.  H:2
WORLDS OF EXILE AND ILLUSION [HAINISH 1,2,3]  LE GUIN, URSULA K.  LP
HAL SPACEJOCK [@HAL SPACEJOCK 1]  HAYNES, SIMON  P
HAL SPACEJOCK - SECOND COURSE [@HAL SPACEJOCK 2]  HAYNES, SIMON  P
DRAGONFRIGATE WIZARD HALCYON BLITHE [HALCYON BLITHE 2]  WARD, JAMES M.  H
ELVENBANE, THE [HALFBLOOD 1]  NORTON, ANDRE/LACKEY, MERCEDES  H:2,P
ELVENBLOOD [HALFBLOOD 2]  NORTON, ANDRE/LACKEY, MERCEDES  H:2
ELVENBORN [HALFBLOOD 3]  NORTON, ANDRE/LACKEY, MERCEDES  H:2
DEAD EASY [HALFLIFE 4]  SIMMONS, WM. MARK  H
DEAD ON MY FEET [HALFLIFE 2]  SIMMONS, WM. MARK  H
ONE FOOT IN THE GRAVE [HALFLIFE 1]  SIMMONS, WM. MARK  C/P:2,P
HALLOWED ISLE, THE [HALLOWED ISLE 1,2]  PAXSON, DIANA L.  P
BOOK OF THE SWORD, THE [@HALLOWED ISLE #1]  PAXSON, DIANA L.  LP
BOOK OF THE SPEAR, THE [@HALLOWED ISLE #2]  PAXSON, DIANA L.  LP
BOOK OF THE CAULDRON, THE [@HALLOWED ISLE #3]  PAXSON, DIANA L.  LP
BOOK OF THE STONE, THE [@HALLOWED ISLE #4]  PAXSON, DIANA L.  H,LP
GHOSTS OF ONYX [@HALO]  NYLUND, ERIC S.  P
HALO: FIRST STRIKE [@HALO]  NYLUND, ERIC S.  P
HALO: THE FALL OF REACH [@HALO]  NYLUND, ERIC S.  P
HALO: THE FLOOD [@HALO]  DIETZ, WILLIAM C.  P
HAMMER AND THE CROSS, THE [HAMMER AND THE CROSS 1]  HARRISON, HARRY  C/P:2,P
KING AND EMPEROR [HAMMER AND THE CROSS 3]  HARRISON, HARRY/HOLM, JOHN  C/P,P,H
ONE KING’S WAY [HAMMER AND THE CROSS 2]  HARRISON, HARRY/HOLM, JOHN  C/P,P
AT ANY PRICE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
BUTCHER’S BILL, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
CAUGHT IN THE CROSSFIRE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,P

COMPLETE HAMMER’S SLAMMERS, THE, VOLUME 1 [HAMMER’S SLAMMERS] DRAKE, DAVID

H

COMPLETE HAMMER’S SLAMMERS, THE, VOLUME 2 [HAMMER’S SLAMMERS] DRAKE, DAVID

H

COUNTING THE COST [HAMMER’S SLAMMERS]  DRAKE, DAVID  P,SFWA-P
HAMMER’S SLAMMERS [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
PAYING THE PIPER [HAMMER’S SLAMMERS]  DRAKE, DAVID  P
ROLLING HOT [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,SFWA-P:2,P
SHARP END, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,SFWA,P,H
SLAMMERS DOWN! [HAMMER’S SLAMMERS]  JOHNSON, TODD  P,SFWA-P
TANK LORDS, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,P
VOYAGE, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,P,H
WARRIOR, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
AT THE EDGE OF SPACE [HANAN REBELLION 1,2]  CHERRYH, C. J.  P
BROTHERS OF EARTH [HANAN REBELLION 1]  CHERRYH, C. J.  H:2,P
HUNTER OF WORLDS [HANAN REBELLION 2]  CHERRYH, C. J.  C/P,P,H
CUTTING EDGE, THE [HANDFUL OF MEN 1]  DUNCAN, DAVE  C/P:2,P
LIVING GOD, THE [HANDFUL OF MEN 4]  DUNCAN, DAVE  C/P,P,H
STRICKEN FIELD, THE [HANDFUL OF MEN 3]  DUNCAN, DAVE  C/P,P,H
UPLAND OUTLAWS [HANDFUL OF MEN 2]  DUNCAN, DAVE  C/P,P,H
CLOUD GATHERER, THE [HARAN OF OAS 1]  TAFF, JOHN F. D.  LP
HEIR TO THE SAND [HARAN OF OAS 2]  TAFF, JOHN F. D.  LP
BLADE OF TYSHALE [HARI MICHAELSON 2]  STOVER, MATTHEW WOODRING  P
HEROES DIE [HARI MICHAELSON 1]  STOVER, MATTHEW WOODRING  H,LP
BECOMING HUMAN [HARMONY OF WORLDS]  FREIREICH, VALERIE J.  P
TESTAMENT [HARMONY OF WORLDS]  FREIREICH, VALERIE J.  C/P,P
OF WAR AND HONOR [@HARRIERS #1]  DICKSON, GORDON R.  PA,C/PA,SFWA-P
BLOOD AND WAR [@HARRIERS #2]  DICKSON, GORDON R.  C/PA,PA
BLAZE OF WRATH [HARRY BORG 3]  HAWKINS, WARD  C/P,P
RED FLAME BURNING [HARRY BORG 1]  HAWKINS, WARD  C/P,P
SWORD OF FIRE [HARRY BORG 2]  HAWKINS, WARD  P
TORCH OF FEAR [HARRY BORG 4]  HAWKINS, WARD  P

HARRY POTTER AND THE CHAMBER OF SECRETS (AMERICAN)

[HARRY POTTER 2] ROWLING, J. K.

H:2

HARRY POTTER AND THE DEATHLY HALLOWS [HARRY POTTER 7]  ROWLING, J. K.  H

HARRY POTTER AND THE GOBLET OF FIRE (AMERICAN)

[HARRY POTTER 4] ROWLING, J. K.

H:2,LP

HARRY POTTER AND THE GOBLET OF FIRE (BRITISH) [HARRY POTTER 4]  ROWLING, J. K.  H
HARRY POTTER AND THE HALF-BLOOD PRINCE [HARRY POTTER 6]  ROWLING, J. K.  H:2

HARRY POTTER AND THE ORDER OF THE PHOENIX (AMERICAN)

[HARRY POTTER 5] ROWLING, J. K.

H

The MITSFS Seriesdex  September 24, 2007  Page 78
TITLE  AUTHOR  CODE

HARRY POTTER AND THE PHILOSOPHER’S STONE (BRITISH)

[HARRY POTTER 1] ROWLING, J. K.

LP

HARRY POTTER AND THE PRISONER OF AZKABAN (AMERICAN)

[HARRY POTTER 3] ROWLING, J. K.

H

HARRY POTTER AND THE SORCERER’S STONE (AMERICAN)

[HARRY POTTER 1] ROWLING, J. K.

C/P,LP,H

HARRY POTTER UND DER STEIN DER WEISEN [HARRY POTTER 1]  ROWLING, J. K.  FH
FLEET OF STARS, THE [HARVEST OF STARS 4]  ANDERSON, POUL  C/P:2,P
HARVEST OF STARS [HARVEST OF STARS 2]  ANDERSON, POUL  H:2,P
HARVEST THE FIRE [HARVEST OF STARS 1]  ANDERSON, POUL  H:2,P
STARS ARE ALSO FIRE, THE [HARVEST OF STARS 3]  ANDERSON, POUL  H,P
STARKADDER [HATHER LAMBISSON 1]  KING, BERNARD  H
VARGR-MOON [HATHER LAMBISSON 2]  KING, BERNARD  H
I, THE SUN [HATTI EMPIRE]  MORRIS, JANET  P
BONES OF HAVEN, THE [HAWK & FISHER 6]  GREEN, SIMON R.  P
GOD KILLER, THE [HAWK & FISHER 3]  GREEN, SIMON R.  C/P:2,P
GUARD AGAINST DISHONOR [HAWK & FISHER 5]  GREEN, SIMON R.  C/P:2,P
HAWK & FISHER [HAWK & FISHER 1]  GREEN, SIMON R.  P
SWORDS OF HAVEN [@HAWK & FISHER 1,2,3]  GREEN, SIMON R.  P
WINNER TAKES ALL [HAWK & FISHER 2]  GREEN, SIMON R.  P
WOLF IN THE FOLD [HAWK & FISHER 4]  GREEN, SIMON R.  C/P:2,P
HAWK ETERNAL, THE [HAWK QUEEN 2]  GEMMELL, DAVID  P
IRONHAND’S DAUGHTER [HAWK QUEEN]  GEMMELL, DAVID  P
HEAVEN CHRONICLES [HEAVEN 1,2]  VINGE, JOAN D.  C/P,P,H
LEGACY [HEAVEN 1]  VINGE, JOAN D.  BS4
OUTCASTS OF HEAVEN BELT, THE [HEAVEN 2]  VINGE, JOAN D.  C/P:2,P
ANNALS OF THE HEECHEE, THE [HEECHEE 4]  POHL, FREDERIK  H:2,P,SFWA
BEYOND THE BLUE EVENT HORIZON [HEECHEE 2]  POHL, FREDERIK  H:2,P,SFWA
BOY WHO WOULD LIVE FOREVER, THE [HEECHEE 6]  POHL, FREDERIK  H
GATEWAY [HEECHEE 1]  POHL, FREDERIK  C/P,P,H
GATEWAY TRIP, THE [HEECHEE 5]  POHL, FREDERIK  C/P,P,H
HEECHEE RENDEZVOUS [HEECHEE 3]  POHL, FREDERIK  C/P,P,H
BYZANTIUM’S CROWN [HEIRS TO BYZANTIUM 1]  SHWARTZ, SUSAN  C/P,P
WOMAN OF FLOWERS, THE [HEIRS TO BYZANTIUM 2]  SHWARTZ, SUSAN  C/P,P
HALL OF WHISPERS [HEIRS TO GNARLSMYRE]  JEFFERIES, MIKE  P
DEATH RIDERS OF HEL [HEL 2]  DRAKE, ASA  P
WARRIOR WITCH OF HEL [HEL 1]  DRAKE, ASA  C/P,P
WEREBEASTS OF HEL [HEL 3]  DRAKE, ASA  P
DARKTRADERS [HELLFLOWER 2]  BES SHAHAR, ELUKI  C/P:2,P
HELLFLOWER [HELLFLOWER 1]  BES SHAHAR, ELUKI  C/P:2,SFWA-P:2,P
ARCHANGEL BLUES [@HELLFLOWER #3]  BES SHAHAR, ELUKI  P
EXODUS [@HELLGATE LONDON #1]  ODOM, MEL  P
HELLICONIA SPRING [HELLICONIA 1]  ALDISS, BRIAN W.  H:2,P,SFWA:2
HELLICONIA SUMMER [HELLICONIA 2]  ALDISS, BRIAN W.  H:2,P,SFWA
HELLICONIA WINTER [HELLICONIA 3]  ALDISS, BRIAN W.  C/P,SFWA-P,H:2,P,SFWA
DEFENDERS, THE [HELMSMAN 5]  BALDWIN, BILL  C/P:2,P
DEFIANCE, THE [HELMSMAN 7]  BALDWIN, BILL  P
GALACTIC CONVOY [HELMSMAN 2]  BALDWIN, BILL  C/P:2,P
HELMSMAN, THE [HELMSMAN 1]  BALDWIN, BILL  C/P:2,P
MERCENARIES, THE [HELMSMAN 4]  BALDWIN, BILL  C/P:2,P
SIEGE, THE [HELMSMAN 6]  BALDWIN, BILL  C/P,P
TROPHY, THE [HELMSMAN 3]  BALDWIN, BILL  C/P:2,P
INTO THE FIRE [HEL’S CRUCIBLE 2/MITHGAR 4]  MCKIERNAN, DENNIS L.  P
INTO THE FORGE [HEL’S CRUCIBLE 1/MITHGAR 3]  MCKIERNAN, DENNIS L.  P
BRETTA MARTYN [HENRY MARTYN 2]  SMITH, L. NEIL  H,P
HENRY MARTYN [HENRY MARTYN 1]  SMITH, L. NEIL  C/P,P

BEOWULF’S CHILDREN [HEOROT 2] NIVEN, LARRY/POURNELLE, JERRY/

BARNES, STEVEN

H:2,P

LEGACY OF HEOROT, THE [HEOROT 1] NIVEN, LARRY/POURNELLE, JERRY/

BARNES, STEVEN

H:2,P

BLOWN [HERALD CHILDE 2]  FARMER, PHILIP JOSE  SR-L
IMAGE OF THE BEAST, THE [HERALD CHILDE 1]  FARMER, PHILIP JOSE  C/P,SR-L
IMAGE OF THE BEAST [HERALD CHILDE 1,2]/BLOWN [HERALD CHILDE 1,2] FARMER, PHILIP JOSE  C/P:2,P
ARROW’S FALL [HERALDS OF VALDEMAR 3]  LACKEY, MERCEDES  C/P:2,P
ARROW’S FLIGHT [HERALDS OF VALDEMAR 2]  LACKEY, MERCEDES  C/P:2,P
ARROWS OF THE QUEEN [HERALDS OF VALDEMAR 1]  LACKEY, MERCEDES  C/P:2,P
QUEEN’S OWN [HERALDS OF VALDEMAR 1,2,3]  LACKEY, MERCEDES  H
VENGEANCE OF HERA, THE [@HERCULES]  BETANCOURT, JOHN GREGORY  P
WRATH OF POSEIDON, THE [@HERCULES]  BETANCOURT, JOHN GREGORY  P
EYE OF THE RAM, THE [@HERCULES THE LEGENDARY JOURNEYS]  BOGGS, TIMOTHY  C/P,P
FIRST CASUALTY, THE [@HERCULES THE LEGENDARY JOURNEYS]  SEIDMAN, DAVID L.  P
DEMON BLUES [HERE BE DEMONS 2]  FRIESNER, ESTHER M.  P
HERE BE DEMONS [HERE BE DEMONS 1]  FRIESNER, ESTHER M.  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 79
TITLE  AUTHOR  CODE
HOORAY FOR HELLYWOOD [HERE BE DEMONS 3]  FRIESNER, ESTHER M.  P,SFWA-P:2
AGAINST THE ODDS [HERIS SERRANO 7]  MOON, ELIZABETH  H,P
CHANGE OF COMMAND [HERIS SERRANO 6]  MOON, ELIZABETH  H:2,P
HERIS SERRANO [HERIS SERRANO 1,2,3]  MOON, ELIZABETH  H
HUNTING PARTY [HERIS SERRANO 1]  MOON, ELIZABETH  C/P:2,SFWA-P,P
ONCE A HERO [HERIS SERRANO 4]  MOON, ELIZABETH  C/P:2,P
RULES OF ENGAGEMENT [HERIS SERRANO 5]  MOON, ELIZABETH  C/P,P,H
SPORTING CHANCE [HERIS SERRANO 2]  MOON, ELIZABETH  C/P:2,P
WINNING COLORS [HERIS SERRANO 3]  MOON, ELIZABETH  C/P:2,SFWA-P,P
SEMPER MARS [@HERITAGE #1]  DOUGLAS, IAN  C/P,P
LUNA MARINE [@HERITAGE #2]  DOUGLAS, IAN  C/P,P
EUROPA STRIKE [@HERITAGE #3]  DOUGLAS, IAN  C/P,P
CONVERGENCE [HERITAGE UNIVERSE 4]  SHEFFIELD, CHARLES  H,P
DIVERGENCE [HERITAGE UNIVERSE 2]  SHEFFIELD, CHARLES  H,P
HERITAGE UNIVERSE, THE [HERITAGE UNIVERSE 1,2,3]  SHEFFIELD, CHARLES  H:2
RESURGENCE [HERITAGE UNIVERSE 5]  SHEFFIELD, CHARLES  H:2
SUMMERTIDE [HERITAGE UNIVERSE 1]  SHEFFIELD, CHARLES  C/P,P,H
TRANSCENDENCE [HERITAGE UNIVERSE 3]  SHEFFIELD, CHARLES  H:2,P
CONVERGENT SERIES [HERITAGE UNIVERSE (REVISED) 1,2]  SHEFFIELD, CHARLES  P
TRANSVERGENCE [HERITAGE UNIVERSE (REVISED) 3,4]  SHEFFIELD, CHARLES  C/P,P
ANGELS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,SFWA-P
CRUSADERS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA
EXPLORERS IN HELL [HEROES IN HELL]  DRAKE, DAVID/MORRIS, JANET  P,SFWA-P
GATES OF HELL, THE [HEROES IN HELL]  CHERRYH, C. J./MORRIS, JANET  C/P,P,H
HEROES IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,C/PA
KINGS IN HELL [HEROES IN HELL]  CHERRYH, C. J./MORRIS, JANET  C/P,P
LEGIONS OF HELL [HEROES IN HELL]  CHERRYH, C. J.  C/P,P
LITTLE HELLIAD, THE [HEROES IN HELL]  MORRIS, JANET/MORRIS, CHRIS  C/P,SFWA-P,P
MASTERS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA
PROPHETS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,C/PA,SFWA-P
REBELS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,C/PA
WAR IN HELL [HEROES IN HELL]  MORRIS, JANET  C/PA,PA,SFWA-P
GAME’S END [HIDDEN REALMS 3]  GREEN, SHARON  P
HIDDEN REALMS, THE [HIDDEN REALMS 1]  GREEN, SHARON  SFWA-P
JARKA RUUS [HIGH DRUID OF SHANNARA 1]  BROOKS, TERRY  H
JARKA RUUS (PROOF) [HIGH DRUID OF SHANNARA 1]  BROOKS, TERRY  SR-LP
STRAKEN [HIGH DRUID OF SHANNARA 3]  BROOKS, TERRY  H
TANEQUIL [HIGH DRUID OF SHANNARA 2]  BROOKS, TERRY  H
KINGS OF THE HIGH FRONTIER [HIGH PILGRIMAGE 1]  KOMAN, VICTOR  H
CAPTIVE SOUL, THE [@HIGHLANDER]  SHERMAN, JOSEPHA  P
ELEMENT OF FIRE, THE [HIGHLANDER]  HENDERSON, JASON  P
EVENING AT JOE’S, AN [HIGHLANDER]  HORVATH, GILLIAN  LPA
MEASURE OF A MAN [@HIGHLANDER]  HOLDER, NANCY  P
PATH, THE [@HIGHLANDER]  NEASON, REBECCA  P
SCIMITAR [HIGHLANDER]  MCCONNELL, ASHLEY  P
SCOTLAND THE BRAVE [@HIGHLANDER]  ROBERSON, JENNIFER  P
SHADOW OF OBSESSION [@HIGHLANDER]  NEASON, REBECCA  P
WHITE SILENCE [HIGHLANDER]  BUCHANAN, GINJER  P
ZEALOT [@HIGHLANDER]  LETTOW, DONNA  P
PASSAGE OF STARS, A [@HIGHROAD #1]  RASMUSSEN, ALIS A.  C/P:2,P
REVOLUTION’S SHORE [@HIGHROAD #2]  RASMUSSEN, ALIS A.  P
PRICE OF RANSOM, THE [@HIGHROAD #3]  RASMUSSEN, ALIS A.  C/P:2,P
AMBER SPYGLASS, THE [HIS DARK MATERIALS 3]  PULLMAN, PHILIP  C/P,P,H

AMBER SPYGLASS, THE (LIMITED EDITION) (CD)

[HIS DARK MATERIALS 1,2,3] PULLMAN, PHILIP

SR-MM

GOLDEN COMPASS, THE [HIS DARK MATERIALS 1]  PULLMAN, PHILIP  C/P:2,P,H
SUBTLE KNIFE, THE [HIS DARK MATERIALS 2]  PULLMAN, PHILIP  C/P,P,H
EARTH WILL SHAKE, THE [@HISTORICAL ILLUMINATUS #1]  WILSON, ROBERT ANTON  C/P:2,LP,H
WIDOW’S SON, THE [@HISTORICAL ILLUMINATUS #2]  WILSON, ROBERT ANTON  C/P,P
NATURE’S GOD [@HISTORICAL ILLUMINATUS #3]  WILSON, ROBERT ANTON  P
OUTLINE OF HISTORY, THE: VOLUME 1 [HISTORY 1]  WELLS, H. G.  [C/RAN-H]
OUTLINE OF HISTORY, THE: VOLUME 2 [HISTORY 2]  WELLS, H. G.  [C/RAN-H]
BOOK OF LOST TALES 1, THE [HISTORY OF MIDDLE-EARTH 1]  TOLKIEN, J. R. R.  H:2,LP
BOOK OF LOST TALES 2, THE [HISTORY OF MIDDLE-EARTH 2]  TOLKIEN, J. R. R.  H:2,P
LAYS OF BELERIAND, THE [HISTORY OF MIDDLE-EARTH 3]  TOLKIEN, J. R. R.  H:2,P
LOST ROAD AND OTHER WRITINGS, THE [HISTORY OF MIDDLE-EARTH 5]  TOLKIEN, J. R. R.  H:2,P
MORGOTH’S RING [HISTORY OF MIDDLE-EARTH 10]  TOLKIEN, J. R. R.  H
PEOPLES OF MIDDLE-EARTH, THE [HISTORY OF MIDDLE-EARTH 12]  TOLKIEN, J. R. R.  H
RETURN OF THE SHADOW, THE [HISTORY OF MIDDLE-EARTH 6]  TOLKIEN, J. R. R.  H
SAURON DEFEATED [HISTORY OF MIDDLE-EARTH 9]  TOLKIEN, J. R. R.  H
SHAPING OF MIDDLE-EARTH, THE [HISTORY OF MIDDLE-EARTH 4]  TOLKIEN, J. R. R.  H,P
TREASON OF ISENGARD [HISTORY OF MIDDLE-EARTH 7]  TOLKIEN, J. R. R.  H
WAR OF THE JEWELS, THE [HISTORY OF MIDDLE-EARTH 11]  TOLKIEN, J. R. R.  H:2

The MITSFS Seriesdex  September 24, 2007  Page 80
TITLE  AUTHOR  CODE
WAR OF THE RING, THE [HISTORY OF MIDDLE-EARTH 8]  TOLKIEN, J. R. R.  H:2
DON’T PANIC [HITCHHIKERS]  GAIMAN, NEIL  R/REF-P
HITCHHIKERS’ GUIDE TO THE GALAXY, THE [HITCHHIKERS 1]  ADAMS, DOUGLAS  C/P,H:2,P

HITCHHIKERS’ GUIDE TO THE GALAXY, THE (5/6 CASSETTES)

[HITCHHIKERS 1] ADAMS, DOUGLAS

SR-MM

HITCHHIKERS’ GUIDE TO THE GALAXY, THE (6 CASSETTES)

[HITCHHIKERS 1] ADAMS, DOUGLAS

SR-MM:2

HITCHHIKERS’ GUIDE TO THE GALAXY, THE (WITH CASABLANCA)

[HITCHHIKERS] ADAMS, DOUGLAS

SR-MM

ILLUSTRATED HITCHHIKER’S GUIDE TO THE GALAXY, THE [HITCHHIKERS] ADAMS, DOUGLAS  A
LIFE, THE UNIVERSE AND EVERYTHING [HITCHHIKERS 3]  ADAMS, DOUGLAS  C/P,P,H:2
MOSTLY HARMLESS [HITCHHIKERS 5]  ADAMS, DOUGLAS  H:2
ORIGINAL HITCHHIKER RADIO SCRIPTS, THE [HITCHHIKERS]  ADAMS, DOUGLAS  H,LP
RESTAURANT AT THE END OF THE UNIVERSE, THE [HITCHHIKERS 2]  ADAMS, DOUGLAS  C/P,P,H

RESTAURANT AT THE END OF THE UNIVERSE, THE (1 CASSETTE)

[HITCHHIKERS 2] ADAMS, DOUGLAS

SR-MM

SCIENCE OF THE HITCHHIKER’S GUIDE TO THE GALAXY, THE [HITCHHIKERS] HANLON, MICHAEL

R/REF-H

SO LONG, AND THANKS FOR ALL THE FISH [HITCHHIKERS 4]  ADAMS, DOUGLAS  C/P,P,H
ULTIMATE HITCHHIKER’S GUIDE, THE [HITCHHIKERS 1,2,3,4,5]  ADAMS, DOUGLAS  H,LP
FALL OF THE WHITE SHIP AVATAR [HOBART FLOYT]  DALEY, BRIAN  C/P:2,P
JINX ON A TERRAN INHERITANCE [HOBART FLOYT 2]  DALEY, BRIAN  C/P:2,P
REQUIEM FOR A RULER OF WORLDS [HOBART FLOYT 1]  DALEY, BRIAN  C/P:2,P
HOKA! [HOKA 1]  ANDERSON, POUL/DICKSON, GORDON R.  C/P,LP,H
HOKA! HOKA! HOKA! [HOKA 2]  ANDERSON, POUL/DICKSON, GORDON R.  P
HOKAS POKAS! [HOKA 3]  ANDERSON, POUL/DICKSON, GORDON R.  P
SOUND AND THE FURRY, THE [HOKA 2,3]  ANDERSON, POUL/DICKSON, GORDON R.  H
LEGEND OF HOLDERLY HALL, THE [HOLDERLY HALL 1]  CAMERON, KATE  P
MUSIC FROM THE PAST [HOLDERLY HALL 5]  CAMERON, KATE  P
PORTRAITS OF THE PAST [HOLDERLY HALL 6]  CAMERON, KATE  P
SHADOWS OF THE PAST [HOLDERLY HALL 2]  CAMERON, KATE  P
VOICES IN THE FOG [HOLDERLY HALL 3]  CAMERON, KATE  P
CONQUEROR’S CHILD, THE [HOLDFAST 4]  CHARNAS, SUZY MCKEE  H,LP
FURIES, THE [HOLDFAST 3]  CHARNAS, SUZY MCKEE  C/P,P,H
MOTHERLINES [HOLDFAST 2]  CHARNAS, SUZY MCKEE  H,P
SLAVE AND THE FREE, THE [HOLDFAST 1,2]  CHARNAS, SUZY MCKEE  LP
WALK TO THE END OF THE WORLD [HOLDFAST 1]  CHARNAS, SUZY MCKEE  C/P:2,P
CALL OF EARTH, THE [HOMECOMING 2]  CARD, ORSON SCOTT  C/P,P,H
EARTHBORN [HOMECOMING 5]  CARD, ORSON SCOTT  C/P,P,H:2
EARTHFALL [HOMECOMING 4]  CARD, ORSON SCOTT  H:2,P
HOMECOMING: EARTH [HOMECOMING 4,5]  CARD, ORSON SCOTT  H:2
HOMECOMING: HARMONY [HOMECOMING 1,2,3]  CARD, ORSON SCOTT  H:2
MEMORY OF EARTH, THE [HOMECOMING 1]  CARD, ORSON SCOTT  C/P,P,H
SHIPS OF EARTH, THE [HOMECOMING 3]  CARD, ORSON SCOTT  H:2,P
ASHES OF VICTORY [HONOR HARRINGTON 9]  WEBER, DAVID  H:2,P
AT ALL COSTS [HONOR HARRINGTON 11]  WEBER, DAVID  H
AT ALL COSTS DISK, THE (CD) [HONOR HARRINGTON]  BAEN PUBLISHING ENTERPRISES  SR-MM
CHANGER OF WORLDS [HONOR HARRINGTON/WORLDS OF HONOR 3]  WEBER, DAVID  HA
CROWN OF SLAVES [HONOR HARRINGTON]  WEBER, DAVID/FLINT, ERIC  H
ECHOES OF HONOR [HONOR HARRINGTON 8]  WEBER, DAVID  H:2,P
FIELD OF DISHONOR [HONOR HARRINGTON 4]  WEBER, DAVID  C/P:2,P
FLAG IN EXILE [HONOR HARRINGTON 5]  WEBER, DAVID  C/P,P,H
HONOR AMONG ENEMIES [HONOR HARRINGTON 6]  WEBER, DAVID  H,P
HONOR OF THE QUEEN, THE [HONOR HARRINGTON 2]  WEBER, DAVID  C/P:2,SFWA-P,P
HONORVERSE DISK, THE (CD) [HONOR HARRINGTON]  BAEN PUBLISHING ENTERPRISES  SR-MM
IN ENEMY HANDS [HONOR HARRINGTON 7]  WEBER, DAVID  H:2,P
MORE THAN HONOR [HONOR HARRINGTON/WORLDS OF HONOR 1]  WEBER, DAVID  C/PA,PA
ON BASILISK STATION [HONOR HARRINGTON 1]  WEBER, DAVID  C/P:2,SFWA-P,P

SERVICE OF THE SWORD, THE [HONOR HARRINGTON/ WORLDS OF HONOR 4] WEBER, DAVID

HA

SHORT VICTORIOUS WAR, THE [HONOR HARRINGTON 3]  WEBER, DAVID  C/P,SFWA-P,P
WAR OF HONOR (HARDCOVER AND CD) [HONOR HARRINGTON 10]  WEBER, DAVID  H:2

WARMASTERS, THE [HONOR HARRINGTON/HAMMER’S SLAMMERS/ BELISARIUS] FAWCETT, BILL

HA

WORLDS OF HONOR [HONOR HARRINGTON/WORLDS OF HONOR 2]  WEBER, DAVID  C/PA,HA,PA
SHADOW OF SAGANAMI, THE [HONOR HARRINGTON SAGANAMI 1]  WEBER, DAVID  H
WHIRLPOOL OF STARS [@HOOK #1]  ZETFORD, TULLY  P
BOOSTED MAN, THE [@HOOK #2]  ZETFORD, TULLY  C/P,P
STAR CITY [@HOOK #3]  ZETFORD, TULLY  P
VIRILITY GENE, THE [@HOOK #4]  ZETFORD, TULLY  C/P,P
ROAD TO HELL, THE [@HORIZON WAR #1/MAGE: THE ASCENSION]  WEINBERG, ROBERT  P
ASCENSION WARRIOR, THE [@HORIZON WAR #2/MAGE: THE ASCENSION]  WEINBERG, ROBERT  P
WAR IN HEAVEN [@HORIZON WAR #3/MAGE: THE ASCENSION]  WEINBERG, ROBERT  P

TITLE  AUTHOR  CODE
BLOWN DEAD [HORN]  SLOANE, BEN  P
HOT ZONE [HORN]  SLOANE, BEN  P
OUTLAND STRIP [HORN]  SLOANE, BEN  P
GREEN FLAMES OF ARIES, THE [@HORRORSCOPE #1]  LORY, ROBERT  P
REVENGE OF TAURUS, THE [@HORRORSCOPE #2]  LORY, ROBERT  P
CURSE OF LEO, THE [@HORRORSCOPE #3]  LORY, ROBERT  P
GEMINI SMILE, GEMINI KILL [@HORRORSCOPE #4]  LORY, ROBERT  C/P,P
DAUGHTER OF ZIR [HORSE GODDESS PEOPLE 2]  TARR, JUDITH  LP
LADY OF HORSES [HORSE GODDESS PEOPLE]  TARR, JUDITH  LP
WHITE MARE’S DAUGHTER [HORSE GODDESS PEOPLE 1]  TARR, JUDITH  H
BILI THE AXE [HORSECLANS 10]  ADAMS, ROBERT  C/P:2,P
CAT OF SILVERY HUE, A [HORSECLANS 4]  ADAMS, ROBERT  C/P:2,P
CHAMPION OF THE LAST BATTLE [HORSECLANS 11]  ADAMS, ROBERT  C/P:2,P
CLAN OF THE CATS, THE [HORSECLANS 18]  ADAMS, ROBERT  P
COMING OF THE HORSECLANS, THE [HORSECLANS 1]  ADAMS, ROBERT  C/P:2,P
DEATH OF A LEGEND, THE [HORSECLANS 8]  ADAMS, ROBERT  C/P:2,P
HORSECLANS ODYSSEY [HORSECLANS 7]  ADAMS, ROBERT  C/P:2,P
HORSES OF THE NORTH [HORSECLANS 13]  ADAMS, ROBERT  P
MADMAN’S ARMY [HORSECLANS 17]  ADAMS, ROBERT  P
MAN CALLED MILO MORAI, A [HORSECLANS 14]  ADAMS, ROBERT  C/P,P
MEMORIES OF MILO MORAI, THE [HORSECLANS 15]  ADAMS, ROBERT  C/P,P
PATRIMONY, THE [HORSECLANS 6]  ADAMS, ROBERT  C/P:2,P
REVENGE OF THE HORSECLANS [HORSECLANS 3]  ADAMS, ROBERT  C/P:2,P
SAVAGE MOUNTAINS, THE [HORSECLANS 5]  ADAMS, ROBERT  C/P:2,P
SWORDS OF THE HORSECLANS [HORSECLANS 2]  ADAMS, ROBERT  C/P:2,P
TALES OF THE HORSECLANS [HORSECLANS 1,2,3]  ADAMS, ROBERT  H
WITCH GODDESS, THE [HORSECLANS 9]  ADAMS, ROBERT  C/P:2,P
WOMAN OF THE HORSECLANS, A [HORSECLANS 12]  ADAMS, ROBERT  P
HORSEGIRL, THE [HORSEGIRL 1]  ASH, CONSTANCE  C/P,P
STALKING HORSE, THE [HORSEGIRL 2]  ASH, CONSTANCE  C/P,P
BEAST MASTER, THE [HOSTEEM STORM 1]  NORTON, ANDRE  C/P,P,H,DD509
BEAST MASTER TEAM [HOSTEEM STORM 3,4]  NORTON, ANDRE/MCCONCHIE, LYN  H
BEAST MASTER’S ARK [HOSTEEM STORM 3]  NORTON, ANDRE/MCCONCHIE, LYN  P
BEAST MASTER’S CIRCUS [HOSTEEM STORM 4]  NORTON, ANDRE/MCCONCHIE, LYN  P
BEAST MASTER’S QUEST [HOSTEEM STORM 5]  NORTON, ANDRE/MCCONCHIE, LYN  H
LORD OF THUNDER [HOSTEEM STORM 2]  NORTON, ANDRE  C/P:2,P
PROFITEER [@HOSTILE TAKEOVER #1]  SWANN, S. ANDREW  P
PARTISAN [@HOSTILE TAKEOVER #2]  SWANN, S. ANDREW  P
REVOLUTIONARY [@HOSTILE TAKEOVER #3]  SWANN, S. ANDREW  C/P,P
STRANGER BY NIGHT [HOT BLOOD]  GELB, JEFF/GARRETT, MICHAEL  PA
HOUND AND THE FALCON, THE [HOUND AND FALCON 1,2,3]  TARR, JUDITH  H:2
ISLE OF GLASS, THE [HOUND AND FALCON #1]  TARR, JUDITH  C/P,@H,SFWA-P,P
GOLDEN HORN, THE [HOUND AND FALCON #2]  TARR, JUDITH  C/P:2,@H,SFWA-P,P
HOUNDS OF GOD, THE [HOUND AND FALCON #3]  TARR, JUDITH  P,@H,SFWA-P
SWORD AND THE EYE, THE [HOUSE OF EIGIN 1]  LEIBER, JUSTIN  C/P,SFWA-P,P
SWORD AND THE TOWER, THE [HOUSE OF EIGIN 2]  LEIBER, JUSTIN  P,SFWA-P
HOUSE OF MOONS [HOUSE OF MOONS 2]  WENTWORTH, K. D.  P
MOONSPEAKER [HOUSE OF MOONS 1]  WENTWORTH, K. D.  P
EONS OF THE NIGHT [@HOWARD LIBRARY #5]  HOWARD, ROBERT E.  P
HOWLING, THE [@HOWLING #1]  BRANDNER, GARY  P
HOWLING II, THE [@HOWLING #2]  BRANDNER, GARY  P
HOWLING III, THE [@HOWLING #3]  BRANDNER, GARY  P
OATH OF SWORDS [HRADANI 2]  WEBER, DAVID  C/P,P
WAR GOD’S OWN, THE [HRADANI 1]  WEBER, DAVID  C/P,P,H
WIND RIDER’S OATH (HARDCOVER AND CD) [HRADANI 3]  WEBER, DAVID  H:2
WIND RIDER’S OATH DISK, THE (CD) [HRADANI]  BAEN PUBLISHING ENTERPRISES  SR-MM
FIRE’S STONE, THE [HUFF 4]  HUFF, TANYA  C/P,P
GATE OF DARKNESS, CIRCLE OF LIGHT [HUFF 3]  HUFF, TANYA  C/P:2,P
OF DARKNESS, LIGHT, AND FIRE (OMNIBUS) [HUFF 3,4]  HUFF, TANYA  P
HULK [HULK]  DAVID, PETER  C/P,P
STALKER FROM THE STARS [HULK]  WEIN, LEN/WOLFMAN, MARV/SILVA, JOSEPH C/P,P
WHAT SAVAGE BEAST [HULK]  DAVID, PETER  P
ALIEN DEBT, THE [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
REBEL DYNASTY 2, THE [HULZEIN DYNASTY]  BUSBY, F. M.  P
REBEL’S QUEST [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
REBELS’ SEED [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
RISSA AND TREGARE [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
RISSA KERGUELEN [HULZEIN DYNASTY]  BUSBY, F. M.  C/P,P,H
STAR REBEL [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
YOUNG RISSA [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
ZELDE M’TANA [HULZEIN DYNASTY]  BUSBY, F. M.  C/P,P
CACHALOT [HUMANX]  FOSTER, ALAN DEAN  H:2,P
DROWNING WORLD [HUMANX]  FOSTER, ALAN DEAN  H

TITLE  AUTHOR  CODE
FLINX’S FOLLY [HUMANX]  FOSTER, ALAN DEAN  P
HOWLING STONES, THE [HUMANX]  FOSTER, ALAN DEAN  H:2,P
MIDWORLD [HUMANX]  FOSTER, ALAN DEAN  H:2,P
NOR CRYSTAL TEARS [HUMANX]  FOSTER, ALAN DEAN  C/P:2,P
SENTENCED TO PRISM [HUMANX]  FOSTER, ALAN DEAN  C/P:2,P
VOYAGE TO THE CITY OF THE DEAD [HUMANX]  FOSTER, ALAN DEAN  C/P:2,P
BLOODHYPE [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
END OF THE MATTER, THE [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
FLINX IN FLUX [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
FLINX’S FOLLY (PROOF) [HUMANX FLINX]  FOSTER, ALAN DEAN  SR-LP
FOR LOVE OF MOTHER-NOT [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
MID-FLINX [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P,P,H
ORPHAN STAR [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
REUNION [HUMANX FLINX]  FOSTER, ALAN DEAN  H
RUNNING FROM THE DEITY [HUMANX FLINX]  FOSTER, ALAN DEAN  H
SLIDING SCALES [HUMANX FLINX]  FOSTER, ALAN DEAN  P
TAR-AIYM KRANG, THE [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
TROUBLE MAGNET [HUMANX FLINX]  FOSTER, ALAN DEAN  H
...WHO NEEDS ENEMIES? [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
DIRGE [HUMANX FOUNDING COMMONWEALTH 2]  FOSTER, ALAN DEAN  H
DIUTURNITY’S DAWN [HUMANX FOUNDING COMMONWEALTH 3]  FOSTER, ALAN DEAN  H
PHYLOGENESIS [HUMANX FOUNDING COMMONWEALTH 1]  FOSTER, ALAN DEAN  H
DELUGE DRIVERS, THE [HUMANX ICERIGGER 3]  FOSTER, ALAN DEAN  C/P:2,P
ICERIGGER [HUMANX ICERIGGER 1]  FOSTER, ALAN DEAN  C/P,P,H
MISSION TO MOULOKIN [HUMANX ICERIGGER 2]  FOSTER, ALAN DEAN  H:2,P
MISSION TO MOULOKIN (PROOF) [HUMANX ICERIGGER 2]  FOSTER, ALAN DEAN  SR-LP
MOULOKIN MISSION, DIE [HUMANX ICERIGGER 2]  FOSTER, ALAN DEAN  FP
DELTA CITY [HUMILITY GARDEN 2]  SAVAGE, FELICITY  P
GARDEN OF SALT [HUMILITY GARDEN 1,2]  SAVAGE, FELICITY  H
HUMILITY GARDEN [HUMILITY GARDEN 1]  SAVAGE, FELICITY  C/P,P
DARKLING PLAIN, A [HUNGRY CITY 4]  REEVE, PHILIP  H
DARKLING PLAIN, A (PROOF) [@HUNGRY CITY 4]  REEVE, PHILIP  SR-LP
INFERNAL DEVICES [HUNGRY CITY 3]  REEVE, PHILIP  H
INFERNAL DEVICES (PROOF) [@HUNGRY CITY 3]  REEVE, PHILIP  SR-LP
MORTAL ENGINES [@HUNGRY CITY 1]  REEVE, PHILIP  H
MORTAL ENGINES (PROOF) [@HUNGRY CITY 1]  REEVE, PHILIP  SR-LP
PREDATOR’S GOLD [HUNGRY CITY 2]  REEVE, PHILIP  H
RING OF CHARON, THE [HUNTED EARTH 1]  ALLEN, ROGER MACBRIDE  C/P,SFWA-P,P,H
SHATTERED SPHERE, THE [HUNTED EARTH 2]  ALLEN, ROGER MACBRIDE  H,P
HUNTER, THE [HUNTER 1]  ESTES, ROSE  C/P,P
HUNTER ON ARENA, THE [HUNTER 2]  ESTES, ROSE  P
HUNTER VICTORIOUS, THE [HUNTER 3]  ESTES, ROSE  P
OMEGA OPERATION, THE [HUNTER 1]  CONWAY, NORMAN  P
INHERIT THE EARTH [@HUNTER THE RECKONING]  WIECK, STEWART  PA
ALIEN CITADEL [HUNTSMAN 3]  HILL, DOUGLAS  H
HUNTSMAN, THE [HUNTSMAN 1]  HILL, DOUGLAS  H
WARRIORS OF THE WASTELAND [HUNTSMAN 2]  HILL, DOUGLAS  H
ENDYMION [HYPERION 3]  SIMMONS, DAN  H:2,P
FALL OF HYPERION, THE [HYPERION 2]  SIMMONS, DAN  C/P:2,P,H
HYPERION [HYPERION 1]  SIMMONS, DAN  C/P:2,P
HYPERION CANTOS [HYPERION 1,2]  SIMMONS, DAN  H:3
RISE OF ENDYMION, THE [HYPERION 4]  SIMMONS, DAN  H:2,P
HARSHINI [HYTHRUN 3]  FALLON, JENNIFER  H
MEDALON [HYTHRUN 1]  FALLON, JENNIFER  H
TREASON KEEP [HYTHRUN 2]  FALLON, JENNIFER  H
WARRIOR [HYTHRUN WOLFBLADE 2]  FALLON, JENNIFER  H
WOLFBLADE [HYTHRUN WOLFBLADE 1]  FALLON, JENNIFER  H

The MITSFS Seriesdex  September 24, 2007  Page 83
TITLE  AUTHOR  CODE
WIPEOUT [@I SPY 4]  TIGER, JOHN  P
DARKNESS UPON THE ICE, A [ICE 3]  FORSTCHEN, WILLIAM R.  C/P,P
FLAME UPON THE ICE, THE [ICE 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
ICE PROPHET [ICE 1]  FORSTCHEN, WILLIAM R.  C/P:2,P

INDEPENDENCE DAY [@ID4] DEVLIN, DEAN/EMMERICH, ROLAND/

MOLSTAD, STEPHEN

P

SILENT ZONE [@ID4]  MOLSTAD, STEPHEN  H,P
WAR IN THE DESERT [@ID4]  MOLSTAD, STEPHEN  C/P,P
EVERFREE [IDLEWILD 3]  SAGAN, NICK  H
IDLEWILD [IDLEWILD 1]  SAGAN, NICK  H
IDLEWILD (PROOF) [IDLEWILD 1]  SAGAN, NICK  SR-LP
LION’S EYE, THE [@ILARIO 1]  GENTLE, MARY  LP
ILIUM [ILIUM 1]  SIMMONS, DAN  H
OLYMPOS [ILIUM 2]  SIMMONS, DAN  H
BINDER’S ROAD, THE [ILLUMINATION 2]  MCGARRY, TERRY  P
ILLUMINATION [ILLUMINATION 1]  MCGARRY, TERRY  H
TRIAD [ILLUMINATION 3]  MCGARRY, TERRY  H
EYE IN THE PYRAMID, THE [ILLUMINATUS! 1]  SHEA, ROBERT/WILSON, ROBERT ANTON  C/P,P
GOLDEN APPLE, THE [ILLUMINATUS! 2]  SHEA, ROBERT/WILSON, ROBERT ANTON  C/P:2,P
ILLUMINATUS! TRILOGY, THE [ILLUMINATUS! 1,2,3]  SHEA, ROBERT/WILSON, ROBERT ANTON  H:3,LP
LEVIATHAN [ILLUMINATUS! 3]  SHEA, ROBERT/WILSON, ROBERT ANTON  C/P:2,P
IMARO [IMARO 1]  SAUNDERS, CHARLES R.  C/P,P
QUEST FOR CUSH, THE [IMARO 2]  SAUNDERS, CHARLES R.  C/P:2,P
TRAIL OF BOHU, THE [IMARO 3]  SAUNDERS, CHARLES R.  P
TOYBOX, THE [@IMMORTAL EYES #1]  CASSADA, JACKIE  P
SHADOWS ON THE HILL [@IMMORTAL EYES #2]  CASSADA, JACKIE  P
CALLING, THE [@IMMORTALS]  ASHLEY, JENNIFER  P
DARKENING, THE [@IMMORTALS]  POPP, ROBIN T.  P
IMMORTALS, THE [IMMORTALS (PIERCE) 1,2,3,4]  PIERCE, TAMORA  H
WILD MAGIC [@IMMORTALS (PIERCE) #1]  PIERCE, TAMORA  P
WOLF-SPEAKER [@IMMORTALS (PIERCE) #2]  PIERCE, TAMORA  P
EMPEROR MAGE [@IMMORTALS (PIERCE) #3]  PIERCE, TAMORA  P
REALMS OF THE GODS, THE [@IMMORTALS (PIERCE) #4]  PIERCE, TAMORA  P
HUNGRE LIKE NO OTHER, A [IMMORTALS AFTER DARK 1]  COLE, KRESLEY  H
NO REST FOR THE WICKED [IMMORTALS AFTER DARK 2]  COLE, KRESLEY  H
STARS AT WAR, THE [@IMPERIAL STARS #1]  POURNELLE, JERRY  C/PA,PA
REPUBLIC AND EMPIRE [@IMPERIAL STARS #2]  POURNELLE, JERRY  C/PA,PA,SFWA-P
CRASH OF EMPIRE, THE [@IMPERIAL STARS #3]  POURNELLE, JERRY  SFWA-P
ASSIGNMENT IN NOWHERE [IMPERIUM 3]  LAUMER, KEITH  C/P:2,P
BEYOND THE IMPERIUM [IMPERIUM 2,3]  LAUMER, KEITH  C/P:2,P
OTHER SIDE OF TIME, THE [IMPERIUM 2]  LAUMER, KEITH  SR-H,C/P:2,P
WORLDS OF THE IMPERIUM [IMPERIUM 1]  LAUMER, KEITH  C/P:2,DF127,SFWA-P,P
WORLDS OF THE IMPERIUM (EXPANDED) [IMPERIUM 1]  LAUMER, KEITH  C/P:2,P
ZONE YELLOW [IMPERIUM 4]  LAUMER, KEITH  C/P,P
DESTROYER GODDESS, THE [IN FIRE FORGED 2]  RESNICK, LAURA  P
WHITE DRAGON, THE [IN FIRE FORGED 1]  RESNICK, LAURA  H,P
AND ETERNITY [INCARNATIONS OF IMMORTALITY 7]  ANTHONY, PIERS  C/P:2,P,H
BEARING AN HOURGLASS [INCARNATIONS OF IMMORTALITY 2]  ANTHONY, PIERS  C/P,P,H
BEING A GREEN MOTHER [INCARNATIONS OF IMMORTALITY 5]  ANTHONY, PIERS  C/P,P,H
FOR LOVE OF EVIL [INCARNATIONS OF IMMORTALITY 6]  ANTHONY, PIERS  C/P:2,P
ON A PALE HORSE [INCARNATIONS OF IMMORTALITY 1]  ANTHONY, PIERS  H:2,P
WIELDING A RED SWORD [INCARNATIONS OF IMMORTALITY 4]  ANTHONY, PIERS  H:2,P,SFWA
WITH A TANGLED SKEIN [INCARNATIONS OF IMMORTALITY 3]  ANTHONY, PIERS  H:2,P,SFWA

COMPLEAT ENCHANTER: THE MAGICAL MISADVENTURES OF HAROLD SHEA, THE [INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/PRATT, FLETCHER

C/P,P,H

COMPLETE COMPLEAT ENCHANTER, THE [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  C/P:2,SFWA-P,P

ENCHANTER COMPLEATED, THE (ORIGINALLY WALL OF SERPENTS)

[INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/PRATT, FLETCHER

SR-P

ENCHANTER REBORN, THE [INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/

STASHEFF, CHRISTOPHER

PA,C/PA,SFWA-P

EXOTIC ENCHANTER, THE [INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/

STASHEFF, CHRISTOPHER

PA

INCOMPLETE ENCHANTER, THE [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  SR-H,C/P:2,P
SIR HAROLD AND THE GNOME KING [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE  LP
WALL OF SERPENTS [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  SR-H,P
EXIT STRATEGY [INCONSTANT MOON 3]  ASKEGREN, PIERCE  P
FALL GIRL [INCONSTANT MOON 2]  ASKEGREN, PIERCE  C/P,P
HUMAN RESOURCE [INCONSTANT MOON 1]  ASKEGREN, PIERCE  P
INDIAN IN THE CUPBOARD, THE [INDIAN IN THE CUPBOARD 1]  BANKS, LYNNE REID  C/P:2,P
INDIANA JONES AND THE DANCE OF THE GIANTS [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE GENESIS DELUGE [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE INTERIOR WORLD [INDIANA JONES]  MACGREGOR, ROB  C/P,P
INDIANA JONES AND THE LAST CRUSADE [INDIANA JONES]  MACGREGOR, ROB  P

The MITSFS Seriesdex  September 24, 2007  Page 84
TITLE  AUTHOR  CODE
INDIANA JONES AND THE SEVEN VEILS [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE SKY PIRATES [INDIANA JONES]  CAIDIN, MARTIN  C/P,P
INDIANA JONES AND THE UNICORN’S LEGACY [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE WHITE WITCH [INDIANA JONES]  CAIDIN, MARTIN  P
AISLING [INDIGO 8]  COOPER, LOUISE  C/P:2,P
NEMESIS [@INDIGO #1]  COOPER, LOUISE  C/P,SFWA-P,P
INFERNO [@INDIGO #2]  COOPER, LOUISE  C/P:2,SFWA-P,P
INFANTA [@INDIGO #3]  COOPER, LOUISE  C/P,SFWA-P,P
NOCTURNE [@INDIGO #4]  COOPER, LOUISE  C/P,SFWA-P,P
TROIKA [@INDIGO #5]  COOPER, LOUISE  C/P:2,P
AVATAR [@INDIGO #6]  COOPER, LOUISE  C/P:2,P
REVENANT [@INDIGO #7]  COOPER, LOUISE  C/P,P
BORDER DISPUTE [INDIW 2]  KERNS, DANIEL R.  P
DOG EAT DOG [INGERSOL 2]  CARROLL, JERRY JAY  H,LP
TOP DOG [INGERSOL 1]  CARROLL, JERRY JAY  C/P,P
ELDEST [INHERITANCE 2]  PAOLINI, CHRISTOPHER  H
ERAGON [INHERITANCE 1]  PAOLINI, CHRISTOPHER  H
INHERITOR, THE [INHERITOR 1]  BRADLEY, MARION ZIMMER  C/P:2,LP
WITCH HILL [INHERITOR 2]  BRADLEY, MARION ZIMMER  C/P:2,SFWA-P,P
GATHERER OF CLOUDS [INITIATE BROTHER 2]  RUSSELL, SEAN  C/P:2,SFWA-P,P
INITIATE BROTHER, THE [INITIATE BROTHER 1]  RUSSELL, SEAN  C/P:2,P
DARKLING WIND, THE [INQUESTOR 4]  SUCHARITKUL, SOMTOW  C/P:2,P

DAWNING SHADOWS: LIGHT ON THE SOUND, THE (REVISED)

[INQUESTOR 1] SUCHARITKUL, SOMTOW

C/P:2,P

DAWNING SHADOWS: THRONE OF MADNESS, THE (REVISED)

[INQUESTOR 2] SUCHARITKUL, SOMTOW

P

LIGHT ON THE SOUND [INQUESTOR 1]  SUCHARITKUL, SOMTOW  C/P,P
THRONE OF MADNESS, THE [INQUESTOR 2]  SUCHARITKUL, SOMTOW  C/P:2,P
UTOPIA HUNTERS [INQUESTOR 3]  SUCHARITKUL, SOMTOW  C/P:2,P
TYRANNY OF THE NIGHT, THE [@INSTRUMENTALITIES OF THE NIGHT #1]  COOK, GLEN  H

LORD OF THE SILENT KINGDOM [@INSTRUMENTALITIES OF THE NIGHT #2] COOK, GLEN

H

INSTRUMENTALITY OF MANKIND, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P:2,P
NORSTRILIA [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P:2,P

NORSTRILIA (COMPLETE AND CORRECTED)

[INSTRUMENTALITY OF MANKIND] SMITH, CORDWAINER

SR-H

PLANET BUYER, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P:2,P
REDISCOVERY OF MAN, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  SR-H
UNDERPEOPLE, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P,P
INTEGRAL TREES, THE [INTEGRAL TREES 1]  NIVEN, LARRY  H:2,P,SFWA
SMOKE RING, THE [INTEGRAL TREES 2]  NIVEN, LARRY  H:2,P,SFWA
INTERFACE [INTERFACE 1]  ADLARD, MARK  C/P,P
MULTIFACE [INTERFACE 3]  ADLARD, MARK  P
VOLTEFACE [INTERFACE 2]  ADLARD, MARK  P
ROBOT WORLD [INTERPLANETARY SPY 3]  MCEVOY, SETH  SCX
INTERSTELLAR PATROL [INTERSTELLAR PATROL]  ANVIL, CHRISTOPHER  P

INTERSTELLAR PATROL II: THE FEDERATION OF HUMANITY [INTERSTELLAR PATROL] ANVIL, CHRISTOPHER

H

INTERVENTION [INTERVENTION 1,2]  MAY, JULIAN  H:2,SFWA
SURVEILLANCE, THE [@INTERVENTION #1]  MAY, JULIAN  C/P:2,P
METACONCERT, THE [@INTERVENTION #2]  MAY, JULIAN  C/P:2,P

INTERZONE: THE FIRST ANTHOLOGY [INTERZONE 1] CLUTE, JOHN/GREENLAND, COLIN/

PRINGLE, DAVID

HA

INTERZONE: THE SECOND ANTHOLOGY [INTERZONE 2] CLUTE, JOHN/PRINGLE, DAVID/

OUNSLEY, SIMON

PA

DARKNESS DESCENDING [INTO THE DARKNESS 2]  TURTLEDOVE, HARRY  C/P,P,H
INTO THE DARKNESS [INTO THE DARKNESS 1]  TURTLEDOVE, HARRY  C/P,P,H
JAWS OF DARKNESS [INTO THE DARKNESS 5]  TURTLEDOVE, HARRY  P
OUT OF THE DARKNESS [INTO THE DARKNESS 6]  TURTLEDOVE, HARRY  P
RULERS OF THE DARKNESS [INTO THE DARKNESS 4]  TURTLEDOVE, HARRY  H
THROUGH THE DARKNESS [INTO THE DARKNESS 3]  TURTLEDOVE, HARRY  C/P,P,H
ARMY OF THE UNDEAD [INVADERS 3]  BERNARD, RAFE  P
DAM OF DEATH [INVADERS]  PEARL, JACK  H
ENEMIES FROM BEYOND [INVADERS 2]  LAUMER, KEITH  P
INVADERS, THE [INVADERS 1]  LAUMER, KEITH  C/P:2,P
INVASION AMERICA [INVASION AMERICA 1]  GOLDEN, CHRISTIE  P
ON THE RUN [INVASION AMERICA 2]  GOLDEN, CHRISTIE  P
LAST ECHO, THE [INVASION: EARTH]  CAVE, PETER  LP
"THEY’RE HERE..." [INVASION OF THE BODY SNATCHERS]  MCCARTHY, KEVIN/GORMAN, ED  R/REF-P
DARK SONATA [IRA VEDDER 4]  BOND, EVELYN  P
GIRL FROM NOWHERE, THE [IRA VEDDER 2]  BOND, EVELYN  P
GRAYMANTLE [IRON ANGEL 2]  MORRESSY, JOHN  C/P:2,P
IRONBRAND [IRON ANGEL 1]  MORRESSY, JOHN  C/P,P

TITLE  AUTHOR  CODE
KINGSBANE [IRON ANGEL 3]  MORRESSY, JOHN  P
MOUNTAINS AND MADNESS [@IRON DRAGONS]  ESTES, ROSE  P,SFWA-P
AND CALL MY KILLER...MODOK! [IRON MAN]  ROTSLER, WILLIAM  P
DARK TIDE, THE [@IRON TOWER #1/MITHGAR 7]  MCKIERNAN, DENNIS L.  C/P:2,SFWA,P
SHADOWS OF DOOM [@IRON TOWER #2/MITHGAR 8]  MCKIERNAN, DENNIS L.  C/P,SFWA,P,H
DARKEST DAY, THE [@IRON TOWER #3/MITHGAR 9]  MCKIERNAN, DENNIS L.  C/P:2,SFWA,P
GLAIVE, THE [IROSHI 2]  OSBORNE, CARY  P
IROSHI [IROSHI 1]  OSBORNE, CARY  C/P,P
PERSEA [IROSHI 3]  OSBORNE, CARY  P
FOSSIL [ISAAC’S UNIVERSE]  CLEMENT, HAL  C/P:2,P
DIPLOMACY GUILD, THE [@ISAAC’S UNIVERSE #1]  GREENBERG, MARTIN H.  C/PA,PA
PHASES IN CHAOS [@ISAAC’S UNIVERSE #2]  GREENBERG, MARTIN H.  PA
UNNATURAL DIPLOMACY [@ISAAC’S UNIVERSE #3]  GREENBERG, MARTIN H.  PA,SFWA-P
FIREBIRD’S VENGEANCE, THE [ISAVALTA 3]  ZETTEL, SARAH  H
SORCERER’S TREASON, A [ISAVALTA 1]  ZETTEL, SARAH  H
SWORD OF THE DECEIVER [ISAVALTA 4]  ZETTEL, SARAH  H
USURPER’S CROWN, THE [ISAVALTA 2]  ZETTEL, SARAH  H
ISIS RISING [ISIS 2]  STEWART, JEAN  LP
KEEPER OF THE ISIS LIGHT, THE [ISIS 1]  HUGHES, MONICA  H,P
WARRIORS OF ISIS [ISIS 3]  STEWART, JEAN  LP
WINGED ISIS [ISIS 4]  STEWART, JEAN  LP
WIZARD OF ISIS [ISIS 5]  STEWART, JEAN  LP
SORCEROUS SEA [ISLAND WARRIOR 3]  SEVARANCE, CAROL  P
STORM CALLER [ISLAND WARRIOR 2]  SEVERANCE, CAROL  P
ACT OF GOD [ISLAND WORLDS 1]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  P
BETWEEN THE STARS [ISLAND WORLDS 3]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  C/P:2,SFWA-P,P
DELTA PAVONIS [ISLAND WORLDS 4]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  C/P,P
ISLAND WORLDS, THE [ISLAND WORLDS 2]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  C/P,P
HAVOC IN ISLANDIA [ISLANDIA]  SAXTON, MARK  H
TWO KINGDOMS, THE [ISLANDIA]  SAXTON, MARK  H:2
AWARE, THE [@ISLES OF GLORY #1]  LARKE, GLENDA  C/P,P
GILFEATHER [@ISLES OF GLORY #2]  LARKE, GLENDA  P
TAINTED, THE [@ISLES OF GLORY #3]  LARKE, GLENDA  P
COMPLETE IVORY, THE [IVORY 1,2,3]  EGAN, DORIS  P
GATE OF IVORY, THE [IVORY 1]  EGAN, DORIS  C/P:2,P
GUILT-EDGED IVORY [IVORY 3]  EGAN, DORIS  C/P:2,P
TWO-BIT HEROES [IVORY 2]  EGAN, DORIS  C/P,P

TITLE  AUTHOR  CODE
HONG ON THE RANGE [JACK HONG]  WU, WILLIAM F.  H
SHADE OF LO MAN GONG, THE [JACK HONG 1]  WU, WILLIAM F.  B
SHAUNESSY FONG [JACK HONG 2]  WU, WILLIAM F.  B
CARDINAL OF THE KREMLIN, THE [JACK RYAN]  CLANCY, TOM  P
DRINK DOWN THE MOON [JACKY ROWAN 2]  DE LINT, CHARLES  C/P:2,SFWA-P:2,P
JACK OF KINROWAN [JACKY ROWAN 1,2]  DE LINT, CHARLES  P
JACK, THE GIANT-KILLER [JACKY ROWAN 1]  DE LINT, CHARLES  C/P,SFWA-P,P,H
JACK THE GIANT KILLER (PROOF) [JACKY ROWAN 1]  DE LINT, CHARLES  SR-LP
JADE DARCY AND THE AFFAIR OF HONOR [@JADE DARCY #1]  GOLDIN, STEPHEN/MASON, MARY  C/P,P
JADE DARCY AND THE ZEN PIRATES [@JADE DARCY #2]  GOLDIN, STEPHEN/MASON, MARY  C/P:2,P
QUAKING LANDS, THE [@JADE DEMONS #1]  VARDEMAN, ROBERT E.  P
FROZEN WAVES, THE [@JADE DEMONS #2]  VARDEMAN, ROBERT E.  P
CRYSTAL CLOUDS, THE [@JADE DEMONS #3]  VARDEMAN, ROBERT E.  P
WHITE FIRE, THE [@JADE DEMONS #4]  VARDEMAN, ROBERT E.  P
DUKE OF URANIUM, THE [JAK JINNAKA 1]  BARNES, JOHN  C/P:2,P
IN THE HALL OF THE MARTIAN KING [JAK JINNAKA 3]  BARNES, JOHN  C/P,P
PRINCESS OF THE AERIE, A [JAK JINNAKA 2]  BARNES, JOHN  P
LOGOS RUN [JAK REBO 2]  DIETZ, WILLIAM C.  H
RUNNER [JAK REBO 1]  DIETZ, WILLIAM C.  P
CHOSEN OF MIDA [JALAV 3]  GREEN, SHARON  C/P,P
CRYSTALS OF MIDA, THE [JALAV 1]  GREEN, SHARON  C/P,P
OATH TO MIDA, AN [JALAV 2]  GREEN, SHARON  C/P,P
TO BATTLE THE GODS [JALAV 5]  GREEN, SHARON  P
WILL OF THE GODS, THE [JALAV 4]  GREEN, SHARON  P
BROKEN CLAW [JAMES BOND 23]  GARDNER, JOHN (BRITISH, 1926-)  P
CASINO ROYALE [JAMES BOND 1]  FLEMING, IAN  C/P,P
COLONEL SUN [JAMES BOND 14]  AMIS, KINGSLEY  P
DIAMONDS ARE FOREVER [JAMES BOND 4]  FLEMING, IAN  P
DOCTOR NO [JAMES BOND 6]  FLEMING, IAN  C/P:2,P
FOR SPECIAL SERVICES [JAMES BOND 16]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
FOR YOUR EYES ONLY [JAMES BOND 8]  FLEMING, IAN  C/P:2,P
FROM RUSSIA, WITH LOVE [JAMES BOND 5]  FLEMING, IAN  C/P,P
GOLDFINGER [JAMES BOND 7]  FLEMING, IAN  C/P,P
ICEBREAKER [JAMES BOND 17]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
LICENSE RENEWED [JAMES BOND 15]  GARDNER, JOHN (BRITISH, 1926-)  P
LIVE AND LET DIE [JAMES BOND 2]  FLEMING, IAN  C/P,P
MAN WITH THE GOLDEN GUN, THE [JAMES BOND 13]  FLEMING, IAN  C/P,P
MOONRAKER [JAMES BOND 3]  FLEMING, IAN  C/P:2,P
NO DEALS, MR. BOND [JAMES BOND 20]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
NOBODY LIVES FOREVER [JAMES BOND 19]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
ON HER MAJESTY’S SECRET SERVICE [JAMES BOND 11]  FLEMING, IAN  C/P,P
ROLE OF HONOR [JAMES BOND 18]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
SCORPIUS [JAMES BOND 21]  GARDNER, JOHN (BRITISH, 1926-)  P
SPY WHO LOVED ME, THE [JAMES BOND 10]  FLEMING, IAN  C/P:2,P
THUNDERBALL [JAMES BOND 9]  FLEMING, IAN  C/P:2,P
WIN, LOSE, OR DIE [JAMES BOND 22]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
YOU ONLY LIVE TWICE [JAMES BOND 12]  FLEMING, IAN  C/P:2,P
WEREWOLF’S REVENGE, THE [JAMES UNDERHILL 2]  JACCOMA, RICHARD  P
WEREWOLF’S TALE, THE [JAMES UNDERHILL 1]  JACCOMA, RICHARD  P
BLOOD & IVORY [JAMETHIEL DREAM-WEAVER]  HODGELL, P. C.  LP
DARK OF THE MOON [JAMETHIEL DREAM-WEAVER 2]  HODGELL, P. C.  H:2,SFWA-P,P
GOD STALK [JAMETHIEL DREAM-WEAVER 1]  HODGELL, P. C.  H:2,P
SEEKER’S MASK [JAMETHIEL DREAM-WEAVER 3]  HODGELL, P. C.  SR-H,H:2
TO RIDE A RATHORN [JAMETHIEL DREAM-WEAVER 4]  HODGELL, P. C.  H
MARS [JAMIE WATERMAN 1]  BOVA, BEN  C/P:2,P,H
RETURN TO MARS [JAMIE WATERMAN 2]  BOVA, BEN  C/P,P,H
ALL THE COLORS OF DARKNESS [JAN DARZEK]  BIGGLE, LLOYD JR.  C/P,P,H:2
SILENCE IS DEADLY [JAN DARZEK]  BIGGLE, LLOYD JR.  H:2,SFWA
THIS DARKENING UNIVERSE [JAN DARZEK]  BIGGLE, LLOYD JR.  H,SFWA
WATCHERS OF THE DARK [JAN DARZEK]  BIGGLE, LLOYD JR.  H:2
WHIRLIGIG OF TIME, THE [JAN DARZEK]  BIGGLE, LLOYD JR.  H
BLACK LEGION OF CALLISTO [JANDAR 2]  CARTER, LIN  C/P:2,P
JANDAR OF CALLISTO [JANDAR 1]  CARTER, LIN  C/P:2,P
LANKAR OF CALLISTO [JANDAR 6]  CARTER, LIN  C/P,P
MAD EMPRESS OF CALLISTO [JANDAR 4]  CARTER, LIN  C/P,P
MIND WIZARDS OF CALLISTO [JANDAR 5]  CARTER, LIN  C/P,P
RENEGADE OF CALLISTO [JANDAR 8]  CARTER, LIN  P
SKY PIRATES OF CALLISTO [JANDAR 3]  CARTER, LIN  C/P,P
YLANA OF CALLISTO [JANDAR 7]  CARTER, LIN  C/P,P
CODE OF CONDUCT [JANI KILIAN 1]  SMITH, KRISTINE  C/P,P,H
CONTACT IMMINENT [JANI KILIAN 4]  SMITH, KRISTINE  C/P,P
CONTACT IMMINENT (PROOF) [JANI KILIAN 4]  SMITH, KRISTINE  SR-P
ENDGAME (PROOF) [JANI KILIAN 5]  SMITH, KRISTINE  SR-P

TITLE  AUTHOR  CODE
LAW OF SURVIVAL [JANI KILIAN 3]  SMITH, KRISTINE  C/P:2,P
RULES OF CONFLICT [JANI KILIAN 2]  SMITH, KRISTINE  C/P,P
CLAN AND CROWN [@JANISSARIES 2]  POURNELLE, JERRY/GREEN, ROLAND  C/P,LP,H
JANISSARIES [@JANISSARIES 1]  POURNELLE, JERRY  C/P,LP,H
LORD OF LANCES [JANISSARIES]  ACRES, MARK  P
STORMS OF VICTORY [@JANISSARIES 3]  POURNELLE, JERRY/GREEN, ROLAND  C/P,SFWA,P,H
TRAN [JANISSARIES]  POURNELLE, JERRY/GREEN, ROLAND  P
BRIDLING CHAOS [JANNA BRILL 1,2,3]  KILLOUGH, LEE  LP
DOPPELGANGER GAMBIT, THE [JANNA BRILL 1]  KILLOUGH, LEE  C/P,P
DRAGON’S TEETH [JANNA BRILL 3]  KILLOUGH, LEE  C/P:2,P
SPIDER PLAY [JANNA BRILL 2]  KILLOUGH, LEE  P
JUDGMENT ON JANUS [JANUS 1/DIPPLE 2]  NORTON, ANDRE  H:2,P
VICTORY ON JANUS [JANUS 2]  NORTON, ANDRE  C/P,P,H
JARAN [JARAN 1]  ELLIOTT, KATE  C/P:2,SFWA-P,P
LAW OF BECOMING, THE [@JARAN #4]  ELLIOTT, KATE  C/P,P
RAVYN’S FLIGHT [JARVED 9 1]  O’SHEA, PATTI  P
DIRTY WORK [JASON COSMO 3]  MCGIRT, DAN  C/P,P
JASON COSMO [JASON COSMO 1]  MCGIRT, DAN  C/P,P
ORPHANAGE [JASON WANDER 1]  BUETTNER, ROBERT  C/P,P
ORPHAN’S DESTINY [JASON WANDER 2]  BUETTNER, ROBERT  P
CHAIN OF COMMAND [JEFFERSON’S WAR]  RANDLE, KEVIN  P
DEATH OF A REGIMENT [JEFFERSON’S WAR 5]  RANDLE, KEVIN  C/P,P
GALACTIC SILVER STAR, THE [JEFFERSON’S WAR 1]  RANDLE, KEVIN  P
JANUARY PLATOON, THE [JEFFERSON’S WAR 4]  RANDLE, KEVIN  C/P,P
LOST COLONY, THE [JEFFERSON’S WAR 3]  RANDLE, KEVIN  C/P,P
PRICE OF COMMAND, THE [JEFFERSON’S WAR 2]  RANDLE, KEVIN  C/P,P
BOOKS OF GREAT ALTA, THE [JENNA 1,2]  YOLEN, JANE  H
HAMMERED [JENNY CASEY 1]  BEAR, ELIZABETH  C/P,P
SCARDOWN [JENNY CASEY 2]  BEAR, ELIZABETH  P
WORLDWIRED [JENNY CASEY 3]  BEAR, ELIZABETH  P
CIRCLE OF ONE [JENNY SIXA 1]  FULLILOVE, ERIC JAMES  C/P,P
STRANGER, THE [JENNY SIXA 2]  FULLILOVE, ERIC JAMES  P
UNCHARTED STARS [JERN MURDOCK 2]  NORTON, ANDRE  H:2,P
ZERO STONE, THE [JERN MURDOCK 1]  NORTON, ANDRE  H:2,P
AMBER CITADEL, THE [JEWELFIRE 1]  WARRINGTON, FREDA  P
OBSIDIAN TOWER, THE [JEWELFIRE 3]  WARRINGTON, FREDA  P
SAPPHIRE THRONE, THE [JEWELFIRE 2]  WARRINGTON, FREDA  P
CHILD OF ELVISH, A [JEWELS OF ELVISH 2]  BERBERICK, NANCY VARIAN  C/P:2,P
JEWELS OF ELVISH, THE [JEWELS OF ELVISH 1]  BERBERICK, NANCY VARIAN  P
HEROING [JIANA 1]  AB HUGH, DAFYDD  C/P,P
WARRIORWARDS [JIANA 2]  AB HUGH, DAFYDD  C/P:3,P
GOBLIN HERO [JIG THE GOBLIN 2]  HINES, JIM C.  P
GOBLINQUEST [JIG THE GOBLIN 1]  HINES, JIM C.  H
MERLIN OF ST. GILLES’ WELL, THE [JOAN OF ARC 1]  CHAMBERLIN, ANN  LP
CAT ON THE EDGE [JOE GREY 1]  MURPHY, SHIRLEY ROUSSEAU  P
CAT RAISE THE DEAD [JOE GREY 3]  MURPHY, SHIRLEY ROUSSEAU  P
CAT UNDER FIRE [JOE GREY 2]  MURPHY, SHIRLEY ROUSSEAU  C/P,P
FORWARD THE MAGE [JOE’S WORLD 2]  FLINT, ERIC/ROACH, RICHARD  C/P,P,H
PHILOSOPHICAL STRANGLER, THE [JOE’S WORLD 1]  FLINT, ERIC  H,P
GHOST OF THE REVELATOR, THE [JOHAN ESBACH 2]  MODESITT, L. E. JR.  H,P
GHOST OF THE WHITE NIGHTS [JOHAN ESBACH 3]  MODESITT, L. E. JR.  P
OF TANGIBLE GHOSTS [JOHAN ESBACH 1]  MODESITT, L. E. JR.  P
DEADLY CYBORGS, THE [JOHN EAGLE EXPEDITER]  EDWARDS, PAUL  P
AFTER DARK [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  C/P,SFWA,P,H
HANGING STONES, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  C/P,SFWA,P,H
JOHN THE BALLADEER [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  P,SFWA-P
LOST AND THE LURKING, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  H:2,SFWA,P
OLD GODS WAKEN, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  C/P,SFWA,P,H
VOICE OF THE MOUNTAIN, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  H,SFWA
WHO FEARS THE DEVIL? [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  SR-P,C/P
SCHOOL OF DARKNESS, THE [JOHN THUNSTONE]  WELLMAN, MANLY WADE  H,SFWA
WHAT DREAMS MAY COME [JOHN THUNSTONE]  WELLMAN, MANLY WADE  H,SFWA
JOHNNY AND THE DEAD [JOHNNY MAXWELL 2]  PRATCHETT, TERRY  H
JOHNNY MAXWELL TRILOGY, THE [JOHNNY MAXWELL 1,2,3]  PRATCHETT, TERRY  H
ONLY YOU CAN SAVE MANKIND [JOHNNY MAXWELL 1]  PRATCHETT, TERRY  H
DANCE OF DEATH [JONATHAN BARRETT 4]  ELROD, P. N.  C/P,P
DEATH AND THE MAIDEN [JONATHAN BARRETT 2]  ELROD, P. N.  C/P:2,P
DEATH MASQUE [JONATHAN BARRETT 3]  ELROD, P. N.  C/P,P
RED DEATH [JONATHAN BARRETT 1]  ELROD, P. N.  P,SFWA-P:2
COLLECTED JORKENS, THE, VOLUME TWO [JORKENS]  DUNSANY, LORD  H
COLLECTED JORKENS, THE, VOLUME THREE [JORKENS]  DUNSANY, LORD  H
FOURTH BOOK OF JORKENS, THE [JORKENS]  DUNSANY, LORD  H
CAVERNS [@JOURNEYS OF MCGILL FEIGHAN #1]  O’DONNELL, KEVIN JR.  C/P:2,P

TITLE  AUTHOR  CODE
REEFS [@JOURNEYS OF MCGILL FEIGHAN #2]  O’DONNELL, KEVIN JR.  C/P,P
LAVA [@JOURNEYS OF MCGILL FEIGHAN #3]  O’DONNELL, KEVIN JR.  C/P:2,P
CLIFFS [@JOURNEYS OF MCGILL FEIGHAN #4]  O’DONNELL, KEVIN JR.  C/P,P
CARNIVORES OF LIGHT AND DARKNESS [JOURNEYS OF THE CATECHIST 1] FOSTER, ALAN DEAN  H
INTO THE THINKING KINGDOMS [JOURNEYS OF THE CATECHIST 2]  FOSTER, ALAN DEAN  P
TRIUMPH OF SOULS, A [JOURNEYS OF THE CATECHIST 3]  FOSTER, ALAN DEAN  H
BAD MOON RISING [@JUDGE DREDD]  BISHOP, DAVID  P
BLACK ATLANTIC [@JUDGE DREDD]  JOWETT, SIMON/EVANS, PETER J.  P
DEATHMASQUES [JUDGE DREDD]  STONE, DAVE  P
FINAL CUT, THE [@JUDGE DREDD]  SMITH, MATTHEW  P
JUDGE DREDD [JUDGE DREDD]  BARRETT, NEAL JR.  P
KINGDOM OF THE BLIND [@JUDGE DREDD]  BISHOP, DAVID  P
SAVAGE AMUSEMENT, THE [JUDGE DREDD]  BISHOP, DAVID  P
ANGEL OF DESTRUCTION [JUDICIARY]  MATTHEWS, SUSAN R.  P
DEVIL AND DEEP SPACE, THE [JUDICIARY 4]  MATTHEWS, SUSAN R.  P
EXCHANGE OF HOSTAGES, AN [JUDICIARY 1]  MATTHEWS, SUSAN R.  C/P,P
HOUR OF JUDGEMENT [JUDICIARY 3]  MATTHEWS, SUSAN R.  C/P,P
PRISONER OF CONSCIENCE [JUDICIARY 2]  MATTHEWS, SUSAN R.  C/P,P
INFOQUAKE [JUMP 225 1]  EDELMAN, DAVID LOUIS  LP
JUMPER [JUMPER 1]  GOULD, STEVEN C.  H:2,P
REFLEX [JUMPER 2]  GOULD, STEVEN C.  H
KOP [JUNO 1]  HAMMOND, WARREN  H
JURASSIC PARK [JURASSIC PARK 1]  CRICHTON, MICHAEL  H:2,P
LOST WORLD, THE [JURASSIC PARK 2]  CRICHTON, MICHAEL  H:2,P
DUSTLAND [JUSTICE 2]  HAMILTON, VIRGINIA  C/P,P
GATHERING, THE [JUSTICE 3]  HAMILTON, VIRGINIA  C/P,P
JUSTICE AND HER BROTHERS [JUSTICE 1]  HAMILTON, VIRGINIA  C/P,P
HERO’S QUEST [@JUSTICE LEAGUE OF AMERICA GREEN LANTERN]  O’NEIL, DENNIS  P
ARTICLE 23 [JUSTIN BELL 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
PROMETHEUS [JUSTIN BELL 3]  FORSTCHEN, WILLIAM R.  C/P,P
STAR VOYAGER ACADEMY [JUSTIN BELL 1]  FORSTCHEN, WILLIAM R.  C/P,P

TITLE  AUTHOR  CODE
CRY WOLF [@K-9 CORPS 3]  VON GUNDEN, KENNETH  P
K-9 CORPS [K-9 CORPS 1]  VON GUNDEN, KENNETH  P
UNDER FIRE [@K-9 CORPS]  VON GUNDEN, KENNETH  C/P,P
K-PAX [K-PAX 1]  BREWER, GENE  P
ON A BEAM OF LIGHT [K-PAX 2]  BREWER, GENE  P
KAI LUNG UNROLLS HIS MAT [KAI LUNG]  BRAMAH, ERNEST  C/P,P
KAI LUNG’S GOLDEN HOURS [KAI LUNG]  BRAMAH, ERNEST  H:2,P
WALLET OF KAI LUNG, THE [KAI LUNG]  BRAMAH, ERNEST  H
SAGA OF LOST EARTHS [KALEVALA 1]  PETAJA, EMIL  C/P:2,P
STAR MILL, THE [KALEVALA 2]  PETAJA, EMIL  C/P:2,P
STOLEN SUN, THE [KALEVALA 3]  PETAJA, EMIL  DG618
TRAMONTANE [KALEVALA 4]  PETAJA, EMIL  DH036
BLACK HOLE OF CARCOSA, THE [KAMUS OF KADIZHAR]  SHIRLEY, JOHN  C/P,P
CIRCLE, CRESCENT, STAR [KANTMORIE 2]  DIBELL, ANSEN  C/P:2,P
PURSUIT OF THE SCREAMER [KANTMORIE 1]  DIBELL, ANSEN  C/P:2,P
SUMMERFAIR [KANTMORIE 3]  DIBELL, ANSEN  P
GOLDEN GRYPHON FEATHER, THE [KAPHTU 1]  PURTILL, RICHARD  C/P:2,P
MIRROR OF HELEN, THE [KAPHTU 3]  PURTILL, RICHARD  P
STOLEN GODDESS, THE [KAPHTU 2]  PURTILL, RICHARD  C/P:2,P
DEEPWOOD [KARAVANS 2]  ROBERSON, JENNIFER  H
KARAVANS [KARAVANS 1]  ROBERSON, JENNIFER  H
BEHOLD THE MAN [KARL GLOGAUER 1]  MOORCOCK, MICHAEL  C/P:2,P
BREAKFAST IN THE RUINS [KARL GLOGAUER 2]  MOORCOCK, MICHAEL  H,P
DAMSEL UNDER STRESS [KATIE CHANDLER 3]  SWENDSON, SHANNA  LP
ENCHANTED, INC. [KATIE CHANDLER 1]  SWENDSON, SHANNA  LP
ONCE UPON STILETTOS [KATIE CHANDLER 2]  SWENDSON, SHANNA  LP
KEDRIGERN AND THE CHARMING COUPLE [KEDRIGERN]  MORRESSY, JOHN  P
KEDRIGERN IN WONDERLAND [KEDRIGERN 3]  MORRESSY, JOHN  P
QUESTING OF KEDRIGERN, THE [KEDRIGERN 2]  MORRESSY, JOHN  C/P,P
REMEMBRANCE FOR KEDRIGERN, A [KEDRIGERN]  MORRESSY, JOHN  P
VOICE FOR PRINCESS, A [KEDRIGERN 1]  MORRESSY, JOHN  C/P,P
DUDGEON & DRAGONS [@KEDRIGERN CHRONICLES #2]  MORRESSY, JOHN  LP
IN THE SERVICE OF DRAGONS [KEEPER MARTIN 5]  STANEK, ROBERT  LP
IN THE SERVICE OF DRAGONS II [KEEPER MARTIN 6]  STANEK, ROBERT  LP
IN THE SERVICE OF DRAGONS III [KEEPER MARTIN 7]  STANEK, ROBERT  LP
IN THE SERVICE OF DRAGONS IV [KEEPER MARTIN 8]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES, THE [KEEPER MARTIN 1]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES II, THE [KEEPER MARTIN 2]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES III, THE [KEEPER MARTIN 3]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES IV, THE [KEEPER MARTIN 4]  STANEK, ROBERT  LP
LONG HOT SUMMONING [KEEPER’S CHRONICLES 3]  HUFF, TANYA  P
SECOND SUMMONING, THE [KEEPER’S CHRONICLES 2]  HUFF, TANYA  C/P:2,P
SUMMON THE KEEPER [KEEPER’S CHRONICLES 1]  HUFF, TANYA  C/P:2,P
FIRE DUKE, THE [@KEEPERS OF THE HIDDEN WAYS 1]  ROSENBERG, JOEL  P
SILVER STONE, THE [@KEEPERS OF THE HIDDEN WAYS #2]  ROSENBERG, JOEL  C/P,P
CRIMSON SKY, THE [@KEEPERS OF THE HIDDEN WAYS #3]  ROSENBERG, JOEL  P
COPPER CROWN, THE [KELTIAD 1]  KENNEALY-MORRISON, PATRICIA  C/P:2,SFWA,P
DEER’S CRY, THE [KELTIAD]  KENNEALY-MORRISON, PATRICIA  C/P,P,H
SILVER BRANCH, THE [KELTIAD 3]  KENNEALY-MORRISON, PATRICIA  P
THRONE OF SCONE, THE [KELTIAD 2]  KENNEALY-MORRISON, PATRICIA  C/P,P,H
HAWK’S GRAY FEATHER, THE [KELTIAD ARTHUR 1]  KENNEALY-MORRISON, PATRICIA  C/P:2,P
HEDGE OF MIST, THE [KELTIAD ARTHUR 3]  KENNEALY-MORRISON, PATRICIA  H,P
OAK ABOVE THE KINGS, THE [KELTIAD ARTHUR 2]  KENNEALY-MORRISON, PATRICIA  C/P,SFWA,P
BLACKMANTLE [KELTIAD MISC 1]  KENNEALY-MORRISON, PATRICIA  H:2,P
CHIMAERA’S COPPER [KELVIN HACKLEBERRY 3]  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P,P,H
DRAGON’S GOLD [KELVIN HACKLEBERRY 1]  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P:2,SFWA-P,P
ORC’S OPAL [KELVIN HACKLEBERRY 4]  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P:2,P
SERPENT’S SILVER [KELVIN HACKLEBERRY 2]  ANTHONY, PIERS/MARGROFF, ROBERT E.  H:2,P
THREE COMPLETE NOVELS [KELVIN HACKLEBERRY 1,2,3]  ANTHONY, PIERS/MARGROFF, ROBERT E.  H
MAJYK BY ACCIDENT [KENDAR GANGLE 1]  FRIESNER, ESTHER M.  C/P,SFWA-P,P
MAJYK BY DESIGN [KENDAR GANGLE]  FRIESNER, ESTHER M.  C/P,P
MAJYK BY HOOK OR CROOK [KENDAR GANGLE 2]  FRIESNER, ESTHER M.  P,SFWA-P
KENT MONTANA AND THE ONCE AND FUTURE THING [KENT MONTANA 3] FENN, LIONEL  C/P,P

KENT MONTANA AND THE REALLY UGLY THINGS FROM MARS [KENT MONTANA 1] FENN, LIONEL

C/P:2,SFWA-P:2,P

KENT MONTANA AND THE REASONABLY INVISIBLE MAN [KENT MONTANA 2] FENN, LIONEL

C/P,P

MARK OF THE MODERATELY VICIOUS VAMPIRE, THE [KENT MONTANA 4]  FENN, LIONEL  P
668: THE NEIGHBOR OF THE BEAST [KENT MONTANA 5]  FENN, LIONEL  C/P,P
BY THE SWORD [KEROWYN’S TALE]  LACKEY, MERCEDES  C/P:2,P
CRUISER DREAMS [KERRION EMPIRE 2]  MORRIS, JANET  H:2,P
DREAM DANCER [KERRION EMPIRE 1]  MORRIS, JANET  H:2,SFWA,P
EARTH DREAMS [KERRION EMPIRE 3]  MORRIS, JANET  C/P,P,H

The MITSFS Seriesdex  September 24, 2007  Page 92
TITLE  AUTHOR  CODE
DAEMONS, INC. [KEVIN BRADLEY AND EVAN DAVIS 1,2]  BACON-SMITH, CAMILLE  H
EYE OF THE DAEMON [KEVIN BRADLEY AND EVAN DAVIS 1]  BACON-SMITH, CAMILLE  C/P,P
EYES OF THE EMPRESS [KEVIN BRADLEY AND EVAN DAVIS 2]  BACON-SMITH, CAMILLE  P
IN STONE’S GRASP [KEVLA OF ARUKAN 2]  GOLDEN, CHRISTIE  LP
ON FIRE’S WINGS [KEVLA OF ARUKAN 1]  GOLDEN, CHRISTIE  LP
INHERITANCE [@KEYS OF POWER #1]  BROWN, SIMON  P
FIRE AND SWORD [@KEYS OF POWER #2]  BROWN, SIMON  P
SOVEREIGN [@KEYS OF POWER #3]  BROWN, SIMON  P
FLAME KEY, THE [KEYS TO PARADISE 1]  MORAN, DANIEL  C/P,P
KEY OF ICE AND STEEL [KEYS TO PARADISE 3]  VARDEMAN, ROBERT E.  P,SFWA-P
SKELETON LORD’S KEY, THE [KEYS TO PARADISE 2]  VARDEMAN, ROBERT E.  C/P,SFWA-P,P
CHARIOTS OF RA, THE [KEYS TO THE DIMENSIONS 7]  BULMER, KENNETH  D10293
DIAMOND CONTESSA, THE [KEYS TO THE DIMENSIONS 8]  BULMER, KENNETH  P
HUNTERS OF JUNDAGAI, THE [KEYS TO THE DIMENSIONS 6]  BULMER, KENNETH  D68310
KEY TO IRUNIUM, THE [KEYS TO THE DIMENSIONS 2]  BULMER, KENNETH  DH020
KEY TO VENUDINE, THE [KEYS TO THE DIMENSIONS 3]  BULMER, KENNETH  DH065
LAND BEYOND THE MAP [KEYS TO THE DIMENSIONS 1]  BULMER, KENNETH  DM111
SHIPS OF DUROSTORUM, THE [KEYS TO THE DIMENSIONS 5]  BULMER, KENNETH  D76096
WIZARDS OF SENCHURIA, THE [KEYS TO THE DIMENSIONS 4]  BULMER, KENNETH  D12140
MISTER MONDAY [@KEYS TO THE KINGDOM #1]  NIX, GARTH  LP
GRIM TUESDAY [@KEYS TO THE KINGDOM #2]  NIX, GARTH  LP
DROWNED WEDNESDAY [@KEYS TO THE KINGDOM #3]  NIX, GARTH  LP
SIR THURSDAY [@KEYS TO THE KINGDOM #4]  NIX, GARTH  LP
WORLD SPIRITS [KHYREN]  BOUCHER-KAPLAN, ALINE  C/P:2,SFWA-P,P
HARPY’S FLIGHT [KI AND VANDIEN 1]  LINDHOLM, MEGAN  C/P,P
LIMBRETH GATE, THE [KI AND VANDIEN 3]  LINDHOLM, MEGAN  C/P,P
LUCK OF THE WHEELS [KI AND VANDIEN 4]  LINDHOLM, MEGAN  C/P,P
WINDSINGERS, THE [KI AND VANDIEN 2]  LINDHOLM, MEGAN  C/P,P
BAALBAK QUEST, THE [KILLSTAR 1]  KELLY, DAVID J.  SR-P
TOWER OF DESPAIR [KILLSTAR 2]  KELLY, DAVID J.  SR-P,C/P
DAHUT [@KING OF YS 3]  ANDERSON, POUL/ANDERSON, KAREN  C/P,SFWA-P,P
DOG AND THE WOLF, THE [@KING OF YS 4]  ANDERSON, POUL/ANDERSON, KAREN  P,SFWA-P
GALLICENAE [@KING OF YS 2]  ANDERSON, POUL/ANDERSON, KAREN  C/P,SFWA-P,P
KING OF YS, THE, VOLUME I [KING OF YS 1,2]  ANDERSON, POUL/ANDERSON, KAREN  H
KING OF YS, THE, VOLUME II [KING OF YS 3,4]  ANDERSON, POUL/ANDERSON, KAREN  H
ROMA MATER [@KING OF YS 1]  ANDERSON, POUL/ANDERSON, KAREN  C/P:2,P
HOOD [KING RAVEN 1]  LAWHEAD, STEPHEN R.  H
FIRE IN HIS HANDS, THE [KING WITHOUT A THRONE 1/RAGNARSON 1]  COOK, GLEN  C/P,P

WITH MERCY TOWARD NONE [KING WITHOUT A THRONE 2/ RAGNARSON 2] COOK, GLEN

C/P:2,P

WRATH OF ASHAR [@KINGDOMS #1]  WELLS, ANGUS  C/P:2,P
USURPER, THE [@KINGDOMS #2]  WELLS, ANGUS  C/P:2,P
WAY BENEATH, THE [@KINGDOMS #3]  WELLS, ANGUS  C/P,P
BLOOD KNIGHT, THE [KINGDOMS OF THORN AND BONE 3]  KEYES, J. GREGORY  H
BRIAR KING, THE [KINGDOMS OF THORN AND BONE 1]  KEYES, J. GREGORY  H
CHARNEL PRINCE, THE [KINGDOMS OF THORN AND BONE 2]  KEYES, J. GREGORY  H
NAME OF THE WIND, THE [KINGKILLER 1]  ROTHFUSS, PATRICK  H
GILDED CHAIN, THE [KING’S BLADES]  DUNCAN, DAVE  H,P
IMPOSSIBLE ODDS [KING’S BLADES]  DUNCAN, DAVE  H,P
JAGUAR KNIGHTS, THE [KING’S BLADES]  DUNCAN, DAVE  H
LORD OF THE FIRE LANDS [KING’S BLADES]  DUNCAN, DAVE  P
PARAGON LOST [@KING’S BLADES]  DUNCAN, DAVE  P
SKY OF SWORDS [KING’S BLADES]  DUNCAN, DAVE  H
SIR STALWART [KING’S DAGGERS 1]  DUNCAN, DAVE  P
FLOATING CASTLE, THE [@KING’S QUEST 1]  MILLS, CRAIG  P
KINGDOM OF SORROW [@KING’S QUEST 2]  MORR, KENYON  P
SEE NO WEEVIL [@KING’S QUEST 3]  MORR, KENYON  P
LAST HUMAN, THE (PROOF) [KING’S TRAMP 3]  DE HAVEN, TOM  SR-LP
WALKER OF WORLDS [@KING’S TRAMP #1]  DE HAVEN, TOM  C/P:2,LP
END-OF-EVERYTHING MAN, THE [@KING’S TRAMP #2]  DE HAVEN, TOM  C/P,P,H
LAST HUMAN, THE [@KING’S TRAMP #3]  DE HAVEN, TOM  C/P,LP
HAWK OF THE WILDERNESS [KIOGA 1]  CHESTER, WILLIAM L.  P
KIOGA OF THE UNKNOWN LAND [KIOGA 4]  CHESTER, WILLIAM L.  C/P,P
KIOGA OF THE WILDERNESS [KIOGA 2]  CHESTER, WILLIAM L.  P
ONE AGAINST A WILDERNESS [KIOGA 3]  CHESTER, WILLIAM L.  P
KITTY AND THE MIDNIGHT HOUR [KITTY NORVILLE 1]  VAUGHN, CARRIE  P
KITTY GOES TO WASHINGTON [KITTY NORVILLE 2]  VAUGHN, CARRIE  P
KITTY TAKES A HOLIDAY [KITTY NORVILLE 3]  VAUGHN, CARRIE  P
AT SWORD’S POINT [@KNIGHTS OF THE BLOOD]  MACMILLAN, SCOTT  C/P,P
VAMPYR-SS [@KNIGHTS OF THE BLOOD 1]  MACMILLAN, SCOTT  C/P,P
DRAGONS REVEALED [KNIGHTS OF THE DRAGONS REVELATIONS 2]  FORBECK, MATT  LP

PROPHECY OF THE DRAGONS [KNIGHTS OF THE DRAGONS REVELATIONS 1] FORBECK, MATT LP

The MITSFS Seriesdex  September 24, 2007  Page 93
TITLE  AUTHOR  CODE

SECRET OF THE SPIRITKEEPER, THE [@KNIGHTS OF THE SILVER DRAGON #1] FORBECK, MATT

LP

RIDDLE IN THE STONE [@KNIGHTS OF THE SILVER DRAGON #2]  SOESBEE, REE  LP
SIGN OF THE SHAPESHIFTER [@KNIGHTS OF THE SILVER DRAGON #3]  DONOVAN, DALE/JOHNS, LINDA  LP
EYE OF FORTUNE [@KNIGHTS OF THE SILVER DRAGON #4]  GRAHAM, DENISE B.  LP
FIGURE IN THE FROST [@KNIGHTS OF THE SILVER DRAGON #5]  PEREZ, LANA  LP
DAGGER OF DOOM, THE [@KNIGHTS OF THE SILVER DRAGON #6]  ROBERTS, KERRY DANIEL  LP
HIDDEN DRAGON, THE [@KNIGHTS OF THE SILVER DRAGON #7]  TRUMBAUER, LISA TRUTKOFF  LP
SILVER SPELL, THE [@KNIGHTS OF THE SILVER DRAGON #8]  BANERJEE, ANJALI  LP
KEY TO THE GRIFFON’S LAIR [@KNIGHTS OF THE SILVER DRAGON #9]  RANSOM, CANDICE  LP
CURSE OF THE LOST GROVE [@KNIGHTS OF THE SILVER DRAGON #10]  GRAHAM, DENISE B.  LP

MYSTERY OF THE WIZARD’S TOMB [@KNIGHTS OF THE SILVER DRAGON #11] PLUMMER, RACHEL

LP

MARK OF THE YUAN-TI [@KNIGHTS OF THE SILVER DRAGON #12]  ROBERTS, KERRY DANIELS  LP
CRASHLANDER [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
DARKER GEOMETRY, A [KNOWN SPACE]  MARTIN, MARK O./BENFORD, GREGORY  P
DOWN IN FLAMES (OUTLINE) [KNOWN SPACE]  NIVEN, LARRY  [C/RAN-H:3]
FLATLANDER [KNOWN SPACE]  NIVEN, LARRY  P
GIFT FROM EARTH, A [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
LONG ARM OF GIL HAMILTON, THE [KNOWN SPACE]  NIVEN, LARRY  SR-P,C/P:2
NEUTRON STAR [KNOWN SPACE]  NIVEN, LARRY  C/P:2,SR-P
PATCHWORK GIRL, THE [KNOWN SPACE]  NIVEN, LARRY  C/P,LP,H
PROTECTOR [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
TALES OF KNOWN SPACE [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
THREE BOOKS OF KNOWN SPACE [KNOWN SPACE]  NIVEN, LARRY  LP
WORLD OF PTAVVS [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
RINGWORLD (ORIGINAL) [KNOWN SPACE RINGWORLD 1]  NIVEN, LARRY  SR-P
RINGWORLD (REVISED) [KNOWN SPACE RINGWORLD 1]  NIVEN, LARRY  C/P,P
RINGWORLD ENGINEERS, THE [KNOWN SPACE RINGWORLD 2]  NIVEN, LARRY  H:2,P
RINGWORLD THRONE, THE [KNOWN SPACE RINGWORLD 3]  NIVEN, LARRY  C/P:2,P
RINGWORLD’S CHILDREN [KNOWN SPACE RINGWORLD 4]  NIVEN, LARRY  H:2
DERVISH IS DIGITAL [KONSTANTIN 2]  CADIGAN, PAT  H
TEA FROM AN EMPTY CUP [KONSTANTIN 1]  CADIGAN, PAT  H,P
TALYN [KORRE]  LISLE, HOLLY  H,P
KOTHAR AND THE CONJURER’S CURSE [KOTHAR 4]  FOX, GARDNER F.  C/P,P
KOTHAR AND THE DEMON QUEEN [KOTHAR 3]  FOX, GARDNER F.  P
KOTHAR AND THE WIZARD SLAYER [KOTHAR 5]  FOX, GARDNER F.  C/P,P
KOTHAR — BARBARIAN SWORDSMAN [KOTHAR 1]  FOX, GARDNER F.  P
KOTHAR OF THE MAGIC SWORD [KOTHAR 2]  FOX, GARDNER F.  P
CITY OF A MILLION LEGENDS [KREN 2]  LICHTENBERG, JACQUELINE  C/P:2,P
MOLT BROTHER [KREN 1]  LICHTENBERG, JACQUELINE  C/P,P
KRIS LONGKNIFE: DEFIANT [KRIS LONGKNIFE 3]  SHEPHERD, MIKE  P
KRIS LONGKNIFE: DESERTER [KRIS LONGKNIFE 2]  SHEPHERD, MIKE  P
KRIS LONGKNIFE: MUTINEER [KRIS LONGKNIFE 1]  SHEPHERD, MIKE  P
KRIS LONGKNIFE: RESOLUTE [KRIS LONGKNIFE 4]  SHEPHERD, MIKE  P

BONES OF ZORA, THE [KRISHNA] DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

H,P

HAND OF ZEI, THE [KRISHNA]  DE CAMP, L. SPRAGUE  C/P,P,H,DF249
HOSTAGE OF ZIR, THE [KRISHNA]  DE CAMP, L. SPRAGUE  H:2,P
HOSTAGE OF ZIR, THE (PROOF) [KRISHNA]  DE CAMP, L. SPRAGUE  SR-LP
PRISONER OF ZHAMANAK, THE [KRISHNA]  DE CAMP, L. SPRAGUE  H,P
QUEEN OF ZAMBA, THE [KRISHNA]  DE CAMP, L. SPRAGUE  C/P:2,P
SEARCH FOR ZEI, THE [KRISHNA]  DE CAMP, L. SPRAGUE  DF249,H

SWORDS OF ZINJABAN, THE [KRISHNA] DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

P

TOWER OF ZANID, THE [KRISHNA]  DE CAMP, L. SPRAGUE  C/P:2,P
KRISPOS OF VIDESSOS [KRISPOS 2]  TURTLEDOVE, HARRY  C/P:2,P
KRISPOS RISING [KRISPOS 1]  TURTLEDOVE, HARRY  C/P:2,P
KRISPOS THE EMPEROR [KRISPOS 3]  TURTLEDOVE, HARRY  P
KING KULL [KULL]  HOWARD, ROBERT E./CARTER, LIN  C/P,P
KULL [KULL]  HOWARD, ROBERT E.  C/P,P
KULL: EXILE OF ATLANTIS [KULL]  HOWARD, ROBERT E.  H
DOUBLE BLIND [KURTIS:P385XL47:RUTH 3]  STAMEY, SARA  P
WILD CARD RUN [KURTIS:P385XL47:RUTH 1]  STAMEY, SARA  P,SFWA-P
WIN, LOSE, DRAW [KURTIS:P385XL47:RUTH 2]  STAMEY, SARA  P
KUSHIEL’S AVATAR [KUSHIEL 3]  CAREY, JACQUELINE  H,P
KUSHIEL’S CHOSEN [KUSHIEL 2]  CAREY, JACQUELINE  H,P
KUSHIEL’S DART [KUSHIEL 1]  CAREY, JACQUELINE  H,P
KUSHIEL’S JUSTICE [KUSHIEL 5]  CAREY, JACQUELINE  H
KUSHIEL’S SCION [KUSHIEL 4]  CAREY, JACQUELINE  H
DAUGHTER OF INDEPENDENCE [KYDAN 3]  BROWN, SIMON  P
EMPIRE’S DAUGHTER [KYDAN 1]  BROWN, SIMON  C/P,P
RIVAL’S SON [KYDAN 2]  BROWN, SIMON  P

TITLE  AUTHOR  CODE
KYRIK AND THE LOST QUEEN [KYRIK 4]  FOX, GARDNER F.  P
KYRIK AND THE WIZARD’S SWORD [KYRIK 3]  FOX, GARDNER F.  P
KYRIK FIGHTS THE DEMON WORLD [KYRIK 2]  FOX, GARDNER F.  P
KYRIK: WARLOCK WARRIOR [KYRIK 1]  FOX, GARDNER F.  P

The MITSFS Seriesdex  September 24, 2007  Page 95
TITLE  AUTHOR  CODE
DEATH OF THE RAVEN [LADY BRANWEN 2]  MEANEY, DEE MORRISON  P
UNKINDNESS OF RAVENS, AN [LADY BRANWEN 1]  MEANEY, DEE MORRISON  P
LADY IN GIL [LADY IN GIL 1]  BRADLEY, REBECCA  C/P,P
LADY PAIN [LADY IN GIL 3]  BRADLEY, REBECCA  P
SCION’S LADY [LADY IN GIL 2]  BRADLEY, REBECCA  P
GALAXY BUILDER, THE [LAFAYETTE O’LEARY 4]  LAUMER, KEITH  C/P:2,P
SHAPE CHANGER, THE [LAFAYETTE O’LEARY 3]  LAUMER, KEITH  C/P,P,H
TIME BENDER, THE [LAFAYETTE O’LEARY 1]  LAUMER, KEITH  C/P:2,P
WORLD SHUFFLER, THE [LAFAYETTE O’LEARY 2]  LAUMER, KEITH  H:2,P
SHADOW AND THE NIGHT, THE [@LAMB AMONG THE STARS #1]  WALLEY, CHRIS  H
DARK FOUNDATIONS, THE [@LAMB AMONG THE STARS #2]  WALLEY, CHRIS  H
DRAGON WAR [LAND BETWEEN THE RIVERS 2]  ACRES, MARK  P
DRAGONSPAWN [LAND BETWEEN THE RIVERS 1]  ACRES, MARK  P
ARMLET OF THE GODS, THE [LAND BEYOND THE GATE 2]  ESHBACH, LLOYD ARTHUR  P
LAND BEYOND THE GATE, THE [LAND BEYOND THE GATE 1]  ESHBACH, LLOYD ARTHUR  C/P,P
SCROLL OF LUCIFER, THE [LAND BEYOND THE GATE 4]  ESHBACH, LLOYD ARTHUR  C/P:2,P
SORCERESS OF SCATH, THE [LAND BEYOND THE GATE 3]  ESHBACH, LLOYD ARTHUR  P
FLIGHT OF FEAR [LAND OF THE GIANTS]  RATHJEN, CARL HENRY  H
HOT SPOT, THE [LAND OF THE GIANTS 2]  LEINSTER, MURRAY  P
LAND OF THE GIANTS [LAND OF THE GIANTS 1]  LEINSTER, MURRAY  C/P:2,P
UNKNOWN DANGER [LAND OF THE GIANTS 3]  LEINSTER, MURRAY  P
SURPRISE GUESTS, THE [LAND OF THE LOST]  GRAHAM, KENNON  H
LAND THAT TIME FORGOT, THE [LAND THAT TIME FORGOT 1]  BURROUGHS, EDGAR RICE  C/P:2,P

LAND THAT TIME FORGOT, THE (OMNIBUS EDITION)

[LAND THAT TIME FORGOT 1,2,3] BURROUGHS, EDGAR RICE

H:2

OUT OF TIME’S ABYSS [LAND THAT TIME FORGOT 3]  BURROUGHS, EDGAR RICE  C/P:2,P
PEOPLE THAT TIME FORGOT, THE [LAND THAT TIME FORGOT 2]  BURROUGHS, EDGAR RICE  C/P:2,P
BLACK UNICORN, THE [@LANDOVER 2]  BROOKS, TERRY  C/P,P,H:2
MAGIC KINGDOM FOR SALE — SOLD! [LANDOVER 1]  BROOKS, TERRY  C/P,H:2,P
TANGLE BOX, THE [LANDOVER 4]  BROOKS, TERRY  C/P,P,H
WITCHES’ BREW [LANDOVER 5]  BROOKS, TERRY  H:2,LP
WIZARD AT LARGE [LANDOVER 3]  BROOKS, TERRY  C/P,H:2,P
PRICE OF BLOOD AND HONOR, THE [LANDUC 3]  WILLEY, ELIZABETH  C/P:2,P
SORCERER AND A GENTLEMAN, A [LANDUC 2]  WILLEY, ELIZABETH  C/P:2,SFWA,P
SORCERER AND A GENTLEMAN, A (PROOF) [LANDUC 2]  WILLEY, ELIZABETH  SR-LP
WELL-FAVORED MAN, THE [LANDUC 1]  WILLEY, ELIZABETH  H,P
LESSER KINDRED, THE [LANEN KAELAR 2]  KERNER, ELIZABETH  P
SONG IN THE SILENCE [LANEN KAELAR 1]  KERNER, ELIZABETH  C/P,P
PRISONER WITHIN, THE [LANNAT 2]  MCQUINN, DONALD E.  P
WITH FULL HONORS [LANNAT 1]  MCQUINN, DONALD E.  P
MAN OF BRONZE, THE [@LARA CROFT TOMB RAIDER]  GARDNER, JAMES ALAN  P
TALISKER [@LAST CLANSMAN #1]  LAU, MILLER  P
DARK THANE [@LAST CLANSMAN #2]  LAU, MILLER  P
LAST HERALD MAGE, THE [LAST HERALD MAGE 1,2,3]  LACKEY, MERCEDES  H:2
MAGIC’S PAWN [LAST HERALD MAGE 1]  LACKEY, MERCEDES  C/P:2,P
MAGIC’S PRICE [LAST HERALD MAGE 3]  LACKEY, MERCEDES  C/P:2,P
MAGIC’S PROMISE [LAST HERALD MAGE 2]  LACKEY, MERCEDES  C/P:2,P
FIREMASK [LAST LEGION 2]  BUNCH, CHRIS  C/P,P
HOMEFALL [LAST LEGION 4]  BUNCH, CHRIS  C/P:2,P
LAST LEGION, THE [LAST LEGION 1]  BUNCH, CHRIS  C/P,P
STORM FORCE [LAST LEGION 3]  BUNCH, CHRIS  C/P,P
DAY OF THE STARWIND [LAST LEGIONARY 3]  HILL, DOUGLAS  H,P
DEATHWING OVER VEYNAA [LAST LEGIONARY 2]  HILL, DOUGLAS  H,P
GALACTIC WARLORD [LAST LEGIONARY 1]  HILL, DOUGLAS  H:2,P
PLANET OF THE WARLORD [LAST LEGIONARY 4]  HILL, DOUGLAS  H
YOUNG LEGIONARY [LAST LEGIONARY 0]  HILL, DOUGLAS  H
VANISHING, THE [@LAST ON EARTH #1]  KAYE, MARILYN  P
CONVERGENCE, THE [@LAST ON EARTH #2]  KAYE, MARILYN  P
RETURN, THE [@LAST ON EARTH #3]  KAYE, MARILYN  P
VILE VILLAGE, THE [@LAST RANGER #7]  SARGENT, CRAIG  P
CUTTHROAT CANNIBALS, THE [@LAST RANGER #8]  SARGENT, CRAIG  P
DAMNED DISCIPLES, THE [@LAST RANGER #9]  SARGENT, CRAIG  P
BEYOND THE PALE [LAST RUNE 1]  ANTHONY, MARK  P
BLOOD OF MYSTERY [LAST RUNE 4]  ANTHONY, MARK  P
DARK REMAINS, THE [LAST RUNE 3]  ANTHONY, MARK  P
FIRST STONE, THE [LAST RUNE 6]  ANTHONY, MARK  P
GATES OF WINTER, THE [LAST RUNE 5]  ANTHONY, MARK  P
GATES OF WINTER, THE (PROOF) [LAST RUNE 5]  ANTHONY, MARK  SR-LP
KEEP OF FIRE, THE [LAST RUNE 2]  ANTHONY, MARK  P
RUNES OF THE EARTH, THE [LAST THOMAS COVENANT 1]  DONALDSON, STEPHEN R.  H
RED DICE [@LAST VAMPIRE #3]  PIKE, CHRISTOPHER  P
GOLDEN HORN, THE [@LAST VIKING #1]  ANDERSON, POUL  P
ROAD OF THE SEA HORSE, THE [@LAST VIKING #2]  ANDERSON, POUL  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 96
TITLE  AUTHOR  CODE
SIGN OF THE RAVEN, THE [@LAST VIKING #3]  ANDERSON, POUL  C/P,P
COMPANIONS [@LAWS OF THE BLOOD 3]  SIZEMORE, SUSAN  C/P,P
DECEPTIONS [@LAWS OF THE BLOOD 4]  SIZEMORE, SUSAN  C/P:2,P
HEROES [@LAWS OF THE BLOOD 5]  SIZEMORE, SUSAN  P
HUNT, THE [@LAWS OF THE BLOOD 1]  SIZEMORE, SUSAN  C/P,P
PARTNERS [@LAWS OF THE BLOOD 2]  SIZEMORE, SUSAN  C/P,P
DEAD MAN ON THE MOON [LCSI]  HARPER, STEVEN  LP
ASCENDING [LEAGUE OF PEOPLES 4]  GARDNER, JAMES ALAN  H,P
EXPENDABLE [LEAGUE OF PEOPLES 1]  GARDNER, JAMES ALAN  H:3,P
HUNTED [LEAGUE OF PEOPLES 3]  GARDNER, JAMES ALAN  H:2,P
RADIANT [LEAGUE OF PEOPLES 6]  GARDNER, JAMES ALAN  H
TRAPPED [LEAGUE OF PEOPLES 5]  GARDNER, JAMES ALAN  H,P
VIGILANT [LEAGUE OF PEOPLES 2]  GARDNER, JAMES ALAN  C/P:2,P

REIGN OF FIRE [LEAR’S DAUGHTERS 2] KELLOGG, MARJORIE BRADLEY/

ROSSOW, WILLIAM B.

P

WAVE AND THE FLAME, THE [LEAR’S DAUGHTERS 1] KELLOGG, MARJORIE BRADLEY/

ROSSOW, WILLIAM B.

P

LEFT BEHIND [@LEFT BEHIND #1]  LAHAYE, TIM/JENKINS, JERRY B.  H:2,LP
TRIBULATION FORCE [@LEFT BEHIND #2]  LAHAYE, TIM/JENKINS, JERRY B.  LP
NICOLAE [@LEFT BEHIND #3]  LAHAYE, TIM/JENKINS, JERRY B.  LP
SOUL HARVEST [@LEFT BEHIND #4]  LAHAYE, TIM/JENKINS, JERRY B.  LP
STAR CORPS [@LEGACY #1]  DOUGLAS, IAN  P
BATTLESPACE [@LEGACY #2]  DOUGLAS, IAN  P
STAR MARINES [@LEGACY #3]  DOUGLAS, IAN  P
ECLIPSE OF THE KAI [@LEGENDS OF LONE WOLF #1]  DEVER, JOE/GRANT, JOHN  P
DARK DOOR OPENS, THE [@LEGENDS OF LONE WOLF #2]  DEVER, JOE/GRANT, JOHN  P
TIDES OF TREACHERY, THE [@LEGENDS OF LONE WOLF #3]  DEVER, JOE/GRANT, JOHN  P
DEMONSTORM [LEGENDS OF THE RAVEN 3]  BARCLAY, JAMES  P
ELFSORROW [LEGENDS OF THE RAVEN 1]  BARCLAY, JAMES  LP
SHADOWHEART [LEGENDS OF THE RAVEN 2]  BARCLAY, JAMES  LP

HONORED ENEMY [@LEGENDS OF THE RIFTWAR #1/RIFTWAR] FEIST, RAYMOND E./

FORSTCHEN, WILLIAM R.

H

BY BLOOD ALONE [LEGION OF THE DAMNED 3]  DIETZ, WILLIAM C.  C/P,P
BY FORCE OF ARMS [LEGION OF THE DAMNED 4]  DIETZ, WILLIAM C.  P
FINAL BATTLE, THE [LEGION OF THE DAMNED 2]  DIETZ, WILLIAM C.  C/P,P
FOR MORE THAN GLORY [LEGION OF THE DAMNED 5]  DIETZ, WILLIAM C.  H,P
FOR THOSE WHO FELL [LEGION OF THE DAMNED 6]  DIETZ, WILLIAM C.  H
LEGION OF THE DAMNED [LEGION OF THE DAMNED 1]  DIETZ, WILLIAM C.  C/P:2,SFWA-P,P
CANTICLE FOR LEIBOWITZ, A [LEIBOWITZ 1]  MILLER, WALTER M. JR.  H:2,LP
SAINT LEIBOWITZ AND THE WILD HORSE WOMAN [LEIBOWITZ 2]  MILLER, WALTER M. JR.  H:2,LP
WAITERS ON THE DANCE [LEMMUS 1]  SAVARIN, JULIAN JAY  H
BEYOND THE OUTER MIRR [LEMMUS #2]  SAVARIN, JULIAN JAY  @SR-P,H
ARCHIVES OF HAVEN, THE [LEMMUS #3]  SAVARIN, JULIAN JAY  @SR-P,H
KING OF THE DEAD [LENS OF THE WORLD 2]  MACAVOY, R. A.  H:2
LENS OF THE WORLD [LENS OF THE WORLD #1]  MACAVOY, R. A.  H:2,@P
BELLY OF THE WOLF, THE [@LENS OF THE WORLD #3]  MACAVOY, R. A.  C/P,P
CHILDREN OF THE LENS [LENSMEN 6]  SMITH, E. E.  C/P,P,H
DRAGON LENSMAN, THE [LENSMEN 7]  KYLE, DAVID A.  C/P:2,P
FIRST LENSMAN [LENSMEN 2]  SMITH, E. E.  H:2,P
GALACTIC PATROL [LENSMEN 3]  SMITH, E. E.  C/P,P,H
GRAY LENSMAN [LENSMEN 4]  SMITH, E. E.  H:2,P
LENSMAN FROM RIGEL [LENSMEN 8]  KYLE, DAVID A.  C/P:2,P
SECOND STAGE LENSMEN [LENSMEN 5]  SMITH, E. E.  C/P,P,H
TRIPLANETARY [LENSMEN 1]  SMITH, E. E.  H:2,P
Z-LENSMAN [LENSMEN 9]  KYLE, DAVID A.  C/P:2,P
DAY OF THE KLESH, THE [LER 3]  FOSTER, M. A.  C/P:2,P
GAMEPLAYERS OF ZAN, THE [LER 1]  FOSTER, M. A.  C/P:2,P
WARRIORS OF DAWN, THE [LER 2]  FOSTER, M. A.  C/P:2,P
BLACKWOOD FARM [LESTAT]  RICE, ANNE  H:2
BLOOD AND GOLD [LESTAT]  RICE, ANNE  H
BLOOD CANTICLE [LESTAT]  RICE, ANNE  H:2
INTERVIEW WITH THE VAMPIRE [LESTAT 1]  RICE, ANNE  C/P:2,P
MEMNOCH THE DEVIL [LESTAT 5]  RICE, ANNE  C/P,P,H
QUEEN OF THE DAMNED, THE [LESTAT 3]  RICE, ANNE  H:2,P
TALE OF THE BODY THIEF, THE [LESTAT 4]  RICE, ANNE  H:2,P
VAMPIRE ARMAND, THE [LESTAT 6]  RICE, ANNE  H,LP
VAMPIRE LESTAT, THE [LESTAT 2]  RICE, ANNE  C/P,P,H
VAMPIRE LESTAT, THE (GRAPHIC NOVEL) [LESTAT 2]  RICE, ANNE/PEROZICH, FAYE  CX

ADVENTURES IN THE LIADEN UNIVERSE 1-5 [LIADEN UNIVERSE 0/ LIADEN ADVENTURES 1,2,3,4,5] LEE, SHARON/MILLER, STEVE

SR-H

AGENT OF CHANGE [LIADEN UNIVERSE 1]  LEE, SHARON/MILLER, STEVE  P
BALANCE OF TRADE [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  H
CALAMITY’S CHILD [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  SR-LP

The MITSFS Seriesdex  September 24, 2007  Page 97
TITLE  AUTHOR  CODE
CARPE DIEM [LIADEN UNIVERSE 3]  LEE, SHARON/MILLER, STEVE  C/P,P
CHANGELING [LIADEN UNIVERSE 0/LIADEN ADVENTURES 6]  LEE, SHARON/MILLER, STEVE  SR-LP
CONFLICT OF HONORS [LIADEN UNIVERSE 2]  LEE, SHARON/MILLER, STEVE  C/P,P
I DARE [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  LP
LIADEN UNIVERSE COMPANION: VOLUME I [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  LP
LOCAL CUSTOM [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  P
LOOSE CANNON [LIADEN UNIVERSE 0/LIADEN ADVENTURES 7]  LEE, SHARON/MILLER, STEVE  SR-LP
PARTNERS IN NECESSITY [LIADEN UNIVERSE 1,2,3]  LEE, SHARON/MILLER, STEVE  H
PILOTS CHOICE [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  H,LP
PLAN B [LIADEN UNIVERSE 4]  LEE, SHARON/MILLER, STEVE  H:2,LP
QUIET KNIVES [LIADEN UNIVERSE 0/LIADEN ADVENTURES 9]  LEE, SHARON/MILLER, STEVE  SR-LP
SCOUT’S PROGRESS [@LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  P
SHADOWS AND SHADES [LIADEN UNIVERSE 0/LIADEN ADVENTURES 8]  LEE, SHARON/MILLER, STEVE  SR-LP
CRYSTAL DRAGON [LIADEN UNIVERSE GREAT MIGRATION 2]  LEE, SHARON/MILLER, STEVE  H
CRYSTAL SOLDIER [LIADEN UNIVERSE GREAT MIGRATION 1]  LEE, SHARON/MILLER, STEVE  H
CASTING FORTUNE [LIAVEK 6]  FORD, JOHN M.  H,SFWA-P
LIAVEK [LIAVEK 1]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIAVEK: FESTIVAL WEEK [LIAVEK 5]  SHETTERLY, WILL/BULL, EMMA  PA
LIAVEK: PLAYERS OF LUCK, THE [LIAVEK 2]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIAVEK: SPELLS OF BINDING [LIAVEK 4]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIAVEK: WIZARD’S ROW [LIAVEK 3]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIFE PROBE [LIFE PROBE 1]  MCCOLLUM, MICHAEL  C/P:2,P
PROCYON’S PROMISE [LIFE PROBE 2]  MCCOLLUM, MICHAEL  C/P:2,P
LIGHT [LIGHT 1]  HARRISON, M. JOHN  H
NOVA SWING [LIGHT 2]  HARRISON, M. JOHN  H
FORGING OF THE SHADOWS, THE [@LIGHTBRINGER #1]  JOHNSON, OLIVER  H,LP
NATIONS OF THE NIGHT, THE [@LIGHTBRINGER #2]  JOHNSON, OLIVER  H,P
LAST STAR AT DAWN, THE [@LIGHTBRINGER #3]  JOHNSON, OLIVER  H,P
LIGHTSTONE, THE [LIGHTSTONE 1]  ZINDELL, DAVID  H
SILVER SWORD, THE [LIGHTSTONE 2]  ZINDELL, DAVID  H
CAVES OF STEEL, THE [LIJE BALEY 1]  ASIMOV, ISAAC  C/P:2,P,H
NAKED SUN, THE [LIJE BALEY 2]  ASIMOV, ISAAC  C/P:3,P,H
ROBOTS OF DAWN, THE [LIJE BALEY 3]  ASIMOV, ISAAC  C/P:2,P,H
BLOOD LINES [LILY YU 3]  WILKS, EILEEN  P
TEMPTING DANGER [LILY YU 1]  WILKS, EILEEN  P
FANGS OF THE HOODED DEMON, THE [LINCOLN BLACKTHORNE 4]  MARSH, GEOFFREY  SFWA-P
KING OF SATAN’S EYES, THE [LINCOLN BLACKTHORNE 1]  MARSH, GEOFFREY  H,SFWA,SFWA-P
PATCH OF THE ODIN SOLDIER, THE [LINCOLN BLACKTHORNE 3]  MARSH, GEOFFREY  H,SFWA,SFWA-P
TAIL OF THE ARABIAN, KNIGHT, THE [LINCOLN BLACKTHORNE 2]  MARSH, GEOFFREY  P,SFWA,SFWA-P
WANDERER’S TALE, THE [LINDORMYN 1]  BILSBOROUGH, DAVID  H
LION’S BLOOD [LION’S BLOOD 1]  BARNES, STEVEN  C/P,P
ZULU HEART [LION’S BLOOD 2]  BARNES, STEVEN  H,P
WOUNDED [LITTLE GODDESS 2]  LANE, AMY  LP
MAD SHIP [LIVESHIP TRADERS 2]  HOBB, ROBIN  C/P,P,H
SHIP OF DESTINY [LIVESHIP TRADERS 3]  HOBB, ROBIN  H
SHIP OF MAGIC [LIVESHIP TRADERS 1]  HOBB, ROBIN  H:2,P
GRANITE SHIELD, THE [LIVING FLAME 3]  PATTON, FIONA  P
PAINTER KNIGHT, THE [LIVING FLAME 2]  PATTON, FIONA  P
STONE PRINCE, THE [LIVING FLAME 1]  PATTON, FIONA  C/P:2,P
CASSILEE [LIVING PLANET 2]  COON, SUSAN  C/P,P
CHIY-UNE [LIVING PLANET 4]  COON, SUSAN  P
RAHNE [LIVING PLANET 1]  COON, SUSAN  P
VIRGIN, THE [LIVING PLANET 3]  COON, SUSAN  C/P,P
TEMPORARY AGENCY [LIVING WORLD 2]  POLLACK, RACHEL  H:2,LP
UNQUENCHABLE FIRE [LIVING WORLD 1]  POLLACK, RACHEL  H,LP
HELVERTI INVASION, THE [LIZARD WAR 2]  DALMAS, JOHN  P
LIZARD WAR, THE [LIZARD WAR 1]  DALMAS, JOHN  C/P:2,P
GUARDIAN OF HONOR [LLADRANA 1]  OWENS, ROBIN D.  LP
PROTECTOR OF THE FLIGHT [LLADRANA 3]  OWENS, ROBIN D.  LP
SORCERESS OF FAITH [LLADRANA 2]  OWENS, ROBIN D.  LP
THIEF OF LLARN [LLARN 2]  FOX, GARDNER F.  C/P:2,P
WARRIOR OF LLARN [LLARN 1]  FOX, GARDNER F.  C/P:2,P

LOGAN [LOGAN 1,2,3] NOLAN, WILLIAM F./

JOHNSON, GEORGE CLAYTON

H,SFWA,LP

LOGAN’S RUN [LOGAN 1] NOLAN, WILLIAM F./

JOHNSON, GEORGE CLAYTON

C/P:2,P

LOGAN’S SEARCH [LOGAN 3]  NOLAN, WILLIAM F.  C/P,P
LOGAN’S WORLD [LOGAN 2]  NOLAN, WILLIAM F.  C/P,P
CALCULATED MAGIC, A [LOGICAL MAGICIAN 2]  WEINBERG, ROBERT  C/P,P
LOGICAL MAGICIAN, A [LOGICAL MAGICIAN 1]  WEINBERG, ROBERT  C/P:2,P
LOIS & CLARK [LOIS & CLARK]  CHERRYH, C. J.  H
CHILDREN OF AMARID [LON TOBYN 1]  COE, DAVID B.  C/P:2,P
EAGLE-SAGE [LON TOBYN 3]  COE, DAVID B.  H,P

TITLE  AUTHOR  CODE
OUTLANDERS, THE [LON TOBYN 2]  COE, DAVID B.  C/P:2,P
THOSE WHO HUNT THE NIGHT [LONDON VAMPIRES 1]  HAMBLY, BARBARA  C/P,P,H:2
TRAVELING WITH THE DEAD [LONDON VAMPIRES 2]  HAMBLY, BARBARA  C/P,P,H
CAPTIVES OF RAAG, THE [LONE WOLF 14]  DEVER, JOE  P
CASTLE DEATH [LONE WOLF 7]  DEVER, JOE  P
CAULDRON OF FEAR, THE [LONE WOLF 9]  DEVER, JOE  P
CAVERNS OF KALTE, THE [LONE WOLF 3]  DEVER, JOE/CHALK, GARY  C/P,P
CHASM OF DOOM, THE [LONE WOLF 4]  DEVER, JOE  P
DARKE CRUSADE, THE [LONE WOLF 15]  DEVER, JOE  P
DAWN OF THE DRAGONS [LONE WOLF]  DEVER, JOE  P
DEATHLORD OF IXIA, THE [LONE WOLF 17]  DEVER, JOE  P
DUNGEONS OF TORGAR, THE [LONE WOLF 10]  DEVER, JOE  P
FIRE ON THE WATER [LONE WOLF 2]  DEVER, JOE  P
FLIGHT FROM THE DARK [LONE WOLF 1]  DEVER, JOE/CHALK, GARY  P
JUNGLE OF HORRORS, THE [LONE WOLF 8]  DEVER, JOE  P
KINGDOMS OF TERROR, THE [LONE WOLF 6]  DEVER, JOE  P
LEGACY OF VASHNA, THE [LONE WOLF 16]  DEVER, JOE  P
MASTERS OF DARKNESS, THE [LONE WOLF 12]  DEVER, JOE  C/P,P
PLAGUE LORDS OF RUEL, THE [LONE WOLF 13]  DEVER, JOE  P
PRISONERS OF TIME, THE [LONE WOLF 11]  DEVER, JOE  P
SHADOW ON THE SAND [LONE WOLF 5]  DEVER, JOE  P
GREY STAR THE WIZARD [@LONE WOLF #1]  PAGE, IAN/DEVER, JOE  P
FORBIDDEN CITY, THE [@LONE WOLF #2]  PAGE, IAN/DEVER, JOE  P
WAR OF THE WIZARDS [@LONE WOLF #4]  PAGE, IAN/DEVER, JOE  P
SHADOW IN SUMMER, A [LONG PRICE 1]  ABRAHAM, DANIEL  H
CALDE OF THE LONG SUN [LONG SUN 3]  WOLFE, GENE  C/P,P,H
EPIPHANY OF THE LONG SUN [LONG SUN 3,4]  WOLFE, GENE  H
EXODUS FROM THE LONG SUN [LONG SUN 4]  WOLFE, GENE  C/P,P,H
LAKE OF THE LONG SUN [LONG SUN 2]  WOLFE, GENE  H,P
LITANY OF THE LONG SUN [LONG SUN 1,2]  WOLFE, GENE  H,LP
NIGHTSIDE THE LONG SUN [LONG SUN 1]  WOLFE, GENE  H,P
DEMON KNIGHT [LONGDIRK 3]  HOOD, KEN  P
DEMON RIDER [LONGDIRK 2]  HOOD, KEN  P
DEMON SWORD [LONGDIRK 1]  HOOD, KEN  P
INTO THE LOOKING GLASS [LOOKING GLASS 1]  RINGO, JOHN  H
LORD DARCY [LORD DARCY 1,2,3]  GARRETT, RANDALL  H
LORD DARCY (BAEN) [LORD DARCY 1,2,3,4]  GARRETT, RANDALL  LP
LORD DARCY INVESTIGATES [LORD DARCY 3]  GARRETT, RANDALL  C/P:2,P
MURDER AND MAGIC [LORD DARCY 1]  GARRETT, RANDALL  C/P:2,P
STUDY IN SORCERY, A [LORD DARCY 6]  KURLAND, MICHAEL  C/P:2,P
TEN LITTLE WIZARDS [LORD DARCY 5]  KURLAND, MICHAEL  C/P:2,SFWA-P,P
TOO MANY MAGICIANS [LORD DARCY 2]  GARRETT, RANDALL  SR-P,C/P:2
GODDESS OF THE ICE REALM [LORD OF THE ISLES 5]  DRAKE, DAVID  H,P
LORD OF THE ISLES [LORD OF THE ISLES 1]  DRAKE, DAVID  C/P:2,P,H
MASTER OF THE CAULDRON [LORD OF THE ISLES 6]  DRAKE, DAVID  P
MISTRESS OF THE CATACOMBS [LORD OF THE ISLES 4]  DRAKE, DAVID  C/P:2,P,H
QUEEN OF DEMONS [LORD OF THE ISLES 2]  DRAKE, DAVID  C/P,P,H:2
SERVANT OF THE DRAGON [LORD OF THE ISLES 3]  DRAKE, DAVID  C/P:3,P,H
ATLAS OF MIDDLE-EARTH, THE [LORD OF THE RINGS]  FONSTAD, KAREN WYNN  R/REF-H
HOBBIT COMPANION, THE [LORD OF THE RINGS]  DAY, DAVID  R/REF-H
LORD OF THE RINGS, THE (12 CASSETTES, BOXED) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
LORD OF THE RINGS, THE (13 CASSETTES, BOXED) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
LORD OF THE RINGS, THE (COLLECTORS EDITION) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-VLH
REALMS OF TOLKIEN [LORD OF THE RINGS]  HARPERPRISM  A
SILMARILLION, VOL. 1, THE (4 CDS) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
SILMARILLION, VOL. 2, THE (4 CDS) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
SILMARILLION, VOL. 3, THE (5 CDS) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
TALE OF THE CHILDREN OF HURIN, THE [LORD OF THE RINGS]  TOLKIEN, J. R. R.  H:2
FELLOWSHIP OF THE RING, THE [@LORD OF THE RINGS #1]  TOLKIEN, J. R. R.  SR-H,C/P,P,H
FELLOWSHIP OF THE RING, THE (ACE) [@LORD OF THE RINGS #1]  TOLKIEN, J. R. R.  SR-P
TWO TOWERS, THE [@LORD OF THE RINGS #2]  TOLKIEN, J. R. R.  H:3,SR-H,P
TWO TOWERS, THE (ACE) [@LORD OF THE RINGS #2]  TOLKIEN, J. R. R.  SR-P
RETURN OF THE KING, THE [@LORD OF THE RINGS #3]  TOLKIEN, J. R. R.  SR-H,H:2,P
RETURN OF THE KING, THE (ACE) [@LORD OF THE RINGS #3]  TOLKIEN, J. R. R.  SR-P
ALIEN REALMS [LORD TEDRIC 4]  SMITH, E. E.  SR-P
LORD TEDRIC [LORD TEDRIC #1]  EKLUND, GORDON  H,@P,@C/P:2
SPACE PIRATES [@LORD TEDRIC #2]  EKLUND, GORDON  P
BLACK KNIGHT OF THE IRON SPHERE [@LORD TEDRIC #3]  EKLUND, GORDON  H,LP,P
KNIGHT OF DARKNESS [LORDS OF AVALON 1]  MACGREGOR, KINLEY  H
SWORD OF DARKNESS [LORDS OF AVALON 2]  MACGREGOR, KINLEY  H
BOOK OF SILENCE, THE [LORDS OF DUS 4]  WATT-EVANS, LAWRENCE  C/P:2,P
LORDS OF DUS, THE [LORDS OF DUS 1,2,3,4]  WATT-EVANS, LAWRENCE  H
LURE OF THE BASILISK, THE [LORDS OF DUS 1]  WATT-EVANS, LAWRENCE  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 99
TITLE  AUTHOR  CODE
SEVEN ALTARS OF DUSARRA, THE [LORDS OF DUS 2]  WATT-EVANS, LAWRENCE  C/P:2,P
SWORD OF BHELEU, THE [LORDS OF DUS 3]  WATT-EVANS, LAWRENCE  C/P:2,P
LORDS OF THE SKY [LORDS OF THE SKY 1]  WELLS, ANGUS  P
TRUTHSAYER’S APPRENTICE, THE [LOREGIVER 1]  CHRISTIAN, DEBORAH  LP
FIRE OPAL, THE [LOST CONTINENT 2]  ASARO, CATHERINE  LP
MISTED CLIFFS, THE [LOST CONTINENT 1]  ASARO, CATHERINE  LP
DAUNTLESS [@LOST FLEET 1]  CAMPBELL, JACK  P
FEARLESS [@LOST FLEET 2]  CAMPBELL, JACK  P
BLAST OFF INTO SPACE [LOST IN SPACE]  PACKER, PETER  SR-S
LOST IN SPACE [LOST IN SPACE]  VAN ARNAM, DAVE/ARCHER, RON  P
LOST IN SPACE [LOST IN SPACE]  VINGE, JOAN D.  P
LOST IN SPACE SCRAPBOOK VOL. TWO, THE [LOST IN SPACE]  ANCHORS, WILLIAM E. JR./MITCHELL, FLINT VLP
PROMISED LAND [LOST IN SPACE]  CADIGAN, PAT  P
VAULT, THE [LOST IN SPACE]  DEWEESE, GENE  P
PASSENGERS, THE [LOST IN SPACE NEW JOURNEYS]  KRULIK, NANCY  LP
BLOOD OF THE TIGER [LOST LANDS 1]  ESTES, ROSE  P
BROTHER TO THE LION [LOST LANDS 2]  ESTES, ROSE  P
SPIRIT OF THE HAWK [LOST LANDS 3]  ESTES, ROSE  P
FIRST DAWN [@LOST MILLENNIUM 1]  MOSCOE, MIKE  C/P,P
LOST DAYS [@LOST MILLENNIUM 3]  MOSCOE, MIKE  C/P,P
SECOND FIRE [@LOST MILLENNIUM 2]  MOSCOE, MIKE  C/P,P
BATTLE HYMN [@LOST REGIMENT 5]  FORSTCHEN, WILLIAM R.  P
DOWN TO THE SEA [@LOST REGIMENT 9]  FORSTCHEN, WILLIAM R.  C/P,P
FATEFUL LIGHTNING [@LOST REGIMENT 4]  FORSTCHEN, WILLIAM R.  P
RALLY CRY [@LOST REGIMENT 1]  FORSTCHEN, WILLIAM R.  H,P
TERRIBLE SWIFT SWORD [@LOST REGIMENT 3]  FORSTCHEN, WILLIAM R.  C/P,P
UNION FOREVER [@LOST REGIMENT 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
NEVER SOUND RETREAT [@LOST REGIMENT #6]  FORSTCHEN, WILLIAM R.  P
BAND OF BROTHERS, A [LOST REGIMENT #7]  FORSTCHEN, WILLIAM R.  P
MEN OF WAR [@LOST REGIMENT #8]  FORSTCHEN, WILLIAM R.  C/P,P
WOUNDHEALER’S STORY [@LOST SWORDS #1]  SABERHAGEN, FRED  H:2,P
FIRST TRIAD, THE [@LOST SWORDS #1,2,3]  SABERHAGEN, FRED  H:2
SIGHTBLINDER’S STORY [@LOST SWORDS #2]  SABERHAGEN, FRED  H:2,P
STONECUTTER’S STORY [@LOST SWORDS #3]  SABERHAGEN, FRED  H:2,SFWA-P,P
FARSLAYER’S STORY [@LOST SWORDS #4]  SABERHAGEN, FRED  H:2,SFWA-P,P
SECOND TRIAD, THE [@LOST SWORDS #4,5,6]  SABERHAGEN, FRED  H:2
COINSPINNER’S STORY [@LOST SWORDS #5]  SABERHAGEN, FRED  C/P,SFWA-P,P,H
MINDSWORD’S STORY [@LOST SWORDS #6]  SABERHAGEN, FRED  C/P,SFWA,P,H
WAYFINDER’S STORY [@LOST SWORDS #7]  SABERHAGEN, FRED  H:2
ENDGAME [@LOST SWORDS #7,8]  SABERHAGEN, FRED  H:2
SHIELDBREAKER’S STORY [@LOST SWORDS #8]  SABERHAGEN, FRED  C/P,P,H
END OF THE THIRD WORLD, THE [LOST WORLD 2]  SOUZA, MARCIO  H,LP
LOST YEARS OF MERLIN, THE [LOST YEARS OF MERLIN 1]  BARRON, T. A.  P
MIRROR OF MERLIN, THE [LOST YEARS OF MERLIN 4]  BARRON, T. A.  P
SEVEN SONGS OF MERLIN, THE [LOST YEARS OF MERLIN 2]  BARRON, T. A.  P
WINGS OF MERLIN [LOST YEARS OF MERLIN 5]  BARRON, T. A.  P
CHALET DIABOLIQUE [LUCIFER COVE 5]  COFFMAN, VIRGINIA  P
DEVIL’S MISTRESS, THE [LUCIFER COVE 1]  COFFMAN, VIRGINIA  P
DEVIL’S VIRGIN, THE [LUCIFER COVE 3]  COFFMAN, VIRGINIA  P
FROM SATAN, WITH LOVE [LUCIFER COVE 6]  COFFMAN, VIRGINIA  P
MASQUE OF SATAN [LUCIFER COVE 4]  COFFMAN, VIRGINIA  P
PRIESTESS OF THE DAMNED [LUCIFER COVE 2]  COFFMAN, VIRGINIA  P
ADVENTURES [LUCIFER JONES 1]  RESNICK, MIKE  P
LUCIFER JONES [LUCIFER JONES 2]  RESNICK, MIKE  C/P,P
JUMP PAY [LUCKY 13TH]  SHELLEY, RICK  P
SIDE SHOW [LUCKY 13TH]  SHELLEY, RICK  P
UNTIL RELIEVED [LUCKY 13TH]  SHELLEY, RICK  P
COMPLETE ADVENTURES OF LUCKY STARR, THE [LUCKY STARR 1,2,3,4,5,6] ASIMOV, ISAAC  H
DAVID STARR, SPACE RANGER [LUCKY STARR 1] ASIMOV, ISAAC C/P:2,P

DAVID STARR, SPACE RANGER [LUCKY STARR 1,2]/ LUCKY STARR AND THE PIRATES OF THE ASTEROIDS [LUCKY STARR 1,2] ASIMOV, ISAAC

P

LUCKY STARR AND THE BIG SUN OF MERCURY [LUCKY STARR 4]  ASIMOV, ISAAC  C/P:2,P
LUCKY STARR AND THE MOONS OF JUPITER [LUCKY STARR 5]  ASIMOV, ISAAC  C/P:2,P

LUCKY STARR AND THE MOONS OF JUPITER [LUCKY STARR 5,6]/ LUCKY STARR AND THE RINGS OF SATURN [LUCKY STARR 5,6] ASIMOV, ISAAC

P

LUCKY STARR AND THE OCEANS OF VENUS [LUCKY STARR 3]  ASIMOV, ISAAC  C/P,P,H

LUCKY STARR AND THE OCEANS OF VENUS [LUCKY STARR 3,4]/ LUCKY STARR AND THE BIG SUN OF MERCURY [LUCKY STARR 3,4] ASIMOV, ISAAC

P

LUCKY STARR AND THE PIRATES OF THE ASTEROIDS [LUCKY STARR 2]  ASIMOV, ISAAC  C/P:2,P
LUCKY STARR AND THE RINGS OF SATURN [LUCKY STARR 6]  ASIMOV, ISAAC  C/P:2,P
PIRATES OF THE ASTEROIDS [LUCKY STARR 2]  ASIMOV, ISAAC  C/P,P
SPACE RANGER [LUCKY STARR 1]  ASIMOV, ISAAC  P
BEYOND THE BURNING LANDS [LUKE 2]  CHRISTOPHER, JOHN  H,P

TITLE  AUTHOR  CODE
PRINCE IN WAITING, THE [LUKE 1]  CHRISTOPHER, JOHN  H:2,P
SWORD OF THE SPIRITS, THE [LUKE 3]  CHRISTOPHER, JOHN  C/P,P,H
GREEN PEARL, THE [LYONESSE 2]  VANCE, JACK  H:2,SFWA-P,LP
MADOUC [LYONESSE 3]  VANCE, JACK  C/P,SFWA-P:2,LP,H
SULDRUN’S GARDEN [LYONESSE 1]  VANCE, JACK  C/P,LP,H
CAUGHT IN CRYSTAL [LYRA 1]  WREDE, PATRICIA C.  C/P:2,SFWA-P,P
DAUGHTER OF WITCHES [LYRA 4]  WREDE, PATRICIA C.  C/P:2,P
HARP OF IMACH THYSSEL, THE [LYRA 5]  WREDE, PATRICIA C.  C/P:2,P
RAVEN RING, THE [LYRA 2]  WREDE, PATRICIA C.  P
SHADOW MAGIC [LYRA 3]  WREDE, PATRICIA C.  C/P:2,P
SHADOWS OVER LYRA [LYRA 1,2,3]  WREDE, PATRICIA C.  H,LP
EMPRESS AND THE ACOLYTE, THE [LYREMOUTH 3]  FLETCHER, JANE  LP
EXILE AND THE SORCERER, THE [LYREMOUTH 1]  FLETCHER, JANE  LP
TRAITOR AND THE CHALICE, THE [LYREMOUTH 2]  FLETCHER, JANE  LP

TITLE  AUTHOR  CODE
CHILDREN OF LLYR, THE [MABINOGI 2]  WALTON, EVANGELINE  C/P:2,P
ISLAND OF THE MIGHTY, THE [MABINOGI 4]  WALTON, EVANGELINE  C/P:2,P
PRINCE OF ANNWN [MABINOGI 1]  WALTON, EVANGELINE  C/P:2,P
SONG OF RHIANNON, THE [MABINOGI 3]  WALTON, EVANGELINE  C/P,SFWA,P
DAY OF THE CHEETAH [MAC MCLANAHAN 2]  BROWN, DALE  H
FLIGHT OF THE OLD DOG [MAC MCLANAHAN 1]  BROWN, DALE  P
DARK PRINCE [MACEDON 2]  GEMMELL, DAVID  LP
LION OF MACEDON [MACEDON 1]  GEMMELL, DAVID  H,LP
BURR IN THE GARDEN OF EDEN, THE [MACNAIR OF MACNAIR 2]  PEIRCE, HAYFORD  LP
NAPOLEON DISENTIMED [MACNAIR OF MACNAIR 1]  PEIRCE, HAYFORD  C/P:2,SFWA-P,P
CHILD OF THE JOURNEY [MADAGASCAR MANIFESTO 2]  BERLINER, JANET/GUTHRIDGE, GEORGE  P
CHILD OF THE LIGHT [MADAGASCAR MANIFESTO 1]  BERLINER, JANET/GUTHRIDGE, GEORGE  P
CHILDREN OF THE DUSK [MADAGASCAR MANIFESTO 3]  BERLINER, JANET/GUTHRIDGE, GEORGE  P
MADAGASCAR MANIFESTO, THE [MADAGASCAR MANIFESTO 1,2,3]  BERLINER, JANET/GUTHRIDGE, GEORGE  LP
ART OF ARROW CUTTING, THE [MAGE AND TAKUMO 1]  DEDMAN, STEPHEN  LP
SHADOWS BITE [MAGE AND TAKUMO 2]  DEDMAN, STEPHEN  H,LP
TOWER OF BABEL [MAGE ASCENSION]  STEELE, JOHN H.  P
STORM BREAKING [MAGE STORMS 3]  LACKEY, MERCEDES  C/P,P,H:2
STORM RISING [MAGE STORMS 2]  LACKEY, MERCEDES  H:2,SFWA,P
STORM WARNING [MAGE STORMS 1]  LACKEY, MERCEDES  H:2,SFWA-P,P
BLACK GRYPHON, THE [MAGE WARS 1]  LACKEY, MERCEDES/DIXON, LARRY  C/P,SFWA,P,H
SILVER GRYPHON, THE [MAGE WARS 3]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
WHITE GRYPHON, THE [MAGE WARS 2]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
WINDS OF CHANGE [MAGE WINDS 2]  LACKEY, MERCEDES  C/P,P,H
WINDS OF FATE [MAGE WINDS 1]  LACKEY, MERCEDES  C/P,P,H
WINDS OF FURY [MAGE WINDS 3]  LACKEY, MERCEDES  H:2,SFWA
AWAKENING, THE [@MAGELORD 1]  MARTIN, THOMAS K.  P
TIME OF MADNESS, THE [@MAGELORD 2]  MARTIN, THOMAS K.  P
BURNING STONE, THE [MAGES OF GARILLON 1]  HARRIS, DEBORAH TURNER  C/P,SFWA,SFWA-P,P
GAUNTLET OF MALICE, THE [MAGES OF GARILLON 2]  HARRIS, DEBORAH TURNER  C/P,SFWA,SFWA-P,LP
SPIRAL OF FIRE [MAGES OF GARILLON 3]  HARRIS, DEBORAH TURNER  C/P,P
MASTER OF THE NIGHT [MAGEVERSE MASTERS 1]  KNIGHT, ANGELA  P
BY HONOR BETRAY’D [MAGEWORLDS 3]  DOYLE, DEBRA/MACDONALD, JAMES D.  P
GATHERING FLAME, THE [MAGEWORLDS 4]  DOYLE, DEBRA/MACDONALD, JAMES D.  P
LONG HUNT, THE [MAGEWORLDS 5]  DOYLE, DEBRA/MACDONALD, JAMES D.  P
PRICE OF THE STARS, THE [MAGEWORLDS 1]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P:2,P
STARPILOT’S GRAVE [MAGEWORLDS 2]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P,P
STARS ASUNDER, THE [MAGEWORLDS 6]  DOYLE, DEBRA/MACDONALD, JAMES D.  H:2
WORKING OF STARS, A [MAGEWORLDS 7]  DOYLE, DEBRA/MACDONALD, JAMES D.  H,P
SEVENTH MOUNTAIN, THE [MAGI 1]  CURTIS, GENE  LP
BURNING CITY, THE [MAGIC 4]  NIVEN, LARRY/POURNELLE, JERRY  H:2,P
BURNING TOWER [MAGIC 5]  NIVEN, LARRY/POURNELLE, JERRY  H:2
MAGIC GOES AWAY, THE [MAGIC 1]  NIVEN, LARRY  H:2,LP
MAGIC MAY RETURN, THE [MAGIC 2]  NIVEN, LARRY  C/PA,LPA
MORE MAGIC [MAGIC 3]  NIVEN, LARRY  LP
JOURNEY BEYOND THE BEYOND [MAGIC LANDS]  STANEK, ROBERT  LP
BY THE SWORD [MAGIC OF THE PLAINS #1]  COSTIKYAN, GREG  C/P,@H:2,P
MAGIC OR NOT? [MAGIC OR NOT? 1]  EAGER, EDWARD  H,LP
WELL-WISHERS, THE [MAGIC OR NOT? 2]  EAGER, EDWARD  LP
JENNIFER MURDLEY’S TOAD [MAGIC SHOP 1]  COVILLE, BRUCE  LP
JEREMY THATCHER, DRAGON HATCHER [MAGIC SHOP 2]  COVILLE, BRUCE  LP
MONSTER’S RING, THE [MAGIC SHOP 3]  COVILLE, BRUCE  C/P,LP
AND PEACE SHALL SLEEP [MAGIC: THE GATHERING]  LYRIS, SONIA ORIN  P
ARENA [MAGIC: THE GATHERING]  FORSTCHEN, WILLIAM R.  C/P:2,P
ASHES OF THE SUN [MAGIC: THE GATHERING]  BRADDOCK, HANOVI  C/P,P
COLORS OF MAGIC, THE [@MAGIC: THE GATHERING]  LEBOW, JESS  PA
CURSED LAND, THE [MAGIC: THE GATHERING]  MCLAREN, TERI  P
DARK LEGACY [@MAGIC: THE GATHERING]  VARDEMAN, ROBERT E.  P
DISTANT PLANES [MAGIC: THE GATHERING]  ICE, KATHY  HA,PA
DRAGONS OF MAGIC, THE [@MAGIC: THE GATHERING]  KING, J. ROBERT  PA
FINAL SACRIFICE [MAGIC: THE GATHERING]  EMERY, CLAYTON  C/P:2,P
MONSTERS OF MAGIC, THE [@MAGIC: THE GATHERING]  KING, J. ROBERT  PA
MYTHS OF MAGIC ANTHOLOGY, THE [@MAGIC: THE GATHERING]  LEBOW, JESS  PA
PRODIGAL SORCERER, THE [MAGIC: THE GATHERING]  SUMNER, MARK  P
RATH AND STORM [@MAGIC: THE GATHERING]  ARCHER, PETER  PA
SECRETS OF MAGIC, THE [@MAGIC: THE GATHERING]  LEBOW, JESS  PA
SHATTERED CHAINS [MAGIC: THE GATHERING]  EMERY, CLAYTON  C/P:2,P
SONG OF TIME [MAGIC: THE GATHERING]  MCLAREN, TERI  P
TAPESTRIES [MAGIC: THE GATHERING]  ICE, KATHY  LP
THRAN, THE [@MAGIC: THE GATHERING]  KING, J. ROBERT  P
WHISPERING WOODS [MAGIC: THE GATHERING]  EMERY, CLAYTON  C/P:2,P
BLOODLINES [@MAGIC: THE GATHERING ARTIFACTS 4]  COLEMAN, LOREN L.  C/P,P
BROTHERS’ WAR, THE [@MAGIC: THE GATHERING ARTIFACTS 1]  GRUBB, JEFF  P

TITLE  AUTHOR  CODE
PLANESWALKER [MAGIC: THE GATHERING ARTIFACTS 2]  ABBEY, LYNN  P
TIME STREAMS [@MAGIC: THE GATHERING ARTIFACTS 3]  KING, J. ROBERT  P
ETERNAL ICE, THE [@MAGIC: THE GATHERING ICE AGE 2]  GRUBB, JEFF  P
GATHERING DARK, THE [@MAGIC: THE GATHERING ICE AGE 1]  GRUBB, JEFF  P
SHATTERED ALLIANCE, THE [@MAGIC: THE GATHERING ICE AGE 3]  GRUBB, JEFF  P
APOCALYPSE [@MAGIC: THE GATHERING INVASION 3]  KING, J. ROBERT  C/P,P
INVASION [@MAGIC: THE GATHERING INVASION 1]  KING, J. ROBERT  C/P:2,P
PLANESHIFT [@MAGIC: THE GATHERING INVASION 2]  KING, J. ROBERT  C/P,P

GUARDIAN:SAVIORS OF KAMIGAWA [@MAGIC: THE GATHERING KAMIGAWA 3] MCGOUGH, SCOTT

P

HERITIC: BETRAYERS OF KAMIGAWA [@MAGIC: THE GATHERING KAMIGAWA 2] MCGOUGH, SCOTT

P

OUTLAW: CHAMPIONS OF KAMIGAWA [@MAGIC: THE GATHERING KAMIGAWA 1] MCGOUGH, SCOTT

C/P,P

HAZEZON [@MAGIC: THE GATHERING LEGENDS 3]  EMERY, CLAYTON  P
JEDIT [@MAGIC: THE GATHERING LEGENDS 2]  EMERY, CLAYTON  P
JOHAN [@MAGIC: THE GATHERING LEGENDS 1]  EMERY, CLAYTON  P
ASSASSIN’S BLADE [@MAGIC: THE GATHERING LEGENDS TWO 1]  MCGOUGH, SCOTT  P
CHAMPION’S TRIAL [@MAGIC: THE GATHERING LEGENDS TWO 3]  MCGOUGH, SCOTT  P
EMPEROR’S FIST [@MAGIC: THE GATHERING LEGENDS TWO 2]  MCGOUGH, SCOTT  P
NEMESIS [@MAGIC: THE GATHERING MASQUERADE 2]  THOMPSON, PAUL B.  P
PROPHECY [@MAGIC: THE GATHERING MASQUERADE 3]  MOORE, VANCE  P
DARKSTEEL EYE, THE [@MAGIC: THE GATHERING MIRRODIN 2]  LEBOW, JESS  P
MOONS OF MIRRODIN, THE [@MAGIC: THE GATHERING MIRRODIN 1]  MCDERMOTT, WILL  P
FIFTH DAWN, THE [@MAGIC: THE GATHERING MIRRODIN #3]  HERNDON, CORY J.  P
CHAINER’S TORMENT [@MAGIC: THE GATHERING ODYSSEY 2]  MCGOUGH, SCOTT  P
JUDGMENT [@MAGIC: THE GATHERING ODYSSEY 3]  MCDERMOTT, WILL  P
ODYSSEY [@MAGIC: THE GATHERING ODYSSEY 1]  MOORE, VANCE  P
LEGIONS [@MAGIC: THE GATHERING ONSLAUGHT 2]  KING, J. ROBERT  P
ONSLAUGHT [@MAGIC: THE GATHERING ONSLAUGHT 1]  KING, J. ROBERT  P
SCOURGE [@MAGIC: THE GATHERING ONSLAUGHT 3]  KING, J. ROBERT  P
DISSENSION [@MAGIC: THE GATHERING RAVNICA 3]  HERNDON, CORY J.  P
GUILDPACT [MAGIC: THE GATHERING RAVNICA 2]  HERNDON, CORY J.  P
RAVNICA: CITY OF GUILDS [MAGIC: THE GATHERING RAVNICA 1]  HERNDON, CORY J.  P
TIME SPIRAL [@MAGIC: THE GATHERING TIME SPIRAL #1]  MCGOUGH, SCOTT  P
PLANAR CHAOS [@MAGIC: THE GATHERING TIME SPIRAL #2]  MCGOUGH, SCOTT/SANDERS, TIMOTHY  P
FUTURE SIGHT [@MAGIC: THE GATHERING TIME SPIRAL #3]  MCGOUGH, SCOTT/DELANEY, JOHN  P
MAGIC TIME [MAGIC TIME]  ZICREE, MARC SCOTT/HAMBLY, BARBARA  H

MAGIC TIME: ANGELFIRE [MAGIC TIME] ZICREE, MARC SCOTT/

BOHNHOFF, MAYA KAATHRYN

H

STEPS UP THE CHIMNEY, THE [MAGICIAN’S HOUSE 1]  CORLETT, WILLIAM  P
MAGICKERS, THE [MAGICKERS 1]  DRAKE, EMILY  P
CURSE OF ARKADY, THE [@MAGICKERS #2]  DRAKE, EMILY  P
DRAGON GUARD, THE [@MAGICKERS #3]  DRAKE, EMILY  P
TOMORROW’S MAGIC [MAGIC’S RETURN 2]  SERVICE, PAMELA F.  H,P
WINTER OF MAGIC’S RETURN [MAGIC’S RETURN 1]  SERVICE, PAMELA F.  H,P
ARMY OF DARKNESS [MAGIRA 2]  WALKER, HUGH  C/P,P
MESSENGERS OF DARKNESS [MAGIRA 3]  WALKER, HUGH  P
WAR-GAMERS’ WORLD [MAGIRA 1]  WALKER, HUGH  P
FEAST OF SOULS [@MAGISTER #1]  FRIEDMAN, C. S.  H
CROWN OF SILENCE, THE [MAGRAVANDIAS 2]  CONSTANTINE, STORM  H:2
SEA DRAGON HEIR [MAGRAVANDIAS 1]  CONSTANTINE, STORM  H:2
WAY OF LIGHT, THE [MAGRAVANDIAS 3]  CONSTANTINE, STORM  H:2
MAGIC & MALICE [MAIRELON AND KIM 1,2]  WREDE, PATRICIA C.  H
MAGICIAN’S WARD [MAIRELON AND KIM 2]  WREDE, PATRICIA C.  H
MAIRELON THE MAGICIAN [MAIRELON AND KIM 1]  WREDE, PATRICIA C.  H,P
LORD VALENTINE’S CASTLE [MAJIPOOR 1]  SILVERBERG, ROBERT  H:2,SFWA,P
LORD VALENTINE’S CASTLE (PROOF) [MAJIPOOR 1]  SILVERBERG, ROBERT  SR-LP
MAJIPOOR CHRONICLES [MAJIPOOR 2]  SILVERBERG, ROBERT  H:2,P
VALENTINE PONTIFEX [MAJIPOOR 3]  SILVERBERG, ROBERT  H:2,P
KING OF DREAMS, THE [MAJIPOOR PRESTIMION 3]  SILVERBERG, ROBERT  H,P
LORD PRESTIMION [MAJIPOOR PRESTIMION 2]  SILVERBERG, ROBERT  H:2,P
SORCERERS OF MAJIPOOR [MAJIPOOR PRESTIMION 1]  SILVERBERG, ROBERT  H:2,P
BLOOD FOLLOWS [MALAZAN]  ERIKSON, STEVEN  SR-LP
BONEHUNTERS, THE [MALAZAN 6]  ERIKSON, STEVEN  LP
DEADHOUSE GATES [MALAZAN 2]  ERIKSON, STEVEN  C/P,LP
GARDENS OF THE MOON [MALAZAN 1]  ERIKSON, STEVEN  C/P,P,H:2
HOUSE OF CHAINS [MALAZAN 4]  ERIKSON, STEVEN  H
MEMORIES OF ICE [MALAZAN 3]  ERIKSON, STEVEN  H,LP
MIDNIGHT TIDES [MALAZAN 5]  ERIKSON, STEVEN  H
DEATH GOD’S DOOM [MALKAR 1]  TUBB, E. C.  LP
SLEEPING CITY, THE [MALKAR 2]  TUBB, E. C.  LP
DEMON LORD OF KARANDA [MALLOREON 3]  EDDINGS, DAVID  H:2,P

The MITSFS Seriesdex  September 24, 2007  Page 103
TITLE  AUTHOR  CODE
GUARDIANS OF THE WEST [MALLOREON 1]  EDDINGS, DAVID  C/P,H,P,SFWA
KING OF THE MURGOS [MALLOREON 2]  EDDINGS, DAVID  H:2,P
SEERESS OF KELL [MALLOREON 5]  EDDINGS, DAVID  C/P,H:2,P
SORCERESS OF DARSHIVA [MALLOREON 4]  EDDINGS, DAVID  H:2,P
MALLWORLD [MALLWORLD]  SUCHARITKUL, SOMTOW  H:2,P
ULTIMATE MALLWORLD, THE [MALLWORLD]  SOMTOW, S. P.  LP

MAMMOTH BOOK OF FANTASTIC SCIENCE FICTION, THE [@MAMMOTH SF 5] WAUGH, CHARLES G./

GREENBERG, MARTIN H.

HA,LPA

MAMMOTH BOOK OF MODERN SCIENCE FICTION, THE [@MAMMOTH SF 6] WAUGH, CHARLES G./

GREENBERG, MARTIN H.

LPA

MAMMOTH BOOK OF VINTAGE SCIENCE FICTION, THE [@MAMMOTH SF 3] WAUGH, CHARLES G./

GREENBERG, MARTIN H.

LPA

ICEBONES [MAMMOTHS]  BAXTER, STEPHEN  H,LP
LONGTUSK [MAMMOTHS 2]  BAXTER, STEPHEN  C/P,P
SILVERHAIR [MAMMOTHS 1]  BAXTER, STEPHEN  C/P,P
ARK OF DOOM [MAN FROM ATLANTIS 4]  WOODLEY, RICHARD  P
DEATH SCOUTS [MAN FROM ATLANTIS 2]  WOODLEY, RICHARD  P
KILLER SPORES [MAN FROM ATLANTIS 3]  WOODLEY, RICHARD  P
MAN FROM ATLANTIS [MAN FROM ATLANTIS 1]  WOODLEY, RICHARD  C/P,P
BACK HOME AT THE O.R.G.Y. [MAN FROM ORGY]  MARK, TED  SR-P
COME BE MY O.R.G.Y. [MAN FROM ORGY]  MARK, TED  SR-P,C/P
SHE-BEAST, THE [@MAN FROM PLANET X #1]  ADAMS, HUNTER  SR-P
TIGER BY THE TAIL [@MAN FROM PLANET X #2]  ADAMS, HUNTER  SR-P
DIRTY, ROTTEN, DEPRIVING RAY, THE [@MAN FROM TOMCAT #4]  KNIGHT, MALLORY T.  P
AFFAIR OF THE GENTLE SABOTEUR, THE [MAN FROM UNCLE]  KEITH, BRANDON  H
AFFAIR OF THE GUNRUNNER’S GOLD, THE [MAN FROM UNCLE]  KEITH, BRANDON  H
COIN OF EL DIABLO AFFAIR, THE [MAN FROM UNCLE]  GIBSON, WALTER B.  VLP
CROSS OF GOLD AFFAIR, THE [MAN FROM UNCLE 14]  DAVIES, FREDRIC  C/P:2,P
MAN FROM U.N.C.L.E., THE [@MAN FROM UNCLE #1]  AVALLONE, MICHAEL  C/P:2,P
DOOMSDAY AFFAIR, THE [@MAN FROM UNCLE #2]  WHITTINGTON, HARRY  C/P:2,P
COPENHAGEN AFFAIR, THE [@MAN FROM UNCLE #3]  ORAM, JOHN  C/P:2,P
DAGGER AFFAIR, THE [@MAN FROM UNCLE #4]  MCDANIEL, DAVID  C/P:2,P

STONE-COLD DEAD IN THE MARKET AFFAIR, THE (BRITISH)

[@MAN FROM UNCLE #4] ORAM, JOHN

SR-P

FINGER IN THE SKY AFFAIR, THE (BRITISH) [@MAN FROM UNCLE #5]  LESLIE, PETER  SR-P
MAD SCIENTIST AFFAIR, THE [@MAN FROM UNCLE #5]  PHILLIFENT, JOHN T.  C/P:2,P
VAMPIRE AFFAIR, THE [@MAN FROM UNCLE #6]  MCDANIEL, DAVID  C/P:2,P
RADIOACTIVE CAMEL AFFAIR, THE [@MAN FROM UNCLE #7]  LESLIE, PETER  C/P:2,P
MONSTER WHEEL AFFAIR, THE [@MAN FROM UNCLE #8]  MCDANIEL, DAVID  C/P:2,P
DIVING DAMES AFFAIR, THE [@MAN FROM UNCLE #9]  LESLIE, PETER  C/P:2,P
ASSASSINATION AFFAIR, THE [@MAN FROM UNCLE #10]  HOLLY, JOAN HUNTER  C/P:2,P
INVISIBILITY AFFAIR, THE [@MAN FROM UNCLE #11]  STRATTON, THOMAS  C/P,P
MIND TWISTER AFFAIR, THE [@MAN FROM UNCLE #12]  STRATTON, THOMAS  C/P:2,P
RAINBOW AFFAIR, THE [@MAN FROM UNCLE #13]  MCDANIEL, DAVID  C/P,P
UTOPIA AFFAIR, THE [@MAN FROM UNCLE #15]  MCDANIEL, DAVID  C/P,P
SPLINTERED SUNGLASSES AFFAIR, THE [@MAN FROM UNCLE #16]  LESLIE, PETER  C/P,P
HOLLOW CROWN AFFAIR, THE [@MAN FROM UNCLE #17]  MCDANIEL, DAVID  C/P:2,P
UNFAIR FARE AFFAIR, THE [@MAN FROM UNCLE #18]  LESLIE, PETER  C/P,P
POWER CUBE AFFAIR, THE [@MAN FROM UNCLE #19]  PHILLIFENT, JOHN T.  P
CORFU AFFAIR, THE [@MAN FROM UNCLE #20]  PHILLIFENT, JOHN T.  P
THINKING MACHINE AFFAIR, THE [@MAN FROM UNCLE #21]  BERNARD, JOEL  C/P,P

STONE-COLD DEAD IN THE MARKET AFFAIR, THE [@MAN FROM UNCLE #22] ORAM, JOHN

C/P,P

FINGER IN THE SKY AFFAIR, THE [@MAN FROM UNCLE #23]  LESLIE, PETER  P
BEST OF ALL POSSIBLE WARS [@MAN-KZIN WARS]  NIVEN, LARRY  PA
CATHOUSE [MAN-KZIN WARS]  ING, DEAN  C/P:2,SFWA-P,P
CHILDREN’S HOUR, THE [MAN-KZIN WARS]  POURNELLE, JERRY/STIRLING, S. M.  C/P:2,SFWA-P,P
DESTINY’S FORGE [@MAN-KZIN WARS]  CHAFE, PAUL C.  H

HOUSES OF THE KZINTI [MAN-KZIN WARS] POURNELLE, JERRY/STIRLING, S. M./

ING, DEAN

LP

INCONSTANT STAR [MAN-KZIN WARS]  ANDERSON, POUL  C/P:2,P
MAN-KZIN WARS, THE [@MAN-KZIN WARS #1]  NIVEN, LARRY  PA,C/PA,SFWA-P
MAN-KZIN WARS 2 [@MAN-KZIN WARS #2]  NIVEN, LARRY  PA,C/PA,SFWA-P:2
MAN-KZIN WARS 3 [@MAN-KZIN WARS #3]  NIVEN, LARRY  PA,C/PA,SFWA-P
MAN-KZIN WARS 4 [@MAN-KZIN WARS #4]  NIVEN, LARRY  PA,C/PA
MAN-KZIN WARS 5 [@MAN-KZIN WARS #5]  NIVEN, LARRY  PA,C/PA,SFWA-P
MAN-KZIN WARS 6 [@MAN-KZIN WARS #6]  NIVEN, LARRY  C/PA,PA
MAN-KZIN WARS 7 [@MAN-KZIN WARS #7]  NIVEN, LARRY  PA,C/PA
CHOOSING NAMES [@MAN-KZIN WARS #8]  NIVEN, LARRY  C/PA,PA
MAN-KZIN WARS 9 [@MAN-KZIN WARS #9]  NIVEN, LARRY  C/PA,PA
WUNDER WAR, THE [@MAN-KZIN WARS #10]  COLEBATCH, HAL  H
MAN-KZIN WARS 11 [@MAN-KZIN WARS #11]  NIVEN, LARRY  HA
EMPEROR AND CLOWN [MAN OF HIS WORD 4]  DUNCAN, DAVE  C/P,P

TITLE  AUTHOR  CODE
FAERY LANDS FORLORN [MAN OF HIS WORD 2]  DUNCAN, DAVE  C/P:2,P
MAGIC CASEMENT [MAN OF HIS WORD 1]  DUNCAN, DAVE  C/P,P
PERILOUS SEAS [MAN OF HIS WORD 3]  DUNCAN, DAVE  C/P,P
MAN PLUS [MAN PLUS 1]  POHL, FREDERIK  C/P,P,H,SFWA
MARS PLUS [MAN PLUS 2]  POHL, FREDERIK/THOMAS, THOMAS T.  H,P
CITY OUTSIDE THE WORLD, THE [MAN WHO LOVED MARS 3]  CARTER, LIN  C/P,P
DOWN TO A SUNLESS SEA [MAN WHO LOVED MARS 4]  CARTER, LIN  C/P,P
MAN WHO LOVED MARS, THE [MAN WHO LOVED MARS 1]  CARTER, LIN  C/P:2,P
VALLEY WHERE TIME STOOD STILL, THE [MAN WHO LOVED MARS 2]  CARTER, LIN  H,P
AEROMANCER [MANCER 4]  CALLANDER, DON  P
AQUAMANCER [MANCER 2]  CALLANDER, DON  P,SFWA-P
GEOMANCER [MANCER 3]  CALLANDER, DON  C/P,SFWA-P:2,P
PYROMANCER [MANCER 1]  CALLANDER, DON  P
MANSE, THE [MANSE 1]  CANTRELL, LISA W.  SFWA-P
TORMENTS [MANSE 2]  CANTRELL, LISA W.  SFWA-P
DAUGHTER OF THE EMPIRE [MARA 1]  FEIST, RAYMOND E./WURTS, JANNY  C/P,P,H
MISTRESS OF THE EMPIRE [MARA 3]  FEIST, RAYMOND E./WURTS, JANNY  H:2,P
SERVANT OF THE EMPIRE [MARA 2]  FEIST, RAYMOND E./WURTS, JANNY  H:2,P
MARATHON [MARATHON 1]  SMITH, DAVID ALEXANDER  P
RENDEZVOUS [MARATHON 2]  SMITH, DAVID ALEXANDER  P
MARCH TO THE SEA [MARCH UPCOUNTRY 2]  WEBER, DAVID/RINGO, JOHN  H,P
MARCH TO THE STARS [MARCH UPCOUNTRY 3]  WEBER, DAVID/RINGO, JOHN  H,P
MARCH UPCOUNTRY [MARCH UPCOUNTRY 1]  WEBER, DAVID/RINGO, JOHN  H
WE FEW [MARCH UPCOUNTRY 4]  WEBER, DAVID/RINGO, JOHN  H
MARCO’S MILLIONS [MARCO 1]  SLEATOR, WILLIAM  H,LP
MARIANNE, THE MADAME, AND THE MOMENTARY GODS [MARIANNE 2]  TEPPER, SHERI S.  C/P:2,P
MARIANNE, THE MAGUS, AND THE MANTICORE [MARIANNE 1]  TEPPER, SHERI S.  C/P:2,P
MARIANNE, THE MATCHBOX, AND THE MALACHITE MOUSE [MARIANNE 3] TEPPER, SHERI S.  C/P,P
BUDAYEEN NIGHTS [MARID AUDRAN]  EFFINGER, GEORGE ALEC  H
EXILE KISS, THE [MARID AUDRAN 3]  EFFINGER, GEORGE ALEC  C/P,LP,H
FIRE IN THE SUN, A [MARID AUDRAN 2]  EFFINGER, GEORGE ALEC  C/P,P,H
WHEN GRAVITY FAILS [MARID AUDRAN 1]  EFFINGER, GEORGE ALEC  C/P,SFWA,P,H
INVISIBLES, THE [MARK HOOD 9]  DARK, JAMES  P
OPERATION OCTOPUS [MARK HOOD 8]  DARK, JAMES  P
SWORD OF GENGHIS KHAN, THE [MARK HOOD 7]  DARK, JAMES  P
THRONE OF SATAN [MARK HOOD 6]  DARK, JAMES  C/P,P
LAST RECALL, THE [MARKED MAN 2]  INGRID, CHARLES  P
MARKED MAN, THE [MARKED MAN 1]  INGRID, CHARLES  C/P:2,P
WEREWOLF’S KISS, THE [MARLEY FAMILY WEREWOLF 1]  SCOTCH, CHERI  P
WEREWOLF’S TOUCH, THE [MARLEY FAMILY WEREWOLF 2]  SCOTCH, CHERI  P
ALANYA TO ALANYA [@MARQ’SSAN #1]  DUCHAMP, L. TIMMEL  LP
RENEGADE [@MARQ’SSAN #2]  DUCHAMP, L. TIMMEL  LP
TSUNAMI [@MARQ’SSAN #3]  DUCHAMP, L. TIMMEL  LP
MARROW [MARROW 1]  REED, ROBERT  H:2,P
WELL OF STARS, THE [MARROW 2]  REED, ROBERT  H
BLUE MARS [MARS 3]  ROBINSON, KIM STANLEY  C/P,P,H
GREEN MARS [MARS 2]  ROBINSON, KIM STANLEY  H:2,LP
GREEN MARS (NOVELLA) [MARS]  ROBINSON, KIM STANLEY  TD1
MARTIANS, THE [MARS]  ROBINSON, KIM STANLEY  H
RED MARS [MARS 1]  ROBINSON, KIM STANLEY  H:2,P
MARS ATTACKS! [@MARS ATTACKS]  GEMS, JONATHAN  P
MARTIAN DEATHTRAP [@MARS ATTACKS]  ARCHER, NATHAN  C/P,P,H
WAR DOGS OF THE GOLDEN HORDE [@MARS ATTACKS]  MURILL, RAY W.  C/P,P
AMAZING SPIDER-MAN, THE [MARVEL COMICS]  LEE, STAN  SCX

BRING ON THE BAD GUYS: ORIGINS OF THE MARVEL COMICS VILLAINS [MARVEL COMICS] LEE, STAN

CX

DOCTOR STRANGE: MASTER OF THE MYSTIC ARTS [MARVEL COMICS]  LEE, STAN  CX
ORIGINS OF MARVEL COMICS [MARVEL COMICS]  LEE, STAN  CX
SON OF ORIGINS OF MARVEL COMICS [MARVEL COMICS]  LEE, STAN  CX
ULTIMATE HULK, THE [MARVEL COMICS]  LEE, STAN/DAVID, PETER  LPA
ULTIMATE SILVER SURFER, THE [MARVEL COMICS]  LEE, STAN  LPA
ULTIMATE SPIDER-MAN, THE [MARVEL COMICS]  LEE, STAN  LPA
ULTIMATE SUPER-VILLAINS, THE [MARVEL COMICS]  LEE, STAN  LPA
SECRET WARS [MARVEL’S FINEST]  SHOOTER, JIM  CX
INADEQUATE ADEPT, THE [MARVIN BREWSTER 2]  HAWKE, SIMON  C/P:2,P
RELUCTANT SORCERER, THE [MARVIN BREWSTER 1]  HAWKE, SIMON  C/P,P
REVELATION [@MASS EFFECT]  KARPYSHYN, DREW  P
BRIDGE OF BIRDS [MASTER LI 1]  HUGHART, BARRY  C/P,P,H
EIGHT SKILLED GENTLEMEN [MASTER LI 3]  HUGHART, BARRY  C/P,LP,H
STORY OF THE STONE, THE [MASTER LI 2]  HUGHART, BARRY  C/P,P,H
STELLAR DEATH PLAN, THE [@MASTERS OF SPACE #1]  VARDEMAN, ROBERT E.  C/P,P
ALIEN WEB, THE [@MASTERS OF SPACE #2]  VARDEMAN, ROBERT E.  P
PLAGUE IN PARADISE, A [@MASTERS OF SPACE #3]  VARDEMAN, ROBERT E.  P

The MITSFS Seriesdex  September 24, 2007  Page 105
TITLE  AUTHOR  CODE
ALBINO KNIFE, THE [MATADOR 5]  PERRY, STEVE  C/P:2,SFWA-P:2,P
BLACK STEEL [MATADOR 6]  PERRY, STEVE  C/P,SFWA-P:2,P
BROTHER DEATH [MATADOR 7]  PERRY, STEVE  P,SFWA-P:2
MACHIAVELLI INTERFACE, THE [MATADOR 3]  PERRY, STEVE  C/P:2,P
MAN WHO NEVER MISSED, THE [MATADOR 1]  PERRY, STEVE  C/P,P
MATADORA [MATADOR 2]  PERRY, STEVE  C/P:2,P
97TH STEP, THE [MATADOR 0]  PERRY, STEVE  C/P:2,SFWA-P,P
OMEGA CAGE, THE [MATADOR 4]  PERRY, STEVE/REAVES, MICHAEL  C/P,SFWA-P,P
DAUGHTERS OF A CORAL DAWN [MATERNAS 1]  FORREST, KATHERINE V.  LP
DAUGHTERS OF AN AMBER MOON [MATERNAS 2]  FORREST, KATHERINE V.  LP
DAUGHTERS OF AN EMERALD DUSK [MATERNAS 3]  FORREST, KATHERINE V.  LP
ANCIENT LEGACY, THE [MATHEW LEWIN 3]  GRAHAM, MITCHELL  P
EMERALD CAVERN, THE [MATHEW LEWIN 2]  GRAHAM, MITCHELL  P
FIFTH RING, THE [MATHEW LEWIN 1]  GRAHAM, MITCHELL  P
DEADLIEST SHOW IN TOWN, THE [MATHEW SWAIN 3]  MCQUAY, MIKE  P
HOT TIME IN OLD TOWN [MATHEW SWAIN 1]  MCQUAY, MIKE  P
WHEN TROUBLE BECKONS [MATHEW SWAIN 2]  MCQUAY, MIKE  P
PAST THE SIZE OF DREAMING [MATILDA BLACK 2]  HOFFMAN, NINA KIRIKI  P
RED HEART OF MEMORIES, A [MATILDA BLACK]  HOFFMAN, NINA KIRIKI  LP
AIRBORN [MATT CRUSE 1]  OPPEL, KENNETH  H,P
CULTURES BEYOND THE EARTH [MATT CRUSE 1]  MARUYAMA, MAGOROH/HARKINS, ARTHUR R/REF-P
SKYBREAKER [MATT CRUSE 2]  OPPEL, KENNETH  H
COLD WAR IN A COUNTRY GARDEN [MATTHEW DILKE 1]  GUTTERIDGE, LINDSAY  C/P:2,P
KILLER PINE [MATTHEW DILKE 2]  GUTTERIDGE, LINDSAY  H,P
BRAZEN RULE, THE [MATTHEWS 1]  BURGAUER, STEVEN  H:2,LP
FORNAX [MATTHEWS 3]  BURGAUER, STEVEN  H,LP
IN THE SHADOW OF OMEN [MATTHEWS 4]  BURGAUER, STEVEN  LP
LAST AMERICAN, THE [MATTHEWS 2]  BURGAUER, STEVEN  H,LP
TREACHERY ON THE DARK SIDE [MATTHEWS 6]  BURGAUER, STEVEN  LP
DESTINED QUEEN, THE [MAURA WOODBURY 2]  HALE, DEBORAH  LP
WIZARD’S WARD, THE [MAURA WOODBURY 1]  HALE, DEBORAH  LP
SUMMER COUNTRY, THE [MAUREEN PIERCE 1]  HETLEY, JAMES A.  LP
WINTER OAK, THE [MAUREEN PIERCE 2]  HETLEY, JAMES A.  LP
FLIGHT OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 2]  TEPPER, SHERI S.  C/P:2,P
SEARCH OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 3]  TEPPER, SHERI S.  C/P:2,P
SONG OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 1]  TEPPER, SHERI S.  C/P:2,P

ADVENTURES IN TIME AND SPACE WITH MAX MERRIWELL [MAX MERRIWELL 4] MURPHY, PAT

H

THERE AND BACK AGAIN [MAX MERRIWELL 1]  MURPHY, PAT  P
WILD ANGEL [MAX MERRIWELL 2]  MURPHY, PAT  P
BRAIDED WORLD, THE [MAXIMUM ICE 2]  KENYON, KAY  P
MAXIMUM ICE [MAXIMUM ICE 1]  KENYON, KAY  P
LASHER [MAYFAIR WITCHES 2]  RICE, ANNE  C/P,P,H
MERRICK [MAYFAIR WITCHES/LESTAT]  RICE, ANNE  H:2
TALTOS [MAYFAIR WITCHES 3]  RICE, ANNE  H:2,P
WITCHING HOUR, THE [MAYFAIR WITCHES 1]  RICE, ANNE  H:2,P
WITCH’S COMPANION, THE [MAYFAIR WITCHES]  RAMSLAND, KATHERINE  R/REF-H
LOVELOCK [MAYFLOWER 1]  CARD, ORSON SCOTT/KIDD, KATHRYN H.  H,P
CRESCENT IN THE SKY [MECHANICAL SKY 1]  MOFFITT, DONALD  C/P,P
GATHERING OF STARS, A [MECHANICAL SKY 2]  MOFFITT, DONALD  P
DYING TIME, THE [@MECHWARRIOR]  GRESSMAN, THOMAS S.  P
GHOST OF WINTER [@MECHWARRIOR]  KENSON, STEPHEN  P
IMMINENT CRISIS [@MECHWARRIOR]  BILLS, RANDALL N.  P
INITIATION TO WAR [@MECHWARRIOR/BATTLETECH]  CHARRETTE, ROBERT N.  P
ROAR OF HONOR [@MECHWARRIOR]  PARDOE, BLAINE LEE  P
BLOOD AVATAR [@MECHWARRIOR DARK AGE/BATTLETECH]  BICK, ILSA J.  P
BLOOD OF THE ISLE [@MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P

BY TEMPTATIONS AND BY WAR [@MECHWARRIOR DARK AGE/ BATTLETECH] COLEMAN, LOREN L.

LP

BY TEMPTATIONS AND BY WAR (PROOF) [@MECHWARRIOR DARK AGE/ BATTLETECH] COLEMAN, LOREN L.

SR-LP

CALL TO ARMS, A [MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
DAUGHTER OF THE DRAGON [@MECHWARRIOR DARK AGE/BATTLETECH]  BICK, ILSA J.  P
DRAGON RISING [@MECHWARRIOR DARK AGE/BATTLETECH]  BICK, ILSA J.  P
FLIGHT OF THE FALCON [MECHWARRIOR DARK AGE/BATTLETECH]  MILAN, VICTOR  P
FORTRESS OF LIES [@MECHWARRIOR DARK AGE/BATTLETECH]  YORK, J. STEVEN  P
FORTRESS REPUBLIC [@MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
GHOST WAR [@MECHWARRIOR DARK AGE 1/BATTLETECH]  STACKPOLE, MICHAEL A.  P
HERETIC’S FAITH [@MECHWARRIOR DARK AGE/BATTLETECH]  BILLS, RANDALL N.  P
HUNTERS OF THE DEEP [@MECHWARRIOR DARK AGE/BATTLETECH]  BILLS, RANDALL N.  P
MASTERS OF WAR [@MECHWARRIOR DARK AGE]  STACKPOLE, MICHAEL A.  P
PATRIOT’S STAND [@MECHWARRIOR DARK AGE/BATTLETECH]  MOSCOE, MIKE  P
PRINCIPLES OF DESOLATION [@MECHWARRIOR DARK AGE/BATTLETECH]  HARDY, JASON M./BILLS, RANDALL N.  P

The MITSFS Seriesdex  September 24, 2007  Page 106
TITLE  AUTHOR  CODE
RENDING OF FALCONS, A [@MECHWARRIOR DARK AGE/BATTLETECH]  MILAN, VICTOR  P
RUINS OF POWER, THE [@MECHWARRIOR DARK AGE/BATTLETECH]  VARDEMAN, ROBERT E.  P
SCORPION JAR, THE [@MECHWARRIOR DARK AGE/BATTLETECH]  HARDY, JASON M.  P
SURRENDER YOUR DREAMS [@MECHWARRIOR DARK AGE/BATTLETECH]  PARDOE, BLAINE LEE  P
SWORD OF SEDITION [@MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
TARGET OF OPPORTUNITY [@MECHWARRIOR DARK AGE/BATTLETECH]  PARDOE, BLAINE LEE  P
TRIAL BY CHAOS [@MECHWARRIOR DARK AGE/BATTLETECH]  YORK, J. STEVEN  P
WOLF HUNTERS [@MECHWARRIOR DARK AGE/BATTLETECH]  KILLIANY, KEVIN  P

SERVICE FOR THE DEAD [@MECHWARRIOR DARK AGE PROVING GROUNDS 3/BATTLETECH] DELRIO, MARTIN

P

SILENCE IN THE HEAVENS, A [@MECHWARRIOR DARK AGE PROVING GROUNDS 1/BATTLETECH] DELRIO, MARTIN

P

TRUTH AND SHADOWS [@MECHWARRIOR DARK AGE PROVING GROUNDS 2/BATTLETECH] DELRIO, MARTIN

P

DOCTOR TO THE STARS [MED]  LEINSTER, MURRAY  C/P:2,P
MED SERIES, THE [MED 1,2,3]  LEINSTER, MURRAY  C/P:2,P
MED SHIP [MED]  LEINSTER, MURRAY  C/P,P
MUTANT WEAPON, THE [MED 1]  LEINSTER, MURRAY  D66525,DD403
S.O.S. FROM THREE WORLDS [MED 2]  LEINSTER, MURRAY  C/P:2,P
THIS WORLD IS TABOO [MED 3]  LEINSTER, MURRAY  C/P:2,P
SWIFTLY TILTING PLANET, A [MEG MURRY 3]  L’ENGLE, MADELEINE  C/P,P,H
WIND IN THE DOOR, A [MEG MURRY 2]  L’ENGLE, MADELEINE  C/P,LP,H
WRINKLE IN TIME, A [MEG MURRY 1]  L’ENGLE, MADELEINE  C/P,P,H:2
WRINKLE IN TIME, A (BRITISH) [MEG MURRY 1]  L’ENGLE, MADELEINE  SR-P
DAWN STAR, THE [MEL DAWNFIELD 2]  ASARO, CATHERINE  LP
NIGHT OF THE LIVING GATOR! [@MELVINGE 3]  LUPOFF, RICHARD  P
NIGHT OF THE LIVING RAT [MELVINGE 2]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P,P
NIGHT OF THE LIVING SHARK [MELVINGE 1]  BISCHOFF, DAVID F.  P
DRAGONBONE CHAIR, THE [MEMORY,SORROW AND THORN 1]  WILLIAMS, TAD  C/P,H:2,P
STONE OF FAREWELL [MEMORY,SORROW AND THORN 2]  WILLIAMS, TAD  H,P
TO GREEN ANGEL TOWER [MEMORY,SORROW AND THORN 3]  WILLIAMS, TAD  H:2,SFWA:2
TO GREEN ANGEL TOWER (PART ONE) [MEMORY,SORROW AND THORN 3]  WILLIAMS, TAD  C/P,P
TO GREEN ANGEL TOWER (PART TWO) [MEMORY,SORROW AND THORN 3]  WILLIAMS, TAD  C/P:2,P
GRAZER CONSPIRACY, THE [MEN IN BLACK]  SMITH, DEAN WESLEY  P
GREEN SALIVA BLUES, THE [MEN IN BLACK]  SMITH, DEAN WESLEY  P
MEN IN BLACK [MEN IN BLACK]  PERRY, STEVE/SOLOMON, ED  C/P:2,P
MEN IN BLACK II [MEN IN BLACK]  FRIESNER, ESTHER M.  P

NIMBLE MAN, THE [@MENAGERIE] GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

P

TEARS OF THE FURIES [@MENAGERIE] GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

P

NAMELESS MAGERY [MENNENKALTENEI 1]  TURNER, DELIA MARSHALL  P
OF SWORDS AND SPELLS [MENNENKALTENEI 2]  TURNER, DELIA MARSHALL  P
WAYS OF MAGIC, THE [MENNENKALTENEI 1,2]  TURNER, DELIA MARSHALL  H
MENNYMS IN THE WILDERNESS [MENNYMS]  WAUGH, SYLVIA  P
MERCHANT PRINCE, THE [MERCHANT PRINCE 1]  SHIMERMAN, ARMIN/SCOTT, MICHAEL  P

OUTRAGEOUS FORTUNE [@MERCHANT PRINCE #2] SHIMERMAN, ARMIN/

YARBRO, CHELSEA QUINN

P

CAPITAL OFFENSE [@MERCHANT PRINCE #3]  SHIMERMAN, ARMIN  P
CLAN CORPORATE, THE [MERCHANT PRINCES 3]  STROSS, CHARLES  H:2
FAMILY TRADE, THE [MERCHANT PRINCES 1]  STROSS, CHARLES  H,P
HIDDEN FAMILY, THE [MERCHANT PRINCES 2]  STROSS, CHARLES  H
BLOOD BOUND [MERCY THOMPSON 2]  BRIGGS, PATRICIA  C/P,P
MOON CALLED [MERCY THOMPSON 1]  BRIGGS, PATRICIA  P
CARESS OF TWILIGHT, A [MEREDITH GENTRY 2]  HAMILTON, LAURELL K.  H,P
KISS OF SHADOWS, A [MEREDITH GENTRY 1]  HAMILTON, LAURELL K.  H
MISTRAL’S KISS [MEREDITH GENTRY 5]  HAMILTON, LAURELL K.  H
SEDUCED BY MOONLIGHT [MEREDITH GENTRY 3]  HAMILTON, LAURELL K.  H
STROKE OF MIDNIGHT, A [MEREDITH GENTRY 4]  HAMILTON, LAURELL K.  H
CRYSTAL ROSE, THE [MERI 3]  BOHNHOFF, MAYA KAATHRYN  P
MERI, THE [MERI 1]  BOHNHOFF, MAYA KAATHRYN  C/P,SFWA-P,P
TAMINY [MERI 2]  BOHNHOFF, MAYA KAATHRYN  P,SFWA-P
BROKEN KINGS, THE [MERLIN CODEX 3]  HOLDSTOCK, ROBERT P.  H
CELTIKA [MERLIN CODEX 1]  HOLDSTOCK, ROBERT P.  P
IRON GRAIL, THE [MERLIN CODEX 2]  HOLDSTOCK, ROBERT P.  P
GUARDIAN OF THE BALANCE [@MERLIN’S DESCENDANTS #1]  RADFORD, IRENE  H,P
GUARDIAN OF THE TRUST [@MERLIN’S DESCENDANTS #2]  RADFORD, IRENE  P
GUARDIAN OF THE VISION [@MERLIN’S DESCENDANTS #3]  RADFORD, IRENE  P
GUARDIAN OF THE PROMISE [@MERLIN’S DESCENDANTS #4]  RADFORD, IRENE  H,P
FESTIVAL MOON [@MEROVINGEN NIGHTS #1]  CHERRYH, C. J.  C/PA,PA
FEVER SEASON [@MEROVINGEN NIGHTS #2]  CHERRYH, C. J.  C/PA,PA
TROUBLED WATERS [@MEROVINGEN NIGHTS #3]  CHERRYH, C. J.  PA,C/PA
SMUGGLER’S GOLD [@MEROVINGEN NIGHTS #4]  CHERRYH, C. J.  PA,C/PA

TITLE  AUTHOR  CODE
DIVINE RIGHT [@MEROVINGEN NIGHTS #5]  CHERRYH, C. J.  PA
FLOOD TIDE [@MEROVINGEN NIGHTS #6]  CHERRYH, C. J.  C/PA,PA
ENDGAME [@MEROVINGEN NIGHTS #7]  CHERRYH, C. J.  C/PA,PA
LOOK INTO THE SUN [MESSENGER 2]  KELLY, JAMES PATRICK  C/P,SFWA-P,P,H
PLANET OF WHISPERS [MESSENGER 1]  KELLY, JAMES PATRICK  C/P,SFWA,SFWA-P,P,H
DARK MESSIAH [MESSIAH STONE 2]  CAIDIN, MARTIN  C/P:2,P
MESSIAH STONE, THE [MESSIAH STONE 1]  CAIDIN, MARTIN  C/P,P
METAPLANETARY [METAPLANETARY 1]  DANIEL, TONY  H
SUPERLUMINAL [METAPLANETARY 2]  DANIEL, TONY  H
BARBARIANS OF MARS [MICHAEL KANE 3]  BRADBURY, EDWARD P.  SR-P,C/P:2
BARBARIANS OF MARS [MICHAEL KANE 3]  MOORCOCK, MICHAEL  P
BLADES OF MARS [MICHAEL KANE 2]  BRADBURY, EDWARD P.  SR-P,C/P:2
BLADES OF MARS [MICHAEL KANE 2]  MOORCOCK, MICHAEL  P
CITY OF THE BEAST, THE [MICHAEL KANE 1]  MOORCOCK, MICHAEL  C/P:2,P
LORD OF THE SPIDERS, THE [MICHAEL KANE 2]  MOORCOCK, MICHAEL  C/P,P
MASTERS OF THE PIT OR BARBARIANS OF MARS [MICHAEL KANE 3]  MOORCOCK, MICHAEL  C/P:2,P
WARRIORS OF MARS [MICHAEL KANE 1]  BRADBURY, EDWARD P.  SR-P
WARRIORS OF MARS [MICHAEL KANE 1]  MOORCOCK, MICHAEL  P
INFINITY CONCERTO, THE [MICHAEL PERRIN 1]  BEAR, GREG  C/P:2,P
SERPENT MAGE, THE [MICHAEL PERRIN 2]  BEAR, GREG  C/P:2,P
SONGS OF EARTH & POWER [MICHAEL PERRIN (REVISED) 1,2]  BEAR, GREG  P
MICROCOLONY, THE [MICRONAUTS 2]  WILLIAMS, GORDON  P
MICRONAUTS, THE [MICRONAUTS 1]  WILLIAMS, GORDON  P
REVOLT OF THE MICRONAUTS [MICRONAUTS 3]  WILLIAMS, GORDON  P
BLUE NOON [@MIDNIGHTERS #3]  WESTERFELD, SCOTT  H
BLUE NOON (PROOF) [@MIDNIGHTERS #3]  WESTERFELD, SCOTT  SR-LP
SEA OF MIST, THE [MIGHT AND MAGIC]  ODOM, MEL  P
DREAMWRIGHT [@MIGHT AND MAGIC #1]  GRAVEL, GEARY  P
SHADOWSMITH, THE [@MIGHT AND MAGIC #2]  GRAVEL, GEARY  C/P,P
MIKE MARS ASTRONAUT [@MIKE MARS #1]  WOLLHEIM, DONALD A.  P
MIKE MARS FLIES THE X-15 [@MIKE MARS #2]  WOLLHEIM, DONALD A.  P
MIKE MARS AT CAPE KENNEDY [@MIKE MARS #3]  WOLLHEIM, DONALD A.  P
MIKE MARS IN ORBIT [MIKE MARS #4]  WOLLHEIM, DONALD A.  H,@P
MIKE MARS FLIES THE DYNA-SOAR [@MIKE MARS #5]  WOLLHEIM, DONALD A.  P
MIKE MARS, SOUTH POLE SPACEMAN [@MIKE MARS #6]  WOLLHEIM, DONALD A.  P
MIKE MARS AND THE MYSTERY SATELLITE [@MIKE MARS #7]  WOLLHEIM, DONALD A.  P
MIKE MARS AROUND THE MOON [@MIKE MARS #8]  WOLLHEIM, DONALD A.  P
BARRAYAR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 2]  BUJOLD, LOIS MCMASTER  C/P:2,SFWA-P,P,H
BORDERS OF INFINITY [MILES VORKOSIGAN 3]  BUJOLD, LOIS MCMASTER  C/P:2,SFWA-P,P
BROTHERS IN ARMS [MILES VORKOSIGAN 2]  BUJOLD, LOIS MCMASTER  C/P:2,P
CETAGANDA [MILES VORKOSIGAN 6]  BUJOLD, LOIS MCMASTER  C/P,P,H
CIVIL CAMPAIGN, A [MILES VORKOSIGAN 9]  BUJOLD, LOIS MCMASTER  H,P
CORDELIA’S HONOR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 1,2]  BUJOLD, LOIS MCMASTER  LP
DIPLOMATIC IMMUNITY [MILES VORKOSIGAN 10]  BUJOLD, LOIS MCMASTER  H,P
KOMARR [MILES VORKOSIGAN 8]  BUJOLD, LOIS MCMASTER  H:2,P
MEMORY [MILES VORKOSIGAN 7]  BUJOLD, LOIS MCMASTER  H:2,P
MIRROR DANCE [MILES VORKOSIGAN 5]  BUJOLD, LOIS MCMASTER  H,SFWA,P
SHARDS OF HONOR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 1]  BUJOLD, LOIS MCMASTER  C/P:2,P,H
VOR GAME, THE [MILES VORKOSIGAN 4]  BUJOLD, LOIS MCMASTER  C/P:2,P
VORKOSIGAN’S GAME [MILES VORKOSIGAN 3,4]  BUJOLD, LOIS MCMASTER  H
WARRIOR’S APPRENTICE, THE [MILES VORKOSIGAN 1]  BUJOLD, LOIS MCMASTER  C/P:2,P
WARRIOR’S APPRENTICE, THE (REVISED) [MILES VORKOSIGAN 1]  BUJOLD, LOIS MCMASTER  H
FRENCHMAN, THE [@MILLENIUM]  HAND, ELIZABETH  C/P,P
IN THE MOOD [MILLENNIUM QUARTET 2]  GRANT, CHARLES L.  H
SYMPHONY [MILLENNIUM QUARTET 1]  GRANT, CHARLES L.  P
ASSASSINS FROM TOMORROW [MIND BROTHERS 2]  HEATH, PETER  C/P:2,P
MEN WHO DIE TWICE [MIND BROTHERS 3]  HEATH, PETER  P
MIND BROTHERS, THE [MIND BROTHERS 1]  HEATH, PETER  C/P:2,P
AMAZONS [MIND MASTERS 4]  ROSS, IAN  C/P,P
RECYCLED SOULS [MIND MASTERS 5]  ROSS, IAN  P
MIND MASTERS, THE [@MIND MASTERS #1]  ROSSMANN, JOHN F.  P
SHAMBALLAH [@MIND MASTERS #2]  ROSSMANN, JOHN F.  P
DEATHKILLER [MINDKILLER 1,2]  ROBINSON, SPIDER  C/P,P
MINDKILLER [MINDKILLER 1]  ROBINSON, SPIDER  C/P,P,H
TIME PRESSURE [MINDKILLER 2]  ROBINSON, SPIDER  C/P,SFWA-P,H,P,SFWA
DOPPELGANGER [MIREI 1]  BRENNAN, MARIE  P
WARRIOR AND WITCH [MIREI 2]  BRENNAN, MARIE  P
RING-A-DING U F O’S, THE [MISS FROM SIS]  TRALINS, ROBERT  P
MISS PICKERELL GOES TO MARS [@MISS PICKERELL]  MACGREGOR, ELLEN  C/P:2,LP
MISS PICKERELL ON THE MOON [@MISS PICKERELL #1]  MACGREGOR, ELLEN/PANTELL, DORA  H:2,P
MISS PICKERELL AND THE WEATHER SATELLITE [@MISS PICKERELL #2]  MACGREGOR, ELLEN/PANTELL, DORA  P
MISS PICKERELL MEETS MR. H.U.M. [@MISS PICKERELL #3]  MACGREGOR, ELLEN/PANTELL, DORA  P
MISS PICKERELL TAKES THE BULL BY THE HORNS [@MISS PICKERELL #4]  MACGREGOR, ELLEN/PANTELL, DORA  P

The MITSFS Seriesdex  September 24, 2007  Page 108
TITLE  AUTHOR  CODE
DEATH OF THE PHOENIX [MISSION CRITICAL]  CHAFE, PAUL C.  P
INVADERS PLAN, THE [@MISSION EARTH #1]  HUBBARD, L. RON  SFWA-P,H,P
BLACK GENESIS [@MISSION EARTH #2]  HUBBARD, L. RON  C/P,SFWA-P,H,P
ENEMY WITHIN, THE [@MISSION EARTH #3]  HUBBARD, L. RON  H,SFWA-P,P
ALIEN AFFAIR, AN [@MISSION EARTH #4]  HUBBARD, L. RON  H,P
FORTUNE OF FEAR [@MISSION EARTH #5]  HUBBARD, L. RON  H,P
FORTUNE OF FEAR (PROOF) [@MISSION EARTH #5]  HUBBARD, L. RON  SR-LP
DEATH QUEST [@MISSION EARTH #6]  HUBBARD, L. RON  H,P
VOYAGE OF VENGEANCE [@MISSION EARTH #7]  HUBBARD, L. RON  C/P,P,H
DISASTER [@MISSION EARTH #8]  HUBBARD, L. RON  C/P,P,H
VILLAINY VICTORIOUS [@MISSION EARTH #9]  HUBBARD, L. RON  H
DOOMED PLANET, THE [@MISSION EARTH #10]  HUBBARD, L. RON  H
CODE NAME: JUDAS [@MISSION IMPOSSIBLE #2]  WALKER, MAX  P
CODE NAME: RAPIER [@MISSION IMPOSSIBLE #3]  WALKER, MAX  P
MISTBORN [MISTBORN 1]  SANDERSON, BRANDON  H
URCHIN OF THE RIDING STARS [@MISTMANTLE #1]  MCALLISTER, M. I.  H
FOREST HOUSE, THE [MISTS OF AVALON 2]  BRADLEY, MARION ZIMMER  H,LP
MISTS OF AVALON, THE [MISTS OF AVALON 1]  BRADLEY, MARION ZIMMER  H:2,LP

PRIESTESS OF AVALON [MISTS OF AVALON 3] BRADLEY, MARION ZIMMER/

PAXSON, DIANA L.

H

DRAGONDOOM [MITHGAR 5]  MCKIERNAN, DENNIS L.  C/P:2,P
DRAGONSTONE, THE [MITHGAR 1]  MCKIERNAN, DENNIS L.  P
EYE OF THE HUNTER, THE [MITHGAR 12]  MCKIERNAN, DENNIS L.  C/P,SFWA,LP
RED SLIPPERS [@MITHGAR]  MCKIERNAN, DENNIS L.  H,P
SILVER CALL, THE (OMNIBUS) [MITHGAR 10,11]  MCKIERNAN, DENNIS L.  P
SILVER WOLF, BLACK FALCON [MITHGAR 13]  MCKIERNAN, DENNIS L.  H
TALES OF MITHGAR [MITHGAR 6]  MCKIERNAN, DENNIS L.  LP
VOYAGE OF THE FOX RIDER [MITHGAR 2]  MCKIERNAN, DENNIS L.  C/P,P
CHAOS MODE [MODE 3]  ANTHONY, PIERS  C/P,P,H
DOOON MODE [MODE 4]  ANTHONY, PIERS  H
FRACTAL MODE [MODE 2]  ANTHONY, PIERS  H:2,SFWA-P:2,P
VIRTUAL MODE [MODE 1]  ANTHONY, PIERS  C/P:2,SFWA-P:2,P
LAST DAY IN LIMBO [MODESTY BLAISE]  O’DONNELL, PETER  SFWA-P
NIGHT OF MORNINGSTAR, THE [MODESTY BLAISE]  O’DONNELL, PETER  SFWA-P
SILVER MISTRESS, THE [MODESTY BLAISE]  O’DONNELL, PETER  SFWA-P
XANADU TALISMAN, THE [@MODESTY BLAISE]  O’DONNELL, PETER  P,SFWA-P
HAWKWOD’S VOYAGE [@MONARCHIES OF GOD #1]  KEARNEY, PAUL  P
HERETIC KINGS, THE [@MONARCHIES OF GOD #2]  KEARNEY, PAUL  P
IRON WARS, THE [@MONARCHIES OF GOD #3]  KEARNEY, PAUL  P
SECOND EMPIRE, THE [@MONARCHIES OF GOD #4]  KEARNEY, PAUL  P
SHIPS FROM THE WEST [@MONARCHIES OF GOD #5]  KEARNEY, PAUL  C/P,P
MONA LISA AWAKENING [MONERE]  SUNNY  LP
AFFAIR OF SORCERERS, AN [MONGO FREDERICKSON 3]  CHESBRO, GEORGE C.  H:2,P
BEASTS OF VALHALLA, THE [MONGO FREDERICKSON 4]  CHESBRO, GEORGE C.  H,P
CITY OF WHISPERING STONE [MONGO FREDERICKSON 2]  CHESBRO, GEORGE C.  C/P,P,H
SHADOW OF A BROKEN MAN [MONGO FREDERICKSON 1]  CHESBRO, GEORGE C.  H,P
CHILDREN OF THE RUNE [MONTE COOK’S ARCANA EVOLVED]  COOK, SUE WEINLEIN  C/PA,PA
DRAGON’S RETURN, THE [MONTE COOK’S ARCANA EVOLVED]  COOK, SUE WEINLEIN  PA
DARE TO GO A-HUNTING [MOON SINGER 4]  NORTON, ANDRE  H:2,SFWA-P,P
EXILES OF THE STARS [MOON SINGER 2]  NORTON, ANDRE  C/P,P,H
FLIGHT IN YIKTOR [MOON SINGER 3]  NORTON, ANDRE  H:2,P
MOON OF THREE RINGS [MOON SINGER 1]  NORTON, ANDRE  C/P:2,P
MOONSINGER [MOON SINGER 1,2]  NORTON, ANDRE  H
MOONRISE [MOONBASE 1]  BOVA, BEN  C/P,P
MOONWAR [MOONBASE 2]  BOVA, BEN  C/P,P
GATHERING DARKNESS [@MOONRUNNER 2]  TOOMBS, JANE  P
UNDER THE SHADOW [@MOONRUNNER 1]  TOOMBS, JANE  C/P:2,P
SEA WITHOUT A SHORE [MOONTIDE AND MAGIC RISE 2]  RUSSELL, SEAN  P
WORLD WITHOUT END [MOONTIDE AND MAGIC RISE 1]  RUSSELL, SEAN  C/P,P
GLASS DRAGONS [MOONWORLDS 2]  MCMULLEN, SEAN  P
VOIDFARER [MOONWORLDS]  MCMULLEN, SEAN  H
VOYAGE OF THE SHADOWMOON [MOONWORLDS 1]  MCMULLEN, SEAN  H
MAN RIDES THROUGH, A [MORDANT’S NEED 2]  DONALDSON, STEPHEN R.  H:2,P
MIRROR OF HER DREAMS, THE [MORDANT’S NEED 1]  DONALDSON, STEPHEN R.  H:2,P,SFWA
MERLIN’S GIFT [MORDRED’S CURSE 2]  MCDOWELL, IAN  P
MORDRED’S CURSE [MORDRED’S CURSE 1]  MCDOWELL, IAN  P
EMPERORS OF THE TWILIGHT [MOREAU 2]  SWANN, S. ANDREW  C/P:2,P
FEARFUL SYMMETRIES [MOREAU 4]  SWANN, S. ANDREW  P
FORESTS OF THE NIGHT [MOREAU 1]  SWANN, S. ANDREW  C/P,SFWA-P,P
SPECTERS OF THE DAWN [MOREAU 3]  SWANN, S. ANDREW  C/P:2,P
BOOK OF MORGAINE, THE [MORGAINE 1,2,3]  CHERRYH, C. J.  H:2
EXILE’S GATE [MORGAINE 4]  CHERRYH, C. J.  H:2,P
FIRES OF AZEROTH [MORGAINE 3]  CHERRYH, C. J.  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 109
TITLE  AUTHOR  CODE
GATE OF IVREL [MORGAINE 1]  CHERRYH, C. J.  C/P:2,P
PUERTA DE IVREL, LA [MORGAINE 1]  CHERRYH, C. J.  FP
WELL OF SHIUAN [MORGAINE 2]  CHERRYH, C. J.  C/P:2,P
MOTHER EARTH [MORGAN OF ALB’NY 2]  MCQUAY, MIKE  P
PURE BLOOD [MORGAN OF ALB’NY 1]  MCQUAY, MIKE  P
DEAD GIRLS’ DANCE, THE [@MORGANVILLE VAMPIRES 2]  CAINE, RACHEL  P
GLASS HOUSES [@MORGANVILLE VAMPIRES 1]  CAINE, RACHEL  P
MORIGU: THE DEAD [MORIGU 1]  PERRY, MARK C.  P
MORIGU: THE DESECRATION [MORIGU 2]  PERRY, MARK C.  P
INCREDIBLE SHRINKING MORK, THE [MORK AND MINDY 2]  WAGNER, ROBIN S.  C/P,P
MORPHODITE, THE [MORPHODITE 1]  FOSTER, M. A.  C/P:2,P
PRESERVER [MORPHODITE 3]  FOSTER, M. A.  C/P:2,P
TRANSFORMER [MORPHODITE 2]  FOSTER, M. A.  C/P:2,P
ANNIHILATION [MORTAL KOMBAT 2]  PREISLER, JEROME  P
MORTAL KOMBAT [MORTAL KOMBAT 1]  DELRIO, MARTIN  P
MORTAL KOMBAT [MORTAL KOMBAT 0]  ROVIN, JEFF  P
GRIPPING HAND, THE [MOTE IN GOD’S EYE 2]  NIVEN, LARRY/POURNELLE, JERRY  H:2,P
MOTE IN GOD’S EYE, THE [MOTE IN GOD’S EYE 1]  NIVEN, LARRY/POURNELLE, JERRY  C/P,P,H:2,SFWA
MOUNTAIN MADE OF LIGHT, THE [MOUNTAIN MADE OF LIGHT 1]  MYERS, EDWARD  P
FIRE AND ICE [@MOUNTAIN MADE OF LIGHT #2]  MYERS, EDWARD  P,SFWA-P
HELL’S GATE [MULTIVERSE 1]  WEBER, DAVID/EVANS, LINDA  H
DEMENTIA [@MUTANT CHRONICLES APOSTLE OF INSANITY 3]  STACKPOLE, MICHAEL A.  P
FRENZY [@MUTANT CHRONICLES APOSTLE OF INSANITY 2]  PRICE, JOHN-ALLEN  P
IN LUNACY [@MUTANT CHRONICLES APOSTLE OF INSANITY 1]  WU, WILLIAM F.  P
MUTANT LEGACY [MUTANT SEASON 4]  HABER, KAREN  P
MUTANT PRIME, THE [MUTANT SEASON 2]  HABER, KAREN  C/P:2,LP
MUTANT SEASON, THE [MUTANT SEASON 1]  HABER, KAREN/SILVERBERG, ROBERT  C/P,LP,P
MUTANT STAR [MUTANT SEASON 3]  HABER, KAREN  P
MUTANTS AMOK [@MUTANTS AMOK #1]  GRANT, MARK  C/P,P
MUTANT HELL [@MUTANTS AMOK #2]  GRANT, MARK  C/P,P
REBEL ATTACK [@MUTANTS AMOK #3]  GRANT, MARK  P
HOLOCAUST HORROR [@MUTANTS AMOK #4]  GRANT, MARK  C/P,P
MY TEACHER FLUNKED THE PLANET [MY TEACHERS 4]  COVILLE, BRUCE  LP
MY TEACHER FRIED MY BRAINS [MY TEACHERS 2]  COVILLE, BRUCE  P
MY TEACHER GLOWS IN THE DARK [MY TEACHERS 3]  COVILLE, BRUCE  P
MY TEACHER IS AN ALIEN [MY TEACHERS 1]  COVILLE, BRUCE  P
PORTS OF CALL [MYRON TANY 1]  VANCE, JACK  H
PORTS OF CALL & LURULU [MYRON TANY 1,2]  VANCE, JACK  H

MYST: THE BOOK OF ATRUS [@MYST 1] MILLER, RAND/MILLER, ROBYN/

WINGROVE, DAVID

C/P,P

MYST: THE BOOK OF D’NI [@MYST 2]  MILLER, RAND/WINGROVE, DAVID  P
MYST: THE BOOK OF TI’ANA [@MYST 3]  MILLER, RAND/WINGROVE, DAVID  H,P
BLACK VESSEL, THE [MYSTARA]  SIMON, MORRIS  P
DARK KNIGHT OF KARAMEIKOS [MYSTARA]  BROWN, TIMOTHY  P
DRAGONKING OF MYSTARA [@MYSTARA DRAGONLORD 2]  GUNNARSSON, THORARINN  P
DRAGONLORD OF MYSTARA [@MYSTARA DRAGONLORD 1]  GUNNARSSON, THORARINN  P
DRAGONMAGE OF MYSTARA [@MYSTARA DRAGONLORD 3]  GUNNARSSON, THORARINN  P
DARKNESS DESCENDS [@MYSTIC REBEL]  SYVERTSEN, RYDER  P
ANOTHER FINE MYTH... [MYTH ADVENTURES 1]  ASPRIN, ROBERT LYNN  C/P:2,LP
CLASS DIS-MYTHED [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  H
HIT OR MYTH [MYTH ADVENTURES 4]  ASPRIN, ROBERT LYNN  C/P:2,LP
LITTLE MYTH MARKER [MYTH ADVENTURES 6]  ASPRIN, ROBERT LYNN  C/P,SFWA-P,P,H
MYTH ADVENTURES [MYTH ADVENTURES 1,2,3,4]  ASPRIN, ROBERT LYNN  H:2
MYTH ALLIANCES (OMNIBUS EDITION) [MYTH ADVENTURES 5,6,7]  ASPRIN, ROBERT LYNN  H:2
MYTH ALLIANCES (SINGLE NOVEL) [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  H,P
MYTH CONCEPTIONS [MYTH ADVENTURES 2]  ASPRIN, ROBERT LYNN  C/P,P,H
MYTH DIRECTIONS [MYTH ADVENTURES 3]  ASPRIN, ROBERT LYNN  C/P,LP,H

MYTH DIRECTIONS [MYTH ADVENTURES 3,4]/ HIT OR MYTH [MYTH ADVENTURES 3,4] ASPRIN, ROBERT LYNN

P

MYTH-GOTTEN GAINS [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  LP
M.Y.T.H. INC. LINK [MYTH ADVENTURES 7/MYTH INC 1]  ASPRIN, ROBERT LYNN  C/P,SFWA-P,H,LP
MYTH-ING PERSONS [MYTH ADVENTURES 5]  ASPRIN, ROBERT LYNN  C/P,P,H
MYTH-ION IMPROBABLE [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN  C/P,P
MYTH-TAKEN IDENTITY [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  LP
MYTH-TOLD TALES [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  P
SOMETHING M.Y.T.H. INC. [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN  C/P,P
SWEET MYTH-TERY OF LIFE [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN  C/P:2,P,H
M.Y.T.H. INC. IN ACTION [MYTH INC 3]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P:2,P,H
MYTH-NOMERS AND IM-PERVECTIONS [MYTH INC 2]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P,P,H
HIGHER MYTHOLOGY [MYTHOLOGY 3]  NYE, JODY LYNN  P
MYTHOLOGY 101 [MYTHOLOGY 1]  NYE, JODY LYNN  P
MYTHOLOGY ABROAD [MYTHOLOGY 2]  NYE, JODY LYNN  C/P,P

TITLE  AUTHOR  CODE
BLOODSTORM [NAGA TEOT 2]  GLADNEY, HEATHER  C/P:2,P
TEOT’S WAR [NAGA TEOT 1]  GLADNEY, HEATHER  P,SFWA-P
BLOOP BOX, THE [NANNY AND THE PROFESSOR 3]  JOHNSTON, WILLIAM  P
NANNY AND THE PROFESSOR [NANNY AND THE PROFESSOR 1]  JOHNSTON, WILLIAM  P
CRESCENT CITY RHAPSODY [NANOTECH]  GOONAN, KATHLEEN ANN  H,P
LIGHT MUSIC [NANOTECH]  GOONAN, KATHLEEN ANN  H
MISSISSIPPI BLUES [NANOTECH VERITY 2]  GOONAN, KATHLEEN ANN  LP
QUEEN CITY JAZZ [NANOTECH VERITY 1]  GOONAN, KATHLEEN ANN  C/P,P
AGAINST THE TIDE OF YEARS [NANTUCKET 2]  STIRLING, S. M.  P
ISLAND IN THE SEA OF TIME [NANTUCKET 1]  STIRLING, S. M.  C/P,P
ON THE OCEANS OF ETERNITY [NANTUCKET 3]  STIRLING, S. M.  C/P:2,P
BURNING TEARS OF SASSURUM [NAPHAR 3]  BAKER, SHARON  P
JOURNEY TO MEMBLIAR [NAPHAR 2]  BAKER, SHARON  C/P,P
QUARRELING, THEY MET THE DRAGON [NAPHAR 1]  BAKER, SHARON  P
HORSE AND HIS BOY, THE [NARNIA 5]  LEWIS, C. S.  C/P:2,P
LAST BATTLE, THE [NARNIA 7]  LEWIS, C. S.  C/P:2,P
LION, THE WITCH AND THE WARDROBE, THE [NARNIA 1]  LEWIS, C. S.  C/P:2,P
MAGICIAN’S NEPHEW, THE [NARNIA 6]  LEWIS, C. S.  C/P:2,P
PRINCE CASPIAN [NARNIA 2]  LEWIS, C. S.  C/P:2,P
SILVER CHAIR, THE [NARNIA 4]  LEWIS, C. S.  C/P:2,P
VOYAGE OF THE "DAWN TREADER", THE [NARNIA 3]  LEWIS, C. S.  C/P:2,P
JUDAS ROSE, THE [NATIVE TONGUE 2]  ELGIN, SUZETTE HADEN  C/P:2,P
NATIVE TONGUE [NATIVE TONGUE 1]  ELGIN, SUZETTE HADEN  C/P:2,P
EARTHSONG [@NATIVE TONGUE #3]  ELGIN, SUZETTE HADEN  C/P:2,P

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 1 [@NAUSICAA PERFECT #1] MIYAZAKI, HAYAO

CX

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 2 [@NAUSICAA PERFECT #2] MIYAZAKI, HAYAO

CX

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 3 [@NAUSICAA PERFECT #3] MIYAZAKI, HAYAO

CX

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 4 [@NAUSICAA PERFECT #4] MIYAZAKI, HAYAO

CX

HOMINIDS [NEANDERTHAL PARALLAX 1]  SAWYER, ROBERT J.  H,P
HUMANS [NEANDERTHAL PARALLAX 2]  SAWYER, ROBERT J.  C/P,P
HYBRIDS [NEANDERTHAL PARALLAX 3]  SAWYER, ROBERT J.  H
KING OF INFINITE SPACE, A [NEAR SPACE]  STEELE, ALLEN  H:2,P
SUMMONER, THE [@NECROMANCER #1]  MARTIN, GAIL Z.  P
BACK FROM THE DEAD [@NECROMUNDA]  KYME, NICK  P
BLOOD ROYAL [@NECROMUNDA]  MCDERMOTT, WILL/RENNIE, GORDON  P
CARDINAL CRIMSON [@NECROMUNDA]  MCDERMOTT, WILL  P
FLESHWORKS [@NECROMUNDA]  SOULBAN, LUCIEN  P
JUNKTION [@NECROMUNDA]  FARRER, MATTHEW  P
LASGUN WEDDING [@NECROMUNDA]  MCDERMOTT, WILL  P
OUTLANDER [@NECROMUNDA]  KEEFE, MATT  P
SALVATION [@NECROMUNDA]  GOTO, C. S.  P
SURVIVAL INSTINCT [@NECROMUNDA]  CHAMBERS, ANDY  P
DEADSPAWN [NECROSCOPE 5]  LUMLEY, BRIAN  P
DEADSPEAK [NECROSCOPE 4]  LUMLEY, BRIAN  P,SFWA-P
LAST AERIE, THE [NECROSCOPE 7/VAMPIRE WORLD 2]  LUMLEY, BRIAN  H
NECROSCOPE [NECROSCOPE 1]  LUMLEY, BRIAN  SFWA-P:2
NECROSCOPE: AVENGERS [NECROSCOPE 13/E-BRANCH 3]  LUMLEY, BRIAN  H
NECROSCOPE: INVADERS [NECROSCOPE 11/E-BRANCH 1]  LUMLEY, BRIAN  P
NECROSCOPE: RESURGENCE [NECROSCOPE 10/LOST YEARS 2]  LUMLEY, BRIAN  P
NECROSCOPE: THE LOST YEARS [NECROSCOPE 9/LOST YEARS 1]  LUMLEY, BRIAN  P
SOURCE, THE [NECROSCOPE 3]  LUMLEY, BRIAN  C/P,SFWA-P:2,P
TOUCH, THE [@NECROSCOPE 14]  LUMLEY, BRIAN  H
VAMPHYRI! [NECROSCOPE 2]  LUMLEY, BRIAN  SFWA-P:2
IN THE CAVES OF EXILE [NEDAO 2]  EMERSON, RU  P,SFWA-P
FIRE CROSSING [NETWORK-CONSORTIUM 2]  FRANKLIN, CHERYL J.  C/P:2,P
GHOST SHADOW [NETWORK-CONSORTIUM 4]  FRANKLIN, CHERYL J.  P
INQUISITOR, THE [NETWORK-CONSORTIUM 3]  FRANKLIN, CHERYL J.  C/P:2,P
LIGHT IN EXILE, THE [NETWORK-CONSORTIUM 1]  FRANKLIN, CHERYL J.  C/P,P
COUNT ZERO [NEUROMANCER 2]  GIBSON, WILLIAM  SFWA-P,H:2,P,SFWA
MONA LISA OVERDRIVE [NEUROMANCER 3]  GIBSON, WILLIAM  C/P,SFWA-P,H:2,P
NEUROMANCER [NEUROMANCER 1]  GIBSON, WILLIAM  H:2,P
GERFALCON [NEUSTRIAN CYCLE 1]  BARRINGER, LESLIE  LP
JORIS OF THE ROCK [NEUSTRIAN CYCLE 2]  BARRINGER, LESLIE  LP
SHY LEOPARDESS [NEUSTRIAN CYCLE 3]  BARRINGER, LESLIE  LP
IRON COUNCIL [NEW CROBUZON]  MIEVILLE, CHINA  H:2
EARTH WINTER [NEW ICE AGE 2]  MORAN, RICHARD  P
EMPIRE OF ICE, THE [NEW ICE AGE 1]  MORAN, RICHARD  H
BOOK OF THE NEW SUN, THE [NEW SUN 1,2,3,4]  WOLFE, GENE  H
CITADEL OF THE AUTARCH, THE [NEW SUN 4]  WOLFE, GENE  C/P,P,H

TITLE  AUTHOR  CODE
CLAW OF THE CONCILIATOR, THE [NEW SUN 2]  WOLFE, GENE  C/P,P,H
SHADOW OF THE TORTURER, THE [NEW SUN 1]  WOLFE, GENE  C/P,P,H
SWORD OF THE LICTOR, THE [NEW SUN 3]  WOLFE, GENE  C/P,SFWA-P,P,H
URTH OF THE NEW SUN, THE [NEW SUN 5]  WOLFE, GENE  C/P,SFWA-P,P,H
PANDORA [NEW TALES 1]  RICE, ANNE  C/P,P,H
VITTORIO THE VAMPIRE [NEW TALES 2]  RICE, ANNE  H:2,P
CITY IN THE STARS, THE [@NEW TOM SWIFT #1]  APPLETON, VICTOR  H,LP
TERROR ON THE MOONS OF JUPITER [@NEW TOM SWIFT #2]  APPLETON, VICTOR  LP
ALIEN PROBE, THE [@NEW TOM SWIFT #3]  APPLETON, VICTOR  H,LP
WAR IN OUTER SPACE, THE [@NEW TOM SWIFT #4]  APPLETON, VICTOR  H,LP
ASTRAL FORTRESS, THE [@NEW TOM SWIFT #5]  APPLETON, VICTOR  C/P,LP
RESCUE MISSION, THE [@NEW TOM SWIFT #6]  APPLETON, VICTOR  C/P,LP
ARK TWO [@NEW TOM SWIFT #7]  APPLETON, VICTOR  H,LP
CRATER OF MYSTERY [@NEW TOM SWIFT #8]  APPLETON, VICTOR  LP
GATEWAY TO DOOM [@NEW TOM SWIFT #9]  APPLETON, VICTOR  LP
INVISIBLE FORCE, THE [@NEW TOM SWIFT #10]  APPLETON, VICTOR  LP
PLANET OF NIGHTMARES [@NEW TOM SWIFT #11]  APPLETON, VICTOR  LP
FORESTS OF THE HEART [NEWFORD]  DE LINT, CHARLES  H
IVORY AND THE HORN, THE [NEWFORD]  DE LINT, CHARLES  C/P,P,H
MEMORY AND DREAM [NEWFORD]  DE LINT, CHARLES  H,P
MOONLIGHT AND VINES [NEWFORD]  DE LINT, CHARLES  C/P,P,H
ONION GIRL, THE [NEWFORD]  DE LINT, CHARLES  H,LP
SOMEPLACE TO BE FLYING [NEWFORD]  DE LINT, CHARLES  C/P:2,P
TAPPING THE DREAM TREE [NEWFORD]  DE LINT, CHARLES  H
TRADER [NEWFORD]  DE LINT, CHARLES  C/P:2,P
WIDDERSHINS [NEWFORD]  DE LINT, CHARLES  H
MONSTER OF MINNESOTA, THE [@NEWS FROM THE EDGE 1]  SUMNER, MARK  P
VAMPIRES OF VERMONT [@NEWS FROM THE EDGE 3]  SUMNER, MARK  P
INSANITY, ILLINOIS [@NEWS FROM THE EDGE #2]  SUMNER, MARK  P
ALIEN TONGUE [NEXT WAVE 2]  LEIGH, STEPHEN  C/P:2,P
MISSING MATTER, THE [NEXT WAVE 3]  MCDONOUGH, THOMAS R.  C/P,P
RED GENESIS [NEXT WAVE 1]  SYKES, S. C.  C/P:2,P
CHALLENGER’S HOPE [NICHOLAS SEAFORT 2]  FEINTUCH, DAVID  C/P:2,P
CHILDREN OF HOPE [NICHOLAS SEAFORT 7]  FEINTUCH, DAVID  H
FISHERMAN’S HOPE [NICHOLAS SEAFORT 4]  FEINTUCH, DAVID  C/P:2,P
MIDSHIPMAN’S HOPE [NICHOLAS SEAFORT 1]  FEINTUCH, DAVID  C/P:2,P
PATRIARCH’S HOPE [NICHOLAS SEAFORT 6]  FEINTUCH, DAVID  C/P,P,H
PRISONER’S HOPE [NICHOLAS SEAFORT 3]  FEINTUCH, DAVID  C/P:2,P
SEAFORT’S HOPE [NICHOLAS SEAFORT 1,2]  FEINTUCH, DAVID  H
VOICES OF HOPE [NICHOLAS SEAFORT 5]  FEINTUCH, DAVID  C/P:2,P
POINT OF DREAMS [NICOLAS RATHE 2]  SCOTT, MELISSA/BARNETT, LISA A.  H
POINT OF HOPES [NICOLAS RATHE 1]  SCOTT, MELISSA/BARNETT, LISA A.  H
FIRSTFLIGHT [NICOLE SHEA 1]  CLAREMONT, CHRIS  C/P:2,SFWA-P,P
GROUNDED! [NICOLE SHEA 2]  CLAREMONT, CHRIS  C/P:2,SFWA-P:2,P
SUNDOWNER [NICOLE SHEA 3]  CLAREMONT, CHRIS  P
MINES OF BEHEMOTH, THE [NIFFT 2]  SHEA, MICHAEL  P
NIFFT THE LEAN [NIFFT 1]  SHEA, MICHAEL  C/P,P
HEART OF VALOR [NIGHT OF THE SOLSTICE 2]  SMITH, L. J.  H
ART OF THE SWORD, THE [@NIGHT-THREADS 5]  EMERSON, RU  C/P,P
CALLING OF THE THREE, THE [NIGHT-THREADS 1]  EMERSON, RU  SFWA-P:2
CRAFT OF LIGHT, THE [@NIGHT-THREADS 4]  EMERSON, RU  P
SCIENCE OF POWER, THE [@NIGHT-THREADS 6]  EMERSON, RU  C/P,P
TWO IN HIDING, THE [@NIGHT-THREADS #2]  EMERSON, RU  C/P,P
ONE LAND, ONE DUKE [@NIGHT-THREADS #3]  EMERSON, RU  C/P,P
NIGHT WHISTLERS, THE [@NIGHT WHISTLERS #1]  TREVOR, DAN  P
TERMINATE [@NIGHT WHISTLERS #3]  TREVOR, DAN  P
LEGEND OF NIGHTFALL, THE [NIGHTFALL 1]  REICHERT, MICKEY ZUCKER  C/P,P
LUCK IN THE SHADOWS [NIGHTRUNNER 1]  FLEWELLING, LYNN  C/P,P
STALKING DARKNESS [NIGHTRUNNER 2]  FLEWELLING, LYNN  C/P,P
TRAITOR’S MOON [NIGHTRUNNER 3]  FLEWELLING, LYNN  C/P,P
CONFEDERATION HANDBOOK, THE [NIGHT’S DAWN]  HAMILTON, PETER F.  P
NAKED GOD, THE [NIGHT’S DAWN 3]  HAMILTON, PETER F.  H
NEUTRONIUM ALCHEMIST [NIGHT’S DAWN 2]  HAMILTON, PETER F.  H
NEUTRONIUM ALCHEMIST 1: CONSOLIDATION [NIGHT’S DAWN 2]  HAMILTON, PETER F.  C/P,P
NEUTRONIUM ALCHEMIST 2: CONFLICT [NIGHT’S DAWN 2]  HAMILTON, PETER F.  P
REALITY DYSFUNCTION [NIGHT’S DAWN 1]  HAMILTON, PETER F.  H
REALITY DYSFUNCTION 1: EMERGENCE [NIGHT’S DAWN 1]  HAMILTON, PETER F.  C/P:2,P
REALITY DYSFUNCTION 2: EXPANSION [NIGHT’S DAWN 1]  HAMILTON, PETER F.  C/P,P
AGENTS OF LIGHT AND DARKNESS [NIGHTSIDE 2]  GREEN, SIMON R.  P
EVERYBODY COMES TO THE NIGHTSIDE [NIGHTSIDE 1,2,3]  GREEN, SIMON R.  H
HELL TO PAY [NIGHTSIDE 7]  GREEN, SIMON R.  P
HEX AND THE CITY [NIGHTSIDE 4]  GREEN, SIMON R.  P
NIGHTINGALE’S LAMENT [NIGHTSIDE 3]  GREEN, SIMON R.  P

The MITSFS Seriesdex  September 24, 2007  Page 113
TITLE  AUTHOR  CODE
PATHS NOT TAKEN [NIGHTSIDE 5]  GREEN, SIMON R.  P
SHARPER THAN A SERPENT’S TOOTH [NIGHTSIDE 6]  GREEN, SIMON R.  P
SOMETHING FROM THE NIGHTSIDE [NIGHTSIDE 1]  GREEN, SIMON R.  C/P,P
GIDEON [@NIGHTWALKERS]  FRANK, JACQUELYN  P
JACOB [@NIGHTWALKERS]  FRANK, JACQUELYN  P
DAY WATCH [NIGHTWATCH 2]  LUKYANENKO, SERGEI  H,LP
NIGHT WATCH [NIGHTWATCH 1]  LUKYANENKO, SERGEI  H,LP
TWILIGHT WATCH [NIGHTWATCH 3]  LUKYANENKO, SERGEI  LP
STRANGELOVE GAMBIT, THE [@NIKOLAI DANTE]  BISHOP, DAVID  P
MRS. FRISBY AND THE RATS OF NIMH [NIMH 1]  O’BRIEN, ROBERT C.  P
R-T, MARGARET, AND THE RATS OF NIMH [NIMH 3]  CONLY, JANE LESLIE  H
ROUSE A SLEEPING CAT [NIMNESTL 1]  CRAWFORD, DAN  P
SURE DEATH OF A MOUSE, THE [NIMNESTL 2]  CRAWFORD, DAN  C/P,P
WILD DOG AND LONE, A [NIMNESTL 3]  CRAWFORD, DAN  P
STAR OF THE MORNING [NINE KINGDOMS]  KURLAND, LYNN  LP
DHAMPIR [NOBLE DEAD SERIES 1 1]  HENDEE, BARB/HENDEE, J. C.  P
REBEL FAY [NOBLE DEAD SERIES 1 5]  HENDEE, BARB/HENDEE, J. C.  H:2
SISTER OF THE DEAD [NOBLE DEAD SERIES 1 3]  HENDEE, BARB/HENDEE, J. C.  P
THIEF OF LIVES [NOBLE DEAD SERIES 1 2]  HENDEE, BARB/HENDEE, J. C.  P
TRAITOR TO THE BLOOD [NOBLE DEAD SERIES 1 4]  HENDEE, BARB/HENDEE, J. C.  H
RESTORATION [NOEL KEDRAN 1]  DALTON, SEAN  P
TERMINATION [NOEL KEDRAN 3]  DALTON, SEAN  C/P,P
TURNCOAT [NOEL KEDRAN 2]  DALTON, SEAN  P
NORBY AND THE COURT JESTER [NORBY]  ASIMOV, JANET/ASIMOV, ISAAC  P
NORBY AND THE OLDEST DRAGON [NORBY]  ASIMOV, JANET/ASIMOV, ISAAC  SFWA-P
NORBY AND YOBO’S GREAT ADVENTURE [NORBY 4]  ASIMOV, JANET/ASIMOV, ISAAC  C/P,SFWA-P:2,P
NORBY CHRONICLES, THE [NORBY 1]  ASIMOV, JANET/ASIMOV, ISAAC  C/P:2,P
NORBY DOWN TO EARTH [NORBY 5]  ASIMOV, JANET/ASIMOV, ISAAC  P,SFWA-P:2
NORBY: ROBOT FOR HIRE [NORBY 2]  ASIMOV, JANET/ASIMOV, ISAAC  C/P:3,P
NORBY THROUGH TIME AND SPACE [NORBY]  ASIMOV, JANET/ASIMOV, ISAAC  C/P:2,P
BEYOND THE TOMORROW MOUNTAINS [NOREN 2]  ENGDAHL, SYLVIA LOUISE  H
DOORS OF THE UNIVERSE, THE [NOREN 3]  ENGDAHL, SYLVIA LOUISE  H
THIS STAR SHALL ABIDE [NOREN 1]  ENGDAHL, SYLVIA LOUISE  H
AMERICAN ZONE, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  LP
GALLATIN DIVERGENCE, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P,P
NAGASAKI VECTOR, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P:2,P
PROBABILITY BROACH, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P:2,P

PROBABILITY BROACH, THE (ORIGINAL)

[NORTH AMERICAN CONFEDERACY] SMITH, L. NEIL

H,LP

TAFLAK LYSANDRA [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  P
TOM PAINE MARU [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P:2,P
VENUS BELT, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P,P
NORTHWORLD [NORTHWORLD 1]  DRAKE, DAVID  C/P:2,SFWA-P:2,P
NORTHWORLD JUSTICE [NORTHWORLD 3]  DRAKE, DAVID  C/P:2,SFWA-P:2,P
NORTHWORLD TRILOGY [NORTHWORLD 1,2,3]  DRAKE, DAVID  C/P,P
NORTHWORLD VENGEANCE [NORTHWORLD 2]  DRAKE, DAVID  C/P,SFWA-P:2,P
BATTERIES NOT INCLUDED [NOT QUITE HUMAN 1]  MCEVOY, SETH  P
TERROR AT PLAY [NOT QUITE HUMAN 5]  MCEVOY, SETH  P
FIRE SANCTUARY [NUALA 1]  KIMBRIEL, KATHERINE ELISKA  C/P,P
FIRES OF NUALA [NUALA 2]  KIMBRIEL, KATHERINE ELISKA  C/P:2,P
HIDDEN FIRES [NUALA 3]  KIMBRIEL, KATHERINE ELISKA  C/P:2,P
CONTEXT [NULAPEIRON 2]  MEANEY, JOHN  H
PARADOX [NULAPEIRON 1]  MEANEY, JOHN  P
RESOLUTION [NULAPEIRON 3]  MEANEY, JOHN  H

TITLE  AUTHOR  CODE
CROWN DISOWNED, A [OAK,YEW,ASH,AND ROWAN 3]  NORTON, ANDRE/MILLER, SASHA  H
DRAGON BLADE [OAK,YEW,ASH,AND ROWAN 4]  NORTON, ANDRE/MILLER, SASHA  H
KNIGHT OR KNAVE [OAK,YEW,ASH,AND ROWAN 2]  NORTON, ANDRE/MILLER, SASHA  H
TO THE KING A DAUGHTER [OAK,YEW,ASH,AND ROWAN 1]  NORTON, ANDRE/MILLER, SASHA  H:2
DARK LORD, THE [OATH OF EMPIRE 4]  HARLAN, THOMAS  H
GATE OF FIRE, THE [OATH OF EMPIRE 2]  HARLAN, THOMAS  H,P
SHADOW OF ARARAT, THE [OATH OF EMPIRE 1]  HARLAN, THOMAS  C/P,P
SHADOW OF ARARAT, THE (PROOF) [OATH OF EMPIRE 1]  HARLAN, THOMAS  SR-LP
STORM OF HEAVEN, THE [OATH OF EMPIRE 3]  HARLAN, THOMAS  H
ASHLING [OBERNEWTYN 3]  CARMODY, ISOBELLE  H
FARSEEKERS, THE [OBERNEWTYN 2]  CARMODY, ISOBELLE  H,P
OBERNEWTYN [OBERNEWTYN 1]  CARMODY, ISOBELLE  P
OBERNEWTYN & THE FARSEEKERS [OBERNEWTYN 1,2]  CARMODY, ISOBELLE  H
OUTSTRETCHED SHADOW, THE [OBSIDIAN 1]  LACKEY, MERCEDES/MALLORY, JAMES  H
TO LIGHT A CANDLE [OBSIDIAN 2]  LACKEY, MERCEDES/MALLORY, JAMES  H
WHEN DARKNESS FALLS [OBSIDIAN 3]  LACKEY, MERCEDES/MALLORY, JAMES  H
DRAGON SOCIETY, THE [OBSIDIAN CHRONICLES 2]  WATT-EVANS, LAWRENCE  H
DRAGON VENOM [OBSIDIAN CHRONICLES 3]  WATT-EVANS, LAWRENCE  H
DRAGON WEATHER [OBSIDIAN CHRONICLES 1]  WATT-EVANS, LAWRENCE  H
OMNIVORE [OF MAN AND MANTA 1]  ANTHONY, PIERS  C/P,P,H
ORN [OF MAN AND MANTA 2]  ANTHONY, PIERS  C/P,P,H
OX [OF MAN AND MANTA 3]  ANTHONY, PIERS  C/P,P,H
HEIR OF AUTUMN [OHNDARIEN 1]  CARWYN, GILES/FAHNESTOCK, TODD  H
MISTRESS OF WINTER [OHNDARIEN 2]  CARWYN, GILES/FAHNESTOCK, TODD  H
GHOST BRIGADES, THE [OLD MAN’S WAR 2]  SCALZI, JOHN  H
LAST COLONY, THE [OLD MAN’S WAR 3]  SCALZI, JOHN  H
OLD MAN’S WAR [OLD MAN’S WAR 1]  SCALZI, JOHN  H:2
SAGAN DIARY, THE [OLD MAN’S WAR]  SCALZI, JOHN  H
HUNTERS’ HAUNT, THE [OMAR 2]  DUNCAN, DAVE  C/P:2,P
PLACE AMONG THE FALLEN, A [@OMARAN #1]  COLE, ADRIAN  H,SFWA,P
THRONE OF FOOLS [@OMARAN #2]  COLE, ADRIAN  P
KING OF LIGHT AND SHADOWS, THE [@OMARAN #3]  COLE, ADRIAN  P
GODS IN ANGER, THE [@OMARAN #4]  COLE, ADRIAN  P
OMEGA SUB [@OMEGA SUB #1]  CAMERON, J. D.  P
COMMAND DECISION [@OMEGA SUB #2]  CAMERON, J. D.  P
CITY OF FEAR [@OMEGA SUB #3]  CAMERON, J. D.  P
BLOOD TIDE [@OMEGA SUB #4]  CAMERON, J. D.  P
DEATH DIVE [@OMEGA SUB #5]  CAMERON, J. D.  P
RAVEN RISING [@OMEGA SUB #6]  CAMERON, J. D.  P
ABOMINATION, THE [OMEN 5]  MCGILL, GORDON  C/P,P
ARMAGEDDON 2000 [OMEN 4]  MCGILL, GORDON  P
DAMIEN — OMEN II [OMEN 2]  HOWARD, JOSEPH  P
FINAL CONFLICT, THE [OMEN 3]  MCGILL, GORDON  P
OMEN, THE [OMEN 1]  SELTZER, DAVID  P
ROYAL FOUR, THE [ONE LAND 2]  STOUT, AMY  P
BARBED ROSE, THE [ONE ROSE 2]  DAYTON, GAIL  LP
COMPASS ROSE, THE [ONE ROSE 1]  DAYTON, GAIL  LP
BEYOND THE GAP [OPENING OF THE WORLD]  TURTLEDOVE, HARRY  H
OPERATION CHAOS [OPERATION CHAOS 1]  ANDERSON, POUL  C/P:2,SFWA,P
OPERATION LUNA [OPERATION CHAOS 2]  ANDERSON, POUL  H,P
ANAZ-VOOHRI [OPERATION PLEIADES]  SCHARTZ, VIJAYA  P
MASKED INVASION, THE [OPERATOR 5 #1]  STEELE, CURTIS  SR-P
INVISIBLE EMPIRE, THE [OPERATOR 5 #2]  STEELE, CURTIS  SR-P,C/P
YELLOW SCOURGE, THE [OPERATOR 5 #3]  STEELE, CURTIS  SR-P
ORACLE BETRAYED, THE [ORACLE 1]  FISHER, CATHERINE  P
BODYGUARD OF LIGHTNING [ORCS FIRST BLOOD 1]  NICHOLLS, STAN  LP
FIRST BLOOD [ORCS FIRST BLOOD 1,2,3]  NICHOLLS, STAN  H
LEGION OF THUNDER [ORCS FIRST BLOOD 2]  NICHOLLS, STAN  LP
ELECTRIC GENE MACHINE, THE [ORGANIC FUTURE]  EASTON, THOMAS A.  LP
GREENHOUSE [ORGANIC FUTURE]  EASTON, THOMAS A.  C/P:2,P
SEEDS OF DESTINY [ORGANIC FUTURE]  EASTON, THOMAS A.  P
SPARROWHAWK [ORGANIC FUTURE]  EASTON, THOMAS A.  C/P,P
TOWER OF THE GODS [ORGANIC FUTURE]  EASTON, THOMAS A.  P
WOODSMAN [ORGANIC FUTURE]  EASTON, THOMAS A.  C/P:2,P
CAT’S GAMBIT [ORIANI 2]  GADALLAH, LESLIE  C/P,P
CAT’S PAWN [ORIANI 1]  GADALLAH, LESLIE  P
BLACK STORM RISING [ORIANUS CREATION 1]  BURTON, D. SHANE  LP
WHEN WORLDS RAGE [ORIANUS CREATION 2]  BURTON, D. SHANE  LP
ORION [ORION 1]  BOVA, BEN  C/P,SFWA-P:2,P,H
ORION AMONG THE STARS [ORION 4]  BOVA, BEN  C/P,SFWA,P,H
ORION AND THE CONQUEROR [ORION 3]  BOVA, BEN  C/P,P
ORION IN THE DYING TIME [ORION 2]  BOVA, BEN  C/P:2,P,H
VENGEANCE OF ORION [ORION 2]  BOVA, BEN  H:2,P

TITLE  AUTHOR  CODE
RING OF ANGELICA, THE [@ORLANDO FURIOSO 1]  ARIOSTO, LUDOVICO  C/P,P
HARLEQUIN’S DANCE [@OROKON #1]  ARDEN, TOM  P
KING AND QUEEN OF SWORDS, THE [@OROKON #2]  ARDEN, TOM  P
SULTAN OF THE MOON AND STARS [@OROKON #3]  ARDEN, TOM  P
SISTERHOOD OF THE BLUE STORM [@OROKON #4]  ARDEN, TOM  H,P
EMPRESS OF THE ENDLESS DREAM [@OROKON #5]  ARDEN, TOM  P
ORON [@ORON #1]  SMITH, DAVID C.  C/P,P
SORCERER’S SHADOW, THE [@ORON #2]  SMITH, DAVID C.  P
MOSUTHA’S MAGIC [@ORON #3]  SMITH, DAVID C.  P
VALLEY OF OGRUM, THE [@ORON #4]  SMITH, DAVID C.  P
GHOST ARMY, THE [@ORON #5]  SMITH, DAVID C.  P
FUGITIVES OF CHAOS [ORPHANS OF CHAOS 2]  WRIGHT, JOHN C.  H
ORPHANS OF CHAOS [ORPHANS OF CHAOS 1]  WRIGHT, JOHN C.  H:2
TITANS OF CHAOS [ORPHANS OF CHAOS 3]  WRIGHT, JOHN C.  H
IN THE NIGHT GARDEN [@ORPHAN’S TALES #1]  VALENTE, CATHERYNNE M  LP
ANCIENT LIGHT [ORTHE 2]  GENTLE, MARY  C/P,P,H
GOLDEN WITCHBREED [ORTHE 1]  GENTLE, MARY  C/P,P,H
AMBER WIZARD, THE [OSSERIAN 1]  FORBES, DAVID  P
LAND LEVIATHAN, THE [OSWALD BASTABLE 2]  MOORCOCK, MICHAEL  C/P,SFWA,P,H
STEEL TSAR, THE [OSWALD BASTABLE 3]  MOORCOCK, MICHAEL  C/P,P
WARLORD OF THE AIR, THE [OSWALD BASTABLE 1]  MOORCOCK, MICHAEL  C/P:2,P
BAD DAY FOR ALI BABA, A [OTHER SINBAD 2]  GARDNER, CRAIG SHAW  C/P,SFWA-P:2,P
LAST ARABIAN NIGHT, THE [OTHER SINBAD 3]  GARDNER, CRAIG SHAW  C/P,SFWA-P,P
OTHER SINBAD, THE [OTHER SINBAD 1]  GARDNER, CRAIG SHAW  C/P,SFWA-P:2,P
CITY OF GOLDEN SHADOW [@OTHERLAND #1]  WILLIAMS, TAD  H,P
RIVER OF BLUE FIRE [@OTHERLAND #2]  WILLIAMS, TAD  H,P
MOUNTAIN OF BLACK GLASS [@OTHERLAND #3]  WILLIAMS, TAD  C/P,P,H
SEA OF SILVER LIGHT [@OTHERLAND #4]  WILLIAMS, TAD  H,P
OTHERS, THE [OTHERS (BONANNO) 1]  BONANNO, MARGARET WANDER  H,SFWA
OTHERWHERE [OTHERS (BONANNO) 2]  BONANNO, MARGARET WANDER  H
OTHERWISE [OTHERS (BONANNO) 3]  BONANNO, MARGARET WANDER  H
DEMON YOU KNOW, THE [OTHERS (WARREN)]  WARREN, CHRISTINE  P
SHE’S NO FAERIE PRINCESS [OTHERS (WARREN) 2]  WARREN, CHRISTINE  P
WOLF AT THE DOOR [OTHERS (WARREN) 1]  WARREN, CHRISTINE  P
CHANGELING [OTHERWORLD]  GALENORN, YASMINE  P
DARKTHUNDER’S WAY [OTHERWORLD 3]  DEITZ, TOM  C/P:2,P
DREAMSEEKER’S ROAD [OTHERWORLD 6]  DEITZ, TOM  H,P
FIRESHAPER’S DOOM [OTHERWORLD 2]  DEITZ, TOM  C/P:2,P
GRYPHON KING, THE [OTHERWORLD]  DEITZ, TOM  C/P:2,P
LANDSLAYER’S LAW [OTHERWORLD]  DEITZ, TOM  P
STONESKIN’S REVENGE [OTHERWORLD 5]  DEITZ, TOM  C/P,P
SUNSHAKER’S WAR [OTHERWORLD 4]  DEITZ, TOM  C/P:2,P
WINDMASTER’S BANE [OTHERWORLD 1]  DEITZ, TOM  C/P:2,P
WITCHLING [OTHERWORLD]  GALENORN, YASMINE  C/P,P
ACROSS THE NIGHTINGALE FLOOR [@OTORI #1]  HEARN, LIAN  H:2
ACROSS THE NIGHTINGALE FLOOR (PROOF) [@OTORI #1]  HEARN, LIAN  SR-LP
GRASS FOR HIS PILLOW [@OTORI #2]  HEARN, LIAN  H:2
BRILLIANCE OF THE MOON [@OTORI #3]  HEARN, LIAN  H
HARSH CRY OF THE HERON, THE [@OTORI #4]  HEARN, LIAN  H
MAX ON EARTH [@OUT OF THIS WORLD #1]  KAYE, MARILYN  P
YANKED! [@OUT OF TIME #1]  KRESS, NANCY  P
TIGER IN THE SKY [@OUT OF TIME #2]  FINCH, SHEILA  P
GAME OF WORLDS, THE [@OUT OF TIME #3]  ALLEN, ROGER MACBRIDE  P
BEWARE THE METAL CHILDREN [@OUTER LIMITS]  PEEL, JOHN  LP
INNOCENT, THE [@OUTER LIMITS]  PEEL, JOHN  LP
INVADERS, THE [@OUTER LIMITS]  PEEL, JOHN  LP
LOST, THE [@OUTER LIMITS]  PEEL, JOHN  LP
ZANTI MISFITS, THE [@OUTER LIMITS]  PEEL, JOHN  LP
BREATH OF SNOW AND ASHES, A [OUTLANDER (GABALDON) 6]  GABALDON, DIANA  H
DRAGONFLY IN AMBER [OUTLANDER (GABALDON) 2]  GABALDON, DIANA  C/P,P
OUTLANDER [OUTLANDER (GABALDON) 1]  GABALDON, DIANA  H:2,P
VOYAGER [OUTLANDER (GABALDON) 3]  GABALDON, DIANA  C/P,P,H
CAPTIVITY [OUTLANDER (SALTERBERG) 1]  SALTERBERG, B. J.  LP
EQUINOX ZERO [@OUTLANDERS]  AXLER, JAMES  P
EVIL ABYSS [@OUTLANDERS]  AXLER, JAMES  P
HELL RISING [@OUTLANDERS]  AXLER, JAMES  P
ICEBLOOD [@OUTLANDERS]  AXLER, JAMES  P
PARALLAX RED [@OUTLANDERS]  AXLER, JAMES  P
SUN LORD [@OUTLANDERS]  AXLER, JAMES  P
SEA OF PLAGUE [@OUTLANDERS HEART OF THE WORLD 2]  AXLER, JAMES  P
TALON AND FANG [OUTLANDERS HEART OF THE WORLD 1]  AXLER, JAMES  P
INTERSTELLAR INTRIGUE (PROOF) [@OUTREACH INVESTIGATIONS #1]  FOOTE, KEITH D.  SR-P
FEAST OF THE KING’S SHADOW (COMPLETE) [OUTREMER 3,4]  BRENCHLEY, CHAZ  P

TITLE  AUTHOR  CODE
TOWER OF THE KING’S DAUGHTER (COMPLETE) [OUTREMER 1,2]  BRENCHLEY, CHAZ  P
DEVIL IN THE DUST, THE [@OUTREMER #1]  BRENCHLEY, CHAZ  P
TOWER OF THE KING’S DAUGHTER [@OUTREMER #2]  BRENCHLEY, CHAZ  C/P,P
DARK WAY TO GLORY, A [@OUTREMER #3]  BRENCHLEY, CHAZ  P
FEAST OF THE KING’S SHADOW [@OUTREMER #4]  BRENCHLEY, CHAZ  P
HAND OF THE KING’S EVIL [@OUTREMER #5]  BRENCHLEY, CHAZ  P
END OF ALL ROADS, THE [@OUTREMER #6]  BRENCHLEY, CHAZ  P
OUTRIDER, THE [@OUTRIDER #1]  HARDING, RICHARD  P
FIRE AND ICE [@OUTRIDER #2]  HARDING, RICHARD  P
BLOOD HIGHWAY [@OUTRIDER #3]  HARDING, RICHARD  P
BAY CITY BURNOUT [@OUTRIDER #4]  HARDING, RICHARD  P
MARVELOUS LAND OF OZ, THE [OZ 2]  BAUM, L. FRANK  H:2,LP
RETURN TO OZ [OZ]  VINGE, JOAN D.  C/P,P
WONDERFUL WIZARD OF OZ, THE [OZ 1]  BAUM, L. FRANK  H:3,LP
WIZARD OF OZ, THE [OZ #1]  BAUM, L. FRANK  H,@P,@C/P:2
LAND OF OZ, THE [OZ #2]  BAUM, L. FRANK  H,@P,@C/P:2
OZMA OF OZ [OZ #3]  BAUM, L. FRANK  H:2,@P,@C/P
DOROTHY AND THE WIZARD IN OZ [OZ #4]  BAUM, L. FRANK  H,@P,@C/P
ROAD TO OZ, THE [OZ #5]  BAUM, L. FRANK  C/P,@P,@C/P,H
EMERALD CITY OF OZ, THE [OZ #6]  BAUM, L. FRANK  H:2,@P,@C/P:2
PATCHWORK GIRL OF OZ, THE [OZ #7]  BAUM, L. FRANK  H,@P
TIK-TOK OF OZ [OZ #8]  BAUM, L. FRANK  C/P,@P,H
SCARECROW OF OZ, THE [OZ #9]  BAUM, L. FRANK  H:2,@P
RINKITINK IN OZ [OZ #10]  BAUM, L. FRANK  C/P,@P,H
LOST PRINCESS OF OZ, THE [OZ #11]  BAUM, L. FRANK  H,@P
TIN WOODMAN OF OZ, THE [OZ #12]  BAUM, L. FRANK  H,@P,@C/P
MAGIC OF OZ, THE [OZ #13]  BAUM, L. FRANK  H,@P
GLINDA OF OZ [OZ #14]  BAUM, L. FRANK  C/P,@P,H
TAKERS, THE [@OZ CHRONICLES #1]  RIDLEY, R. W.  LP
DELON CITY [@OZ CHRONICLES #2]  RIDLEY, R. W.  LP
AND THEN THERE’LL BE FIREWORKS [OZARK 3]  ELGIN, SUZETTE HADEN  C/P,SFWA,P,H
GRAND JUBILEE, THE [OZARK 2]  ELGIN, SUZETTE HADEN  C/P,SFWA,P,H
OZARK TRILOGY, THE [OZARK 1,2,3]  ELGIN, SUZETTE HADEN  H:2
TWELVE FAIR KINGDOMS [OZARK 1]  ELGIN, SUZETTE HADEN  C/P,SFWA,P

TITLE  AUTHOR  CODE
BOY FROM THE BURREN, THE [PAINTER 1]  GILLULY, SHEILA  P
DEED OF PAKSENARRION, THE [PAKSENARRION 1,2,3]  MOON, ELIZABETH  H,SFWA
DIVIDED ALLEGIANCE [PAKSENARRION 2]  MOON, ELIZABETH  C/P:3,SFWA-P,P
OATH OF GOLD [PAKSENARRION 3]  MOON, ELIZABETH  C/P:2,P
SHEEPFARMER’S DAUGHTER [PAKSENARRION 1]  MOON, ELIZABETH  C/P:3,SFWA-P:2,P
PALADINS [PALADINS 1]  ROSENBERG, JOEL  P
KNIGHT MOVES [@PALADINS #2]  ROSENBERG, JOEL  H
PANDORA’S CHILDREN [PANDORA 2]  LANCE, KATHRYN  P
PANDORA’S GENES [PANDORA 1]  LANCE, KATHRYN  C/P,P
MASK AND THE SORCERESS, THE [PANDRAGORE 2]  JONES, DENNIS  H
STONE AND THE MAIDEN, THE [PANDRAGORE 1]  JONES, DENNIS  C/P,P,H
IMPERIUM WITHOUT END [@PANGAEA #1]  MASON, LISA  P
IMPERIUM AFIRE [@PANGAEA #2]  MASON, LISA  P
BOOK OF THE BEAST, THE [PARADYS 2]  LEE, TANITH  H
BOOK OF THE DAMNED [PARADYS 1]  LEE, TANITH  LP
BOOK OF THE DEAD, THE [PARADYS 3]  LEE, TANITH  H,LP
BOOK OF THE MAD, THE [PARADYS 4]  LEE, TANITH  H,LP
SECRET BOOKS OF PARADYS I & II, THE [PARADYS 1,2]  LEE, TANITH  H:2
SECRET BOOKS OF PARADYS III & IV, THE [PARADYS 3,4]  LEE, TANITH  H
COMPLETE PARATIME, THE [PARATIME POLICE]  PIPER, H. BEAM  H,LP
GREAT KINGS’ WAR [PARATIME POLICE]  GREEN, ROLAND J./CARR, JOHN F.  C/P:2,P
GUNPOWDER GOD [PARATIME POLICE]  PIPER, H. BEAM  P
LORD KALVAN OF OTHERWHEN [PARATIME POLICE]  PIPER, H. BEAM  C/P,P,H
PARATIME [PARATIME POLICE]  PIPER, H. BEAM  C/P:2,P
ASH OCK [PARATWA 2]  HINZ, CHRISTOPHER  H,P
LIEGE-KILLER [PARATWA 1]  HINZ, CHRISTOPHER  C/P:2,P,H
PARATWA, THE [PARATWA 3]  HINZ, CHRISTOPHER  SFWA
CRASH DELUXE [PARRISH PLESSIS 3]  DE PIERRES, MARIANNE  P
NYLON ANGEL [PARRISH PLESSIS 1]  DE PIERRES, MARIANNE  P
CRYSTALS OF AIR AND WATER [PARSINA 3]  GOLDIN, STEPHEN  P
SHRINE OF THE DESERT MAGE [PARSINA 1]  GOLDIN, STEPHEN  P
STORYTELLER AND THE JAHN, THE [PARSINA 2]  GOLDIN, STEPHEN  P
BLOOD AND DREAMS [PARSIVAL 4]  MONACO, RICHARD  P
FINAL QUEST, THE [PARSIVAL 3]  MONACO, RICHARD  C/P,SFWA,P
GRAIL WAR, THE [PARSIVAL 2]  MONACO, RICHARD  C/P,P
PARSIVAL OR A KNIGHT’S TALE [PARSIVAL 1]  MONACO, RICHARD  C/P,P,H
PATTERNS OF CHAOS, VOLUME ONE, THE [PATTERNS OF CHAOS 1,2]  INGRID, CHARLES  P
PATTERNS OF CHAOS, VOLUME TWO, THE [PATTERNS OF CHAOS 3,4]  INGRID, CHARLES  P
RADIUS OF DOUBT [@PATTERNS OF CHAOS #1]  INGRID, CHARLES  P
PATH OF FIRE [@PATTERNS OF CHAOS #2]  INGRID, CHARLES  P
DOWNFALL MATRIX, THE [@PATTERNS OF CHAOS #3]  INGRID, CHARLES  P
SOUL FIRE [@PATTERNS OF CHAOS #4]  INGRID, CHARLES  P
IN YOUR DREAMS [PAUL CARPENTER 2]  HOLT, TOM  H
PORTABLE DOOR, THE [PAUL CARPENTER 1]  HOLT, TOM  H
CUP OF THANATOS, THE [PAUL HOLTON 2]  HUNT, CHARLOTTE  P
GILDED SARCOPHAGUS, THE [PAUL HOLTON 1]  HUNT, CHARLOTTE  P
AGAINST ALL ENEMIES [PAUL SINCLAIR 4]  HEMRY, JOHN G.  P
BURDEN OF PROOF [PAUL SINCLAIR 2]  HEMRY, JOHN G.  P
JUST DETERMINATION, A [PAUL SINCLAIR 1]  HEMRY, JOHN G.  P
RULE OF EVIDENCE [PAUL SINCLAIR 3]  HEMRY, JOHN G.  P
STAR BORN [PAX]  NORTON, ANDRE  C/P:2,P,DD299
STARS ARE OURS!, THE [PAX]  NORTON, ANDRE  C/P,H,DD121,P
MOUNTAIN WALKS, THE [PEACE COMPANY]  GREEN, ROLAND J.  P
PEACE COMPANY [PEACE COMPANY]  GREEN, ROLAND J.  P
THESE GREEN FOREIGN HILLS [PEACE COMPANY]  GREEN, ROLAND J.  C/P,P
RING OF FIVE DRAGONS, THE [PEARL 1]  LUSTBADER, ERIC VAN  H:2,P
RING OF FIVE DRAGONS, THE (PROOF) [PEARL 1]  LUSTBADER, ERIC VAN  SR-LP
VEIL OF A THOUSAND TEARS, THE [PEARL 2]  LUSTBADER, ERIC VAN  H
VEIL OF A THOUSAND TEARS, THE (PROOF) [PEARL 2]  LUSTBADER, ERIC VAN  H:2,SR-LP
DOORS OF DEATH AND LIFE [PEARL OF IMMORTALITY 2]  CLOUGH, BRENDA W.  H
HOW LIKE A GOD [PEARL OF IMMORTALITY 1]  CLOUGH, BRENDA W.  H,P
PEGASUS IN FLIGHT [PEGASUS 2]  MCCAFFREY, ANNE  H:2,SR-H,P
PEGASUS IN SPACE [PEGASUS 3]  MCCAFFREY, ANNE  H:2,P
TO RIDE PEGASUS [PEGASUS 1]  MCCAFFREY, ANNE  C/P:2,P
WINGS OF PEGASUS, THE [PEGASUS 1,2]  MCCAFFREY, ANNE  H
AMBUSH OF SHADOWS, AN [PELBAR 5]  WILLIAMS, PAUL O.  C/P:2,P
BREAKING OF NORTHWALL, THE [PELBAR 1]  WILLIAMS, PAUL O.  C/P:2,P
DOME IN THE FOREST, THE [PELBAR 3]  WILLIAMS, PAUL O.  C/P:3,P
ENDS OF THE CIRCLE, THE [PELBAR 2]  WILLIAMS, PAUL O.  C/P:2,P
FALL OF THE SHELL, THE [PELBAR 4]  WILLIAMS, PAUL O.  C/P:2,P
SONG OF THE AXE, THE [PELBAR 6]  WILLIAMS, PAUL O.  C/P:2,P
SWORD OF FORBEARANCE, THE [PELBAR 7]  WILLIAMS, PAUL O.  P

The MITSFS Seriesdex  September 24, 2007  Page 120
TITLE  AUTHOR  CODE

AT THE EARTH’S CORE [PELLUCIDAR 1,2,3]/ PELLUCIDAR [PELLUCIDAR 1,2,3]/ TANAR OF PELLUCIDAR [PELLUCIDAR 1,2,3] BURROUGHS, EDGAR RICE

LP

AT THE EARTH’S CORE [PELLUCIDAR #1]  BURROUGHS, EDGAR RICE  H:2,@P
PELLUCIDAR [PELLUCIDAR #2]  BURROUGHS, EDGAR RICE  C/P,@P,H
TANAR OF PELLUCIDAR [PELLUCIDAR #3]  BURROUGHS, EDGAR RICE  C/P:2,@P
BACK TO THE STONE AGE [PELLUCIDAR #5]  BURROUGHS, EDGAR RICE  C/P,@P,H
LAND OF TERROR [PELLUCIDAR #6]  BURROUGHS, EDGAR RICE  C/P:2,@P
SAVAGE PELLUCIDAR [PELLUCIDAR #7]  BURROUGHS, EDGAR RICE  C/P:2,@P
POWER AND THE PROPHET, THE [PELMEN THE POWERSHAPER 3]  HUGHES, ROBERT DON  C/P:2,P
PROPHET OF LAMATH, THE [PELMEN THE POWERSHAPER 1]  HUGHES, ROBERT DON  C/P:2,P
WIZARD IN WAITING, THE [PELMEN THE POWERSHAPER 2]  HUGHES, ROBERT DON  C/P:2,P
BODDEKKER’S DEMONS [PEMBROKE HALL 2]  FAUST, JOE CLIFFORD  P
FERMAN’S DEVILS [PEMBROKE HALL 1]  FAUST, JOE CLIFFORD  P
HANDLING IT [PEMBROKE HALL 1,2]  FAUST, JOE CLIFFORD  H
GRAIL [PENDRAGON 5]  LAWHEAD, STEPHEN R.  H,P
PENDRAGON [PENDRAGON 4]  LAWHEAD, STEPHEN R.  C/P,P
TALIESIN [@PENDRAGON #1]  LAWHEAD, STEPHEN R.  C/P:2,P
MERLIN [@PENDRAGON #2]  LAWHEAD, STEPHEN R.  P
ARTHUR [@PENDRAGON #3]  LAWHEAD, STEPHEN R.  P

PEOPLE OF THE EARTH [PEOPLE OF THE WOLF 1] GEAR, W. MICHAEL/

GEAR, KATHLEEN O’NEAL

C/P,P

PEOPLE OF THE FIRE [PEOPLE OF THE WOLF 2] GEAR, W. MICHAEL/

GEAR, KATHLEEN O’NEAL

SFWA-P

DON OF THE DEAD [PEPPER MARTIN]  DANIELS, CASEY  P
ODALISQUE [PERCHERON 1]  MCINTOSH, FIONA  LP
LIGHTNING THIEF, THE [@PERCY JACKSON & THE OLYMPIANS #1]  RIORDAN, RICK  H
OUT OF THE SILENT PLANET [PERELANDRA 1]  LEWIS, C. S.  C/P,P,H
PERELANDRA [PERELANDRA 2]  LEWIS, C. S.  C/P:2,P,H
THAT HIDEOUS STRENGTH [PERELANDRA 3]  LEWIS, C. S.  C/P:2,P
TORTURED PLANET, THE (PERELANDRA 3 ABRIDGED) [PERELANDRA 3]  LEWIS, C. S.  C/P:2,P
ALL THE WEYRS OF PERN [PERN]  MCCAFFREY, ANNE  H:2,P
ATLAS OF PERN, THE [PERN]  FONSTAD, KAREN WYNN  R/REF-H,C/REF-H:3
DOLPHIN’S BELL, THE [PERN]  MCCAFFREY, ANNE  H:2
DOLPHINS OF PERN, THE [PERN]  MCCAFFREY, ANNE  H:2,P
DRAGONLOVER’S GUIDE TO PERN, THE [PERN]  NYE, JODY LYNN/MCCAFFREY, ANNE  R/REF-H
DRAGON’S FIRE [PERN]  MCCAFFREY, ANNE/MCCAFFREY, TODD  H
DRAGON’S KIN [PERN]  MCCAFFREY, ANNE/MCCAFFREY, TODD  H
DRAGON’S KIN (PROOF) [PERN]  MCCAFFREY, ANNE/MCCAFFREY, TODD  SR-LP
DRAGONSBLOOD [PERN]  MCCAFFREY, TODD  H
DRAGONSDAWN [PERN]  MCCAFFREY, ANNE  H:2,P
DRAGONSEYE [PERN]  MCCAFFREY, ANNE  H:2,P
FIRST FALL [PERN]  MCCAFFREY, ANNE  C/P,P,H:2
GIFT OF DRAGONS, A [PERN]  MCCAFFREY, ANNE  H
MASTERHARPER OF PERN, THE [PERN]  MCCAFFREY, ANNE  H:2,P
MORETA: DRAGONLADY OF PERN [PERN]  MCCAFFREY, ANNE  H:2,P
NERILKA’S STORY [PERN]  MCCAFFREY, ANNE  H:2,P
RENEGADES OF PERN, THE [PERN]  MCCAFFREY, ANNE  H:2,P
RESCUE RUN [PERN]  MCCAFFREY, ANNE  H:2
SKIES OF PERN, THE [PERN]  MCCAFFREY, ANNE  H,P
DRAGONFLIGHT [PERN DRAGONRIDERS 1]  MCCAFFREY, ANNE  C/P:2,P
DRAGONQUEST [PERN DRAGONRIDERS 2]  MCCAFFREY, ANNE  C/P:2,P
DRAGONRIDERS OF PERN [PERN DRAGONRIDERS 1,2,3]  MCCAFFREY, ANNE  H:2
WHITE DRAGON, THE [PERN DRAGONRIDERS 3]  MCCAFFREY, ANNE  C/P,P,H
DRAGONDRUMS [PERN HARPER HALL 3]  MCCAFFREY, ANNE  H:2,P
DRAGONSINGER [PERN HARPER HALL 2]  MCCAFFREY, ANNE  H:2,P
DRAGONSONG [PERN HARPER HALL 1]  MCCAFFREY, ANNE  H:2,P
HARPER HALL OF PERN, THE [PERN HARPER HALL 1,2,3]  MCCAFFREY, ANNE  H:2
AGENTS OF DESTRUCTION [PERRY RHODAN 134]  BRAND, KURT  SR-LP
ATOM FIRE ON MECHANICA [PERRY RHODAN 121]  DARLTON, CLARK  SR-LP
BEASTS BELOW, THE [PERRY RHODAN 128]  MAHR, KURT  SR-LP
BETWEEN THE GALAXIES [PERRY RHODAN 119]  MAHR, KURT  SR-LP
BLITZKRIEG GALACTICA [PERRY RHODAN 129]  BRAND, KURT  SR-LP
DEADMEN SHOULDN’T DIE [PERRY RHODAN 132]  DARLTON, CLARK  SR-LP
FORTRESS IN TIME [PERRY RHODAN 123]  MAHR, KURT  SR-LP
GUNS OF EVERBLACK, THE [PERRY RHODAN 126]  SCHEER, K. H.  SR-LP
HUMANS KEEP OUT! [PERRY RHODAN 135]  VOLTZ, WILLIAM  SR-LP
IN THE CENTER OF THE GALAXY (NOVEL) [PERRY RHODAN]  ACE  PA
KILLERS FROM HYPERSPACE [PERRY RHODAN 120]  VOLTZ, WILLIAM  SR-LP
PERIL UNLIMITED [PERRY RHODAN 130]  BRAND, KURT  SR-LP

PERRY RHODAN: DIE GROESSTE SCIENCE FICTION-SERIE DER WELT (CD)

[PERRY RHODAN] PABEL MOEWIG KG SR-MM

The MITSFS Seriesdex  September 24, 2007  Page 121
TITLE  AUTHOR  CODE

PHANTOM HORDE [PERRY RHODAN 137]/ STAR MAN #1 THROUGH #5 [PERRY RHODAN 137] MASTER PUBLICATIONS

SR-LP

ROBOT INVITATION, THE [PERRY RHODAN 136]  SCHEER, K. H.  SR-LP
ROBOTS, BOMBS AND MUTANTS [PERRY RHODAN 125]  VOLTZ, WILLIAM  SR-LP
SENTINELS OF SOLITUDE [PERRY RHODAN 127]  DARLTON, CLARK  SR-LP
SINISTER POWER, THE [PERRY RHODAN 124]  BRAND, KURT  SR-LP
STATION OF THE INVISIBLES [PERRY RHODAN 133]  MAHR, KURT  SR-LP
VOLUNTEER FOR FRAGO [PERRY RHODAN 122]  BRAND, KURT  SR-LP
WORLD WITHOUT MERCY [PERRY RHODAN 131]  VOLTZ, WILLIAM  SR-LP
ENTERPRISE STARDUST [@PERRY RHODAN #1]  ACE  PA,C/PA
OPERATION "ASTREE" (FRENCH) [@PERRY RHODAN #1]  SCHEER, K. H.  FPR
RADIANT DOME, THE [@PERRY RHODAN #2]  ACE  PA,C/PA
TERRE A PEUR, LA (FRENCH) [@PERRY RHODAN #2]  SCHEER, K. H.  FPR
GALACTIC ALARM [@PERRY RHODAN #3]  ACE  PA,C/PA
MILICE DES MUTANTS, LA (FRENCH) [@PERRY RHODAN #3]  SCHEER, K. H.  FPR
SCHATZKAMMER DER STERNE [@PERRY RHODAN #3]  BRAND, KURT  FPR
AARDE VALT AAN!, DE (DUTCH) [@PERRY RHODAN #4]  SCHEER, K. H./DARLTON, CLARK  FPR
BASES SUR VENUS (FRENCH) [@PERRY RHODAN #4]  SCHEER, K. H.  FPR
INVASION FROM SPACE [@PERRY RHODAN #4]  ACE  PA,C/PA
ATOOM-ALARM! (DUTCH) [@PERRY RHODAN #5]  MAHR, KURT  FPR
VAINQUEURS DE VEGA, LES (FRENCH) [@PERRY RHODAN #5]  SCHEER, K. H./DARLTON, CLARK  FPR
VEGA SECTOR, THE [@PERRY RHODAN #5]  ACE  PA,C/PA
SECRET OF THE TIME VAULT, THE [@PERRY RHODAN #6]  ACE  PA,C/PA
TOCHTER DES ROBOTERS, DIE [@PERRY RHODAN #6]  MAHR, KURT  FPR
DIE ZEITSPRINGER [@PERRY RHODAN #7]  EWERS, H. G.  FPR
FORTERESSE DES SIX LUNES, LA (FRENCH) [@PERRY RHODAN #7]  SCHEER, K. H./DARLTON, CLARK  FPR
FORTRESS OF THE SIX MOONS [@PERRY RHODAN #7]  ACE  PA,C/PA
AM RAND DES BLAUEN NEBELS [@PERRY RHODAN #8]  KNEIFEL, HANS  FPR
GALACTIC RIDDLE, THE [@PERRY RHODAN #8]  ACE  PA,C/PA
QUEST THROUGH SPACE AND TIME [@PERRY RHODAN #9]  ACE  PA,C/PA
GHOSTS OF GOL, THE [@PERRY RHODAN #10]  ACE  PA,C/PA
PLANET OF THE DYING SUN, THE [@PERRY RHODAN #11]  ACE  PA,C/PA
QUETE COSMIQUE, LA (FRENCH) [@PERRY RHODAN #11]  SCHEER, K. H./DARLTON, CLARK  FPR
REBELS OF TUGLAN, THE [@PERRY RHODAN #12]  ACE  PA,C/PA
IMMORTAL UNKNOWN, THE [@PERRY RHODAN #13]  ACE  PA,C/PA
VENUS IN DANGER [@PERRY RHODAN #14]  ACE  PA,C/PA
ESCAPE TO VENUS [@PERRY RHODAN #15]  ACE  PA,C/PA
PHANTOM-STATION [@PERRY RHODAN #16]  EWERS, H. G.  FPR
SECRET BARRIER X [@PERRY RHODAN #16]  ACE  C/PA,PA
VENUS TRAP, THE [@PERRY RHODAN #17]  ACE  PA
MENACE OF THE MUTANT MASTER [@PERRY RHODAN #18]  ACE  C/PA,PA
MUTANTS VS. MUTANTS [@PERRY RHODAN #19]  ACE  PA
THRALL OF HYPNO, THE [@PERRY RHODAN #20]  ACE  C/PA,PA
COSMIC DECOY, THE [@PERRY RHODAN #21]  ACE  PA,C/PA
FLEET OF THE SPRINGERS, THE [@PERRY RHODAN #22]  ACE  PA,C/PA
PERIL ON ICE PLANET [@PERRY RHODAN #23]  ACE  PA,C/PA
INFINITY FLIGHT [@PERRY RHODAN #24]  ACE  PA,C/PA
SNOWMAN IN FLAMES [@PERRY RHODAN #25]  ACE  PA,C/PA
COSMIC TRAITOR [@PERRY RHODAN #26]  ACE  PA,C/PA
PLANET OF THE GODS [@PERRY RHODAN #27]  ACE  PA,C/PA
GUCKY UND DIE MORDWESPEN [@PERRY RHODAN #28]  DARLTON, CLARK  FPR
PLAGUE OF OBLIVION, THE [@PERRY RHODAN #28]  ACE  PA,C/PA
WORLD GONE MAD, A [@PERRY RHODAN #29]  ACE  C/PA,PA
TO ARKON! [@PERRY RHODAN #30]  ACE  PA,C/PA
REALM OF THE TRI-PLANETS [@PERRY RHODAN #31]  ACE  PA,C/PA
CHALLENGE OF THE UNKNOWN [@PERRY RHODAN #32]  ACE  C/PA,PA
SCHATTEN DES KRISTALLENEN TODES, DIE [@PERRY RHODAN #32]  KNEIFEL, HANS  FPR
GIANT’S PARTNER, THE [@PERRY RHODAN #33]  ACE  PA,C/PA
SOS: SPACESHIP TITAN! [@PERRY RHODAN #34]  ACE  PA,C/PA
BEWARE THE MICROBOTS [@PERRY RHODAN #35]  ACE  PA,C/PA
MAN AND MONSTER [@PERRY RHODAN #36]  ACE  PA,C/PA
EPIDEMIC CENTER: ARALON [@PERRY RHODAN #37]  ACE  C/PA,PA
PROJECT: EARTHSAVE [@PERRY RHODAN #38]  ACE  PA,C/PA
SILENCE OF GOM, THE [@PERRY RHODAN #39]  ACE  PA,C/PA
RED EYE OF BETELGEUSE [@PERRY RHODAN #40]  ACE  PA,C/PA
EARTH DIES, THE [@PERRY RHODAN #41]  ACE  PA,C/PA
TIME’S LONELY ONE [@PERRY RHODAN #42]  ACE  PA,C/PA
LIFE HUNT [@PERRY RHODAN #43]  ACE  PA,C/PA
PFLANZEN DES TODES, DIE [@PERRY RHODAN #43]  KNEIFEL, HANS  FPR
PSEUDO ONE, THE [@PERRY RHODAN #44]  ACE  PA,C/PA
UNKNOWN SECTOR: MILKY WAY [@PERRY RHODAN #45]  ACE  PA,C/PA
AGAIN: ATLAN! [@PERRY RHODAN #46]  ACE  PA,C/PA
HOLLENTANZ DER MARIONETTEN [@PERRY RHODAN #47]  KNEIFEL, HANS  FPR

The MITSFS Seriesdex  September 24, 2007  Page 122
TITLE  AUTHOR  CODE
SHADOW OF THE MUTANT MASTER [@PERRY RHODAN #47]  ACE  PA,C/PA
DEAD LIVE, THE [@PERRY RHODAN #48]  ACE  PA,C/PA
SOLAR ASSASSINS [@PERRY RHODAN #49]  ACE  PA,C/PA
ATTACK FROM THE UNSEEN [@PERRY RHODAN #50]  ACE  PA,C/PA
RETURN FROM THE VOID [@PERRY RHODAN #51]  ACE  PA,C/PA
FORTRESS ATLANTIS [@PERRY RHODAN #52]  ACE  C/PA:2,PA
SPYBOT! [@PERRY RHODAN #53]  ACE  PA,C/PA
BLUE DWARFS, THE [@PERRY RHODAN #54]  ACE  PA,C/PA
MICRO-TECHS, THE [@PERRY RHODAN #55]  ACE  C/PA,PA
PRISONER OF TIME [@PERRY RHODAN #56]  ACE  C/PA:2,PA
TOUCH OF ETERNITY, A [@PERRY RHODAN #57]  ACE  C/PA,PA
GUARDIANS, THE [@PERRY RHODAN #58]  ACE  C/PA:2,PA
VERBOTENE SONNENSYSTEM, DAS [@PERRY RHODAN #58]  DARLTON, CLARK  FPR
INTERLUDE ON SILIKO 5 [@PERRY RHODAN #59]  ACE  C/PA,PA
DIMENSION SEARCH [@PERRY RHODAN #60]  ACE  PA
DEATH WAITS IN SEMISPACE [@PERRY RHODAN #61]  ACE  C/PA,PA
PLANETENKONIG, DER [@PERRY RHODAN #61]  MAHR, KURT  FPR
LAST DAYS OF ATLANTIS, THE [@PERRY RHODAN #62]  ACE  PA
TIGRIS LEAPS, THE [@PERRY RHODAN #63]  ACE  PA,C/PA
AMBASSADORS FROM AURIGEL, THE [@PERRY RHODAN #64]  ACE  PA,C/PA
MANNER FUR LACERTAE [@PERRY RHODAN #64]  DARLTON, CLARK  FPR
DIE WELT DER GLUCKSELIGEN [@PERRY RHODAN #65]  EWERS, H. G.  FPR
RENEGADES OF THE FUTURE [@PERRY RHODAN #65]  ACE  C/PA,PA
HORROR, THE [@PERRY RHODAN #66]  ACE  C/PA,PA
SUPERNOVA [@PERRY RHODAN #66]  MAHR, KURT  FPR
CRIMSON UNIVERSE [@PERRY RHODAN #67]  ACE  PA
ENDLOSE ALPTRAUM, DER [@PERRY RHODAN #67]  VLCEK, ERNST  FPR
UNDER THE STARS OF DRUUFON [@PERRY RHODAN #68]  ACE  C/PA,PA
BONDS OF ETERNITY, THE [@PERRY RHODAN #69]  ACE  PA
THORA’S SACRIFICE [@PERRY RHODAN #70]  ACE  PA
ATOM HELL OF GRAUTIER, THE [@PERRY RHODAN #71]  ACE  PA
CAVES OF THE DRUUFS [@PERRY RHODAN #72]  ACE  C/PA,PA
SPACESHIP OF ANCESTORS [@PERRY RHODAN #73]  ACE  PA,C/PA
CHECKMATE: UNIVERSE [@PERRY RHODAN #74]  ACE  C/PA,PA
PLANET TOPIDE, PLEASE REPLY! [@PERRY RHODAN #75]  ACE  C/PA,PA
RECRUITS FOR ARKON [@PERRY RHODAN #76]  ACE  C/PA,PA
CONFLICT CENTER: NAATOR [@PERRY RHODAN #77]  ACE  C/PA,PA
POWER KEY [@PERRY RHODAN #78]  ACE  PA
SLEEPERS, THE [@PERRY RHODAN #79]  ACE  PA
COLUMBUS AFFAIR, THE [@PERRY RHODAN #80]  ACE  PA
PUCKY’S GREATEST HOUR [@PERRY RHODAN #81]  ACE  PA
ATLAN IN DANGER [@PERRY RHODAN #82]  ACE  C/PA,PA
ERNST ELLERT RETURNS! [@PERRY RHODAN #83]  ACE  PA
SECRET MISSION: MOLUK [@PERRY RHODAN #84]  ACE  PA
ENEMY IN THE DARK [@PERRY RHODAN #85]  ACE  PA
SATANS UNIVERSUM [@PERRY RHODAN #85]  VLCEK, ERNST  FPR
BLAZING SUN [@PERRY RHODAN #86]  ACE  C/PA:2,PA
STARLESS REALM, THE [@PERRY RHODAN #87]  ACE  PA
MYSTERY OF THE ANTI, THE [@PERRY RHODAN #88]  ACE  C/PA,PA
POWER’S PRICE [@PERRY RHODAN #89]  ACE  C/PA,PA
UNLEASHED POWERS [@PERRY RHODAN #90]  ACE  C/PA,PA
FRIEND TO MANKIND [@PERRY RHODAN #91]  ACE  PA
TARGET STAR, THE [@PERRY RHODAN #92]  ACE  C/PA:2,PA
VAGABOND OF SPACE [@PERRY RHODAN #93]  ACE  C/PA:2,PA
ACTION: DIVISION 3 [@PERRY RHODAN #94]  ACE  C/PA,PA
PLASMA MONSTER, THE [@PERRY RHODAN #95]  ACE  C/PA,PA
HORN: GREEN [@PERRY RHODAN #96]  ACE  C/PA,PA
PHANTOM FLEET [@PERRY RHODAN #97]  ACE  PA
IDOL FROM PASSA, THE [@PERRY RHODAN #98]  ACE  PA
BLUE SYSTEM, THE [@PERRY RHODAN #99]  ACE  C/PA,PA
DESERT OF DEATH’S DOMAIN [@PERRY RHODAN #100]  ACE  C/PA,PA
BLOCKADE: LEPSO [@PERRY RHODAN #101]  ACE  PA
SPOOR OF THE ANTIS [@PERRY RHODAN #102]  ACE  C/PA,PA
FALSE FRONT [@PERRY RHODAN #103]  ACE  PA
MAN WITH TWO FACES, THE [@PERRY RHODAN #104]  ACE  C/PA,PA
WONDERFLOWER OF UTIK [@PERRY RHODAN #105]  ACE  PA
CALLER FROM ETERNITY [@PERRY RHODAN #106]  ACE  PA
EMPEROR AND THE MONSTER, THE [@PERRY RHODAN #107]  ACE  PA
DUEL UNDER THE DOUBLE SUN [@PERRY RHODAN #108]  ACE  PA

STOLEN SPACEFLEET, THE [@PERRY RHODAN #109]/ SGT. ROBOT [PERRY RHODAN 110] ACE

PA

SEEDS OF RUIN [@PERRY RHODAN #111]/ PLANET MECHANICA [PERRY RHODAN 112] ACE C/PA,PA

The MITSFS Seriesdex  September 24, 2007  Page 123
TITLE  AUTHOR  CODE

HERITAGE OF THE LIZARD PEOPLE [@PERRY RHODAN #113]/ DEATH’S DEMAND [PERRY RHODAN 114] ACE

C/PA,PA

SABOTEURS IN A-1 [@PERRY RHODAN #115]/ PSYCHO DUEL, THE [PERRY RHODAN 116] ACE

C/PA,PA

SAVIOR OF THE EMPIRE [@PERRY RHODAN #117]/ SHADOWS ATTACK, THE [PERRY RHODAN 118] ACE

PA

DIE TODESMASCHINE [@PERRY RHODAN #119]  EWERS, H. G.  FPR
INSELN DER BEWAHRUNG, DIE [@PERRY RHODAN #130]  KNEIFEL, HANS  FPR
ABTEILUNG PSI [@PERRY RHODAN #131]  MAHR, KURT  FPR
DIE MACHT DER SCHATTEN [@PERRY RHODAN #133]  EWERS, H. G.  FPR
GROSSEN VON AGOPP, DIE [@PERRY RHODAN #138]  FRANCIS, H. G.  FPR
SIE KAMEN VON KENDARKOS [@PERRY RHODAN #142]  DARLTON, CLARK  FPR
FLUCHT DER ANDROIDEN [@PERRY RHODAN #147]  KNEIFEL, HANS  FPR
BRENNENDEN SCHIFFE, DIE [@PERRY RHODAN ANDROMEDA 1]  ANTON, UWE  FPR
ARK OF THE STARS [@PERRY RHODAN LEMURA #1]  BORSCH, FRANK  P
DRITTE MACHT, DIE [@PERRY RHODAN SILBERBANDE #1]  PABEL MOEWIG KG  FPR
ALTMUTANTEN, DIE [@PERRY RHODAN SILBERBANDE #65]  PABEL MOEWIG KG  FPR
ANTI-UNIVERSUM [@PERRY RHODAN SILBERBANDE #68]  PABEL MOEWIG KG  FPR
HYPERSEUCHE, DIE [@PERRY RHODAN SILBERBANDE #69]  PABEL MOEWIG KG  FPR
GEHIRN IN FESSELN [@PERRY RHODAN SILBERBANDE #70]  PABEL MOEWIG KG  FPR
ERBE DER YULOCS, DAS [@PERRY RHODAN SILBERBANDE #71]  PABEL MOEWIG KG  FPR
KONTAKTE MIT DER EWIGKEIT [@PERRY RHODAN SILBERBANDE #72]  PABEL MOEWIG KG  FPR
SCHACH DER FINSTERNIS [@PERRY RHODAN SILBERBANDE #73]  PABEL MOEWIG KG  FPR
KONZIL DER SIEBEN [@PERRY RHODAN SILBERBANDE #74]  PABEL MOEWIG KG  FPR
LAREN, DIE [@PERRY RHODAN SILBERBANDE #75]  PABEL MOEWIG KG  FPR
RAUMSCHIFF ERDE [@PERRY RHODAN SILBERBANDE #76]  PABEL MOEWIG KG  FPR
IM MAHLSTROM DER STERNE [@PERRY RHODAN SILBERBANDE #77]  PABEL MOEWIG KG  FPR
SUCHE NACH DER ERDE [@PERRY RHODAN SILBERBANDE #78]  PABEL MOEWIG KG  FPR
MENSCHHEIT AM SCHEIDEWEG [@PERRY RHODAN SILBERBANDE #80]  PABEL MOEWIG KG  FPR
APHILE [@PERRY RHODAN SILBERBANDE #81]  PABEL MOEWIG KG  FPR

POWER LINES [PETAYBEE 2] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2,P

POWER PLAY [PETAYBEE 3] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2,P

POWERS THAT BE [PETAYBEE 1] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2,P

PETER AND THE SHADOW THIEVES [PETER 2]  BARRY, DAVE/PEARSON, RIDLEY  H
PETER AND THE STARCATCHERS [PETER 1]  BARRY, DAVE/PEARSON, RIDLEY  H
COVENANT OF JUSTICE, A [PHAESTOR 2]  GERROLD, DAVID  C/P,P
UNDER THE EYE OF GOD [PHAESTOR 1]  GERROLD, DAVID  C/P:2,P
CITIZEN PHAID [PHAID 2]  FARREN, MICK  C/P:2,P
PHAID THE GAMBLER [PHAID 1]  FARREN, MICK  C/P:2,P
GOLDEN CIRCLE, THE [PHANTOM 5]  SHAWN, WILLIAM S.  P
HYDRA MONSTER, THE [PHANTOM 8]  SHAWN, WILLIAM S.  P
KILLER’S TOWN [PHANTOM 9]  FALK, LEE  P
MYSTERY OF THE SEA HORSE, THE [PHANTOM 7]  SHAWN, WILLIAM S.  P
PHANTOM AND THE SCORPIA MENACE, THE [PHANTOM 3]  COPPER, BASIL  P
STORY OF THE VEILED LADY, THE [PHANTOM 4]  SHAWN, WILLIAM S.  P
AUGUSTA, THE FIRST [@PHENWICK #1]  KIMBROUGH, KATHERYN  P
JANE, THE COURAGEOUS [@PHENWICK #2]  KIMBROUGH, KATHERYN  P
MARGARET, THE FAITHFUL [@PHENWICK #3]  KIMBROUGH, KATHERYN  P
PATRICIA, THE BEAUTIFUL [@PHENWICK #4]  KIMBROUGH, KATHERYN  P
RACHEL, THE POSESSESED [@PHENWICK #5]  KIMBROUGH, KATHERYN  P
SUSANNAH, THE RIGHTEOUS [@PHENWICK #6]  KIMBROUGH, KATHERYN  P
REBECCA, THE MYSTERIOUS [@PHENWICK #7]  KIMBROUGH, KATHERYN  P
JOANNE, THE UNPREDICTABLE [@PHENWICK #8]  KIMBROUGH, KATHERYN  P
OLIVIA, THE TORMENTED [@PHENWICK #9]  KIMBROUGH, KATHERYN  P
HARRIET, THE HAUNTED [@PHENWICK #10]  KIMBROUGH, KATHERYN  P
NANCY, THE DARING [@PHENWICK #11]  KIMBROUGH, KATHERYN  P
MARCIA, THE INNOCENT [@PHENWICK #12]  KIMBROUGH, KATHERYN  P
KATE, THE CURIOUS [@PHENWICK #13]  KIMBROUGH, KATHERYN  P
ILENE, THE SUPERSTITIOUS [@PHENWICK #14]  KIMBROUGH, KATHERYN  P
MILLIJOY, THE DETERMINED [@PHENWICK #15]  KIMBROUGH, KATHERYN  P
BARBARA, THE VALIANT [@PHENWICK #16]  KIMBROUGH, KATHERYN  P
RUTH, THE UNSUSPECTING [@PHENWICK #17]  KIMBROUGH, KATHERYN  P
OPHELIA, THE ANXIOUS [@PHENWICK #18]  KIMBROUGH, KATHERYN  P
DOROTHY, THE TERRIFIED [@PHENWICK #19]  KIMBROUGH, KATHERYN  P
ANN, THE GENTLE [@PHENWICK #20]  KIMBROUGH, KATHERYN  P
NELLIE, THE OBVIOUS [@PHENWICK #21]  KIMBROUGH, KATHERYN  P
ISABELLE, THE FRANTIC [@PHENWICK #22]  KIMBROUGH, KATHERYN  P
EVELYN, THE AMBITIOUS [@PHENWICK #23]  KIMBROUGH, KATHERYN  P
POLLY, THE WORRIED [@PHENWICK #25]  KIMBROUGH, KATHERYN  P
YVONNE, THE CONFIDENT [@PHENWICK #26]  KIMBROUGH, KATHERYN  P

The MITSFS Seriesdex  September 24, 2007  Page 124
TITLE  AUTHOR  CODE
JOYCE, THE BELOVED [@PHENWICK #27]  KIMBROUGH, KATHERYN  P
AUGUSTA, THE SECOND [@PHENWICK #28]  KIMBROUGH, KATHERYN  P
CAROL, THE PURSUED [@PHENWICK #29]  KIMBROUGH, KATHERYN  P
KATHERINE, THE RETURNED [@PHENWICK #30]  KIMBROUGH, KATHERYN  P
PEGGY, THE CONCERNED [@PHENWICK #31]  KIMBROUGH, KATHERYN  P
OLGA, THE DISILLUSIONED [@PHENWICK #32]  KIMBROUGH, KATHERYN  P
URSULA, THE PROUD [@PHENWICK #34]  KIMBROUGH, KATHERYN  P
LETITIA, THE DREAMER [@PHENWICK #35]  KIMBROUGH, KATHERYN  P
LAURA, THE IMPERILED [@PHENWICK #38]  KIMBROUGH, KATHERYN  P
IRIS, THE BEWITCHED [@PHENWICK #39]  KIMBROUGH, KATHERYN  P
CONSCIOUSNESS PLAGUE, THE [PHIL D’AMATO 2]  LEVINSON, PAUL  H
PIXEL EYE, THE [PHIL D’AMATO 3]  LEVINSON, PAUL  H
SILK CODE, THE [PHIL D’AMATO 1]  LEVINSON, PAUL  H:2,P
DARK MESSIAH [@PHOENIX #1]  ALEXANDER, DAVID  P
METALSTORM [@PHOENIX #4]  ALEXANDER, DAVID  P
HOUSE OF THE WOLF [PHOENIX LEGACY 3]  WREN, M. K.  C/P:2,P
SHADOW OF THE SWAN [PHOENIX LEGACY 2]  WREN, M. K.  C/P:2,P
SWORD OF THE LAMB [PHOENIX LEGACY 1]  WREN, M. K.  C/P:2,P
THIEVES OF LIGHT [PHOTON]  HUDSON, MICHAEL  P
FOR THE GLORY [@PHOTON #1]  PETERS, DAVID  P
HIGH STAKES [@PHOTON #2]  PETERS, DAVID  P
IN SEARCH OF MOM [@PHOTON #3]  PETERS, DAVID  P
THIS IS YOUR LIFE, BHODILI [@PHOTON #4]  PETERS, DAVID  P
EXILE [@PHOTON #5]  PETERS, DAVID  P
SKIN DEEP [@PHOTON #6]  PETERS, DAVID  P
NO PHULE LIKE AN OLD PHULE [PHULE 5]  ASPRIN, ROBERT LYNN/HECK, PETER J.  P
PHULE AND HIS MONEY, A [PHULE 3]  ASPRIN, ROBERT LYNN/HECK, PETER J.  C/P,P
PHULE ME TWICE [PHULE 4]  ASPRIN, ROBERT LYNN/HECK, PETER J.  C/P:2,P
PHULE’S COMPANY [PHULE 1]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P:2,P
PHULE’S ERRAND [PHULE 6]  ASPRIN, ROBERT LYNN/HECK, PETER J.  C/P,P
PHULE’S PARADISE [PHULE 2]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P:2,P
AFTER THE FACT [PILGRIM 2]  SABERHAGEN, FRED  C/P,SFWA-P,P
PILGRIM [PILGRIM 1,2]  SABERHAGEN, FRED  P
PYRAMIDS [PILGRIM 1]  SABERHAGEN, FRED  C/P:2,P
EYES OF GOD, THE [PINCH 2]  KREIGHBAUM, MARK  P
PALACE [PINCH 1]  KERR, KATHARINE/KREIGHBAUM, MARK  P
DRAGON’S BLOOD [PIT DRAGONS 1]  YOLEN, JANE  C/P,P
HEART’S BLOOD [PIT DRAGONS 2]  YOLEN, JANE  C/P,P
PIT DRAGON TRILOGY, THE [PIT DRAGONS 1,2,3]  YOLEN, JANE  H
SENDING OF DRAGONS, A [PIT DRAGONS 3]  YOLEN, JANE  H,P
TORMENT [@PLANE SCAPE]  VALLESE, RAY/VALLESE, VALERIE  P
BLOOD HOSTAGES [@PLANESCAPE BLOOD WARS #1]  KING, J. ROBERT  P
ABYSSAL WARRIORS [@PLANESCAPE BLOOD WARS #2]  KING, J. ROBERT  P
MOUNTAIN OF STOLEN DREAMS [@PLANET BUILDERS #1]  TALLIS, ROBYN  C/P,P
NIGHT OF GHOSTS AND LIGHTNING [@PLANET BUILDERS #2]  TALLIS, ROBYN  C/P,P
REBEL FROM ALPHORION [@PLANET BUILDERS #3]  TALLIS, ROBYN  P
ZERO-SUM GAMES [@PLANET BUILDERS #5]  TALLIS, ROBYN  P
NIGHT OF TWO NEW MOONS [@PLANET BUILDERS #6]  TALLIS, ROBYN  P
FIRE IN THE SKY [@PLANET BUILDERS #10]  TALLIS, ROBYN  P
PLANET DWELLER, THE [PLANET DWELLER 1]  PALMER, JANE  P
COLONY [@PLANET OF THE APES]  QUICK, W. T.  P
FALL, THE [@PLANET OF THE APES]  QUICK, W. T.  P
MAN THE FUGITIVE [@PLANET OF THE APES #1]  EFFINGER, GEORGE ALEC  C/P:2,P
ESCAPE TO TOMORROW [@PLANET OF THE APES #2]  EFFINGER, GEORGE ALEC  C/P,P
JOURNEY INTO TERROR [@PLANET OF THE APES #3]  EFFINGER, GEORGE ALEC  C/P,P
DEATH OF SLEEP, THE [PLANET PIRATES 2]  MCCAFFREY, ANNE/NYE, JODY LYNN  C/P,P,H
GENERATION WARRIORS [PLANET PIRATES 3]  MCCAFFREY, ANNE/MOON, ELIZABETH  C/P:2,SFWA-P,P

PLANET PIRATES, THE [PLANET PIRATES 1,2,3] MCCAFFREY, ANNE/MOON, ELIZABETH/

NYE, JODY LYNN

LP

SASSINAK [PLANET PIRATES 1]  MCCAFFREY, ANNE/MOON, ELIZABETH  C/P:2,SFWA-P,P
PLANETFALL [PLANETFALL 1]  COVER, ARTHUR BYRON  C/P:2,P
STATIONFALL [PLANETFALL 2]  COVER, ARTHUR BYRON  C/P,P
GODPLAYERS [PLAYERS IN THE CONTEST OF WORLDS 1]  BRODERICK, DAMIEN  LP
K-MACHINES [PLAYERS IN THE CONTEST OF WORLDS 2]  BRODERICK, DAMIEN  LP
MOTHER OF PLENTY [PLENTY 3]  GREENLAND, COLIN  P
SEASONS OF PLENTY [PLENTY 2]  GREENLAND, COLIN  C/P,P
TAKE BACK PLENTY [PLENTY 1]  GREENLAND, COLIN  C/P:2,P
ADVERSARY, THE [PLIOCENE EXILE 4]  MAY, JULIAN  H:2,P
GOLDEN TORC, THE [PLIOCENE EXILE 2]  MAY, JULIAN  C/P,P,H
MANY-COLORED LAND, THE [PLIOCENE EXILE 1]  MAY, JULIAN  C/P,P,H

MANY-COLORED LAND, THE [PLIOCENE EXILE 1,2]/ GOLDEN TORC, THE [PLIOCENE EXILE 1,2] MAY, JULIAN

H:2

NONBORN KING, THE [PLIOCENE EXILE 3]  MAY, JULIAN  C/P:2,SFWA,P

The MITSFS Seriesdex  September 24, 2007  Page 125
TITLE  AUTHOR  CODE

NONBORN KING, THE [PLIOCENE EXILE 3,4]/ ADVERSARY, THE [PLIOCENE EXILE 3,4] MAY, JULIAN

H:2

CHANGELING [POL DETSON 1]  ZELAZNY, ROGER  C/P,H,LP
MADWAND [POL DETSON 2]  ZELAZNY, ROGER  C/P,H,LP
WIZARD WORLD [POL DETSON 1,2]  ZELAZNY, ROGER  C/P:2,P
POLAR CITY BLUES [POLAR CITY 1]  KERR, KATHARINE  C/P,P
POLAR CITY NIGHTMARE [POLAR CITY 2]  KERR, KATHARINE/DANIEL, KATE  LP
PRADOR MOON [POLITY]  ASHER, NEAL  LP
RUNCIBLE TALES [POLITY]  ASHER, NEAL  SR-LP
BRASS MAN [POLITY GRIDLINKED 3]  ASHER, NEAL  P
GRIDLINKED [POLITY GRIDLINKED 1]  ASHER, NEAL  C/P,P,H
LINE OF POLITY, THE [POLITY GRIDLINKED 2]  ASHER, NEAL  C/P,P
POLITY AGENT [POLITY GRIDLINKED 4]  ASHER, NEAL  H
SKINNER, THE [POLITY SKINNER 1]  ASHER, NEAL  H:2,LP
VOYAGE OF THE SABLE KEECH, THE [POLITY SKINNER 2]  ASHER, NEAL  H
HIDDEN SAINT, THE [@POLTERGEIST THE LEGACY]  HAUTALA, RICK  P
GUST FRONT [POSLEEN 2]  RINGO, JOHN  P
HELL’S FAIRE (HARDCOVER AND CD) [POSLEEN 4]  RINGO, JOHN  H:2
HELL’S FAIRE (PROOF) [POSLEEN 4]  RINGO, JOHN  H,SR-LP
HYMN BEFORE BATTLE, A [POSLEEN 1]  RINGO, JOHN  C/P,P
POSLEEN DISK, THE (CD) [POSLEEN]  BAEN PUBLISHING ENTERPRISES  SR-MM
WATCH ON THE RHINE [POSLEEN]  RINGO, JOHN/KRATMAN, TOM  H
WHEN THE DEVIL DANCES [POSLEEN 3]  RINGO, JOHN  H
YELLOW EYES [POSLEEN]  RINGO, JOHN/KRATMAN, TOM  H
CHILDREN OF ENCHANTMENT [POWER AND THE PATTERN 2]  BUSH, ANNE KELLEHER  P
DAUGHTER OF PROPHECY [POWER AND THE PATTERN 1]  BUSH, ANNE KELLEHER  C/P,P
BLOODLOVER [POWER OF THE BOOD 4]  KILPATRICK, NANCY  LP
BIG GAME [PREDATOR]  SCHOFIELD, SANDY  P
COLD WAR [@PREDATOR 2]  ARCHER, NATHAN  C/P,P
CONCRETE JUNGLE [@PREDATOR 1]  ARCHER, NATHAN  P
PREDATOR [PREDATOR]  MONETTE, PAUL  C/P,P
PREDATOR 2 [PREDATOR]  HAWKE, SIMON  P
PRETERNATURAL [PRETERNATURAL 1]  BONANNO, MARGARET WANDER  C/P,P
PRETERNATURAL 3 [PRETERNATURAL 3]  BONANNO, MARGARET WANDER  H:2
PRETERNATURAL TOO: GYRE [PRETERNATURAL 2]  BONANNO, MARGARET WANDER  H,LP
HOUSE OF CTHULHU, THE [PRIMAL LAND 1]  LUMLEY, BRIAN  SR-VLP
SORCERY IN SHAD [PRIMAL LAND 3]  LUMLEY, BRIAN  H
TARRA KHASH: HROSSAK! [PRIMAL LAND 2]  LUMLEY, BRIAN  H
AVATARS, THE [@PRIMAL RAGE]  VORNHOLT, JOHN  P
TARGET EARTH [@PRIMORTALS]  PERRY, STEVE  P
DARKNESS THAT COMES BEFORE, THE [@PRINCE OF NOTHING #1]  BAKKER, R. SCOTT  H
WARRIOR-PROPHET, THE [@PRINCE OF NOTHING #2]  BAKKER, R. SCOTT  H
THOUSANDFOLD THOUGHT, THE [@PRINCE OF NOTHING #3]  BAKKER, R. SCOTT  H
FORGING THE RUNES [PRINCE OF THE SIDHE 2]  SHERMAN, JOSEPHA  P
SHATTERED OATH, THE [PRINCE OF THE SIDHE 1]  SHERMAN, JOSEPHA  P
CHINDI [PRISCILLA HUTCHINS 3]  MCDEVITT, JACK  H:2
DEEPSIX [PRISCILLA HUTCHINS 2]  MCDEVITT, JACK  H
ENGINES OF GOD, THE [PRISCILLA HUTCHINS 1]  MCDEVITT, JACK  C/P:2,P
OMEGA [PRISCILLA HUTCHINS 4]  MCDEVITT, JACK  H
PRISONER, THE [PRISONER]  DISCH, THOMAS M.  C/P,P
PRISONER #2, THE [PRISONER]  MCDANIEL, DAVID  C/P,P
PRISONER #3, THE [PRISONER]  STINE, HANK  C/P:2,P
PRISONER’S DILEMMA, THE [PRISONER]  BLUM, JONATHAN/BOOTH, RUPERT  LP
NIGHTMARE SESSION [@PRIVATE SCHOOL #1]  CHARLES, STEVEN  P
ACADEMY OF TERROR [@PRIVATE SCHOOL #2]  CHARLES, STEVEN  P
WITCH’S EYE [@PRIVATE SCHOOL #3]  CHARLES, STEVEN  P
SKELETON KEY [@PRIVATE SCHOOL #4]  CHARLES, STEVEN  P
ENEMY WITHIN, THE [@PRIVATE SCHOOL #5]  CHARLES, STEVEN  P
LAST ALIEN, THE [@PRIVATE SCHOOL #6]  CHARLES, STEVEN  P
EMPIRE BUILDERS [PRIVATEERS 2]  BOVA, BEN  P
PRIVATEERS [PRIVATEERS 1]  BOVA, BEN  C/P:2,SFWA-P,P,H
COUNTERPROBE [PROBE 2]  DOUGLAS, CAROLE NELSON  H:2,SFWA-P,P

PROBE [PROBE 1] DOUGLAS, CAROLE NELSON C/P,SFWA-P:2,H,LP,

SFWA

CRY REPUBLIC [PROCURATOR 3]  MITCHELL, KIRK  C/P,SFWA-P,P
NEW BARBARIANS [PROCURATOR 2]  MITCHELL, KIRK  P
PROCURATOR [PROCURATOR 1]  MITCHELL, KIRK  P

LOST WORLD, THE [PROF. CHALLENGER 1]/ POISON BELT, THE [PROF. CHALLENGER 1] DOYLE, ARTHUR CONAN

SR-H

WHEN THE WORLD SCREAMED AND OTHER STORIES [PROF. CHALLENGER 2] DOYLE, ARTHUR CONAN

LP

DOOMSDAY ON AJIAT [PROFESSOR JAMESON 5]  JONES, NEIL R.  C/P:2,P
PLANET OF THE DOUBLE SUN, THE [PROFESSOR JAMESON 1]  JONES, NEIL R.  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 126
TITLE  AUTHOR  CODE
SPACE WAR [PROFESSOR JAMESON 3]  JONES, NEIL R.  C/P:2,P
SUNLESS WORLD, THE [PROFESSOR JAMESON 2]  JONES, NEIL R.  C/P:2,P
TWIN WORLDS [PROFESSOR JAMESON 4]  JONES, NEIL R.  C/P:2,P
CURSE OF THE GIANT HOGWEED, THE [PROFESSOR SHANDY 5]  MACLEOD, CHARLOTTE  P
BLOOD AND IRON [PROMETHEAN AGE 1]  BEAR, ELIZABETH  LP
WHISKEY AND WATER [PROMETHEAN AGE 2]  BEAR, ELIZABETH  LP
FIRST TEST [@PROTECTOR OF THE SMALL 1]  PIERCE, TAMORA  H
LADY KNIGHT [@PROTECTOR OF THE SMALL 4]  PIERCE, TAMORA  H
PAGE [@PROTECTOR OF THE SMALL 2]  PIERCE, TAMORA  H
SQUIRE [@PROTECTOR OF THE SMALL 3]  PIERCE, TAMORA  H:2
PROTEUS COMBINED [PROTEUS]  SHEFFIELD, CHARLES  P
PROTEUS IN THE UNDERWORLD [PROTEUS 3]  SHEFFIELD, CHARLES  C/P,P,H
PROTEUS MANIFEST [PROTEUS 1,2]  SHEFFIELD, CHARLES  H
PROTEUS UNBOUND [PROTEUS 2]  SHEFFIELD, CHARLES  C/P,P
SIGHT OF PROTEUS [PROTEUS 1]  SHEFFIELD, CHARLES  C/P:2,P
PREDATOR AND PREY [@PROWLERS 3]  GOLDEN, CHRISTOPHER  P
WILD THINGS [@PROWLERS 4]  GOLDEN, CHRISTOPHER  P
BLACK CAULDRON, THE [PRYDAIN 2]  ALEXANDER, LLOYD  C/P:2,LP,H
BOOK OF THREE, THE [PRYDAIN 1]  ALEXANDER, LLOYD  C/P:2,LP,H
CASTLE OF LLYR, THE [PRYDAIN 3]  ALEXANDER, LLOYD  C/P:3,LP
FOUNDLING, THE [PRYDAIN 6]  ALEXANDER, LLOYD  H,P
HIGH KING, THE [PRYDAIN 5]  ALEXANDER, LLOYD  C/P:2,LP
PRYDAIN CHRONICLES, THE [PRYDAIN 1,2,3,4,5,6]  ALEXANDER, LLOYD  H
TARAN WANDERER [PRYDAIN 4]  ALEXANDER, LLOYD  C/P:3,LP
MIND FORCE WARRIOR [@PSI-MAN #1]  DAVID, PETER  P
DEATHSCAPE [@PSI-MAN #2]  DAVID, PETER  P
MAIN STREET D.O.A. [@PSI-MAN #3]  DAVID, PETER  P
CHAOS KID, THE [@PSI-MAN #4]  DAVID, PETER  C/P,P
STALKER [@PSI-MAN #5]  DAVID, PETER  P
HAVEN [@PSI-MAN #6]  DAVID, PETER  P
HENDRA’S BOOK [@PSI PATROL 2]  BENOIT, HENDRA  P
MAX’S BOOK [@PSI PATROL 3]  HURLEY, MAXWELL  P
SAL’S BOOK [@PSI PATROL 1]  LIQUORI, SAL  P
PROJECT AURA [PSYCHIC WARRIOR]  DOHERTY, ROBERT  P
PSYCHO [PSYCHO 1]  BLOCH, ROBERT  SFWA-P
PSYCHO HOUSE [PSYCHO 3]  BLOCH, ROBERT  SFWA-P
PSYCHO II [PSYCHO 2]  BLOCH, ROBERT  SFWA-P
PSYCHODROME [PSYCHODROME 1]  HAWKE, SIMON  C/P,SFWA-P,P
SHAPECHANGER SCENARIO, THE [PSYCHODROME 2]  HAWKE, SIMON  C/P,SFWA-P,P
PSYCHAMOK [PSYCHOMECH 3]  LUMLEY, BRIAN  C/P,P
PSYCHOMECH [PSYCHOMECH 1]  LUMLEY, BRIAN  C/P,P
PSYCHOSPHERE [PSYCHOMECH 2]  LUMLEY, BRIAN  C/P,P
PTOLIA [PTOLIA 1]  BAKER, FRED  P
DEADSTOCK [PUNKTOWN]  THOMAS, JEFFREY  P
PUNKTOWN (REVISED) [PUNKTOWN]  THOMAS, JEFFREY  H

PURRFECT CRIME [PURRFECT SERIES] WAUGH, CAROL-LYNN ROSSEL/

GREENBERG, MARTIN/ASIMOV, ISAAC

PA

LAUGHTER OF CARTHAGE, THE [PYAT 2]  MOORCOCK, MICHAEL  H

The MITSFS Seriesdex  September 24, 2007  Page 127
TITLE  AUTHOR  CODE
SEEKING THE MYTHICAL FUTURE [Q 1]  HOYLE, TREVOR  C/P,P
THROUGH THE EYE OF TIME [Q 2]  HOYLE, TREVOR  C/P,P
QUAG KEEP [QUAG KEEP 1]  NORTON, ANDRE  H:2,P
RETURN TO QUAG KEEP [QUAG KEEP 2]  NORTON, ANDRE/RABE, JEAN  H:2
KEEPING IT REAL [@QUANTUM GRAVITY #1]  ROBSON, JUSTINA  LP
ANGELS UNAWARE [QUANTUM LEAP]  STORM, L. ELIZABETH  P
DOUBLE OR NOTHING [@QUANTUM LEAP]  HENDERSON, C. J.  P
FOREKNOWLEDGE [QUANTUM LEAP]  DEFILIPPIS  P
HEAT WAVE [QUANTUM LEAP]  KENT, MELANIE  P
INDEPENDENCE [QUANTUM LEAP]  PEEL, JOHN  P
KNIGHTS OF THE MORNINGSTAR [QUANTUM LEAP]  RAWN, MELANIE  C/P:2,P
MAKING OF QUANTUM LEAP, THE [QUANTUM LEAP]  SCHUSTER, HAL  R/REF-P
MIRROR’S EDGE [@QUANTUM LEAP]  DAVIS, CAROL/REESE, ESTHER D.  P
OBSESSIONS [@QUANTUM LEAP]  DAVIS, CAROL  P
ODYSSEY [@QUANTUM LEAP]  WALTON, BARBARA E.  P
PRELUDE [QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,P
PULITZER [QUANTUM LEAP]  STORM, L. ELIZABETH  C/P,P
QUANTUM LEAP [QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,SFWA-P,P
RANDOM MEASURES [@QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,P
SEARCH AND RESCUE [QUANTUM LEAP]  CRANDALL, MELISSA  C/P,P
TOO CLOSE FOR COMFORT [@QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P:2,P
WALL, THE [@QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,SFWA-P:2,P
FIFTH QUARTER [QUARTER SINGERS 2]  HUFF, TANYA  C/P:2,P
NO QUARTER [QUARTER SINGERS 3]  HUFF, TANYA  C/P:2,P
QUARTERED SEA, THE [QUARTER SINGERS 4]  HUFF, TANYA  C/P,P
SING THE FOUR QUARTERS [QUARTER SINGERS 1]  HUFF, TANYA  C/P:2,P
HEADS [QUEEN OF ANGELS 2]  BEAR, GREG  C/P,P,H
MOVING MARS [QUEEN OF ANGELS 3]  BEAR, GREG  H:2,P
QUEEN OF ANGELS [QUEEN OF ANGELS 1]  BEAR, GREG  C/P:2,P
SLANT [QUEEN OF ANGELS 4]  BEAR, GREG  C/P:2,P
COLLAPSIUM, THE [QUEENDOM OF SOL 1]  MCCARTHY, WIL  H
LOST IN TRANSMISSION [QUEENDOM OF SOL 3]  MCCARTHY, WIL  C/P,P
MONARCHS OF SOL, THE [QUEENDOM OF SOL 1,2]  MCCARTHY, WIL  H
TO CRUSH THE MOON [QUEENDOM OF SOL 4]  MCCARTHY, WIL  P
WELLSTONE, THE [QUEENDOM OF SOL 2]  MCCARTHY, WIL  P
BELDAN’S FIRE [QUEEN’S QUARTER 3]  SNYDER, MIDORI  C/P:2,P
NEW MOON [QUEEN’S QUARTER 1]  SNYDER, MIDORI  C/P:2,P
SADAR’S KEEP [QUEEN’S QUARTER 2]  SNYDER, MIDORI  C/P:2,SFWA-P,P
AGNES DAY [QUEST FOR THE WHITE DUCK 3]  FENN, LIONEL  C/P,SFWA-P,P
BLOOD RIVER DOWN [QUEST FOR THE WHITE DUCK 1]  FENN, LIONEL  P,SFWA-P
WEB OF DEFEAT [QUEST FOR THE WHITE DUCK 2]  FENN, LIONEL  C/P,P
BEYOND THE STARS [QUEST FOR TOMORROW 4]  SHATNER, WILLIAM  H
DELTA SEARCH [@QUEST FOR TOMORROW 1]  SHATNER, WILLIAM  P
IN ALIEN HANDS [@QUEST FOR TOMORROW 2]  SHATNER, WILLIAM  P
SHADOW PLANET [QUEST FOR TOMORROW 5]  SHATNER, WILLIAM  H
STEP INTO CHAOS [@QUEST FOR TOMORROW 3]  SHATNER, WILLIAM  P
IMMORTAL GAME [QUEST TRIAD 3]  NILES, DOUGLAS  P
SUITORS DUEL [QUEST TRIAD 2]  NILES, DOUGLAS  P
CITY OF CRYSTAL SHADOW [QUESTIONER 2]  SCHMIDT, DENNIS  P
DARK PARADISE [QUESTIONER 3]  SCHMIDT, DENNIS  P
LABYRINTH [QUESTIONER 1]  SCHMIDT, DENNIS  P
BRIDGE OF SOULS [QUICKENING 3]  MCINTOSH, FIONA  LP
MYRREN’S GIFT [QUICKENING 1]  MCINTOSH, FIONA  H,LP
BLOOD AND MEMORY [@QUICKENING #2]  MCINTOSH, FIONA  C/P,LP

DEMONS AT RAINBOW BRIDGE, THE [QUINTARA 1] CHALKER, JACK L.

C/P,SFWA-P:2,H:2,P,

SFWA

NINETY TRILLION FAUSTS, THE [QUINTARA 3]  CHALKER, JACK L.  H:2,SFWA-P,P
RUN TO CHAOS KEEP, THE [QUINTARA 2]  CHALKER, JACK L.  SFWA-P:2,H:2,P,SFWA:2

The MITSFS Seriesdex  September 24, 2007  Page 129
TITLE  AUTHOR  CODE
NAVIGATOR’S SINDROME [RABELAIS 1]  CARR, JAYGE  H,SFWA
RABELAISIAN REPRISE [RABELAIS 3]  CARR, JAYGE  H,SFWA
TREASURE IN THE HEART OF THE MAZE, THE [RABELAIS 2]  CARR, JAYGE  H,SFWA
DEAD WITCH WALKING [RACHEL MORGAN 1]  HARRISON, KIM  P
EVERY WHICH WAY BUT DEAD [RACHEL MORGAN 3]  HARRISON, KIM  C/P,P
FISTFUL OF CHARMS, A [RACHEL MORGAN 4]  HARRISON, KIM  C/P,P
FOR A FEW DEMONS MORE [RACHEL MORGAN 5]  HARRISON, KIM  H:2
GOOD, THE BAD, AND THE UNDEAD, THE [RACHEL MORGAN 2]  HARRISON, KIM  P
NEW LIFE FOR OLD [RACK 5]  JAMES, LAURENCE  SR-P
EARTH LIES SLEEPING [@RACK #1]  JAMES, LAURENCE  P
WAR ON ALEPH [@RACK #2]  JAMES, LAURENCE  P
BACKFLASH [@RACK #3]  JAMES, LAURENCE  P
PLANET OF THE BLIND [@RACK #4]  JAMES, LAURENCE  P
RADIANT DAWN [RADIANT DAWN 1]  GOODFELLOW, CODY  LP
RAVENOUS DUSK [RADIANT DAWN 2]  GOODFELLOW, CODY  LP
ARC OF THE DREAM [RADIX 3]  ATTANASIO, A. A.  C/P,P
IN OTHER WORLDS [RADIX 2]  ATTANASIO, A. A.  C/P,P
LAST LEGENDS OF EARTH, THE [RADIX 4]  ATTANASIO, A. A.  C/P:2,SFWA,LP
RADIX [RADIX 1]  ATTANASIO, A. A.  C/P:2,P
MOREVI [RAFE AND ASKANA]  LEE, LISA/MORRIS, TEE  LP
RAHXEPHON, VOL. 1 [RAHXEPHON 1]  IZUBUCHI, YUTAKA/BONES  SCX
RAHXEPHON, VOL. 2 [RAHXEPHON 2]  IZUBUCHI, YUTAKA/BONES  SCX
BRIGHT MESSENGERS [RAMA]  LEE, GENTRY  C/P,P
DOUBLE FULL MOON NIGHT [RAMA]  LEE, GENTRY  H,P
GARDEN OF RAMA, THE [RAMA 3]  CLARKE, ARTHUR C./LEE, GENTRY  H:2,P
RAMA II [RAMA 2]  CLARKE, ARTHUR C./LEE, GENTRY  C/P,P,H
RAMA REVEALED [RAMA 4]  CLARKE, ARTHUR C./LEE, GENTRY  H:2,P
RENDEZVOUS WITH RAMA [RAMA 1]  CLARKE, ARTHUR C.  H:2,P
PRINCE OF AYODHYA (PROOF) [RAMAYANA 1]  BANKER, ASHOK K.  SR-LP
PRINCE OF AYODHYA [@RAMAYANA #1]  BANKER, ASHOK K.  P
ORION ARM [RAMPART WORLDS 2]  MAY, JULIAN  C/P,P
PERSEUS SPUR [RAMPART WORLDS 1]  MAY, JULIAN  C/P,P
SAGITTARIUS WHORL [RAMPART WORLDS 3]  MAY, JULIAN  C/P,P
EXCALIBUR ALTERNATIVE, THE [RANKS OF BRONZE]  WEBER, DAVID  H,P
FOREIGN LEGIONS [RANKS OF BRONZE]  DRAKE, DAVID  HA,PA
RANKS OF BRONZE [RANKS OF BRONZE]  DRAKE, DAVID  C/P:2,P
CLAN GROUND [RATHA 2]  BELL, CLARE  H
PEOPLE OF THE SKY [RATHA 5]  BELL, CLARE  C/P,SFWA-P,P,H
RATHA’S CREATURE [RATHA 1]  BELL, CLARE  C/P,SFWA,P
RATS, BATS, & VATS [RATS & BATS 1]  FREER, DAVE/FLINT, ERIC  H
RATS, THE BATS & THE UGLY, THE [RATS & BATS 2]  FLINT, ERIC/FREER, DAVE  H
TIME OF GHOSTS, A [@RAVEN 2]  KIRK, RICHARD  SR-P
SWORDSMISTRESS OF CHAOS [@RAVEN #1]  KIRK, RICHARD  SR-P
LORDS OF THE SHADOWS [@RAVEN #4]  KIRK, RICHARD  SR-P
TIME OF DYING, A [@RAVEN #5]  KIRK, RICHARD  SR-P,C/P
BARONESS OF BLOOD [RAVENLOFT]  BERGSTROM, ELAINE  P
CARNIVAL OF FEAR [RAVENLOFT]  KING, J. ROBERT  P
DANCE OF THE DEAD [@RAVENLOFT]  GOLDEN, CHRISTIE  P
DEATH OF A DARKLORD [@RAVENLOFT]  HAMILTON, LAURELL K.  C/P,LP
ENEMY WITHIN, THE [@RAVENLOFT]  GOLDEN, CHRISTIE  P
HEART OF MIDNIGHT [@RAVENLOFT]  KING, J. ROBERT  P
I, STRAHD [@RAVENLOFT]  ELROD, P. N.  C/P:2,P
KING OF THE DEAD [@RAVENLOFT]  DEWEESE, GENE  P
KNIGHT OF THE BLACK ROSE [RAVENLOFT]  LOWDER, JAMES  P
LORD OF THE NECROPOLIS [RAVENLOFT]  DEWEESE, GENE  P
MORDENHEIM [RAVENLOFT]  WILLIAMSON, CHET  P
SCHOLAR OF DECAY [@RAVENLOFT]  HUFF, TANYA  P
SHADOWBORN [@RAVENLOFT]  CONNORS, WILLIAM W./BEBRIS, CARRIE A.  P

SPECTRE OF THE BLACK ROSE [RAVENLOFT] LOWDER, JAMES/

WHITNEY-ROBINSON, VORONICA

C/P,P

TALES OF RAVENLOFT [RAVENLOFT]  THOMSEN, BRIAN M.  C/PA:2,PA
TAPESTRY OF DARK SOULS [@RAVENLOFT]  BERGSTROM, ELAINE  P
TO SLEEP WITH EVIL [RAVENLOFT]  CARDARELLE, ANDRIA  P
TOWER OF DOOM [@RAVENLOFT]  ANTHONY, MARK  P
VAMPIRE OF THE MISTS [@RAVENLOFT]  GOLDEN, CHRISTIE  C/P:2,P
DEMON WITCH [@RAVENSCLIFF #2]  HUNTINGTON, GEOFFREY  P
DINOSAUR CONQUEST [RAY BRADBURY’S DINOSAURS 6]  LEIGH, STEPHEN  P
DINOSAUR EMPIRE [RAY BRADBURY’S DINOSAURS 5]  LEIGH, STEPHEN/MILLER, JOHN J.  P
DINOSAUR PLANET [RAY BRADBURY’S DINOSAURS 2]  LEIGH, STEPHEN  P,SFWA-P
DINOSAUR SAMURAI [RAY BRADBURY’S DINOSAURS 3]  LEIGH, STEPHEN/MILLER, JOHN L.  P
DINOSAUR WARRIORS [RAY BRADBURY’S DINOSAURS 4]  LEIGH, STEPHEN  P
DINOSAUR WORLD [RAY BRADBURY’S DINOSAURS 1]  LEIGH, STEPHEN  P
FAR SIDE OF THE STARS, THE [RCN 3]  DRAKE, DAVID  H:2

TITLE  AUTHOR  CODE
FAR SIDE OF THE STARS, THE (PROOF) [RCN 3]  DRAKE, DAVID  SR-LP
LT. LEARY, COMMANDING [RCN 2]  DRAKE, DAVID  C/P,P,H
SOME GOLDEN HARBOR [RCN 5]  DRAKE, DAVID  H
WAY TO GLORY, THE [RCN 4]  DRAKE, DAVID  H
WITH THE LIGHTNINGS [RCN 1]  DRAKE, DAVID  H:2,P
FIRESHIPS [REACHES 3]  DRAKE, DAVID  C/P,P,H
IGNITING THE REACHES [REACHES 1]  DRAKE, DAVID  C/P,P,H
THROUGH THE BREACH [REACHES 2]  DRAKE, DAVID  C/P,P,H
LIONESS, THE [REALMS OF ARKANIA THALIONMEL 1]  KRAMER, INA  P
SACRIFICE, THE [REALMS OF ARKANIA THALIONMEL 2]  KRAMER, INA  P
ENEMY REBORN, AN [REALMS OF CHAOS 2]  STACKPOLE, MICHAEL A.  P
HERO BORN, A [REALMS OF CHAOS 1]  STACKPOLE, MICHAEL A.  P
TOUCHED BY MAGIC [REANDN 1]  DURGIN, DORANNA  P
WOLF JUSTICE [REANDN 2]  DURGIN, DORANNA  P
CHAOS BALANCE, THE [RECLUCE 7]  MODESITT, L. E. JR.  H:2,P
COLORS OF CHAOS [RECLUCE 9]  MODESITT, L. E. JR.  H:2,P
DEATH OF CHAOS, THE [RECLUCE 5]  MODESITT, L. E. JR.  H:2,P
FALL OF ANGELS [RECLUCE 6]  MODESITT, L. E. JR.  H:2,P
MAGIC ENGINEER, THE [RECLUCE 3]  MODESITT, L. E. JR.  P
MAGIC OF RECLUCE, THE [RECLUCE 1]  MODESITT, L. E. JR.  H:2,P
MAGI’I OF CYADOR [RECLUCE 10]  MODESITT, L. E. JR.  H:2,P
ORDER WAR, THE [RECLUCE 4]  MODESITT, L. E. JR.  H,P
ORDERMASTER [RECLUCE 13]  MODESITT, L. E. JR.  H
SCION OF CYADOR [RECLUCE 11]  MODESITT, L. E. JR.  H:2
TOWERS OF THE SUNSET, THE [RECLUCE 2]  MODESITT, L. E. JR.  C/P:2,P
WELLSPRING OF CHAOS [RECLUCE 12]  MODESITT, L. E. JR.  H
WHITE ORDER, THE [RECLUCE 8]  MODESITT, L. E. JR.  H:2,P
CAPACITY [RECURSION 2]  BALLANTYNE, TONY  P
DIVERGENCE [RECURSION 3]  BALLANTYNE, TONY  P
RECUSRION [RECURSION 1]  BALLANTYNE, TONY  P
BETTER THAN LIFE [@RED DWARF 2]  NAYLOR, GRANT  C/P:2,P
INFINITY WELCOMES CAREFUL DRIVERS [@RED DWARF 1]  NAYLOR, GRANT  C/P:2,SFWA-P,P
MAN IN THE RUBBER MASK, THE [RED DWARF]  LLEWELLYN, ROBERT  R/REF-P
RED DWARF [RED DWARF 1,2]  NAYLOR, GRANT  H
RED DWARF LOG NO. 1996 [RED DWARF]  ALEXANDER, PAUL  [C/RAN-H]
KNIGHT’S DAWN [@RED PAVILIONS #1]  HUNTER, KIM  LP
WIZARD’S FUNERAL [@RED PAVILIONS #2]  HUNTER, KIM  P
RING OF IKRIBU, THE [@RED SONJA #1]  SMITH, DAVID C./TIERNEY, RICHARD L.  C/P,P
DEMON NIGHT [@RED SONJA #2]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
WHEN HELL LAUGHS [@RED SONJA #3]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
ENDITHOR’S DAUGHTER [@RED SONJA #4]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
AGAINST THE PRINCE OF HELL [@RED SONJA #5]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
STAR OF DOOM [@RED SONJA #6]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
RED LIGHTNING [RED THUNDER 2]  VARLEY, JOHN  H
RED THUNDER [RED THUNDER 1]  VARLEY, JOHN  H
BELLMAKER, THE [REDWALL]  JACQUES, BRIAN  H,P
HIGH RHULAIN [@REDWALL]  JACQUES, BRIAN  P
LEGEND OF LUKE, THE [REDWALL]  JACQUES, BRIAN  LP
LONG PATROL, THE [REDWALL]  JACQUES, BRIAN  P
LORD BROCKTREE [REDWALL]  JACQUES, BRIAN  H
MARIEL OF REDWALL [REDWALL]  JACQUES, BRIAN  C/P,P,H
MARLFOX [REDWALL]  JACQUES, BRIAN  P
MARTIN THE WARRIOR [REDWALL]  JACQUES, BRIAN  P
MATTIMEO [REDWALL]  JACQUES, BRIAN  P
MOSSFLOWER [REDWALL]  JACQUES, BRIAN  C/P,P
OUTCAST OF REDWALL, THE [REDWALL]  JACQUES, BRIAN  P
PEARLS OF LUTRA [REDWALL]  JACQUES, BRIAN  P
RAKKETY TAM [REDWALL]  JACQUES, BRIAN  LP
REDWALL [REDWALL]  JACQUES, BRIAN  C/P,P
SALAMANDASTRON [REDWALL]  JACQUES, BRIAN  C/P,P
TAGGERUNG [REDWALL]  JACQUES, BRIAN  P
KALIF’S WAR, THE [REGIMENT 3]  DALMAS, JOHN  C/P,SFWA-P,P
REGIMENT, THE [REGIMENT 1]  DALMAS, JOHN  C/P:2,P
REGIMENT’S WAR, THE [REGIMENT 4]  DALMAS, JOHN  C/P,SFWA-P,P
THREE-CORNERED WAR, THE [REGIMENT 5]  DALMAS, JOHN  C/P,P
WHITE REGIMENT, THE [REGIMENT 2]  DALMAS, JOHN  C/P:2,SFWA-P,P
MANIFOLD: ORIGIN [REID MALENFANT 3]  BAXTER, STEPHEN  C/P,P,H
MANIFOLD: SPACE [REID MALENFANT 2]  BAXTER, STEPHEN  C/P,P,H:2
MANIFOLD: TIME [REID MALENFANT 1]  BAXTER, STEPHEN  C/P,P,H
PHASE SPACE [REID MALENFANT]  BAXTER, STEPHEN  H
REALM OF LIGHT [REIGN OF SHADOWS 3]  CHESTER, DEBORAH  P
REIGN OF SHADOWS [REIGN OF SHADOWS 1]  CHESTER, DEBORAH  C/P,P
SHADOW WAR [REIGN OF SHADOWS 2]  CHESTER, DEBORAH  P

The MITSFS Seriesdex  September 24, 2007  Page 131
TITLE  AUTHOR  CODE
PATH OF BLOOD [REISIL 3]  FRANCIS, DIANA PHARAOH  C/P,P
PATH OF FATE [REISIL 1]  FRANCIS, DIANA PHARAOH  C/P,P
PATH OF HONOR [REISIL 2]  FRANCIS, DIANA PHARAOH  C/P,P
CHARLATAN, THE [RELAMS OF ARKANIA]  KIESOW, ULRICH  P
RELIC [RELIC 1]  PRESTON, DOUGLAS/CHILD, LINCOLN  P
RELIQUARY [RELIC 2]  PRESTON, DOUGLAS/CHILD, LINCOLN  H
CLOCKS OF IRAZ, THE [RELUCTANT KING 2]  DE CAMP, L. SPRAGUE  C/P:2,P
GOBLIN TOWER, THE [RELUCTANT KING 1]  DE CAMP, L. SPRAGUE  C/P:2,P
HONORABLE BARBARIAN, THE [RELUCTANT KING]  DE CAMP, L. SPRAGUE  H:2,P
RELUCTANT KING, THE [RELUCTANT KING 1,2,3]  DE CAMP, L. SPRAGUE  H:2
UNBEHEADED KING, THE [RELUCTANT KING 3]  DE CAMP, L. SPRAGUE  C/P,SFWA,P,H
REMEMBER GETTYSBURG! [REMEMBER! 2]  RANDLE, KEVIN/CORNETT, ROBERT  P
REMEMBER THE ALAMO! [REMEMBER! 1]  RANDLE, KEVIN/CORNETT, ROBERT  P
REMEMBER THE LITTLE BIGHORN! [REMEMBER! 3]  RANDLE, KEVIN/CORNETT, ROBERT  P
CHILD OF THUNDER [RENSHAI 3]  REICHERT, MICKEY ZUCKER  C/P,P
LAST OF THE RENSHAI, THE [RENSHAI 1]  REICHERT, MICKEY ZUCKER  C/P:2,P
WESTERN WIZARD, THE [RENSHAI 2]  REICHERT, MICKEY ZUCKER  C/P,P
BEYOND RAGNAROK [RENSHAI CHRONICLES 1]  REICHERT, MICKEY ZUCKER  P,SFWA
CHILDREN OF WRATH, THE [RENSHAI CHRONICLES 3]  REICHERT, MICKEY ZUCKER  P
PRINCE OF DEMONS [RENSHAI CHRONICLES 2]  REICHERT, MICKEY ZUCKER  P
ALL THE RAGE [REPAIRMAN JACK]  WILSON, F. PAUL  H
CONSPIRACIES [REPAIRMAN JACK]  WILSON, F. PAUL  H
CRISSCROSS [REPAIRMAN JACK]  WILSON, F. PAUL  P
ENEMY OF THE STATE, AN [REPAIRMAN JACK]  WILSON, F. PAUL  C/P,SFWA,P,H
HAUNTED AIR, THE [REPAIRMAN JACK]  WILSON, F. PAUL  P
HEALER [REPAIRMAN JACK]  WILSON, F. PAUL  C/P,P,H,SFWA
HOSTS [REPAIRMAN JACK]  WILSON, F. PAUL  P
INFERNAL [REPAIRMAN JACK]  WILSON, F. PAUL  H
LEGACIES [REPAIRMAN JACK]  WILSON, F. PAUL  P
WHEELS WITHIN WHEELS [REPAIRMAN JACK]  WILSON, F. PAUL  C/P,SFWA,P,H
BROKEN GOD, THE [REQUIEM FOR HOMO SAPIENS 2]  ZINDELL, DAVID  C/P:2,P
NEVERNESS [REQUIEM FOR HOMO SAPIENS 1]  ZINDELL, DAVID  C/P:2,P,H
WAR IN HEAVEN [REQUIEM FOR HOMO SAPIENS 4]  ZINDELL, DAVID  P
WILD, THE [REQUIEM FOR HOMO SAPIENS 3]  ZINDELL, DAVID  P
ZERO HOUR [@RESIDENT EVIL #0]  PERRY, S. D.  P
UMBRELLA CONSPIRACY, THE [@RESIDENT EVIL #1]  PERRY, S. D.  P
CALIBAN COVE [@RESIDENT EVIL #2]  PERRY, S. D.  P
CITY OF THE DEAD [@RESIDENT EVIL #3]  PERRY, S. D.  P
UNDERWORLD [@RESIDENT EVIL #4]  PERRY, S. D.  P
NEMESIS [@RESIDENT EVIL #5]  PERRY, S. D.  P
CODE: VERONICA [@RESIDENT EVIL #6]  PERRY, S. D.  P
GALVESTON [RESURRECTION MAN 2]  STEWART, SEAN  LP
NIGHT WATCH, THE [RESURRECTION MAN 3]  STEWART, SEAN  C/P:2,P
RESURRECTION MAN [RESURRECTION MAN 1]  STEWART, SEAN  C/P,LP
ENVOY TO NEW WORLDS [RETIEF]  LAUMER, KEITH  C/P:2,DF223,P
GALACTIC DIPLOMAT [RETIEF]  LAUMER, KEITH  C/P,P
RETIEF! [RETIEF]  LAUMER, KEITH  P
RETIEF: AMBASSADOR TO SPACE [RETIEF]  LAUMER, KEITH  H,P
RETIEF AND THE PANGALACTIC PAGEANT OF PULCHRITUDE [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF AND THE RASCALS [RETIEF]  LAUMER, KEITH  P,SFWA-P
RETIEF AND THE WARLORDS [RETIEF]  LAUMER, KEITH  C/P,P,H
RETIEF AT LARGE [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF: DIPLOMAT AT ARMS [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF: EMISSARY TO THE STARS [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF: EMISSARY TO THE STARS (EXPANDED COLLECTION) [RETIEF]  LAUMER, KEITH  P

RETIEF: EMISSARY TO THE STARS (MORE EXPANDED COLLECTION)

[RETIEF] LAUMER, KEITH

C/P,P

RETIEF: ENVOY TO NEW WORLDS [RETIEF]  LAUMER, KEITH  P
RETIEF IN THE RUINS [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF OF THE CDT [RETIEF]  LAUMER, KEITH  C/P,P,H,SR-H,SFWA
RETIEF TO THE RESCUE [RETIEF]  LAUMER, KEITH  C/P,P,H
RETIEF UNBOUND [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF’S PEACE [RETIEF]  KEITH, WILLIAM H. JR.  P
RETIEF’S RANSOM [RETIEF]  LAUMER, KEITH  C/P,SR-H,P,H
RETIEF’S WAR [RETIEF]  LAUMER, KEITH  C/P:2,P
RETURN OF RETIEF, THE [RETIEF]  LAUMER, KEITH  C/P,P
REWARD FOR RETIEF [RETIEF]  LAUMER, KEITH  C/P:2,SFWA-P,P
BURIED DEEP [RETRIEVAL ARTIST 4]  RUSCH, KRISTINE KATHRYN  C/P,P
CONSEQUENCES [RETRIEVAL ARTIST 3]  RUSCH, KRISTINE KATHRYN  P
DISAPPEARED, THE [RETRIEVAL ARTIST 1]  RUSCH, KRISTINE KATHRYN  P
EXTREMES [RETRIEVAL ARTIST 2]  RUSCH, KRISTINE KATHRYN  P
PALOMA [RETRIEVAL ARTIST 5]  RUSCH, KRISTINE KATHRYN  P
BRING IT ON [RETRIEVERS]  GILMAN, LAURA ANNE  LP

TITLE  AUTHOR  CODE
BURNING BRIDGES [RETRIEVERS]  GILMAN, LAURA ANNE  LP
CURSE THE DARK [RETRIEVERS]  GILMAN, LAURA ANNE  H,LP
STAYING DEAD [RETRIEVERS]  GILMAN, LAURA ANNE  H,LP
VISIONS FROM NOWHERE [@RETURN TO THE PLANET OF THE APES #1]  ARROW, WILLIAM  P

ESCAPE FROM TERROR LAGOON [@RETURN TO THE PLANET OF THE APES #2] ARROW, WILLIAM

P

MAN, THE HUNTED ANIMAL [@RETURN TO THE PLANET OF THE APES #3]  ARROW, WILLIAM  P
ABSOLUTION GAP [REVELATION SPACE 3]  REYNOLDS, ALASTAIR  H
CENTURY RAIN [REVELATION SPACE]  REYNOLDS, ALASTAIR  H:2
CHASM CITY [REVELATION SPACE]  REYNOLDS, ALASTAIR  H
DIAMOND DOGS, TURQUOISE DAYS [REVELATION SPACE]  REYNOLDS, ALASTAIR  H
REDEMPTION ARK [REVELATION SPACE 2]  REYNOLDS, ALASTAIR  H:2
REVELATION SPACE [REVELATION SPACE 1]  REYNOLDS, ALASTAIR  C/P,P,H
FUTURE FORWARD [REWIND TO YESTERDAY 2]  PFEFFER, SUSAN BETH  H
NAVIGATOR OF RHADA, THE [RHADA 3]  GILMAN, ROBERT CHAM  C/P,P,H
REBEL OF RHADA, THE [RHADA 1]  GILMAN, ROBERT CHAM  H:2,P
STARKAHN OF RHADA, THE [RHADA 4]  GILMAN, ROBERT CHAM  C/P,P,H
WARLOCK OF RHADA, THE [RHADA 2]  GILMAN, ROBERT CHAM  C/P,P
TOWER OF RAVENS, THE [@RHIANNON’S RIDE #1]  FORSYTH, KATE  P
SHINING CITY, THE [@RHIANNON’S RIDE #2]  FORSYTH, KATE  P
HEART OF STARS, THE [@RHIANNON’S RIDE #3]  FORSYTH, KATE  P
ROCKET’S SHADOW, THE [@RICK BRANT #1]  BLAINE, JOHN  H:2
LOST CITY, THE [@RICK BRANT #2]  BLAINE, JOHN  H:2
SEA GOLD [@RICK BRANT #3]  BLAINE, JOHN  H
100 FATHOMS UNDER [@RICK BRANT #4]  BLAINE, JOHN  H
WHISPERING BOX MYSTERY, THE [@RICK BRANT #5]  BLAINE, JOHN  H:2
PHANTOM SHARK, THE [@RICK BRANT #6]  BLAINE, JOHN  H:2
SMUGGLER’S REEF [@RICK BRANT #7]  BLAINE, JOHN  H
CAVES OF FEAR, THE [@RICK BRANT #8]  BLAINE, JOHN  H
STAIRWAY TO DANGER [@RICK BRANT #9]  BLAINE, JOHN  H:2
GOLDEN SKULL, THE [@RICK BRANT #10]  BLAINE, JOHN  H:2
WAILING OCTOPUS, THE [@RICK BRANT #11]  BLAINE, JOHN  H
ELECTRONIC MIND READER, THE [@RICK BRANT #12]  BLAINE, JOHN  H:2
SCARLET LAKE MYSTERY, THE [@RICK BRANT #13]  BLAINE, JOHN  H
PIRATES OF SHAN, THE [@RICK BRANT #14]  BLAINE, JOHN  H
BLUE GHOST MYSTERY, THE [@RICK BRANT #15]  BLAINE, JOHN  H
EGYPTIAN CAT MYSTERY, THE [@RICK BRANT #16]  BLAINE, JOHN  H:2
FLAMING MOUNTAIN, THE [@RICK BRANT #17]  BLAINE, JOHN  H
FLYING STINGAREE, THE [@RICK BRANT #18]  BLAINE, JOHN  H
RUBY RAY MYSTERY, THE [@RICK BRANT #19]  BLAINE, JOHN  H
VEILED RAIDERS, THE [@RICK BRANT #20]  BLAINE, JOHN  H
ROCKET JUMPER [@RICK BRANT #21]  BLAINE, JOHN  H
DEADLY DUTCHMAN, THE [@RICK BRANT #22]  BLAINE, JOHN  H
DANGER BELOW! [@RICK BRANT #23]  BLAINE, JOHN  H
MAELSTROM [RIFTERS 2]  WATTS, PETER  H,P
STARFISH [RIFTERS 1]  WATTS, PETER  H,P
BEHEMOTH: B-MAX [RIFTERS 3.0]  WATTS, PETER  H
BEHEMOTH: SEPPUKU [RIFTERS 3.1]  WATTS, PETER  H
DARKNESS AT SETHANON, A [RIFTWAR 4]  FEIST, RAYMOND E.  C/P,P,H,SFWA
KING’S BUCCANEER, THE [RIFTWAR]  FEIST, RAYMOND E.  H:2,P
MAGICIAN [RIFTWAR 1,2]  FEIST, RAYMOND E.  H,SFWA:2
MAGICIAN: APPRENTICE [RIFTWAR 1]  FEIST, RAYMOND E.  C/P:3,P
MAGICIAN: MASTER [RIFTWAR 2]  FEIST, RAYMOND E.  C/P:2,P
PRINCE OF THE BLOOD [RIFTWAR]  FEIST, RAYMOND E.  C/P,P,H,SFWA
RAYMOND E. FEIST’S WORLD OF KRONDOR (CD) [RIFTWAR]  SIERRA STUDIOS  SR-MM
SILVERTHORN [RIFTWAR 3]  FEIST, RAYMOND E.  C/P,P,H,SFWA
KRONDOR: TEAR OF THE GODS [RIFTWAR LEGACY 3]  FEIST, RAYMOND E.  P
KRONDOR: THE ASSASSINS [RIFTWAR LEGACY 2]  FEIST, RAYMOND E.  P
KRONDOR: THE BETRAYAL [RIFTWAR LEGACY 1]  FEIST, RAYMOND E.  H:2,P
CROSSOVER [@RIFTWORLD 1]  MCCAY, BILL  C/P:2,SFWA-P,P
VILLAINS [@RIFTWORLD 2]  MCCAY, BILL  C/P,P
MIDNIGHT FALCON [RIGANTE 2]  GEMMELL, DAVID  P
RAVENHEART [RIGANTE 3]  GEMMELL, DAVID  H,P
STORMRIDER [RIGANTE 4]  GEMMELL, DAVID  P
SWORD IN THE STORM [RIGANTE 1]  GEMMELL, DAVID  P
TALES OF THE RIGANTE [RIGANTE 1,2]  GEMMELL, DAVID  H
DANGEROUS GAMES [RILEY JENSON 4]  ARTHUR, KERI  P
FULL MOON RISING [RILEY JENSON 1]  ARTHUR, KERI  P
KISSING SIN [RILEY JENSON 2]  ARTHUR, KERI  P
TEMPTING EVIL [RILEY JENSON 3]  ARTHUR, KERI  P
BLOOD OF THE LAND, THE [RILSIN 2]  BRENNAN, NOEL-ANNE  P
SWORD OF THE LAND, THE [RILSIN 1]  BRENNAN, NOEL-ANNE  P
LORDS OF THE MIDDLE DARK [@RINGS OF THE MASTER #1]  CHALKER, JACK L.  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 133
TITLE  AUTHOR  CODE
PIRATES OF THE THUNDER [@RINGS OF THE MASTER #2]  CHALKER, JACK L.  C/P:2,P
WARRIORS OF THE STORM [@RINGS OF THE MASTER #3]  CHALKER, JACK L.  C/P:2,P
MASKS OF THE MARTYRS [@RINGS OF THE MASTER #4]  CHALKER, JACK L.  C/P:2,P
BORN IN FIRE [RISING STARS 1]  COVER, ARTHUR BYRON  P
RISING STARS: BORN IN FIRE [RISING STARS]  STRACZYNSKI, J. MICHAEL  CX
RISING STARS: POWER [RISING STARS]  STRACZYNSKI, J. MICHAEL  CX
RISING STARS: FIRE AND ASH [@RISING STARS #3]  STRACZYNSKI, J. MICHAEL  CX
BENEATH THE VAULTED HILLS [RIVER INTO DARKNESS 1]  RUSSELL, SEAN  C/P,P
COMPASS OF THE SOUL, THE [RIVER INTO DARKNESS 2]  RUSSELL, SEAN  P
FOREST OF THE NIGHT [RIVERRUN 2]  SOMTOW, S. P.  P
RIVERRUN [RIVERRUN 1]  SOMTOW, S. P.  P
RIVERRUN TRILOGY [RIVERRUN 1,2,3]  SOMTOW, S. P.  LP
RIVERS OF WAR, THE [RIVERS OF WAR 1]  FLINT, ERIC  H
DARK DESIGN, THE [RIVERWORLD 3]  FARMER, PHILIP JOSE  C/P,P,H
DARK DESIGN, THE (PROOF) [RIVERWORLD 3]  FARMER, PHILIP JOSE  SR-LP
FABULOUS RIVERBOAT, THE [RIVERWORLD 2]  FARMER, PHILIP JOSE  C/P,P,H
GODS OF RIVERWORLD [RIVERWORLD 5]  FARMER, PHILIP JOSE  H:2,P
MAGIC LABYRINTH, THE [RIVERWORLD 4]  FARMER, PHILIP JOSE  C/P,P,H,SFWA
QUEST TO RIVERWORLD [RIVERWORLD]  FARMER, PHILIP JOSE  PA
RIVERWORLD AND OTHER STORIES [RIVERWORLD]  FARMER, PHILIP JOSE  C/P:2,P
TALES OF RIVERWORLD [RIVERWORLD]  FARMER, PHILIP JOSE  C/PA,PA
TO YOUR SCATTERED BODIES GO [RIVERWORLD 1]  FARMER, PHILIP JOSE  C/P,P,H
MASTER OF BORANGA [@RO-LAN #1]  SIROTA, MIKE  P
SHROUDED WALKS OF BORANGA, THE [@RO-LAN #2]  SIROTA, MIKE  P
JOURNEY TO MESHARRA [@RO-LAN #3]  SIROTA, MIKE  P
DEMONS OF ZAMMAR, THE [@RO-LAN #4]  SIROTA, MIKE  P
ODYSSEY [ROBOT CITY 1]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
PRODIGY [ROBOT CITY 4]  COVER, ARTHUR BYRON  C/P:2,SFWA-P,P
REFUGE [ROBOT CITY 5]  CHILSON, ROBERT  C/P:2,SFWA-P,P
SUSPICION [ROBOT CITY #2]  MCQUAY, MIKE  C/P:2,P
CYBORG [@ROBOT CITY #3]  WU, WILLIAM F.  C/P:2,SFWA-P,P
PERIHELION [@ROBOT CITY #6]  WU, WILLIAM F.  C/P:2,SFWA-P,P
ESCAPE FROM ROBOTROPOLIS [ROBOT ODYSSEY 1]  D’IGNAZIO, FRED  H:2
GENESIS [@ROBOTECH #1]  MCKINNEY, JACK  C/P:2,P
ROBOTECH [@ROBOTECH #1,2,3]  MCKINNEY, JACK  P
BATTLE CRY [@ROBOTECH #2]  MCKINNEY, JACK  P
HOMECOMING [@ROBOTECH #3]  MCKINNEY, JACK  P
BATTLEHYMN [@ROBOTECH #4]  MCKINNEY, JACK  P
FORCE OF ARMS [@ROBOTECH #5]  MCKINNEY, JACK  P
DOOMSDAY [@ROBOTECH #6]  MCKINNEY, JACK  P
SOUTHERN CROSS [@ROBOTECH #7]  MCKINNEY, JACK  C/P,P
METAL FIRE [@ROBOTECH #8]  MCKINNEY, JACK  C/P,P
FINAL NIGHTMARE, THE [@ROBOTECH #9]  MCKINNEY, JACK  C/P,P
INVID INVASION [@ROBOTECH #10]  MCKINNEY, JACK  C/P,P
METAMORPHOSIS [@ROBOTECH #11]  MCKINNEY, JACK  C/P,P
SYMPHONY OF LIGHT [@ROBOTECH #12]  MCKINNEY, JACK  P
END OF THE CIRCLE, THE [@ROBOTECH #18]  MCKINNEY, JACK  C/P,P
ZENTRAEDI REBELLION, THE [@ROBOTECH #19]  MCKINNEY, JACK  P
MASTERS’ GAMBIT, THE [@ROBOTECH #20]  MCKINNEY, JACK  C/P,P
BEFORE THE INVID STORM [@ROBOTECH #21]  MCKINNEY, JACK  C/P,P
ALLIANCE [ROBOTS AND ALIENS 4]  OLTION, JERRY  P,SFWA-P:2
CHANGELING [ROBOTS AND ALIENS 1]  LEIGH, STEPHEN  C/P:2,P
HUMANITY [ROBOTS AND ALIENS 6]  OLTION, JERRY  P
INTRUDER [ROBOTS AND ALIENS 3]  THURSTON, ROBERT  C/P,SFWA-P:2,P
MAVERICK [ROBOTS AND ALIENS 5]  BETHKE, BRUCE  P,SFWA-P:2
RENEGADE [ROBOTS AND ALIENS 2]  SCOTTEN, CORDELL  C/P:3,SFWA-P,P
DICTATOR [@ROBOTS IN TIME 4]  WU, WILLIAM F.  C/P,P
EMPEROR [@ROBOTS IN TIME 5]  WU, WILLIAM F.  P
INVADER [@ROBOTS IN TIME 6]  WU, WILLIAM F.  P
MARAUDER [@ROBOTS IN TIME 2]  WU, WILLIAM F.  P,SFWA-P
PREDATOR [@ROBOTS IN TIME 1]  WU, WILLIAM F.  P,SFWA-P
WARRIOR [@ROBOTS IN TIME 3]  WU, WILLIAM F.  P

MAROONED ON EDEN [ROCHEWORLD 3] FORWARD, ROBERT L./

FORWARD, MARGARET DODSON

C/P,SFWA-P,P

OCEAN UNDER THE ICE [ROCHEWORLD 4] FORWARD, ROBERT L./

FORWARD, MARTHA DODSON

C/P,P

RESCUED FROM PARADISE [ROCHEWORLD 5] FORWARD, ROBERT L./

FULLER, JULIE FORWARD

C/P,P

RETURN TO ROCHEWORLD [ROCHEWORLD 2] FORWARD, ROBERT L./

FULLER, JULIE FORWARD

C/P:2,SFWA-P,P

ROCHEWORLD [ROCHEWORLD 1]  FORWARD, ROBERT L.  C/P:2,SFWA-P,P
BOLD AS LOVE [ROCK AND ROLL REICH 1]  JONES, GWYNETH  LP
CASTLES MADE OF SAND [ROCK AND ROLL REICH 2]  JONES, GWYNETH  LP

TITLE  AUTHOR  CODE
ALIENS ATE MY HOMEWORK [ROD ALLBRIGHT 1]  COVILLE, BRUCE  P
I LEFT MY SNEAKERS IN DIMENSION X [ROD ALLBRIGHT 2]  COVILLE, BRUCE  LP
RODERICK [RODERICK THE ROBOT 1]  SLADEK, JOHN  C/P:2,SFWA-P,P
RODERICK AT RANDOM [RODERICK THE ROBOT 2]  SLADEK, JOHN  P
KING, THE [RODRIGO OF CALEDON 2]  FEINTUCH, DAVID  H
STILL, THE [RODRIGO OF CALEDON 1]  FEINTUCH, DAVID  C/P:2,LP
WHO CENSORED ROGER RABBIT? [ROGER RABBIT 1]  WOLF, GARY K.  C/P:2,P
WHO P-P-P-PLUGGED ROGER RABBIT? [ROGER RABBIT 2]  WOLF, GARY K.  H
CHAOS AND AMBER [ROGER ZELAZNY’S DAWN OF AMBER 2]  BETANCOURT, JOHN GREGORY  H
DAWN OF AMBER, THE [ROGER ZELAZNY’S DAWN OF AMBER 1]  BETANCOURT, JOHN GREGORY  H:2
SHADOWS OF AMBER [@ROGER ZELAZNY’S DAWN OF AMBER 4]  BETANCOURT, JOHN GREGORY  H
TO RULE IN AMBER [ROGER ZELAZNY’S DAWN OF AMBER 3]  BETANCOURT, JOHN GREGORY  H
BLOOD RELATIVE [@ROGUE TROOPER]  SWALLOW, JAMES  P
CRUCIBLE [@ROGUE TROOPER]  RENNIE, GORDON  P
WIZARD AND A WARLORD, A [ROGUE WIZARD 7]  STASHEFF, CHRISTOPHER  H
WIZARD IN A FEUD, A [ROGUE WIZARD 9]  STASHEFF, CHRISTOPHER  H,P
WIZARD IN BEDLAM, A [ROGUE WIZARD 2]  STASHEFF, CHRISTOPHER  H:2,SFWA,P
WIZARD IN CHAOS, A [ROGUE WIZARD 5]  STASHEFF, CHRISTOPHER  H
WIZARD IN MIDGARD, A [ROGUE WIZARD 6]  STASHEFF, CHRISTOPHER  H
WIZARD IN MIND, A [ROGUE WIZARD 1]  STASHEFF, CHRISTOPHER  P
WIZARD IN PEACE, A [ROGUE WIZARD 4]  STASHEFF, CHRISTOPHER  P
WIZARD IN THE WAY, A [ROGUE WIZARD 8]  STASHEFF, CHRISTOPHER  H
WIZARD IN WAR, A [ROGUE WIZARD 3]  STASHEFF, CHRISTOPHER  H,P
CHARLEMAGNE’S CHAMPION [ROLAND 1]  VAN ASTEN, GAIL  C/P,P
DARK SWORD’S LOVER, THE [ROLAND 2]  VAN ASTEN, GAIL  C/P,P
WILL OF THE WANDERER, THE [@ROSE OF THE PROPHET #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
PALADIN OF THE NIGHT, THE [@ROSE OF THE PROPHET #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
PROPHET OF AKHRAN, THE [@ROSE OF THE PROPHET #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
LONG SHOT FOR ROSINANTE [ROSINANTE 2]  GILLILAND, ALEXIS A.  C/P:2,P
PIRATES OF ROSINANTE, THE [ROSINANTE 3]  GILLILAND, ALEXIS A.  C/P:2,P
REVOLUTION FROM ROSINANTE, THE [ROSINANTE 1]  GILLILAND, ALEXIS A.  P
PRINCESS OF ROUMANIA, A [ROUMANIA 1]  PARK, PAUL  P
TOURMALINE, THE [ROUMANIA 2]  PARK, PAUL  H
DESTRUCTION OF THE BOOKS, THE [ROVER 2]  ODOM, MEL  H
LORD OF THE LIBRARIES [ROVER 3]  ODOM, MEL  P
QUEST FOR THE TRILOGY, THE [ROVER 4]  ODOM, MEL  H
ROVER, THE [ROVER 1]  ODOM, MEL  P
DAMIA [ROWAN 2]  MCCAFFREY, ANNE  C/P,SFWA-P:2,P,H
DAMIA’S CHILDREN [ROWAN 3]  MCCAFFREY, ANNE  C/P,SFWA-P,P,H
ROWAN, THE [ROWAN 1]  MCCAFFREY, ANNE  C/P,SFWA-P,H:2,P,SFWA
TOWER AND THE HIVE, THE [ROWAN 4]  MCCAFFREY, ANNE  H,P
ELF QUEEN & THE KING [RUIN MIST]  STANEK, ROBERT  LP
ELF QUEEN & THE KING II [RUIN MIST]  STANEK, ROBERT  LP
FIELDS OF HONOR [RUIN MIST 3]  STANEK, ROBERT  LP
KEEPER MARTIN’S TALE [RUIN MIST 1]  STANEK, ROBERT  LP
KINGDOM ALLIANCE [RUIN MIST 2]  STANEK, ROBERT  LP
MARK OF THE DRAGON [RUIN MIST 4]  STANEK, ROBERT  LP
ELF QUEEN’S QUEST [RUIN MIST DARK PATH 1]  STANEK, ROBERT  LP
PRINCESS OF THE CHAMELN, A [RULERS OF HYLOR 1]  WILDER, CHERRY  H,P
SUMMER’S KING, THE [RULERS OF HYLOR 3]  WILDER, CHERRY  H:2,P
YORATH THE WOLF [RULERS OF HYLOR 2]  WILDER, CHERRY  H,SFWA,P
BROKEN BLADE [RUNE BLADE 3]  MARSTON, ANN  P
KINGMAKER’S SWORD [RUNE BLADE 1]  MARSTON, ANN  P
WESTERN KING, THE [RUNE BLADE 2]  MARSTON, ANN  P
HIDDEN STARS, THE [RUNE OF UNMAKING 1]  HOWARD, MADELINE  H,LP
BROTHERHOOD OF THE WOLF [RUNELORDS 2]  FARLAND, DAVID  H:2,P
LAIR OF BONES, THE [RUNELORDS 4]  FARLAND, DAVID  H
RUNELORDS, THE [RUNELORDS 1]  FARLAND, DAVID  C/P:2,P
SONS OF THE OAK [RUNELORDS 5]  FARLAND, DAVID  H
WIZARDBORN [RUNELORDS 3]  FARLAND, DAVID  P
LOST RUNES, THE [RUNESPELL 2]  WELCH, JANE  P
RUNES OF SORCERY, THE [RUNESPELL 3]  WELCH, JANE  P
RUNES OF WAR, THE [RUNESPELL 1]  WELCH, JANE  P
HAWKMOON [RUNESTAFF 1,2,3,4]  MOORCOCK, MICHAEL  H
JEWEL IN THE SKULL, THE [RUNESTAFF 1]  MOORCOCK, MICHAEL  C/P:2,P
MAD GOD’S AMULET, THE [RUNESTAFF 2]  MOORCOCK, MICHAEL  C/P:2,P
RUNESTAFF, THE [RUNESTAFF 4]  MOORCOCK, MICHAEL  C/P:2,P
SECRET OF THE RUNESTAFF, THE [RUNESTAFF 4]  MOORCOCK, MICHAEL  P
SORCERER’S AMULET [RUNESTAFF 2]  MOORCOCK, MICHAEL  P
SWORD OF THE DAWN [RUNESTAFF 3]  MOORCOCK, MICHAEL  C/P:2,P
DARK DIVIDE [RUNESWORD 5]  ACRES, MARK  C/P,P
DREAMSTONE, THE [RUNESWORD 3]  RIVKIN, J. F.  C/P,P
HORRIBLE HUMES [RUNESWORD 4]  BILLIAS, STEPHEN  C/P,P

TITLE  AUTHOR  CODE
OUTCASTS [RUNESWORD 1]  EMERY, CLAYTON  C/P,P
SKRYLING’S BLADE [RUNESWORD 2]  ESTES, ROSE/WHAM, TOM  C/P,SFWA-P:2,P
STONE OF TIME, THE [RUNESWORD 6]  ESTES, ROSE/WHAM, TOM  P
NEW AMERICA [RUSTUM]  ANDERSON, POUL  C/P:2,P
ORBIT UNLIMITED [RUSTUM]  ANDERSON, POUL  C/P:2,P
GATE OF IVORY, GATE OF HORN [RYHOPE WOOD 3]  HOLDSTOCK, ROBERT P.  P
HOLLOWING, THE [RYHOPE WOOD 2]  HOLDSTOCK, ROBERT P.  H,P
MYTHAGO WOOD [RYHOPE WOOD 1]  HOLDSTOCK, ROBERT P.  H:2,SFWA,P

The MITSFS Seriesdex  September 24, 2007  Page 137
TITLE  AUTHOR  CODE
EIGHT SPELLS A WEEK [SABRINA THE TEENAGE WITCH]  POCKET BOOKS  PA
SABRINA GOES TO ROME [@SABRINA THE TEENAGE WITCH]  ODOM, MEL  P
SABRINA THE TEENAGE WITCH [@SABRINA THE TEENAGE WITCH #1]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
SHOWDOWN AT THE MALL [@SABRINA THE TEENAGE WITCH #2]  GALLAGHER, DIANA G.  P
GOOD SWITCH, BAD SWITCH [@SABRINA THE TEENAGE WITCH #3]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
HALLOWEEN HAVOC [@SABRINA THE TEENAGE WITCH #4]  GALLAGHER, DIANA G.  P
SANTA’S LITTLE HELPER [@SABRINA THE TEENAGE WITCH #5]  DUBOWSKI, CATHY EAST  P
BEN THERE, DONE THAT [@SABRINA THE TEENAGE WITCH #6]  LOCKE, JOSEPH  P
ALL YOU NEED IS A LOVE SPELL [@SABRINA THE TEENAGE WITCH #7]  REISFELD, RANDI  P
SALEM ON TRIAL [@SABRINA THE TEENAGE WITCH #8]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
DOG’S LIFE, A [@SABRINA THE TEENAGE WITCH #9]  DUBOWSKI, CATHY EAST  P
LOTSA LUCK [@SABRINA THE TEENAGE WITCH #10]  GALLAGHER, DIANA G.  P
PRISONER OF CABIN 13 [@SABRINA THE TEENAGE WITCH #11]  VORNHOLT, JOHN  P
ALL THAT GLITTERS [@SABRINA THE TEENAGE WITCH #12]  GARTON, RAY  P
GO FETCH! [@SABRINA THE TEENAGE WITCH #13]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
SPYING EYES [@SABRINA THE TEENAGE WITCH #14]  HOLDER, NANCY  P
HARVEST MOON [@SABRINA THE TEENAGE WITCH #15]  ODOM, MEL  P

NOW YOU SEE HER, NOW YOU DON’T [@SABRINA THE TEENAGE WITCH #16] GALLAGHER, DIANA G.

P

I’LL ZAP MANHATTAN [@SABRINA THE TEENAGE WITCH #18]  ODOM, MEL  P
SHAMROCK SHENANIGANS [@SABRINA THE TEENAGE WITCH #19]  GALLAGHER, DIANA G.  P
AGE OF AQUARIUMS [@SABRINA THE TEENAGE WITCH #20]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
PROM TIME [@SABRINA THE TEENAGE WITCH #21]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
WITCHOPOLY [@SABRINA THE TEENAGE WITCH #22]  VORNHOLT, JOHN  P
BRIDAL BEDLAM [@SABRINA THE TEENAGE WITCH #23]  GALLAGHER, DIANA G.  P
SCARABIAN NIGHTS [@SABRINA THE TEENAGE WITCH #24]  HOLDER, NANCY  P
WHILE THE CAT’S AWAY [@SABRINA THE TEENAGE WITCH #25]  BATRAE, MARGOT  P
CAT TV [@SABRINA THE TEENAGE WITCH SALEM #1]  DUBOWSKI, MARK  P
TEACHER’S PET [@SABRINA THE TEENAGE WITCH SALEM #2]  BARNES-SVARNEY, PATRICIA  P
YOU’RE HISTORY [@SABRINA THE TEENAGE WITCH SALEM #3]  STRICKLAND, BRAD/STRICKLAND, BARBARA P
KING OF CATS, THE [@SABRINA THE TEENAGE WITCH SALEM #4]  DUBOWSKI, MARK  P
DOG DAY AFTERNOON [@SABRINA THE TEENAGE WITCH SALEM #5]  GALLAGHER, DIANA G.  P
PSYCHIC KITTY [@SABRINA THE TEENAGE WITCH SALEM #6]  DUBOWSKI, CATHY EAST  P
CAT BY THE TAIL [@SABRINA THE TEENAGE WITCH SALEM #7]  VERNEY, SARAH J.  P
SHADOW ON THE STONES [SACRED STONES 3]  CALDECOTT, MOYRA  P
TALL STONES, THE [SACRED STONES 1]  CALDECOTT, MOYRA  P
TEMPLE OF THE SUN, THE [SACRED STONES 2]  CALDECOTT, MOYRA  H,P
CURSE OF SAGAMORE, THE [SAGAMORE 1]  DALKEY, KARA  P
SWORD OF SAGAMORE, THE [SAGAMORE 2]  DALKEY, KARA  C/P,P
BETTER IN THE DARK [SAINT-GERMAIN 8]  YARBRO, CHELSEA QUINN  H,LP
BLOOD GAMES [SAINT-GERMAIN 3]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
BLOOD ROSES [SAINT-GERMAIN 12]  YARBRO, CHELSEA QUINN  LP
COME TWILIGHT [SAINT-GERMAIN 14]  YARBRO, CHELSEA QUINN  H
COMMUNION BLOOD [SAINT-GERMAIN 13]  YARBRO, CHELSEA QUINN  H
DARK OF THE SUN [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H
DARKER JEWELS [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H,LP
FEAST IN EXILE, A [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H
HOTEL TRANSYLVANIA [SAINT-GERMAIN 1]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
MANSIONS OF DARKNESS [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  LP
MIDNIGHT HARVEST [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H
OUT OF THE HOUSE OF LIFE [SAINT-GERMAIN 7]  YARBRO, CHELSEA QUINN  H
PALACE, THE [SAINT-GERMAIN 2]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
PATH OF THE ECLIPSE [SAINT-GERMAIN 4]  YARBRO, CHELSEA QUINN  H,SFWA,SFWA-P,P
SAINT-GERMAIN CHRONICLES, THE [SAINT-GERMAIN 6]  YARBRO, CHELSEA QUINN  H,P
TEMPTING FATE [SAINT-GERMAIN 5]  YARBRO, CHELSEA QUINN  H:2
WRIT IN BLOOD [SAINT-GERMAIN 11]  YARBRO, CHELSEA QUINN  H
CANDLE FOR D’ARTAGNAN, A [SAINT-GERMAIN OLIVIA 3]  YARBRO, CHELSEA QUINN  H:2,LP
CRUSADER’S TORCH [SAINT-GERMAIN OLIVIA 2]  YARBRO, CHELSEA QUINN  C/P,SFWA-P,P,H:2
FLAME IN BYZANTIUM, A [SAINT-GERMAIN OLIVIA 1]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
BEYOND GRAVITY [SALLY FORTH 3]  LEIBER, JUSTIN  C/P:2,SFWA-P,P
BEYOND HUMANITY [SALLY FORTH 2]  LEIBER, JUSTIN  C/P:2,P
BEYOND REJECTION [SALLY FORTH 1]  LEIBER, JUSTIN  C/P,P,H
SAM GUNN FOREVER [SAM GUNN]  BOVA, BEN  P
SAM GUNN OMNIBUS, THE [SAM GUNN]  BOVA, BEN  H
SAM GUNN, UNLIMITED [SAM GUNN]  BOVA, BEN  P
ALIEN BOUNTY [SAM MCCADE 3]  DIETZ, WILLIAM C.  C/P,P
GALACTIC BOUNTY [SAM MCCADE 1]  DIETZ, WILLIAM C.  P
IMPERIAL BOUNTY [SAM MCCADE 2]  DIETZ, WILLIAM C.  P
MCCADE ON THE RUN [SAM MCCADE 3,4]  DIETZ, WILLIAM C.  LP
MCCADE’S BOUNTY [SAM MCCADE 4]  DIETZ, WILLIAM C.  P
SPACE FOR HIRE [SAM SPACE]  NOLAN, WILLIAM F.  C/P:2,P
3 FOR SPACE [SAM SPACE]  NOLAN, WILLIAM F.  LP
SAMANTHA SLADE: MONSTER-SITTER [@SAMANTHA SLADE #1]  SMITH, SUSAN  P

The MITSFS Seriesdex  September 24, 2007  Page 138
TITLE  AUTHOR  CODE
CONFESSIONS OF A TEENAGE FROG [@SAMANTHA SLADE #2]  SMITH, SUSAN  P
OUR FRIEND: PUBLIC NUISANCE #1 [@SAMANTHA SLADE #3]  SMITH, SUSAN  P
TERRORS OF ROCK AND ROLL, THE [@SAMANTHA SLADE #4]  SMITH, SUSAN  P
ALLELUIA FILES, THE [SAMARIA]  SHINN, SHARON  LP
ANGEL-SEEKER [SAMARIA]  SHINN, SHARON  H,P
ANGELICA [SAMARIA]  SHINN, SHARON  H
ARCHANGEL [SAMARIA 1]  SHINN, SHARON  C/P:2,P
JOVAH’S ANGEL [SAMARIA 2]  SHINN, SHARON  H,P
IF I NEVER GET BACK [SAMUEL CLEMENS FOWLER 1]  BROCK, DARRYL  H
TWO IN THE FIELD [SAMUEL CLEMENS FOWLER 2]  BROCK, DARRYL  LP
ADVENTURES OF SAMURAI CAT, THE [SAMURAI CAT 1]  ROGERS, MARK E.  SFWA,LP
MORE ADVENTURES OF SAMURAI CAT [SAMURAI CAT 2]  ROGERS, MARK E.  SFWA
SAMURAI CAT GOES TO HELL [SAMURAI CAT 5]  ROGERS, MARK E.  LP
SAMURAI CAT IN THE REAL WORLD [SAMURAI CAT 3]  ROGERS, MARK E.  VLP
SWORD OF SAMURAI CAT, THE [SAMURAI CAT 4]  ROGERS, MARK E.  LP
SOLAR KILL [@SAND WARS #1]  INGRID, CHARLES  C/P:2,P
LASERTOWN BLUES [@SAND WARS #2]  INGRID, CHARLES  C/P,P
CELESTIAL HIT LIST [@SAND WARS #3]  INGRID, CHARLES  C/P,P
ALIEN SALUTE [@SAND WARS #4]  INGRID, CHARLES  C/P,P
RETURN FIRE [@SAND WARS #5]  INGRID, CHARLES  C/P,P
CHALLENGE MET [@SAND WARS #6]  INGRID, CHARLES  C/P:2,P
ABSOLUTE SANDMAN, VOL. 1 [SANDMAN]  GAIMAN, NEIL  CX
DREAM HUNTERS, THE [@SANDMAN]  GAIMAN, NEIL  CX
SANDMAN BOOK OF DREAMS, THE [SANDMAN]  GAIMAN, NEIL/KRAMER, ED  LPA
PRELUDES & NOCTURNES [@SANDMAN #1]  GAIMAN, NEIL  H,CX
DOLL’S HOUSE, THE [@SANDMAN #2]  GAIMAN, NEIL  H:2,CX
DREAM COUNTRY [@SANDMAN #3]  GAIMAN, NEIL  H,CX
SEASON OF MISTS [@SANDMAN #4]  GAIMAN, NEIL  H,CX
GAME OF YOU, A [@SANDMAN #5]  GAIMAN, NEIL  H,CX
FABLES & REFLECTIONS [@SANDMAN #6]  GAIMAN, NEIL  H,CX
BRIEF LIVES [@SANDMAN #7]  GAIMAN, NEIL  H:2,CX
WORLDS’ END [@SANDMAN #8]  GAIMAN, NEIL  H,CX
KINDLY ONES, THE [@SANDMAN #9]  GAIMAN, NEIL  H,CX
WAKE, THE [@SANDMAN #10]  GAIMAN, NEIL  CX
PROMISE, THE [SANDWRITER 2]  HUGHES, MONICA  H
SWORD OF STRAW, THE [SANGREAL 2]  HEMINGWAY, AMANDA  LP
RETURN OF SANTIAGO, THE [SANTIAGO 2]  RESNICK, MIKE  H
SANTIAGO [SANTIAGO 1]  RESNICK, MIKE  C/P:2,SFWA-P,P
LORD OF EMPERORS [SARANTINE MOSAIC 2]  KAY, GUY GAVRIEL  C/P,P,H
SAILING TO SARANTIUM [SARANTINE MOSAIC 1]  KAY, GUY GAVRIEL  H
SAILING TO SARANTIUM (BRITISH) [SARANTINE MOSAIC 1]  KAY, GUY GAVRIEL  C/P,P,H
WEREWOLF WALKS TONIGHT, THE [@SATAN SLEUTH #2]  AVALLONE, MICHAEL  P
SATELLITE NIGHT FEVER [SATELLITE NIGHT 3]  HOPKINS, JACK  C/P,P
SATELLITE NIGHT NEWS [SATELLITE NIGHT 1]  HOPKINS, JACK  C/P,P
SATELLITE NIGHT SPECIAL [SATELLITE NIGHT 2]  HOPKINS, JACK  C/P,P
CAPTIVES OF THE SAVAGE EMPIRE [SAVAGE EMPIRE 3]  LORRAH, JEAN  C/P,P
DRAGON LORD OF THE SAVAGE EMPIRE [SAVAGE EMPIRE 2]  LORRAH, JEAN  C/P,P
EMPRESS UNBORN [SAVAGE EMPIRE 7]  LORRAH, JEAN  C/P,P
FLIGHT TO THE SAVAGE EMPIRE [SAVAGE EMPIRE 4]  LORRAH, JEAN/HOWLETT, WINSTON A.  C/P,P
SAVAGE EMPIRE [SAVAGE EMPIRE 1]  LORRAH, JEAN  C/P:2,P
SORCERERS OF THE FROZEN ISLES [SAVAGE EMPIRE 5]  LORRAH, JEAN  C/P,P
CHRONICLES OF SCAR, THE [SCAR 1]  SARTI, RON  C/P,P
LEGACY OF THE ANCIENTS [@SCAR #2]  SARTI, RON  P
LANTERNS OF GOD, THE [@SCAR #3]  SARTI, RON  P
SHADOW [@SCAVENGER #1]  PARKER, K. J.  LP
PATTERN [@SCAVENGER #2]  PARKER, K. J.  LP
MEMORY [@SCAVENGER #3]  PARKER, K. J.  LP
BASTARD KING, THE [@SCEPTER OF MERCY #1]  CHERNENKO, DAN  LP
CHERNAGOR PIRATES, THE [@SCEPTER OF MERCY #2]  CHERNENKO, DAN  LP
SCEPTER’S RETURN, THE [@SCEPTER OF MERCY #3]  CHERNENKO, DAN  LP
ILL MET BY MOONLIGHT [SCEPTER’D ISLE 2]  LACKEY, MERCEDES/GELLIS, ROBERTA  H
ILL MET BY MOONLIGHT (PROOF) [SCEPTER’D ISLE 2]  LACKEY, MERCEDES/GELLIS, ROBERTA  SR-LP
THIS SCEPTER’D ISLE [SCEPTER’D ISLE 1]  LACKEY, MERCEDES/GELLIS, ROBERTA  P
SCHRODINGER’S CAT TRILOGY [SCHRODINGER’S CAT 1,2,3]  WILSON, ROBERT ANTON  H,LP
UNIVERSE NEXT DOOR, THE [@SCHRODINGER’S CAT 1]  WILSON, ROBERT ANTON  C/P,P
TRICK TOP HAT, THE [@SCHRODINGER’S CAT #2]  WILSON, ROBERT ANTON  C/P:2,P
HOMING PIGEONS [@SCHRODINGER’S CAT #3]  WILSON, ROBERT ANTON  C/P,P

SCIENCE OF DISCWORLD, THE (REVISED) [SCIENCE OF DISCWORLD 1/ DISCWORLD] PRATCHETT, TERRY/STEWART, IAN/

COHEN, JACK

P

GLOBE, THE [@SCIENCE OF DISCWORLD #2] PRATCHETT, TERRY/STEWART, IAN/

COHEN, JACK

P

DARWIN’S WATCH [@SCIENCE OF DISCWORLD #3/DISCWORLD] PRATCHETT, TERRY/STEWART, IAN/

COHEN, JACK H,P

TITLE  AUTHOR  CODE
DRAGON’S BLOOD [SCORPIO 4]  MCDONOUGH, ALEX  C/P,P
DRAGON’S CLAW [SCORPIO 6]  MCDONOUGH, ALEX  C/P,P
DRAGON’S EYE [SCORPIO 5]  MCDONOUGH, ALEX  P
SCORPIO [SCORPIO 1]  MCDONOUGH, ALEX  C/P:2,P
SCORPIO DESCENDING [SCORPIO 3]  MCDONOUGH, ALEX  C/P,P
SCORPIO RISING [SCORPIO 2]  MCDONOUGH, ALEX  C/P,P
BROTHERS IN ARMS [SCOTT ST. ANDREW 1]  WEAVER, BEN  P
PATRIOTS IN ARMS [SCOTT ST. ANDREW 3]  WEAVER, BEN  P
REBELS IN ARMS [SCOTT ST. ANDREW 2]  WEAVER, BEN  P
FORSAKEN EARTH, THIS [@SEA BEGGARS 2]  KEARNEY, PAUL  LP
MARK OF RAN, THE [@SEA BEGGARS 1]  KEARNEY, PAUL  LP
GODBOND [SEA KING 3]  SPRINGER, NANCY  P,SFWA-P
MADBOND [SEA KING 1]  SPRINGER, NANCY  P,SFWA-P
MINDBOND [SEA KING 2]  SPRINGER, NANCY  P
CHILD OF EARTH [SEA OF GRASS 1]  GERROLD, DAVID  LP
ANCIENT, THE [@SEAQUEST DSV 3]  BISCHOFF, DAVID F.  C/P,P
FIRE BELOW [@SEAQUEST DSV 2]  COSTELLO, MATTHEW J.  P
SEAQUEST DSV: THE NOVEL [@SEAQUEST DSV 1]  DUANE, DIANE/MORWOOD, PETER  P
CITY OF IRON [@SEARCHERS #1]  WILLIAMSON, CHET  P
EMPIRE OF DUST [@SEARCHERS #2]  WILLIAMSON, CHET  C/P,P
SIEGE OF STONE [@SEARCHERS #3]  WILLIAMSON, CHET  C/P:2,P
VOYAGE TO ENEH [SEAS OF KILMOYN 1]  GREEN, ROLAND J.  H,P
COMPASS REACH [SECANTIS 1]  TIEDEMANN, MARK W.  H,LP
METAL OF NIGHT [SECANTIS 2]  TIEDEMANN, MARK W.  LP
OTHER WAYS [SECANTIS]  TIEDEMANN, MARK W.  SR-LP
PEACE & MEMORY [SECANTIS 3]  TIEDEMANN, MARK W.  LP
BUCHANAN CAMPAIGN, THE [SECOND COMMONWEALTH 1]  SHELLEY, RICK  P
FIRES OF COVENTRY, THE [SECOND COMMONWEALTH 2]  SHELLEY, RICK  P
RETURN TO CAMEREIN [SECOND COMMONWEALTH 3]  SHELLEY, RICK  P
BLOOD [SECOND ETHER 1]  MOORCOCK, MICHAEL  H
FABULOUS HARBORS [SECOND ETHER 2]  MOORCOCK, MICHAEL  H,LP
WAR AMONGST THE ANGELS, THE [SECOND ETHER 3]  MOORCOCK, MICHAEL  H,LP
FOUNDATION AND CHAOS [SECOND FOUNDATION 2]  BEAR, GREG  C/P,P,H
FOUNDATION’S FEAR [SECOND FOUNDATION 1]  BENFORD, GREGORY  H:2,P
FOUNDATION’S TRIUMPH [SECOND FOUNDATION 3]  BRIN, DAVID  C/P,P,H
EYE OF THE LABYRINTH [SECOND SONS 2]  FALLON, JENNIFER  P
LION OF SENET, THE [SECOND SONS 1]  FALLON, JENNIFER  P
LORD OF THE SHADOWS [SECOND SONS 3]  FALLON, JENNIFER  P
FACES UNDER WATER [@SECRET BOOKS OF VENUS #1]  LEE, TANITH  H
INITIATION, THE [SECRET CIRCLE 1]  SMITH, L. J.  P
HIDDEN LAND, THE [SECRET COUNTRY 2]  DEAN, PAMELA C.  C/P,P
SECRET COUNTRY, THE [SECRET COUNTRY 1]  DEAN, PAMELA C.  P
WHIM OF THE DRAGON, THE [SECRET COUNTRY 3]  DEAN, PAMELA C.  C/P,P
COURAGE OF FALCONS [SECRET TEXTS 3]  LISLE, HOLLY  H,LP
DIPLOMACY OF WOLVES [SECRET TEXTS 1]  LISLE, HOLLY  C/P:2,LP
VENGEANCE OF DRAGONS [SECRET TEXTS 2]  LISLE, HOLLY  C/P,P
ALIEN EMERGENCIES [SECTOR GENERAL 4,5,6]  WHITE, JAMES  LP
AMBULANCE SHIP [SECTOR GENERAL 4]  WHITE, JAMES  C/P:2,P
CODE BLUE — EMERGENCY [SECTOR GENERAL 7]  WHITE, JAMES  C/P:2,P
DOUBLE CONTACT [SECTOR GENERAL 12]  WHITE, JAMES  C/P,P
FINAL DIAGNOSIS [SECTOR GENERAL 10]  WHITE, JAMES  H
GALACTIC GOURMET, THE [SECTOR GENERAL 9]  WHITE, JAMES  H,P
GENOCIDAL HEALER, THE [SECTOR GENERAL 8]  WHITE, JAMES  H:2,P
HOSPITAL STATION [SECTOR GENERAL 1]  WHITE, JAMES  C/P:2,P
MAJOR OPERATION [SECTOR GENERAL 3]  WHITE, JAMES  C/P:2,P
MIND CHANGER [SECTOR GENERAL 11]  WHITE, JAMES  H
SECTOR GENERAL [SECTOR GENERAL 5]  WHITE, JAMES  C/P:2,P
STAR HEALER [SECTOR GENERAL 6]  WHITE, JAMES  C/P:2,P
STAR SURGEON [SECTOR GENERAL 2]  WHITE, JAMES  C/P:2,P
ALDEBARAN CAMPAIGN, THE [SEEDS OF WAR 2]  RANDLE, KEVIN/CORNETT, ROBERT  C/P,P
AQUARIAN ATTACK, THE [SEEDS OF WAR 3]  RANDLE, KEVIN/CORNETT, ROBERT  C/P:2,P
SEEDS OF WAR [SEEDS OF WAR 1]  RANDLE, KEVIN/CORNETT, ROBERT  P
PRINCESS OF THE STARS [SEIKAI CREST OF THE STARS #1]  MORIOKA, HIROYUKI  LP
MASTER OF CHAOS, THE [SENTIENCE 2]  ADAMS, TERRY A.  C/P:2,P
SENTIENCE [SENTIENCE 1]  ADAMS, TERRY A.  C/P:2,P
DEVIL’S HAND, THE [@SENTINELS #1/ROBOTECH 13]  MCKINNEY, JACK  C/P,P
SENTINELS, THE [@SENTINELS #1,2,3/ROBOTECH 13,14,15]  MCKINNEY, JACK  P
DARK POWERS [@SENTINELS #2/ROBOTECH 14]  MCKINNEY, JACK  P
WORLD KILLERS [@SENTINELS #4/ROBOTECH 16]  MCKINNEY, JACK  C/P,P
RUBICON [@SENTINELS #5/ROBOTECH 17]  MCKINNEY, JACK  C/P,P
MAGYK [SEPTIMUS HEAP 1]  SAGE, ANGIE  H
RAVEN’S SHADOW [SERAPH 1]  BRIGGS, PATRICIA  P
RAVEN’S STRIKE [SERAPH 2]  BRIGGS, PATRICIA  P

The MITSFS Seriesdex  September 24, 2007  Page 140
TITLE  AUTHOR  CODE
SERENITY [SERENITY]  WHEDON, JOSS  R/REF-H

SERENITY 1 [SERENITY 1] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

CX

SERENITY 2 [SERENITY 2] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

CX

SERENITY 3 [SERENITY 3] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

CX

SERENITY: THOSE LEFT BEHIND [SERENITY 1,2,3] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

H

BURNING CROWN, THE [SERPENT’S EGG 2]  MCCURDY, J. FITSGERALD  LP
SERPENT’S EGG, THE [SERPENT’S EGG 1]  MCCURDY, J. FITSGERALD  LP
RAGE OF A DEMON KING [SERPENTWAR 3]  FEIST, RAYMOND E.  C/P,P,H
RISE OF A MERCHANT PRINCE [SERPENTWAR 2]  FEIST, RAYMOND E.  H:2,P
RISE OF A MERCHANT PRINCE (PROOF) [SERPENTWAR 2]  FEIST, RAYMOND E.  SR-LP
SHADOW OF A DARK QUEEN [SERPENTWAR 1]  FEIST, RAYMOND E.  C/P,P,H
SHARDS OF A BROKEN CROWN [SERPENTWAR 4]  FEIST, RAYMOND E.  H:2,P
BORN TO RUN [SERRATED EDGE 1]  LACKEY, MERCEDES/DIXON, LARRY  C/P:2,SFWA-P,P
CHROME BORNE, THE [SERRATED EDGE 1,4]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P
CHROME CIRCLE [SERRATED EDGE 4]  LACKEY, MERCEDES/DIXON, LARRY  C/P:2,P
WHEELS OF FIRE [SERRATED EDGE 2]  LACKEY, MERCEDES/SHEPHERD, MARK  C/P:2,SFWA-P,P
WHEN THE BOUGH BREAKS [SERRATED EDGE 3]  LACKEY, MERCEDES/LISLE, HOLLY  C/P:2,SFWA-P,P
CENTER OF THE CIRCLE, THE [SERVANTS OF ARK 2]  WYLIE, JONATHAN  C/P,P
FIRST NAMED, THE [SERVANTS OF ARK 1]  WYLIE, JONATHAN  C/P,P
MAGE-BORN CHILD, THE [SERVANTS OF ARK 3]  WYLIE, JONATHAN  C/P:2,P
DRIVE TO THE EAST [@SETTLING ACCOUNTS 2]  TURTLEDOVE, HARRY  H
GRAPPLE, THE [@SETTLING ACCOUNTS 3]  TURTLEDOVE, HARRY  H
RETURN ENGAGEMENT [SETTLING ACCOUNTS 1]  TURTLEDOVE, HARRY  H
GATES OF HEAVEN, THE [SEVEN BROTHERS 3]  BENJAMIN, CURT  H
PRINCE OF DREAMS, THE [SEVEN BROTHERS 2]  BENJAMIN, CURT  P
PRINCE OF SHADOW, THE [SEVEN BROTHERS 1]  BENJAMIN, CURT  C/P,P
CIRCLE AT CENTER [SEVEN CIRCLES 1]  NILES, DOUGLAS  LP
WORLD FALL [SEVEN CIRCLES 2]  NILES, DOUGLAS  LP
GODDESS WORLDWEAVER, THE [@SEVEN CIRCLES #3]  NILES, DOUGLAS  LP
CHILDREN OF THE WIND, THE [SEVEN CITADELS 2]  HARRIS, GERALDINE  P
DEAD KINGDOM, THE [SEVEN CITADELS 3]  HARRIS, GERALDINE  P
PRINCE OF THE GODBORN [SEVEN CITADELS 1]  HARRIS, GERALDINE  P
SEVENTH GATE, THE [SEVEN CITADELS 4]  HARRIS, GERALDINE  C/P,P
HIDDEN EMPIRE [SEVEN SUNS 1]  ANDERSON, KEVIN J.  C/P,P
SCATTERED SUNS [SEVEN SUNS 4]  ANDERSON, KEVIN J.  H

VEILED ALLIANCES [@SEVEN SUNS] ANDERSON, KEVIN J./TERANISHI, ROBERT/

BROOME, WENDY

CX

FOREST OF STARS, A [@SEVEN SUNS #2]  ANDERSON, KEVIN J.  H,P
FOREST OF STARS, A (PROOF) [@SEVEN SUNS #2]  ANDERSON, KEVIN J.  SR-LP
HORIZON STORMS [@SEVEN SUNS #3]  ANDERSON, KEVIN J.  H
WIZARD AT HOME, THE [@SEVEN TOWERS]  SHELLEY, RICK  P
WIZARD AT MECQ, THE [@SEVEN TOWERS]  SHELLEY, RICK  P
TRIAL OF THE SEVENTH CARRIER [SEVENTH CARRIER 4]  ALBANO, PETER  P
COMING OF WISDOM, THE [SEVENTH SWORD 2]  DUNCAN, DAVE  C/P:2,P
DESTINY OF THE SWORD, THE [SEVENTH SWORD 3]  DUNCAN, DAVE  C/P:2,P
RELUCTANT SWORDSMAN, THE [SEVENTH SWORD 1]  DUNCAN, DAVE  C/P:2,P
CHILD OF THE PROPHECY [SEVENWATERS 3]  MARILLIER, JULIET  H
DAUGHTER OF THE FOREST [SEVENWATERS 1]  MARILLIER, JULIET  H
SON OF THE SHADOWS [SEVENWATERS 2]  MARILLIER, JULIET  H
RESTORATION [SEYONNE 3]  BERG, CAROL  P
REVELATION [SEYONNE 2]  BERG, CAROL  P
TRANSFORMATION [SEYONNE 1]  BERG, CAROL  P
SHADOWKILL [SHADITH’S QUEST 3/DIADEM 12]  CLAYTON, JO  C/P,P
SHADOWPLAY [SHADITH’S QUEST 1/DIADEM 10]  CLAYTON, JO  C/P:2,P
SHADOWSPEER [SHADITH’S QUEST 2/DIADEM 11]  CLAYTON, JO  C/P:2,P
DAGGER’S EDGE [SHADOW 4]  LOGSTON, ANNE  P
DAGGER’S POINT [SHADOW 5]  LOGSTON, ANNE  P
SHADOW [SHADOW 1]  LOGSTON, ANNE  C/P:2,SFWA-P:2,P
SHADOW DANCE [SHADOW 3]  LOGSTON, ANNE  P
SHADOW HUNT [SHADOW 2]  LOGSTON, ANNE  P
MANKIND WITCH, A [SHADOW OF THE LION 3]  FREER, DAVE  H

SHADOW OF THE LION, THE [SHADOW OF THE LION 1] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

H:2,P

THIS ROUGH MAGIC [SHADOW OF THE LION 2] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

H,P

THIS ROUGH MAGIC (PROOF) [SHADOW OF THE LION 2] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

SR-LP

GATHERING DARK, THE [SHADOW SAGA]  GOLDEN, CHRISTOPHER  C/P,P
SHADOW DAWN [SHADOW WAR 2]  CLAREMONT, CHRIS/LUCAS, GEORGE  H:2

The MITSFS Seriesdex  September 24, 2007  Page 141
TITLE  AUTHOR  CODE
SHADOW MOON [SHADOW WAR 1]  LUCAS, GEORGE/CLAREMONT, CHRIS  C/P,P,H
SHADOW STAR [SHADOW WAR 3]  CLAREMONT, CHRIS/LUCAS, GEORGE  P
WIND AFTER TIME, THE [@SHADOW WARRIOR #1]  BUNCH, CHRIS  C/P,P
HUNT THE HEAVENS [@SHADOW WARRIOR #2]  BUNCH, CHRIS  P
DARKNESS OF GOD, THE [@SHADOW WARRIOR #3]  BUNCH, CHRIS  P
BURNING GODDESS, THE [@SHADOW WORLD 1]  HAMMELL, IAN  P
CITY OF ASSASSINS [@SHADOW WORLD 3]  HAMMELL, IAN  P
CLOCK STRIKES SWORD [@SHADOW WORLD 2]  HAMMELL, IAN  P
STORMRIDERS [@SHADOW WORLD 4]  HAMMELL, IAN  P
SHADOWKINGS [@SHADOWKINGS #1]  COBLEY, MICHAEL  LP
ECHO OF ETERNITY [SHADOWLEAGUE 3]  FUREY, MAGGIE  C/P,P
HEART OF MYRIAL, THE [SHADOWLEAGUE 1]  FUREY, MAGGIE  P
SPIRIT OF THE STONE, THE [SHADOWLEAGUE 2]  FUREY, MAGGIE  P
SHADOWMARCH [SHADOWMARCH 1]  WILLIAMS, TAD  H

SHADOWMARCH, VOLUME 1, SPECIAL SNEAK PREVIEW #2 (EXCERPT, CHAPTERS 3 AND 4) [SHADOWMARCH 1] WILLIAMS, TAD

SR-LP

SHADOWMARCH, VOLUME 1, SPECIAL SNEAK PREVIEW #3 (EXCERPT, CHAPTERS 5 AND 6) [SHADOWMARCH 1] WILLIAMS, TAD

SR-LP

SHADOWPLAY [SHADOWMARCH 2]  WILLIAMS, TAD  H
AFTERSHOCK [@SHADOWRUN]  RABE, JEAN/HELFERS, JOHN  P
BEYOND THE PALE [SHADOWRUN 30/DRAGON HEART 3]  KOKE, JAK  P
BLACK MADONNA [@SHADOWRUN 20]  SARGENT, CARL/GASCOIGNE, MARC  P
BLOOD SPORT [SHADOWRUN 29]  SMEDMAN, LISA  P
BURNING BRIGHT [@SHADOWRUN 15]  DOWD, TOM  P
BURNING TIME, THE [@SHADOWRUN 40]  KENSON, STEPHEN  P

CHOOSE YOUR ENEMIES CAREFULLY [SHADOWRUN 2/ SECRETS OF POWER 2] CHARRETTE, ROBERT N.

C/P:2,P

CROSSROADS [SHADOWRUN 36]  KENSON, STEPHEN  P
DEAD AIR [@SHADOWRUN 22]  KOKE, JAK  P
DROPS OF CORRUPTION [SHADOWRUN]  HARDY, JASON M.  P
FADE TO BLACK [@SHADOWRUN 13]  SMITH, NYX  P
FIND YOUR OWN TRUTH [SHADOWRUN 3/SECRETS OF POWER 3]  CHARRETTE, ROBERT N.  C/P,P
FISTFUL OF DATA, A [@SHADOWRUN]  DEDMAN, STEPHEN  P
FOREVER DRUG, THE [SHADOWRUN 37]  SMEDMAN, LISA  P
HEADHUNTERS [@SHADOWRUN 27]  ODOM, MEL  P
HOUSE OF THE SUN [@SHADOWRUN 17]  FINDLEY, NIGEL D.  P
INTO THE SHADOWS [SHADOWRUN 7]  WEISMAN, JOSEPH K.  C/PA,PA
JUST COMPENSATION [SHADOWRUN 19]  CHARRETTE, ROBERT N.  C/P:2,P
LONE WOLF [@SHADOWRUN 12]  FINDLEY, NIGEL D.  P
LUCIFER DECK, THE [SHADOWRUN 23]  SMEDMAN, LISA  P
NEVER DEAL WITH A DRAGON [SHADOWRUN 1/SECRETS OF POWER 1]  CHARRETTE, ROBERT N.  C/P:2,P
NEVER TRUST AN ELF [SHADOWRUN 6]  CHARRETTE, ROBERT N.  P
NIGHT’S PAWN [@SHADOWRUN 10]  DOWD, TOM  C/P,P
NOSFERATU [SHADOWRUN 14]  SARGENT, CARL/GASCOIGNE, MARC  C/P,P
POISON AGENDAS [SHADOWRUN]  KENSON, STEPHEN  P
PREYING FOR KEEPS [@SHADOWRUN 21]  ODOM, MEL  P
PSYCHOTROPE [@SHADOWRUN 33]  SMEDMAN, LISA  P
RAGNAROCK [@SHADOWRUN 38]  KENSON, STEPHEN  P
RUN HARD, DIE FAST [@SHADOWRUN 35]  ODOM, MEL  P
SHADOWBOXER [SHADOWRUN 25]  POLLOTTA, NICK  P
SHADOWPLAY [@SHADOWRUN 9]  FINDLEY, NIGEL D.  C/P,P
STEEL RAIN [@SHADOWRUN 24]  SMITH, NYX  P
STRANGER SOULS [SHADOWRUN 26/DRAGON HEART 1]  KOKE, JAK  P
STREETS OF BLOOD [@SHADOWRUN 8]  SARGENT, CARL/GASCOIGNE, MARC  C/P,P
STRIPER ASSASSIN [SHADOWRUN 11]  SMITH, NYX  C/P,P
TAILS YOU LOSE [@SHADOWRUN 39]  SMEDMAN, LISA  P
TECHNOBABEL [@SHADOWRUN 31]  KENSON, STEPHEN  P
TERMINUS EXPERIMENT, THE [@SHADOWRUN 34]  BOND, JONATHAN E./KOKE, JAK  P
2XS [SHADOWRUN 4]  FINDLEY, NIGEL D.  C/P:2,P
WHO HUNTS THE HUNTER [@SHADOWRUN 16]  SMITH, NYX  C/P,P
WOLF AND RAVEN [@SHADOWRUN 32]  STACKPOLE, MICHAEL A.  P
WORLDS WITHOUT END [@SHADOWRUN 18/IMMORTALS 3]  SPECTOR, CAROLINE  C/P,P
CHANGELING [@SHADOWRUN #5]  KUBASIK, CHRIS  C/P,SFWA-P,P
CLOCKWORK ASYLUM [@SHADOWRUN DRAGON HEART 2]  KOKE, JAK  P
BORN TO RUN [@SHADOWRUN KELLAN COLT 1]  KENSON, STEPHEN  P
FALLEN ANGELS [@SHADOWRUN KELLAN COLT 2]  KENSON, STEPHEN  P
ANGEL SOULS AND DEVIL HEARTS [SHADOWS 2]  GOLDEN, CHRISTOPHER  C/P,P
OF MASQUES AND MARTYRS [SHADOWS 3]  GOLDEN, CHRISTOPHER  P
OF SAINTS AND SHADOWS [SHADOWS 1]  GOLDEN, CHRISTOPHER  C/P,P
BURNING GROUND, THE [SHADOWSONG 2]  CLAYTON, JO  C/P:2,P
FIRE IN THE SKY [@SHADOWSONG #1]  CLAYTON, JO  C/P:2,P
CRYSTAL HEAT [@SHADOWSONG #3]  CLAYTON, JO  C/P,P
EARTH OF FIRE, SKY OF ICE [SHALRODAN SAGA 1]  HOTCHKISS, ROBERT R.  LP

TITLE  AUTHOR  CODE
SHAMER’S SIGNET, THE [SHAMER 2]  KAABERSOL, LENE  H
EMPRESS OF LIGHT [SHANJI INTERSTELLAR EMPIRE 2]  GLASS, JAMES C.  C/P,P
SHANJI [SHANJI INTERSTELLAR EMPIRE 1]  GLASS, JAMES C.  P
ELFSTONES OF SHANNARA, THE [SHANNARA 2]  BROOKS, TERRY  C/P,LP,H
FIRST KING OF SHANNARA [SHANNARA 0]  BROOKS, TERRY  H:2,P
SWORD OF SHANNARA, THE [SHANNARA 1]  BROOKS, TERRY  C/P,LP,H,SFWA:2
WISHSONG OF SHANNARA, THE [SHANNARA 3]  BROOKS, TERRY  H:2,P
DRUID OF SHANNARA, THE [SHANNARA HERITAGE 2]  BROOKS, TERRY  H:2,P
ELF QUEEN OF SHANNARA, THE [SHANNARA HERITAGE 3]  BROOKS, TERRY  H:2,P
SCIONS OF SHANNARA, THE [SHANNARA HERITAGE 1]  BROOKS, TERRY  H:2,P
TALISMANS OF SHANNARA, THE [SHANNARA HERITAGE 4]  BROOKS, TERRY  H:2
GARDEN OF THE SHAPED, THE [SHAPER EXILE 1]  FINCH, SHEILA  C/P:2,P
SHAPER’S LEGACY [SHAPER EXILE 2]  FINCH, SHEILA  C/P:2,P
SHAPING THE DAWN [SHAPER EXILE 3]  FINCH, SHEILA  C/P:2,SFWA-P,P
BEGUILEMENT [@SHARING KNIFE 1]  BUJOLD, LOIS MCMASTER  H
LEGACY [@SHARING KNIFE 2]  BUJOLD, LOIS MCMASTER  H
CATSEYE [SHATTERED LIGHT 3]  FORSTCHEN, WILLIAM R./DEMAREST, JAKI  P
DARK LIGHT [SHATTERED LIGHT 2]  YARBRO, CHELSEA QUINN  P
RUBY FLAMES [SHATTERED LIGHT 1]  KNAAK, RICHARD A.  P
BEYOND THE ’ZONE [SHATTERZONE 1]  STARK, ED  P
RIVER OF GOD, THE [SHATTERZONE 3]  FARSHTEY, GREG  P
SHATTERED AND OTHER STORIES [SHATTERZONE]  FARSHTEY, GREG  LPA
SOLE SURVIVOR [SHATTERZONE 2]  HENSLEY, SHANE LACY  P
CITY WHO FOUGHT, THE [SHELLPEOPLE 4]  MCCAFFREY, ANNE/STIRLING, S. M.  H:2,P
PARTNERSHIP [SHELLPEOPLE 2]  MCCAFFREY, ANNE/BALL, MARGARET  C/P,SFWA-P,P,H
PARTNERSHIP (EXCERPT) [SHELLPEOPLE 2]  MCCAFFREY, ANNE/BALL, MARGARET  SR-LP
SHIP AVENGED, THE [SHELLPEOPLE 7]  STIRLING, S. M.  C/P,P,H
SHIP ERRANT, THE [SHELLPEOPLE 6]  NYE, JODY LYNN  C/P,P
SHIP WHO SANG, THE [SHELLPEOPLE 1]  MCCAFFREY, ANNE  C/P,P,H
SHIP WHO SEARCHED, THE [SHELLPEOPLE 3]  MCCAFFREY, ANNE/LACKEY, MERCEDES  C/P,P,H
SHIP WHO WON, THE [SHELLPEOPLE 5]  MCCAFFREY, ANNE/NYE, JODY LYNN  C/P,P,H
ROBIN AND THE KING [SHERWOOD 2]  GODWIN, PARKE  C/P,P
SHERWOOD [SHERWOOD 1]  GODWIN, PARKE  P,SFWA-P
DARWIN’S CHILDREN [SHEVA 2]  BEAR, GREG  H
DARWIN’S RADIO [SHEVA 1]  BEAR, GREG  H:2,P
HERO STRIKES BACK, THE [SHIELD AND SOURCE 2]  MOORE, MOIRA J.  P
RESENTING THE HERO [SHIELD AND SOURCE 1]  MOORE, MOIRA J.  P
ASCENSION FACTOR, THE [SHIP 4]  HERBERT, FRANK/RANSOM, BILL  C/P,SFWA-P,P,H
DESTINATION: VOID [SHIP 1]  HERBERT, FRANK  C/P:2,P
JESUS INCIDENT, THE [SHIP 2]  HERBERT, FRANK/RANSOM, BILL  C/P,P,H,SFWA
LAZARUS EFFECT, THE [SHIP 3]  HERBERT, FRANK/RANSOM, BILL  C/P,P,H
BETWEEN DARKNESS AND LIGHT [SHOLAN ALLIANCE 7]  NORMAN, LISANNE  P
BETWEEN DARKNESS AND LIGHT (PROOF) [SHOLAN ALLIANCE 7]  NORMAN, LISANNE  SR-LP
DARK NADIR [SHOLAN ALLIANCE 5]  NORMAN, LISANNE  P
FIRE MARGINS [SHOLAN ALLIANCE 3]  NORMAN, LISANNE  P
FORTUNE’S WHEEL [SHOLAN ALLIANCE 2]  NORMAN, LISANNE  P
RAZOR’S EDGE [SHOLAN ALLIANCE 4]  NORMAN, LISANNE  P
STRONGHOLD RISING [SHOLAN ALLIANCE 6]  NORMAN, LISANNE  P
TURNING POINT [SHOLAN ALLIANCE 1]  NORMAN, LISANNE  C/P,P
SANDS OF KALAVEN, THE [@SHUNLAR 2]  HELLER, CAROL  P
STONES OF STIGA, THE [@SHUNLAR 3]  HELLER, CAROL  P
GATES OF VENSUNOR, THE [@SHUNLAR #1]  HELLER, CAROL  P
CHAMPIONS OF THE SIDHE [SIDHE 2]  FLINT, KENNETH C.  C/P:2,P
MASTER OF THE SIDHE [SIDHE 3]  FLINT, KENNETH C.  C/P:2,P
RIDERS OF THE SIDHE, THE [SIDHE 1]  FLINT, KENNETH C.  C/P:2,P
SILVER’S BANE [SIDHE AND SHADOWLANDS 2]  KELLEHER, ANNE  LP
SIGNAL SHATTERED, A [SIGNAL TO NOISE 2]  NYLUND, ERIC S.  H,P
SIGNAL TO NOISE [SIGNAL TO NOISE 1]  NYLUND, ERIC S.  C/P,P
EMPRESS OF EARTH, THE [SILENCE LEIGH 3]  SCOTT, MELISSA  C/P:2,SFWA-P,P
FIVE-TWELFTHS OF HEAVEN [SILENCE LEIGH 1]  SCOTT, MELISSA  C/P:2,P
ROADS OF HEAVEN, THE [SILENCE LEIGH 1,2,3]  SCOTT, MELISSA  H:2
SILENCE IN SOLITUDE [SILENCE LEIGH 2]  SCOTT, MELISSA  C/P:2,P
DREAMER [SILENT EMPIRE 1]  HARPER, STEVEN  P
NIGHTMARE [SILENT EMPIRE 2]  HARPER, STEVEN  P
OFFSPRING [SILENT EMPIRE 4]  HARPER, STEVEN  P
TRICKSTER [SILENT EMPIRE 3]  HARPER, STEVEN  P
CARNELIAN THRONE, THE [SILISTRA 4]  MORRIS, JANET  C/P,P
GOLDEN SWORD, THE [SILISTRA 2]  MORRIS, JANET  C/P,P
HIGH COUCH OF SILISTRA [SILISTRA 1]  MORRIS, JANET  C/P:2,P
WIND FROM THE ABYSS [SILISTRA 3]  MORRIS, JANET  C/P:3,P
RETURNING CREATION [SILISTRA (REVISED) 1]  MORRIS, JANET  C/P,P
MAGIC’S SILKEN SNARE [SILKEN MAGIC 1]  GILLIGAN, ELIZABETH  P
SILKEN SHROUD, THE [SILKEN MAGIC 2]  GILLIGAN, ELIZABETH  P

TITLE  AUTHOR  CODE
PATH OF THE BOLD [SILVER AGE SENTINELS]  LOWDER, JAMES  LPA
PATH OF THE JUST [SILVER AGE SENTINELS]  LOWDER, JAMES  LPA
TREK TO KRAGGEN-COR [@SILVER CALL #1/MITHGAR 10]  MCKIERNAN, DENNIS L.  C/P,SFWA,P,H
BREGA PATH, THE [@SILVER CALL #2/MITHGAR 11]  MCKIERNAN, DENNIS L.  C/P,SFWA,P,H
METALLIC LOVE [SILVER METAL LOVER 2]  LEE, TANITH  P
METALLIC LOVE (PROOF) [SILVER METAL LOVER 2]  LEE, TANITH  SR-P
SILVER METAL LOVER, THE [SILVER METAL LOVER 1]  LEE, TANITH  H:2,P
GOLD FALCON, THE [@SILVER WYRM #1]  KERR, KATHARINE  H
SPIRIT STONE, THE [@SILVER WYRM #2]  KERR, KATHARINE  H
MISTRESS OF AMBIGUITIES [SILVERGLASS 4]  RIVKIN, J. F.  P
SILVERGLASS [SILVERGLASS 1]  RIVKIN, J. F.  C/P:2,P
WEB OF WIND [SILVERGLASS 2]  RIVKIN, J. F.  P
WITCH OF RHOSTSHYL [SILVERGLASS 3]  RIVKIN, J. F.  C/P,P
AMBROV KEON [SIME/GEN 7]  LORRAH, JEAN  C/P:2,P
CHANNEL’S DESTINY [SIME/GEN 5]  LORRAH, JEAN/LICHTENBERG, JACQUELINE C/P:2,SFWA,P
FIRST CHANNEL [SIME/GEN 3]  LORRAH, JEAN/LICHTENBERG, JACQUELINE C/P,P,H,SFWA
HOUSE OF ZEOR [SIME/GEN 1]  LICHTENBERG, JACQUELINE  C/P,P,H,SFWA
MAHOGANY TRINROSE [SIME/GEN 4]  LICHTENBERG, JACQUELINE  C/P:2,SFWA,P
RENSIME [SIME/GEN 6]  LICHTENBERG, JACQUELINE  C/P,SFWA,P,H
UNTO ZEOR, FOREVER [SIME/GEN 2]  LICHTENBERG, JACQUELINE  C/P,P,H,SFWA
ZELEROD’S DOOM [SIME/GEN 8]  LICHTENBERG, JACQUELINE/LORRAH, JEAN C/P:2,P
PORTERO METHOD, THE [@SIMS #2]  WILSON, F. PAUL  SR-H
IRON SUNRISE [SINGULARITY SKY 2]  STROSS, CHARLES  H
SINGULARITY SKY [SINGULARITY SKY 1]  STROSS, CHARLES  C/P,P,H
TIMELIKE DIPLOMACY [SINGULARITY SKY 1,2]  STROSS, CHARLES  H
SIR APROPOS OF NOTHING [SIR APROPOS OF NOTHING 1]  DAVID, PETER  C/P,P
WOAD TO WUIN, THE [SIR APROPOS OF NOTHING 2]  DAVID, PETER  H,P
ANGRY ANGEL, THE [SISTERS OF THE NIGHT]  YARBRO, CHELSEA QUINN  H,LP
CIRCLE OF THE MOON [SISTERS OF THE RAVEN 2]  HAMBLY, BARBARA  LP
SISTERS OF THE RAVEN [SISTERS OF THE RAVEN 1]  HAMBLY, BARBARA  H:2,LP
MACE OF SOULS, THE [SIX KINGDOMS]  FERGUSSON, BRUCE  C/P,P,H
SHADOW OF HIS WINGS, THE [SIX KINGDOMS]  FERGUSSON, BRUCE  H,SFWA,P
CYBORG IV [SIX MILLION DOLLAR MAN 6]  CAIDIN, MARTIN  H,@P
HIGH CRYSTAL [SIX MILLION DOLLAR MAN 3]  CAIDIN, MARTIN  H:2,@P
WINE, WOMEN, AND WAR [@SIX MILLION DOLLAR MAN #1]  JAHN, MICHAEL  C/P,P
SOLID GOLD KIDNAPPING [@SIX MILLION DOLLAR MAN #2]  RICHARDS, EVAN  P
PILOT ERROR [@SIX MILLION DOLLAR MAN #4]  BARBREE, JAY  P
RESCUE OF ATHENA ONE, THE [@SIX MILLION DOLLAR MAN #5]  JAHN, MICHAEL  C/P,P
GRANTVILLE GAZETTE, VOL. 1 [SIXTEEN THIRTY-TWO]  FLINT, ERIC  PA
GRANTVILLE GAZETTE, VOL. 2 [SIXTEEN THIRTY-TWO]  FLINT, ERIC  HA
GRANTVILLE GAZETTE, VOL. 3 [SIXTEEN THIRTY-TWO]  FLINT, ERIC  HA
RING OF FIRE (PROOF) [SIXTEEN THIRTY-TWO]  FLINT, ERIC  SR-LPA
1635: THE CANNON LAW [SIXTEEN THIRTY-TWO]  FLINT, ERIC/DENNIS, ANDREW  H
1634: THE GALILEO AFFAIR [SIXTEEN THIRTY-TWO 3]  FLINT, ERIC/DENNIS, ANDREW  H:2
1633 [SIXTEEN THIRTY-TWO 2]  WEBER, DAVID/FLINT, ERIC  H
1632 [SIXTEEN THIRTY-TWO 1]  FLINT, ERIC  H
1634: THE RAM REBELLION [SIXTEEN THIRTY-TWO]  FLINT, ERIC/DEMARCE, VIRGINIA  H
WITCH, WITCH, BURNING BRIGHT [SIXTH SENSE 1]  BLOCH, JOHN W.  C/P,P
BOOK OF SKAITH, THE [SKAITH 1,2,3]  BRACKETT, LEIGH  H:2
GINGER STAR, THE [SKAITH 1]  BRACKETT, LEIGH  C/P:2,P
HOUNDS OF SKAITH, THE [SKAITH 2]  BRACKETT, LEIGH  C/P:2,P
REAVERS OF SKAITH, THE [SKAITH 3]  BRACKETT, LEIGH  C/P:2,P
SKEEN’S LEAP [SKEEN 1]  CLAYTON, JO  C/P:2,P
SKEEN’S RETURN [SKEEN 2]  CLAYTON, JO  C/P:2,P
SKEEN’S SEARCH [SKEEN 3]  CLAYTON, JO  C/P:2,P
SWORDS OF NIGHT AND DAY, THE [SKILGANNON THE DAMNED]  GEMMELL, DAVID  H
ASCENDANT SUN [SKOLIAN EMPIRE 5]  ASARO, CATHERINE  H,P
CATCH THE LIGHTNING [SKOLIAN EMPIRE 2]  ASARO, CATHERINE  C/P:2,P
LAST HAWK, THE [SKOLIAN EMPIRE 3]  ASARO, CATHERINE  C/P,P,H
MOON’S SHADOW, THE [SKOLIAN EMPIRE 8]  ASARO, CATHERINE  H
PRIMARY INVERSION [SKOLIAN EMPIRE 1]  ASARO, CATHERINE  C/P,P
QUANTUM ROSE, THE [SKOLIAN EMPIRE 6]  ASARO, CATHERINE  H,P
RADIANT SEAS, THE [SKOLIAN EMPIRE 4]  ASARO, CATHERINE  C/P,P,H
SKYFALL [SKOLIAN EMPIRE 9]  ASARO, CATHERINE  P
SPHERICAL HARMONIC [SKOLIAN EMPIRE 7]  ASARO, CATHERINE  H
FINAL KEY, THE [SKOLIAN TRIAD 2]  ASARO, CATHERINE  H
FALL OF THE SKY LORDS, THE [SKY LORDS 3]  BROSNAN, JOHN  P
SKY LORDS, THE [SKY LORDS 1]  BROSNAN, JOHN  SFWA
BLACK LYNX, THE [SKYLA 2]  BOYER, ELIZABETH H.  P
CLAN OF THE WARLORD, THE [SKYLA 1]  BOYER, ELIZABETH H.  P
KEEPER OF CATS [SKYLA 3]  BOYER, ELIZABETH H.  C/P,P
SKYLARK DUQUESNE [SKYLARK 4]  SMITH, E. E.  C/P:2,P
SKYLARK OF SPACE, THE [SKYLARK 1]  SMITH, E. E.  C/P:2,P,H

The MITSFS Seriesdex  September 24, 2007  Page 144
TITLE  AUTHOR  CODE
SKYLARK OF VALERON [SKYLARK 3]  SMITH, E. E.  C/P,P,H:2
SKYLARK THREE [SKYLARK 2]  SMITH, E. E.  C/P:2,P,H
FIRST BATTLE [SKYRIDER 2]  MICHAELS, MELISA C.  P,SFWA-P
FLOATER FACTOR [SKYRIDER 5]  MICHAELS, MELISA C.  P,SFWA-P
LAST WAR [SKYRIDER 3]  MICHAELS, MELISA C.  P,SFWA-P
PIRATE PRINCE [SKYRIDER 4]  MICHAELS, MELISA C.  P,SFWA-P
SKIRMISH [SKYRIDER 1]  MICHAELS, MELISA C.  C/P,SFWA-P:2,P
PARADOX ALLEY [SKYWAY 3]  DE CHANCIE, JOHN  C/P,P
RED LIMIT FREEWAY [SKYWAY 2]  DE CHANCIE, JOHN  C/P:3,P
STARRIGGER [SKYWAY 1]  DE CHANCIE, JOHN  C/P:2,P
SLAVE MASTERS [SLAVE TRADE 2]  WRIGHT, SUSAN  P
SLAVE TRADE [SLAVE TRADE 1]  WRIGHT, SUSAN  P
SLAVES UNCHAINED [SLAVE TRADE 3]  WRIGHT, SUSAN  P
PERSISTENCE OF MEMORY, THE [@SLOW WORLD #1]  RIPLEY, KAREN  P
WARDEN OF HORSES, THE [@SLOW WORLD #2]  RIPLEY, KAREN  P
ALCHEMIST OF TIME, THE [@SLOW WORLD #3]  RIPLEY, KAREN  P
VERY BIG BANG!, A [SLUGGY FREELANCE]  ABRAMS, PETE  CX
YIPPY SKIPPY, THE EVIL! [SLUGGY FREELANCE 5]  ABRAMS, PETE  SCX
CITY [@SMALLVILLE]  GRAYSON, DEVIN  P
CURSE [@SMALLVILLE]  GRANT, ALAN  P
DRAGON [@SMALLVILLE]  GRANT, ALAN  P
HAUNTINGS [@SMALLVILLE]  HOLDER, NANCY  P
SHADOWS [@SMALLVILLE]  GALLAGHER, DIANA G.  P
SHADOWS (PROOF) [@SMALLVILLE]  GALLAGHER, DIANA G.  SR-P,C/P
SILENCE [@SMALLVILLE]  HOLDER, NANCY  P
STRANGE VISITORS [@SMALLVILLE]  STERN, ROGER  P
WHODUNNIT [@SMALLVILLE]  SMITH, DEAN WESLEY  P
BOY WHO WAS THROWN AWAY, THE [SNOW-EYES 2]  SMITH, STEPHANIE A.  C/P,P
SNOW-EYES [SNOW-EYES 1]  SMITH, STEPHANIE A.  C/P,P,H
SNOW QUEEN, THE [SNOW QUEEN 1]  VINGE, JOAN D.  C/P,SFWA-P,H,P,SFWA
SUMMER QUEEN, THE [SNOW QUEEN 3]  VINGE, JOAN D.  H:2,P
TANGLED UP IN BLUE [SNOW QUEEN 4]  VINGE, JOAN D.  H:2
WORLD’S END [SNOW QUEEN 2]  VINGE, JOAN D.  SFWA-P,H:2,P,SFWA
KINGDOM RIVER [SNOWFALL 2]  SMITH, MITCHELL  P
MOONRISE [SNOWFALL 3]  SMITH, MITCHELL  P
SNOWFALL [SNOWFALL 1]  SMITH, MITCHELL  P
FIRST CASUALTY, THE [SOCIETY OF HUMANITY/UNITY PARTY 1]  MOSCOE, MIKE  C/P,P
PRICE OF PEACE, THE [SOCIETY OF HUMANITY/UNITY PARTY 2]  MOSCOE, MIKE  C/P,P
THEY ALSO SERVE [SOCIETY OF HUMANITY/UNITY PARTY 3]  MOSCOE, MIKE  C/P,P
DEPTHS OF TIME, THE [SOLACE 1]  ALLEN, ROGER MACBRIDE  H,LP
OCEAN OF YEARS, THE [SOLACE 2]  ALLEN, ROGER MACBRIDE  P
SHORES OF TOMORROW, THE [SOLACE 3]  ALLEN, ROGER MACBRIDE  P
DERELICT FOR TRADE [SOLAR QUEEN 5]  NORTON, ANDRE/SMITH, SHERWOOD  C/P,P
MIND FOR TRADE, A [SOLAR QUEEN 6]  NORTON, ANDRE/SMITH, SHERWOOD  P
PLAGUE SHIP [SOLAR QUEEN 2]  NORTH, ANDREW  DD345
PLAGUE SHIP [SOLAR QUEEN 2]  NORTON, ANDRE  C/P,P,H
POSTMARKED THE STARS [SOLAR QUEEN 3]  NORTON, ANDRE  C/P,P,H
SARGASSO OF SPACE [SOLAR QUEEN 1]  NORTH, ANDREW  DD249
SARGASSO OF SPACE [SOLAR QUEEN 1]  NORTON, ANDRE  C/P:2,P
SOLDIER OF ARETE [SOLDIER OF THE MIST 2]  WOLFE, GENE  H:2,SFWA-P,P
SOLDIER OF THE MIST [SOLDIER OF THE MIST 1]  WOLFE, GENE  H:2,P
FOREST MAGE [SOLDIER SON 2]  HOBB, ROBIN  H
SHAMAN’S CROSSING [SOLDIER SON 1]  HOBB, ROBIN  H
HAND OF KANE, THE [SOLOMON KANE 2]  HOWARD, ROBERT E.  P
MOON OF SKULLS, THE [SOLOMON KANE 1]  HOWARD, ROBERT E.  C/P,P
SAVAGE TALES OF SOLOMON KANE, THE [SOLOMON KANE]  HOWARD, ROBERT E.  H
SKULLS IN THE STARS [SOLOMON KANE 4]  HOWARD, ROBERT E.  P
SOLOMON KANE [SOLOMON KANE 3]  HOWARD, ROBERT E.  P
ECLIPSE [SONG CALLED YOUTH 1]  SHIRLEY, JOHN  C/P,LP
ECLIPSE CORONA [SONG CALLED YOUTH 3]  SHIRLEY, JOHN  C/P,P
ECLIPSE PENUMBRA [SONG CALLED YOUTH 2]  SHIRLEY, JOHN  P
PARADISE WAR, THE [SONG OF ALBION #1]  LAWHEAD, STEPHEN R.  H,@P
SILVER HAND, THE [@SONG OF ALBION #2]  LAWHEAD, STEPHEN R.  P
ENDLESS KNOT, THE [SONG OF ALBION #3]  LAWHEAD, STEPHEN R.  H,@P
CELESTIAL STEAM LOCOMOTIVE, THE [SONG OF EARTH 1]  CONEY, MICHAEL G.  H:2,SFWA
GODS OF THE GREATAWAY [SONG OF EARTH 2]  CONEY, MICHAEL G.  H:2
GODS OF THE GREATAWAY (PROOF) [SONG OF EARTH 2]  CONEY, MICHAEL G.  SR-LP

ART OF GEORGE R.R. MARTIN’S A SONG OF ICE AND FIRE, THE [SONG OF ICE AND FIRE] WOOD, BRIAN

A

CLASH OF KINGS, A [SONG OF ICE AND FIRE 2]  MARTIN, GEORGE R. R.  C/P,LP,H:2
CLASH OF KINGS, A (BRITISH) [SONG OF ICE AND FIRE 2]  MARTIN, GEORGE R. R.  H
FEAST FOR CROWS, A [SONG OF ICE AND FIRE 4]  MARTIN, GEORGE R. R.  H
GAME OF THRONES, A [SONG OF ICE AND FIRE 1]  MARTIN, GEORGE R. R.  C/P:2,P,H

The MITSFS Seriesdex  September 24, 2007  Page 145
TITLE  AUTHOR  CODE
STORM OF SWORDS, A [SONG OF ICE AND FIRE 3]  MARTIN, GEORGE R. R.  H:2,P
STORM OF SWORDS, A (BRITISH) [SONG OF ICE AND FIRE 3]  MARTIN, GEORGE R. R.  H

ARMS OF THE KRAKEN, THE (NOVELLA FROM DRAGON#305)

[SONG OF ICE AND FIRE 4.01] MARTIN, GEORGE R. R.

H

REVENGE OF THE VALKYRIE [SONG OF THE DWARVES 2]  GUNNARSSON, THORARINN  C/P,P
SONG OF THE DWARVES [SONG OF THE DWARVES 1]  GUNNARSSON, THORARINN  P
ALANNA [SONG OF THE LIONESS 1]  PIERCE, TAMORA  P
IN THE HAND OF THE GODDESS [SONG OF THE LIONESS 2]  PIERCE, TAMORA  P
LIONESS RAMPANT [SONG OF THE LIONESS 4]  PIERCE, TAMORA  H,P
WOMAN WHO RIDES LIKE A MAN, THE [SONG OF THE LIONESS 3]  PIERCE, TAMORA  H,P
PHANTOM BANJO [@SONGKILLER #1]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
PICKING THE BALLAD’S BONES [@SONGKILLER #2]  SCARBOROUGH, ELIZABETH ANN  C/P,P
STRUM AGAIN? [@SONGKILLER #3]  SCARBOROUGH, ELIZABETH ANN  P
REQUIEM OF STARS [SONGS OF THE STELLAR WIND 1]  HICKMAN, TRACY  P
DARKEST HEART [SONJA BLUE]  COLLINS, NANCY A.  H
DOZEN BLACK ROSES, A [SONJA BLUE]  COLLINS, NANCY A.  LP
MIDNIGHT BLUE [SONJA BLUE 1,2,3]  COLLINS, NANCY A.  H
SONJA BLUE COLLECTION, THE [SONJA BLUE 1,2,3]  COLLINS, NANCY A.  LP
SUNGLASSES AFTER DARK [SONJA BLUE 1]  COLLINS, NANCY A.  C/P,P
HEIR TO THE KING [@SOONE 1]  SOMERS, JAMES  LP
SACRED POOL, THE [SORCERER’S TALE 1]  DOUGLAS, L. WARREN  H
VEIL OF YEARS, THE [SORCERER’S TALE 2]  DOUGLAS, L. WARREN  H,P
SORCERY SPELL BOOK, THE [SORCERY!]  JACKSON, STEVE  [C/RAN-P]
SHAMUTANTI HILLS, THE [@SORCERY! #1]  JACKSON, STEVE  [C/RAN-P:2]
KHARE — CITYPORT OF TRAPS [@SORCERY! #2]  JACKSON, STEVE  [C/RAN-P:2]
SEVEN SERPENTS, THE [@SORCERY! #3]  JACKSON, STEVE  [C/RAN-P]
CROWN OF KINGS, THE [@SORCERY! #4]  JACKSON, STEVE  [C/RAN-P]
BLUE MAGIC [SOUL DRINKER 2]  CLAYTON, JO  C/P:2,P
DRINKER OF SOULS [SOUL DRINKER 1]  CLAYTON, JO  C/P:2,P
GATHERING OF STONES, A [SOUL DRINKER 3]  CLAYTON, JO  C/P:2,P
SOUL DRINKER, THE [SOUL DRINKER 1,2,3]  CLAYTON, JO  H:2
SPIRITS OF FLUX AND ANCHOR [@SOUL RIDER #1]  CHALKER, JACK L.  C/P:2,P
EMPIRES OF FLUX AND ANCHOR [@SOUL RIDER #2]  CHALKER, JACK L.  C/P:2,P
MASTERS OF FLUX AND ANCHOR [@SOUL RIDER #3]  CHALKER, JACK L.  C/P:2,SFWA-P,P
BIRTH OF FLUX AND ANCHOR, THE [@SOUL RIDER #4]  CHALKER, JACK L.  C/P:2,SFWA-P,P
CHILDREN OF FLUX AND ANCHOR [@SOUL RIDER #5]  CHALKER, JACK L.  C/P:2,SFWA-P,P
DREAMBUILDER [SOULSMITH 2]  DEITZ, TOM  P,SFWA-P
SOULSMITH [SOULSMITH 1]  DEITZ, TOM  P
WORDWRIGHT [SOULSMITH 3]  DEITZ, TOM  P,SFWA-P
ALL TOGETHER DEAD [SOUTHERN VAMPIRES 7]  HARRIS, CHARLAINE  H:2
CLUB DEAD [SOUTHERN VAMPIRES 3]  HARRIS, CHARLAINE  P
DEAD AS A DOORNAIL [SOUTHERN VAMPIRES 5]  HARRIS, CHARLAINE  H
DEAD BY DAY [SOUTHERN VAMPIRES 4,5]  HARRIS, CHARLAINE  H
DEAD IN DIXIE [SOUTHERN VAMPIRES 1,2,3]  HARRIS, CHARLAINE  H
DEAD TO THE WORLD [SOUTHERN VAMPIRES 4]  HARRIS, CHARLAINE  H
DEAD UNTIL DARK [SOUTHERN VAMPIRES 1]  HARRIS, CHARLAINE  P
DEFINITELY DEAD [SOUTHERN VAMPIRES 6]  HARRIS, CHARLAINE  H:2
LIVING DEAD IN DALLAS [SOUTHERN VAMPIRES 2]  HARRIS, CHARLAINE  C/P:2,P
JOURNEY INTO THE VOID [SOVEREIGN STONE 3]  WEIS, MARGARET/HICKMAN, TRACY  H
JOURNEY INTO THE VOID (PROOF) [SOVEREIGN STONE 3]  WEIS, MARGARET/HICKMAN, TRACY  SR-LP
WELL OF DARKNESS [SOVEREIGN STONE 1]  WEIS, MARGARET/HICKMAN, TRACY  H
GUARDIANS OF THE LOST [@SOVEREIGN STONE #2]  WEIS, MARGARET/HICKMAN, TRACY  H,P
ALIEN SEED [SPACE: 1999 7]  TUBB, E. C.  C/P:2,P
ANDROID PLANET [SPACE: 1999 8]  RANKINE, JOHN  C/P,P
ASTRAL QUEST [SPACE: 1999 6]  RANKINE, JOHN  P
BREAKAWAY [SPACE: 1999 1]  TUBB, E. C.  C/P:2,P
COLLISION COURSE [SPACE: 1999 4]  TUBB, E. C.  C/P,P
ETERNITY UNBOUND [@SPACE: 1999]  LATHAM, WILLIAM  LP
FORSAKEN, THE [@SPACE: 1999]  MUIR, JOHN KENNETH  P
LUNAR ATTACK [SPACE: 1999 5]  RANKINE, JOHN  C/P,P
MOON ODYSSEY [SPACE: 1999 2]  RANKINE, JOHN  C/P:2,P
PHOENIX OF MEGARON [SPACE: 1999 10]  RANKINE, JOHN  P
RESURRECTION [@SPACE: 1999]  LATHAM, WILLIAM  P
ROGUE PLANET [SPACE: 1999 9]  TUBB, E. C.  C/P,P
SPACE GUARDIANS, THE [SPACE: 1999 3]  BALL, BRIAN N.  C/P:2,P
SURVIVAL [@SPACE: 1999]  BALL, BRIAN N.  P
SPACE: 1999 YEAR TWO [SPACE: 1999 YEAR 2]  BUTTERWORTH, MICHAEL  SR-VLH
PLANETS OF PERIL [@SPACE: 1999 YEAR 2 #1]  BUTTERWORTH, MICHAEL  C/P,P
MIND-BREAKS OF SPACE [@SPACE: 1999 YEAR 2 #2]  BUTTERWORTH, MICHAEL  C/P,P
SPACE-JACKERS, THE [@SPACE: 1999 YEAR 2 #3]  BUTTERWORTH, MICHAEL  P
PSYCHOMORPH, THE [@SPACE: 1999 YEAR 2 #4]  BUTTERWORTH, MICHAEL  C/P,P
TIME FIGHTERS, THE [@SPACE: 1999 YEAR 2 #5]  BUTTERWORTH, MICHAEL  P
EDGE OF THE INFINITE, THE [@SPACE: 1999 YEAR 2 #6]  BUTTERWORTH, MICHAEL  P

The MITSFS Seriesdex  September 24, 2007  Page 146
TITLE  AUTHOR  CODE
DEMOLITION WINTER [SPACE: ABOVE AND BEYOND 2]  TELEP, PETER  C/P,P
SPACE: ABOVE AND BEYOND [SPACE: ABOVE AND BEYOND 1]  TELEP, PETER  P
SPACE ANGEL [SPACE ANGEL 1]  ROBERTS, JOHN MADDOX  C/P:2,P
SPACER: WINDOW OF THE MIND [SPACE ANGEL 2]  ROBERTS, JOHN MADDOX  P
HIGH MOON [@SPACE COPS 3]  DUANE, DIANE/MORWOOD, PETER  C/P:2,P
KILL STATION [@SPACE COPS 1]  DUANE, DIANE/MORWOOD, PETER  C/P:2,P
MINDBLAST [@SPACE COPS 2]  DUANE, DIANE/MORWOOD, PETER  C/P,P
FASTER THAN LIGHT [@SPACE HAWKS #1]  PACKARD, EDWARD  P
ALIEN INVADERS [@SPACE HAWKS #2]  PACKARD, EDWARD  P
CHILDREN OF THE NIGHT [SPACE MAVERICKS 2]  KRING, MICHAEL  P
SPACE MAVERICKS, THE [SPACE MAVERICKS 1]  KRING, MICHAEL  P
MERCHANTS’ WAR, THE [SPACE MERCHANTS 2]  POHL, FREDERIK  C/P,P,H
SPACE MERCHANTS, THE [SPACE MERCHANTS 1]  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P

2001: A SPACE ODYSSEY THE FINAL CHAPTERS (LP RECORD)

[SPACE ODYSSEY] CLARKE, ARTHUR C.

SR-MM

2001: A SPACE ODYSSEY [@SPACE ODYSSEY #1]  CLARKE, ARTHUR C.  C/P,SFWA-P,H:2,SR-H,P
2001: A SPACE ODYSSEY (REVISED) [@SPACE ODYSSEY #1]  CLARKE, ARTHUR C.  P
2010: ODYSSEY TWO [@SPACE ODYSSEY #2]  CLARKE, ARTHUR C.  H:2,P
2010: ODYSSEY TWO (PROOF) [@SPACE ODYSSEY #2]  CLARKE, ARTHUR C.  SR-LP
2061: ODYSSEY THREE [@SPACE ODYSSEY #3]  CLARKE, ARTHUR C.  H:2,P
3001: THE FINAL ODYSSEY [@SPACE ODYSSEY #4]  CLARKE, ARTHUR C.  H:2,P
3001: THE FINAL ODYSSEY (PROOF) [@SPACE ODYSSEY #4]  CLARKE, ARTHUR C.  SR-LP
DEITY-FATHER, THE [@SPACE PRECINCT #1]  BISCHOFF, DAVID F.  P
DEMON WING [@SPACE PRECINCT #2]  BISCHOFF, DAVID F.  P
ALIEN ISLAND [@SPACE PRECINCT #3]  BISCHOFF, DAVID F.  C/P,P
OF ALIEN BONDAGE [@SPACEWAYS #1]  CLEVE, JOHN  P
CORUNDUM’S WOMAN [@SPACEWAYS #2]  CLEVE, JOHN  P
ESCAPE FROM MACHO [@SPACEWAYS #3]  CLEVE, JOHN  C/P:2,P
SATANA ENSLAVED [@SPACEWAYS #4]  CLEVE, JOHN  P
MASTER OF MISFIT [@SPACEWAYS #5]  CLEVE, JOHN  P
PURRFECT PLUNDER [@SPACEWAYS #6]  CLEVE, JOHN  P
MANHUNTRESS, THE [@SPACEWAYS #7]  CLEVE, JOHN  P
UNDER TWIN SUNS [@SPACEWAYS #8]  CLEVE, JOHN  P
IN QUEST OF QALARA [@SPACEWAYS #9]  CLEVE, JOHN  P
YOKE OF SHEN, THE [@SPACEWAYS #10]  CLEVE, JOHN  P
ICEWORLD CONNECTION, THE [@SPACEWAYS #11]  CLEVE, JOHN  P
STAR SLAVER [@SPACEWAYS #12]  CLEVE, JOHN  P
JONUTA RISING! [@SPACEWAYS #13]  CLEVE, JOHN  P
ASSIGNMENT: HELLHOLE [@SPACEWAYS #14]  CLEVE, JOHN  P
STARSHIP SAPPHIRE [@SPACEWAYS #15]  CLEVE, JOHN  P
PLANET MURDERER, THE [@SPACEWAYS #16]  CLEVE, JOHN  P
CARNADYNE HORDE, THE [@SPACEWAYS #17]  CLEVE, JOHN  P
RACE ACROSS THE STARS [@SPACEWAYS #18]  CLEVE, JOHN  P
KING OF THE SLAVERS [@SPACEWAYS #19]  CLEVE, JOHN  C/P,P
DRAGON’S DAGGER, THE [SPEARWIELDER 2]  SALVATORE, R. A.  C/P,P
DRAGONSLAYER’S RETURN [SPEARWIELDER 3]  SALVATORE, R. A.  P
WOODS OUT BACK, THE [SPEARWIELDER 1]  SALVATORE, R. A.  C/P:2,SFWA-P,P
DEEP STRIKE [SPEC OPS SQUAD 2]  SHELLEY, RICK  C/P,P
HOLDING THE LINE [SPEC OPS SQUAD 1]  SHELLEY, RICK  P
SUCKER PUNCH [@SPEC OPS SQUAD 3]  SHELLEY, RICK  P
SURVIVAL [@SPECIES IMPERATIVE #1]  CZERNEDA, JULIE E.  H
MIGRATION [@SPECIES IMPERATIVE #2]  CZERNEDA, JULIE E.  H
REGENERATION [@SPECIES IMPERATIVE #3]  CZERNEDA, JULIE E.  H
SILVERADO [@SPECTROS #1]  WINTERS, LOGAN  P
HUNT THE BEAST DOWN [@SPECTROS #2]  WINTERS, LOGAN  P
NATCHEZ [@SPECTROS #3]  WINTERS, LOGAN  P
SILVER CANYON, THE [@SPECTROS #4]  WINTERS, LOGAN  P
IN THE SEA NYMPH’S LAIR [SPELL WARS 2]  HALE, F. J.  P
OGRE CASTLE [SPELL WARS 1]  HALE, F. J.  P
RADIANT DRAGON, THE [SPELLJAMMER CLOAKMASTER 4]  CUNNINGHAM, ELAINE  C/P,P
BEYOND THE MOONS [@SPELLJAMMER CLOAKMASTER #1]  COOK, DAVID  C/P:2,P
INTO THE VOID [@SPELLJAMMER CLOAKMASTER #2]  FINDLEY, NIGEL D.  C/P,P
MAELSTROM’S EYE, THE [@SPELLJAMMER CLOAKMASTER #3]  MOORE, ROGER E.  P
BROKEN SPHERE, THE [@SPELLJAMMER CLOAKMASTER #5]  FINDLEY, NIGEL D.  P
ULTIMATE HELM, THE [@SPELLJAMMER CLOAKMASTER #6]  HOWARD, RUSS T.  C/P,P
BOOKS OF THE KEEPERS, THE [SPELLKEY 3]  DOWNER, ANN  H
SPELLKEY TRILOGY, THE [SPELLKEY (REVISED) 1,2,3]  DOWNER, ANN  P
CHORUS SKATING [SPELLSINGER 8]  FOSTER, ALAN DEAN  C/P,P
DAY OF THE DISSONANCE, THE [SPELLSINGER 3]  FOSTER, ALAN DEAN  C/P:2,P
HOUR OF THE GATE, THE [SPELLSINGER 2]  FOSTER, ALAN DEAN  C/P:2,P
MOMENT OF THE MAGICIAN, THE [SPELLSINGER 4]  FOSTER, ALAN DEAN  C/P,P,H
PATHS OF THE PERAMBULATOR, THE [SPELLSINGER 5]  FOSTER, ALAN DEAN  C/P,P,H
SEASON OF THE SPELLSONG [SPELLSINGER 1,2,3]  FOSTER, ALAN DEAN  H:2

The MITSFS Seriesdex  September 24, 2007  Page 147
TITLE  AUTHOR  CODE
SON OF SPELLSINGER [SPELLSINGER 7]  FOSTER, ALAN DEAN  C/P:2,SFWA-P,P
SPELLSINGER [SPELLSINGER 1]  FOSTER, ALAN DEAN  C/P:2,P
SPELLSINGER’S SCHERZO [SPELLSINGER 4,5,6]  FOSTER, ALAN DEAN  H
TIME OF THE TRANSFERENCE, THE [SPELLSINGER 6]  FOSTER, ALAN DEAN  C/P,P,H
DARKSONG RISING [SPELLSONG 3]  MODESITT, L. E. JR.  C/P,P,H
SHADOW SORCERESS, THE [SPELLSONG 4]  MODESITT, L. E. JR.  C/P,P,H
SHADOWSINGER [SPELLSONG 5]  MODESITT, L. E. JR.  H
SOPRANO SORCERESS, THE [SPELLSONG 1]  MODESITT, L. E. JR.  H,P
SPELLSONG WAR, THE [SPELLSONG 2]  MODESITT, L. E. JR.  H,P
DEATH REIGN OF THE VAMPIRE KING [@SPIDER #1]  STOCKBRIDGE, GRANT  P
MASTER OF MEN! [@SPIDER #2]  STOCKBRIDGE, GRANT  P
CITY DESTROYER, THE [@SPIDER #3]  STOCKBRIDGE, GRANT  P
DEATH AND THE SPIDER [@SPIDER #4]  STOCKBRIDGE, GRANT  P
GATHERING OF THE SINISTER SIX, THE [SPIDER-MAN]  CASTRO, ADAM-TROY  P
MAYHEM IN MANHATTAN [SPIDER-MAN]  WEIN, LEN/WOLFMAN, MARV  P
MIDNIGHT JUSTICE [SPIDER-MAN]  DELRIO, MARTIN  C/P,P
SPIDER-MAN [SPIDER-MAN]  DAVID, PETER  P
SPIDER-MAN: CARNAGE IN NEW YORK [SPIDER-MAN]  MICHELINIE, DAVID/SMITH, DEAN WESLEY  P
SPIDER-MAN: THE VENOM FACTOR [SPIDER-MAN]  DUANE, DIANE  P
DELTA, THE [SPIDER WORLD 4]  WILSON, COLIN  P
DESERT, THE [SPIDER WORLD 1]  WILSON, COLIN  P
FORTRESS, THE [SPIDER WORLD 3]  WILSON, COLIN  P
TOWER, THE [SPIDER WORLD 2]  WILSON, COLIN  P
OCTOPUS AGENDA, THE [@SPIDERMAN]  DUANE, DIANE  H
WANTED: DEAD OR ALIVE [@SPIDERMAN]  GARDNER, CRAIG SHAW  H
PRINCESS OF DHAGABAD, THE [SPIRITS OF ANCIENT SANDS 1]  KASHINA, ANNA  H
SQUIRE [SQUIRE 1]  TELEP, PETER  C/P,P
SQUIRE’S BLOOD [SQUIRE 2]  TELEP, PETER  P
SQUIRE’S HONOR [SQUIRE 3]  TELEP, PETER  P
VAMPIRE IN MOSCOW [ST. JOHN THE PURSUER]  HENRICK, RICHARD  P
JOURNEY TO YESTERDAY [STACY GORDON 1]  SHIPLETT, JUNE LUND  P
RETURN TO YESTERDAY [STACY GORDON 2]  SHIPLETT, JUNE LUND  P

ADVENTURES OF THE STAINLESS STEEL RAT, THE [STAINLESS STEEL RAT 1,2,3] HARRISON, HARRY

C/P,P,H

STAINLESS STEEL RAT, THE [STAINLESS STEEL RAT 1]  HARRISON, HARRY  C/P,P
STAINLESS STEEL RAT FOR PRESIDENT, THE [STAINLESS STEEL RAT 5]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT GETS DRAFTED, THE [STAINLESS STEEL RAT 6]  HARRISON, HARRY  H:2,P
STAINLESS STEEL RAT GOES TO HELL, THE [STAINLESS STEEL RAT 8]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT IS BORN, A [STAINLESS STEEL RAT 0]  HARRISON, HARRY  H:2,P
STAINLESS STEEL RAT JOINS THE CIRCUS, THE [STAINLESS STEEL RAT 9]  HARRISON, HARRY  H

STAINLESS STEEL RAT JOINS THE CIRCUS, THE (PROOF)

[STAINLESS STEEL RAT 9] HARRISON, HARRY

SR-LP

STAINLESS STEEL RAT SAVES THE WORLD, THE [STAINLESS STEEL RAT 2]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT SINGS THE BLUES, THE [STAINLESS STEEL RAT 7]  HARRISON, HARRY  H:2,P
STAINLESS STEEL RAT WANTS YOU!, THE [STAINLESS STEEL RAT 4]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT’S REVENGE, THE [STAINLESS STEEL RAT 3]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL VISIONS [STAINLESS STEEL RAT]  HARRISON, HARRY  C/P,P,H
YOU CAN BE THE STAINLESS STEEL RAT [STAINLESS STEEL RAT]  HARRISON, HARRY  P
MONSTERS AND MAGICIANS [STAIRWAY TO FOREVER 2]  ADAMS, ROBERT  C/P:2,P
STAIRWAY TO FOREVER [STAIRWAY TO FOREVER 1]  ADAMS, ROBERT  C/P,SFWA-P,P
CALL TO ARMS [STAR COMMANDOS 9]  GRIFFIN, P. M.  C/P,P
COLONY IN PERIL [STAR COMMANDOS 2]  GRIFFIN, P. M.  P,SFWA-P
DEATH PLANET [STAR COMMANDOS 4]  GRIFFIN, P. M.  P
FIRE PLANET [STAR COMMANDOS 7]  GRIFFIN, P. M.  P
JUNGLE ASSAULT [STAR COMMANDOS 8]  GRIFFIN, P. M.  P
MIND SLAVER [STAR COMMANDOS 5]  GRIFFIN, P. M.  P
MISSION UNDERGROUND [STAR COMMANDOS 3]  GRIFFIN, P. M.  C/P:2,P
RETURN TO WAR [STAR COMMANDOS 6]  GRIFFIN, P. M.  C/P,P
STAR COMMANDOS [STAR COMMANDOS 1]  GRIFFIN, P. M.  C/P:2,P
STAR CONTROL INTERBELLUM [@STAR CONTROL]  QUICK, W. T.  P
DRAGON RIGGER [STAR DRAGONS 2]  CARVER, JEFFREY A.  C/P,P,H
DRAGONS IN THE STARS [STAR DRAGONS 1]  CARVER, JEFFREY A.  C/P,P
ALIEN DEATH FLEET [@STAR FRONTIER #1]  HUDSON, EDWARD S.  C/P:2,P
CYBORG KING, THE [STAR HAWKS 2]  GOULART, RON  P
EMPIRE 99 [STAR HAWKS 1]  GOULART, RON  C/P,P
LOST KING, THE [@STAR OF THE GUARDIANS #1]  WEIS, MARGARET  C/P:2,P
KING’S TEST [@STAR OF THE GUARDIANS #2]  WEIS, MARGARET  C/P:2,P
KING’S SACRIFICE [@STAR OF THE GUARDIANS #3]  WEIS, MARGARET  C/P:2,P
GHOST LEGION [@STAR OF THE GUARDIANS #4]  WEIS, MARGARET  C/P:2,P
HUNG OUT [STAR OF THE GUARDIANS MAG FORCE 7]  WEIS, MARGARET/PERRIN, DON  P

KNIGHTS OF THE BLACK EARTH, THE [STAR OF THE GUARDIANS MAG FORCE 7] WEIS, MARGARET/PERRIN, DON

C/P:2,P

ROBOT BLUES [STAR OF THE GUARDIANS MAG FORCE 7]  WEIS, MARGARET/PERRIN, DON  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 148
TITLE  AUTHOR  CODE
STAR PRECINCT [@STAR PRECINCT #1]  RANDLE, KEVIN/DRISCOLL, RICHARD  P
MINDSLAYER [@STAR PRECINCT #2]  RANDLE, KEVIN/DRUSCOLL, RICHARD  C/P,P
INSIDE JOB [@STAR PRECINCT #3]  RANDLE, KEVIN/DRISCOLL, RICHARD  C/P,P
MOTHER OF STORMS [@STAR REQUIEM #1]  COLE, ADRIAN  P,SFWA-P
THIEF OF DREAMS [@STAR REQUIEM #2]  COLE, ADRIAN  P,SFWA-P
WARLORD OF HEAVEN [@STAR REQUIEM #3]  COLE, ADRIAN  P,SFWA-P
LABYRINTH OF WORLDS [@STAR REQUIEM #4]  COLE, ADRIAN  C/P,SFWA-P,P
DOG FROM HELL, THE [@STAR RISK,LTD. 4]  BUNCH, CHRIS  P
DOUBLECROSS PROGRAM, THE [STAR RISK,LTD. 3]  BUNCH, CHRIS  P
GANGSTER CONSPIRACY, THE [@STAR RISK,LTD.]  PERRY, STEVE/PERRY, DAL  P
SCOUNDREL WORLDS, THE [STAR RISK,LTD. 2]  BUNCH, CHRIS  P
STAR RISK, LTD. [STAR RISK,LTD. 1]  BUNCH, CHRIS  P
SCORPION SHARDS [STAR SHARDS 1]  SHUSTERMAN, NEAL  P
SHATTERED SKY [STAR SHARDS 3]  SHUSTERMAN, NEAL  P
THIEF OF SOULS [STAR SHARDS 2]  SHUSTERMAN, NEAL  P
LAST PLANET, THE [STAR SOLDIERS 2]  NORTON, ANDRE  DD096,C/P:2,P
STAR GUARD [STAR SOLDIERS 1]  NORTON, ANDRE  C/P,H:2,P,DD199
STAR RANGERS [STAR SOLDIERS 2]  NORTON, ANDRE  C/P,P,H
STAR SOLDIERS [STAR SOLDIERS 1,2]  NORTON, ANDRE  P
SHAMAN, THE [@STAR STONE #1]  STASHEFF, CHRISTOPHER  P
SAGE, THE [@STAR STONE #2]  STASHEFF, CHRISTOPHER  H,P
ABODE OF LIFE, THE [STAR TREK]  CORREY, LEE  C/P:3,P
ACROSS THE UNIVERSE [STAR TREK]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  C/P,P
AMAZING STORIES, THE [STAR TREK]  STAR TREK BOOKS  LPA
AMOK TIME [STAR TREK]  STURGEON, THEODORE  SR-S
ARTICLES OF THE FEDERATION [STAR TREK]  DECANDIDO, KEITH R. A.  P
ASSIGNMENT: ETERNITY [STAR TREK]  COX, GREG  C/P,P
BADLANDS 1, THE [STAR TREK]  WRIGHT, SUSAN  C/P:2,P
BADLANDS 2, THE [STAR TREK]  WRIGHT, SUSAN  C/P,P
BALANCE OF TERROR [STAR TREK]  SCHNEIDER, PAUL  SR-S
BATTLESTATIONS! [STAR TREK]  CAREY, DIANE  C/P:2,P
BEST AND THE BRIGHTEST, THE [STAR TREK]  WRIGHT, SUSAN  P
BEST DESTINY [STAR TREK]  CAREY, DIANE  C/P,P,H
BETTER MAN, THE [STAR TREK]  WEINSTEIN, HOWARD  C/P:2,P
BEYOND UHURA [STAR TREK]  NICHOLS, NICHELLE  R/REF-H
BLACK FIRE [STAR TREK]  COOPER, SONNI  C/P:3,P
BLOODTHIRST [STAR TREK]  DILLARD, J. M.  C/P:2,P
BRAVE AND THE BOLD, BOOK ONE [@STAR TREK]  DECANDIDO, KEITH R. A.  P
BRAVE AND THE BOLD, BOOK TWO [@STAR TREK]  DECANDIDO, KEITH R. A.  P
BURNING DREAMS [STAR TREK]  BONANNO, MARGARET WANDER  P
CACOPHONY (CASSETTE) [STAR TREK]  MOLLOY, J. J.  SR-MM

CAPTAIN’S BLOOD [STAR TREK] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

H

CAPTAIN’S DAUGHTER, THE [STAR TREK]  DAVID, PETER  C/P:2,P

CAPTAIN’S GLORY [STAR TREK] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

H

CAPTAIN’S PERIL [@STAR TREK] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

CHAIN OF ATTACK [STAR TREK]  DEWEESE, GENE  C/P:3,P
CITY ON THE EDGE OF FOREVER, THE [STAR TREK]  ELLISON, HARLAN  SR-S
CITY ON THE EDGE OF FOREVER (FOTONOVEL) [STAR TREK]  RODDENBERRY, GENE  [C/RAN-P]
CONSTELLATIONS [STAR TREK]  PALMIERI, MARCO  LPA
CONVERSATIONAL KLINGON (CASSETTE) [STAR TREK]  OKRAND, MARC  SR-MM
CORONA [STAR TREK]  BEAR, GREG  C/P:3,P
COVENANT OF THE CROWN, THE [STAR TREK]  WEINSTEIN, HOWARD  C/P,P,H
CRISIS ON CENTAURUS [STAR TREK]  FERGUSON, BRAD  C/P:2,P
CROSSROAD [STAR TREK]  HAMBLY, BARBARA  C/P:2,P
CRY OF THE ONLIES, THE [STAR TREK]  KLASS, JUDY  C/P:2,P
DARK PASSIONS BOOK ONE [STAR TREK]  WRIGHT, SUSAN  C/P,P
DARK PASSIONS BOOK TWO [STAR TREK]  WRIGHT, SUSAN  C/P,P
DEADLY YEARS, THE (FOTONOVEL) [STAR TREK]  HARMON, DAVID P.  [C/RAN-P]
DEATH COUNT [STAR TREK]  GRAF, L. A.  C/P:2,P
DEATH’S ANGEL [STAR TREK]  SKY, KATHLEEN  C/P:2,P
DEBT OF HONOR [STAR TREK]  CLAREMONT, CHRIS  CX
DEEP DOMAIN [STAR TREK]  WEINSTEIN, HOWARD  C/P:3,P
DEMONS [STAR TREK]  DILLARD, J. M.  C/P:3,P
DEVIL IN THE DARK, THE [STAR TREK]  COON, GENE L.  SR-S
DEVIL WORLD [STAR TREK]  EKLUND, GORDON  C/P:2,P

DISINHERITED, THE [STAR TREK] DAVID, PETER/FRIEDMAN, MICHAEL JAN/

GREENBERGER, ROBERT C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 149
TITLE  AUTHOR  CODE
DOCTOR’S ORDERS [STAR TREK]  DUANE, DIANE  C/P:3,P
DOUBLE, DOUBLE [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:2,P
DREADNOUGHT! [STAR TREK]  CAREY, DIANE  C/P:3,P
DREAMS OF THE RAVEN [STAR TREK]  CARTER, CARMEN  C/P:2,P
DUTY, HONOR, REDEMPTION [@STAR TREK]  MCINTYRE, VONDA N.  LP
DWELLERS IN THE CRUCIBLE [STAR TREK]  BONANNO, MARGARET WANDER  C/P:3,P
ENEMY UNSEEN [STAR TREK]  MITCHELL, V. E.  C/P:3,P
ENGINES OF DESTINY [@STAR TREK]  DEWEESE, GENE  P
ENTERPRISE: THE FIRST ADVENTURE [STAR TREK]  MCINTYRE, VONDA N.  C/P:3,P
ENTROPY EFFECT, THE [STAR TREK]  MCINTYRE, VONDA N.  C/P:3,P
ERRAND OF MERCY [STAR TREK]  COON, GENE L.  SR-S
EUGENICS WARS, VOLUME 1 [STAR TREK]  COX, GREG  P
EUGENICS WARS, VOLUME 2 [STAR TREK]  COX, GREG  H,P
EX MACHINA [@STAR TREK]  BENNETT, CHRISTOPHER L.  P
FACES OF FIRE [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:2,P
FATE OF THE PHOENIX, THE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
FEARFUL SUMMONS, THE [STAR TREK]  FLINN, DENNY MARTIN  C/P:2,P

FEDERATION [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P,H

FINAL FRONTIER [STAR TREK]  CAREY, DIANE  C/P:2,P
FINAL NEXUS, THE [STAR TREK]  DEWEESE, GENE  C/P:2,P
FINAL REFLECTION, THE [STAR TREK]  FORD, JOHN M.  C/P:3,P
FIRESTORM [STAR TREK]  GRAF, L. A.  C/P:2,P
FIRST FRONTIER [STAR TREK]  CAREY, DIANE/KIRKLAND, DR. JAMES I.  C/P:2,P
FLAG FULL OF STARS, A [STAR TREK]  FERGUSON, BRAD  C/P:3,P
FROM THE DEPTHS [STAR TREK]  MILAN, VICTOR  C/P:2,P
GALACTIC WHIRLPOOL, THE [STAR TREK]  GERROLD, DAVID  C/P:3,P
GARTH OF IZAR [STAR TREK]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  P
GEMINI [STAR TREK]  BARR, MIKE W.  P
GHOST-WALKER [STAR TREK]  HAMBLY, BARBARA  C/P:3,P
GREAT STARSHIP RACE, THE [STAR TREK]  CAREY, DIANE  C/P:3,P
HEART OF THE SUN [STAR TREK]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  C/P,P
HOME IS THE HUNTER [STAR TREK]  KRAMER-ROLLS, DANA  C/P:2,P
HOW MUCH FOR JUST THE PLANET? [STAR TREK]  FORD, JOHN M.  C/P,P,H
ICE TRAP [STAR TREK]  GRAF, L. A.  C/P:2,P
IDIC EPIDEMIC, THE [STAR TREK]  LORRAH, JEAN  C/P:2,P
IN THE NAME OF HONOR [STAR TREK]  WARD, DAYTON  P
ISHMAEL [STAR TREK]  HAMBLY, BARBARA  C/P:3,P
JOY MACHINE, THE [STAR TREK]  GUNN, JAMES E.  C/P,P
KILLING TIME (ORIGINAL) [STAR TREK]  VAN HISE, DELLA  C/P:2,SR-P
KILLING TIME (REVISED) [STAR TREK]  VAN HISE, DELLA  C/P:3,P
KLINGON DICTIONARY, THE [STAR TREK]  OKRAND, MARC  R/REF-P
KLINGON DICTIONARY, THE (WITH 1992 ADDENDUM) [STAR TREK]  OKRAND, MARC  R/REF-P,C/REF-H
KLINGON GAMBIT, THE [STAR TREK]  VARDEMAN, ROBERT E.  C/P:3,P
KOBAYASHI MARU, THE [STAR TREK]  ECKLAR, JULIA  C/P:2,P
LAST ROUNDUP, THE [@STAR TREK]  GOLDEN, CHRISTIE  P
LEGACY [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:3,P
LIFE SIGNS [STAR TREK]  JENKINS, SUSAN/JENKINS, ROBERT  R/REF-H
LIVES OF DAX, THE [@STAR TREK]  PALMIERI, MARCO  LPA
LOST YEARS, THE [STAR TREK]  DILLARD, J. M.  C/P,P,H:2

MAKING OF STAR TREK, THE [STAR TREK] WHITFIELD, STEPHEN E./ RODDENBERRY, GENE

R/REF-P,C/REF-P:2

MAKING OF STAR TREK DEEP SPACE NINE, THE [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

R/REF-H

MAKING OF STAR TREK II: THE WRATH OF KHAN, THE [STAR TREK]  ASHERMAN, ALLAN  R/REF-P
MAKING OF STAR TREK: THE MOTION PICTURE, THE [STAR TREK]  SACKETT, SUSAN/RODDENBERRY, GENE  R/REF-P,C/REF-H:2

MEMORY PRIME [STAR TREK] REEVES-STEVENS, GARFIELD/ REEVES-STEVENS, JUDITH

C/P:3,P

MENAGERIE, THE [STAR TREK]  RODDENBERRY, GENE  SR-S
MIND MELD [STAR TREK]  VORNHOLT, JOHN  C/P,P
MINDSHADOW [STAR TREK]  DILLARD, J. M.  C/P:2,P
MIRROR, MIRROR [STAR TREK]  BIXBY, JEROME  SR-S
MISSION TO HORATIUS [STAR TREK]  REYNOLDS, MACK  H:2
MUDD IN YOUR EYE [STAR TREK]  OLTION, JERRY  C/P:2,P
MUDD’S ANGELS [STAR TREK]  LAWRENCE, J. A.  C/P:2,P
MUDD’S WOMEN [STAR TREK]  RODDENBERRY, GENE/KANDEL, S.  SR-S
MUTINY ON THE ENTERPRISE [STAR TREK]  VARDEMAN, ROBERT E.  C/P:3,P
NITPICKER’S GUIDE FOR CLASSIC TREKKERS, THE [STAR TREK]  FARRAND, PHIL  R/REF-H
OFFICIAL STAR TREK QUIZ BOOK, THE [STAR TREK]  MAGLIO, MITCHELL  [C/RAN-P]
PANDORA PRINCIPLE, THE [STAR TREK]  CLOWES, CAROLYN  C/P:3,P
PATRIAN TRANSGRESSION, THE [STAR TREK]  HAWKE, SIMON  C/P:2,P
PAWNS AND SYMBOLS [STAR TREK]  LARSON, MAJLISS  C/P:2,P
PERRY’S PLANET [STAR TREK]  HALDEMAN, JACK C. II  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 150
TITLE  AUTHOR  CODE
PHYSICS OF STAR TREK, THE [STAR TREK]  KRAUSS, LAWRENCE M.  R/REF-P,C/REF-H
PIECE OF THE ACTION, A [STAR TREK]  HARMON, DAVID P./COON, GENE L.  SR-S
PLANET OF JUDGMENT [STAR TREK]  HALDEMAN, JOE  C/P:2,P
POWER KLINGON (CASSETTE) [STAR TREK]  LEVINE, BARRY/OKRAND, MARC  SR-MM
PRICE OF THE PHOENIX, THE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P

PRIME DIRECTIVE [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

H:2,P

PROBE [STAR TREK]  BONANNO, MARGARET WANDER  C/P,P,H
PROMETHEUS DESIGN, THE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
RECOVERY [STAR TREK]  DILLARD, J. M.  C/P,P
RENEGADE [STAR TREK]  DEWEESE, GENE  C/P:2,P
RIFT, THE [STAR TREK]  DAVID, PETER  C/P:3,P

RINGS OF TAUTEE, THE [STAR TREK] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

RULES OF ENGAGEMENT [STAR TREK]  MORWOOD, PETER  C/P:2,P
SANCTUARY [STAR TREK]  VORNHOLT, JOHN  C/P:3,P
SAND AND STARS [STAR TREK]  POCKET BOOKS  LPA
SAREK [STAR TREK]  CRISPIN, A. C.  H:2,P
SHADOW LORD [STAR TREK]  YEP, LAURENCE  C/P:3,P
SHADOWS ON THE SUN [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:2,P
SHELL GAME [STAR TREK]  CRANDALL, MELISSA  C/P:2,P
SHORE LEAVE [STAR TREK]  STURGEON, THEODORE  SR-S
SPACE SEED [STAR TREK]  COON, GENE L./RODDENBERRY, GENE  SR-S

SPOCK, MESSIAH! [STAR TREK] COGSWELL, THEODORE R./ SPANO, CHARLES A.

C/P:3,P

SPOCK MUST DIE! [STAR TREK]  BLISH, JAMES  C/P:2,P
SPOCK’S WORLD [STAR TREK]  DUANE, DIANE  H:2,P
STAR TREK [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 2 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 3 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 4 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 5 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 6 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 7 [STAR TREK]  BLISH, JAMES  C/P:3,P
STAR TREK 8 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 9 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 10 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 11 [STAR TREK]  BLISH, JAMES  C/P,P,H
STAR TREK 12 [STAR TREK]  BLISH, JAMES/LAWRENCE, J. A.  C/P:2,P
STAR TREK (VIEWMASTER STEREO PICTURES) [STAR TREK]  ANONYMOUS  SR-MM

STAR TREK ALIENS & ARTIFACTS [STAR TREK] WESTMORE, MICHAEL/SIMS, ALAN/

LOOK, BRADLEY M./BIRNES, WILLIAM J.

R/REF-P

STAR TREK CATALOG [STAR TREK]  TURNBULL, GERRY  [C/RAN-H]
STAR TREK COMPENDIUM, THE (REVISED 1986) [STAR TREK]  ASHERMAN, ALLEN  R/REF-H
STAR TREK COMPENDIUM, THE (REVISED 1993) [STAR TREK]  ASHERMAN, ALLEN  R/REF-H
STAR TREK FEDERATION PASSPORT [STAR TREK]  DILLARD, J. M.  B
STAR TREK FEDERATION TRAVEL GUIDE [STAR TREK]  FRIEDMAN, MICHAEL JAN  B
STAR TREK GENERATIONS [STAR TREK]  DILLARD, J. M.  H
STAR TREK GUIDE, THE [STAR TREK]  RODDENBERRY, GENE  SR-S
STAR TREK II SHORT STORIES [STAR TREK]  ROTSLER, WILLIAM  P
STAR TREK II: THE WRATH OF KHAN [STAR TREK]  MCINTYRE, VONDA N.  C/P,P,H
STAR TREK III SHORT STORIES [STAR TREK]  ROTSLER, WILLIAM  P
STAR TREK III: THE SEARCH FOR SPOCK [STAR TREK]  MCINTYRE, VONDA N.  H:2,P
STAR TREK IV: THE VOYAGE HOME [STAR TREK]  MCINTYRE, VONDA N.  C/P:2,P,H

STAR TREK IV: THE VOYAGE HOME [STAR TREK] MEERSON, STEVE/KRIKES, PETER/

BENNETT, HARVE/MEYER, NICHOLAS

SR-S

STAR TREK IV: THE VOYAGE HOME (FOTONOVEL) [STAR TREK]  LERANGIS, PETER  LP
STAR TREK LOG 1 [STAR TREK]  FOSTER, ALAN DEAN  C/P:3,P
STAR TREK LOG ONE LOG TWO LOG THREE [STAR TREK]  FOSTER, ALAN DEAN  P
STAR TREK LOG 2 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 3 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 4 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG FOUR LOG FIVE LOG SIX [STAR TREK]  FOSTER, ALAN DEAN  P
STAR TREK LOG 5 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 6 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 7 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG SEVEN LOG EIGHT LOG NINE [STAR TREK]  FOSTER, ALAN DEAN  P
STAR TREK LOG 8 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 9 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 10 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK QUOTE JOURNAL [STAR TREK]  PARAMOUNT PICTURES  [C/RAN-H]
STAR TREK READER, THE [STAR TREK]  BLISH, JAMES  H:2
STAR TREK READER II, THE [STAR TREK]  BLISH, JAMES  H

The MITSFS Seriesdex  September 24, 2007  Page 151
TITLE  AUTHOR  CODE
STAR TREK READER III, THE [STAR TREK]  BLISH, JAMES  H
STAR TREK READER IV, THE [STAR TREK]  BLISH, JAMES  H
STAR TREK SPACEFLIGHT CHRONOLOGY [STAR TREK]  GOLDSTEIN, STAN/GOLDSTEIN, FRED  VLP

STAR TREK: STRANGE NEW WORLDS [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

HA:2,LPA

STAR TREK: STRANGE NEW WORLDS II [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

C/PA,PA

STAR TREK: STRANGE NEW WORLDS III [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS IV [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS V [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS VI [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS VII [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./ KASSIN, ELISA J./BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS 8 [STAR TREK] SMITH, DEAN WESLEY/KASSIN, ELISA J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS 9 [STAR TREK] SMITH, DEAN WESLEY/KASSIN, ELISA J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS 10 [STAR TREK]  SMITH, DEAN WESLEY/BLOCK, PAULA M.  LPA
STAR TREK: THE CLASSIC EPISODES 1 [STAR TREK]  BLISH, JAMES/LAWRENCE, J. A.  P
STAR TREK: THE CLASSIC EPISODES 2 [STAR TREK]  BLISH, JAMES  P
STAR TREK: THE CLASSIC EPISODES 3 [STAR TREK]  BLISH, JAMES/LAWRENCE, J. A.  P
STAR TREK: THE MAGNA [STAR TREK]  SHINSEI, SHINSEI  SCX
STAR TREK: THE MOTION PICTURE [STAR TREK]  RODDENBERRY, GENE  C/P,P,H
STAR TREK: THE MOTION PICTURE [STAR TREK]  RODDENBERRY, GENE/LIVINGSTON, HAROLD SR-S
STAR TREK: THE NEW VOYAGES [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  SR-PA,C/PA:2
STAR TREK: THE NEW VOYAGES 2 [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  SR-PA,C/PA:2
STAR TREK V: THE FINAL FRONTIER [STAR TREK]  DILLARD, J. M.  C/P:2,P
STAR TREK VI: THE UNDISCOVERED COUNTRY [STAR TREK]  DILLARD, J. M.  C/P:2,P
STARFLEET ACADEMY [STAR TREK]  CAREY, DIANE  C/P,P
STARFLEET YEAR ONE [STAR TREK]  FRIEDMAN, MICHAEL JAN  P
STARLESS WORLD, THE [STAR TREK]  EKLUND, GORDON  C/P:2,LP
STARSHIP TRAP, THE [STAR TREK]  GILDEN, MEL  C/P:2,P
STRANGERS FROM THE SKY [STAR TREK]  BONANNO, MARGARET WANDER  C/P:2,P
TALES FROM THE CAPTAIN’S TABLE [STAR TREK]  DECANDIDO, KEITH R. A.  LPA
TALES OF THE DOMINION WAR [@STAR TREK]  DECANDIDO, KEITH R. A.  LPA
TEARS OF THE SINGERS, THE [STAR TREK]  SNODGRASS, MELINDA M.  C/P:3,P
THREE-MINUTE UNIVERSE, THE [STAR TREK]  PAUL, BARBARA  C/P:2,P
TIME FOR YESTERDAY [STAR TREK]  CRISPIN, A. C.  C/P:2,P
TIMETRAP [STAR TREK]  DVORKIN, DAVID  C/P:2,P
TO REIGN IN HELL [STAR TREK]  COX, GREG  P
TRAITOR WINDS [STAR TREK]  GRAF, L. A.  C/P:2,P
TRANSFORMATIONS (CASSETTE) [STAR TREK]  STERN, DAVE  SR-MM
TREK: THE UNAUTHORIZED A–Z [STAR TREK]  SCHUSTER, HAL/RATHBONE, WENDY  R/REF-P,C/REF-P
TREK TO MADWORLD [STAR TREK]  GOLDIN, STEPHEN  C/P:2,P
TREKTOONS [STAR TREK]  LISTER, MARK  SCX
TRELLISANE CONFRONTATION, THE [STAR TREK]  DVORKIN, DAVID  C/P:3,P
TRIANGLE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
TROUBLE WITH TRIBBLES, THE [STAR TREK]  GERROLD, DAVID  R/REF-P,C/REF-P:2
TWILIGHT’S END [STAR TREK]  OLTION, JERRY  C/P,P
UHURA’S SONG [STAR TREK]  KAGAN, JANET  C/P:3,SFWA-P,P
ULTIMATE UNAUTHORIZED STAR TREK QUIZ BOOK, THE [STAR TREK]  BLY, ROBERT W.  [C/RAN-P]
VOYAGES OF IMAGINATION [STAR TREK]  AYERS, JEFF  R/REF-H
VULCAN! [STAR TREK]  SKY, KATHLEEN  C/P,P
VULCAN ACADEMY MURDERS, THE [STAR TREK]  LORRAH, JEAN  C/P:2,P
VULCAN’S FORGE [@STAR TREK]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  P
VULCAN’S GLORY [STAR TREK]  FONTANA, D. C.  C/P:2,P
VULCAN’S HEART [STAR TREK]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  P
WEB OF THE ROMULANS [STAR TREK]  MURDOCK, M. S.  C/P:3,P
WHERE NO MAN HAS GONE BEFORE [STAR TREK]  PEEPLES, SAMUEL A.  SR-S
WHY YOU SHOULD NEVER BEAM DOWN IN A RED SHIRT [STAR TREK]  BLY, ROBERT W.  R/REF-P,C/REF-P

WILLIAM SHATNER’S PERSONAL ACCOUNT OF THE MAKING OF STAR TREK 5 [STAR TREK] SHATNER, LISABETH

R/REF-P,C/REF-H

WINDOWS ON A LOST WORLD [STAR TREK]  MITCHELL, V. E.  C/P:2,P
WORLD WITHOUT END [STAR TREK]  HALDEMAN, JOE  C/P:2,P

WORLDS IN COLLISION [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

LP

WOUNDED SKY, THE [STAR TREK]  DUANE, DIANE  C/P:2,P
WRATH OF KHAN, THE (PHOTOSTORY) [STAR TREK]  ANOBILE, RICHARD J.  P
YESTERDAY’S SON [STAR TREK]  CRISPIN, A. C.  C/P:2,P
DUJONIAN’S HOARD [STAR TREK CAPTAIN’S TABLE 2/STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 152
TITLE  AUTHOR  CODE
FIRE SHIP [STAR TREK CAPTAIN’S TABLE 4/STAR TREK VOYAGER]  CAREY, DIANE  C/P,P

MIST, THE [STAR TREK CAPTAIN’S TABLE 3/STAR TREK DS9] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P:2,P

ONCE BURNED [STAR TREK CAPTAIN’S TABLE 5/ STAR TREK NEW FRONTIER] DAVID, PETER

C/P,P

WAR DRAGONS [STAR TREK CAPTAIN’S TABLE 1]  GRAF, L. A.  C/P:2,P
WHERE SEA MEETS SKY [STAR TREK CAPTAIN’S TABLE 6]  OLTION, JERRY  C/P:2,P
FIRE AND THE ROSE, THE [STAR TREK CRUCIBLE]  GEORGE, DAVID R. III  P
PROVENANCE OF SHADOWS [STAR TREK CRUCIBLE]  GEORGE, DAVID R. III  P
STAR TO EVERY WANDERING, THE [STAR TREK CRUCIBLE]  GEORGE, DAVID R. III  P
ANCIENT BLOOD [STAR TREK DAY OF HONOR 1/STAR TREK TNG]  CAREY, DIANE  C/P:2,P
ARMAGEDDON SKY [STAR TREK DAY OF HONOR 2/STAR TREK DS9]  GRAF, L. A.  C/P:2,P
HER KLINGON SOUL [STAR TREK DAY OF HONOR 3/STAR TREK VOYAGER] FRIEDMAN, MICHAEL JAN  C/P:2,P

TREATY’S LAW [STAR TREK DAY OF HONOR 4] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

BEHIND ENEMY LINES [@STAR TREK DOMINION WAR #1/STAR TREK TNG]  VORNHOLT, JOHN  P
CALL TO ARMS... [@STAR TREK DOMINION WAR #2/STAR TREK DS9]  CAREY, DIANE  C/P,P

TUNNEL THROUGH THE STARS [@STAR TREK DOMINION WAR #3/ STAR TREK TNG] VORNHOLT, JOHN

P

...SACRIFICE OF ANGELS [@STAR TREK DOMINION WAR #4/ STAR TREK DS9] CAREY, DIANE

P

ANTIMATTER [STAR TREK DS9]  VORNHOLT, JOHN  C/P:2,P
AVATAR, BOOK ONE [@STAR TREK DS9]  PERRY, S. D.  P
AVATAR, BOOK TWO [@STAR TREK DS9]  PERRY, S. D.  P
BETRAYAL [STAR TREK DS9]  TILTON, LOIS  C/P:2,P
BIG GAME, THE [STAR TREK DS9]  SCHOFIELD, SANDY  C/P:2,P
BIG GAME, THE (PROOF) [STAR TREK DS9]  SCHOFIELD, SANDY  SR-VLP
BLOODLETTER [STAR TREK DS9]  JETER, K. W.  C/P:2,P
DEVIL IN THE SKY [STAR TREK DS9]  COX, GREG/BETANCOURT, JOHN GREGORY  C/P:2,P
EMISSARY [STAR TREK DS9]  DILLARD, J. M.  C/P:2,P
EMISSARY (2 CASSETTES) [STAR TREK DS9]  DILLARD, J. M.  SR-MM
FALLEN HEROES [STAR TREK DS9]  AB HUGH, DAFYDD  C/P:2,P
FALLEN HEROES (PROOF) [STAR TREK DS9]  AB HUGH, DAFYDD  SR-VLP
FAR BEYOND THE STARS [@STAR TREK DS9]  BARNES, STEVEN  P
FERENGI RULES OF ACQUISITION, THE [STAR TREK DS9]  BEHR, IRA STEVEN  R/REF-P,C/REF-P
HEART OF THE WARRIOR, THE [STAR TREK DS9]  BETANCOURT, JOHN GREGORY  C/P:2,P
HOLLOW MEN [@STAR TREK DS9]  MCCORMACK, UNA  P
LAERTIAN GAMBLE, THE [STAR TREK DS9]  SHECKLEY, ROBERT  P
LEFT HAND OF DESTINY BOOK ONE, THE [@STAR TREK DS9]  HERTZLER, J. G./LANG, JEFFREY  P
LEFT HAND OF DESTINY BOOK TWO, THE [@STAR TREK DS9]  HERTZLER, J. G./LANG, JEFFREY  P
LEGENDS OF THE FERENGI [STAR TREK DS9]  QUARK  LP

LONG NIGHT, THE [STAR TREK DS9] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P:2,P

OBJECTIVE: BAJOR [STAR TREK DS9]  PEEL, JOHN  C/P,P
PRISONERS OF PEACE [@STAR TREK DS9]  PEEL, JOHN  H,LP
PROPHECY AND CHANGE [STAR TREK DS9]  PALMIERI, MARCO  LPA
PROUD HELIOS [STAR TREK DS9]  SCOTT, MELISSA  C/P:3,P
RISING SON [@STAR TREK DS9]  PERRY, S. D.  P
SARATOGA [STAR TREK DS9]  FRIEDMAN, MICHAEL JAN  P
SEARCH, THE [STAR TREK DS9]  CAREY, DIANE  C/P,P
SIEGE, THE [STAR TREK DS9]  DAVID, PETER  C/P:2,P
SPACE CAMP [STAR TREK DS9]  PEDERSEN, TED  LP
STAR GHOST, THE [@STAR TREK DS9]  STRICKLAND, BRAD  LP
STAR TREK DEEP SPACE NINE CALENDAR (1994) [STAR TREK DS9]  CALENDAR  CAL
STATION RAGE [STAR TREK DS9]  CAREY, DIANE  C/P,P
STITCH IN TIME, A [STAR TREK DS9]  ROBINSON, ANDREW J.  C/P,P
STOWAWAYS [STAR TREK DS9]  STRICKLAND, BRAD  H,LP
TEMPEST, THE [STAR TREK DS9]  WRIGHT, SUSAN  C/P,P
34TH RULE, THE [STAR TREK DS9]  SHIMERMAN, ARMIN/GEORGE III, DAVID R.  C/P,P
TRIAL BY ERROR [STAR TREK DS9]  GARLAND, MARK A.  C/P,P
TRIALS AND TRIBBLE-ATIONS [STAR TREK DS9]  CAREY, DIANE  P
TWIST OF FAITH [STAR TREK DS9]  POCKET BOOKS  LPA
UNITY [@STAR TREK DS9]  PERRY, S. D.  P
VALHALLA [STAR TREK DS9]  ARCHER, NATHAN  C/P,P
VENGEANCE [@STAR TREK DS9]  AB HUGH, DAFYDD  P
WARCHILD [STAR TREK DS9]  FRIESNER, ESTHER M.  C/P:2,P
WARCHILD (PROOF) [STAR TREK DS9]  FRIESNER, ESTHER M.  SR-VLP
WARPATH [@STAR TREK DS9]  MACK, DAVID  P
WARPED [STAR TREK DS9]  JETER, K. W.  H,P
WAY OF THE WARRIOR, THE [STAR TREK DS9]  CAREY, DIANE  P
WHAT YOU LEAVE BEHIND [@STAR TREK DS9]  CAREY, DIANE  P

WRATH OF THE PROPHETS [STAR TREK DS9] DAVID, PETER/FRIEDMAN, MICHAEL JAN/

GREENBERGER, ROBERT C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 153
TITLE  AUTHOR  CODE

FALL OF TEROK NOR [@STAR TREK DS9 MILLENIUM #1] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

WAR OF THE PROPHETS, THE [@STAR TREK DS9 MILLENIUM #2] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

INFERNO [@STAR TREK DS9 MILLENIUM #3] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

TWILIGHT [STAR TREK DS9 MISSION GAMMA #1]  GEORGE, DAVID R. III  C/P,P
THIS GRAY SPIRIT [STAR TREK DS9 MISSION GAMMA #2]  JARMAN, HEATHER  C/P,P
CATHEDRAL [@STAR TREK DS9 MISSION GAMMA #3]  MARTIN, MICHAEL A./MANGELS, ANDY  C/P,P
LESSER EVIL [STAR TREK DS9 MISSION GAMMA #4]  SIMPSON, ROBERT  C/P,P
CONQUERED, THE [@STAR TREK DS9 REBELS #1]  AB HUGH, DAFYDD  P
COURAGEOUS, THE [@STAR TREK DS9 REBELS #2]  AB HUGH, DAFYDD  C/P,P
LIBERATED, THE [@STAR TREK DS9 REBELS #3]  AB HUGH, DAFYDD  P
ABYSS [STAR TREK DS9 SECTION 31]  WEDDLE, DAVID/LANG, JEFFREY  P

CARDASSIA: LOTUS FLOWER, THE [STAR TREK DS9 WORLDS 1]/ ANDOR: PARADIGM [STAR TREK DS9 WORLDS 1] MCCORMACK, UNA/JARMAN, HEATHER

C/P,P

FERENGINAR: SATISFACTION IS NOT GUARANTEED [STAR TREK DS9 WORLDS 3]/ DOMINION, THE: OLYMPUS DESCENDING [STAR TREK DS9 WORLDS 3] DECANDIDO, KEITH R. A./ GEORGE, DAVID R. III

C/P,P

TRILL: UNJOINED [STAR TREK DS9 WORLDS 2]/ BAJOR: FRAGMENTS AND OMENS [STAR TREK DS9 WORLDS 2] MARTIN, MICHAEL A./MANGELS, ANDY/

KYM, J. NOAH

C/P,P

BROKEN BOW [@STAR TREK ENTERPRISE]  CAREY, DIANE  P

BY THE BOOK [STAR TREK ENTERPRISE] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

DAEDALUS’S CHILDREN, PART TWO [STAR TREK ENTERPRISE]  STERN, DAVE  P
DEADALUS [STAR TREK ENTERPRISE]  STERN, DAVE  P
ENTERPRISE: THE EXPANSE [STAR TREK ENTERPRISE]  DILLARD, J. M.  LP
GOOD THAT MEN DO, THE [STAR TREK ENTERPRISE]  MANGELS, ANDY/MARTIN, MICHAEL A.  P
LAST FULL MEASURE [STAR TREK ENTERPRISE]  MARTIN, MICHAEL A./MANGELS, ANDY  P
ROSETTA [STAR TREK ENTERPRISE]  STERN, DAVE  P
SHOCKWAVE [STAR TREK ENTERPRISE]  RUDITIS, PAUL  H,@P
SURAK’S SOUL [STAR TREK ENTERPRISE]  DILLARD, J. M.  P
WHAT PRICE HONOR? [STAR TREK ENTERPRISE]  STERN, DAVE  P
SEEDS OF RAGE [@STAR TREK ERRAND OF FURY #1]  RYAN, KEVIN  P
DEMANDS OF HONOR [@STAR TREK ERRAND OF FURY #2]  RYAN, KEVIN  P
EDGE OF THE SWORD, THE [@STAR TREK ERRAND OF VENGEANCE #1]  RYAN, KEVIN  P
KILLING BLOW [@STAR TREK ERRAND OF VENGEANCE #2]  RYAN, KEVIN  P
RIVER OF BLOOD [@STAR TREK ERRAND OF VENGEANCE #3]  RYAN, KEVIN  P
ONE SMALL STEP [STAR TREK GATEWAYS #1]  WRIGHT, SUSAN  P
CHAINMAIL [STAR TREK GATEWAYS #2/STAR TREK CHALLENGER]  CAREY, DIANE  P
DOORS INTO CHAOS [STAR TREK GATEWAYS #3/STAR TREK TNG]  GREENBERGER, ROBERT  P

DEMONS OF AIR AND DARKNESS [STAR TREK GATEWAYS #4/ STAR TREK DS9] DECANDIDO, KEITH R. A.

P

NO MAN’S LAND [STAR TREK GATEWAYS #5/STAR TREK VOYAGER]  GOLDEN, CHRISTIE  P
COLD WARS [STAR TREK GATEWAYS #6/STAR TREK NEW FRONTIER]  DAVID, PETER  P
WHAT LAY BEYOND [STAR TREK GATEWAYS #7]  PARAMOUNT PICTURES  HA,PA
ENEMY TERRITORY [@STAR TREK I.K.S. GORKON 3]  DECANDIDO, KEITH R. A.  P
GOOD DAY TO DIE, A [@STAR TREK I.K.S. GORKON 1]  DECANDIDO, KEITH R. A.  P
HONOR BOUND [@STAR TREK I.K.S. GORKON 2]  DECANDIDO, KEITH R. A.  P
FINAL FURY, THE [STAR TREK INVASION! 4/STAR TREK VOYAGER]  AB HUGH, DAFYDD  C/P,P
FIRST STRIKE [STAR TREK INVASION! 1]  CAREY, DIANE  C/P,P

SOLDIERS OF FEAR, THE [STAR TREK INVASION! 2/STAR TREK TNG] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

TIME’S ENEMY [STAR TREK INVASION! 3/STAR TREK DS9]  GRAF, L. A.  C/P,P
JANUS GATE, THE [STAR TREK JANUS GATE 1,2,3]  GRAF, L. A.  H
PRESENT TENSE [@STAR TREK JANUS GATE #1]  GRAF, L. A.  P
FUTURE IMPERFECT [@STAR TREK JANUS GATE #2]  GRAF, L. A.  P
PAST PROLOGUE [@STAR TREK JANUS GATE #3]  GRAF, L. A.  P

STAR TREK: THE KEY COLLECTION [STAR TREK KEY COLLECTION 1] RODDENBERRY, GENE/ZACCARA, NEVIO/

GIOLITTI, ALBERTO

CX

STAR TREK: THE KEY COLLECTION: VOLUME 2 [STAR TREK KEY COLLECTION 2] RODDENBERRY, GENE/GIOLITTI, ALBERTO

CX

STAR TREK: THE KEY COLLECTION VOLUME 3 [STAR TREK KEY COLLECTION 3] GOLD KEY COMICS

CX

STAR TREK: THE KEY COLLECTION: VOLUME 4 [STAR TREK KEY COLLECTION 4] RODDENBERRY, GENE/ZACCARA, NEVIO/

GIOLITTI, ALBERTO

CX

STAR TREK: THE KEY COLLECTION: VOLUME 5 [STAR TREK KEY COLLECTION 5] RODDENBERRY, GENE/MCWILLIAMS, AL/ ZACCARA, NEVIO/GIOLITTI, ALBERTO

CX

SUNDERED, THE [@STAR TREK LOST ERA (2298 AD)]  MARTIN, MICHAEL A./MANGELS, ANDY  P
SERPENTS AMONG THE RUINS [@STAR TREK LOST ERA (2311 AD)]  GEORGE, DAVID R. III  P
ART OF THE IMPOSSIBLE, THE [@STAR TREK LOST ERA (2328-46 AD)]  DECANDIDO, KEITH R. A.  P
WELL OF SOULS [@STAR TREK LOST ERA (2336 AD)]  BICK, ILSA J.  P
DENY THY FATHER [@STAR TREK LOST ERA (2355-57 AD)]  MARIOTTE, JEFF  P
CATALYST OF SORROWS [@STAR TREK LOST ERA (2360 AD)]  BONANNO, MARGARET WANDER  P

The MITSFS Seriesdex  September 24, 2007  Page 154
TITLE  AUTHOR  CODE

DARK VICTORY [STAR TREK MIRROR 2] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

GLASS EMPIRES [STAR TREK MIRROR]  POCKET BOOKS  LPA
OBSIDIAN ALLIANCES [STAR TREK MIRROR]  POCKET BOOKS  LPA

PRESERVER [STAR TREK MIRROR 3] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

SPECTRE [STAR TREK MIRROR 1] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

C/P,P

REPUBLIC [@STAR TREK MY BROTHER’S KEEPER #1]  FRIEDMAN, MICHAEL JAN  P
CONSTITUTION [@STAR TREK MY BROTHER’S KEEPER #2]  FRIEDMAN, MICHAEL JAN  P
ENTERPRISE [@STAR TREK MY BROTHER’S KEEPER #3]  FRIEDMAN, MICHAEL JAN  P
WAGON TRAIN TO THE STARS [@STAR TREK NEW EARTH #1]  CAREY, DIANE  P
BELLE TERRE [@STAR TREK NEW EARTH #2]  SMITH, DEAN WESLEY/CAREY, DIANE  C/P,P
ROUGH TRAILS [@STAR TREK NEW EARTH #3]  GRAF, L. A.  C/P,P
FLAMING ARROW, THE [@STAR TREK NEW EARTH #4]  OLTION, KATHY/OLTION, JERRY  C/P,P

THIN AIR [@STAR TREK NEW EARTH #5] RUSCH, KRISTINE KATHRYN/ SMITH, DEAN WESLEY

C/P:2,P

CHALLENGER [@STAR TREK NEW EARTH #6]  CAREY, DIANE  C/P,P
AFTER THE FALL [STAR TREK NEW FRONTIER 15]  DAVID, PETER  H
GODS ABOVE [@STAR TREK NEW FRONTIER 13]  DAVID, PETER  P
MISSING IN ACTION [STAR TREK NEW FRONTIER 16]  DAVID, PETER  H
NO LIMITS [@STAR TREK NEW FRONTIER]  DAVID, PETER  LPA
STONE AND ANVIL [@STAR TREK NEW FRONTIER 14]  DAVID, PETER  P
HOUSE OF CARDS [@STAR TREK NEW FRONTIER #1]  DAVID, PETER  C/P,P
INTO THE VOID [@STAR TREK NEW FRONTIER #2]  DAVID, PETER  C/P,P
TWO-FRONT WAR, THE [@STAR TREK NEW FRONTIER #3]  DAVID, PETER  C/P,P
END GAME [@STAR TREK NEW FRONTIER #4]  DAVID, PETER  C/P,P
MARTYR [@STAR TREK NEW FRONTIER #5]  DAVID, PETER  P
FIRE ON HIGH [@STAR TREK NEW FRONTIER #6]  DAVID, PETER  P
QUIET PLACE, THE [@STAR TREK NEW FRONTIER #7]  DAVID, PETER  P
DARK ALLIES [@STAR TREK NEW FRONTIER #8]  DAVID, PETER  P
EXCALIBUR: REQUIEM [@STAR TREK NEW FRONTIER #9]  DAVID, PETER  P
EXCALIBUR: RENAISSANCE [@STAR TREK NEW FRONTIER #10]  DAVID, PETER  P
EXCALIBUR: RESTORATION [@STAR TREK NEW FRONTIER #11]  DAVID, PETER  P
ASHES OF EDEN, THE [STAR TREK ODYSSEY 1]  SHATNER, WILLIAM  H,P
AVENGER [STAR TREK ODYSSEY 3]  SHATNER, WILLIAM  P

RETURN, THE [STAR TREK ODYSSEY 2] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

EMPTY CHAIR, THE [STAR TREK RIHANNSU 5]  DUANE, DIANE  P
HONOR BLADE [STAR TREK RIHANNSU 4]  DUANE, DIANE  C/P,P
MY ENEMY, MY ALLY [STAR TREK RIHANNSU 1]  DUANE, DIANE  C/P:3,P
RIHANNSU [STAR TREK RIHANNSU 1,2,3,4]  DUANE, DIANE  H
ROMULAN WAY, THE [STAR TREK RIHANNSU 2]  DUANE, DIANE/MORWOOD, PETER  C/P:2,P
SWORDHUNT [STAR TREK RIHANNSU 3]  DUANE, DIANE  C/P:2,P
CASE OF THE COLONIST’S CORPSE, THE [@STAR TREK SAMUEL COGLEY]  INGERSOLL, BOB/ISABELLA, TONY  P
AFTERMATH [STAR TREK S.C.E. 8]  POCKET BOOKS  LPA
FOUNDATIONS [@STAR TREK S.C.E. 5]  WARD, DAYTON/DILMORE, KEVIN  P

HAVE TECH, WILL TRAVEL [@STAR TREK S.C.E. #1] DECANDIDO, KEITH R. A./DILMORE, KEVIN/ GOLDEN, CHRISTIE/SMITH, DEAN WESLEY/ WARD, DAYTON

P

MIRACLE WORKERS [@STAR TREK S.C.E. #2] DECANDIDO, KEITH R. A./DILMORE, KEVIN/

MACK, DAVID/WARD, DAYTON

P

SOME ASSEMBLY REQUIRED [@STAR TREK S.C.E. #3] BRODEUR, GREG/

CIENCIN, SCOTT/DECANDIDO, KEITH R. A./ GALANTER, DAVE/JOLLEY, DAN/ ROSENBERG, AARON

P

NO SURRENDER [@STAR TREK S.C.E. #4] COLLINS, MIKE/

EDGINTON, IAN/GREENBARGER, ROBERT/ HAUMAN, GLENN/MARIOTTE, JEFF

P

WILDFIRE [@STAR TREK S.C.E. #6] DECANDIDO, KEITH R. A./MACK, DAVID/ YORK, J. STEVEN/YORK, CHRISTINA F.

P

BREAKDOWNS [@STAR TREK S.C.E. #7] CIENCIN, SCOTT/DECANDIDO, KEITH R. A./ DILMORE, KEVIN/JARMAN, HEATHER/ WARD, DAYTON

P

CLOAK [STAR TREK SECTION 31]  PERRY, S. D.  C/P,P
CRISIS ON VULCAN [@STAR TREK STARFLEET ACADEMY #1]  STRICKLAND, BRAD/STRICKLAND, BARBARA LP
AFTERSHOCK [@STAR TREK STARFLEET ACADEMY #2]  VORNHOLT, JOHN  LP
CADET KIRK [@STAR TREK STARFLEET ACADEMY #3]  CAREY, DIANE  LP

STAR LOST, THE [STAR TREK,THE NEXT GENERATION] FRIEDMAN, MICHAEL JAN/KRAUSE, PETER/

MARCO, PABLO CX

The MITSFS Seriesdex  September 24, 2007  Page 155
TITLE  AUTHOR  CODE
ORION’S HOUNDS [STAR TREK TITAN]  BENNETT, CHRISTOPHER L.  P
RED KING, THE [STAR TREK TITAN]  MANGELS, ANDY/MARTIN, MICHAEL A.  P
TAKING WING [@STAR TREK TITAN 1]  MARTIN, MICHAEL A./MANGELS, ANDY  P
ALL GOOD THINGS [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  H:2,P
BALANCE OF POWER [STAR TREK TNG]  AB HUGH, DAFYDD  C/P:2,P
BATTLE OF BETAZED, THE [STAR TREK TNG]  DOUGLAS, CHARLOTTE/KEARNEY, SUSAN  C/P,P
BEGINNINGS [STAR TREK TNG]  CARLIN, MIKE  CX
BIRTHRIGHT [STAR TREK TNG]  BRAGA, BRANDON  SR-S
BLAZE OF GLORY [STAR TREK TNG]  HAWKE, SIMON  C/P:2,P
BOOGEYMEN [STAR TREK TNG]  GILDEN, MEL  C/P:2,P
BURIED AGE, THE [@STAR TREK TNG]  BENNETT, CHRISTOPHER L.  P
CALL TO DARKNESS, A [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P:2,P
CAPTAIN’S HONOR, THE [STAR TREK TNG]  DVORKIN, DAVID/DVORKIN, DANIEL  C/P:2,P
CHAINS OF COMMAND [STAR TREK TNG]  MCCAY, BILL/FLOOD, ELOISE  C/P:3,P
CHILDREN OF HAMLIN, THE [STAR TREK TNG]  CARTER, CARMEN  C/P:3,P
CONTAMINATION [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
CROSSOVER [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  P
DARK MIRROR [STAR TREK TNG]  DUANE, DIANE  H:2,P
DEATH IN WINTER [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  H
DEATH OF PRINCES, THE [STAR TREK TNG]  PEEL, JOHN  C/P:2,P
DEBTORS’ PLANET [STAR TREK TNG]  THOMPSON, W. R.  C/P:2,P
DESCENT [STAR TREK TNG]  CAREY, DIANE  C/P,P
DEVIL’S HEART, THE [STAR TREK TNG]  CARTER, CARMEN  C/P,P,H
DIPLOMATIC IMPLAUSIBILITY [STAR TREK TNG]  DECANDIDO, KEITH R. A.  P
DO COMETS DREAM? [@STAR TREK TNG]  SOMTOW, S. P.  P

DOOMSDAY WORLD [STAR TREK TNG] CARTER, CARMEN/

DAVID, PETER/FRIEDMAN, MICHAEL JAN/ GREENBERGER, ROBERT

C/P:2,P

DRAGON’S HONOR [STAR TREK TNG]  JOHNSON, KIJ/COX, GREG  C/P:2,P

DYSON SPHERE [@STAR TREK TNG] PELLEGRINO, CHARLES/ ZEBROWSKI, GEORGE

C/P:2,P

ENCOUNTER AT FARPOINT [STAR TREK TNG]  GERROLD, DAVID  C/P:2,P
EXILES [STAR TREK TNG]  WEINSTEIN, HOWARD  C/P:2,P
EYES OF THE BEHOLDERS, THE [STAR TREK TNG]  CRISPIN, A. C.  C/P:2,P
FOREIGN FOES [STAR TREK TNG]  GALANTER, DAVE/BRODEUR, GREG  C/P:2,P
FORGOTTEN WAR, THE [@STAR TREK TNG]  FORSTCHEN, WILLIAM R.  C/P,P
FORTUNE’S LIGHT [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P:2,P
FURY SCORNED, A [STAR TREK TNG]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  C/P:2,P
GEMWORLD, BOOK ONE [@STAR TREK TNG]  VORNHOLT, JOHN  C/P,P
GEMWORLD, BOOK TWO [@STAR TREK TNG]  VORNHOLT, JOHN  C/P,P
GENESIS FORCE [STAR TREK TNG]  VORNHOLT, JOHN  P
GENESIS WAVE BOOK ONE, THE [STAR TREK TNG]  VORNHOLT, JOHN  P
GENESIS WAVE BOOK TWO, THE [STAR TREK TNG]  VORNHOLT, JOHN  P
GENESIS WAVE BOOK THREE, THE [STAR TREK TNG]  VORNHOLT, JOHN  H,P
GHOST SHIP [STAR TREK TNG]  CAREY, DIANE  C/P:2,P
GROUNDED [STAR TREK TNG]  BISCHOFF, DAVID F.  C/P:2,P
GUISES OF THE MIND [STAR TREK TNG]  NEASON, REBECCA  C/P:2,P
GULLIVER’S FUGITIVES [STAR TREK TNG]  SHAREE, KEITH  C/P:2,P
HARD RAIN, A [STAR TREK TNG]  SMITH, DEAN WESLEY  C/P,P
HERE THERE BE DRAGONS [STAR TREK TNG]  PEEL, JOHN  C/P:3,P
I, Q [STAR TREK TNG]  DE LANCIE, JOHN/DAVID, PETER  P
IMBALANCE [STAR TREK TNG]  MITCHELL, V. E.  C/P:2,P
IMMORTAL COIL [STAR TREK TNG]  LANG, JEFFREY  P
IMZADI [STAR TREK TNG]  DAVID, PETER  H:2,P
IMZADI FOREVER [STAR TREK TNG]  DAVID, PETER  LP
INFILTRATOR [STAR TREK TNG]  THOMPSON, W. R.  C/P,P
INTELLIVORE [STAR TREK TNG]  DUANE, DIANE  C/P:2,P
INTO THE NEBULA [STAR TREK TNG]  DEWEESE, GENE  C/P:2,P
KAHLESS [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  P
LAST STAND, THE [STAR TREK TNG]  FERGUSON, BRAD  C/P:3,P
MAKE IT SO [STAR TREK TNG]  ROBERTS, WESS/ROSS, BILL  H
MASKS [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
METAMORPHOSIS [STAR TREK TNG]  LORRAH, JEAN  C/P:2,P
NEMESIS [STAR TREK TNG]  VORNHOLT, JOHN  LP
NIGHTSHADE [STAR TREK TNG]  HAMILTON, LAURELL K.  C/P:2,P

NITPICKER’S GUIDE FOR NEXT GENERATION TREKKERS, THE [STAR TREK TNG] FARRAND, PHIL

R/REF-H

PEACEKEEPERS, THE [STAR TREK TNG]  DEWEESE, GENE  C/P:3,P
PERCHANCE TO DREAM [STAR TREK TNG]  WEINSTEIN, HOWARD  C/P:2,P
PLANET X [@STAR TREK TNG/X-MEN]  FRIEDMAN, MICHAEL JAN  C/P,P
POSSESSION [STAR TREK TNG]  DILLARD, J. M./O’MALLEY, KATHLEEN  C/P:2,P
POWER HUNGRY [STAR TREK TNG]  WEINSTEIN, HOWARD  C/P:3,P
Q-IN-LAW [STAR TREK TNG]  DAVID, PETER  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 156
TITLE  AUTHOR  CODE
Q-SPACE [STAR TREK TNG]  COX, GREG  C/P,P
Q-SQUARED [STAR TREK TNG]  DAVID, PETER  H:2,P
Q-STRIKE [@STAR TREK TNG]  COX, GREG  C/P,P
Q-ZONE [STAR TREK TNG]  COX, GREG  C/P:2,P
RELICS [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P,P
REQUIEM [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN/RYAN, KEVIN  C/P:3,P
REUNION [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P:2,P
ROCK AND A HARD PLACE, A [STAR TREK TNG]  DAVID, PETER  C/P:2,P
ROGUE SAUCER [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
ROMULAN PRIZE, THE [STAR TREK TNG]  HAWKE, SIMON  C/P:2,P
ROMULAN STRATAGEM, THE [STAR TREK TNG]  GREENBERGER, ROBERT  C/P:2,P
SHIP OF THE LINE [STAR TREK TNG]  CAREY, DIANE  P
SINS OF COMMISSION [STAR TREK TNG]  WRIGHT, SUSAN  C/P:2,P
SPARTACUS [STAR TREK TNG]  MANCOUR, T. L.  C/P:3,P
STAR TREK: FIRST CONTACT [STAR TREK TNG]  DILLARD, J. M.  H:2
STAR TREK INSURRECTION [STAR TREK TNG]  DILLARD, J. M.  H

STAR TREK KLINGON [@STAR TREK TNG] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

STAR TREK: NEMESIS [STAR TREK TNG]  DILLARD, J. M.  H

STAR TREK THE NEXT GENERATION TECHNICAL MANUAL [STAR TREK TNG] STERNBACH, RICK/OKUDA, MICHAEL

R/REF-H,C/REF-H

STAR TREK THE NEXT GENERATION TECHNICAL MANUAL (FIFTH SEASON EDITION) [STAR TREK TNG] STERNBACH, RICK/OKUDA, MICHAEL

R/REF-H

STRIKE ZONE [STAR TREK TNG]  DAVID, PETER  C/P:3,P
SURVIVORS [STAR TREK TNG]  LORRAH, JEAN  C/P:2,P
TO STORM HEAVEN [STAR TREK TNG]  FRIESNER, ESTHER M.  C/P,P
TOOTH AND CLAW [STAR TREK TNG]  DURGIN, DORANNA  P
TREK: THE NEXT GENERATION REVISED EDITION [STAR TREK TNG]  VAN HISE, JAMES  R/REF-H
TREKS NOT TAKEN [STAR TREK TNG]  BOYETT, STEVEN R.  LP
20TH CENTURY COMPUTERS AND HOW THEY WORKED [STAR TREK TNG]  FLYNN, JENNIFER  R/REF-H
UNIFICATION [STAR TREK TNG]  TAYLOR, JERI  C/P:3,P
VENDETTA [STAR TREK TNG]  DAVID, PETER  C/P:2,P
WAR DRUMS [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
DOUBLE OR NOTHING [@STAR TREK TNG DOUBLE HELIX 5]  DAVID, PETER  P

FIRST VIRTUE, THE [@STAR TREK TNG DOUBLE HELIX 6] FRIEDMAN, MICHAEL JAN/ GOLDEN, CHRISTIE

P

INFECTION [@STAR TREK TNG DOUBLE HELIX 1]  BETANCOURT, JOHN GREGORY  P
QUARANTINE [@STAR TREK TNG DOUBLE HELIX 4]  VORNHOLT, JOHN  P
RED SECTOR [@STAR TREK TNG DOUBLE HELIX 3]  CAREY, DIANE  P

VECTORS [@STAR TREK TNG DOUBLE HELIX 2] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

DEAD ZONE [@STAR TREK TNG MAXIMUM WARP #1]  GALANTER, DAVE/BRODEUR, GREG  P
FOREVER DARK [@STAR TREK TNG MAXIMUM WARP #2]  GALANTER, DAVE/BRODEUR, GREG  P
ROGUE [@STAR TREK TNG SECTION 31]  MANGELS, ANDY/MARTIN, MICHAEL A.  P
WORF’S FIRST ADVENTURE [@STAR TREK TNG STARFLEET ACADEMY #1]  DAVID, PETER  H:3,LP
LINE OF FIRE [@STAR TREK TNG STARFLEET ACADEMY #2]  DAVID, PETER  H,LP
SURVIVAL [@STAR TREK TNG STARFLEET ACADEMY #3]  DAVID, PETER  H,LP
CAPTURE THE FLAG [@STAR TREK TNG STARFLEET ACADEMY #4]  VORNHOLT, JOHN  LP
ATLANTIS STATION [@STAR TREK TNG STARFLEET ACADEMY #5]  MITCHELL, V. E.  H,LP

MYSTERY OF THE MISSING CREW [@STAR TREK TNG STARFLEET ACADEMY #6] FRIEDMAN, MICHAEL JAN

H,LP

STARFALL [@STAR TREK TNG STARFLEET ACADEMY #8]  STRICKLAND, BRAD/STRICKLAND, BARBARA LP
NOVA COMMAND [@STAR TREK TNG STARFLEET ACADEMY #9]  STRICKLAND, BRAD/STRICKLAND, BARBARA LP
LOYALTIES [@STAR TREK TNG STARFLEET ACADEMY #10]  BARNES-SVARNEY, PATRICIA  LP
CROSSFIRE [@STAR TREK TNG STARFLEET ACADEMY #11]  VORNHOLT, JOHN  LP
BREAKAWAY [@STAR TREK TNG STARFLEET ACADEMY #12]  WEISS, BOBBI JG/WEISS, DAVID CODY  H,P
ENIGMA [@STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
MAKER [@STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
OBLIVION [@STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
VALIANT, THE [STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
GAUNTLET [@STAR TREK TNG STARGAZER #1]  FRIEDMAN, MICHAEL JAN  P
PROGENITOR [@STAR TREK TNG STARGAZER #2]  FRIEDMAN, MICHAEL JAN  P
THREE [@STAR TREK TNG STARGAZER #3]  FRIEDMAN, MICHAEL JAN  P
TIME FOR WAR, A TIME FOR PEACE, A [STAR TREK TNG TIME TO 9]  DECANDIDO, KEITH R. A.  P
TIME TO BE BORN, A [@STAR TREK TNG TIME TO 1]  VORNHOLT, JOHN  P
TIME TO DIE, A [@STAR TREK TNG TIME TO 2]  VORNHOLT, JOHN  P
TIME TO HARVEST, A [@STAR TREK TNG TIME TO 4]  WARD, DAYTON/DILMORE, KEVIN  P
TIME TO HATE, A [@STAR TREK TNG TIME TO 6]  GREENBERGER, ROBERT  P
TIME TO HEAL, A [STAR TREK TNG TIME TO 8]  MACK, DAVID  P
TIME TO KILL, A [@STAR TREK TNG TIME TO 7]  MACK, DAVID  P
TIME TO LOVE, A [@STAR TREK TNG TIME TO 5]  GREENBERGER, ROBERT  P
TIME TO SOW, A [@STAR TREK TNG TIME TO 3]  WARD, DAYTON/DILMORE, KEVIN  P
HARBINGER [STAR TREK VANGUARD]  MACK, DAVID  P
REAP THE WHIRLWIND [@STAR TREK VANGUARD]  MACK, DAVID  P

The MITSFS Seriesdex  September 24, 2007  Page 157
TITLE  AUTHOR  CODE
SUMMON THE THUNDER [STAR TREK VANGUARD]  WARD, DAYTON/DILMORE, KEVIN  P
BATTLE LINES [@STAR TREK VOYAGER]  GALANTER, DAVE/BRODEUR, GREG  C/P:2,P
BLACK SHORE, THE [STAR TREK VOYAGER]  COX, GREG  C/P:2,P
BLESS THE BEASTS [STAR TREK VOYAGER]  HABER, KAREN  P
CARETAKER [STAR TREK VOYAGER]  GRAF, L. A.  C/P,P
CHRYSALIS [@STAR TREK VOYAGER]  WILSON, DAVID NIALL  P
CYBERSONG [STAR TREK VOYAGER]  LEWITT, SHARIANN  P
DEATH OF A NEUTRON STAR [STAR TREK VOYAGER]  KOTANI, ERIC/SMITH, DEAN WESLEY  C/P:2,P
DISTANT SHORES [STAR TREK VOYAGER]  PALMIERI, MARCO  LPA

ECHOES [@STAR TREK VOYAGER] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN/ HOFFMAN, NINA KIRIKI

C/P,P

ENDGAME [STAR TREK VOYAGER]  CAREY, DIANE  LP
EQUINOX [STAR TREK VOYAGER]  CAREY, DIANE  P

ESCAPE, THE [STAR TREK VOYAGER] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

FLASHBACK [STAR TREK VOYAGER]  CAREY, DIANE  C/P:2,P
GARDEN, THE [STAR TREK VOYAGER]  SCOTT, MELISSA  P
GHOST OF A CHANCE [STAR TREK VOYAGER]  GARLAND, MARK A./MCGRAW, CHARLES G. C/P:2,P
INCIDENT AT ARBUK [STAR TREK VOYAGER]  BETANCOURT, JOHN GREGORY  P
MAROONED [STAR TREK VOYAGER]  GOLDEN, CHRISTIE  C/P,P
MOSAIC [@STAR TREK VOYAGER]  TAYLOR, JERI  C/P,P
MURDERED SUN, THE [STAR TREK VOYAGER]  GOLDEN, CHRISTIE  P
NANOTECH WAR, THE [STAR TREK VOYAGER]  PIZIKS, STEVEN  P
PATHWAYS [@STAR TREK VOYAGER]  TAYLOR, JERI  C/P,P
RAGNAROK [STAR TREK VOYAGER]  ARCHER, NATHAN  P
SEVEN OF NINE [STAR TREK VOYAGER]  GOLDEN, CHRISTIE  C/P:2,P
VIOLATIONS [STAR TREK VOYAGER]  WRIGHT, SUSAN  C/P:2,P
CLOAK AND DAGGER [@STAR TREK VOYAGER DARK MATTERS #1]  GOLDEN, CHRISTIE  P
GHOST DANCE [@STAR TREK VOYAGER DARK MATTERS #2]  GOLDEN, CHRISTIE  P
SHADOW OF HEAVEN [@STAR TREK VOYAGER DARK MATTERS #3]  GOLDEN, CHRISTIE  P
FARTHER SHORE, THE [@STAR TREK VOYAGER HOMECOMING 2]  GOLDEN, CHRISTIE  P
HOMECOMING [@STAR TREK VOYAGER HOMECOMING 1]  GOLDEN, CHRISTIE  P

SHADOW [@STAR TREK VOYAGER SECTION 31] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

ENEMY OF MY ENEMY [@STAR TREK VOYAGER SPIRIT WALK 2]  GOLDEN, CHRISTIE  P
OLD WOUNDS [@STAR TREK VOYAGER SPIRIT WALK 1]  GOLDEN, CHRISTIE  P
QUARANTINE [@STAR TREK VOYAGER STARFLEET ACADEMY #3]  BARNES-SVARNEY, PATRICIA  LP
COHESION [STAR TREK VOYAGER STRING THEORY 1]  LANG, JEFFREY  P
FUSION [STAR TREK VOYAGER STRING THEORY 2]  BEYER, KIRSTEN  P
EVOLUTION [@STAR TREK VOYAGER STRING THEORY #3]  JARMAN, HEATHER  P
EXILES [STAR TREK VULCAN’S SOUL 2]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  H
EXODUS [@STAR TREK VULCAN’S SOUL #1]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  P
EPIPHANY [@STAR TREK VULCAN’S SOUL #3]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  H
ALLEGIANCE [@STAR WARS]  ZAHN, TIMOTHY  H:3
APPROACHING STORM, THE [@STAR WARS]  FOSTER, ALAN DEAN  H,P
ART OF STAR WARS, THE [STAR WARS]  TITELMAN, CAROL  A
ART OF STAR WARS: EPISODE V, THE [STAR WARS]  CALL, DEBORAH  A
ART OF STAR WARS: EPISODE VI, THE [STAR WARS]  LUCASFILM LTD.  A
CHILDREN OF THE JEDI [STAR WARS]  HAMBLY, BARBARA  H:2,P
CLOAK OF DECEPTION [@STAR WARS]  LUCENO, JAMES  P
COURTSHIP OF PRINCESS LEIA, THE [STAR WARS]  WOLVERTON, DAVE  H:2,P
CRYSTAL STAR, THE [@STAR WARS]  MCINTYRE, VONDA N.  C/P:2,P
DARK FORCE RISING [STAR WARS]  ZAHN, TIMOTHY  H:2,P
DARK LORD: THE RISE OF DARTH VADER [STAR WARS]  LUCENO, JAMES  H
DARKSABER [@STAR WARS]  ANDERSON, KEVIN J.  H:2,P
DARTH MAUL: SHADOW HUNTER [STAR WARS]  REAVES, MICHAEL  P

EMPIRE STRIKES BACK, THE:

NATIONAL PUBLIC RADIO DRAMATIZATION, THE [STAR WARS] DALEY, BRIAN

LP

FALL OF THE REPUBLIC (STORY TREATMENT) [STAR WARS]  FLYNN, JOHN L.  SR-S
HEIR TO THE EMPIRE [STAR WARS]  ZAHN, TIMOTHY  H:2,P
I, JEDI [@STAR WARS]  STACKPOLE, MICHAEL A.  C/P,P,H
JABBA’S PALACE POP-UP BOOK [STAR WARS]  ANDERSON, KEVIN J./MOESTA, REBECCA  SR-MM
LABYRINTH OF EVIL [@STAR WARS]  LUCENO, JAMES  H
LANDO CALRISSIAN ADVENTURES, THE [STAR WARS]  SMITH, L. NEIL  P
LANDO CALRISSIAN AND THE FLAMEWIND OF OSEON [STAR WARS]  SMITH, L. NEIL  P
LANDO CALRISSIAN AND THE MINDHARP OF SHARU [STAR WARS]  SMITH, L. NEIL  C/P,P
LANDO CALRISSIAN AND THE STARCAVE OF THONBOKA [STAR WARS]  SMITH, L. NEIL  P
LAST COMMAND, THE [STAR WARS]  ZAHN, TIMOTHY  H:2,P
NEW REBELLION, THE [@STAR WARS]  RUSCH, KRISTINE KATHRYN  C/P,P
OUTBOUND FLIGHT [@STAR WARS]  ZAHN, TIMOTHY  H
PLANET OF TWILIGHT [@STAR WARS]  HAMBLY, BARBARA  P
REVENGE OF THE SITH [@STAR WARS]  STOVER, MATTHEW WOODRING  H:2

The MITSFS Seriesdex  September 24, 2007  Page 158
TITLE  AUTHOR  CODE
ROGUE PLANET [@STAR WARS]  BEAR, GREG  P
SHADOWS OF THE EMPIRE [@STAR WARS]  PERRY, STEVE  H:2,P
SPECTER OF THE PAST [STAR WARS]  ZAHN, TIMOTHY  P
SPLINTER OF THE MIND’S EYE [STAR WARS]  FOSTER, ALAN DEAN  H:2,P
STAR WARS [STAR WARS]  LUCAS, GEORGE  C/P,H,SR-P
STAR WARS [STAR WARS]  THOMAS, ROY/CHAYKIN, HOWARD  SCX
STAR WARS (RADIO SHOW) (7 CASSETTES) [STAR WARS]  LUCAS, GEORGE  SR-MM
STAR WARS BLUEPRINTS [STAR WARS]  BALLANTINE BOOKS  A
STAR WARS BOOK ONE [STAR WARS]  DAVIDS, PAUL/DAVIDS, HOLLACE  H
STAR WARS EPISODE 1 JOURNAL: DARTH MAUL [STAR WARS]  WATSIN, JUDE  LP

STAR WARS: NATIONAL PUBLIC RADIO DRAMATIZATION, THE [STAR WARS] DALEY, BRIAN

LP

STAR WARS SKETCHBOOK, THE [STAR WARS]  JOHNSTON, JOE  A

STAR WARS: THE SCRIPTS [STAR WARS] LUCAS, GEORGE/BRACKETT, LEIGH/

KASDAN, LAWRENCE

SR-S

STAR WARS TRILOGY, THE [STAR WARS]  BALLANTINE BOOKS  LP
SURVIVOR’S QUEST [STAR WARS]  ZAHN, TIMOTHY  H
TALES FROM JABBA’S PALACE [STAR WARS]  ANDERSON, KEVIN J.  PA,C/PA
TALES FROM THE EMPIRE [@STAR WARS]  SCHWEIGHOFER, PETER  PA
TALES FROM THE MOS EISLEY CANTINA [STAR WARS]  ANDERSON, KEVIN J.  C/PA,PA
TALES FROM THE NEW REPUBLIC [@STAR WARS]  SCHWEIGHOFER, PETER/CAREY, CRAIG  PA
TALES OF THE BOUNTY HUNTERS [STAR WARS]  ANDERSON, KEVIN J.  PA
TATOOINE GHOST [@STAR WARS]  DENNING, TROY  P
TRUCE AT BAKURA, THE [@STAR WARS]  TYERS, KATHY  C/P,P,H
VECTOR PRIME (CD) [STAR WARS]  LUCAS BOOKS  SR-MM
VISION OF THE FUTURE [STAR WARS]  ZAHN, TIMOTHY  C/P,P
BEFORE THE STORM [@STAR WARS BLACK FLEET CRISIS 1]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
SHIELD OF LIES [STAR WARS BLACK FLEET CRISIS 2]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
TYRANT’S TEST [@STAR WARS BLACK FLEET CRISIS 3]  KUBE-MCDOWELL, MICHAEL P.  C/P,P
BOUNTY HUNTER WARS, THE [STAR WARS BOUNTY HUNTER WARS 1,2,3]  JETER, K. W.  H:2
HARD MERCHANDISE [@STAR WARS BOUNTY HUNTER WARS 3]  JETER, K. W.  C/P,P
MANDALORIAN ARMOR, THE [@STAR WARS BOUNTY HUNTER WARS 1]  JETER, K. W.  C/P,P
SLAVE SHIP [@STAR WARS BOUNTY HUNTER WARS 2]  JETER, K. W.  C/P,P
CESTUS DECEPTION, THE [@STAR WARS CLONE WARS]  BARNES, STEVEN  P
JEDI TRIAL [@STAR WARS CLONE WARS]  SHERMAN, DAVID/CRAGG, DAN  H
SHATTERPOINT [@STAR WARS CLONE WARS]  STOVER, MATTHEW WOODRING  P
BATTLE SURGEONS [@STAR WARS CLONE WARS MEDSTAR 1]  REAVES, MICHAEL/PERRY, STEVE  P
MEDSTAR [@STAR WARS CLONE WARS MEDSTAR 1,2]  REAVES, MICHAEL/PERRY, STEVE  H
DARK RENDEZVOUS [@STAR WARS CLONE WARS YODA]  STEWART, SEAN  P
AMBUSH AT CORELLIA [STAR WARS CORELLIAN 1]  ALLEN, ROGER MACBRIDE  C/P:2,P
ASSAULT AT SELONIA [STAR WARS CORELLIAN 2]  ALLEN, ROGER MACBRIDE  C/P:2,P
SHOWDOWN AT CENTERPOINT [STAR WARS CORELLIAN 3]  ALLEN, ROGER MACBRIDE  C/P:2,P
REBEL AGENT [@STAR WARS DARK FORCES 2]  DIETZ, WILLIAM C.  H
SOLDIER FOR THE EMPIRE [@STAR WARS DARK FORCES 1]  DIETZ, WILLIAM C.  H
DARK NEST [@STAR WARS DARK NEST 1,2,3]  DENNING, TROY  H
PATH OF DESTRUCTION [@STAR WARS DARTH BANE]  KARPYSHYN, DREW  H

RUINS OF DANTOOINE [@STAR WARS GALAXIES] WHITNEY-ROBINSON, VORONICA/

BLACKMAN, HADEN

P

EATEN ALIVE [@STAR WARS GALAXY OF FEAR #1]  WHITMAN, JOHN  P
CITY OF THE DEAD [@STAR WARS GALAXY OF FEAR #2]  WHITMAN, JOHN  P
PLANET PLAGUE [@STAR WARS GALAXY OF FEAR #3]  WHITMAN, JOHN  P
NIGHTMARE MACHINE, THE [@STAR WARS GALAXY OF FEAR #4]  WHITMAN, JOHN  P
GHOST OF THE JEDI [@STAR WARS GALAXY OF FEAR #5]  WHITMAN, JOHN  P
HAN SOLO ADVENTURES, THE [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P
HAN SOLO AND THE LOST LEGACY [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P,H
HAN SOLO AT STARS’ END [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P,H
HAN SOLO’S REVENGE [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P,H
HUTT GAMBIT, THE [@STAR WARS HAN SOLO 2]  CRISPIN, A. C.  P
PARADISE SNARE, THE [@STAR WARS HAN SOLO 1]  CRISPIN, A. C.  C/P,P
REBEL DAWN [@STAR WARS HAN SOLO 3]  CRISPIN, A. C.  P
CHAMPIONS OF THE FORCE [@STAR WARS JEDI ACADEMY 3]  ANDERSON, KEVIN J.  C/P:2,P
DARK APPRENTICE [@STAR WARS JEDI ACADEMY 2]  ANDERSON, KEVIN J.  C/P:2,P
JEDI SEARCH [@STAR WARS JEDI ACADEMY 1]  ANDERSON, KEVIN J.  C/P:2,P
RETURN OF THE DARK SIDE [@STAR WARS LAST OF THE JEDI #6]  WATSON, JUDE  LP
BETRAYAL [STAR WARS LEGACY OF THE FORCE 1]  ALLSTON, AARON  H
BLOODLINES [@STAR WARS LEGACY OF THE FORCE 2]  TRAVISS, KAREN  P
EXILE [@STAR WARS LEGACY OF THE FORCE 4]  ALLSTON, AARON  H,P
SACRIFICE [STAR WARS LEGACY OF THE FORCE 5]  TRAVISS, KAREN  H
TEMPEST [@STAR WARS LEGACY OF THE FORCE 3]  DENNING, TROY  P
ATTACK OF THE CLONES [STAR WARS MOVIE 2]  SALVATORE, R. A.  H
EMPIRE STRIKES BACK, THE [STAR WARS MOVIE 5]  GLUT, DONALD F.  H:2,P
EMPIRE STRIKES BACK THE SCREENPLAY, THE [STAR WARS MOVIE 5]  BRACKETT, LEIGH/KASDAN, LAWRENCE  SR-S
NEW HOPE, A [STAR WARS MOVIE 4]  LUCAS, GEORGE  P

The MITSFS Seriesdex  September 24, 2007  Page 159
TITLE  AUTHOR  CODE
PHANTOM MENACE, THE [STAR WARS MOVIE 1]  BROOKS, TERRY  H:2,P
RETURN OF THE JEDI [STAR WARS MOVIE 6]  KAHN, JAMES  C/P,P,H
BALANCE POINT [@STAR WARS NEW JEDI ORDER]  TYERS, KATHY  P
DARK JOURNEY [@STAR WARS NEW JEDI ORDER]  CUNNINGHAM, ELAINE  P
DESTINY’S WAY [@STAR WARS NEW JEDI ORDER]  WILLIAMS, WALTER JON  P
FINAL PROPHECY, THE [@STAR WARS NEW JEDI ORDER]  KEYES, J. GREGORY  C/P,P
STAR BY STAR [@STAR WARS NEW JEDI ORDER]  DENNING, TROY  P
TRAITOR [@STAR WARS NEW JEDI ORDER]  STOVER, MATTHEW WOODRING  P
UNIFYING FORCE, THE [@STAR WARS NEW JEDI ORDER]  LUCENO, JAMES  P

UNIFYING FORCE, THE (HARDCOVER AND CD)

[@STAR WARS NEW JEDI ORDER] LUCENO, JAMES

H

VECTOR PRIME [@STAR WARS NEW JEDI ORDER]  SALVATORE, R. A.  P

AGENTS OF CHAOS [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS 1,2] LUCENO, JAMES

H

HERO’S TRIAL [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS #1]  LUCENO, JAMES  P
JEDI ECLIPSE [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS #2]  LUCENO, JAMES  P
DARK TIDE [@STAR WARS NEW JEDI ORDER DARK TIDE 1,2]  STACKPOLE, MICHAEL A.  H
ONSLAUGHT [@STAR WARS NEW JEDI ORDER DARK TIDE #1]  STACKPOLE, MICHAEL A.  P
RUIN [@STAR WARS NEW JEDI ORDER DARK TIDE #2]  STACKPOLE, MICHAEL A.  P
EDGE OF VICTORY [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY 1,2]  KEYES, J. GREGORY  H
CONQUEST [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY #1]  KEYES, J. GREGORY  P
REBIRTH [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY #2]  KEYES, J. GREGORY  P
ENEMY LINES [@STAR WARS NEW JEDI ORDER ENEMY LINES 1,2]  ALLSTON, AARON  H
REBEL DREAM [@STAR WARS NEW JEDI ORDER ENEMY LINES 1]  ALLSTON, AARON  P
REBEL STAND [@STAR WARS NEW JEDI ORDER ENEMY LINES 2]  ALLSTON, AARON  P
FORCE HERETIC [@STAR WARS NEW JEDI ORDER FORCE HERETIC 1,2,3]  WILLIAMS, SEAN/DIX, SHANE  H
REMNANT [@STAR WARS NEW JEDI ORDER FORCE HERETIC #1]  WILLIAMS, SEAN/DIX, SHANE  P
REFUGEE [@STAR WARS NEW JEDI ORDER FORCE HERETIC #2]  WILLIAMS, SEAN/DIX, SHANE  P
REUNION [@STAR WARS NEW JEDI ORDER FORCE HERETIC #3]  WILLIAMS, SEAN/DIX, SHANE  P
HARD CONTACT [@STAR WARS REPUBLIC COMMANDO 1]  TRAVISS, KAREN  C/P,P
TRIPLE ZERO [@STAR WARS REPUBLIC COMMANDO 2]  TRAVISS, KAREN  P
BACTRA WAR, THE [@STAR WARS X-WING 4]  STACKPOLE, MICHAEL A.  P
IRON FIST [@STAR WARS X-WING 6]  ALLSTON, AARON  C/P,P
ISARD’S REVENGE [@STAR WARS X-WING 8]  STACKPOLE, MICHAEL A.  C/P,P
KRYTOS TRAP, THE [@STAR WARS X-WING 3]  STACKPOLE, MICHAEL A.  C/P,P
ROGUE SQUADRON [@STAR WARS X-WING 1]  STACKPOLE, MICHAEL A.  C/P,P
SOLO COMMAND [@STAR WARS X-WING 7]  ALLSTON, AARON  P
STARFIGHTERS OF ADUMAR [@STAR WARS X-WING 9]  ALLSTON, AARON  C/P,P
WEDGE’S GAMBLE [@STAR WARS X-WING 2]  STACKPOLE, MICHAEL A.  C/P,P
WRAITH SQUADRON [@STAR WARS X-WING 5]  ALLSTON, AARON  P
CRISIS AT CRYSTAL REEF [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
DARKEST KNIGHT [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
DELUSIONS OF GRANDEUR [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
DIVERSITY ALLIANCE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
EMPEROR’S PLAGUE, THE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
HEIRS OF THE FORCE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
JEDI BOUNTY [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
JEDI UNDER SIEGE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
LIGHTSABERS [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P:2,P
LOST ONES, THE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P:2,P
RETURN TO ORD MANTELL [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
SHADOW ACADEMY [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
SHARDS OF ALDERAAN [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
TROUBLE ON CLOUD CITY [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
MIDDLE OF NOWHERE, THE [STAR WOLF 2]  GERROLD, DAVID  C/P,P
STARHUNT [STAR WOLF 0]  GERROLD, DAVID  P
TALES OF THE STAR WOLF [STAR WOLF 1,2,3]  GERROLD, DAVID  H
VOYAGE OF THE STAR WOLF [STAR WOLF 1]  GERROLD, DAVID  C/P,P,H
YESTERDAY’S CHILDREN [STAR WOLF 0]  GERROLD, DAVID  C/P:2,P
YESTERDAY’S CHILDREN (REVISED, PROOF) [STAR WOLF 0]  GERROLD, DAVID  SR-LP
YESTERDAY’S CHILDREN (REVISED) [STAR WOLF 0]  GERROLD, DAVID  P
GENERATION GAP, THE [@STAR WRECK #1]  REWOLINSKI, LEAH  C/P,P
UNDISCOVERED NURSING HOME, THE [@STAR WRECK #5]  REWOLINSKI, LEAH  P
GEEK SPACE NINE [@STAR WRECK #6]  REWOLINSKI, LEAH  P
SPACE, THE FIDO FRONTIER [@STAR WRECK #7]  REWOLINSKI, LEAH  P
ANCESTOR’S WORLD [STARBRIDGE 6]  CRISPIN, A. C./KING, T. JACKSON  P
VOICES OF CHAOS [STARBRIDGE 7]  CRISPIN, A. C./EMERSON, RU  P
STARBRIDGE [@STARBRIDGE #1]  CRISPIN, A. C.  C/P:2,SFWA-P:2,P
SILENT DANCES [@STARBRIDGE #2]  CRISPIN, A. C./O’MALLEY, KATHLEEN  C/P:2,SFWA-P:2,P
SHADOW WORLD [@STARBRIDGE #3]  CRISPIN, A. C./ELLIOTT, JANNEAN  C/P:2,SFWA-P:2,P
SERPENT’S GIFT [@STARBRIDGE #4]  CRISPIN, A. C./MARSHALL, DEBORAH A.  C/P,SFWA-P:2,P
SILENT SONGS [@STARBRIDGE #5]  CRISPIN, A. C./O’MALLEY, KATHLEEN  P,SFWA-P
CULT OF LOVING KINDNESS, THE [STARBRIDGE CHRONICLES 3]  PARK, PAUL  H,P

TITLE  AUTHOR  CODE
SOLDIERS OF PARADISE [STARBRIDGE CHRONICLES 1]  PARK, PAUL  C/P,SFWA,P,H
SUGAR FESTIVAL, THE [STARBRIDGE CHRONICLES 1,2]  PARK, PAUL  H:2
SUGAR RAIN [STARBRIDGE CHRONICLES 2]  PARK, PAUL  C/P,P,H
QUEEN OF BLADES [@STARCRAFT]  ROSENBERG, AARON  P
LIBERTY’S CRUSADE [@STARCRAFT #1]  GRUBB, JEFF  P
SHADOW OF THE XEL’NAGA [@STARCRAFT #2]  MESTA, GABRIEL  P
SPEED OF DARKNESS [@STARCRAFT #3]  HICKMAN, TRACY  P
FIRSTBORN [@STARCRAFT DARK TEMPLAR #1]  GOLDEN, CHRISTIE  P
NOVA [@STARCRAFT GHOST]  DECANDIDO, KEITH R. A.  P
DIVISION OF THE SPOILS [STARCRUISER SHENANDOAH 2]  GREEN, ROLAND J.  P
PAINFUL FIELD, THE [STARCRUISER SHENANDOAH 5]  GREEN, ROLAND J.  P
SQUADRON ALERT [STARCRUISER SHENANDOAH 1]  GREEN, ROLAND J.  C/P:2,P
WARRIORS FOR THE WORKING DAY [STARCRUISER SHENANDOAH 6]  GREEN, ROLAND J.  P
SUM OF THINGS, THE [@STARCRUISER SHENANDOAH #3]  GREEN, ROLAND J.  P
VAIN COMMAND [@STARCRUISER SHENANDOAH #4]  GREEN, ROLAND J.  P
STAR DANCERS, THE [STARDANCE 1,2]  ROBINSON, SPIDER/ROBINSON, JEANNE  P
STARDANCE [STARDANCE 1]  ROBINSON, SPIDER/ROBINSON, JEANNE  H:2,P
STARDANCE (NOVELLA) [STARDANCE 1]  ROBINSON, SPIDER/ROBINSON, JEANNE  H
STARDANCE (PROOF) [STARDANCE 1]  ROBINSON, SPIDER/ROBINSON, JEANNE  SR-LP
STARMIND [STARDANCE 3]  ROBINSON, SPIDER/ROBINSON, JEANNE  C/P,P,H:2
STARSEED [STARDANCE 2]  ROBINSON, SPIDER/ROBINSON, JEANNE  C/P,SFWA-P:2,P,H
BEYOND VARALLAN [STARDOC 2]  VIEHL, S. L.  C/P,P
ENDURANCE [STARDOC 3]  VIEHL, S. L.  C/P,P
ETERNITY ROW [@STARDOC]  VIEHL, S. L.  C/P,P
PLAGUE OF MEMORY [@STARDOC]  VIEHL, S. L.  P
REBEL ICE [@STARDOC 5]  VIEHL, S. L.  P
SHOCKBALL [STARDOC 4]  VIEHL, S. L.  C/P,P
STARDOC [STARDOC 1]  VIEHL, S. L.  C/P,P
BLADE DANCER [STARDOC JORY RASK]  VIEHL, S. L.  H,P
GRIDRUNNER [STAR*DRIVE]  REID, THOMAS M.  C/P,P
ON THE VERGE [STAR*DRIVE 1]  GREEN, ROLAND J.  P
STARFALL [STAR*DRIVE]  GREENBERG, MARTIN H.  PA
TWO OF MINDS [STAR*DRIVE]  KEITH, WILLIAM H. JR.  P
ZERO POINT [STAR*DRIVE]  BAKER, RICHARD  P
NIGHTFALL AT ALGEMRON [STAR*DRIVE HARBINGER 3]  DUANE, DIANE  P
STARRISE AT CORRIVALE [STAR*DRIVE HARBINGER 1]  DUANE, DIANE  P
STORM AT ELDALA [STAR*DRIVE HARBINGER 2]  DUANE, DIANE  P
METAPHASE [STARFARERS 3]  MCINTYRE, VONDA N.  C/P:2,P
NAUTILUS [STARFARERS 4]  MCINTYRE, VONDA N.  C/P,P
STARFARERS [STARFARERS 1]  MCINTYRE, VONDA N.  C/P:2,SFWA-P,P
TRANSITION [STARFARERS 2]  MCINTYRE, VONDA N.  C/P:2,P
CRUSADE [STARFIRE 2]  WEBER, DAVID/WHITE, STEVE  C/P:2,SFWA-P,P
EXODUS [STARFIRE 5]  WHITE, STEVE/MEIER, SHIRLEY  H
IN DEATH GROUND [STARFIRE 3]  WEBER, DAVID/WHITE, STEVE  C/P,P
INSURRECTION [STARFIRE 1]  WEBER, DAVID/WHITE, STEVE  C/P:2,SFWA-P,P
SHIVA OPTION, THE [STARFIRE 4]  WEBER, DAVID/WHITE, STEVE  C/P,P
PASSAGE AT ARMS [STARFISHERS]  COOK, GLEN  H,P
SHADOWLINE [STARFISHERS 1]  COOK, GLEN  P
STARFISHERS [STARFISHERS 2]  COOK, GLEN  C/P,P
STARS’ END [STARFISHERS 3]  COOK, GLEN  C/P,P
BLOOD CONTACT [@STARFIST 4]  SHERMAN, DAVID/CRAGG, DAN  P
FIRESTORM [@STARFIST 12]  SHERMAN, DAVID/CRAGG, DAN  H
FIRST TO FIGHT [@STARFIST 1]  SHERMAN, DAVID/CRAGG, DAN  C/P:2,P
FLASHFIRE [@STARFIST 11]  SHERMAN, DAVID/CRAGG, DAN  H
HANGFIRE [@STARFIST 6]  SHERMAN, DAVID/CRAGG, DAN  C/P,P
KINGDOM’S FURY [@STARFIST 8]  SHERMAN, DAVID/CRAGG, DAN  P
KINGDOM’S SWORDS [@STARFIST 7]  SHERMAN, DAVID/CRAGG, DAN  C/P,P
LAZARUS RISING [@STARFIST 9]  SHERMAN, DAVID/CRAGG, DAN  H,P
SCHOOL OF FIRE [@STARFIST 2]  SHERMAN, DAVID/CRAGG, DAN  P
STEEL GAUNTLET [@STARFIST 3]  SHERMAN, DAVID/CRAGG, DAN  P
TECHNOKILL [@STARFIST 5]  SHERMAN, DAVID/CRAGG, DAN  P
WORLD OF HURT, A [@STARFIST 10]  SHERMAN, DAVID/CRAGG, DAN  P
WORLD OF HURT, A (PROOF) [@STARFIST 10]  SHERMAN, DAVID/CRAGG, DAN  SR-LP
BACKSHOT [@STARFIST FORCE RECON 1]  SHERMAN, DAVID/CRAGG, DAN  P
POINTBLANK [@STARFIST FORCE RECON 2]  SHERMAN, DAVID/CRAGG, DAN  P
REBELLION [STARGATE]  MCCAY, BILL  C/P:2,P
RECONNAISSANCE [STARGATE]  MCCAY, BILL  P
RESISTANCE [STARGATE]  MCCAY, BILL  C/P,P
RETALIATION [STARGATE]  MCCAY, BILL  C/P,P
RETRIBUTION [STARGATE]  MCCAY, BILL  P
STARGATE [STARGATE]  DEVLIN, DEAN/EMMERICH, ROLAND  P
STARGATE (ADAPTED FOR YOUNGER READERS) [STARGATE]  BLACK, SHEILA  LP
RISING [STARGATE ATLANTIS #1]  MALCOLM, SALLY  P

The MITSFS Seriesdex  September 24, 2007  Page 161
TITLE  AUTHOR  CODE
RELIQUARY [STARGATE ATLANTIS #2]  WELLS, MARTHA  P

CHOSEN, THE [STARGATE ATLANTIS #3] WHITELAW, SONNY/

CHRISTENSEN, ELIZABETH

P

EXOGENESIS [@STARGATE ATLANTIS #5] WHITELAW, SONNY/

CHRISTENSEN, ELIZABETH

P

ENTANGLEMENT [@STARGATE ATLANTIS #6]  WELLS, MARTHA  P
TRIAL BY FIRE [@STARGATE SG-1 #1]  BAUER, SABINE C.  P
SACRIFICE MOON [@STARGATE SG-1 #2]  FORTUNE, JULIE  P
MATTER OF HONOR, A [@STARGATE SG-1 #3]  MALCOLM, SALLY  P
CITY OF THE GODS [@STARGATE SG-1 #4]  WHITELAW, SONNY  P
COST OF HONOR, THE [@STARGATE SG-1 #5]  MALCOLM, SALLY  P
SIREN SONG [@STARGATE SG-1 #6]  SCOTT, HOLLY/DUNCAN, JAMES  P
SURVIVAL OF THE FITTEST [@STARGATE SG-1 #7]  BAUER, SABINE C.  P
ALLIANCES [@STARGATE SG-1 #8]  MILLER, KAREN  P
ROSWELL [@STARGATE SG-1 #9]  WHITELAW, SONNY/FALLON, JENNIFER  P
FIRST AMENDMENT, THE [@STARGATE SG-1 (MCCONNELL) 3]  MCCONNELL, ASHLEY  P
MORPHEUS FACTOR, THE [@STARGATE SG-1 (MCCONNELL) 4]  MCCONNELL, ASHLEY  P
PRICE YOU PAY, THE [@STARGATE SG-1 (MCCONNELL) 2]  MCCONNELL, ASHLEY  P
STARGATE SG-1 [@STARGATE SG-1 (MCCONNELL) 1]  MCCONNELL, ASHLEY  C/P,P
STARGAZER (LARGE PAPERBACK AND CASSETTE) [STARGAZER 1]  HAUSMAN, GERALD  SR-MM
DRAGON CIRCLE, THE [@STARGODS #2]  RADFORD, IRENE  P
BATTLE AT ZERO POINT [@STARHAWK 4]  MALONEY, MACK  P
FOURTH EMPIRE, THE [@STARHAWK 3]  MALONEY, MACK  P
PLANET AMERICA [@STARHAWK 2]  MALONEY, MACK  P
STARHAWK [@STARHAWK 1]  MALONEY, MACK  P
STORM OVER SATURN [@STARHAWK 5]  MALONEY, MACK  P
SPACE HAWKS [@STARHAWKS #1]  DALTON, SEAN  C/P:2,P
CODE NAME PEREGRINE [@STARHAWKS #2]  DALTON, SEAN  C/P,P
BEYOND THE VOID [@STARHAWKS #3]  DALTON, SEAN  C/P,P
ROSTMA LURE, THE [@STARHAWKS #4]  DALTON, SEAN  C/P,P
DESTINATION: MUTINY [@STARHAWKS #5]  DALTON, SEAN  C/P:2,P
SALUKAN GAMBIT [@STARHAWKS #6]  DALTON, SEAN  P
GALACTIC WARRIORS [STARHOUNDS 2]  BISCHOFF, DAVID F.  C/P:2,P
INFINITE BATTLE, THE [STARHOUNDS 1]  BISCHOFF, DAVID F.  C/P:2,P
MACROCOSMIC CONFLICT, THE [STARHOUNDS 3]  BISCHOFF, DAVID F.  C/P,P
BLUEBLOODS [STARHUNTERS 3]  DRAKE, DAVID  C/PA,PA,SFWA-P
MEN HUNTING THINGS [STARHUNTERS 1]  DRAKE, DAVID  SFWA-P,C/PA,PA
THINGS HUNTING MEN [STARHUNTERS 2]  DRAKE, DAVID  PA,SFWA-P
FALLING STARS [STARS 4]  FLYNN, MICHAEL  H
FIRESTAR [STARS 1]  FLYNN, MICHAEL  H:2,P
LODESTAR [STARS 3]  FLYNN, MICHAEL  H,P
ROGUE STAR [STARS 2]  FLYNN, MICHAEL  P
STARS & STRIPES FOREVER [STARS & STRIPES 1]  HARRISON, HARRY  C/P,P,H
STARS & STRIPES IN PERIL [STARS & STRIPES 2]  HARRISON, HARRY  H,P
STARS & STRIPES TRIUMPHANT [STARS & STRIPES 3]  HARRISON, HARRY  P
FIRST ACTION [@STARSEA INVADERS 1]  STINE, G. HARRY  C/P,P
SECOND CONTACT [@STARSEA INVADERS 2]  STINE, G. HARRY  C/P,P
THIRD ENCOUNTER [@STARSEA INVADERS 3]  STINE, G. HARRY  P
MANTLE OF KENDIS-DAI, THE [@STARSHIELD 1]  WEIS, MARGARET/HICKMAN, TRACY  P
NIGHTSWORD [@STARSHIELD 3]  WEIS, MARGARET/HICKMAN, TRACY  H
SENTINELS [@STARSHIELD 2]  WEIS, MARGARET/HICKMAN, TRACY  H:2
STARSHIP:PIRATE [STARSHIP 2]  RESNICK, MIKE  H
RETURN FROM THE DEAD [@STARSHIP ORPHEUS #1]  JADE, SYMON  P
COSMIC CARNAGE [@STARSHIP ORPHEUS #2]  JADE, SYMON  P
ALTER EVIL [@STARSHIP ORPHEUS #3]  JADE, SYMON  P
STARSHIP: MUTINY [STARSHIP THEODORE ROOSEVELT 1]  RESNICK, MIKE  H
COMPANY OF STARS, A [@STARSHIP TROUPERS #1]  STASHEFF, CHRISTOPHER  C/P,P,H
WE OPEN ON VENUS [@STARSHIP TROUPERS #2]  STASHEFF, CHRISTOPHER  C/P:2,P
SLIGHT DETOUR, A [@STARSHIP TROUPERS #3]  STASHEFF, CHRISTOPHER  C/P:2,P
TASK FORCE MARS [@STARSTRIKE]  DOCKERY, KEVIN/NILES, DOUGLAS  P
STARWOLF [STARWOLF 1,2,3]  HAMILTON, EDMOND  C/P:2,P
WEAPON FROM BEYOND, THE [@STARWOLF #1]  HAMILTON, EDMOND  C/P:2,P
CLOSED WORLDS, THE [@STARWOLF #2]  HAMILTON, EDMOND  C/P,P
WORLD OF THE STARWOLVES [@STARWOLF #3]  HAMILTON, EDMOND  C/P,P
BATTLE OF THE RING [STARWOLVES 2]  GUNNARSSON, THORARINN  P,SFWA-P
STARWOLVES, THE [STARWOLVES 1]  GUNNARSSON, THORARINN  C/P,P
STARWOLVES DREADNOUGHT [STARWOLVES 4]  GUNNARSSON, THORARINN  P
TACTICAL ERROR [STARWOLVES 3]  GUNNARSSON, THORARINN  P
JAGGED STEELE [STEELE 4]  MASTERS, J. D.  P
STEELE [STEELE 1]  MASTERS, J. D.  P
COMING OF STEELEYE, THE [STEELEYE 1]  DUNN, SAUL  SR-P
STEELEYE: WATERSPACE [STEELEYE]  DUNN, SAUL  SR-P
OUTSKIRTER’S SECRET, THE [STEERSWOMAN 2]  KIRSTEIN, ROSEMARY  C/P,P

TITLE  AUTHOR  CODE
STEERSWOMAN, THE [STEERSWOMAN 1]  KIRSTEIN, ROSEMARY  C/P:2,P
LONE STAR [STELLAR RANGER 2]  PERRY, STEVE  P
STELLAR RANGER [STELLAR RANGER 1]  PERRY, STEVE  C/P,P
COURT OF A THOUSAND SUNS, THE [STEN 3]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
EMPIRE’S END [STEN 8]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
FLEET OF THE DAMNED [STEN 4]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
RETURN OF THE EMPEROR, THE [STEN 6]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
REVENGE OF THE DAMNED [STEN 5]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
STEN [STEN 1]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
VORTEX [STEN 7]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
WOLF WORLDS, THE [STEN 2]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
CHASE THE MORNING [STEPHEN FISHER 1]  ROHAN, MICHAEL SCOTT  P,SFWA
CLOUD CASTLES [STEPHEN FISHER 3]  ROHAN, MICHAEL SCOTT  P
GATES OF NOON, THE [STEPHEN FISHER 2]  ROHAN, MICHAEL SCOTT  C/P,SFWA,P
STINKER FROM SPACE [STINKER 1]  SERVICE, PAMELA F.  P
STINKER’S RETURN [STINKER 2]  SERVICE, PAMELA F.  SFWA
STONE & SEA [STONE 2]  EDWARDS, GRAHAM  P
STONE & SKY [STONE 1]  EDWARDS, GRAHAM  P
STONE & SUN [STONE 3]  EDWARDS, GRAHAM  P
CHOSEN, THE [STONE DANCE OF THE CHAMELEON 1]  PINTO, RICARDO  C/P,P,H
STANDING DEAD, THE [STONE DANCE OF THE CHAMELEON 2]  PINTO, RICARDO  H
DRAGON’S EYE [STONEFORT,MAINE 1]  HETLEY, JAMES A.  LP
DRAGON’S TEETH [STONEFORT,MAINE 2]  HETLEY, JAMES A.  LP
BLOODSTONE [STONES OF POWER 5]  GEMMELL, DAVID  C/P,P
GHOST KING [STONES OF POWER 1]  GEMMELL, DAVID  C/P,P
JERUSALEM MAN, THE [STONES OF POWER 3]  GEMMELL, DAVID  P,SFWA-P
LAST SWORD OF POWER [@STONES OF POWER #2]  GEMMELL, DAVID  P
WOLF IN SHADOW [@STONES OF POWER #3]  GEMMELL, DAVID  P
LAST GUARDIAN, THE [@STONES OF POWER #4]  GEMMELL, DAVID  P
BLACK SHIELDS, THE [STORMLANDS 2]  ROBERTS, JOHN MADDOX  C/P:2,P
ISLANDER, THE [STORMLANDS 1]  ROBERTS, JOHN MADDOX  C/P:2,SFWA-P,P
POISONED LANDS, THE [STORMLANDS 3]  ROBERTS, JOHN MADDOX  C/P,P
QUEENS OF LAND AND SEA [STORMLANDS 5]  ROBERTS, JOHN MADDOX  P
STEEL KINGS, THE [STORMLANDS 4]  ROBERTS, JOHN MADDOX  P
UNICORN DILEMMA, THE [STRAND 2]  LEE, JOHN  P
UNICORN PEACE, THE [STRAND 4]  LEE, JOHN  P
UNICORN QUEST, THE [STRAND 1]  LEE, JOHN  C/P,P
UNICORN SOLUTION, THE [STRAND 3]  LEE, JOHN  C/P:2,SFWA-P,P
UNICORN WAR, THE [STRAND 5]  LEE, JOHN  P
MAZE OF MOONLIGHT [STRANDS OF STARLIGHT 2]  BAUDINO, GAEL  C/P,P
SHROUD OF SHADOW [STRANDS OF STARLIGHT 3]  BAUDINO, GAEL  C/P:2,P
SPIRES OF SPIRIT [STRANDS OF STARLIGHT]  BAUDINO, GAEL  P
STRANDS OF STARLIGHT [STRANDS OF STARLIGHT 1]  BAUDINO, GAEL  C/P:2,P
STRANDS OF SUNLIGHT [STRANDS OF STARLIGHT 4]  BAUDINO, GAEL  C/P,P
BAD TIMING [STRONTIUM DOG]  LEVENE, REBECCA  P
AFFINITY TRAP, THE [STRUCTURE 1]  SKETCHLEY, MARTIN  LP
DESTINY MASK, THE [STRUCTURE 2]  SKETCHLEY, MARTIN  LP
KILLING OF WORLDS, THE [SUCCESSION 2]  WESTERFELD, SCOTT  P
RISEN EMPIRE, THE [SUCCESSION 1]  WESTERFELD, SCOTT  P
SUCCESSION [SUCCESSION 1,2]  WESTERFELD, SCOTT  H
SEASON OF SHADOWS [@SUMMERLANDS #1]  FOXXE, ELLEN  P
SEASON OF STORMS [@SUMMERLANDS #2]  FOXXE, ELLEN  P
MAGICIANS OF NIGHT, THE [SUN-CROSS 2]  HAMBLY, BARBARA  C/P:2,P
RAINBOW ABYSS, THE [SUN-CROSS 1]  HAMBLY, BARBARA  C/P:2,P
SUN-CROSS [SUN-CROSS 1,2]  HAMBLY, BARBARA  H:2
BROKEN CROWN, THE [@SUN SWORD #1]  WEST, MICHELLE  P
UNCROWNED KING, THE [@SUN SWORD #2]  WEST, MICHELLE  P
SHINING COURT, THE [@SUN SWORD #3]  WEST, MICHELLE  C/P,P
SEA OF SORROWS [@SUN SWORD #4]  WEST, MICHELLE  P
RIVEN SHIELD, THE [@SUN SWORD #5]  WEST, MICHELLE  P
SUN SWORD, THE [@SUN SWORD #6]  WEST, MICHELLE  C/P,P
CHAINS OF DARKNESS, CHAINS OF LIGHT [SUNDERED 4]  SAGARA, MICHELLE  P
CHAINS OF DARKNESS, CHAINS OF LIGHT [SUNDERED 4]  WEST, MICHELLE  LP
CHILDREN OF THE BLOOD [SUNDERED 2]  SAGARA, MICHELLE  P,LP
INTO THE DARK LANDS [SUNDERED 1]  SAGARA, MICHELLE  P,LP
LADY OF MERCY [SUNDERED 3]  SAGARA, MICHELLE  P
LADY OF MERCY [SUNDERED 3]  WEST, MICHELLE  LP
BANEWREAKER [@SUNDERING #1]  CAREY, JACQUELINE  H
GODSLAYER [@SUNDERING #2]  CAREY, JACQUELINE  H
BENEATH AN OPAL MOON [SUNSET WARRIOR 4]  LUSTBADER, ERIC VAN  H,SFWA,P
DAI-SAN [SUNSET WARRIOR 3]  LUSTBADER, ERIC VAN  C/P,SFWA,P,H
SHALLOWS OF NIGHT [SUNSET WARRIOR 2]  LUSTBADER, ERIC VAN  C/P,SFWA,P,H
SUNSET WARRIOR, THE [SUNSET WARRIOR 1]  LUSTBADER, ERIC VAN  C/P:2,SFWA,P

The MITSFS Seriesdex  September 24, 2007  Page 163
TITLE  AUTHOR  CODE
STARS WILL SPEAK, THE [SUNSPACE 2]  ZEBROWSKI, GEORGE  H
SUNSPACER [SUNSPACE 1]  ZEBROWSKI, GEORGE  LP
DARK HAND OF MAGIC, THE [SUNWOLF AND STARHAWK 3]  HAMBLY, BARBARA  C/P,P,H
LADIES OF MANDRIGYN, THE [SUNWOLF AND STARHAWK 1]  HAMBLY, BARBARA  C/P:3,P
UNSCHOOLED WIZARD, THE [SUNWOLF AND STARHAWK 1,2]  HAMBLY, BARBARA  H:2
WITCHES OF WENSHAR, THE [SUNWOLF AND STARHAWK 2]  HAMBLY, BARBARA  C/P:2,P
PROTECTOR OF THE REALM [@SUPREME CONSTELLATIONS #1]  BROOKE, GUN  LP
REBEL’S QUEST [@SUPREME CONSTELLATIONS #2]  BROOKE, GUN  LP
BLOOD QUEST [SURVIVAL 2000 1]  MCPHEE, JAMES  P
FROZEN FIRE [@SURVIVAL 2000 3]  MCPHEE, JAMES  P
MID-WAKE [@SURVIVALIST]  AHERN, JERRY  P
TOTAL WAR, THE [@SURVIVALIST #1]  AHERN, JERRY  P
NIGHTMARE BEGINS, THE [@SURVIVALIST #2]  AHERN, JERRY  P
QUEST, THE [@SURVIVALIST #3]  AHERN, JERRY  P
DOOMSAYER, THE [@SURVIVALIST #4]  AHERN, JERRY  P
WEB, THE [@SURVIVALIST #5]  AHERN, JERRY  P
SAVAGE HORDE, THE [@SURVIVALIST #6]  AHERN, JERRY  P
PROPHET, THE [@SURVIVALIST #7]  AHERN, JERRY  P
END IS COMING, THE [@SURVIVALIST #8]  AHERN, JERRY  P
EARTH FIRE [@SURVIVALIST #9]  AHERN, JERRY  P
AWAKENING, THE [@SURVIVALIST #10]  AHERN, JERRY  P
REPRISAL, THE [@SURVIVALIST #11]  AHERN, JERRY  P
REBELLION, THE [@SURVIVALIST #12]  AHERN, JERRY  P
PURSUIT [@SURVIVALIST #13]  AHERN, JERRY  P
TERROR, THE [@SURVIVALIST #14]  AHERN, JERRY  P
OVERLORD [@SURVIVALIST #15]  AHERN, JERRY  P
ARSENAL, THE [@SURVIVALIST #16]  AHERN, JERRY  P
ORDEAL, THE [@SURVIVALIST #17]  AHERN, JERRY  P
SHADOW ROADS, THE [SWANS’ WAR 3]  RUSSELL, SEAN  H
ONE KINGDOM, THE [@SWANS’ WAR #1]  RUSSELL, SEAN  H
ISLE OF BATTLE, THE [@SWANS’ WAR #2]  RUSSELL, SEAN  H
FORBIDDEN [SWORD & SORCERY DEAD GOD 3]  BYERS, RICHARD LEE  P
FORSAKEN [SWORD & SORCERY DEAD GOD 1]  BYERS, RICHARD LEE  P
FORSWORN [SWORD & SORCERY DEAD GOD 2]  BYERS, RICHARD LEE  P
HEIR OF RENGARTH [SWORD AND CIRCLET 2]  DOUGLAS, CAROLE NELSON  H:2,SFWA-P,P
KEEPERS OF EDANVANT [SWORD AND CIRCLET 1]  DOUGLAS, CAROLE NELSON  H:2,SFWA-P,P
SEVEN OF SWORDS [SWORD AND CIRCLET 3]  DOUGLAS, CAROLE NELSON  C/P,SFWA-P,P,H
CLOUDBEARER’S SHADOW [SWORD IN EXILE 1]  MARSTON, ANN  P
KING OF SHADOWS [SWORD IN EXILE 2]  MARSTON, ANN  P
SWORD AND SHADOW [SWORD IN EXILE 3]  MARSTON, ANN  P
DEVLIN’S HONOR [SWORD OF CHANGE 2]  BRAY, PATRICIA  C/P,P
DEVLIN’S HONOR (PROOF) [SWORD OF CHANGE 2]  BRAY, PATRICIA  SR-P
DEVLIN’S JUSTICE [SWORD OF CHANGE 3]  BRAY, PATRICIA  C/P,P
DEVLIN’S LUCK [SWORD OF CHANGE 1]  BRAY, PATRICIA  C/P,P
EARTHLY CROWN, AN [@SWORD OF HEAVEN #1]  ELLIOTT, KATE  C/P:2,P
HIS CONQUERING SWORD [@SWORD OF HEAVEN #2]  ELLIOTT, KATE  C/P:2,SFWA-P,P
REAP THE WHIRLWIND [SWORD OF KNOWLEDGE 3]  CHERRYH, C. J./LACKEY, MERCEDES  C/P:2,SFWA-P:2,P

SWORD OF KNOWLEDGE, THE [SWORD OF KNOWLEDGE 1,2,3] CHERRYH, C. J./LACKEY, MERCEDES/

ASIRE, NANCY/FISH, LESLIE

LP

WIZARD SPAWN [SWORD OF KNOWLEDGE 2]  CHERRYH, C. J./ASIRE, NANCY  C/P:2,SFWA-P:2,P
DIRGE FOR SABIS, A [SWORD OF KNOWLEDGE #1]  CHERRYH, C. J./FISH, LESLIE  C/P:2,SFWA-P,P
CAVERN OF BLACK ICE, A [SWORD OF SHADOWS 1]  JONES, J. V.  H,P
FORTRESS OF GREY ICE, A [SWORD OF SHADOWS 2]  JONES, J. V.  H
BLOOD OF THE FOLD [SWORD OF TRUTH 3]  GOODKIND, TERRY  C/P:2,P,H
CHAINFIRE [SWORD OF TRUTH 9]  GOODKIND, TERRY  H
DEBT OF BONES [SWORD OF TRUTH 10]  GOODKIND, TERRY  H
FAITH OF THE FALLEN [SWORD OF TRUTH 6]  GOODKIND, TERRY  C/P,P,H
NAKED EMPIRE [SWORD OF TRUTH 8]  GOODKIND, TERRY  H
PHANTOM [SWORD OF TRUTH 11]  GOODKIND, TERRY  H
PILLARS OF CREATION, THE [SWORD OF TRUTH 7]  GOODKIND, TERRY  H:2
SOUL OF THE FIRE [SWORD OF TRUTH 5]  GOODKIND, TERRY  H:2,P
STONE OF TEARS [SWORD OF TRUTH 2]  GOODKIND, TERRY  C/P:2,P
TEMPLE OF THE WINDS [SWORD OF TRUTH 4]  GOODKIND, TERRY  H:2,P
WIZARD’S FIRST RULE [SWORD OF TRUTH 1]  GOODKIND, TERRY  C/P:2,SFWA-P,P
CHALICE, THE [@SWORD,RING AND CHALICE 3]  CHESTER, DEBORAH  P
RING, THE [@SWORD,RING AND CHALICE 2]  CHESTER, DEBORAH  P
SWORD, THE [@SWORD,RING AND CHALICE 1]  CHESTER, DEBORAH  P
KING BETRAYED, THE [SWORD,RING AND CHALICE PHERESA 3]  CHESTER, DEBORAH  P
KING IMPERILLED, THE [SWORD,RING AND CHALICE PHERESA 4]  CHESTER, DEBORAH  P
QUEEN’S GAMBIT, THE [SWORD,RING AND CHALICE PHERESA 1]  CHESTER, DEBORAH  C/P,P
QUEEN’S KNIGHT, THE [SWORD,RING AND CHALICE PHERESA 2]  CHESTER, DEBORAH  P
ARMORY OF SWORDS, AN [SWORDS]  SABERHAGEN, FRED  PA
SWORDS TRILOGY, THE [SWORDS 1,2,3]  MOORCOCK, MICHAEL  C/P,P,H

TITLE  AUTHOR  CODE
FIRST BOOK OF SWORDS, THE [@SWORDS #1]  SABERHAGEN, FRED  C/P,LP,H
COMPLETE BOOK OF SWORDS, THE [@SWORDS #1,2,3]  SABERHAGEN, FRED  H:2
SECOND BOOK OF SWORDS, THE [@SWORDS #2]  SABERHAGEN, FRED  C/P,P,LP,H
THIRD BOOK OF SWORDS, THE [@SWORDS #3]  SABERHAGEN, FRED  C/P,SFWA-P,P,H
BEASTS OF THE MIST, THE [SWORDS OF RAEMLLYN 5]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P:2,P
BLOOD FOUNTAIN [SWORDS OF RAEMLLYN 3]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P:2,P
DEATH’S ACOLYTE [SWORDS OF RAEMLLYN 4]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  P
FOR CROWN AND KINGDOM [SWORDS OF RAEMLLYN 6]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  P
TO DEMONS BOUND [SWORDS OF RAEMLLYN 1]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P,P
YOKE OF MAGIC, A [SWORDS OF RAEMLLYN 2]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P,P
FALL OF THE KINGS, THE [SWORDSPOINT 2]  KUSHNER, ELLEN/SHERMAN, DELIA  C/P,LP
PRIVILEGE OF THE SWORD, THE [SWORDSPOINT]  KUSHNER, ELLEN  H,LP
SWORDSPOINT [SWORDSPOINT 1]  KUSHNER, ELLEN  C/P,SFWA-P,P,H
DEVIL & DAN COOLEY, THE [SYMPATHY FOR THE DEVIL 2]  LISLE, HOLLY  C/P,P
HELL ON HIGH [SYMPATHY FOR THE DEVIL 3]  LISLE, HOLLY/NOLAN, TED  P
SYMPATHY FOR THE DEVIL [SYMPATHY FOR THE DEVIL 1]  LISLE, HOLLY  C/P,P
DESTINY [SYMPHONY OF AGES 3]  HAYDON, ELIZABETH  H,P
ELEGY FOR A LOST STAR [SYMPHONY OF AGES 5]  HAYDON, ELIZABETH  H,P
PROPHECY [SYMPHONY OF AGES 2]  HAYDON, ELIZABETH  H,P
REQUIEM FOR THE SUN [SYMPHONY OF AGES 4]  HAYDON, ELIZABETH  H
REQUIEM FOR THE SUN (PROOF) [SYMPHONY OF AGES 4]  HAYDON, ELIZABETH  SR-LP
RHAPSODY [SYMPHONY OF AGES 1]  HAYDON, ELIZABETH  H,P
SINGLE COMBAT [SYSTEMIC SHOCK 2]  ING, DEAN  C/P:2,P
SYSTEMIC SHOCK [SYSTEMIC SHOCK 1]  ING, DEAN  C/P:2,P
WILD COUNTRY [SYSTEMIC SHOCK 3]  ING, DEAN  C/P,SFWA-P,P

TITLE  AUTHOR  CODE
LOST AND FOUND [TAKEN 1]  FOSTER, ALAN DEAN  H
ALTERED CARBON [TAKESHI KOVACS 1]  MORGAN, RICHARD  H
BROKEN ANGELS [TAKESHI KOVACS 2]  MORGAN, RICHARD  H,LP
WOKEN FURIES [TAKESHI KOVACS 3]  MORGAN, RICHARD  H
DOOR INTO FIRE, THE [TALE OF THE FIVE 1]  DUANE, DIANE  P
DOOR INTO FIRE, THE (REVISED) [TALE OF THE FIVE 1]  DUANE, DIANE  C/P,SFWA-P,LP
DOOR INTO SHADOW, THE [TALE OF THE FIVE 2]  DUANE, DIANE  H:2,SFWA-P,LP
DOOR INTO SUNSET, THE [TALE OF THE FIVE 3]  DUANE, DIANE  SR-P,H
VULGMASTER, THE [@TALES FROM THE ONE-EYED CROW]  MCKIERNAN, DENNIS L.  C/P:2,SCX,SFWA-P
RIDDLE AND THE RUNE, THE [TALES OF GOM]  CHETWIN, GRACE  P
STARSTONE, THE [TALES OF GOM]  CHETWIN, GRACE  P
DRAGON QUEEN, THE [TALES OF GUINEVERE 1]  BORCHARDT, ALICE  H,P
RAVEN WARRIOR, THE [TALES OF GUINEVERE 2]  BORCHARDT, ALICE  H
CULAI HERITAGE, THE [TALES OF THE BARD 1,2,3]  SCOTT, MICHAEL  LP
DEATH’S LAW [TALES OF THE BARD #3]  SCOTT, MICHAEL  LP
IN THE TOMB OF THE BISHOP [@TALES OF THE WEREWOLF CLAN #1]  MUNN, H. WARNER  SR-H
MASTER GOES HOME, THE [@TALES OF THE WEREWOLF CLAN #2]  MUNN, H. WARNER  SR-H
ICE FROM SPACE [TALES OF TOMORROW 4]  HOUSTON, DAVID  P
INVADERS AT GROUND ZERO [TALES OF TOMORROW 1]  HOUSTON, DAVID  P
RED DUST [TALES OF TOMORROW 2]  HOUSTON, DAVID  P
SUBSTANCE X [TALES OF TOMORROW 3]  HOUSTON, DAVID  P
CUP OF CLAY [TALISWOMAN 1]  DOUGLAS, CAROLE NELSON  H:2,P
SEED UPON THE WIND [TALISWOMAN 2]  DOUGLAS, CAROLE NELSON  H:2
MOONHEART [TAMSON HOUSE 1]  DE LINT, CHARLES  C/P:2,P
SPIRITWALK [TAMSON HOUSE 2]  DE LINT, CHARLES  C/P,P,H
DOMES OF FIRE [TAMULI 1]  EDDINGS, DAVID  H:2,LP
HIDDEN CITY, THE [TAMULI 3]  EDDINGS, DAVID  H:2
SHINING ONES, THE [TAMULI 2]  EDDINGS, DAVID  H:2,P
SCHEMES OF DRAGONS, THE [TANAGARAN 2]  SMEDS, DAVE  P
SORCERY WITHIN, THE [TANAGARAN 1]  SMEDS, DAVE  C/P,P
BLACK UNICORN [TANAQUIL 1]  LEE, TANITH  H:2,SFWA,P
GOLD UNICORN [TANAQUIL 2]  LEE, TANITH  C/P,P,H
RED UNICORN [TANAQUIL 3]  LEE, TANITH  H,P
FIRE GET [TAORMIN 1]  FRANKLIN, CHERYL J.  C/P:2,P
FIRE LORD [TAORMIN 2]  FRANKLIN, CHERYL J.  C/P:2,P
TAROT [TAROT 1,2,3]  ANTHONY, PIERS  LP
GOD OF TAROT [TAROT #1]  ANTHONY, PIERS  C/P:2,@P
VISION OF TAROT [TAROT #2]  ANTHONY, PIERS  C/P:2,@P
FAITH OF TAROT [TAROT #3]  ANTHONY, PIERS  C/P:2,@P
FANTASY SHOWCASE TAROT, THE [TAROT DECK]  PELZ, BRUCE  SR-PA
DARK HEART OF TIME, THE [@TARZAN]  FARMER, PHILIP JOSE  P
TARZAN AND THE SILVER GLOBE [TARZAN]  WERPER, BARTON  SR-P
TARZAN AND THE SNAKE PEOPLE [TARZAN]  WERPER, BARTON  SR-P
TARZAN AND THE VALLEY OF GOLD [TARZAN]  LEIBER, FRITZ  P
TARZAN AND THE WELL OF SLAVES [TARZAN]  NILES, DOUGLAS  P
TARZAN: THE EPIC ADVENTURES [TARZAN]  SALVATORE, R. A.  P
TARZAN TWINS, THE [TARZAN]  BURROUGHS, EDGAR RICE  H
TARZAN OF THE APES [TARZAN #1]  BURROUGHS, EDGAR RICE  H:2,@P
RETURN OF TARZAN, THE [TARZAN #2]  BURROUGHS, EDGAR RICE  H:2,@P
BEASTS OF TARZAN, THE [TARZAN #3]  BURROUGHS, EDGAR RICE  C/P,@P,H
SON OF TARZAN, THE [TARZAN #4]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN AND THE JEWELS OF OPAR [TARZAN #5]  BURROUGHS, EDGAR RICE  H:2,@P
JUNGLE TALES OF TARZAN [TARZAN #6]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN THE UNTAMED [TARZAN #7]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN THE TERRIBLE [TARZAN #8]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN AND THE GOLDEN LION [TARZAN #9]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN AND THE ANT-MEN [TARZAN #10]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN, LORD OF THE JUNGLE [TARZAN #11]  BURROUGHS, EDGAR RICE  C/P,@P,H:2
TARZAN AND THE LOST EMPIRE [TARZAN #12]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN AT THE EARTH’S CORE [TARZAN #13/PELLUCIDAR 4]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN THE INVINCIBLE [TARZAN #14]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN TRIUMPHANT [TARZAN #15]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN AND THE CITY OF GOLD [TARZAN #16]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN AND THE LION MAN [TARZAN #17]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN AND THE LEOPARD MEN [TARZAN #18]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN’S QUEST [TARZAN #19]  BURROUGHS, EDGAR RICE  @P
TARZAN AND THE FORBIDDEN CITY [TARZAN #20]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN THE MAGNIFICENT [TARZAN #21]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN AND "THE FOREIGN LEGION" [TARZAN #22]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN AND THE MADMAN [TARZAN #23]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN AND THE CASTAWAYS [TARZAN #24]  BURROUGHS, EDGAR RICE  C/P:2,@P
JOURNEY INTO THE GATEWAY OF DIMENSIONS [TAVAN 1]  BURTON, JOSH MICHAEL  LP
FOOL’S ERRAND [TAWNY MAN 1]  HOBB, ROBIN  C/P,P

The MITSFS Seriesdex  September 24, 2007  Page 166
TITLE  AUTHOR  CODE
FOOL’S FATE [TAWNY MAN 3]  HOBB, ROBIN  H
GOLDEN FOOL [TAWNY MAN 2]  HOBB, ROBIN  H,P
LADY AND THE TIGER, THE [TAYLOR’S ARK 3]  NYE, JODY LYNN  C/P,P
MEDICINE SHOW [TAYLOR’S ARK 2]  NYE, JODY LYNN  P
TAYLOR’S ARK [TAYLOR’S ARK 1]  NYE, JODY LYNN  P
CHILDREN OF THE SERPENT GATE [TEARS OF ARTAMON 3]  ASH, SARAH  H
LORD OF SNOW AND SHADOWS [TEARS OF ARTAMON 1]  ASH, SARAH  C/P,P
LORD OF SNOW AND SHADOWS (PROOF) [TEARS OF ARTAMON 1]  ASH, SARAH  SR-LP
PRISONER OF THE IRON TOWER [TEARS OF ARTAMON 2]  ASH, SARAH  H
ALL THINGS UNDER THE MOON [TEDDY LONDON,P.I.]  MORGAN, ROBERT  P
WISDOM TEETH [TEETH #1,2,3]  HUCKELBERY, TIMOTHY LEE  WRD
TEK KILL [TEK 8]  SHATNER, WILLIAM  P
TEK MONEY [TEK 7]  SHATNER, WILLIAM  H,P
TEK NET [TEK 9]  SHATNER, WILLIAM  H,P
TEK POWER [TEK 6]  SHATNER, WILLIAM  P
TEK SECRET [TEK 5]  SHATNER, WILLIAM  P
TEK VENGEANCE [TEK 4]  SHATNER, WILLIAM  H,SFWA-P:2,P
TEKLAB [TEK 3]  SHATNER, WILLIAM  C/P,SFWA-P,P
TEKLORDS [TEK 2]  SHATNER, WILLIAM  C/P:2,SFWA-P:2,P
TEKWAR [TEK 1]  SHATNER, WILLIAM  C/P,SFWA-P:2,P,H
FLAMESONG [TEKUMEL 2]  BARKER, M. A. R.  C/P,P
MAN OF GOLD, THE [TEKUMEL 1]  BARKER, M. A. R.  C/P,P
CAPTAIN, THE [@TELNARIAN 2]  NORMAN, JOHN  C/P:2,P
CHIEFTAIN, THE [@TELNARIAN 1]  NORMAN, JOHN  C/P,P
KING, THE [@TELNARIAN 3]  NORMAN, JOHN  P

HUB: DANGEROUS TERRITORY, THE [TELZEY] SCHMITZ, JAMES H./FLINT, ERIC/

GORDON, GUY

C/P,P

LION GAME, THE [TELZEY 3]  SCHMITZ, JAMES H.  C/P:2,P
TELZEY AMBERDON [TELZEY]  SCHMITZ, JAMES H./FLINT, ERIC  P
TELZEY TOY, THE [TELZEY 2]  SCHMITZ, JAMES H.  C/P:2,P

T’NT: TELZEY & TRIGGER [TELZEY] SCHMITZ, JAMES H./FLINT, ERIC/

GORDON, GUY

C/P,P

TRIGGER & FRIENDS [TELZEY]  SCHMITZ, JAMES H./FLINT, ERIC  C/P,P
UNIVERSE AGAINST HER, THE [TELZEY 1]  SCHMITZ, JAMES H.  C/P:2,P
BLACK POWDER WAR [TEMERAIRE 3]  NOVIK  C/P:2,P
HIS MAJESTY’S DRAGON [TEMERAIRE 1]  NOVIK, NAOMI  P
THRONE OF JADE [TEMERAIRE 2]  NOVIK, NAOMI  C/P,P
TEMPS [@TEMPS 1]  GAIMAN, NEIL/STEWART, ALEX  SR-PA
BLACK JADE ROAD, THE [TEN THOUSAND WILLOWS 2]  GRANT, KATHRYN  C/P,P
PHOENIX BELLS, THE [TEN THOUSAND WILLOWS 1]  GRANT, KATHRYN  P,SFWA-P
WILLOW GARDEN, THE [TEN THOUSAND WILLOWS 3]  GRANT, KATHRYN  C/P,P
ENTERING TENEBREA [@TENEBREA #1]  DAWSON, ROXANN/GRAHAM, DANIEL  P
TENEBREA’S HOPE [@TENEBREA #2]  DAWSON, ROXANN/GRAHAM, DANIEL  P
TENEBREA RISING [@TENEBREA #3]  DAWSON, ROXANN/GRAHAM, DANIEL  P

TENTH PLANET, THE [@TENTH PLANET 1] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

TENTH PLANET: FINAL ASSAULT, THE [@TENTH PLANET 3] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

TENTH PLANET: OBLIVION, THE [@TENTH PLANET 2] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

TERMINATOR, THE [TERMINATOR]  FRAKES, RANDALL/WISHER, BILL  P
TERMINATOR 2 [TERMINATOR]  FRAKES, RANDALL  C/P,P
TERMINATOR 3: RISE OF THE MACHINES [TERMINATOR]  HAGBERG, DAVID  P
TERMINATOR: THE BURNING EARTH [TERMINATOR]  FORTIER, RON  CX
T2: INFILTRATOR [TERMINATOR T 2 1]  STIRLING, S. M.  H
T2: RISING STORM [TERMINATOR T 2 2]  STIRLING, S. M.  P
T2: THE FUTURE WAR [TERMINATOR T 2 3]  STIRLING, S. M.  P
HOUR OF THE WOLF [TERMINATOR 2]  TIEDEMANN, MARK W.  LP
DARK FUTURES [@TERMINATOR 2 NEW JOHN CONNOR 1]  BLACKFORD, RUSSELL  P
TIMES OF TROUBLE [@TERMINATOR 2 NEW JOHN CONNOR 3]  BLACKFORD, RUSSELL  P
EVIL HOUR, AN [@TERMINATOR 2 NEW JOHN CONNOR #2]  BLACKFORD, RUSSELL  P
TERMINATOR DREAMS [TERMINATOR 3]  ALLSTON, AARON  P
CALLIPYGIA [TERRA MAGICA 4]  CARTER, LIN  P
DRAGONROUGE [TERRA MAGICA 2]  CARTER, LIN  C/P,P
KESRICK [TERRA MAGICA 1]  CARTER, LIN  C/P:2,P
MANDRICARDO [TERRA MAGICA 3]  CARTER, LIN  P
ENEMY OF MY ENEMY [TERRA NOVA]  OHLANDER, BEN/DRAKE, DAVID  C/P,P
WARRIOR CHALLENGED, THE [TERRILIAN 4]  GREEN, SHARON  C/P:2,P
WARRIOR ENCHAINED, THE [TERRILIAN 2]  GREEN, SHARON  C/P:2,P
WARRIOR REARMED, THE [TERRILIAN 3]  GREEN, SHARON  C/P:2,P
WARRIOR VICTORIOUS, THE [TERRILIAN 5]  GREEN, SHARON  P
WARRIOR WITHIN, THE [TERRILIAN 1]  GREEN, SHARON  C/P:2,P
TESSERACTS NINE [TESSERACTS 9]  HOPKINSON, NALO/RYMAN, GEOFF  LPA

TITLE  AUTHOR  CODE
PANDORA DIRECTIVE, THE [@TEX MURPHY]  CONNERS, AARON  P
UNDER A KILLING MOON [TEX MURPHY]  CONNERS, AARON  P
AYES OF TEXAS, THE [TEXAS 1]  DA CRUZ, DANIEL  C/P,P
TEXAS ON THE ROCKS [TEXAS 2]  DA CRUZ, DANIEL  C/P,P
TEXAS TRIUMPHANT [TEXAS 3]  DA CRUZ, DANIEL  C/P,P
THERE WILL BE WAR [@THERE WILL BE WAR #1]  POURNELLE, JERRY  PA,C/PA
MEN OF WAR [@THERE WILL BE WAR #2]  POURNELLE, JERRY  C/PA,PA
BLOOD AND IRON [@THERE WILL BE WAR #3]  POURNELLE, JERRY  C/PA,PA
DAY OF THE TYRANT [@THERE WILL BE WAR #4]  POURNELLE, JERRY  PA,C/PA,SFWA-P
WARRIOR! [@THERE WILL BE WAR #5]  POURNELLE, JERRY  C/PA,PA,SFWA-P
GUNS OF DARKNESS [@THERE WILL BE WAR #6]  POURNELLE, JERRY  PA,C/PA,SFWA-P
CALL TO BATTLE! [@THERE WILL BE WAR #7]  POURNELLE, JERRY  PA,SFWA-P
ARMAGEDDON! [@THERE WILL BE WAR #8]  POURNELLE, JERRY  PA,SFWA-P
LOVE MACHINE, THE [@THESE LAWLESS WORLDS #1]  COMSTOCK, JARROD  P
SCALES OF JUSTICE [@THESE LAWLESS WORLDS #2]  COMSTOCK, JARROD  P
DAGGER [THIEVES’ WORLD]  DRAKE, DAVID  C/P,SFWA-P,P,H
LYTHANDE [THIEVES’ WORLD]  BRADLEY, MARION ZIMMER  C/P:2,P
PRICE OF VICTORY, THE [THIEVES’ WORLD 10,11,12]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  H:2
SANCTUARY [THIEVES’ WORLD]  ABBEY, LYNN  H
SHADOWS OF SANCTUARY (BRITISH) [THIEVES’ WORLD 3]  ASPRIN, ROBERT LYNN  PA
THIEVES’ WORLD (BRITISH) [THIEVES’ WORLD 1]  ASPRIN, ROBERT LYNN  PA
TURNING POINTS [THIEVES’ WORLD]  ABBEY, LYNN  HA
TURNING POINTS (PROOF) [THIEVES’ WORLD]  ABBEY, LYNN  SR-LPA
THIEVES’ WORLD [@THIEVES’ WORLD #1]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
TALES FROM THE VULGAR UNICORN [@THIEVES’ WORLD #2]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
SHADOWS OF SANCTUARY [@THIEVES’ WORLD #3]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
STORM SEASON [@THIEVES’ WORLD #4]  ASPRIN, ROBERT LYNN  PA,C/PA
FACE OF CHAOS, THE [@THIEVES’ WORLD #5]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
WINGS OF OMEN [@THIEVES’ WORLD #6]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
DEAD OF WINTER, THE [@THIEVES’ WORLD #7]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
SOUL OF THE CITY [@THIEVES’ WORLD #8]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
BLOOD TIES [@THIEVES’ WORLD #9]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
AFTERMATH [@THIEVES’ WORLD #10]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
UNEASY ALLIANCES [@THIEVES’ WORLD #11]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
STEALER’S SKY [@THIEVES’ WORLD #12]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA,SFWA-P
SHADOW OF SORCERY, THE [THIEVES’ WORLD SHADOWSPAWN]  OFFUTT, ANDREW J.  P,SFWA-P:2
SHADOWSPAWN [THIEVES’ WORLD SHADOWSPAWN]  OFFUTT, ANDREW J.  H:2,P
BEYOND SANCTUARY [THIEVES’ WORLD TEMPUS 1]  MORRIS, JANET  H:2,P
BEYOND THE VEIL [THIEVES’ WORLD TEMPUS 2]  MORRIS, JANET  C/P,P,H
BEYOND WIZARDWALL [THIEVES’ WORLD TEMPUS]  MORRIS, JANET  H:2,SFWA-P,P
CITY AT THE EDGE OF TIME [THIEVES’ WORLD TEMPUS 5]  MORRIS, JANET/MORRIS, CHRIS  C/P:2,SFWA-P,P
STORM SEED [THIEVES’ WORLD TEMPUS 7]  MORRIS, JANET/MORRIS, CHRIS  P
TEMPUS [THIEVES’ WORLD TEMPUS 4]  MORRIS, JANET  C/P:2,P
TEMPUS UNBOUND [THIEVES’ WORLD TEMPUS 6]  MORRIS, JANET/MORRIS, CHRIS  C/P:2,P
ACROSS THE THLASSA MEY [THLASSA MEY 4]  MCCARTY, DENNIS  C/P,P
BIRTH OF THE BLADE, THE [THLASSA MEY 5]  MCCARTY, DENNIS  C/P,P
FLIGHT TO THLASSA MEY [THLASSA MEY 1]  MCCARTY, DENNIS  C/P,P
LORDS OF THLASSA MEY [THLASSA MEY 3]  MCCARTY, DENNIS  C/P,P
WARRIORS OF THLASSA MEY [THLASSA MEY 2]  MCCARTY, DENNIS  P
ILLEARTH WAR, THE [THOMAS COVENANT 2]  DONALDSON, STEPHEN R.  C/P,H,LP
LORD FOUL’S BANE [THOMAS COVENANT 1]  DONALDSON, STEPHEN R.  C/P,H,LP
ONE TREE, THE [THOMAS COVENANT 5]  DONALDSON, STEPHEN R.  C/P,P,H
POWER THAT PRESERVES, THE [THOMAS COVENANT 3]  DONALDSON, STEPHEN R.  C/P,LP,H:2
WHITE GOLD WIELDER [THOMAS COVENANT 6]  DONALDSON, STEPHEN R.  C/P,P,H
WOUNDED LAND, THE [THOMAS COVENANT 4]  DONALDSON, STEPHEN R.  C/P,P,H
SWORD OF POYANA [THORGRIM 2]  BAILEY, GERALD EARL  P
SWORD OF THE NURLINGAS [THORGRIM 1]  BAILEY, GERALD EARL  P
ARMIES OF MEMORY, THE [THOUSAND CULTURES 4]  BARNES, JOHN  H
EARTH MADE OF GLASS [THOUSAND CULTURES 2]  BARNES, JOHN  C/P,P,H
MERCHANTS OF SOULS, THE [THOUSAND CULTURES 3]  BARNES, JOHN  H,P
MILLION OPEN DOORS, A [THOUSAND CULTURES 1]  BARNES, JOHN  C/P,P,H
DEATH AND THRAXAS [THRAXAS]  SCOTT, MARTIN  P
THRAXAS [THRAXAS]  SCOTT, MARTIN  SR-P,C/P
GOLD COAST, THE [THREE CALIFORNIAS]  ROBINSON, KIM STANLEY  H,SFWA-P,P
PACIFIC EDGE [THREE CALIFORNIAS]  ROBINSON, KIM STANLEY  H,P
WILD SHORE, THE [THREE CALIFORNIAS]  ROBINSON, KIM STANLEY  C/P:2,LP
GREEN KNIGHT, THE [THREE DAMOSELS 1]  CHAPMAN, VERA  C/P,P
KING ARTHUR’S DAUGHTER [THREE DAMOSELS 3]  CHAPMAN, VERA  C/P:2,P
KING’S DAMOSEL, THE [THREE DAMOSELS 2]  CHAPMAN, VERA  C/P,P
BALSHAZZAR’S SERPENT [THREE KINGS 1]  CHALKER, JACK L.  H
KASPAR’S BOX [THREE KINGS 3]  CHALKER, JACK L.  P
KASPAR’S BOX (PROOF) [THREE KINGS 3]  CHALKER, JACK L.  SR-LP
MELCHIOR’S FIRE [THREE KINGS 2]  CHALKER, JACK L.  H

TITLE  AUTHOR  CODE
PALINDOR [THREE LANDS 1]  EVANS, D. R.  LP
PHENDRIC [THREE LANDS 3]  EVANS, D. R.  LP
SHADOW [THREE LANDS 2]  EVANS, D. R.  LP
IN THE EMPIRE OF SHADOW [THREE WORLDS 2]  WATT-EVANS, LAWRENCE  LP
OUT OF THIS WORLD [THREE WORLDS 1]  WATT-EVANS, LAWRENCE  H:2,P
REIGN OF THE BROWN MAGICIAN, THE [THREE WORLDS 3]  WATT-EVANS, LAWRENCE  P
WORLDS OF SHADOW [THREE WORLDS 1,2,3]  WATT-EVANS, LAWRENCE  H
WHALES ON STILTS [THRILLING TALES 1]  ANDERSON, M. T.  H
BOOK OF KANTELA, THE [THRONE OF SHERRAN 1]  GREEN, ROLAND J./MURRAY, FRIEDA A.  C/P,SFWA-P,LP
EYRE AFFAIR, THE [THURSDAY NEXT 1]  FFORDE, JASPER  LP
LOST IN A GOOD BOOK [THURSDAY NEXT 2]  FFORDE, JASPER  LP
SOMETHING ROTTEN [THURSDAY NEXT 4]  FFORDE, JASPER  H
WELL OF LOST PLOTS, THE [THURSDAY NEXT 3]  FFORDE, JASPER  H
PRINCE OF FIRE AND ASHES [TIELMARAN 3]  REIMANN, KATYA  P
TREMOR IN THE BITTER EARTH, A [TIELMARAN 2]  REIMANN, KATYA  H,P
WIND FROM A FOREIGN SKY [TIELMARAN 1]  REIMANN, KATYA  H,P
SWORD-BORN [TIGER 5]  ROBERSON, JENNIFER  P
SWORD-BREAKER [TIGER 4]  ROBERSON, JENNIFER  C/P:2,SFWA-P:2,P
SWORD-DANCER [TIGER 1]  ROBERSON, JENNIFER  C/P:2,P
SWORD-MAKER [TIGER 3]  ROBERSON, JENNIFER  C/P:2,P
SWORD-SINGER [TIGER 2]  ROBERSON, JENNIFER  C/P:2,P
SWORD-SWORN [TIGER 6]  ROBERSON, JENNIFER  H,P
FIRE MUSIC [TIGUAFAYA 2]  GRAY, JULIA  C/P,P
ICE MAGE [TIGUAFAYA 1]  GRAY, JULIA  P
FROM TIME TO TIME [TIME AND AGAIN 2]  FINNEY, JACK  H,LP
TIME AND AGAIN [TIME AND AGAIN 1]  FINNEY, JACK  H:2,LP
TIME GATE [@TIME GATE #1]  SILVERBERG, ROBERT/FAWCETT, BILL  PA,C/PA
DANGEROUS INTERFACES [@TIME GATE #2]  SILVERBERG, ROBERT/FAWCETT, BILL  PA,C/PA,SFWA-P
ALBINO’S DANCER, THE [TIME HUNTER 9]  SMITH, DALE  LP
CLOCKWORK WOMAN, THE [@TIME HUNTER 3]  BOTT, CLAIRE  LP
DEUS LE VOLT [TIME HUNTER 8]  MILLER, JON DE BURGH  SR-H
ECHOES [@TIME HUNTER 6]  MCLAUGHLIN, IAIN/BARTLETT, CLAIRE  LP
KITSUNE [@TIME HUNTER 4]  CATTON, JOHN PAUL  LP
PECULIAR LIVES [@TIME HUNTER 7]  PURSER-HALLARD, PHILIP  LP
SEVERED MAN, THE [@TIME HUNTER 5]  MANN, GEORGE  LP
TUNNEL AT THE END OF THE LIGHT, THE [@TIME HUNTER 2]  PETRUCHA, STEFAN  LP
WINNING SIDE, THE [@TIME HUNTER 1]  PARKIN, LANCE  LP
CHILD OF TOMORROW [TIME KEEPER 2]  BARTHOLOMEW, BARBARA  P
TIME KEEPER, THE [TIME KEEPER 1]  BARTHOLOMEW, BARBARA  P
WHEN DREAMERS CEASE TO DREAM [TIME KEEPER 3]  BARTHOLOMEW, BARBARA  P
SECOND TIME AROUND [TIME LOTTERY 2]  MOSER, NANCY  LP
TIME LOTTERY [TIME LOTTERY 1]  MOSER, NANCY  LP
SECRET OF THE KNIGHTS [@TIME MACHINE #1]  GASPERINI, JIM  C/P,P
SEARCH FOR DINOSAURS [@TIME MACHINE #2]  BISCHOFF, DAVID F.  P
SAIL WITH PIRATES [@TIME MACHINE #4]  GASPERINI, JIM  P
INITIATE, THE [TIME MASTER 1]  COOPER, LOUISE  C/P:2,SFWA-P,P
MASTER, THE [TIME MASTER 3]  COOPER, LOUISE  C/P:2,SFWA-P,P
OUTCAST, THE [TIME MASTER 2]  COOPER, LOUISE  C/P:2,SFWA-P,P
STAR ASCENDANT [TIME MASTER 0]  COOPER, LOUISE  C/P,P,H
SUNSTORM [TIME ODYSSEY 2]  CLARKE, ARTHUR C./BAXTER, STEPHEN  H
TIME’S EYE (CD) [TIME ODYSSEY 1]  CLARKE, ARTHUR C./BAXTER, STEPHEN  SR-MM
TIME’S EYE (HARDCOVER AND CD) [TIME ODYSSEY 1]  CLARKE, ARTHUR C./BAXTER, STEPHEN  H
TIME’S EYE (PROOF) [TIME ODYSSEY 1]  CLARKE, ARTHUR C./BAXTER, STEPHEN  SR-LP
HOUSE OF REEDS [TIME OF THE SIXTH SUN 2]  HARLAN, THOMAS  H
WASTELAND OF FLINT [TIME OF THE SIXTH SUN 1]  HARLAN, THOMAS  H
HAMMER AND ANVIL [TIME OF TROUBLES 2]  TURTLEDOVE, HARRY  P
STOLEN THRONE, THE [TIME OF TROUBLES 1]  TURTLEDOVE, HARRY  P
THOUSAND CITIES, THE [TIME OF TROUBLES 3]  TURTLEDOVE, HARRY  P
VANISHED [@TIME POLICE #1]  NORWOOD, WARREN  C/P,P
TRAPPED! [@TIME POLICE #2]  NORWOOD, WARREN  P
STRANDED [@TIME POLICE #3]  NORWOOD, WARREN/ODOM, MEL  C/P,P
FOR KING & COUNTRY [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  H,P
HOUSE THAT JACK BUILT, THE [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  P
RIPPING TIME [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  P
TIME SCOUT [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  C/P:2,P
WAGERS OF SIN [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  P
TIME STATION BERLIN [TIME STATION 3]  EVANS, DAVID  P
TIME STATION LONDON [TIME STATION 1]  EVANS, DAVID  C/P,P
TIME STATION PARIS [TIME STATION 2]  EVANS, DAVID  P
SHIP THAT SAILED THE TIME STREAM, THE [TIME STREAM 1]  EDMONDSON, G. C.  DM109
SHIP THAT SAILED THE TIME STREAM, THE (REVISED) [TIME STREAM 1]  EDMONDSON, G. C.  C/P:3,P
TO SAIL THE CENTURY SEA [TIME STREAM 2]  EDMONDSON, G. C.  C/P:2,P
GLORY’S END [TIME TOURS 2]  BARON, NICK  P

The MITSFS Seriesdex  September 24, 2007  Page 169
TITLE  AUTHOR  CODE
ROBIN HOOD AMBUSH, THE [TIME TOURS 1]  WU, WILLIAM F.  P
ATLANTIS ENDGAME [TIME TRADERS]  NORTON, ANDRE/SMITH, SHERWOOD  P
DEFIANT AGENTS, THE [TIME TRADERS 3]  NORTON, ANDRE  C/P,P,H

DEFIANT AGENTS, THE (REVISED) [TIME TRADERS 3,4]/ KEY OUT OF TIME (REVISED) [TIME TRADERS 3,4] NORTON, ANDRE

P

ECHOES IN TIME [TIME TRADERS 6]  NORTON, ANDRE/SMITH, SHERWOOD  H
FIREHAND [TIME TRADERS 5]  NORTON, ANDRE/GRIFFIN, P. M.  H:2
GALACTIC DERELICT [TIME TRADERS 2]  NORTON, ANDRE  C/P:2,P
KEY OUT OF TIME [TIME TRADERS 4]  NORTON, ANDRE  C/P:2,P
TIME TRADERS, THE [TIME TRADERS 1]  NORTON, ANDRE  C/P:2,P

TIME TRADERS (REVISED) [TIME TRADERS 1,2]/ GALACTIC DERELICT (REVISED) [TIME TRADERS 1,2] NORTON, ANDRE

P

TIME TRADERS III [TIME TRADERS 3,4]  NORTON, ANDRE/SMITH, SHERWOOD  H
PIECES OF EIGHT [TIME TRAP 3]  DALTON, SEAN  P
SHOWDOWN [TIME TRAP 2]  DALTON, SEAN  P
TIME TRAP [TIME TRAP 1]  DALTON, SEAN  P
MICRONAUTS [TIME TRAVELER 3]  LYONS, STEVE  P
TIME TUNNEL, THE [TIME TUNNEL 1]  LEINSTER, MURRAY  C/P:2,P
TIMESLIP! [TIME TUNNEL 2]  LEINSTER, MURRAY  C/P:2,P
TIME OF THE FOX [TIME WARRIOR 1]  COSTELLO, MATTHEW J.  C/P,P
HOUR OF THE SCORPION [@TIME WARRIOR #2]  COSTELLO, MATTHEW J.  P
FUSE POINT [TIME WARRIORS #1]  NORTH, DAVID  C/P,P
FORBIDDEN REGION [@TIME WARRIORS #2]  NORTH, DAVID  C/P,P
BLOOD TIES [@TIMECOP 3]  PARKINSON, DAN  P
SCAVENGER, THE [@TIMECOP 2]  PARKINSON, DAN  P
VIPER’S SPAWN [@TIMECOP 1]  PARKINSON, DAN  P
TIMEDIVER’S DAWN [TIMEDIVERS 2]  MODESITT, L. E. JR.  C/P:2,P
TIMEGOD, THE [TIMEDIVERS 1]  MODESITT, L. E. JR.  P
TIMEGODS’ WORLD [TIMEDIVERS 1,2]  MODESITT, L. E. JR.  H
CLOUDS OF MAGELLAN [TIMELAPSE 2]  NIGHBERT, DAVID E.  H,SFWA
PATTON’S SPACESHIP [@TIMELINE WARS #1]  BARNES, JOHN  C/P:2,P
WASHINGTON’S DIRIGIBLE [@TIMELINE WARS #2]  BARNES, JOHN  C/P:2,P
CAESAR’S BICYCLE [@TIMELINE WARS #3]  BARNES, JOHN  C/P,P
RASHANYN DARK [@TIMEQUEST #1]  TEDFORD, WILLIAM  C/P,P
HYDRABYSS RED [@TIMEQUEST #2]  TEDFORD, WILLIAM  P
NEMYDIA DEEP [@TIMEQUEST #3]  TEDFORD, WILLIAM  P
UNION FIRES [@TIMERAIDER]  BARNES, JOHN  P
WARTIDE [@TIMERAIDER]  BARNES, JOHN  P
EMPEROR [@TIME’S TAPESTRY #1]  BAXTER, STEPHEN  H
SECOND TIME AROUND [@TIMESHARE 2]  DANN, JOSHUA  P
TIME FOR WAR, A [@TIMESHARE 3]  DANN, JOSHUA  P
TIMESHARE [@TIMESHARE 1]  DANN, JOSHUA  P
ARGONAUT AFFAIR, THE [TIMEWARS 7]  HAWKE, SIMON  P
CLEOPATRA CRISIS, THE [TIMEWARS 11]  HAWKE, SIMON  P
DRACULA CAPER, THE [TIMEWARS 8]  HAWKE, SIMON  C/P,P
HELLFIRE REBELLION, THE [TIMEWARS 10]  HAWKE, SIMON  C/P,P
IVANHOE GAMBIT, THE [TIMEWARS 1]  HAWKE, SIMON  C/P:2,P
KHYBER CONNECTION, THE [TIMEWARS 6]  HAWKE, SIMON  P
LILLIPUT LEGION, THE [TIMEWARS 9]  HAWKE, SIMON  C/P,P
NAUTILUS SANCTION, THE [TIMEWARS 5]  HAWKE, SIMON  C/P:2,P
PIMPERNEL PLOT, THE [TIMEWARS 3]  HAWKE, SIMON  C/P:2,P
SIX-GUN SOLUTION, THE [TIMEWARS 12]  HAWKE, SIMON  C/P,P
TIMEKEEPER CONSPIRACY, THE [TIMEWARS 2]  HAWKE, SIMON  C/P:2,P
ZENDA VENDETTA, THE [TIMEWARS 4]  HAWKE, SIMON  C/P:2,P
VAMPIRE JUNCTION [TIMOTHY VALENTINE 1]  SOMTOW, S. P.  C/P:2,P
VANITAS [TIMOTHY VALENTINE 3]  SOMTOW, S. P.  H
GODS AWAKEN, THE [TIMURA 3]  COLE, ALLAN  P
WIZARD OF THE WINDS [TIMURA 1]  COLE, ALLAN  LP
WOLVES OF THE GODS [TIMURA 2]  COLE, ALLAN  P
TINKER [TINKER 1]  SPENCER, WEN  P
WOLF WHO RULES [TINKER 2]  SPENCER, WEN  H:2
HOUSE OF GAIAN [TIR ALAINN 3]  BISHOP, ANNE  P
PILLARS OF THE WORLD, THE [TIR ALAINN 1]  BISHOP, ANNE  C/P,P
SHADOWS AND LIGHT [TIR ALAINN 2]  BISHOP, ANNE  C/P,P
KING’S NAME, THE [TIR TANAGIRI 2]  WALTON, JO  H
KING’S PEACE, THE [TIR TANAGIRI 1]  WALTON, JO  H
PRIZE IN THE GAME, THE [TIR TANAGIRI 3]  WALTON, JO  H,P
PRIZE IN THE GAME, THE (PROOF) [TIR TANAGIRI 3]  WALTON, JO  SR-LP
AKIMA’S STORY [@TITAN A.E.]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
CALE’S STORY [@TITAN A.E.]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
LORD OF THE CROOKED PATHS [TITANS 1]  ADKINS, PATRICK H.  P
MASTER OF THE FEARFUL DEPTHS [TITANS 2]  ADKINS, PATRICK H.  P
SONS OF THE TITANS [TITANS 3]  ADKINS, PATRICK H.  P

TITLE  AUTHOR  CODE
CLOCK OF DREAMS, THE [TITUS CROW]  LUMLEY, BRIAN  P
COMPLEAT CROW, THE [TITUS CROW]  LUMLEY, BRIAN  LP
TITUS CROW VOLUME THREE [TITUS CROW 5,6]  LUMLEY, BRIAN  LP
TRANSITION OF TITUS CROW, THE [TITUS CROW 1]  LUMLEY, BRIAN  C/P,P
HOMEWORLD [TO THE STARS 1]  HARRISON, HARRY  C/P:2,P
STARWORLD [TO THE STARS 3]  HARRISON, HARRY  C/P:2,P
TO THE STARS [TO THE STARS 1,2,3]  HARRISON, HARRY  H:2,P
WHEELWORLD [TO THE STARS 2]  HARRISON, HARRY  C/P:2,P
BONE DOLL’S TWIN, THE [TOBIN OF SKALA 1]  FLEWELLING, LYNN  P
HIDDEN WARRIOR [TOBIN OF SKALA 2]  FLEWELLING, LYNN  C/P,P
ORACLE’S QUEEN, THE [TOBIN OF SKALA 3]  FLEWELLING, LYNN  C/P,P
FUGITIVE WORLDS, THE [TOLLER MARAQUINE 3]  SHAW, BOB  P
RAGGED ASTRONAUTS, THE [TOLLER MARAQUINE 1]  SHAW, BOB  H:2,SFWA,P
WOODEN SPACESHIPS, THE [TOLLER MARAQUINE 2]  SHAW, BOB  C/P:2,SFWA-P,P
DANGER IN DEEP SPACE [TOM CORBETT 2]  ROCKWELL, CAREY  H
ON THE TRAIL OF THE SPACE PIRATES [TOM CORBETT 3]  ROCKWELL, CAREY  H:2
REVOLT ON VENUS, THE [TOM CORBETT 5]  ROCKWELL, CAREY  H:2
ROBOT ROCKET, THE [TOM CORBETT 8]  ROCKWELL, CAREY  H:2
SABOTAGE IN SPACE [TOM CORBETT 7]  ROCKWELL, CAREY  H
SPACE PIONEERS, THE [TOM CORBETT 4]  ROCKWELL, CAREY  H
STAND BY FOR MARS! [TOM CORBETT 1]  ROCKWELL, CAREY  H
TOM CORBETT: A TRIP TO THE MOON [TOM CORBETT]  MARTIN, MARCIA  BLB
TREACHERY IN OUTER SPACE [TOM CORBETT 6]  ROCKWELL, CAREY  H
TOM SWIFT AND HIS MOTOR-CYCLE [@TOM SWIFT #1]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS MOTOR-BOAT [@TOM SWIFT #2]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS AIRSHIP [@TOM SWIFT #3]  APPLETON, VICTOR  H
TOM SWIFT AND HIS SUBMARINE BOAT [@TOM SWIFT #4]  APPLETON, VICTOR  H
TOM SWIFT AND HIS ELECTRIC RUNABOUT [@TOM SWIFT #5]  APPLETON, VICTOR  H
TOM SWIFT AND HIS WIRELESS MESSAGE [@TOM SWIFT #6]  APPLETON, VICTOR  H
TOM SWIFT AMONG THE DIAMOND MAKERS [@TOM SWIFT #7]  APPLETON, VICTOR  H
TOM SWIFT IN THE CAVES OF ICE [@TOM SWIFT #8]  APPLETON, VICTOR  H
TOM SWIFT AND HIS SKY RACER [@TOM SWIFT #9]  APPLETON, VICTOR  H
TOM SWIFT AND HIS ELECTRIC RIFLE [@TOM SWIFT #10]  APPLETON, VICTOR  H:2
TOM SWIFT IN THE CITY OF GOLD [@TOM SWIFT #11]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AIR GLIDER [@TOM SWIFT #12]  APPLETON, VICTOR  H
TOM SWIFT IN CAPTIVITY [@TOM SWIFT #13]  APPLETON, VICTOR  H
TOM SWIFT AND HIS WIZARD CAMERA [@TOM SWIFT #14]  APPLETON, VICTOR  H
TOM SWIFT AND HIS GREAT SEARCHLIGHT [@TOM SWIFT #15]  APPLETON, VICTOR  H
TOM SWIFT AND HIS GIANT CANNON [@TOM SWIFT #16]  APPLETON, VICTOR  H
TOM SWIFT AND HIS PHOTO TELEPHONE [@TOM SWIFT #17]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS AERIAL WARSHIP [@TOM SWIFT #18]  APPLETON, VICTOR  H
TOM SWIFT AND HIS BIG TUNNEL [@TOM SWIFT #19]  APPLETON, VICTOR  H:2
TOM SWIFT IN THE LAND OF WONDERS [@TOM SWIFT #20]  APPLETON, VICTOR  H
TOM SWIFT AND HIS WAR TANK [@TOM SWIFT #21]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AIR SCOUT [@TOM SWIFT #22]  APPLETON, VICTOR  H
TOM SWIFT AND HIS UNDERSEA SEARCH [@TOM SWIFT #23]  APPLETON, VICTOR  H
TOM SWIFT AMONG THE FIRE FIGHTERS [@TOM SWIFT #24]  APPLETON, VICTOR  H
TOM SWIFT AND HIS ELECTRIC LOCOMOTIVE [@TOM SWIFT #25]  APPLETON, VICTOR  H
TOM SWIFT AND HIS FLYING BOAT [@TOM SWIFT #26]  APPLETON, VICTOR  H
TOM SWIFT AND HIS GREAT OIL GUSHER [@TOM SWIFT #27]  APPLETON, VICTOR  H
TOM SWIFT AND HIS CHEST OF SECRETS [@TOM SWIFT #28]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS AIRLINE EXPRESS [@TOM SWIFT #29]  APPLETON, VICTOR  H:2
TOM SWIFT CIRCLING THE GLOBE [@TOM SWIFT #30]  APPLETON, VICTOR  H
TOM SWIFT AND HIS TALKING PICTURES [@TOM SWIFT #31]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS HOUSE ON WHEELS [@TOM SWIFT #32]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS BIG DIRIGIBLE [@TOM SWIFT #33]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS SKY TRAIN [@TOM SWIFT #34]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS GIANT MAGNET [@TOM SWIFT #35]  APPLETON, VICTOR  H
TOM SWIFT AND HIS TELEVISION DETECTOR [@TOM SWIFT #36]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS OCEAN AIRPORT [@TOM SWIFT #37]  APPLETON, VICTOR  H
TOM SWIFT AND HIS PLANET STONE [@TOM SWIFT #38]  APPLETON, VICTOR  H
BLACK DRAGON, THE [@TOM SWIFT (ARCHWAY) #1]  APPLETON, VICTOR  P
NEGATIVE ZONE, THE [@TOM SWIFT (ARCHWAY) #2]  APPLETON, VICTOR  P
CYBORG KICKBOXER [@TOM SWIFT (ARCHWAY) #3]  APPLETON, VICTOR  P
DNA DISASTER, THE [@TOM SWIFT (ARCHWAY) #4]  APPLETON, VICTOR  P
MONSTER MACHINE [@TOM SWIFT (ARCHWAY) #5]  APPLETON, VICTOR  P
AQUATECH WARRIORS [@TOM SWIFT (ARCHWAY) #6]  APPLETON, VICTOR  P
MOONSTALKER [@TOM SWIFT (ARCHWAY) #7]  APPLETON, VICTOR  P
MICROBOTS, THE [@TOM SWIFT (ARCHWAY) #8]  APPLETON, VICTOR  P
FIRE BIKER [@TOM SWIFT (ARCHWAY) #9]  APPLETON, VICTOR  P
MIND GAMES [@TOM SWIFT (ARCHWAY) #10]  APPLETON, VICTOR  P
MUTANT BEACH [@TOM SWIFT (ARCHWAY) #11]  APPLETON, VICTOR  P
DEATH QUAKE [@TOM SWIFT (ARCHWAY) #12]  APPLETON, VICTOR  P

TITLE  AUTHOR  CODE
TOM SWIFT AND HIS FLYING LAB [@TOM SWIFT JR. #1]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS JETMARINE [@TOM SWIFT JR. #2]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS ROCKET SHIP [@TOM SWIFT JR. #3]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS GIANT ROBOT [@TOM SWIFT JR. #4]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS ATOMIC EARTH BLASTER [@TOM SWIFT JR. #5]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS OUTPOST IN SPACE [@TOM SWIFT JR. #6]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS DIVING SEACOPTER [@TOM SWIFT JR. #7]  APPLETON, VICTOR II  H:2
TOM SWIFT IN THE CAVES OF NUCLEAR FIRE [@TOM SWIFT JR. #8]  APPLETON, VICTOR II  H:2
TOM SWIFT ON THE PHANTOM SATELLITE [@TOM SWIFT JR. #9]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS ULTRASONIC CYCLOPLANE [@TOM SWIFT JR. #10]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS DEEP-SEA HYDRODOME [@TOM SWIFT JR. #11]  APPLETON, VICTOR II  H:2
TOM SWIFT IN THE RACE TO THE MOON [@TOM SWIFT JR. #12]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS SPACE SOLARTRON [@TOM SWIFT JR. #13]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS ELECTRONIC RETROSCOPE [@TOM SWIFT JR. #14]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS SPECTROMARINE SELECTOR [@TOM SWIFT JR. #15]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE COSMIC ASTRONAUTS [@TOM SWIFT JR. #16]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE VISITOR FROM PLANET X [@TOM SWIFT JR. #17]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE ELECTRONIC HYDROLUNG [@TOM SWIFT JR. #18]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS TRIPHIBIAN ATOMICAR [@TOM SWIFT JR. #19]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS MEGASCOPE SPACE PROBER [@TOM SWIFT JR. #20]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE ASTEROID PIRATES [@TOM SWIFT JR. #21]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS REPELATRON SKYWAY [@TOM SWIFT JR. #22]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS AQUATOMIC TRACKER [@TOM SWIFT JR. #23]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS 3-D TELEJECTOR [@TOM SWIFT JR. #24]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS POLAR-RAY DYNASPHERE [@TOM SWIFT JR. #25]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS SONIC BOOM TRAP [@TOM SWIFT JR. #26]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS SUBOCEAN GEOTRON [@TOM SWIFT JR. #27]  APPLETON, VICTOR II  H
TOM SWIFT AND THE MYSTERY COMET [@TOM SWIFT JR. #28]  APPLETON, VICTOR II  H
TOM SWIFT AND THE CAPTIVE PLANETOID [@TOM SWIFT JR. #29]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS G-FORCE INVERTER [@TOM SWIFT JR. #30]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS DYNA-4 CAPSULE [@TOM SWIFT JR. #31]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS COSMOTRON EXPRESS [@TOM SWIFT JR. #32]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE GALAXY GHOSTS [@TOM SWIFT JR. #33]  APPLETON, VICTOR II  H
STARS COMPEL, THE [TOMMASO 2]  ROESSNER, MICHAELA  P
STARS DISPOSE, THE [TOMMASO 1]  ROESSNER, MICHAELA  C/P,P
GOLDEN NAGINATA, THE [TOMOE GOZEN 2]  SALMONSON, JESSICA AMANDA  C/P:2,P
THOUSAND SHRINE WARRIOR [TOMOE GOZEN 3]  SALMONSON, JESSICA AMANDA  C/P:2,P
TOMOE GOZEN [TOMOE GOZEN 1]  SALMONSON, JESSICA AMANDA  C/P:2,P
SMOKE AND ASHES [TONY FOSTER 3]  HUFF, TANYA  H:2
SMOKE AND MIRRORS [TONY FOSTER 2]  HUFF, TANYA  H
SMOKE AND SHADOWS [TONY FOSTER 1]  HUFF, TANYA  H
ROGUE POWERS [TORCH OF HONOR 2]  ALLEN, ROGER MACBRIDE  C/P,P
TORCH OF HONOR, THE [TORCH OF HONOR 1]  ALLEN, ROGER MACBRIDE  C/P:2,P
DRAGONS OVER ENGLAND [TORG POSSIBILITY WARS]  WEST END BOOKS  PA
STRANGE TALES FROM THE NILE EMPIRE [TORG POSSIBILITY WARS]  WEST END BOOKS  PA
STORM KNIGHTS [@TORG POSSIBILITY WARS #1]  SLAVICSEK, BILL/TRAMONTANA, C. J.  P
DARK REALM, THE [@TORG POSSIBILITY WARS #2]  KAUFMAN, DOUGLAS  C/P,P
NIGHTMARE DREAM, THE [@TORG POSSIBILITY WARS #3]  CASPIAN, JONATHA ARIADNE  C/P,P
LUCK OF BRIN’S FIVE, THE [TORIN 1]  WILDER, CHERRY  H,P
NEAREST FIRE, THE [TORIN 2]  WILDER, CHERRY  H,P
TAPESTRY WARRIORS, THE [TORIN 3]  WILDER, CHERRY  H
BETTER PART OF VALOR, THE [TORIN KERR 2]  HUFF, TANYA  P
CONFEDERATION OF VALOR, A [TORIN KERR 1,2]  HUFF, TANYA  P
VALOR’S CHOICE [TORIN KERR 1]  HUFF, TANYA  P
DANCERS OF ARUN, THE [TORNOR 2]  LYNN, ELIZABETH A.  C/P,P,H,SFWA
NORTHERN GIRL, THE [TORNOR 3]  LYNN, ELIZABETH A.  C/P:2,SFWA,P
WATCHTOWER [TORNOR 1]  LYNN, ELIZABETH A.  C/P,P,H,SFWA
MYRIAD, THE [@TOUR OF THE MERRIMACK #1]  MELUCH, R. M.  H
WOLF STAR [@TOUR OF THE MERRIMACK #2]  MELUCH, R. M.  H
THOUSAND WORDS FOR STRANGER, A [TRADE PACT 1]  CZERNEDA, JULIE E.  H:2,P
TIES OF POWER [@TRADE PACT #2]  CZERNEDA, JULIE E.  C/P,P,H
TO TRADE THE STARS [@TRADE PACT #3]  CZERNEDA, JULIE E.  P
FIRST, YOU FIGHT [@TRAVELER #1]  DRUMM, D. B.  P
KINGDOM COME [@TRAVELER #2]  DRUMM, D. B.  C/P,P
STALKERS, THE [@TRAVELER #3]  DRUMM, D. B.  P
TO KILL A SHADOW [@TRAVELER #4]  DRUMM, D. B.  P
ROAD WAR [@TRAVELER #5]  DRUMM, D. B.  P
BORDER WAR [@TRAVELER #6]  DRUMM, D. B.  P
ROAD GHOST, THE [@TRAVELER #7]  DRUMM, D. B.  P
TERMINAL ROAD [@TRAVELER #8]  DRUMM, D. B.  P
STALKING TIME, THE [@TRAVELER #9]  DRUMM, D. B.  P
HELL ON EARTH [@TRAVELER #10]  DRUMM, D. B.  P
CHILDREN’S CRUSADE, THE [@TRAVELER #11]  DRUMM, D. B.  P

The MITSFS Seriesdex  September 24, 2007  Page 172
TITLE  AUTHOR  CODE
PREY, THE [@TRAVELER #12]  DRUMM, D. B.  P
GHOST DANCERS [@TRAVELER #13]  DRUMM, D. B.  P
GATEWAY TO THE STARS [@TRAVELLER 1]  ASKEGREN, PIERCE  P
SCHISM [@TRIAD #1]  ASARO, CATHERINE  H
HEARTWOOD [TRICKSTER’S GAME 1]  CAMPBELL, BARBARA  P
DAY OF THE TRIFFIDS, THE (AMERICAN) [TRIFFIDS 1]  WYNDHAM, JOHN  C/P,P,H
DAY OF THE TRIFFIDS, THE (BRITISH) [TRIFFIDS 1]  WYNDHAM, JOHN  SR-P,C/P
NIGHT OF THE TRIFFIDS, THE [TRIFFIDS 2]  CLARK, SIMON  P
EMPERY [TRIGON DISUNITY 3]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,SFWA-P,P
EMPRISE [TRIGON DISUNITY 1]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
ENIGMA [TRIGON DISUNITY 2]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P

BLACK TRILLIUM [TRILLIUM 1] BRADLEY, MARION ZIMMER/MAY, JULIAN/

NORTON, ANDRE

C/P,P,H

BLOOD TRILLIUM [TRILLIUM 2]  MAY, JULIAN  H:2,P
GOLDEN TRILLIUM [TRILLIUM 3]  NORTON, ANDRE  H:2,P
LADY OF THE TRILLIUM [TRILLIUM 4]  BRADLEY, MARION ZIMMER  C/P,P,H
SKY TRILLIUM [TRILLIUM]  MAY, JULIAN  H,P

DANCE OF THE VOODOO HANDBAG, THE [TRILOGY THAT DARE NOT SPEAK ITS NAME 2] RANKIN, ROBERT

P

SPROUT MASK REPLICA [TRILOGY THAT DARE NOT SPEAK ITS NAME 1]  RANKIN, ROBERT  P
WAITING FOR GODALMING [TRILOGY THAT DARE NOT SPEAK ITS NAME 3] RANKIN, ROBERT  P
ESCAPE FROM CHINA [TRIO]  MONTGOMERY, R. A.  P
CITY OF GOLD AND LEAD, THE [TRIPODS 2]  CHRISTOPHER, JOHN  C/P:2,P
POOL OF FIRE, THE [TRIPODS 3]  CHRISTOPHER, JOHN  C/P:2,P
WHEN THE TRIPODS CAME [TRIPODS 0]  CHRISTOPHER, JOHN  C/P,SFWA-P,P
WHITE MOUNTAINS, THE [TRIPODS 1]  CHRISTOPHER, JOHN  C/P,P,H
FORTRESS IN THE EYE OF TIME [TRISTEN 1]  CHERRYH, C. J.  H:2,P
FORTRESS IN THE EYE OF TIME (PROOF) [TRISTEN 1]  CHERRYH, C. J.  SR-LP
FORTRESS OF DRAGONS [TRISTEN 4]  CHERRYH, C. J.  C/P,P,H
FORTRESS OF EAGLES [TRISTEN 2]  CHERRYH, C. J.  H:2,P
FORTRESS OF ICE [TRISTEN 5]  CHERRYH, C. J.  H
FORTRESS OF OWLS [TRISTEN 3]  CHERRYH, C. J.  C/P,P,H
LEGEND OF THE FIREFISH, THE [TROPHY CHASE 1]  POLIVKA, GEORGE BRYAN  LP
HELENE [TROS OF SAMOTHRACE 4]  MUNDY, TALBOT  SR-P
HELMA [TROS OF SAMOTHRACE 2]  MUNDY, TALBOT  SR-P,C/P:2
LIAFAIL [TROS OF SAMOTHRACE 3]  MUNDY, TALBOT  SR-P,C/P:2

TROS OF SAMOTHRACE [TROS OF SAMOTHRACE 1]/ LUD OF LUNDUN [TROS OF SAMOTHRACE 1] MUNDY, TALBOT

P

LORD OF THE SILVER BOW [@TROY 1]  GEMMELL, DAVID  LP
SHIELD OF THUNDER [@TROY 2]  GEMMELL, DAVID  H
DARKWITCH RISING [TROY GAME 3]  DOUGLASS, SARA  H
DRUID’S SWORD [TROY GAME 4]  DOUGLASS, SARA  H
HADES’ DAUGHTER [TROY GAME #1]  DOUGLASS, SARA  H:2,@P
GODS’ CONCUBINE [@TROY GAME #2]  DOUGLASS, SARA  P
DERVISH DAUGHTER [TRUE GAME 5]  TEPPER, SHERI S.  C/P,SFWA-P,P
END OF THE GAME, THE [TRUE GAME 4,5,6]  TEPPER, SHERI S.  H
JINIAN FOOTSEER [TRUE GAME 4]  TEPPER, SHERI S.  C/P,SFWA-P,P
JINIAN STAR-EYE [TRUE GAME 6]  TEPPER, SHERI S.  C/P,SFWA-P,P
KING’S BLOOD FOUR [TRUE GAME 1]  TEPPER, SHERI S.  C/P:2,P
NECROMANCER NINE [TRUE GAME 2]  TEPPER, SHERI S.  C/P:2,P
TRUE GAME, THE [TRUE GAME 1,2,3]  TEPPER, SHERI S.  H:2
WIZARD’S ELEVEN [TRUE GAME 3]  TEPPER, SHERI S.  C/P,P
ACCESS DENIED [TURING HOPPER 3]  ANDREWS, DONNA  P
CLICK HERE FOR MURDER [TURING HOPPER 2]  ANDREWS, DONNA  P
DELETE ALL SUSPECTS [TURING HOPPER 4]  ANDREWS, DONNA  H
DARK MOON DEFENDER [TWELVE HOUSES]  SHINN, SHARON  H
MYSTIC AND RIDER [TWELVE HOUSES]  SHINN, SHARON  H
THIRTEENTH HOUSE, THE [TWELVE HOUSES]  SHINN, SHARON  H
MUSTAPHA AND HIS WISE DOG [TWELVE KINGDOMS 1]  FRIESNER, ESTHER M.  C/P:2,P
SPELLS OF MORTAL WEAVING [TWELVE KINGDOMS 2]  FRIESNER, ESTHER M.  C/P,P
WATER KING’S LAUGHTER, THE [TWELVE KINGDOMS 4]  FRIESNER, ESTHER M.  C/P:2,P
WITCHWOOD CRADLE, THE [TWELVE KINGDOMS 3]  FRIESNER, ESTHER M.  P
CLOAK OF NIGHT AND DAGGERS, THE [TWELVE TREASURES 3]  EDGHILL, ROSEMARY  P
CUP OF MORNING SHADOWS, THE [TWELVE TREASURES 2]  EDGHILL, ROSEMARY  C/P:2,P
SWORD OF MAIDEN’S TEARS, THE [TWELVE TREASURES 1]  EDGHILL, ROSEMARY  C/P,P
MELTDOWN [@TWENTY NINETY-NINE #5]  PEEL, JOHN  LP
AVENGING SPIRIT, THE [TWILIGHT 10]  STEVENSON, E.  P
BLOOD RED ROSES [TWILIGHT 8]  ARMSTRONG, SARAH  P
DANCE OF DEATH [TWILIGHT 19]  KASEM, LOU  P
DEADLY SLEEP [TWILIGHT 1]  COWAN, DALE  P
DEMON TREE [TWILIGHT 9]  DANIEL, COLIN  P
FAMILY CRYPT [TWILIGHT 20]  TRAINOR, JOSEPH  P
FATAL ATTRACTION [TWILIGHT 4]  HOWE, IMOGEN  P

The MITSFS Seriesdex  September 24, 2007  Page 173
TITLE  AUTHOR  CODE
FOOTPRINTS OF THE DEAD [TWILIGHT 14]  CALLAHAN, JAY  P
PLAY TO LIVE [TWILIGHT 7]  VELEY, CHARLES  P
POWER, THE [TWILIGHT 6]  HAYNES, BETSY  P
STORM CHILD [TWILIGHT 17]  NETTER, SUSAN  P
VOICES IN THE DARK [TWILIGHT 2]  HAYNES, JAMES  P
INITIATION, THE [TWILIGHT #3]  BRUNN, ROBERT  P
BLINK OF THE MIND [TWILIGHT #5]  FRANCIS, DOROTHY BRENNER  P
WATERY GRAVE [TWILIGHT #18]  TRAINOR, JOSEPH  P
EVIL ON THE BAYOU [TWILIGHT #21]  CUSICK, RICHIE TANKERSLEY  P
HAUNTED DOLLHOUSE, THE [TWILIGHT #22]  BLAKE, SUSAN  P
FIRST NAME, THE [@TWILIGHT OF THE GODS #1]  SCHMIDT, DENNIS  C/P:2,P
GROA’S OTHER EYE [@TWILIGHT OF THE GODS #2]  SCHMIDT, DENNIS  C/P:2,P
THREE TRUMPS SOUNDING [@TWILIGHT OF THE GODS #3]  SCHMIDT, DENNIS  C/P,SFWA-P,P
SHADES OF NIGHT, FALLING [@TWILIGHT ZONE 1]  MILLER, JOHN J.  LP

CHANGELINGS [TWINS OF PETAYBEE 1] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2

MAELSTROM [TWINS OF PETAYBEE 2] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

CALL TO ARMS, A [TWO-EDGED SWORD 3]  MARTIN, THOMAS K.  P
MATTER OF HONOR, A [TWO-EDGED SWORD 2]  MARTIN, THOMAS K.  P
TWO-EDGED SWORD, A [TWO-EDGED SWORD 1]  MARTIN, THOMAS K.  P
MANJINN MOON [TY MERRICK 3]  VITOLA, DENISE  P
OPALITE MOON [TY MERRICK 2]  VITOLA, DENISE  P
QUANTUM MOON [TY MERRICK 1]  VITOLA, DENISE  P
RADON FILE, THE [TY MERRICK 5]  VITOLA, DENISE  P
RED SKY FILE, THE [TY MERRICK 4]  VITOLA, DENISE  P
RIDERS OF LEVIATHAN [TYPHON 2]  ANZETTI, TONI  P
TYPHON’S CHILDREN [TYPHON 1]  ANZETTI, TONI  P
GRAND DESIGN, THE [TYRANTS AND KINGS 2]  MARCO, JOHN  C/P,LP
JACKAL OF NAR, THE [TYRANTS AND KINGS 1]  MARCO, JOHN  C/P:2,LP
SAINTS OF THE SWORD, THE [TYRANTS AND KINGS 3]  MARCO, JOHN  H,LP

TITLE  AUTHOR  CODE
FLESH HUNTERS [@UFO #1]  MIALL, ROBERT  P
SPORTING BLOOD [@UFO #2]  MIALL, ROBERT  P
ALIEN TASTE [UKIAH OREGON 1]  SPENCER, WEN  C/P,P
BITTER WATERS [UKIAH OREGON 3]  SPENCER, WEN  C/P,P
DOG WARRIOR [UKIAH OREGON 4]  SPENCER, WEN  C/P:2,P
TAINTED TRAIL [UKIAH OREGON 2]  SPENCER, WEN  C/P,P
FORGE OF VIRTUE, THE [ULTIMA]  ABBEY, LYNN  C/P,P
TEMPER OF WISDOM, THE [ULTIMA]  ABBEY, LYNN  C/P,P
MACHINATIONS [@ULTIMA TECHNOCRAT WAR #1]  ANDREWS, AUSTEN  P
MASQUERADE [@ULTIMA TECHNOCRAT WAR #2]  ANDREWS, AUSTEN  P
MAELSTROM [@ULTIMA TECHNOCRAT WAR #3]  ANDREWS, AUSTEN  P
DREAMS OF STONE [@UNBALANCED EARTH #1]  WYLIE, JONATHAN  P
LIGHTLESS KINGDOM, THE [@UNBALANCED EARTH #2]  WYLIE, JONATHAN  C/P,P
AGE OF CHAOS, THE [@UNBALANCED EARTH #3]  WYLIE, JONATHAN  C/P,P
UNDERSEA CITY [UNDERSEA 3]  POHL, FREDERIK/WILLIAMSON, JACK  H:2,P
UNDERSEA FLEET [UNDERSEA 2]  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P
UNDERSEA QUEST [UNDERSEA 1]  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P
UNDERSEA TRILOGY, THE [UNDERSEA 1,2,3]  POHL, FREDERIK/WILLIAMSON, JACK  SFWA-P
BLOOD ENEMY [@UNDERWORLD 0]  COX, GREG  P
UNDERWORLD [@UNDERWORLD 1]  COX, GREG  P
UNDERWORLD EVOLUTION [@UNDERWORLD 2]  COX, GREG  P
DRAGONNE’S EG [UNEXPECTED DRAGON 3]  BROWN, MARY  P
MASTER OF MANY TREASURES [UNEXPECTED DRAGON 2]  BROWN, MARY  P
PIGS DON’T FLY [UNEXPECTED DRAGON 1]  BROWN, MARY  C/P:2,P
UNEXPECTED DRAGON, THE [UNEXPECTED DRAGON 1,2,3]  BROWN, MARY  H
UNLIKELY ONES, THE [UNEXPECTED DRAGON]  BROWN, MARY  H:2,SFWA-P,P
UNICORN AND DRAGON [@UNICORN AND DRAGON #1]  ABBEY, LYNN  H:2,LP
CONQUEST [@UNICORN AND DRAGON #2]  ABBEY, LYNN  LP
DOMINION’S REACH [UNIFIED WORLDS 3]  THORNLEY, DIANN  H,P
ECHOES OF ISSEL [UNIFIED WORLDS 2]  THORNLEY, DIANN  P
GANWOLD’S CHILD [UNIFIED WORLDS 1]  THORNLEY, DIANN  H,P
DOWNBELOW STATION [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
FINITY’S END [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
FORTY THOUSAND IN GEHENNA [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
MERCHANTER’S LUCK [UNION/ALLIANCE]  CHERRYH, C. J.  C/P,P,H
RIMRUNNERS [UNION/ALLIANCE]  CHERRYH, C. J.  C/P,P,H
TRIPOINT [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
CYTEEN [UNION/ALLIANCE CYTEEN 1,2,3]  CHERRYH, C. J.  H
CYTEEN: THE BETRAYAL [UNION/ALLIANCE CYTEEN 1]  CHERRYH, C. J.  C/P:2,P
CYTEEN: THE REBIRTH [UNION/ALLIANCE CYTEEN 2]  CHERRYH, C. J.  C/P:2,P
CYTEEN: THE VINDICATION [UNION/ALLIANCE CYTEEN 3]  CHERRYH, C. J.  C/P:2,P
HARD CRASH [UNREAL 1]  HUGHES, RYAN  P
PROPHET’S POWER [UNREAL 2]  SMITH, DEAN WESLEY  P
BRIGHTNESS REEF [UPLIFT 4]  BRIN, DAVID  H:2,P
EARTHCLAN [UPLIFT 2,3]  BRIN, DAVID  H:2
HEAVEN’S REACH [UPLIFT 6]  BRIN, DAVID  H:2,P
INFINITY’S SHORE [UPLIFT 5]  BRIN, DAVID  H:2,P
STARTIDE RISING [UPLIFT 2]  BRIN, DAVID  C/P:2,P
STARTIDE RISING (REVISED) [UPLIFT 2]  BRIN, DAVID  C/P,P,H
SUNDIVER [UPLIFT 1]  BRIN, DAVID  C/P:2,P
UPLIFT WAR, THE [UPLIFT 3]  BRIN, DAVID  C/P:2,P
KRUTON INTERFACE, THE [U.S.S. REPULSE 1]  DE CHANCIE, JOHN  C/P,P
HOPESHIP [USS SKIPJACK 3]  LANG, SIMON  P
TIMESLIDE [USS SKIPJACK 2]  LANG, SIMON  P
TRUMPETS OF TAGAN, THE [USS SKIPJACK 1]  LANG, SIMON  P

The MITSFS Seriesdex  September 24, 2007  Page 177
TITLE  AUTHOR  CODE
ALIEN SWORDMASTER, THE [V]  SUCHARITKUL, SOMTOW  C/P:2,SFWA-P,P
BELOW THE THRESHOLD [V]  WOLD, ALLEN L.  C/P,SFWA-P,P
CHICAGO CONVERSION, THE [V]  PROCTOR, GEO. W.  C/P,P
CRIVIT EXPERIMENT, THE [V]  WOLD, ALLEN L.  C/P,P
DEATH TIDE [V]  CRISPIN, A. C./MARSHALL, DEBORAH A.  C/P,P
EAST COAST CRISIS [V]  WEINSTEIN, HOWARD/CRISPIN, A. C.  C/P,P
FLORIDA PROJECT, THE [V]  SULLIVAN, TIM  C/P,P
NEW ENGLAND RESISTANCE, THE [V]  SULLIVAN, TIM  C/P,P
OREGON INVASION, THE [V]  TANNEHILL, JAYNE  C/P,SFWA-P,P
PATH TO CONQUEST [V]  WEINSTEIN, HOWARD  C/P,SFWA-P,P
PRISONERS AND PAWNS [V]  WEINSTEIN, HOWARD  C/P,SFWA-P,P
PURSUIT OF DIANA, THE [V]  WOLD, ALLEN L.  C/P,SFWA-P,P
SANCTION, THE [V]  TAGGERT, BRIAN  SR-S
SYMPHONY OF TERROR [V]  SUCHARITKUL, SOMTOW  SFWA-P
TEXAS RUN, THE [V]  PROCTOR, GEO. W.  C/P,P
V [V]  CRISPIN, A. C.  C/P:2,P
BRIGHTLY BURNING [VALDEMAR]  LACKEY, MERCEDES  H:2,P
CROSSROADS [VALDEMAR]  LACKEY, MERCEDES  PA
EXILE’S HONOR [VALDEMAR]  LACKEY, MERCEDES  H,P
EXILE’S VALOR [VALDEMAR]  LACKEY, MERCEDES  H:2
EXILE’S VALOR (PROOF) [VALDEMAR]  LACKEY, MERCEDES  SR-LP
SUN IN GLORY [VALDEMAR]  LACKEY, MERCEDES  PA
SWORD OF ICE [VALDEMAR]  LACKEY, MERCEDES  C/PA,PA
TAKE A THIEF [VALDEMAR]  LACKEY, MERCEDES  H:2
VALDEMAR COMPANION, THE [VALDEMAR]  HELFERS, JOHN/LITTLE, DENISE  R/REF-H
BRONZE KING, THE [VALENTINE MARSH 1]  CHARNAS, SUZY MCKEE  H,SFWA
GOLDEN THREAD, THE [VALENTINE MARSH 3]  CHARNAS, SUZY MCKEE  LP
SILVER GLOVE, THE [VALENTINE MARSH 2]  CHARNAS, SUZY MCKEE  H
WAY OF THE WOLF [@VAMPIRE EARTH #1]  KNIGHT, E. E.  P
CHOICE OF THE CAT [@VAMPIRE EARTH #2]  KNIGHT, E. E.  P
TALE OF THE THUNDERBOLT [@VAMPIRE EARTH #3]  KNIGHT, E. E.  P
VALENTINE’S RISING [@VAMPIRE EARTH #4]  KNIGHT, E. E.  P
VALENTINE’S EXILE [@VAMPIRE EARTH #5]  KNIGHT, E. E.  H
CHILL IN THE BLOOD, A [@VAMPIRE FILES 7]  ELROD, P. N.  H
DARK SLEEP, THE [VAMPIRE FILES 8]  ELROD, P. N.  H:2
LADY CRYMSYN [@VAMPIRE FILES 9]  ELROD, P. N.  H
SONG IN THE DARK [@VAMPIRE FILES 11]  ELROD, P. N.  H
BLOODLIST [@VAMPIRE FILES #1]  ELROD, P. N.  C/P:2,P
LIFEBLOOD [@VAMPIRE FILES #2]  ELROD, P. N.  C/P:2,P
BLOODCIRCLE [@VAMPIRE FILES #3]  ELROD, P. N.  C/P,P
ART IN THE BLOOD [@VAMPIRE FILES #4]  ELROD, P. N.  C/P,P
FIRE IN THE BLOOD [@VAMPIRE FILES #5]  ELROD, P. N.  C/P:2,P
BLOOD ON THE WATER [@VAMPIRE FILES #6]  ELROD, P. N.  C/P:2,P
HUNTED, THE [VAMPIRE HUNTRESS 3]  BANKS, L. A.  LP
MINION [VAMPIRE HUNTRESS 1]  BANKS, L. A.  LP
MINION (SPECIAL HUNTRESS EDITION) [VAMPIRE HUNTRESS]  BANKS, L. A.  P
DANCE WITH A VAMPIRE (PROOF) [@VAMPIRE KISSES #4]  SHCREIBER, ELLEN  SR-LP
BITTER BLOOD [VAMPIRE LEGACY 2]  TAYLOR, KAREN E.  P
BLOOD SECRETS [VAMPIRE LEGACY 1]  TAYLOR, KAREN E.  P
RESURRECTION [VAMPIRE LEGACY]  TAYLOR, KAREN E.  C/P,P
BLOODSTALK [@VAMPIRELLA #1]  GOULART, RON  C/P,P
ON ALIEN WINGS [@VAMPIRELLA #2]  GOULART, RON  C/P,P
DEADWALK [@VAMPIRELLA #3]  GOULART, RON  P
BLOOD WEDDING [@VAMPIRELLA #4]  GOULART, RON  P
DEATHGAME [@VAMPIRELLA #5]  GOULART, RON  P
LADY OF THE SERPENTS, THE [@VAMPYRICON 2]  CLEGG, DOUGLAS  H

MAZIRIAN THE MAGICIAN [@VANCE INTEGRAL EDITION #1/ DYING EARTH 1] VANCE, JACK

SR-H

WORLD THINKER AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #2] VANCE, JACK

SR-H

GADGET STORIES [@VANCE INTEGRAL EDITION #3]  VANCE, JACK  SR-H

RAPPAREE, THE [@VANCE INTEGRAL EDITION #4]/ BIG PLANET [@VANCE INTEGRAL EDITION #4]/ VANDALS OF THE VOID [@VANCE INTEGRAL EDITION #4] VANCE, JACK

SR-H

SON OF THE TREE AND OTHER STORIES [@VANCE INTEGRAL EDITION #5]  VANCE, JACK  SR-H
GOLDEN GIRL AND OTHER STORIES [@VANCE INTEGRAL EDITION #6]  VANCE, JACK  SR-H

GOLD AND IRON [@VANCE INTEGRAL EDITION #7]/ CLARGES [@VANCE INTEGRAL EDITION #7]/ LANGUAGES OF PAO, THE [@VANCE INTEGRAL EDITION #7] VANCE, JACK

SR-H

HOUSES OF ISZM AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #8] VANCE, JACK SR-H

The MITSFS Seriesdex  September 24, 2007  Page 178
TITLE  AUTHOR  CODE

MIRACLE WORKERS, THE [@VANCE INTEGRAL EDITION #9]/ DRAGON MASTERS, THE [@VANCE INTEGRAL EDITION #9]/ LAST CASTLE, THE [@VANCE INTEGRAL EDITION #9] VANCE, JACK

SR-H

FLESH MASK, THE [@VANCE INTEGRAL EDITION #10]/ STRANGE PEOPLE, QUEER NOTIONS [@VANCE INTEGRAL EDITION #10]/ BIRD ISLAND [@VANCE INTEGRAL EDITION #10] VANCE, JACK

SR-H

HOUSE ON LILY STREET, THE [@VANCE INTEGRAL EDITION #11]/ VIEW FROM CHICKWEE’S WINDOW, THE [@VANCE INTEGRAL EDITION #11] VANCE, JACK

SR-H

BAD RONALD [@VANCE INTEGRAL EDITION #12]/ DARK OCEAN, THE [@VANCE INTEGRAL EDITION #12] VANCE, JACK

SR-H

FOX VALLEY MURDERS, THE [@VANCE INTEGRAL EDITION #13]/ PLEASANT GROVE MURDERS, THE [@VANCE INTEGRAL EDITION #13] VANCE, JACK

SR-H

MAN IN THE CAGE, THE [@VANCE INTEGRAL EDITION #14]/ DEADLY ISLES, THE [@VANCE INTEGRAL EDITION #14] VANCE, JACK

SR-H

CUGEL THE CLEVER [@VANCE INTEGRAL EDITION #15]  VANCE, JACK  SR-H
BLUE WORLD (VIE), THE [@VANCE INTEGRAL EDITION #16]  VANCE, JACK  SR-H
MOON MOTH AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #17]  VANCE, JACK  SR-H
SPACE OPERA (VIE) [@VANCE INTEGRAL EDITION #18]  VANCE, JACK  SR-H

MAGNIFICENT SHOWBOATS OF THE LOWER VISSEL RIVER, LUNE XXIII SOUTH, BIG PLANET, THE [@VANCE INTEGRAL EDITION #19] VANCE, JACK

SR-H

EMPHYRIO (VIE) [@VANCE INTEGRAL EDITION #20]  VANCE, JACK  SR-H
TSCHAI [@VANCE INTEGRAL EDITION #21]  VANCE, JACK  SR-H

STAR KING (VIE), THE [@VANCE INTEGRAL EDITION #22/ DEMON PRINCES 1] VANCE, JACK

SR-H

KILLING MACHINE (VIE), THE [@VANCE INTEGRAL EDITION #23/ DEMON PRINCES 2] VANCE, JACK

SR-H

PALACE OF LOVE (VIE), THE [@VANCE INTEGRAL EDITION #24/ DEMON PRINCES 3] VANCE, JACK

SR-H

FACE, THE (VIE) [@VANCE INTEGRAL EDITION #25/DEMON PRINCES 4]  VANCE, JACK  SR-H

BOOK OF DREAMS, THE (VIE) [@VANCE INTEGRAL EDITION #26/ DEMON PRINCES 5] VANCE, JACK

SR-H

DURDANE [@VANCE INTEGRAL EDITION #27]  VANCE, JACK  SR-H
DOMAINS OF KORYPHON, THE [@VANCE INTEGRAL EDITION #28]  VANCE, JACK  SR-H

TRULLION: ALASTOR 2262 (VIE) [@VANCE INTEGRAL EDITION #29/ ALASTOR 1] VANCE, JACK

SR-H

MARUNE: ALASTOR 933 (VIE) [@VANCE INTEGRAL EDITION #30/ ALASTOR 2] VANCE, JACK

SR-H

WYST: ALASTOR 1716 (VIE) [@VANCE INTEGRAL EDITION #31/ALASTOR 3]  VANCE, JACK  SR-H

DOGTOWN TOURIST AGENCY, THE [@VANCE INTEGRAL EDITION #32]/ FREITZKE’S TURN [@VANCE INTEGRAL EDITION #32] VANCE, JACK

SR-H

MASKE: THAERY (VIE) [@VANCE INTEGRAL EDITION #33]  VANCE, JACK  SR-H

RHIALTO THE MARVELLOUS (VIE) [@VANCE INTEGRAL EDITION #34/ DYING EARTH 4] VANCE, JACK

SR-H

CUGEL: THE SKYBREAK [@VANCE INTEGRAL EDITION #35]/ SPATTERLIGHT [@VANCE INTEGRAL EDITION #35] VANCE, JACK

SR-H

SULDRUN’S GARDEN (VIE) [@VANCE INTEGRAL EDITION #36/LYONESSE 1]  VANCE, JACK  SR-H
GREEN PEARL, THE (VIE) [@VANCE INTEGRAL EDITION #37/LYONESSE 2]  VANCE, JACK  SR-H
MADOUC (VIE) [@VANCE INTEGRAL EDITION #38/LYONESSE 3]  VANCE, JACK  SR-H
ARAMINTA STATION (VIE) [@VANCE INTEGRAL EDITION #39/CADWAL 1]  VANCE, JACK  SR-H
ECCE AND OLD EARTH (VIE) [@VANCE INTEGRAL EDITION #40/CADWAL 2]  VANCE, JACK  SR-H
THROY (VIE) [@VANCE INTEGRAL EDITION #41/CADWAL 3]  VANCE, JACK  SR-H
NIGHT LAMP (VIE) [@VANCE INTEGRAL EDITION #42]  VANCE, JACK  SR-H
PORTS OF CALL (VIE) [@VANCE INTEGRAL EDITION #43/MYRON TANY 1]  VANCE, JACK  SR-H

WILD THYME AND VIOLETS, OTHER UNPUBLISHED WORKS, AND ADDENDA [@VANCE INTEGRAL EDITION #44] VANCE, JACK

SR-H

VANG: THE BATTLEMASTER, THE [VANG]  ROWLEY, CHRISTOPHER  P
VANG: THE MILITARY FORM, THE [VANG]  ROWLEY, CHRISTOPHER  C/P,P
SON OF THE HERO [@VARAYAN #1]  SHELLEY, RICK  P
HERO OF VARAY, THE [@VARAYAN #2]  SHELLEY, RICK  P
HERO KING, THE [@VARAYAN #3]  SHELLEY, RICK  P
SKEWED THRONE, THE [VARIS 1]  PALMATIER, JOSHUA  H
VARNEY THE VAMPYRE I [VARNEY #1]  RYMER, JAMES  LP
VARNEY THE VAMPYRE II [VARNEY #2]  RYMER, JAMES  LP
BRIDE OF THE VAMPYRE [@VARNEY #4]  RYMER, JAMES  LP
FATE OF THE VAMPYRE, THE [@VARNEY THE VAMPIRE #5]  RYMER, JAMES  LP
COMMAND DECISION [VATTA’S WAR 4]  MOON, ELIZABETH  H:2
ENGAGING THE ENEMY [VATTA’S WAR 3]  MOON, ELIZABETH  H:2
MARQUE AND REPRISAL [VATTA’S WAR 2]  MOON, ELIZABETH  H,P
TRADING IN DANGER [VATTA’S WAR 1]  MOON, ELIZABETH  H
MYTH HUNTERS, THE [VEIL 1]  GOLDEN, CHRISTOPHER  LP
EROS ASCENDING [VELVET COMET #1]  RESNICK, MIKE  C/P,@P,H
EROS AT ZENITH [VELVET COMET #2]  RESNICK, MIKE  C/P:2,@P,H

TITLE  AUTHOR  CODE
EROS DESCENDING [VELVET COMET #3]  RESNICK, MIKE  C/P,@P
EROS AT NADIR [VELVET COMET #4]  RESNICK, MIKE  C/P,@P
SECRET BOOKS OF VENUS I & II, THE [VENUS 1,2]  LEE, TANITH  H
SECRET BOOKS OF VENUS III & IV, THE [VENUS 3,4]  LEE, TANITH  H
PIRATES OF VENUS [VENUS #1]  BURROUGHS, EDGAR RICE  C/P:2,@P
LOST ON VENUS [VENUS #2]  BURROUGHS, EDGAR RICE  C/P:2,@P
CARSON OF VENUS [VENUS #3]  BURROUGHS, EDGAR RICE  C/P:2,@P
ESCAPE ON VENUS [VENUS #4]  BURROUGHS, EDGAR RICE  C/P:2,@P
VENUS PRESERVED [@VENUS #4]  LEE, TANITH  LP
WIZARD OF VENUS, THE [VENUS #5]  BURROUGHS, EDGAR RICE  @P
CHILD OF VENUS [VENUS OF DREAMS 3]  SARGENT, PAMELA  P
VENUS OF DREAMS [VENUS OF DREAMS 1]  SARGENT, PAMELA  C/P:2,P
VENUS OF SHADOWS [VENUS OF DREAMS 2]  SARGENT, PAMELA  C/P,P,H
BREAKING STRAIN [@VENUS PRIME #1]  PREUSS, PAUL  C/P:2,P
MAELSTROM [@VENUS PRIME #2]  PREUSS, PAUL  C/P:2,P
HIDE AND SEEK [@VENUS PRIME #3]  PREUSS, PAUL  C/P:2,P
MEDUSA ENCOUNTER, THE [@VENUS PRIME #4]  PREUSS, PAUL  C/P:3,P
DIAMOND MOON, THE [@VENUS PRIME #5]  PREUSS, PAUL  C/P:2,P
SHINING ONES, THE [@VENUS PRIME #6]  PREUSS, PAUL  C/P,P,H
CAIN’S LAND [VERESHCHAGIN 3]  FREZZA, ROBERT  P
FIRE IN A FARAWAY PLACE [VERESHCHAGIN 2]  FREZZA, ROBERT  P
SMALL COLONIAL WAR, A [VERESHCHAGIN 1]  FREZZA, ROBERT  C/P:2,P
DARKLOST [VICTOR RENQUIST 1]  FARREN, MICK  H
MORE THAN MORTAL [VICTOR RENQUIST 2]  FARREN, MICK  H
BLOOD DEBT [VICTORY NELSON 5]  HUFF, TANYA  C/P:2,P
BLOOD LINES [VICTORY NELSON 3]  HUFF, TANYA  C/P:2,P
BLOOD PACT [VICTORY NELSON 4]  HUFF, TANYA  C/P:2,P
BLOOD PRICE [VICTORY NELSON 1]  HUFF, TANYA  C/P:2,P
BLOOD TRAIL [VICTORY NELSON 2]  HUFF, TANYA  C/P:2,P
BRIDGE OF THE SEPARATOR [VIDESSOS]  TURTLEDOVE, HARRY  H
EMPEROR FOR THE LEGION, AN [VIDESSOS 2]  TURTLEDOVE, HARRY  C/P:2,P
LEGION OF VIDESSOS, THE [VIDESSOS 3]  TURTLEDOVE, HARRY  C/P:2,P
MISPLACED LEGION, THE [VIDESSOS 1]  TURTLEDOVE, HARRY  C/P:2,P
SWORDS OF THE LEGION [VIDESSOS 4]  TURTLEDOVE, HARRY  C/P:2,P
VIDESSOS BESIEGED [VIDESSOS 5]  TURTLEDOVE, HARRY  P
SHADOW ON THE GLASS, A [@VIEW FROM THE MIRROR #1]  IRVINE, IAN  C/P:2,P
TOWER ON THE RIFT, THE [@VIEW FROM THE MIRROR #2]  IRVINE, IAN  C/P,P
DARK IS THE MOON [@VIEW FROM THE MIRROR #3]  IRVINE, IAN  P
WAY BETWEEN THE WORLDS, THE [@VIEW FROM THE MIRROR #4]  IRVINE, IAN  P
IN ANOTHER COUNTRY [VINTAGE SEASON 2]  SILVERBERG, ROBERT  SFWA-TD18,TD18
VINTAGE SEASON [VINTAGE SEASON 1]  MOORE, C. L.  SFWA-TD18,TD18
THROUGH VIOLET EYES [VIOLETS 1]  WOODWORTH, STEPHEN  H
WITH RED HANDS [VIOLETS 2]  WOODWORTH, STEPHEN  H
SUN OF SUNS [VIRGA 1]  SCHROEDER, KARL  H
FLOATING GODS, THE [VIRICONIUM]  HARRISON, M. JOHN  P
PASTEL CITY, THE [VIRICONIUM 1]  HARRISON, M. JOHN  H:2,SFWA,P
STORM OF WINGS, A [VIRICONIUM 2]  HARRISON, M. JOHN  C/P,SFWA,P
STORM OF WINGS, A (BRITISH) [VIRICONIUM 2]  HARRISON, M. JOHN  SR-P
VIRICONIUM [VIRICONIUM]  HARRISON, M. JOHN  LP
VIRICONIUM NIGHTS [VIRICONIUM]  HARRISON, M. JOHN  P
CHI [VIRTUAL REALITY 3]  BESHER, ALEXANDER  H
MIR [VIRTUAL REALITY 2]  BESHER, ALEXANDER  H
RIM [VIRTUAL REALITY 1]  BESHER, ALEXANDER  P
NO LIMITS [@VIRTUAL WORLD]  FINDLEY, NIGEL D.  P
ANACKIRE [VIS 2]  LEE, TANITH  C/P:2,P
STORM LORD, THE [VIS 1]  LEE, TANITH  C/P,P
WARS OF VIS, THE [VIS 1,2]  LEE, TANITH  H
CREATION DESCENDING [@VISION OF BEASTS #1]  LOVEJOY, JACK  C/P,P
SECOND KINGDOM, THE [@VISION OF BEASTS #2]  LOVEJOY, JACK  C/P,SFWA-P,P
BROTHERHOOD OF DIABLO, THE [@VISION OF BEASTS #3]  LOVEJOY, JACK  P,SFWA-P
ANGEL FIRE EAST [VOID & WORD 3]  BROOKS, TERRY  H
KNIGHT OF THE WORD, A [VOID & WORD 2]  BROOKS, TERRY  H,P
RUNNING WITH THE DEMON [VOID & WORD 1]  BROOKS, TERRY  H,P
BROTHEL IN ROSENSTRASSE, THE [VON BEK]  MOORCOCK, MICHAEL  LP
CITY IN THE AUTUMN STARS, THE [VON BEK 2]  MOORCOCK, MICHAEL  C/P,SFWA,SFWA-P,P
VON BEK (EXPANDED OMNIBUS EDITION) [VON BEK]  MOORCOCK, MICHAEL  LP
VON BEK (OMNIBUS EDITION) [VON BEK]  MOORCOCK, MICHAEL  P
WAR HOUND AND THE WORLD’S PAIN, THE [VON BEK 1]  MOORCOCK, MICHAEL  C/P,P,H
HELL HEART [@VOR THE MAELSTROM]  VARDEMAN, ROBERT E.  P
OPERATION SIERRA-75 [@VOR THE MAELSTROM]  GRESSMAN, THOMAS S.  P
OATHBLOOD [VOWS AND HONOR 3]  LACKEY, MERCEDES  C/P:2,P
OATHBOUND, THE [VOWS AND HONOR 1]  LACKEY, MERCEDES  C/P:2,P
OATHBREAKERS [VOWS AND HONOR 2]  LACKEY, MERCEDES  C/P:2,P

TITLE  AUTHOR  CODE
ANTRAX [VOYAGE OF THE JERLE SHANNARA 2/SHANNARA JERLE]  BROOKS, TERRY  H,P
MORGAWR [@VOYAGE OF THE JERLE SHANNARA 3/SHANNARA JERLE]  BROOKS, TERRY  H
ILSE WITCH [@VOYAGE OF THE JERLE SHANNARA #1/SHANNARA JERLE]  BROOKS, TERRY  H
VOYAGERS [@VOYAGERS #1]  BOVA, BEN  C/P:2,SFWA-P,P,H
ALIEN WITHIN, THE [@VOYAGERS #2]  BOVA, BEN  C/P,SFWA-P,P,H:2
STAR BROTHERS [@VOYAGERS #3]  BOVA, BEN  C/P:2,SFWA-P,P,H

The MITSFS Seriesdex  September 24, 2007  Page 181
TITLE  AUTHOR  CODE
THUNDERBIRD FALLS [WALKER 2]  MURPHY, C. E.  H:2,LP
URBAN SHAMAN [WALKER 1]  MURPHY, C. E.  LP
COYOTE DREAMS [@WALKER #3]  MURPHY, C. E.  LP
WANDOR’S FLIGHT [WANDOR 4]  GREEN, ROLAND J.  P
WANDOR’S JOURNEY [WANDOR 2]  GREEN, ROLAND J.  P
WANDOR’S RIDE [WANDOR 1]  GREEN, ROLAND J.  P
WANDOR’S VOYAGE [WANDOR 3]  GREEN, ROLAND J.  C/P,P
DAY FOR DAMNATION, A [WAR AGAINST THE CHTORR 2]  GERROLD, DAVID  C/P:2,P
MATTER FOR MEN, A [WAR AGAINST THE CHTORR 1]  GERROLD, DAVID  C/P,P
RAGE FOR REVENGE, A [WAR AGAINST THE CHTORR 3]  GERROLD, DAVID  C/P:2,P
SEASON FOR SLAUGHTER, A [WAR AGAINST THE CHTORR 4]  GERROLD, DAVID  C/P,P

DAY FOR DAMNATION, A (ABRIDGED)

[WAR AGAINST THE CHTORR (ABRIDGED) 2] GERROLD, DAVID

C/P,LP

MATTER FOR MEN, A (ABRIDGED)

[WAR AGAINST THE CHTORR (ABRIDGED) 1] GERROLD, DAVID

C/P,LP

WAR AGAINST THE CHTORR: INVASION, THE [WAR AGAINST THE CHTORR (ABRIDGED) 1,2] GERROLD, DAVID

H:2

ADVANCE AND RETREAT [WAR BETWEEN THE PROVINCES 3]  TURTLEDOVE, HARRY  H
MARCHING THROUGH PEACHTREE [WAR BETWEEN THE PROVINCES 2]  TURTLEDOVE, HARRY  H
SENTRY PEAK [WAR BETWEEN THE PROVINCES 1]  TURTLEDOVE, HARRY  H,P
CITY IN THE GLACIER, THE [WAR OF POWERS 2]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P:2,P
DEMON OF THE DARK ONES [WAR OF POWERS 6]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P:2,P
DESTINY STONE, THE [WAR OF POWERS 3]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P,P
FALLEN ONES, THE [WAR OF POWERS 4]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P,P
IN THE SHADOW OF OMIZANTRIM [WAR OF POWERS 5]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P,P
SUNDERED REALM, THE [WAR OF POWERS 1]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P:2,P
MISTS OF EVERNESS [WAR OF THE DREAMING 2]  WRIGHT, JOHN C.  H
IRON LORDS, THE [WAR OF THE GODS ON EARTH 1]  OFFUTT, ANDREW J.  C/P:2,P
LADY OF THE SNOWMIST, THE [WAR OF THE GODS ON EARTH 3]  OFFUTT, ANDREW J.  C/P:2,P
SHADOWS OUT OF HELL [WAR OF THE GODS ON EARTH 2]  OFFUTT, ANDREW J.  C/P,P
SERPENT AND THE ROSE, THE [WAR OF THE ROSE 1]  BRYAN, KATHLEEN  LP
DEMON IN THE MIRROR [WAR OF THE WIZARDS 1]  OFFUTT, ANDREW J./LYON, RICHARD K.  C/P,P
EYES OF SARSIS, THE [WAR OF THE WIZARDS 2]  OFFUTT, ANDREW J./LYON, RICHARD K.  P
WEB OF THE SPIDER [WAR OF THE WIZARDS 3]  OFFUTT, ANDREW J./LYON, RICHARD K.  P
RESURRECTION, THE [WAR OF THE WORLDS 1]  DILLARD, J. M.  C/P,P
CYBORG AND THE SORCERERS, THE [WAR SURPLUS 1]  WATT-EVANS, LAWRENCE  C/P:2,P
WIZARD AND THE WAR MACHINE, THE [WAR SURPLUS 2]  WATT-EVANS, LAWRENCE  C/P,P
ACACIA [WAR WITH THE MEIN 1]  DURHAM, DAVID ANTHONY  H
ACACIA (PROOF) [WAR WITH THE MEIN 1]  DURHAM, DAVID ANTHONY  SR-LP

BLOOD FEUDS [@WAR WORLD] STIRLING, S. M./TARR, JUDITH/

SHWARTZ, SUSAN/TURTLEDOVE, HARRY

C/P,SFWA-P,P

BLOOD VENGEANCE [@WAR WORLD] STIRLING, S. M./TARR, JUDITH/

SHWARTZ, SUSAN/TURTLEDOVE, HARRY

P,SFWA-P

BURNING EYE, THE [@WAR WORLD #1]  POURNELLE, JERRY  PA,C/PA,SFWA-P
DEATH’S HEAD REBELLION [@WAR WORLD #2]  POURNELLE, JERRY  PA,C/PA
SAURON DOMINION [@WAR WORLD #3]  POURNELLE, JERRY  SFWA-P,C/PA,PA
INVASION [@WAR WORLD #4]  POURNELLE, JERRY  PA
FAR STARS WAR, THE [@WAR YEARS #1]  FAWCETT, BILL  PA
SIEGE OF ARISTA, THE [@WAR YEARS #2]  FAWCETT, BILL  C/PA,PA
JUPITER WAR, THE [@WAR YEARS #3]  FAWCETT, BILL  PA
WARBOTS [@WARBOTS #1]  STINE, G. HARRY  C/P,P
OPERATION STEEL BAND [@WARBOTS #2]  STINE, G. HARRY  C/P,P
BASTAARD REBELLION, THE [@WARBOTS #3]  STINE, G. HARRY  C/P:2,P
SIERRA MADRE [@WARBOTS #4]  STINE, G. HARRY  P
OPERATION IRON FIST [@WARBOTS #7]  STINE, G. HARRY  P
FORCE OF ARMS [@WARBOTS #8]  STINE, G. HARRY  P
GOBLIN MARKET [WARCHILD 2]  BOWES, RICHARD  P
WARCHILD [WARCHILD 1]  BOWES, RICHARD  C/P,P

WARCRAFT ARCHIVE, THE [WARCRAFT] KNAAK, RICHARD A./GOLDEN, CHRISTIE/

GRUBB, JEFF/METZEN, CHRIS

H

DAY OF THE DRAGON [@WARCRAFT #1]  KNAAK, RICHARD A.  P
LORD OF THE CLANS [@WARCRAFT #2]  GOLDEN, CHRISTIE  P
LAST GUARDIAN, THE [@WARCRAFT #3]  GRUBB, JEFF  P
DRAGON HUNT [@WARCRAFT SUNWELL 1]  KNAAK, RICHARD A.  SCX
SHADOWS OF ICE [@WARCRAFT SUNWELL 2]  KNAAK, RICHARD A.  SCX
WELL OF ETERNITY, THE [@WARCRAFT WAR OF THE ANCIENTS #1]  KNAAK, RICHARD A.  P
DEMON SOUL, THE [@WARCRAFT WAR OF THE ANCIENTS #2]  KNAAK, RICHARD A.  P
SUNDERING, THE [@WARCRAFT WAR OF THE ANCIENTS #3]  KNAAK, RICHARD A.  P
DRAGON BLOOD [WARD OF HUROG 2]  BRIGGS, PATRICIA  C/P,P
DRAGON BONES [WARD OF HUROG 1]  BRIGGS, PATRICIA  C/P,P
FREEWARE [WARE 3]  RUCKER, RUDY  LP
MOLDIES & MEATBOPS [WARE 1,2,3]  RUCKER, RUDY  H
REALWARE [WARE 4]  RUCKER, RUDY  LP

TITLE  AUTHOR  CODE
SOFTWARE [WARE 1]  RUCKER, RUDY  C/P:2,P
WETWARE [WARE 2]  RUCKER, RUDY  C/P:2,P
BURNING SHORE, THE [@WARHAMMER]  EARL, ROBERT  P

CALL OF CHAOS, THE [WARHAMMER] GASCOIGNE, MARC/JONES, ANDY/

DUNN, CHRISTIAN

CX

COLD HAND OF BETRAYAL, THE [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  P
CORRUPTED, THE [@WARHAMMER]  EARL, ROBERT  P
DEAD AND THE DAMNED, THE [WARHAMMER]  GREEN, JONATHAN  P
DOMINION [WARHAMMER]  SAVILE, STEVEN  P
ENEMY WITHIN, THE [@WARHAMMER]  BYERS, RICHARD LEE  P
FELL CARGO [@WARHAMMER]  ABNETT, DAN  P
FORGED IN BATTLE [@WARHAMMER]  HUNTER, JUSTIN  P
GRUDGE BEARER [@WARHAMMER]  THORPE, GAV  P
GUARDIANS OF THE FOREST [@WARHAMMER]  MCNEILL, GRAHAM  P
HAMMERS OF ULRIC [WARHAMMER]  ABNETT, DAN/VINCENT, NIK/WALLIS, JAMES P
IGNORANT ARMIES [WARHAMMER]  PRINGLE, DAVID  SR-LP
INVASION! [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
KAL JERICO [WARHAMMER]  RENNIE, GORDON/KOPINSKI, KARL  CX
LAUGHTER OF DARK GODS, THE [WARHAMMER]  PRINGLE, DAVID  PA
LORDS OF VALOUR [WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
MAGESTORM [@WARHAMMER]  GREEN, JONATHAN  P
MARK OF HERESY [@WARHAMMER]  WALLIS, JAMES  P
MURDER IN MARIENBURG, A [@WARHAMMER]  BISHOP, DAVID  P
NECROMANCER [@WARHAMMER]  GREEN, JONATHAN  P
REALM OF CHAOS [WARHAMMER]  GASCOIGNE, MARC/JONES, ANDY  PA
RED THIRST [WARHAMMER]  PRINGLE, DAVID  LP
RIDERS OF THE DEAD [@WARHAMMER]  ABNETT, DAN  H
SAVAGE CITY [@WARHAMMER]  EARL, ROBERT  P
STAR OF ERENGRAD [@WARHAMMER]  MCINTOSH, NEIL  P
SWORDS OF THE EMPIRE [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
TALES FROM THE TEN-TAILED CAT [WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  CX
TALES OF THE OLD WORLD [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  LPA
VERMINTIDE [@WARHAMMER]  LEE, BRUNO  P
WAY OF THE DEAD [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
WILD KINGDOMS [@WARHAMMER]  EARL, ROBERT  C/P,P
WINE OF DREAMS, THE [WARHAMMER]  CRAIG, BRIAN  P
WITCH HUNTER [@WARHAMMER]  WERNER, C. L.  P
WITCH KILLER [WARHAMMER]  WERNER, C. L.  P
WOLF RIDERS [WARHAMMER]  PRINGLE, DAVID  SR-LP
HONOR OF THE GRAVE [@WARHAMMER ANGELIKA FLEISCHER 1]  LAWS, ROBIN D.  P
LIAR’S PEAK [@WARHAMMER ANGELIKA FLEISCHER 3]  LAWS, ROBIN D.  P
SACRED FLESH [@WARHAMMER ANGELIKA FLEISCHER 2]  LAWS, ROBIN D.  P
BROKEN LANCE, THE [@WARHAMMER BLACK HEARTS 2]  LONG, NATHAN  P
VALNIR’S BANE [@WARHAMMER BLACK HEARTS 1]  LONG, NATHAN  P
DEATH’S CITY [@WARHAMMER BLOOD ON THE REIK 2]  MITCHELL, SANDY  P
DEATH’S MESSENGER [@WARHAMMER BLOOD ON THE REIK 1]  MITCHELL, SANDY  C/P,P
IN THE EYE OF TERROR [WARHAMMER BLOODQUEST 2]  RENNIE, GORDON/MACNEIL, COLIN  CX
BLOOD & STEEL [@WARHAMMER BRUNNER 2]  WERNER, C. L.  P
BLOOD OF THE DRAGON [@WARHAMMER BRUNNER 3]  WERNER, C. L.  P
BORN OF BLOOD [WARHAMMER DARKBLADE #1]  ABNETT, DAN  CX
BEASTS IN VELVET [WARHAMMER GENEVIEVE 3]  YEOVIL, JACK  P
DRACHENFELS [WARHAMMER GENEVIEVE 1]  YEOVIL, JACK  P
GENEVIEVE UNDEAD [WARHAMMER GENEVIEVE 2]  YEOVIL, JACK  P
SILVER NAILS [WARHAMMER GENEVIEVE 4]  YEOVIL, JACK  P
GILEAD’S BLOOD [WARHAMMER GILEAD]  ABNETT, DAN/VINCENT, NIK  P
BEASTSLAYER [WARHAMMER GOTREK & FELIX 5]  KING, WILLIAM  P
DAEMONSLAYER [WARHAMMER GOTREK & FELIX 3]  KING, WILLIAM  P
DRAGONSLAYER [WARHAMMER GOTREK & FELIX 4]  KING, WILLIAM  P
GIANTSLAYER [WARHAMMER GOTREK & FELIX 7]  KING, WILLIAM  P

GOTREK & FELIX: THE FIRST OMNIBUS [@WARHAMMER GOTREK & FELIX 1,2,3] KING, WILLIAM

H

GOTREK & FELIX: THE SECOND OMNIBUS [@WARHAMMER GOTREK & FELIX 4,5,6] KING, WILLIAM

H

SKAVENSLAYER [WARHAMMER GOTREK & FELIX 2]  KING, WILLIAM  P
TROLLSLAYER [WARHAMMER GOTREK & FELIX 1]  KING, WILLIAM  P
VAMPIRESLAYER [WARHAMMER GOTREK & FELIX 6]  KING, WILLIAM  P
CHAOS CHILD [WARHAMMER INQUISITION WAR 3]  WATSON, IAN  P
DRACO [WARHAMMER INQUISITION WAR 1]  WATSON, IAN  P
HARLEQUIN [WARHAMMER INQUISITION WAR 2]  WATSON, IAN  P
INQUISITION WAR, THE [@WARHAMMER INQUISITION WAR 1,2,3]  WATSON, IAN  LP
HUNT FOR DEFAY, THE [WARHAMMER INQUISITOR ASCENDANT]  ABNETT, DAN  CX
KONRAD [WARHAMMER KONRAD 1]  FERRING, DAVID  C/P,LP
KONRAD SAGA, THE [@WARHAMMER KONRAD 1,2,3]  FERRING, DAVID  H

TITLE  AUTHOR  CODE
SHADOWBREED [WARHAMMER KONRAD 2]  FERRING, DAVID  LP
WARBLADE [WARHAMMER KONRAD 3]  FERRING, DAVID  P
BLOODSTORM [@WARHAMMER MALUS DARKBLADE 2]  ABNETT, DAN/LEE, MIKE  P
DAEMON’S CURSE, THE [@WARHAMMER MALUS DARKBLADE 1]  ABNETT, DAN/LEE, MIKE  P
REAPER OF SOULS [@WARHAMMER MALUS DARKBLADE]  ABNETT, DAN/LEE, MIKE  P
WARPSWORD [@WARHAMMER MALUS DARKBLADE]  ABNETT, DAN/LEE, MIKE  P
MARK OF CHAOS [@WARHAMMER MARK OF CHAOS]  REYNOLDS, ANTHONY  P
MARK OF DAMNATION [@WARHAMMER MARK OF CHAOS]  WALLIS, JAMES  P
PLAGUE DAEMON [WARHAMMER ORFEO 2]  CRAIG, BRIAN  LP
STORM WARRIORS [WARHAMMER ORFEO 3]  CRAIG, BRIAN  H,P
ZARAGOZ [WARHAMMER ORFEO 1]  CRAIG, BRIAN  LP
BLADES OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 2]  THORPE, GAV  P
CLAWS OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 1]  THORPE, GAV  P
HEART OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 3]  THORPE, GAV  P
KEEPERS OF THE FLAME [@WARHAMMER STEFAN KUMANSKY 3]  MCINTOSH, NEIL  P
TAINT OF EVIL [@WARHAMMER STEFAN KUMANSKY 2]  MCINTOSH, NEIL  P
RETRIBUTION [@WARHAMMER VON CARSTEIN 3]  SAVILE, STEVEN  P
AMBASSADOR, THE [@WARHAMMER VON VELTEN 1]  MCNEILL, GRAHAM  P
URSUN’S TEETH [@WARHAMMER VON VELTEN 2]  MCNEILL, GRAHAM  P
WITCH FINDER [@WARHAMMER WITCH HUNTER 1]  WERNER, C. L.  P
ZAVANT [WARHAMMER ZAVANT KONNIGER]  RENNIE, GORDON  P
ANGELS OF DARKNESS [@WARHAMMER 40K]  THORPE, GAV  P
ARMOUR OF CONTEMPT, THE [@WARHAMMER 40K]  ABNETT, DAN  H
BRINGERS OF DEATH [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
BROTHERS OF THE SNAKE [@WARHAMMER 40K]  ABNETT, DAN  H
CHAPTER WAR [@WARHAMMER 40K]  COUNTER, BEN  P
CRUCIBLE OF WAR [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
DAEMON WORLD [@WARHAMMER 40K]  COUNTER, BEN  P
DARK IMPERIUM [WARHAMMER 40K]  GASCOIGNE, MARC/JONES, ANDY  PA
DEATH WORLD [@WARHAMMER 40K]  LYONS, STEVE  P
DEATHWING [WARHAMMER 40K]  JONES, NEIL/PRINGLE, DAVID  PA
DOUBLE EAGLE [@WARHAMMER 40K]  ABNETT, DAN  P
DUTY CALLS [@WARHAMMER 40K]  MITCHELL, SANDY  P
ELDAR PROPHECY [@WARHAMMER 40K]  GOTO, C. S.  P
EXECUTION HOUR [WARHAMMER 40K]  RENNIE, GORDON  P
EYE OF TERROR [WARHAMMER 40K]  BAYLEY, BARRINGTON J.  P
FIFTEEN HOURS [@WARHAMMER 40K]  SCANLON, MITCHEL  P
FIRE WARRIOR [@WARHAMMER 40K]  SPURRIER, SIMON  P
INTO THE MAELSTROM [WARHAMMER 40K]  GASCOIGNE, MARC/JONES, ANDY  PA
IRON HANDS [@WARHAMMER 40K]  GREEN, JONATHAN  P
LAST CHANCERS, THE [@WARHAMMER 40K]  THORPE, GAV  LP
LET THE GALAXY BURN [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  LPA
LORD OF THE NIGHT [@WARHAMMER 40K]  SPURRIER, SIMON  P
PAWNS OF CHAOS [WARHAMMER 40K]  CRAIG, BRIAN  P
REBEL WINTER [WARHAMMER 40K]  PARKER, STEVE  P
ROGUE STAR [@WARHAMMER 40K]  HOARE, ANDY  P
SHADOW POINT [@WARHAMMER 40K]  RENNIE, GORDON  P
SONS OF FENRIS [@WARHAMMER 40K]  LIGHTNER, LEE  P
STATUS: DEADZONE [WARHAMMER 40K]  GASCOIGNE, MARC/JONES, ANDY  PA
STORM OF IRON [WARHAMMER 40K]  MCNEILL, GRAHAM  P
TALES FROM THE DARK MILLENNIUM [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
WHAT PRICE VICTORY [WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
WORDS OF BLOOD [WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
CONQUEST OF ARMAGEDDON [@WARHAMMER 40K BLACK TEMPLARS 2]  GREEN, JONATHAN  P
CRUSADE FOR ARMAGEDDON [@WARHAMMER 40K BLACK TEMPLARS 1]  GREEN, JONATHAN  P
DEUS ENCARMINE [@WARHAMMER 40K BLOOD ANGELS 1]  SWALLOW, JAMES  P
DEUS SANGUINIUS [@WARHAMMER 40K BLOOD ANGELS 2]  SWALLOW, JAMES  P
CAVES OF ICE [@WARHAMMER 40K CIAPHAS CAIN 2]  MITCHELL, SANDY  P

CIAPHAS CAIN: HERO OF THE IMPERIUM [@WARHAMMER 40K CIAPHAS CAIN 1,2,3] MITCHELL, SANDY

LP

DEATH OR GLORY [@WARHAMMER 40K CIAPHAS CAIN 4]  MITCHELL, SANDY  P
FOR THE EMPEROR [@WARHAMMER 40K CIAPHAS CAIN 1]  MITCHELL, SANDY  P
TRAITOR’S HAND, THE [@WARHAMMER 40K CIAPHAS CAIN 3]  MITCHELL, SANDY  P

HERETIC SAINT [@WARHAMMER 40K DAEMONIFUGE] CAMPBELL, JIM/SPURRIER, SIMON/ RENNIE, GORDON/WALKER, KEV

SCX

ASCENSION [@WARHAMMER 40K DAWN OF WAR 2]  GOTO, C. S.  P
TEMPEST [@WARHAMMER 40K DAWN OF WAR 3]  GOTO, C. S.  P
WARRIOR BLOOD [@WARHAMMER 40K DEATHWATCH 1]  GOTO, C. S.  P
WARRIOR COVEN [@WARHAMMER 40K DEATHWATCH 2]  GOTO, C. S.  P
EISENHORN [@WARHAMMER 40K EISENHORN 1,2,3]  ABNETT, DAN  H,LP
HERETICUS [WARHAMMER 40K EISENHORN 3]  ABNETT, DAN  P
MALLEUS [WARHAMMER 40K EISENHORN 2]  ABNETT, DAN  P
XENOS [WARHAMMER 40K EISENHORN 1]  ABNETT, DAN  P

The MITSFS Seriesdex  September 24, 2007  Page 184
TITLE  AUTHOR  CODE
FARSEER [WARHAMMER 40K ELDAR]  KING, WILLIAM  P
FIRST AND ONLY [WARHAMMER 40K GAUNT’S GHOSTS 1]  ABNETT, DAN  P
FOUNDING, THE [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  LP
GHOSTMAKER [WARHAMMER 40K GAUNT’S GHOSTS 2]  ABNETT, DAN  P
HIS LAST COMMAND [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  H
HONOUR GUARD [WARHAMMER 40K GAUNT’S GHOSTS 4]  ABNETT, DAN  P
NECROPOLIS [WARHAMMER 40K GAUNT’S GHOSTS 3]  ABNETT, DAN  P
SABBAT MARTYR [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  P
STRAIGHT SILVER [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  P
TRAITOR GENERAL [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  P
GUNS OF TANITH, THE [@WARHAMMER 40K GAUNT’S GHOSTS #5]  ABNETT, DAN  P
DARK ADEPTUS [@WARHAMMER 40K GREY KNIGHTS 2]  COUNTER, BEN  P
GREY KNIGHTS [@WARHAMMER 40K GREY KNIGHTS 1]  COUNTER, BEN  P
FALSE GODS [@WARHAMMER 40K HORUS HERESY 2]  MCNEILL, GRAHAM  P
FLIGHT OF THE EISENSTEIN, THE [@WARHAMMER 40K HORUS HERESY 4]  SWALLOW, JAMES  P
GALAXY IN FLAMES [@WARHAMMER 40K HORUS HERESY 3]  COUNTER, BEN  P
HORUS RISING [@WARHAMMER 40K HORUS HERESY 1]  ABNETT, DAN  P
ANNIHILATION SQUAD [WARHAMMER 40K LAST CHANCERS 3]  THORPE, GAV  P
KILL TEAM [WARHAMMER 40K LAST CHANCERS 2]  THORPE, GAV  P
13TH LEGION [WARHAMMER 40K LAST CHANCERS 1]  THORPE, GAV  P
RAVENOR [@WARHAMMER 40K RAVENOR 1]  ABNETT, DAN  P
RAVENOR RETURNED [@WARHAMMER 40K RAVENOR 2]  ABNETT, DAN  H
RAVENOR ROGUE [@WARHAMMER 40K RAVENOR 3]  ABNETT, DAN  H
CROSSFIRE [@WARHAMMER 40K SHIRA CALPURNIA 1]  FARRER, MATTHEW  P
LEGACY [@WARHAMMER 40K SHIRA CALPURNIA 2]  FARRER, MATTHEW  P
FAITH & FIRE [@WARHAMMER 40K SISTERS OF BATTLE 1]  SWALLOW, JAMES  P
BLEEDING CHALICE, THE [@WARHAMMER 40K SOUL DRINKERS 2]  COUNTER, BEN  P
CRIMSON TEARS [@WARHAMMER 40K SOUL DRINKERS 3]  COUNTER, BEN  P
SOUL DRINKER [@WARHAMMER 40K SOUL DRINKERS 1]  COUNTER, BEN  P
GREY HUNTER [WARHAMMER 40K SPACE WOLF 3]  KING, WILLIAM  P
RAGNAR’S CLAW [WARHAMMER 40K SPACE WOLF 2]  KING, WILLIAM  P
SPACE WOLF [WARHAMMER 40K SPACE WOLF 1]  KING, WILLIAM  P
SPACE WOLF OMNIBUS, THE [@WARHAMMER 40K SPACE WOLF 1,2,3]  KING, WILLIAM  H
WOLFBLADE [@WARHAMMER 40K SPACE WOLF 4]  KING, WILLIAM  P
DEAD SKY, BLACK SUN [@WARHAMMER 40K ULTRAMARINES 3]  MCNEILL, GRAHAM  P
NIGHTBRINGER [WARHAMMER 40K ULTRAMARINES 1]  MCNEILL, GRAHAM  P

ULTRAMARINES OMNIBUS, THE [@WARHAMMER 40K ULTRAMARINES 1,2,3] MCNEILL, GRAHAM

LP

WARRIORS OF ULTRAMAR [WARHAMMER 40K ULTRAMARINES 2]  MCNEILL, GRAHAM  P
PRINCE OF ILL LUCK, THE [WARHORSE OF ESDRAGON 1]  DEXTER, SUSAN  P
TRUE KNIGHT, THE [WARHORSE OF ESDRAGON 3]  DEXTER, SUSAN  P
WIND-WITCH, THE [WARHORSE OF ESDRAGON 2]  DEXTER, SUSAN  C/P:2,P
KILLING FIELDS [WARKEEP 2030]  KASNER, MICHAEL  P
ESCAPE VELOCITY [WARLOCK 0]  STASHEFF, CHRISTOPHER  C/P:2,P
KING KOBOLD [WARLOCK 2]  STASHEFF, CHRISTOPHER  C/P:2,P
KING KOBOLD REVIVED [WARLOCK 2]  STASHEFF, CHRISTOPHER  C/P:2,P
ODD WARLOCK OUT [WARLOCK 7,8,9]  STASHEFF, CHRISTOPHER  H
TO THE MAGIC BORN [WARLOCK 0,1]  STASHEFF, CHRISTOPHER  H:2
WARLOCK AND SON [WARLOCK 11]  STASHEFF, CHRISTOPHER  C/P,SFWA-P:2
WARLOCK ENLARGED, THE [WARLOCK 2,3,4]  STASHEFF, CHRISTOPHER  H:2
WARLOCK ENRAGED, THE [WARLOCK 4]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK HERETICAL, THE [WARLOCK 7]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK IN SPITE OF HIMSELF, THE [WARLOCK 1]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK INSANE, THE [WARLOCK 9]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK IS MISSING, THE [WARLOCK 6]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK ROCK, THE [WARLOCK 10]  STASHEFF, CHRISTOPHER  C/P,SFWA-P:2,P
WARLOCK UNLOCKED, THE [WARLOCK 3]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK WANDERING, THE [WARLOCK 5]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK’S COMPANION, THE [WARLOCK 8]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK’S LAST RIDE, THE [WARLOCK 12]  STASHEFF, CHRISTOPHER  P
WARLOCK’S NIGHT OUT, THE [WARLOCK 5,6]  STASHEFF, CHRISTOPHER  H:2
HERE BE MONSTERS [WARLOCK’S HEIRS 5]  STASHEFF, CHRISTOPHER  P
M’LADY WITCH [WARLOCK’S HEIRS 2]  STASHEFF, CHRISTOPHER  C/P,P
QUICKSILVER’S KNIGHT [WARLOCK’S HEIRS 3]  STASHEFF, CHRISTOPHER  P
SPELL-BOUND SCHOLAR, THE [WARLOCK’S HEIRS]  STASHEFF, CHRISTOPHER  P
WIZARD IN ABSENTIA, A [WARLOCK’S HEIRS 1]  STASHEFF, CHRISTOPHER  C/P:2,SFWA-P:2,P
WARLORD, THE [@WARLORD #1]  FROST, JASON  P
CUTTHROAT, THE [@WARLORD #2]  FROST, JASON  P
BADLAND [@WARLORD #3]  FROST, JASON  P
PRISONLAND [@WARLORD #4]  FROST, JASON  P
TERMINAL ISLAND [@WARLORD #5]  FROST, JASON  P
KILLER’S KEEP [@WARLORD #6]  FROST, JASON  P
WANDERER [WARRIOR 2]  MCQUINN, DONALD E.  P

The MITSFS Seriesdex  September 24, 2007  Page 185
TITLE  AUTHOR  CODE
WARRIOR [WARRIOR 1]  MCQUINN, DONALD E.  C/P,H,LP
WITCH [WARRIOR 3]  MCQUINN, DONALD E.  C/P,LP
SORCERER’S BLOOD [@WARRIOR OF VENGEANCE #1]  COE, ROSS ANTON  P
TRAILS OF PERIL [@WARRIOR OF VENGEANCE #2]  COE, ROSS ANTON  P
SUNSET [@WARRIORS NEW PROPHECY #6]  HUNTER, ERIN  H
SILVER LAKE, THE [@WARRIORS OF ESTAVIA #1]  PATTON, FIONA  H
WARRIORS OF SPIDER, THE [WARRIORS OF SPIDER 1]  GEAR, W. MICHAEL  C/P:2,P
WAY OF SPIDER, THE [WARRIORS OF SPIDER 2]  GEAR, W. MICHAEL  C/P:2,P
WEB OF SPIDER, THE [WARRIORS OF SPIDER 3]  GEAR, W. MICHAEL  C/P:2,P
CURSE OF THE MISTWRAITH [WARS OF LIGHT AND SHADOW 1]  WURTS, JANNY  C/P:3,SFWA,P

FUGITIVE PRINCE [WARS OF LIGHT AND SHADOW 3/ ALLIANCE OF LIGHT 1] WURTS, JANNY

H,P

GRAND CONSPIRACY [WARS OF LIGHT AND SHADOW 4/ ALLIANCE OF LIGHT 2] WURTS, JANNY

H

PERIL’S GATE [WARS OF LIGHT AND SHADOW 5/ALLIANCE OF LIGHT 3]  WURTS, JANNY  H

SHIPS OF MERIOR (COMPLETE) [WARS OF LIGHT AND SHADOW 2/ SHIPS OF MERIOR 1,2] WURTS, JANNY

H

SHIPS OF MERIOR (PARTIAL) [WARS OF LIGHT AND SHADOW 2/ SHIPS OF MERIOR 1] WURTS, JANNY

C/P:2,P

WARHOST OF VASTMARK [WARS OF LIGHT AND SHADOW 2/ SHIPS OF MERIOR 2] WURTS, JANNY

C/P,P

WARSTRIDER [@WARSTRIDER 1]  KEITH, WILLIAM H. JR.  P,SFWA-P
WARSTRIDER: BATTLEMIND [@WARSTRIDER 6]  KEITH, WILLIAM H. JR.  P
WARSTRIDER: JACKERS [@WARSTRIDER 3]  KEITH, WILLIAM H. JR.  P
WARSTRIDER: NETLINK [@WARSTRIDER 5]  KEITH, WILLIAM H. JR.  C/P,P
WARSTRIDER: REBELLION [@WARSTRIDER 2]  KEITH, WILLIAM H. JR.  P,SFWA-P
WARSTRIDER: SYMBIONTS [@WARSTRIDER 4]  KEITH, WILLIAM H. JR.  P
RESURRECTION [@WARTORN 1]  ASPRIN, ROBERT LYNN/DEL CARLO, ERIC  P
AFTERMATH [WASTEWORLD 1]  BARTON, JAMES  P
RESURRECTION [WASTEWORLD 2]  BARTON, JAMES  P
HIDDEN WORLD, THE [WATCHERS 2]  GORDON, STUART  P
MEETING OF THE WATERS, THE [@WATCHERS #1]  MOR, CAISEAL  P
KING OF SLEEP, THE [@WATCHERS #2]  MOR, CAISEAL  P
RAVEN GAME, THE [@WATCHERS #3]  MOR, CAISEAL  P
CLOUD WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
DARK WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
FIRE WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
STORM WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
ASCENSION [@WATER #1]  DALKEY, KARA  LP
O GREENEST BRANCH! [@WATER! #1]  BAUDINO, GAEL  P
DOVE LOOKED IN, THE [@WATER! #2]  BAUDINO, GAEL  P
REUNION [@WATER #2]  DALKEY, KARA  LP
BRANCH AND CROWN [@WATER! #3]  BAUDINO, GAEL  C/P,P
BREACH IN THE WATERSHED, A [WATERSHED 1]  NILES, DOUGLAS  H,LP
DARKENHEIGHT [WATERSHED 2]  NILES, DOUGLAS  P
WAR OF THREE WATERS [WATERSHED 3]  NILES, DOUGLAS  H,LP
TALES FROM WATERSHIP DOWN [WATERSHIP DOWN 2]  ADAMS, RICHARD  C/P,P
WATERSHIP DOWN [WATERSHIP DOWN 1]  ADAMS, RICHARD  C/P,P
CRUSADER [WAYFARER REDEMPTION 6]  DOUGLASS, SARA  H
ENCHANTER [WAYFARER REDEMPTION 2]  DOUGLASS, SARA  H
PILGRIM [WAYFARER REDEMPTION 5]  DOUGLASS, SARA  H
SINNER [WAYFARER REDEMPTION 4]  DOUGLASS, SARA  P
STARMAN [WAYFARER REDEMPTION 3]  DOUGLASS, SARA  H:2
WAYFARER REDEMPTION, THE [WAYFARER REDEMPTION 1]  DOUGLASS, SARA  P
ECHOES OF CHAOS, THE [@WEAPONS OF CHAOS #1]  VARDEMAN, ROBERT E.  C/P:2,SFWA-P,P
EQUATIONS OF CHAOS [@WEAPONS OF CHAOS #2]  VARDEMAN, ROBERT E.  P
COLORS OF CHAOS [@WEAPONS OF CHAOS #3]  VARDEMAN, ROBERT E.  P
CHILL FACTOR [WEATHER WARDEN 3]  CAINE, RACHEL  C/P,P
FIRESTORM [WEATHER WARDEN 5]  CAINE, RACHEL  P
HEAT STROKE [WEATHER WARDEN 2]  CAINE, RACHEL  P
ILL WIND [WEATHER WARDEN 1]  CAINE, RACHEL  P
WINDFALL [WEATHER WARDEN 4]  CAINE, RACHEL  P
DREAMCASTLE [WEB,THE]  BOWKETT, STEPHEN  P
FALL OF ATLANTIS, THE [WEB 1,2]  BRADLEY, MARION ZIMMER  C/P:2,P
GULLIVERZONE [WEB,THE]  BAXTER, STEPHEN  P
LIGHTSTORM [WEB,THE]  HAMILTON, PETER F.  P
SORCERESS [WEB,THE]  FUREY, MAGGIE  P
UNTOUCHABLE [WEB,THE]  BROWN, ERIC  P
WEB OF DARKNESS [WEB 2]  BRADLEY, MARION ZIMMER  C/P:2,LP
WEB OF LIGHT [WEB 1]  BRADLEY, MARION ZIMMER  C/P:2,P
BEHOLDER’S EYE [WEB SHIFTERS 1]  CZERNEDA, JULIE E.  C/P,P
CHANGING VISION [@WEB SHIFTERS #2]  CZERNEDA, JULIE E.  C/P:2,P
HIDDEN IN SIGHT [WEB SHIFTERS #3]  CZERNEDA, JULIE E.  P

The MITSFS Seriesdex  September 24, 2007  Page 186
TITLE  AUTHOR  CODE
MEMORIES END [@WEB WARRIORS #1]  LUCENO, JAMES  P
DIMENSION X [@WEB WARRIORS #2]  LUCENO, JAMES  P
NIGHTSHADE [WEIRD HEROES 4]  KING, TAPPAN/MEACHAM, BETH  P
OZ ENCOUNTER, THE [WEIRD HEROES 5]  WOLFMAN, MARV/WHITE, TED  C/P,P
ZOMBIE SURF COMMANDOS FROM MARS! [WEIRD ZONE 1]  ABBOTT, TONY  LP
BEYOND, THE [WELL-BUILT CITY 3]  FORD, JEFFREY  H
MEMORANDA [WELL-BUILT CITY 2]  FORD, JEFFREY  P
PHYSIOGNOMY, THE [WELL-BUILT CITY 1]  FORD, JEFFREY  C/P:2,LP
EXILES AT THE WELL OF SOULS [WELL WORLD 2]  CHALKER, JACK L.  C/P:2,P
GHOST OF THE WELL OF SOULS [WELL WORLD]  CHALKER, JACK L.  P
MIDNIGHT AT THE WELL OF SOULS [WELL WORLD 1]  CHALKER, JACK L.  C/P:3,P
QUEST FOR THE WELL OF SOULS [WELL WORLD 3]  CHALKER, JACK L.  C/P:3,P
RETURN OF NATHAN BRAZIL, THE [WELL WORLD 4]  CHALKER, JACK L.  C/P:2,P
SEA IS FULL OF STARS, THE [WELL WORLD 6]  CHALKER, JACK L.  C/P,P
TWILIGHT AT THE WELL OF SOULS [WELL WORLD 5]  CHALKER, JACK L.  C/P:2,P
WATCHERS AT THE WELL, THE [WELL WORLD 6,7,8]  CHALKER, JACK L.  H:2

ECHOES OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 1] CHALKER, JACK L.

C/P,LP,H

GODS OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 3] CHALKER, JACK L.  H:2,LP

SHADOW OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 2] CHALKER, JACK L.

H:2,LP

ALLY [WESS’HAR WARS 5]  TRAVISS, KAREN  P
CITY OF PEARL [WESS’HAR WARS 1]  TRAVISS, KAREN  P
CITY OF PEARL (PROOF) [WESS’HAR WARS 1]  TRAVISS, KAREN  SR-P
CROSSING THE LINE [WESS’HAR WARS 2]  TRAVISS, KAREN  P
MATRIARCH [WESS’HAR WARS 4]  TRAVISS, KAREN  P
WORLD BEFORE, THE [WESS’HAR WARS 3]  TRAVISS, KAREN  P
DARK SLEEPER [WESTERN LIGHTS 1]  BARLOUGH, JEFFREY E.  LP
HOUSE IN THE HIGH WOOD, THE [WESTERN LIGHTS 2]  BARLOUGH, JEFFREY E.  LP
STRANGE CARGO [WESTERN LIGHTS 3]  BARLOUGH, JEFFREY E.  LP
VOICES [WESTERN SHORE 2]  LE GUIN, URSULA K.  H
BEGGAR QUEEN, THE [WESTMARK 3]  ALEXANDER, LLOYD  P
KESTREL, THE [WESTMARK 2]  ALEXANDER, LLOYD  P
WESTMARK [WESTMARK 1]  ALEXANDER, LLOYD  C/P,P
EARTHSTONE, THE [WESTRIA 4]  PAXSON, DIANA L.  P,SFWA-P
JEWEL OF FIRE, THE [WESTRIA 7]  PAXSON, DIANA L.  C/P,P
LADY OF DARKNESS [WESTRIA 2]  PAXSON, DIANA L.  C/P,P
LADY OF LIGHT [WESTRIA 1]  PAXSON, DIANA L.  P
MISTRESS OF THE JEWELS, THE [WESTRIA 1,2]  PAXSON, DIANA L.  C/P,P
SEA STAR, THE [WESTRIA 5]  PAXSON, DIANA L.  P,SFWA-P
SILVERHAIR THE WANDERER [WESTRIA 3]  PAXSON, DIANA L.  P
WIND CRYSTAL, THE [WESTRIA 6]  PAXSON, DIANA L.  C/P,SFWA-P,P

WHAT MIGHT HAVE BEEN I AND II [WHAT MIGHT HAVE BEEN 1,2] BENFORD, GREGORY/

GREENBERG, MARTIN H.

HA

WHAT MIGHT HAVE BEEN III AND IV [WHAT MIGHT HAVE BEEN 3,4] BENFORD, GREGORY/

GREENBERG, MARTIN H.

HA

ALTERNATE EMPIRES [@WHAT MIGHT HAVE BEEN #1] BENFORD, GREGORY/

GREENBERG, MARTIN H.

C/PA,PA,SFWA-P

ALTERNATE HEROES [@WHAT MIGHT HAVE BEEN #2] BENFORD, GREGORY/

GREENBERG, MARTIN H.

C/PA,PA

ALTERNATE AMERICAS [@WHAT MIGHT HAVE BEEN #4] BENFORD, GREGORY/

GREENBERG, MARTIN H.

C/PA,PA

DARKLIGHT GRIMOIRE, THE [WHEEL 3]  FRANKOS, STEVEN  P
HEART OF SPARRILL, THE [WHEEL 2]  FRANKOS, STEVEN  P
JEWEL OF EQUILIBRANT, THE [WHEEL 1]  FRANKOS, STEVEN  P
CROSSROADS OF TWILIGHT [WHEEL OF TIME 10]  JORDAN, ROBERT  H
CROWN OF SWORDS, A [WHEEL OF TIME 7]  JORDAN, ROBERT  H:2,P
DRAGON REBORN, THE [WHEEL OF TIME 3]  JORDAN, ROBERT  C/P,P,H:2
DRAGON REBORN, THE (PROOF) [WHEEL OF TIME 3]  JORDAN, ROBERT  SR-LP
EYE OF THE WORLD, THE [WHEEL OF TIME 1]  JORDAN, ROBERT  C/P:2,H,LP,SFWA
EYE OF THE WORLD, THE (CHAPTERS 1-18) [@WHEEL OF TIME 1]  JORDAN, ROBERT  C/P,P
FIRES OF HEAVEN, THE [WHEEL OF TIME 5]  JORDAN, ROBERT  C/P,SFWA,P,H:2
GREAT HUNT, THE [WHEEL OF TIME 2]  JORDAN, ROBERT  H:2,P,SFWA:2
KNIFE OF DREAMS [WHEEL OF TIME 11]  JORDAN, ROBERT  H:2
LORD OF CHAOS [WHEEL OF TIME 6]  JORDAN, ROBERT  H:2,P
NEW SPRING [WHEEL OF TIME 0]  JORDAN, ROBERT  H
PATH OF DAGGERS, THE [WHEEL OF TIME 8]  JORDAN, ROBERT  H:2,P
SHADOW RISING, THE [WHEEL OF TIME 4]  JORDAN, ROBERT  C/P:2,P
WHEEL OF TIME, THE (SYNOPSIS) [WHEEL OF TIME 1,2]  JORDAN, ROBERT  SR-L
WINTER’S HEART [WHEEL OF TIME 9]  JORDAN, ROBERT  H:2,P

WORLD OF ROBERT JORDAN’S THE WHEEL OF TIME, THE [WHEEL OF TIME] JORDAN, ROBERT/PATTERSON, TERESA

R/REF-H

AFTER WORLDS COLLIDE [WHEN WORLDS COLLIDE 2]  BALMER, EDWIN/WYLIE, PHILIP  H
AFTER WORLDS COLLIDE [WHEN WORLDS COLLIDE 2]  WYLIE, PHILIP/BALMER, EDWIN  C/P,P,H:2

The MITSFS Seriesdex  September 24, 2007  Page 187
TITLE  AUTHOR  CODE
WHEN WORLDS COLLIDE [WHEN WORLDS COLLIDE 1]  WYLIE, PHILIP/BALMER, EDWIN  C/P,P,H:2
AWAKENING DREAM, THE [WHISPERING HILLS 5]  CAMERON, KATE  P
CURSE OF WHISPERING HILLS, THE [WHISPERING HILLS 2]  CAMERON, KATE  P
ECHOES OF EVIL [WHISPERING HILLS 6]  CAMERON, KATE  P
SHADOWS ON THE MOON [WHISPERING HILLS 3]  CAMERON, KATE  P
2 L8 4 U (PROOF) [@WICKED DEAD #1]  PETRUCHA, STEFAN/PENDLETON, THOMAS  SR-LP
GATHERING OF WIDOWMAKERS, A [WIDOWMAKER 4]  RESNICK, MIKE  H
WIDOWMAKER, THE [WIDOWMAKER 1]  RESNICK, MIKE  C/P,P
WIDOWMAKER REBORN, THE [WIDOWMAKER 2]  RESNICK, MIKE  C/P,P
WIDOWMAKER UNLEASHED, THE [WIDOWMAKER 3]  RESNICK, MIKE  P
WIDOWMAKERS, THE [WIDOWMAKER 1,2,3]  RESNICK, MIKE  H:2
TURN OF THE CARDS [WILD CARDS 12]  MILAN, VICTOR  C/P:2,P
WILD CARDS [@WILD CARDS #1]  MARTIN, GEORGE R. R.  HA,PA
ACES HIGH [@WILD CARDS #2]  MARTIN, GEORGE R. R.  C/PA,PA
JOKERS WILD [@WILD CARDS #3]  MARTIN, GEORGE R. R.  C/PA,PA
ACES ABROAD [@WILD CARDS #4]  MARTIN, GEORGE R. R.  C/PA,PA
DOWN AND DIRTY [@WILD CARDS #5]  MARTIN, GEORGE R. R.  C/PA,PA
ACE IN THE HOLE [@WILD CARDS #6]  MARTIN, GEORGE R. R.  C/PA,PA
DEAD MAN’S HAND [@WILD CARDS #7]  MARTIN, GEORGE R. R./MILLER, JOHN J.  C/P:2,P
ONE-EYED JACKS [@WILD CARDS #8]  MARTIN, GEORGE R. R.  PA,C/PA
JOKERTOWN SHUFFLE [@WILD CARDS #9]  MARTIN, GEORGE R. R.  C/PA,PA
DOUBLE SOLITAIRE [@WILD CARDS #10]  SNODGRASS, MELINDA M.  C/P:2,P
DEALER’S CHOICE [@WILD CARDS #11]  MARTIN, GEORGE R. R.  PA,C/PA
DEATH DRAWS FIVE [@WILD CARDS #17]  MILLER, JOHN J.  H
BLACK TRUMP [@WILD CARDS NEW CYCLE 3]  MARTIN, GEORGE R. R.  C/PA,PA,SFWA-P
CARD SHARKS [@WILD CARDS NEW CYCLE 1]  MARTIN, GEORGE R. R.  PA,C/PA,SFWA-P
MARKED CARDS [@WILD CARDS NEW CYCLE 2]  MARTIN, GEORGE R. R.  PA
WILD HUNT: CHILD OF FIRE, THE [WILD HUNT 2]  FOXE, JOCELIN  P
WILD HUNT: VENGEANCE MOON, THE [WILD HUNT 1]  FOXE, JOCELIN  P
WILD MAGIC [WILD MAGIC 1]  CLAYTON, JO  C/P:2,P
WILDFIRE [@WILD MAGIC #2]  CLAYTON, JO  C/P:2,SFWA-P,P
MAGIC WARS, THE [@WILD MAGIC #3]  CLAYTON, JO  C/P,P
GOLDEN CAT, THE [WILD ROAD 2]  KING, GABRIEL  P
WILD ROAD, THE [WILD ROAD 1]  KING, GABRIEL  C/P,P
ACROSS THE FAR MOUNTAIN [@WILDERNESS OF FOUR #1]  HANCOCK, NIEL  C/P,P
PLAINS OF THE SEA, THE [@WILDERNESS OF FOUR #2]  HANCOCK, NIEL  C/P,P
ON THE BOUNDARIES OF DARKNESS [@WILDERNESS OF FOUR #3]  HANCOCK, NIEL  P
ROAD TO THE MIDDLE ISLANDS, THE [@WILDERNESS OF FOUR #4]  HANCOCK, NIEL  P
ALL NIGHT AWAKE [WILLIAM SHAKESPEARE 2]  HOYT, SARAH A.  P
ILL MET BY MOONLIGHT [WILLIAM SHAKESPEARE 1]  HOYT, SARAH A.  P
WIND SINGER, THE [WIND ON FIRE 1]  NICHOLSON, WILLIAM  H
BRIDGE OF DAWN, THE [WINDAMEIR CIRCLE 4]  HANCOCK, NIEL  P
FIRES OF WINDAMEIR, THE [WINDAMEIR CIRCLE 1]  HANCOCK, NIEL  P
SEA OF SILENCE, THE [WINDAMEIR CIRCLE 2]  HANCOCK, NIEL  P
WANDERER’S RETURN, A [WINDAMEIR CIRCLE 3]  HANCOCK, NIEL  P
IMAGE OF VOICES, AN [@WINDHOVER TAPES #1]  NORWOOD, WARREN  C/P,P
FLEXING THE WARP [@WINDHOVER TAPES #2]  NORWOOD, WARREN  C/P:2,P
FIZE OF THE GABRIEL RATCHETS [@WINDHOVER TAPES #3]  NORWOOD, WARREN  C/P:2,P
PLANET OF FLOWERS [@WINDHOVER TAPES #4]  NORWOOD, WARREN  C/P:2,P
DARKMAGE [WINDROSE 1,2]  HAMBLY, BARBARA  H:2
DOG WIZARD [WINDROSE 3]  HAMBLY, BARBARA  C/P,P,H
SILENT TOWER, THE [WINDROSE 1]  HAMBLY, BARBARA  C/P:2,P
SILICON MAGE, THE [WINDROSE 2]  HAMBLY, BARBARA  C/P:2,P
BONDS OF VENGEANCE [WINDS OF THE FORELANDS 3]  COE, DAVID B.  H
RULES OF ASCENSION [WINDS OF THE FORELANDS 1]  COE, DAVID B.  H
SEEDS OF BETRAYAL [WINDS OF THE FORELANDS 2]  COE, DAVID B.  P
SHAPERS OF DARKNESS [WINDS OF THE FORELANDS 4]  COE, DAVID B.  H
WEAVERS OF WAR [WINDS OF THE FORELANDS 5]  COE, DAVID B.  H
ACTION STATIONS [@WING COMMANDER 6]  FORSTCHEN, WILLIAM R.  C/P,P

END RUN [@WING COMMANDER 2] STASHEFF, CHRISTOPHER/ FORSTCHEN, WILLIAM R.

P,SFWA-P

FALSE COLORS [@WING COMMANDER 7]  FORSTCHEN, WILLIAM R./KEITH, ANDREW  C/P,P
FLEET ACTION [@WING COMMANDER 3]  FORSTCHEN, WILLIAM R.  P,SFWA-P
FREEDOM FLIGHT [WING COMMANDER 1]  LACKEY, MERCEDES/GUON, ELLEN  C/P:2,SFWA-P,P
HEART OF THE TIGER [@WING COMMANDER 4]  FORSTCHEN, WILLIAM R./KEITH, ANDREW  P
PILGRIM STARS [@WING COMMANDER]  TELEP, PETER  P
PRICE OF FREEDOM, THE [@WING COMMANDER 5]  FORSTCHEN, WILLIAM R./OHLANDER, BEN  P
WING COMMANDER [WING COMMANDER]  TELEP, PETER  P
CIRCLE WAR [WINGMAN 2]  MALONEY, MACK  P
FINAL STORM, THE [WINGMAN 6]  MALONEY, MACK  P
FREEDOM EXPRESS [WINGMAN 7]  MALONEY, MACK  P
LUCIFER CRUSADE, THE [WINGMAN 3]  MALONEY, MACK  P
TWISTED CROSS, THE [WINGMAN 5]  MALONEY, MACK  P

The MITSFS Seriesdex  September 24, 2007  Page 188
TITLE  AUTHOR  CODE
WINGMAN [WINGMAN 1]  MALONEY, MACK  P
AT WINTER’S END [WINTER 1]  SILVERBERG, ROBERT  H:2,P
NEW SPRINGTIME, THE [WINTER 2]  SILVERBERG, ROBERT  C/P,P,H:2
QUEEN OF SPRINGTIME, THE [WINTER 2]  SILVERBERG, ROBERT  LP
MOUNTAINS OF CHANNADRAN, THE [WINTER KING’S WAR 3]  DEXTER, SUSAN  C/P,P
RING OF ALLAIRE, THE [WINTER KING’S WAR 1]  DEXTER, SUSAN  C/P:2,P
SWORD OF CALANDRA, THE [WINTER KING’S WAR 2]  DEXTER, SUSAN  C/P,P
SINGER AND THE SEA, THE [WINTER OF THE WORLD]  ROHAN, MICHAEL SCOTT  P
ANVIL OF ICE, THE [@WINTER OF THE WORLD #1]  ROHAN, MICHAEL SCOTT  C/P,P
FORGE IN THE FOREST, THE [@WINTER OF THE WORLD #2]  ROHAN, MICHAEL SCOTT  P
HAMMER OF THE SUN, THE [@WINTER OF THE WORLD #3]  ROHAN, MICHAEL SCOTT  H,P
BRANDER’S BOOK [WINTER WORLD 4]  MILLS, C. J.  C/P,P
EGIL’S BOOK [WINTER WORLD 2]  MILLS, C. J.  C/P,P
KIT’S BOOK [WINTER WORLD 3]  MILLS, C. J.  C/P,P
WINTER WORLD [WINTER WORLD 1]  MILLS, C. J.  C/P,P
DRAGONSBANE [WINTERLANDS 1]  HAMBLY, BARBARA  C/P,P,H
DRAGONSHADOW [WINTERLANDS 2]  HAMBLY, BARBARA  H
DRAGONSTAR [WINTERLANDS 4]  HAMBLY, BARBARA  H
KNIGHT OF THE DEMON QUEEN [WINTERLANDS 3]  HAMBLY, BARBARA  P
AESTIVAL TIDE [WINTERLONG 2]  HAND, ELIZABETH  C/P,P
WINTERLONG [WINTERLONG 1]  HAND, ELIZABETH  C/P:2,P
UNLEASHED IN SPACE [WISHBONE 3]  STEELE, ALEXANDER  LP
CIARA’S SONG [WITCH WORLD]  NORTON, ANDRE/MCCONCHIE, LYN  P
LORE OF THE WITCH WORLD [WITCH WORLD]  NORTON, ANDRE  C/P:2,P
SILVER MAY TARNISH [WITCH WORLD]  NORTON, ANDRE/MCCONCHIE, LYN  H
TALES OF THE WITCH WORLD [WITCH WORLD]  NORTON, ANDRE  SR-HA,HA
TALES OF THE WITCH WORLD 2 [WITCH WORLD]  NORTON, ANDRE  PA,SFWA-P
TALES OF THE WITCH WORLD 3 [WITCH WORLD]  NORTON, ANDRE  C/PA,PA
GATE OF THE CAT, THE [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P,SFWA,P,H
LOST LANDS OF THE WITCH WORLD [WITCH WORLD ESTCARP 1,2,3]  NORTON, ANDRE  H
SORCERESS OF THE WITCH WORLD [WITCH WORLD ESTCARP 3]  NORTON, ANDRE  C/P:2,P
THREE AGAINST THE WITCH WORLD [WITCH WORLD ESTCARP 1]  NORTON, ANDRE  C/P:2,P
TREY OF SWORDS [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P,P,H
’WARE HAWK [WITCH WORLD ESTCARP]  NORTON, ANDRE  H:2,P
WARLOCK OF THE WITCH WORLD [WITCH WORLD ESTCARP 2]  NORTON, ANDRE  C/P:2,P
WEB OF THE WITCH WORLD [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P:2,P
WITCH WORLD [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P:2,P
HORN CROWN [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  H:2,P
JARGOON PARD, THE [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P,P,H
SONGSMITH [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE/CRISPIN, A. C.  H:2,P
SPELL OF THE WITCH WORLD [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P:2,P
YEAR OF THE UNICORN [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P:2,P
ZARSTHOR’S BANE [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P:2,P
CRYSTAL GRYPHON, THE [WITCH WORLD HIGH HALLACK GRYPHON]  NORTON, ANDRE  C/P:2,SFWA-P,P
GRYPHON IN GLORY [WITCH WORLD HIGH HALLACK GRYPHON]  NORTON, ANDRE  H:2,P
GRYPHON’S EYRIE [WITCH WORLD HIGH HALLACK GRYPHON]  NORTON, ANDRE/CRISPIN, A. C.  H:2,SFWA-P:2,P
KEY OF THE KEPLAN, THE [WITCH WORLD SECRETS 1]  NORTON, ANDRE/MCCONCHIE, LYN  P
MAGESTONE, THE [WITCH WORLD SECRETS]  NORTON, ANDRE/SCHAUB, MARY H.  C/P,P
WARDING OF WITCH WORLD, THE [WITCH WORLD SECRETS]  NORTON, ANDRE  H:2,P

FLIGHT OF VENGEANCE [WITCH WORLD TURNING 2] NORTON, ANDRE/GRIFFIN, P. M./

SCHAUB, MARY H.

P

ON WINGS OF MAGIC [WITCH WORLD TURNING 3] NORTON, ANDRE/MATHEWS, PATRICIA/

MILLER, SASHA

H:2

STORMS OF VICTORY [WITCH WORLD TURNING 1]  NORTON, ANDRE/GRIFFIN, P. M.  C/P,SFWA,P
DEMONS [@WITCHBLADE]  BARON, MIKE  P
TALONS [@WITCHBLADE]  DE CHANCIE, JOHN  P
TALONS [@WITCHBLADE]/TERRIBLE BEAUTY, A [@WITCHBLADE]  DE CHANCIE, JOHN/MILLER, JOHN J.  P
TERRIBLE BEAUTY, A [@WITCHBLADE]  MILLER, JOHN J.  P
WITCHES OF CHISWICK, THE [WITCHES OF CHISWICK 1]  RANKIN, ROBERT  P
CURSED TOWERS, THE [WITCHES OF EILEANAN 3]  FORSYTH, KATE  C/P:2,P
FATHOMLESS CAVES, THE [WITCHES OF EILEANAN 6]  FORSYTH, KATE  C/P,P
FORBIDDEN LAND, THE [WITCHES OF EILEANAN 4]  FORSYTH, KATE  C/P,P
POOL OF TWO MOONS, THE [WITCHES OF EILEANAN 2]  FORSYTH, KATE  C/P:2,P
SKULL OF THE WORLD, THE [WITCHES OF EILEANAN 5]  FORSYTH, KATE  C/P,P
WITCHES OF EILEANAN, THE [WITCHES OF EILEANAN 1]  FORSYTH, KATE  C/P:2,P
WITCHES OF KARRES, THE [WITCHES OF KARRES 1]  SCHMITZ, JAMES H.  H:2,P

WIZARD OF KARRES, THE [WITCHES OF KARRES 2] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

H

WIZ BIZ, THE [WIZ ZUMWALT 1,2]  COOK, RICK  C/P,P
WIZARDRY COMPILED, THE [WIZ ZUMWALT 2]  COOK, RICK  C/P:2,SFWA-P:2,P
WIZARDRY CONSULTED, THE [WIZ ZUMWALT 4]  COOK, RICK  C/P:2,P
WIZARDRY CURSED, THE [WIZ ZUMWALT 3]  COOK, RICK  C/P:2,SFWA-P,P
WIZARDRY QUESTED, THE [WIZ ZUMWALT 5]  COOK, RICK  P

TITLE  AUTHOR  CODE
WIZARD’S BANE [WIZ ZUMWALT 1]  COOK, RICK  C/P:2,SFWA-P,P
FAITHFUL TRAITOR, THE [WIZARD AND DRAGON 2]  HUGHES, ROBERT DAN  C/P,P
FORGING OF THE DRAGON, THE [WIZARD AND DRAGON 1]  HUGHES, ROBERT DON  C/P:2,P
CRUSADING WIZARD, THE [WIZARD IN RHYME 7]  STASHEFF, CHRISTOPHER  C/P,P
FELINE WIZARD, THE [WIZARD IN RHYME 8]  STASHEFF, CHRISTOPHER  C/P,P
HAUNTED WIZARD, THE [WIZARD IN RHYME 6]  STASHEFF, CHRISTOPHER  P
HER MAJESTY’S WIZARD [WIZARD IN RHYME 1]  STASHEFF, CHRISTOPHER  C/P,P,H
MY SON, THE WIZARD [WIZARD IN RHYME 5]  STASHEFF, CHRISTOPHER  C/P,LP
OATHBOUND WIZARD, THE [WIZARD IN RHYME 2]  STASHEFF, CHRISTOPHER  C/P,P,H
SECULAR WIZARD, THE [WIZARD IN RHYME 4]  STASHEFF, CHRISTOPHER  C/P,P,H
WITCH DOCTOR, THE [WIZARD IN RHYME 3]  STASHEFF, CHRISTOPHER  C/P,P,H
KNIGHT, THE [WIZARD KNIGHT 1]  WOLFE, GENE  H
WIZARD, THE [WIZARD KNIGHT 2]  WOLFE, GENE  H
BAD SPELL IN YURT, A [WIZARD OF YURT 1]  BRITTAIN, C. DALE  P,SFWA-P
DAUGHTER OF MAGIC [WIZARD OF YURT 5]  BRITTAIN, C. DALE  P
MAGE QUEST [WIZARD OF YURT 3]  BRITTAIN, C. DALE  C/P,SFWA-P,P
WITCH AND THE CATHEDRAL, THE [WIZARD OF YURT 4]  BRITTAIN, C. DALE  P
WOOD NYMPH AND THE CRANKY SAINT, THE [WIZARD OF YURT 2]  BRITTAIN, C. DALE  C/P,SFWA-P,P
LORDS OF THE SWORD [@WIZARD WAR]  COOK, HUGH  P,SFWA-P:2
WIZARD WAR [@WIZARD WAR #1]  COOK, HUGH  C/P:2,P
QUESTING HERO, THE [@WIZARD WAR #2]  COOK, HUGH  C/P:2,P
HERO’S RETURN, THE [@WIZARD WAR #3]  COOK, HUGH  C/P:2,P
ORACLE, THE [@WIZARD WAR #4]  COOK, HUGH  P
LEAGUE OF THE CRIMSON CRESCENT, THE [WIZARDRY]  REAGEN, JAMES E.  P
CURSE OF SLAGFID, THE [WIZARD’S WAR 2]  BOYER, ELIZABETH H.  P
DRAGON’S CARBUNCLE, THE [WIZARD’S WAR 3]  BOYER, ELIZABETH H.  C/P,P
LORD OF CHAOS, THE [WIZARD’S WAR 4]  BOYER, ELIZABETH H.  C/P:2,P
TROLL’S GRINDSTONE, THE [WIZARD’S WAR 1]  BOYER, ELIZABETH H.  C/P,P
SHADOW SHAIA, THE [WIZENBEAK 2]  GILLILAND, ALEXIS A.  P
WIZENBEAK [WIZENBEAK 1]  GILLILAND, ALEXIS A.  C/P:2,LP
DRAGONS OF THE RHINE, THE [WODAN’S CHILDREN 2]  PAXSON, DIANA L.  C/P,P
LORD OF HORSES, THE [WODAN’S CHILDREN 3]  PAXSON, DIANA L.  P
WOLF AND THE RAVEN, THE [WODAN’S CHILDREN 1]  PAXSON, DIANA L.  H,P
NIGHT OF THE WOLF [WOLF RICE 1]  BORCHARDT, ALICE  P
SILVER WOLF, THE [WOLF RICE 2]  BORCHARDT, ALICE  C/P:2,P
WOLF KING, THE [WOLF RICE 3]  BORCHARDT, ALICE  P
LOST PRINCE, THE [WOLFKINGS]  WOOD, BRIDGET  P
WOLFKING [WOLFKINGS]  WOOD, BRIDGET  C/P,LP
GRAYHEART [WOLFWALKER 6]  HARPER, TARA K.  C/P:2,P
SHADOW LEADER [WOLFWALKER 2]  HARPER, TARA K.  C/P,P
SILVER MOONS, BLACK STEEL [WOLFWALKER 5]  HARPER, TARA K.  P
STORM RUNNER [WOLFWALKER 3]  HARPER, TARA K.  C/P,P
WOLF IN NIGHT [WOLFWALKER 7]  HARPER, TARA K.  P
WOLF’S BANE [WOLFWALKER 4]  HARPER, TARA K.  C/P,P
WOLFWALKER [WOLFWALKER 1]  HARPER, TARA K.  C/P:2,P
WOMAN OF THE IRON PEOPLE, A [WOMAN OF THE IRON PEOPLE 1,2]  ARNASON, ELEANOR  H
IN THE LIGHT OF SIGMA DRACONIS [@WOMAN OF THE IRON PEOPLE #1]  ARNASON, ELEANOR  P
CHANGING WOMEN [@WOMAN OF THE IRON PEOPLE #2]  ARNASON, ELEANOR  P,SFWA-P
SARAH [WOMEN OF GENESIS 1]  CARD, ORSON SCOTT  H
BITTEN [WOMEN OF THE OTHERWORLD 1]  ARMSTRONG, KELLEY  P
BROKEN [WOMEN OF THE OTHERWORLD]  ARMSTRONG, KELLEY  P
HAUNTED [WOMEN OF THE OTHERWORLD]  ARMSTRONG, KELLEY  P
NO HUMANS INVOLVED [WOMEN OF THE OTHERWORLD]  ARMSTRONG, KELLEY  H
STOLEN [WOMEN OF THE OTHERWORLD 2]  ARMSTRONG, KELLEY  P

DIME STORE MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 1] ARMSTRONG, KELLEY

C/P:2,P

INDUSTRIAL MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 2] ARMSTRONG, KELLEY

C/P,P

WITCH MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 1,2] ARMSTRONG, KELLEY  H
CYBERNETIC WALRUS, THE [@WONDERLAND GAMBIT #1]  CHALKER, JACK L.  C/P,SFWA-P,LP,H:2
MARCH HARE NETWORK, THE [@WONDERLAND GAMBIT #2]  CHALKER, JACK L.  H:2,P
HOT-WIRED DODO, THE [@WONDERLAND GAMBIT #3]  CHALKER, JACK L.  H:2,P
BENEATH THE WEB [WOODEN SWORD 2]  ABBEY, LYNN  P
WOODEN SWORD, THE [WOODEN SWORD 1]  ABBEY, LYNN  SFWA-P:2
REWARDS AND FAIRIES [@WORKS OF KIPLING #25]  KIPLING, RUDYARD  SR-H,H
GODS OLD AND DARK [WORLD GATES 3]  LISLE, HOLLY  C/P,P
MEMORY OF FIRE [WORLD GATES 1]  LISLE, HOLLY  C/P,P
WRECK OF HEAVEN, THE [WORLD GATES 2]  LISLE, HOLLY  P
IN THE ICE KING’S PALACE [WORLD IN AMBER 2]  ORR, A.  H,SFWA
WORLD IN AMBER, THE [WORLD IN AMBER 1]  ORR, A.  H:2,SFWA-P,P
DARKFALL [WORLD OF ADEN THUNDERSCAPE]  HENSLEY, SHANE LACY  P
INDOMITABLE THUNDER [@WORLD OF ADEN THUNDERSCAPE]  ACRES, MARK  P
SENTINEL, THE [@WORLD OF ADEN THUNDERSCAPE]  MCKEONE, DIXIE LEE  P

The MITSFS Seriesdex  September 24, 2007  Page 190
TITLE  AUTHOR  CODE
CITY OF DARKNESS: UNSEEN [WORLD OF DARKNESS]  KELLY, ERIN/WIECK, STEWART  C/PA,PA
HOUSE OF SECRETS [WORLD OF DARKNESS]  MOORE, JAMES A./MURPHY, KEVIN  P
POMEGRANATES FULL AND FINE [@WORLD OF DARKNESS]  BASSINGTHWAITE, DON  P
QUINTESSENTIAL WORLD OF DARKNESS, THE [WORLD OF DARKNESS]  WIECK, STEWART/BRANSCOME, ANNA  LPA
STRANGE CITY [WORLD OF DARKNESS]  KRAUSE, STALEY/WIECK, STEWART  C/PA,PA
WYRM WOLF [WORLD OF DARKNESS]  VAN BELKOM, EDO  P
ASSAMITE [@WORLD OF DARKNESS DARK AGES CLAN #2]  PETRUCHA, STEFAN  P
LASOMBRA [@WORLD OF DARKNESS DARK AGES CLAN #5]  WILSON, DAVID NIALL  P

ASHES AND ANGEL WINGS [@WORLD OF DARKNESS DEMON THE FALLEN #1] STOLZE, GREG

P

WRECKAGE OF PARADISE, THE [@WORLD OF DARKNESS DEMON THE FALLEN #3] STOLZE, GREG

P

TO SIFT THROUGH BITTER ASHES [@WORLD OF DARKNESS GRAILS COVENANT #1] WILSON, DAVID NIALL

P

TO SPEAK IN LIFELESS TONGUES [@WORLD OF DARKNESS GRAILS COVENANT #2] WILSON, DAVID NIALL

P

TO DREAM OF DREAMERS LOST [@WORLD OF DARKNESS GRAILS COVENANT #3] WILSON, DAVID NIALL

P

SUCH PAIN [@WORLD OF DARKNESS MAGE 1]  BASSINGTHWAITE, DON  C/P,P
HAUNTING THE DEAD [@WORLD OF DARKNESS ORPHEUS]  BOULLE, PHILIPPE  PA
VAMPIRE [@WORLD OF DARKNESS PREDATOR & PREY #1]  BOWEN, CARL  P
JUDGE [@WORLD OF DARKNESS PREDATOR & PREY #2]  FLEMING, GHERBOD  P
WEREWOLF [@WORLD OF DARKNESS PREDATOR & PREY #3]  FLEMING, GHERBOD  P
JURY [@WORLD OF DARKNESS PREDATOR & PREY #4]  FLEMING, GHERBOD  P
MAGE [@WORLD OF DARKNESS PREDATOR & PREY #5]  BOWEN, CARL  P
EXECUTIONER [@WORLD OF DARKNESS PREDATOR & PREY #6]  FLEMING, GHERBOD  P
SILVER CROWN, THE [WORLD OF DARKNESS RAGE]  BRIDGES, WILLIAM  C/P,P
CALL TO BATTLE [WORLD OF DARKNESS RAGE JAY NO-NAME 1]  MURRAY, DOUG  P
BEAST WITHIN, THE [WORLD OF DARKNESS VAMPIRE]  WIECK, STEWART  PA
BLOOD ON THE SUN [WORLD OF DARKNESS VAMPIRE]  HERBERT, BRIAN/LANDIS, MARIE  P
BLOOD RELATIONS [@WORLD OF DARKNESS VAMPIRE]  MURRAY, DOUG  P
DARK PRINCE [WORLD OF DARKNESS VAMPIRE]  HERBER, KEITH  C/P,P
NETHERWORLD [WORLD OF DARKNESS VAMPIRE]  BYERS, RICHARD LEE  P
PRINCE OF THE CITY [WORLD OF DARKNESS VAMPIRE]  HERBER, KEITH  P
EMBRACE, THE [WORLD OF DARKNESS VAMPIRE DIARY]  WEINBERG, ROBERT/REIN-HAGEN, MARK  SR-H

DEVIL’S ADVOCATE, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #1] FLEMING, GHERBOD

P

WINNOWING, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #2] FLEMING, GHERBOD

P

DARK PROPHECY [@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #3] FLEMING, GHERBOD

P

ANTHOLOGY [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  WIECK, STEWART  PA
ASSAMITE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  FLEMING, GHERBOD  P
GANGREL [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  FLEMING, GHERBOD  P
GIOVANNI [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  ACHILLI, JUSTIN  P
NOSFERATU [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  FLEMING, GHERBOD  P
RAVNOS [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  RYAN, KATHLEEN  P
SETITE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  RYAN, KATHLEEN  P
TOREADOR [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  WIECK, STEWART  P
TZIMISCE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  GRIFFIN, ERIC  P

FALL OF ATLANTA, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #1] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

EYE OF GEHENNA, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #2] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

BLOODY SEPTEMBER [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #3] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

END GAMES [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #4] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

BLOODWAR [@WORLD OF DARKNESS VAMPIRE MASQUERADE RED DEATH 1] WEINBERG, ROBERT

P

MADNESS OF PRIESTS, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN AGE #2] BOULLE, PHILIPPE

P

WOUNDED KING, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN AGE #3] BOULLE, PHILIPPE

P

BONE GNAWERS [WORLD OF DARKNESS WEREOLF 4]/ STARGAZERS [WORLD OF DARKNESS WEREOLF 4] BRIDGES, BILL/ACHILLI, JUSTIN

PA

CHILDREN OF GAIA [WORLD OF DARKNESS WEREOLF 5]/ UKTENA [WORLD OF DARKNESS WEREOLF 5] BYERS, RICHARD LEE/PETRUCHA, STEFAN

PA

SILENT STRIDERS [WORLD OF DARKNESS WEREOLF 2]/ BLACK FURIES [WORLD OF DARKNESS WEREOLF 2] BOWEN, CARL/FLEMING, GHERBOD

PA

CONSPICUOUS CONSUMPTION [@WORLD OF DARKNESS WEREWOLF 2]  VON ALLMEN, STEWART  P
HELL-STORM [@WORLD OF DARKNESS WEREWOLF]  MOORE, JAMES A.  P

RED TALONS [WORLD OF DARKNESS WEREWOLF 3]/ FIANNA [WORLD OF DARKNESS WEREWOLF 3] BOULLE, PHILIPPE/GRIFFIN, ERIC PA

TITLE  AUTHOR  CODE
WATCHER [@WORLD OF DARKNESS WEREWOLF]  GRANT, CHARLES L.  C/P,P
WHEN WILL YOU RAGE? [@WORLD OF DARKNESS WEREWOLF]  WIECK, STEWART  PA
BEYOND THE SHROUD [WORLD OF DARKNESS WRAITH]  HAUTALA, RICK  H
CARAVAN OF SHADOWS [WORLD OF DARKNESS WRAITH]  BYERS, RICHARD LEE  P
DEATH AND DAMNATION [WORLD OF DARKNESS WRAITH]  KRAUSE, STALEY/WIECK, STEWART  PA
RISEN, THE [WORLD OF DARKNESS WRAITH]  DITCHBURN, ELIZABETH/GROVE, HEATHER  SR-VLH
SINS OF THE FATHERS [WORLD OF DARKNESS WRAITH]  CHUPP, SAM  P
WRAITH: THE OBLIVION (2ND EDITION) [WORLD OF DARKNESS WRAITH]  DANSKY, RICHARD E.  [C/RAN-H]

HERALDS OF THE STORM [@WORLD OF DARKNESS YEAR OF THE SCARAB #1] BATES, ANDREW

P

LAY DOWN WITH LIONS [@WORLD OF DARKNESS YEAR OF THE SCARAB #2] BATES, ANDREW

P

LAND OF THE DEAD [@WORLD OF DARKNESS YEAR OF THE SCARAB #3]  BATES, ANDREW  P
BEHIND THE WALLS OF TERRA [WORLD OF TIERS 4]  FARMER, PHILIP JOSE  C/P:2,P
GATES OF CREATION, THE [WORLD OF TIERS 2]  FARMER, PHILIP JOSE  C/P:2,P
LAVALITE WORLD, THE [WORLD OF TIERS 5]  FARMER, PHILIP JOSE  C/P:2,P
MAKER OF UNIVERSES, THE [WORLD OF TIERS 1]  FARMER, PHILIP JOSE  C/P:2,P
MORE THAN FIRE [WORLD OF TIERS 6]  FARMER, PHILIP JOSE  H
PRIVATE COSMOS, A [WORLD OF TIERS 3]  FARMER, PHILIP JOSE  C/P:2,P
WORLD OF TIERS 1, THE [WORLD OF TIERS 1,2]  FARMER, PHILIP JOSE  H:2
WORLD OF TIERS 2, THE [WORLD OF TIERS 3,4,5]  FARMER, PHILIP JOSE  H:2
CYCLE OF HATRED [WORLD OF WARCRAFT]  DECANDIDO, KEITH R. A.  P
RISE OF THE HORDE [WORLD OF WARCRAFT]  GOLDEN, CHRISTIE  P
WORLDS [WORLDS 1]  HALDEMAN, JOE  C/P,SFWA,P,H
WORLDS APART [WORLDS 2]  HALDEMAN, JOE  C/P,P,SR-H,H
WORLDS ENOUGH AND TIME [WORLDS 3]  HALDEMAN, JOE  C/P,P,H
PLASM [WORLDS OF CHTHON]  PLATT, CHARLES  C/P,P
WIZARDS & WARRIORS [WORLDS OF POWER 5]  MILES, ELLEN  P
COLONIZATION: AFTERSHOCKS [WORLDWAR 7]  TURTLEDOVE, HARRY  H,P
COLONIZATION: DOWN TO EARTH [WORLDWAR 6]  TURTLEDOVE, HARRY  C/P,P
COLONIZATION: SECOND CONTACT [WORLDWAR 5]  TURTLEDOVE, HARRY  H
WORLDWAR: IN THE BALANCE [@WORLDWAR 1]  TURTLEDOVE, HARRY  C/P,P,H:2
WORLDWAR: STRIKING THE BALANCE [@WORLDWAR 4]  TURTLEDOVE, HARRY  C/P,P,H
WORLDWAR: TILTING THE BALANCE [@WORLDWAR 2]  TURTLEDOVE, HARRY  H:2,P
WORLDWAR: UPSETTING THE BALANCE [@WORLDWAR 3]  TURTLEDOVE, HARRY  C/P,P
HOT SLEEP: THE WORTHING CHRONICLE [WORTHING]  CARD, ORSON SCOTT  C/P:2,P
WORTHING CHRONICLE, THE [WORTHING]  CARD, ORSON SCOTT  C/P:2,P
WORTHING SAGA, THE [WORTHING]  CARD, ORSON SCOTT  C/P:2,SFWA-P,P
BEWITCHMENTS OF LOVE AND HATE, THE [WRAETHTHU 2]  CONSTANTINE, STORM  P
ENCHANTMENTS OF FLESH AND SPIRIT, THE [WRAETHTHU 1]  CONSTANTINE, STORM  SFWA-P
FULFILMENTS OF FATE AND DESIRE, THE [WRAETHTHU 3]  CONSTANTINE, STORM  P,SFWA-P
WRAETHTU [WRAETHTHU 1,2,3]  CONSTANTINE, STORM  LP
WRAITHS OF WILL AND PLEASURE, THE [WRAETHTHU HISTORIES 1]  CONSTANTINE, STORM  H
DIFFICULTY WITH DWARVES, A [WUNTVOR 1]  GARDNER, CRAIG SHAW  C/P:2,P
DISAGREEMENT WITH DEATH, A [WUNTVOR 3]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P,P
EXCESS OF ENCHANTMENTS, AN [WUNTVOR 2]  GARDNER, CRAIG SHAW  C/P:2,P
WANDERINGS OF WUNTVOR, THE [WUNTVOR 1,2,3]  GARDNER, CRAIG SHAW  H:2
DERIK’S BANE [WYNDHAM WEREWOLF]  DAVIDSON, MARYJANICE  P
LAST WIZARD, THE [WYRDRUNE 8]  HAWKE, SIMON  C/P,P
NINE LIVES OF CATSEYE GOMEZ, THE [WYRDRUNE]  HAWKE, SIMON  C/P,P
SAMURAI WIZARD, THE [WYRDRUNE 5]  HAWKE, SIMON  C/P,P
WIZARD OF CAMELOT, THE [WYRDRUNE 0]  HAWKE, SIMON  C/P:2,P
WIZARD OF 4TH STREET, THE [WYRDRUNE 1]  HAWKE, SIMON  C/P:2,P
WIZARD OF LOVECRAFT’S CAFE, THE [WYRDRUNE 7]  HAWKE, SIMON  C/P:2,P
WIZARD OF RUE MORGUE, THE [WYRDRUNE 4]  HAWKE, SIMON  C/P:2,P
WIZARD OF SANTA FE, THE [WYRDRUNE 6]  HAWKE, SIMON  C/P,P
WIZARD OF SUNSET STRIP, THE [WYRDRUNE 3]  HAWKE, SIMON  C/P:2,SFWA-P,P
WIZARD OF WHITECHAPEL, THE [WYRDRUNE 2]  HAWKE, SIMON  C/P:2,P

The MITSFS Seriesdex  September 24, 2007  Page 193
TITLE  AUTHOR  CODE
UFO DEFENSE [@X-COM]  DUANE, DIANE  P
ANTIBODIES [X-FILES]  ANDERSON, KEVIN J.  H
GOBLINS [X-FILES]  GRANT, CHARLES L.  C/P:2,P
GOBLINS [X-FILES]/WHIRLWIND [X-FILES]  GRANT, CHARLES L.  H,LP
GROUND ZERO [X-FILES]  ANDERSON, KEVIN J.  P
REGENERATION [@X-FILES]  OWENS, EVERETT  H
RUINS [X-FILES]  ANDERSON, KEVIN J.  H,P
SKIN [@X-FILES]  MEZRICH, BEN  H
TRUTH IS OUT THERE [X-FILES]  LOWRY, BRIAN  R/REF-P,C/REF-H
WHIRLWIND [X-FILES]  GRANT, CHARLES L.  C/P,P
X FILES FIGHT THE FUTURE, THE [X-FILES]  CARTER, CHRIS/HAND, ELIZABETH  H,P

X-TREME POSSIBILITIES [X-FILES] CORNELL, PAUL/DAY, MARTIN/

TOPPING, KEITH

R/REF-P

CODENAME WOLVERINE [X-MEN]  GOLDEN, CHRISTOPHER  H
LAW OF THE JUNGLE [@X-MEN]  SMEDS, DAVE  P
SABRETOOTH UNLEASHED [X-MEN]  KAMIDA, VICKI  [C/RAN-P]

X-MEN [X-MEN] RUSCH, KRISTINE KATHRYN/ SMITH, DEAN WESLEY

P

X-MEN THE BROOD SAGA, PART 1 [X-MEN]  CLAREMONT, CHRIS  SCX
CASTLE ROOGNA (PROOF) [XANTH 3]  ANTHONY, PIERS  SR-LP
MAGIC OF XANTH, THE [XANTH 1,2,3]  ANTHONY, PIERS  H
PIERS ANTHONY’S VISUAL GUIDE TO XANTH [XANTH]  ANTHONY, PIERS/NYE, JODY LYNN  R/REF-P,C/REF-H:2
RETURN TO CENTAUR [XANTH]  PINI, RICHARD  CX
SPELL FOR CHAMELEON, A [@XANTH #1]  ANTHONY, PIERS  C/P:2,P
SOURCE OF MAGIC, THE [@XANTH #2]  ANTHONY, PIERS  C/P:2,P
CASTLE ROOGNA [@XANTH #3]  ANTHONY, PIERS  C/P:2,P
CENTAUR AISLE [@XANTH #4]  ANTHONY, PIERS  C/P,P,H
OGRE, OGRE [@XANTH #5]  ANTHONY, PIERS  C/P:2,P
NIGHT MARE [@XANTH #6]  ANTHONY, PIERS  C/P,P,H
DRAGON ON A PEDESTAL [@XANTH #7]  ANTHONY, PIERS  C/P,P,H
CREWEL LYE [@XANTH #8]  ANTHONY, PIERS  C/P:2,P
GOLEM IN THE GEARS [@XANTH #9]  ANTHONY, PIERS  C/P,P,H
VALE OF THE VOLE [@XANTH #10]  ANTHONY, PIERS  C/P:2,P
HEAVEN CENT [@XANTH #11]  ANTHONY, PIERS  H:2,P
MAN FROM MUNDANIA [@XANTH #12]  ANTHONY, PIERS  C/P:2,P
ISLE OF VIEW [@XANTH #13]  ANTHONY, PIERS  H:2,SFWA,P
QUESTION QUEST [@XANTH #14]  ANTHONY, PIERS  C/P:2,P
COLOR OF HER PANTIES, THE [@XANTH #15]  ANTHONY, PIERS  C/P,SFWA-P,P,H
DEMONS DON’T DREAM [@XANTH #16]  ANTHONY, PIERS  H:2,P
HARPY THYME [@XANTH #17]  ANTHONY, PIERS  C/P,P,H
GEIS OF THE GARGOYLE [@XANTH #18]  ANTHONY, PIERS  C/P,P
ROC AND A HARD PLACE [@XANTH #19]  ANTHONY, PIERS  C/P:2,P
YON ILL WIND [XANTH #20]  ANTHONY, PIERS  H,@P,@C/P
FAUN & GAMES [XANTH #21]  ANTHONY, PIERS  H,@P,@C/P
ZOMBIE LOVER [XANTH #22]  ANTHONY, PIERS  H
XONE OF CONTENTION [XANTH #23]  ANTHONY, PIERS  H
DASTARD, THE [XANTH #24]  ANTHONY, PIERS  C/P,P,H
SWELL FOOP [XANTH #25]  ANTHONY, PIERS  H,P
UP IN A HEAVAL [XANTH #26]  ANTHONY, PIERS  H
CURRANT EVENTS [@XANTH #28]  ANTHONY, PIERS  H
PET PEEVE [@XANTH #29]  ANTHONY, PIERS  P
STORK NAKED [@XANTH #30]  ANTHONY, PIERS  H
FLUX [XEELEE]  BAXTER, STEPHEN  C/P,P,H
RAFT [XEELEE]  BAXTER, STEPHEN  C/P:2,SFWA-P,P
RIDING THE ROCK [XEELEE]  BAXTER, STEPHEN  SR-LP
RING [XEELEE]  BAXTER, STEPHEN  C/P:2,P
TIMELIKE INFINITY [XEELEE]  BAXTER, STEPHEN  C/P,P
BATTLE ON! [XENA]  COX, GREG  R/REF-P
EMPTY THRONE, THE [@XENA]  EMERSON, RU  P
FURTHER ADVENTURES OF XENA WARRIOR PRINCESS [XENA]  GREENBERG, MARTIN H.  PA
GO QUEST, YOUNG MAN [@XENA]  EMERSON, RU  P
HUNTRESS AND THE SPHINX, THE [@XENA]  EMERSON, RU  P
PROPHECY OF DARKNESS [@XENA]  HOWARD, STELLA  P
THIEF OF HERMES, THE [XENA]  EMERSON, RU  P
ADULTHOOD RITES [XENOGENESIS 2]  BUTLER, OCTAVIA E.  H:2,P
DAWN [XENOGENESIS 1]  BUTLER, OCTAVIA E.  H:2,P
IMAGO [XENOGENESIS 3]  BUTLER, OCTAVIA E.  C/P,P,H
ULTIMA GENESI (ITALIAN) [XENOGENESIS 1]  BUTLER, OCTAVIA E.  FP
XENOGENESIS [XENOGENESIS 1,2,3]  BUTLER, OCTAVIA E.  H:2
DARK CRUSADE, THE [XOR 3]  HUNT, WALTER H.  P
DARK PATH, THE [XOR 2]  HUNT, WALTER H.  H
DARK WING, THE [XOR 1]  HUNT, WALTER H.  C/P,P

TITLE  AUTHOR  CODE
RAGE IN HEAVEN, A (PART 1) [YAMATO]  KATO, KEN  P
RAGE IN HEAVEN, A (COMPLETE) [@YAMATO #1]  KATO, KEN  H
RAGE IN HEAVEN, A (PART 2) [@YAMATO #1]  KATO, KEN  C/P,P
WAY OF THE WARRIOR, THE (PART 2) [@YAMATO #2]  KATO, KEN  P
MAGIC STUDY [YELINA 2]  SNYDER, MARIA V.  H
POISON STUDY [YELINA 1]  SNYDER, MARIA V.  H
CASTLE OF HAPE, THE [YNELL 3]  MURPHY, SHIRLEY ROUSSEAU  H,P
CAVES OF FIRE AND ICE [YNELL 4]  MURPHY, SHIRLEY ROUSSEAU  C/P,P
JOINING OF THE STONE, THE [YNELL 5]  MURPHY, SHIRLEY ROUSSEAU  H,P
RING OF FIRE, THE [YNELL 1]  MURPHY, SHIRLEY ROUSSEAU  H,P
WOLF BELL, THE [YNELL 2]  MURPHY, SHIRLEY ROUSSEAU  C/P,P,H
YNGLING, THE [YNGLING 1]  DALMAS, JOHN  C/P:2,P
YNGLING AND THE CIRCLE OF POWER, THE [YNGLING 2]  DALMAS, JOHN  C/P:2,SFWA-P,P
YNGLING IN YAMATO, THE [YNGLING]  DALMAS, JOHN  C/P,P
ORC WARS, THE [@YNGLING #1,2]  DALMAS, JOHN  C/P:2,SFWA-P,P
BASTION OF DARKNESS [YNIS AIELLE 3]  SALVATORE, R. A.  P
ECHOES OF THE FOURTH MAGIC [YNIS AIELLE 1]  SALVATORE, R. A.  P
WITCH’S DAUGHTER, THE [YNIS AIELLE 2]  SALVATORE, R. A.  P
GIRL FROM EROSPHERE, THE [YOLANDA 1]  VERSEAU, DOMINIQUE  SR-L
SLAVES OF SPACE [YOLANDA 2]  VERSEAU, DOMINIQUE  SR-L
HOBBY [YOUNG MERLIN 2]  YOLEN, JANE  H
HIGH WIZARDRY [YOUNG WIZARDS 3]  DUANE, DIANE  H:2,P
SO YOU WANT TO BE A WIZARD [YOUNG WIZARDS 1]  DUANE, DIANE  C/P,P,H
SUPPORT YOUR LOCAL WIZARD [YOUNG WIZARDS 1,2,3]  DUANE, DIANE  H
WIZARD ABROAD, A [YOUNG WIZARDS 4]  DUANE, DIANE  C/P,P
WIZARD ALONE, A [YOUNG WIZARDS 6]  DUANE, DIANE  H
WIZARDS AT WAR [YOUNG WIZARDS 8]  DUANE, DIANE  H
WIZARD’S DILEMMA, THE [YOUNG WIZARDS 5]  DUANE, DIANE  H
WIZARD’S HOLIDAY [YOUNG WIZARDS 7]  DUANE, DIANE  H

TITLE  AUTHOR  CODE
DOOMSDAY BRUNETTE, THE [ZACHARY NIXON JOHNSON 2]  ZAKOUR, JOHN/GANEM, LAWRENCE  C/P,P
FROST-HAIRED VIXEN, THE [ZACHARY NIXON JOHNSON 4]  ZAKOUR, JOHN  P
PLUTONIUM BLONDE, THE [ZACHARY NIXON JOHNSON 1]  ZAKOUR, JOHN/GANEM, LAWRENCE  C/P,P
RADIOACTIVE REDHEAD, THE [ZACHARY NIXON JOHNSON 3]  ZAKOUR, JOHN/GANEM, LAWRENCE  P
EARTH-SHAKER, THE [ZARKON 4]  CARTER, LIN  H,SFWA
INVISIBLE DEATH [ZARKON 2]  CARTER, LIN  H:2,SFWA,P
NEMESIS OF EVIL, THE [ZARKON 1]  CARTER, LIN  P,SFWA
VOLCANO OGRE, THE [ZARKON 3]  CARTER, LIN  C/P,SFWA,H
PRISONER OF ZENDA, THE [ZENDA 1]  HOPE, ANTHONY  SR-H
RUPERT OF HENTZAU [ZENDA 2]  HOPE, ANTHONY  SR-H
FISH DINNER IN MEMISON, A [ZIMIAMVIA 2]  EDDISON, E. R.  SR-H,C/P:2,P
MEZENTIAN GATE, THE [ZIMIAMVIA 3]  EDDISON, E. R.  C/P:2,P
MISTRESS OF MISTRESSES [ZIMIAMVIA 1]  EDDISON, E. R.  SR-H,C/P:2,P
WORM OUROBOROS, THE [ZIMIAMVIA]  EDDISON, E. R.  H:3,P
ZIMIAMVIA [ZIMIAMVIA 1,2,3]  EDDISON, E. R.  LP
SCENT OF SHADOWS, THE [ZODIAC 1]  PETTERSON, VICKI  P
TASTE OF NIGHT, THE [ZODIAC 2]  PETTERSON, VICKI  P
WRAITH [@ZOE MARTINIQUE]  WELDON, PHAEDRA  LP
BLIND FIRE [@ZONE #2]  ROUCH, JAMES  P
NIGHTMARE OF GOD, THE [ZORACHUS 2]  ROGERS, MARK E.  P
ZORACHUS [ZORACHUS 1]  ROGERS, MARK E.  P
CAVERN OF DOOM, THE [ZORK 3]  MERETZKY, S. ERIC  C/P,P
CONQUEST AT QUENDOR [ZORK 4]  MERETZKY, S. ERIC  SFWA-P
FORCES OF KRILL, THE [ZORK 1]  MERETZKY, S. ERIC  P
LOST CITY OF ZORK, THE [ZORK]  BAILEY, ROBIN WAYNE  C/P:2,P
MALIFESTRO QUEST, THE [ZORK 2]  MERETZKY, S. ERIC  P
ZORK CHRONICLES, THE [ZORK]  EFFINGER, GEORGE ALEC  C/P:2,P
TALES OF ZOTHIQUE [ZOTHIQUE]  SMITH, CLARK ASHTON  LP
ZOTHIQUE [ZOTHIQUE]  SMITH, CLARK ASHTON  SR-P,C/P:2